A Complete Bibliography of Publications in the Journal of Statistical Planning and Inference:

2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

(69, 17, 4) [Pav01]. (c, γ) [AKK09]. (D_{t}, C) [TV01]. (m, n) [HDM09]. (q − q + 1) [SV01b]. (r, s) [HDM09]. (v, 5, λ) [CGS02]. 1 [BZ00, Chu01, Hai07, KM00a, LS02, Mor09]. $120.00 [Roy07]. 13 [Top01]. 2 [BRO^{+}09, CH05a, CLLZ06, HZN04, KM00a, LMM01, SB04, TTC06]. 2 − (27, 9, 4) [LT96, LT00]. 2 − (69, 17, 4) [Top01]. 2^{(n_{1}+n_{2})−(k_{1}+k_{2})} [YLLZ06]. 2^{k} [BR09]. 2^{m} [Kuw02, SST02]. 2^{m−k} [Jac04]. 2^{n} [Ge01a]. 2^{n}u^{1} [Ge01b]. 2_{IV}^{m−(m−k)} [YB08]. 2 ≤ t ≤ 4 [DM02]. 2 × C [KL07]. 3 [BRS00, CKMS02, ITA^{+}09, NB08, vT01]. $39.95 [McF01]. 3 × m_{2} × m_{3} [CKM06]. 4 [LT08b]. 5 [BRS00]. $74.95 [Lah00]. $84.95 [Ko00]. $94.50 [Lu02]. $94.95 [Dea00]. A [BdM09, GRP02, Jac02, Yan08]. a^{n}b^{1} [XP01]. α [Kag08, Kör00]. c [FC03, HC03, HS00b]. C(α) [BB01b, Pal03]. C_{2} [SA02]. C_{2}^{b} [SA02]. T^{2} [SM01]. χ^{2} [Gil05, KS09a]. COV(X_{j}^{2}, X_{p}^{2}) [Hod04]. D [BJJV08, BR09, CN02b, CKM06, DH08, FWW06, HC03, JLN08, KM08, KS06, LM08, MW08a, OP04, Orr08, RO05a, RGR09, WMSY06, Wu03, Yan05, Zho08, vBP03, MG08]. D(r^{m}(r + 1), r^{m}(r + 1); p) [ZDLZ02]. D_{s} [RVZRZ07, Yan05]. E [MR07, YO09]. E(χ^{2}) [AFH07]. E(s^{2}) [Bul07b, BR08, DDCC08, KMS08, KMS09a, LZ00, LRD07, RB07]. ED_{p} [HYP02]. F [Bat04, DPA08, HDM09, HS09, LL09, WRV05, Wit01, YJS09, BH02]. G [CH08, JHCH05, LDT04]. G_{7} [Ge01a]. gh [Fie04]. k [AJMP09, ATS02, BS07a, CC09, Che03a, CL08, CH08, ENAZ09, GT08b, HA03, JHCH05, KP09a, KK01, KK05, Ker04, Kwo01, MCDC08, Oht00, QT07, Raq04, Ree00, SAC02, War08, YJS09]. K_{4} − e [LT08b]. K_{r} × K_{c} [FHJ^{+}04]. k_{v} [Ge01a, ZY02]. L [Ani10, ES04, GR08a, Hos06b, Hos07a, Hos07b, Jon04, KOS05, MA08, PR04, vdWD01, Tan09a, Tan09b]. L^{∞} [ANR00]. L_{1} [Lou00, Maj07, SW07]. L_{2} [Che06a, ZN08]. L_{p} [HL04, LC01b, Lin08]. λ [Cha01b, LR02, LT08b]. λ = 1 [IT10, ITA^{+}09]. λ ≥ 2 [LT08b, Yin00]. λK_{n} [LT08b]. λK_{v} [KDT06]. lim sup [ALRV01]. M [Arc01, Ars04, BGS03, BLV00, CC07b, CZ09b, DHP06, HM01, Hlá03, Hwa04, Jin07, KS00, SMS08, Sto05, TZ05, WR02, Wie00, Zha00, DdRB00, HZ06, IJZ09, MvdM03, Tan09a, Tan09b, Wyn08, XL06b]. m = 4, 5, 6 [LLvRZ00]. m^{4} [yS05]. MR(3, k; v) [Ree00]. N [Aga02, Bak09, BR08, RSS02, Sen02, Che03a, CH08, DR05, FHKMS02, HW09, IJZ09, JHCH05, SW07, XP01, YJS09, ZW03b]. n^{r} [OK00]. n ≤ 2k [JHCH05, CH08]. P [IVM03, Raq04, DR05, Ioa07, KM03, MP08, MS07a, SB05, WW06, YLH04]. p, q [LLLH05]. P [X ≤ Y] [SD01]. PG(3, q) [BE01]. PG(q − q − 3, q) [SV01b]. Φ [Cao09, CP02b, JW09, PPZ02, PP05, PPdCP06]. πps [Air00, AR05, Ros00]. q [Cha05, CD07b, Kem02c, Kem05, KK09c, SV01b, WZ01a]. R [MW07b, HW09]. r = λ_{1} + 1 [SK02]. R^{2} [HS08a]. ρ [Che07, GN09, LL07b, vdWD01]. S [Ber02c, CG03]. S(2, 4, 52) [LMM01]. σ [MNOP05]. t [DvdW05, DGGJ09, Fab09, Fan05, HCS04, LL07c, MRC03, OK00, Sar06, Wit01, YH09, Zha09b, ZH08]. T (2, 3, v) [BKTR08]. T (K_{1,2p}) [LK08]. t − (t + 8, t + 2, 4) [DM02]. T^{2} [KI08, SM01]. t_{ν} [JB08, JB09]. τ [Che07, GN09, LL07b]. t ≥ 2 [MRC03]. θb [Lia09]. × [BCS07, Beh08, Che04d, FS02a, KM08, SJ08]. U [BW08, Chr04, DR01a, FM00b, FM00c, GH02, HH05, RS04b, SS02e, TS08, ZS08, CD07a]. U|(U + v = m) [Kem03]. V [BR09, DH08, QW05, Ree00]. V (m, t) [LLvRZ00, CvRZ01, CZ02]. v ≡ 3 mod 6 [DS02b]. v ≤ 32 [BH01b]. W [BC03a]. W_{p} [Has09]. X [SP05]. X_{(1)} [Sch04a]. X_{(n)} [Sch04a]. Z_{4} [BRS00].

-aberration [LDT04]. -Analogues [Kem02c]. -arc [SV01b]. -ary [CD07b, FHKMS02]. -Bhaskar [HS00b]. -circulant [CL08]. -class [Oht00]. -cluster [KK09c]. -codes [KM00a, BRS00]. -Colored [BRS00]. -confluent [Kem05]. -contamination [AKK09]. -cube [Ker04]. -cycles [LT08b]. -dependence [BdM09]. -dependent [Hai07]. -Designs [BRS00, KM00a, MRC03, Pav01, vT01]. -deviation [Che06a]. -dimensional [CH05a]. -discrepancy [ZN08]. -distance [Lou00]. -distances [Bak09]. -distributions [Cha05, JB09]. -divergence [JW09, PPZ02]. -divergences [CP02b, PP05, PPdCP06]. -efficient [BJJV08]. -estimates [Ars04, BLV00]. -estimation [BGS03, CZ09b, DHP06, HM01, Jin07, WR02, Wie00, Zha00]. -estimator [Hlá03]. -estimators [Arc01, Ber02c, CC07b, HL04, KS00, SMS08]. -factorizations [Chu01]. -factors [LS02]. -family [CGS02]. -field [MNOP05]. -fold [LR02, LT08b]. -identity [WZ01a]. -information [Fab09]. -level [CG03, CLLZ06, NB08, SB04]. -match [HA03]. -means [KP09a, QT07]. -metric [KS09a]. -mixture [YO09]. -moments [ES04, Hos06b, Hos07a, Hos07b, Jon04]. -norm [Raq04]. -norms [SW07]. -NRBDs [Mor09]. -optimal [FC03, HC03, RB07, AFH07, Bul07b, BR08, BR09, CN02b, CKM06, DDCC08, DH08, FWW06, GRP02, JLN08, KM08, KS06, KMS08, KMS09a, LM08, MW08a, MR07, Orr08, RVZRZ07, RGR09, TV01, WMSY06, Wu03, Yan05, Yan08, Zho08, vBP03]. -optimality [QKO08, LC01b, OP04]. -optimum [RO05a]. -out-of- [Che03a, CH08, HDM09, HW09, JHCH05, YJS09]. -packings [Ge01a]. -parameter [HZN04]. -processes [SS02e]. -quantiles [ZS08]. -record [AJMP09]. -regression [Sto05]. -resolvability [Kag08]. -robust [TZ05]. -rotational [BZ00, Mor09]. -run [yS05, RSS02]. -Sample [MCDC08, DvdW05, ENAZ09, KK01, War08]. -smoothers [Hwa04]. -spacings [HZ06, MvdM03, XL06b]. -squared [BH02]. -SSA [GO07]. -statistic [Ani10, GH02, MA08]. -statistics [FM00b, BW08, Chr04, DR01a, FM00c, GR08a, HH05, KOS05, PR04, RS04b, TS08]. -step [TTC06]. -symmetry [MW07b]. -systems [LMM01]. -test [Kör00, WRV05]. -tests [Zha09b, HS09]. -th [BS07a, Tan09b]. -triple [LK08]. -tuple [GT08b]. -type [HCS04]. -value [MP08, MS07a, YLH04]. -values [KM03, SB05, IVM03]. -way [Wyn08]. -wise [CKMS02, OK00].

/M/1 [SA02].

0 [Cha01a, Cha02a, Con02a, Con02b, Dea00, Ko00, Lah00, Lee01a, Lin01b, Lu02, Mas01, Sin06b, Wes00]. 0-387-98629-4 [Cha01a]. 0-387-98674-X [Mas01]. 0-387-98718-5 [Con02a]. 0-471-13773-1 [Lu02]. 0-471-15410-5 [Ko00]. 0-471-25184-4 [Wes00]. 0-471-29414-4 [Dea00]. 0-471-31711-X [Lee01a]. 0-471-34912-7 [Lah00]. 0-471-49600-6 [Roy07]. 0-471-82668-5 [Cha02a]. 0-521-35564-8 [Lin01b]. 0-521-58611-9 [Lin01b]. 0-521-63201-3 [Con02b]. 0-8204-7348-0 [Sin06b].

1 [Aga02, AWL04, Con04, KKLD07, Lu02, McF01, SA02]. 1-110 [Ano01n]. 1-120 [Ano01o]. 1-162 [Ano04-52]. 1-166 [Ano03-49]. 1-168 [Ano04-53]. 1-170 [Ano00x]. 1-176 [Ano00z, Ano00y]. 1-178 [Ano00-27]. 1-180 [Ano03-50]. 1-206 [Ano04-54]. 1-208 [Ano05-64]. 1-210 [Ano02v]. 1-222 [Ano04-61, Ano05-55]. 1-224 [Ano04-55]. 1-228 [Ano00-28]. 1-238 [Ano02w].

1-250 [Ano04-56]. 1-282 [Ano00-30, Ano00-29]. 1-288 [Ano00-31]. 1-302 [Ano05-56]. 1-308 [Ano01p]. 1-310 [Ano06-39]. 1-312 [Ano01q]. 1-326 [Ano01r, Ano05-57]. 1-328 [Ano02x]. 1-332 [Ano03-55]. 1-348 [Ano03-56, Ano05-58]. 1-354 [Ano00-32]. 1-362 [Ano07-39]. 1-374 [Ano03-51]. 1-378 [Ano02y]. 1-400 [Ano04-62]. 1-472 [Ano02z]. 1-58025-357-1 [McF01]. 1-90 [Ano02-27]. 100th [Sri02b]. 102 [Rao02a, Sri02a]. 1099-1514 [Ano07-40]. 111-224 [Ano01s]. 1197-1568 [Ano06-40]. 1227-1520 [Ano08-37]. 124 [FCCN05]. 1257-1574 [Ano09-38]. 128 [DGL07]. 129 [Gov09]. 136 [EN07]. 137 [EJ09, Ten09]. 138 [Ani10, BD09]. 139 [Fox09, IT10, LT09b, Tan09a]. 144 [Gov09]. 1521-1918 [Ano08-38]. 163-340 [Ano04-57]. 167-388 [Ano03-52]. 171-356 [Ano00-33]. 1711-2088 [Ano07-41]. 177-350 [Ano00-34]. 177-376 [Ano00-35]. 1821-2116 [Ano09-39]. 1919-2252 [Ano08-39]. 1987 [Moh02b]. 1997 [Suk02]. 1998 [Nie02a].

2 [YHA09b]. 2- [EKTR00]. 201-416 [Ano01t]. 207-416 [Ano04-58]. 2089-2534 [Ano07-42]. 209-418 [Ano05-59]. 211-502 [Ano02-28]. 2117-2510 [Ano09-40]. 221pp [Sin06b]. 223-562 [Ano05-60]. 225-450 [Ano04-59]. 2253-2592 [Ano08-40]. 239-504 [Ano02-29]. 2407-2846 [Ano06-41]. 245-530 [Ano05-61]. 251-490 [Ano04-63]. 2511-2862 [Ano09-41]. 258pp [Con02a]. 2733-3040 [Ano07-43]. 283-450 [Ano02-30]. 283-482 [Ano00-36]. 283-630 [Ano00-37]. 2847-3292 [Ano08-41]. 2847-3318 [Ano06-42]. 2863-3392 [Ano09-42]. 2k [But09]. 2m [Jac06, LJ07a, Tan07]. 2m-k [Jac06, LJ07a]. 2SHI [Nam00].

3 [Con02b]. 3-631-52934-1 [Sin06b]. 301-440 [Ano01u]. 303-606 [Ano05-62]. 311-526 [Ano06-43]. 327-664 [Ano05-63]. 3319-3734 [Ano06-44]. 333-538 [Ano03-53]. 338pp [McF01]. 3393-3738 [Ano09-43]. 349-702 [Ano03-54]. 363-628 [Ano07-44]. 3635-4198 [Ano08-42]. 3735-4118 [Ano06-45]. 375-718 [Ano03-57]. 382pp [Lu02]. 3947-4240 [Ano09-44].

4 [Cha01a, Dea00, Wes00]. 401-594 [Ano04-60]. 4119-4474 [Ano06-46]. 411pp [Con02b]. 442 [FR08]. 4th [Ano05a].

5 [Cha02a, Con02a, Ko00]. 527-1196 [Ano06-47]. 539-816 [Ano08-43]. 56 [LT00]. 593pp [Lee01a]. 5th [Che04a].

679-1256 [Ano09-45]. 6th [Ano09a].

7 [Lah00]. 72 [Bed00].

81 [Jan01]. 81-678 [Ano09-46]. 817-1226 [Ano08-44]. 83 [FM00b]. 86 [FR08]. 8th [Ano09a].

9 [Lin01b]. 91 [TLBJ11]. 91-446 [Ano02-31]. 96 [Fre02].

A-optimal [CDD01]. AA2024 [RWF04]. AA2024-T3 [RWF04]. abandon [KSS01]. aberration [AZ04b, AYZ06, FZ04, Jac04, LDT04, Qin03, TD03, Tan07, XL06a]. Absorption [Raw02]. ACC [CLA09]. Accelerated [Ano09-48, OP00, BCN04, BM06e, ES00, FL09, LS01, LuV09, OP04, VM04, WJ08, Yu06]. acceleration [UF09]. according [CC09]. account [SG08]. accumulation [AFF09, Fin09]. Accuracy [Gil05, KV07, Llo05]. achieve [Pal09]. acids [OCR^{+}05]. Acknowledgement [Ano03a]. action [KZ03a, Pav01]. Active [KS03a, HW09, ZZL07]. activity [AC04a]. actuarial [BJPZ08]. acute [Sta06]. acyclic [CS03a, Gil06, KKM09]. Adaptation [Ber07, She02]. adapting [CD03b]. Adaptive [AIK00, Bla06, Con09, DF06, DFW08, Efr07a, EBZ08, GJ05, HS08b, Kra09, MPS05, Pen00, Aic01, BB04a, Bel07, BBP06, Cai02, cIC01, cICY04, CBG04, CBD05, CD07b, CG00, CW05, DL03, DH09c, FM00a, GMPZ06, Ger05, HOS06a, HS07a, HRZ06, IBL06, Jeo00, KA06, KBG09, Kus03, LC09a, LP06, MP00, Phi02, Pla09, PS09, PG07b, Sta06, WBvdL07, YCB06, YW08, Yua08]. Adaptively [MY04]. added [CR02]. Addendum [EJ09]. addition [DDvdL08]. additional [AH08, Har09, Koh06]. Additive [GFDMG04, ACW02, Che09, FK09, JL09a, MRCN09, SA03, XY06, YSH03, Csá02]. additivity [Gol01, KSS01, Sud01]. adequacy [SW09]. adjacent [DW00, WS08, Wri08]. adjust [BCR08]. adjusted [FCCN04, FCCN05, RMS02, SSP03, SN09, dSSdL08, Vas03]. adjustment [Che06c, GM01a, Lee01b, LdC06, MMM07]. Adjustments [PS06, Bra06]. Admissibilities [WN06]. Admissibility [NWS01, XF00, Tak09, Cao09]. Admissible [MT07a, ZW03b, HT09a, cIWsL05]. admitting [MNOP05, Pav01]. admixture [PS05]. Adrian [Lin01b]. advances [Dob02, GT06]. advantages [HML00]. adverse [LHOP01]. aeronautic [MCM^{+}07]. Affine [VOK^{+}03, CK08, Kag08, LT96, LT00]. affinely [AT09]. AFT [AGP09, CDB^{+}09]. after [Din03, DW06, Lee05, MP00, Sch04b, Tan09a, Tan09b]. against [Ani10, BP07, EKT06, ENAZ09, FW07, KSH02, LY02, MA08, MML02, Par00a, Pou01, SSB02]. age [Ahm05a, Ver01]. ageing [CZ09a, DL06b]. aggregate [Beh08, KA08]. aggregation [Azo09, FCS05, Nau02]. aging [Ahm01, AAEK06, Fin09]. agreement [CN07, Cho08]. AIC [HK03, YSF03, YO05a]. AIDS [FS01a]. air [SSC07]. al [IT10]. Albert [Con02a]. ALC [CLA09]. alert [DG09a, LSL06]. alert-delay [DG09a]. algebra [GPR03a, Zha01a]. Algebraic [PR08a, Wyn08]. algebras [DHW00, Sri05b, Sri07a]. algorithm [Ars04, BRCK03, BB03b, DDvdL08, GGS09a, JCS05, KS03a, KMS09a, MPS05, NS00, QT07, RBV00, SA03, SY09, ZZF09, vdLP03]. Algorithmic [YLS00, BW00b]. Algorithms [SR05, Gar02, HS08b, LS06, LZ07, Roy02, Sch05a, SvTW00, TMM09, WPV04, WC00]. alias [Kuw02]. aliased [Kuw02]. aliasing [BY09, Wyn08]. Aligned [EOPR05]. alignment [HLM08]. allied [SS02e]. Allocation [Kus03, Aic01, BB04a, BB04b, BB06a, BK02, GK01, GHSS02, HW09, LRZ02, MP00, Wie05, ZH09]. allocations [YJS09]. alloy [RWF04]. alloys [ZWHF07]. Almost [ALRV01, Bag04, Mis02, WW09]. alternate [WZ01a]. alternating [YCB06]. Alternative [Cho07b, TK06, CP02a, DG09b, IJZ09, KSH02, LL09, Lin07, PW02a, Pfa03, Pou01, SSM00, SA07, YW09a]. Alternatives [Eks08, Ani10, BP07, BCR00, CA04, CS02c, DL04, DN03, FJT09, GKM07, Gua07, JÜ08, Mey06, MFSW07, MA08, MML02, NPF09, PW02b, RW06, WL01]. aluminum [RWF04, ZWHF07]. always [Bos06]. Aman [Lin01b]. ambiguity [Man02]. amino [OCR^{+}05]. among [DMMS07, FW07, HK05, KK02, KS03b, SN08, TKP06, Wu07]. among-group [HK05]. amounts [KA08]. analog [SM01]. analogue [IK06, LW07]. Analogues [Kem02c]. analyse [VdBS01]. Analysis [ÅvR06, BBKN07, BR02, CN02a, Cha03a, Con02b, IJK08, KB00, LHOP01, OCR^{+}05, PPX09, Sah08, AC02, Ahm05a, Ant08, AVGRZ00, BZZH09, Bar03, BM02b, Beh08, Ber02a, BG09, BBIAV00, BPAK04, CK08, CA04, CCH08, CM00a, Che03a, CGY07, CZ09c, CA03, Con04, Cua02, DG09a, Fab09, FTY04, FR07, FS01a, FS02b, FHW08, Gar02, GK02, GML06, GP04, GMM03a, GMM03b, GKM07, Gol01, GS08, GO07, GNP07, HS05, HKP08, HCS04, HMW04, HKZ08, HC06a, HLH09, HN04, IL08, KMPP02, KB08, KKB09a, KK09b, Kir07a, KY07, Ko00, KV06, Kot06, KS07d, LL05a, LX07a, LLLH05, LP01, LTW08, LL07c, LRD07, Mar07, MDM06, Muk06, NA09, Neu00, ND05, NBH05, OM07, PW06a, PW06b, PVHK09, Pav03, PP06b, PS05, PMS^{+}01, RDS08, RS07b, Sam07, SSO07, Sen02, SA02, SK07b]. analysis [Sta04, SSY05, Ste02, Sud01, SN04, Swi02, TMP06, TL08, TP03, VEM07, Vel03a, VIBAV06, VS00, WW02a, WG03a, Whi05, WYL09b, YZ09, ZKW01, ZjLS05, ZG05, ZC03, ZGH08, dL06a, vEKO00, Lu02]. Analytic [KKM09, Xin01, MBA^{+}08]. Analytical [FM00a, GN09, DD08, GH07]. Analyzing [NK06, GACS07, MH00]. ancillarity [Kag02]. Anderson [Atk09, Bor09, Jon09, Khu09, LOV08, Luc09, Par09b, Pie09, BBR04, CSX08]. Anderson-Cook [Atk09, Bor09, Jon09, Khu09, Luc09, Par09b, Pie09]. Andrews [KN02]. angles [EPS08]. angular [KS08a, RS04c]. animals [BPAK04]. annealing [KMS08]. Anniversary [Sri02b]. annotated [PS02]. Announcement [Ano07a, Ano07b, Sri04]. ANOVA [Bat04, GKK08, GHSN00, KD07, OM07, WRV05, WB05]. ANOVA-like [OM07]. Antieigenvalues [Kha03]. antieigenvectors [Kha03]. any [Ma09]. Appeals [Goo08]. applicable [HSS08]. Application [BCC02, MW00b, SK08, BGP00, BF07, BGL01, BK09c, Bon08, BB06b, CBG04, CBD05, CTL04, Che03b, DFS08, DH09a, DdHL06, FPvdVvH06, GKDN06, Ge01a, GPL09, IS02b, KM07a, KK09b, KTMS04, KT07, LT09b, LT09c, MDM06, MCM^{+}07, PS02, RdCCB09, RSS04, VEM07, WH08, WL01, CGG09, Pro02]. Applications [Ano06b, Ahm01, AM04b, ALRV01, AK03, AK05, BTK08, BL08a, BJPZ08, BDHS06, CF04, Che00b, CCL06, CJRV08, DGGJ06, DDvdL08, EN09, FZL00, FR01, FS02a, FCM07, FHMS02, GG07, HC03, HB00, Hay06, Hil09, HT06, HLH09, Jan01, Jia03b, KR03, KN02, Koh00, KS07d, KMS09b, LL05a, LX07a, LLL09, LB06, LLP09, Mei07b, MP05, MW07b, Nie07, NSV02, Pal03, Per00, Qiu01, Ral02, SGK09, Sri07a, dBJM09]. applied [CLE09, PLLS08, SA02, Roy07, Ko00, Lee01a, McF01]. Applying [KMPP02]. appraisal [Rao02b, SS02b]. Appreciation [Sri07b, SV01a, Zac05a]. approach [AdL05, Aga02, AP05, AH01, Ani10, Bon08, CGG09, CS04a, Che08, Cho08, CBC^{+}07, CO02, DMP04, DP08c, Dub07, EOPR05, ETS02, EKK06, FG06, GSCG09, GL02b, GPW01, Hai03, HCBK05, KK09a, KKB09a, Ken00, KK09d, KY07, KK03, KLX09, KN03, Kwo01, Lee01b, LT02, Len09, LM00a, LW09, MT06, Mee05, Mel05a, MS07a, MA08, Moj07, MM06, Mor09, OP04, Pfa03, PRBAC08, Pip09, RM05c, SAC02, Sen02, Sen07b, SR00, sS06a, SA07, SdMSL04, UF09, Van05, VBF03, Xin01, Yao08, YGJ05, Zha09c, ZjLS05, dHD^{+}09, Wes00]. approaches [TWH^{+}05, WC09b, Whi05]. Approximate [GGS09b, JI04, RM07, WW06, FW03, KK01, WRV05]. Approximating [CS00, Lij03, KT08b]. Approximation [DGR07, FZ06, Xue04, AA08, And05, AWL04, DT05, KD07, LCM07, LHS00, LBM00, MPS09, Roo03, Sel00, YZZ09]. Approximations [BGM03, DdHL06, Hor00, AC07, CP05b, CC02, Con04, DFS08, Gil05, GC07, KT02, KR07, Lee05, MW05, Sta09a, YZ09]. April [Ano00-28, Ano00-30, Ano02-28, Ano02z, Ano03-51, Ano04-57, Ano05-64, Ano06-40, Ano07-40, Ano08-44, Ano09-38]. arbitrarily [For09]. arbitrary [CT09]. arc [SV01b]. ARCH [HL04, Laï05, TT08, TL08]. Archimedean [BNW05]. area [Blu09, BL03, Con09, DS03, EN09, FMN09, Jia03a, Lar03, Mai01, MCDC08, WYL09a, YR03]. areas [ZC03]. ARFIMA [AM04c]. argument [Goo08]. ARHD [MP07a]. arising [GM01a]. arithmetic [Var02]. ARMA [FZ00, KSL03, LLLH05, LP04, SL00, UV07]. Armitage [KL07]. array [CMyS02, EKK06, GHJ^{+}07]. arrays [BCN06, BM08, CKMS02, DL07, EKK06, EKL07, FHM00, HY08, Hil09, KM00b, MJS00, Ngu08, RSS02, ySD03, ySK05, yS07, WC09a]. artificial [Heu08]. ary [CD07b, FHKMS02]. ascertained [Dav05]. aspects [Bul07a, Kem02b, LRZ02, MS07b, RS07b, SB06, SS02b]. assays [CDD01]. Assessing [Ahm05b, EMP06, GK01, KHL07, KG05, Tsa06, Zha04a, DPA08, PS03]. assessment [Aus05, CB06, CN07, Cho08, FCM07, GS08, Sta06]. assessments [WPV04]. assets [TW03]. assisted [Dub07, Kot05]. associated [Chr04, DR01a, KV02, MR01, MS02d, Pom00, Swi02, XY08, Xin09]. Association [SZ06, Gup03, HO03, HLH09, OKD06]. assumption [GACS07, GR08b]. assumptions [CvdL01, GPC01]. assured [LRZ02]. ASTA [ETS02]. asymmetric [AVGQ05, BS05c, MY04, ySD03]. asymmetry [CHP08]. Asymptotic [Arc01, AD01, BS08, BL05c, BRO^{+}09, CCL03, Cha00, CZ09b, Che05, CLR05, DEH00a, DH06, DR01a, FR01, FJT09, Fuj02, FHW08, Gar05, GKK04, HP06a, Hlá03, Jan04, KK03, Lia00a, LJ09, Mar07, Oga09, Pla09, QJ00, RL06, Rob09a, Ruk02, SZ07, WvEZ04, WT08, WA07, ZKW01, Zha04b, ZGH08, ABK00, AC07, BV06, BP01, BK07, Bos09, CC07b, Cha02a, Cha02c, Che06a, CP05b, DB08, FCCN01, Fer02b, FM00b, FM00c, GM01b, GSC07, HH01, Haw01, HP00b, HCBK05, HL09b, IS02b, JL09a, JS00, Kle03, Kör00, KV05, LLC02, Lee05, Mae05, MG08, PWW08, PF02, Pfa03, Qi07, SA07, TT07, Ten07, Ten09, WW02c, ZL08, Zie03]. Asymptotical [EY08]. Asymptotically [BD05, DMMS07, GFM05, HRZ06, Jeo00, Nak09, YW08, EJ07, EJ09, Oga09, Pou01, SK07a, XM09, YHZ05, YL09]. Asymptotics [BL06, Har09, KQ02, KD00, LS04, STP02, YCB06, ARV03, BC03a, CS02b, Tak09, Tan01a, HPS07]. attainable [DS06, DS08]. attraction [NPA06]. attributes [BS05c]. AUCs [TVYS09]. augmentation [DTW06]. August [Ano00-33, Ano00-31, Ano03-54, Ano04-56, Ano05-60, Ano06-41, Ano08-40, Ano09-41]. Author [Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02h, Ano02i, Ano02b, Ano02c, Ano02d, Ano02f, Ano02g, Ano02e, Ano03c, Ano03d, Ano03j, Ano03e, Ano03f, Ano03g, Ano03h, Ano03b, Ano03i, Ano04a, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04b, Ano04h, Ano04i, Ano05b, Ano05c, Ano05d, Ano05e, Ano05j, Ano05f, Ano05g, Ano05h, Ano05i, Ano06a, Mat01a]. auto [BG09]. auto-mutual [BG09]. autocorrelated [GP04, PK01, SYZ07]. autocorrelation [PR06, UV07]. autocorrelations [SG08]. Automatic [RV07, YL09, OK08]. automorphisms [Top01]. automotive [GPV09]. autoregression [STP02]. autoregressions [DL05]. autoregressive [AYH09, APdR02, Azo09, BRO^{+}09, Con09, GT08a, GB07, GC09, HPS07, HKPS08, KV06, Laï05, NLP06, PT08, Phi02, RBN09, SN04, VEM07, ZBD07]. autosomal [GPC01]. auxiliary [CL01, FBK03, GS07d, QW01]. availabilities [Par00a]. Availability [CX05, SS00, TLBJ11]. average [Heu08, JR03b, KV06, MW08b, Mun00, SW08, SCW09, WW05, YZZ09, Zha01b, vBLO04]. averages [Hal07, Koh00]. averse [MT07a]. Avoiding [Gol01, Sud01, BD08c]. axial [SM05a].

Back [Arn03]. Backfitting [PW06c]. background [GGS03]. bad [GS00]. Baer [Zhu01]. bagging [FH07]. Bahadur [DN03, ZS08]. Bailey [Kem02b]. Bailey-Daum [Kem02b]. balance [Aic01]. Balanced [FFH00, FHM00, FHKMS02, AGGZ01, AJMP09, AYH09, AEK09b, Bag04, BJ06, Bat04, Cao09, CCW05, CKMS02, DT05, DL07, DDM09, GKK08, GTT^{+}07, HB00, HJ02, IS01, KS07e, Kuw02, Len02, LTI00, LL08, MPGS09, ME04, MJK01, MMM06, Mut05, NN09, NS00, OK00, PCK07, PM01, Ruk07, Udd05, WRV05, WS08, Wri08, YLH04, ZCS09]. balls [BK02]. BAMS [KRV06]. Banach [DNR09]. band [HLAK09]. bandit [CP05b]. bands [EHM09, Fre08, HLW07, HHZ08, SSW09]. bandwidth [APdR02, EPdRV02, PLK^{+}06, SC05b]. bargaining [Hai03]. Barry [Lu02]. Bartlett [Bra06, GM01a]. Bartlett-type [Bra06]. based [AZ04a, AJMP09, AH03, ABK00, AC07, ALS07, ARV03, Arn09, BE00, BS04, BNP06, BTK09, Bar08, BAA08, BN09, BK09a, CYLC01, CA00a, CL09b, CvdL01, DLS09, Det02, DGLS05, DGL07, DBDG09, DF06, DFW08, Eks08, ES00, Ery03, FGT08, FWB08, FMT04, FZ03, FG02, GHL03, GNP07, GK08, GS09b, HYP02, HY08, Hlá03, HB04, HLP06, Jia03a, JL04, JMJ09, JL09b, KK09a, KP09a, KL06, Kal07, KC05, KP09b, Kot05, KLX09, Lai09, LF07, Lee01b, LSW06, Li03, LDT04, LLL09, LLFF00, LJ09, MCDC08, Mas09, MS02b, MLK09, MP07b, Mei05, Mei07a, MN08b, MR00, MMM06, MBG^{+}07, NvdL07, Oga09, OP05, PSV06, PPdCP06, Par09a, PS04b, PKZ06, PvdL04, PSW00, Raq02, SA03, SSO07, Ser04, Ste04b, SSH08, SZW08, Sun09, SS02e, Tan09a, Tan09b, Tha08, TS08]. based [TVYS09, UV07, Vie09, VG07, WL02, WM08, WLCI09, WWN05, WZ09, YHA09b, YB06, Zad02, Zha04a, Zha06, Zha09d, Zog00, dHD^{+}09, vdH05, BBdR00]. baseline [Che06c, DL03, Sub00]. basic [CLR05, Fuj02]. basis [AP05, AKI08, Bak09, GN03, KK09b, RA06]. Basu [Das07a]. bathtub [BLZ08]. bathtub-shaped [BLZ08]. Bayes [Lia00b, AJMP09, BGM03, BL02b, BL03, Che08, CG08b, DR03a, DL04, DLM02, DS01, DK04b, EL04, FG06, GS02, GS07b, GKK08, GL02b, GL05b, HT09a, HC07a, HC06b, KZ03a, KZ03b, KI00, KSK00, LGL05, LG06, Lia00c, Lia06a, Lia09, MB00a, MS06, Mas03, Mez03, NWS01, Pen00, Pen03, Rao00, SZ08a, Sen00, SH09, SVZ07, SK07b, TP03, VM04, WD04, YR03]. Bayes/frequentist [EL04]. Bayesian [Fox09, LHL09, PW06a, Roy07, AKJ09, AH03, AH00, AVGRZ00, AC06, AWL04, ALW07, Aus05, BF07, Bar03, BM02b, BM03, Ber01a, BS09, BBIAV00, BH05, BD08c, BMW09, BL08b, CG01, CLA09, CC07a, CS04a, Che03b, CJK04, CL09b, Con04, Cow03, CBC^{+}07, Dau04, De 04, EC03, EOGQ09, ES03, FD02, FED09, FR02, FI05, GK02, GH03, GML06, GHMR06, GMM03a, Gli08, GS08, GR09a, GR09b, GRV08, GG09a, GPW01, HDM09, HW08, Huz05, IOQ09, JLN08, KMPP02, Ken00, KS09a, KM04, KK09d, KV06, KS07c, Kot06, KS07d, KBG09, LDC09, LLLH05, LM00a, LL07c, Lin01a, MPS03, MW03, Mat01b, Mee05, MCM^{+}07, ML03, MBG^{+}07, NDF03, OK08, Ong05, PvZ09, PW06b, Par08, PZY09, Pot07, Qui06, RSD06, RGR09, RM07, SG08, Sas01, SC04, SB06, SK07a, SR05, SSY05]. Bayesian [SN04, TGV02, TCD08, Tat05, Tri08, VIBAV06, VG03, WG03a, WHS07, XC06, XR09, YO09, Yin09, ZG08, dL06a]. Bayesian-type [AH00]. Bayesianism [Wil03]. Bayesics [Arn03]. BCH [FS01b]. be [Bat04, For09, HK04]. Bechhofer [Fab00]. Beder [BK09d]. Begun [ZY01]. behavior [Ahm05a, ALRV01, ABK00, Ars04, BRO^{+}09, CT04, KS00, Ros07, Ruk02]. behaviour [BT04, GM01b, Ten07, Ten09]. Behrens [BMP02, DMMS07, KI08, LW07, Van05]. being [For09]. beliefs [PH05]. Bellman [Hyr07]. Bellman-Harris [Hyr07]. belt [ZF01]. Benjamini [KHC02]. bent [CLR05]. bent-cable [CLR05]. Bera [Ano01i]. Berkson [BL02a, KS08b, WC08]. Berlekamp [Xin01]. Bernoulli [Che08, CP05b, HR04]. Bernstein [BCC02, DH09b]. Bernstein-von [DH09b]. Berry [AL08, Ben03, BW08, LCM07]. Best [WB05, BR03b, BNP06, BM05a, BM06b, Che08, GFM05, Hay07, HRZ06, HC06b, Jia03a, LL07a, Liu09, Nak00, Nay00, PSW00, TS08, WA07]. Beta [GC07, HLZ05, Jon02, KvdV08]. Beta-Riesz [HLZ05]. Better [GKM07, Bos06, Che08, GL02b]. between [AL08, AMY02, Air00, CC00, Che07, DH06, DPA08, FN07, Gan00, GKDN06, GN09, GF00, GS00, KA08, KGM05, LF07, LLC09, LSL06, Liu00b, LFTT08, MMM07, PSV06, QL06, Sch04a, TVYS09, VV05, WYL09b, Zha01b, ZTQ04, ZQ05, ZQ07, ZN08, Zog00, vdH05]. Bhaskar [CGS02, HS00b]. BHHJ [MLK09]. Bi [BGG00]. Bi-embeddings [BGG00]. Bias [AZ04a, Ago09, Bat06, CL06, DGGG08, GM04, Pen04, SJ09, Sar06, YSF03, Zha06, AR05, AK03, BFGV08, BI02, BC07, BS01, But06, CP05a, CB06, EN02, FM00b, FM00c, GK01, HH09, LP09, NdC08, UF09, ZC03, vEKO00]. Bias-corrected [Pen04, Zha06]. Bias-reduced [DGGG08]. bias-robustness [AK03]. biased [Ant08, Bon08, CV09, Che06d, CA00b, DI09, DH09c, Efr04, Fan00, Gil04, LP01, MdUÁ04, Ong05, WS09, Wie00]. Biases [KSL03, DW06, GHJ^{+}07, ZMJW08]. BIB [Cha01b]. BIBD [MR07, RM05a]. BIBDs [BT04, LZ00]. Bibliography [Lah00, PS02]. BIBRC}s [Udd07]. BIC [CG06]. Bickel [Ten09, LJ07b, Ten07]. bifaced [DG01]. bijective [Fer02a]. bilinear [WZ01b]. bimodal [BK05b, EOGQ09]. binary [Ber02a, Cho07b, CG00, Duc00, ED07, FS01c, GIA04, GACS07, Ger05, Gov05, Gov09, JL09b, KKM09, Lin08, LLC09, LZ09b, MPP07, MS01, MP07c, NA09, NdC08, PP02a, PH05, PG07a, PMS^{+}01, SR03, SR05, SBKF01, Ste00, Yan05, Zha01b]. Binomial [LBM00, YLH04, AA08, BL05b, BL08b, CD06, Cho07a, DR05, LL07a, LM03, PP09, SD01, Sta09a, SNM08, TR08, Wan06, WWEL08, ZBZ00, ZTQ04, ZQ05, ZQ07, ZW03b]. binormality [GG09a]. bioassay [BK07, DTM08]. bioassays [SR03]. bioequivalence [FCM07, MW08b, WH01]. bioinformatics [Sen08]. biological [Fin09, FL08, Rao02b]. biomarker [Sim08]. biomarkers [CBC^{+}07, TVYS09]. biomass [BF07]. biomedical [DDvdL08]. biometric [RO05b]. biotechnical [CDB^{+}09]. bipartite [MJK01]. biplot [MRaPGVVV02, PLLS08]. Birth [CKS07a, Dau04, KHL07, KM07c]. birth-death [KM07c]. Birth-multiple [CKS07a]. birthday [Das05]. bispectra [HW05]. bispectral [Soe02]. Bivariate [HNW08, ASD09, BES05, Bon09, CH05b, GHMR06, Gup03, HP00b, HL09b, Ili03, Ioa07, JB08, KP05, LB06, LN03, LB08b, LT09b, LT09c, MP08, MP07e, MW05, PQ07, sS06a, SY08a, WG03a, WC09b, ZM06]. Blackwell [FM00a]. Blackwellization [MW00a]. blind [Gov05, LZ09b]. Block [Chi03, DS02b, dSBFG07, Hai00, Kir07a, ABZ00, AGGZ01, Bag04, BDH05, CDD01, CCW05, DDM09, FKL^{+}02, Gan00, GRP02, HM02b, HJ02, ITA^{+}09, IT10, Jac02, Jac03, JM08, KM01, LX07b, LS02, LR02, McL08, MJK01, MMM06, Mut05, NB08, OJKM02, PM01, RD08, Udd00, Udd08, WH05, Yin00, ZC07, ZC09, ZCS09]. blocked [AZ04b, CLLZ06, Dey09, Jac06, LJ07a, McL08]. Blocking [Jac09, AYZ06, AKJ09, Tan07, XL06a, ZP00]. blocks [SS01]. blocksize [MS02c]. blockwise [Efr07b]. blockwise-shrinkage [Efr07b]. BLUE [IPS08, Liu00b]. Blum [MW05]. BMAP} [KKLD07]. BMAP}/ [KKLD07]. board [Ano07p, Ano07o]. boards [Cha02b]. Boltzmann [Nek00]. Book [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Cha01a, Cha02a, Con02a, Con02b, Dea00, Ko00, Lah00, LHL09, Lee01a, Lin01b, Lu02, Mas01, McF01, Nay01, Roy07, Sin06b, Wes00]. Boolean [And05]. boosting [DT08]. Bootstrap [Con04, FI05, KP08a, LHS00, MFSW07, MR00, RMS02, VFVF02, APR02, BR03a, BC03c, Cha02c, DR08a, DG02, ENG00, GRC09, HLM08, Hlá03, HS08b, IJZ09, MM06, Pal09, PP02b, Shi08, vdLP03, Lah00]. bootstrapped [ALRV01, Bab04, GH02]. Bootstrapping [CA00a, LGMPS03]. Borror [Atk09, Bor09, Jon09, Khu09, Luc09, Par09b, Pie09]. both [DR05, HT08, MY00, Rao02c, Tom09, WDE05]. bound [Ago09, Ben03, DNK02, Din03, Efr07b, Efr09, FS01c, MS05]. boundaries [GJ03, GM05, KdMSY09, Shi08, Zac05b]. boundary [Jan04, KA06, Tam02, ZK09a]. Bounded [MZ07, NH09, Pro06]. Bounding [Blu09]. Bounds [BCP03, LB00, LBM06, PR04, AGK00, BLBR00, BBW06, Bie08, BW08, Dan05a, DZ06, DL07, FTY08, Fin03, GO03, GBC^{+}05, GR08a, IVM03, KOS05, KKM09, Kle03, Lia00b, LBCM06, LBR08, Pap01, Pet07, RRS06, Raq04, Raq05, Raq07, Ryc01, TSN07, ZN08]. Box [But06, CL06, FWB08, LLLH05]. Bradley [CH05a]. Bradley-Terry [CH05a]. branching [HNW08, Hyr07, KD01, Qi07]. Breakdown [MN03]. breakthroughs [AG07]. breast [FPvdVvH06]. brief [CD01]. broad [dC02]. broken [GB04]. broken-line [GB04]. Brownian [CGPP06, LMD^{+}01, LL05a, LS01, Tan08b]. BSAs [ZC07]. BSECs [ZC09]. Budget [TV04]. budgets [FFP09]. Buehler [KL06]. bulk [AC06, SA02]. Burden [Sen08]. Burr [SP05]. Bush [KKK03]. Bush-type [KKK03]. busy [Aga02, SA02].

C [Khu09, Par09b, Pie09, ALW07]. cable [CLR05]. calculating [AIK00]. calculations [Shi05]. calculus [AA08]. Calibrated [FMT05]. calibration [BBIAV00, DK04b, DL09b, GB00, JI07, KKM01, MS02b, MS02a, RMMA07, ST06, TSR03, TWH^{+}05, VG03, WMES07]. Cambridge [Con02b, Lin01b]. Cameron [Con02b]. can [Bat04, For09, GHJ^{+}07, HK04]. cancer [CLS09, FPvdVvH06]. Candelabra [MRC02]. Canonical [TL08, FHW08, HMW04, HLH09, SG05]. cap [ZF01]. capability [Ahm05b]. capacities [AK03]. capacity [BK02, LA08]. capital [Ban03]. caps [DÖ01]. capture [CYLC01, HH07, WHS07]. capture-recapture [CYLC01, HH07, WHS07]. CAR [Wal04]. carbon [FFP09]. carcinogenicity [NHZ04]. care [Ber01a, Gas07, Lin01a, Mat01b, Sas01]. Carlen [KLR07]. Carlo [And05, Aus05, BM05b, CG01, DEH00a, DPR06, MW00a, Tan08a]. carried [SJ09]. carryover [AH02, BS07b, HY08]. Cary [McF01]. Case [CZ09c, AL08, AFF09, CC04, CC06, CC08, CLR05, DH09a, ENAZ09, FM05b, Huz05, LWL^{+}09, LP01, Lia06a, MJS00, Mun00, PZ08, QRC01, QT07, RD08, SW01, Yan00, YHZ05, Zha06]. Case-cohort [CZ09c]. case-control [CC04, CC06, CC08, LWL^{+}09, LP01, PZ08, SW01, Zha06]. cases [But09, KX07, LCM07, PQ09]. catalog [LDT04]. catastrophe [CKS07a]. catastrophes [KM07c]. categorical [BI02, BdM09, BG09, FR07, MY00, PM08a, SS02a, WC07]. categories [Bor00]. category [Ber01b]. Caterpillar [GO07]. Cauchy [Fre07a]. causal [KKM09, NvdL05, NvdL07]. causation [SK08]. cause [TKP06]. cautious [DTM08]. cell [AAJ00]. censored [AW09a, AP07, Alv05, BX07b, BX07a, BD08a, BD09, BH09a, BKZ09, CG01, CC00, Che03a, CL01, CY04b, DMN01, EHM09, FZ06, FI05, GHL03, HSV09, HN04, IJK08, JI04, Jin07, KHP06, Kun07, Let08, Li07, LWB09, Lop09, MdUÁ04, Par00b, QJ00, SY08a, Sub01, SZW08, Sun03, TR05, TKP06, WYL09a, Xu09, Yin09, ZZF09, ZP04]. censoring [AH01, BE06, BS08, BH08a, BTK08, Bha07, Bor04, Bur08, Dup09, Fer04, GB09, HCBK05, LN03, LB08b, NB05, Raq05, sS06a, SGU05, WJ08, WLC08, Yan00, ZjLS05, vEKO00]. Censorship [dUÁGMCS03, Bel07, Dik00, DGS05, IPdUÁ08, KS07b, Li03, Xue04]. census [DM02]. center [Bos06]. centering [GHJ^{+}07]. Central [AAJ00, GM05, KV02, AEK09b, DGS05, Hil09, LY09, MT00, Saj00, SCW09, SJ08, TL09b]. centrality [FPR00b, LY09]. centroid [JMJ09]. centuries [FR07]. certain [AK03, CP02a, Kem03, Kem05, MJK01, PS02, Zha01a, ZY02]. chain [CT09, DZ06, FCS05, Kem05, MW00a, Tan08a]. chains [ANN08, AT02, CT02, FFH00, FI05, TL09b, WW02b]. Challenor [Cha09]. Chambers [LGMPS03]. Change [GR01, GV05, HN04, Sri04, AH00, BD05, Din03, DW06, Dup09, Fer01, FJT09, Gen01, GHH00, Gua07, GS09b, Jar07, Kir07a, MGB05, NJP05, PK06, RS04a, Ruk02, SY08b, Tan09a, Tan09b, Wil03, Wu06]. change-of-variance [Gen01]. change-point [BD05, Fer01, FJT09, GS09b, MGB05, Ruk02, SY08b]. change-points [Gua07]. changeover [JR03b]. changes [AEK09a, BGH09, GP07a, HS00a, HHKS04, HH05, HS00c, HPS07, HKPS08, NJP05, Vex06, Viš03]. channel [Kak06]. Chao [Ber05]. Characteristic [TL09a, Bon08, KT08b, LZ09a, Mei05, PZ04, WM08]. characteristics [NA09, ODY08]. Characterization [GK08, Qin03, SS02e, ZB09a, AK03, ARS01, CS03a, GK00, Hos06b, Kag01a, Kag02, MT02]. Characterizations [BAA08, BN09, BC09, Kem03, MO07, PNRdA03, Pap05, SSH08, AT09, BS04, CH05b, GM01c, HC07b, LH08, MML02]. Characterizing [RS04c]. charts [AKN04, AKN06, AK08, BTK09]. Chaudhuri [QRC01]. check [LS08]. Checking [NW05, SW09, CP02b, KN04, KS08b, PP09, SS03, WPV04]. chemical [BW05]. chemistry [Sri07a]. chemotherapies [VdBS01]. Chen [BC03a]. Chen-Shapiro [BC03a]. Chernick [Lah00]. Chernoff [Ano05-65, EOPR05, Zac05a]. Chernoff-Savage [EOPR05]. Chi [LY02, dUÁGMCS03, Bet01, LY09, Oga09, SGK09]. chi-square [Oga09, SGK09]. Chi-squared [LY02, dUÁGMCS03, LY09]. China [Che00a]. Choice [PvdL04, AKN06, APdR02, ASC05, BS05c, DMMS07, Man02, RW06, SB04, VG03]. choosing [Gan00, vdH05, War08]. chop [HH08]. Christine [Bor09, Jon09, Khu09, Luc09, Par09b]. CHSS [BCN04, LGV09]. circle [YW04]. Circuit [Goo08]. circulant [CL08]. circular [AYH09, BS09]. claim [BJZ09, KA08]. claims [TW03]. clarifying [YSB09]. class [AAEK06, AJMP09, ALS04, AK05, AVGQ05, BM06c, BC09, BP02b, CD06, Cha05, CIC08, Dan05b, DF07, EOGQ09, Hod04, HS00c, IS01, Kal07, MS06, MN08a, Oht00, OS06, PP06a, PSG07, PW02a, RL09, TAGF07, TMP06, Yun00, Zha01a, Zhu01, dC02]. classes [AM04b, CZ09a, Che06a, FS01b, Gil06, GH01, Jon06, KS06, PS02, Swi02, Udd07, UG02]. classical [AL08, ADW03, AK05, ACS07, FCW00, GN09, GH00, TT09]. Classification [BM08, But05, FPR00a, LT96, LT00, LR09, CFS05, DDvdL08, GPL09, GHSN00, IS01, JMJ09, KHL07, PG07a, Par09a, Qu06, RK05, YMJ08, vdW04]. classifications [HH01]. classifier [Zaf02b]. classifiers [Sim08]. Classifying [SM07]. clear [CLLZ06, YLLZ06, YB08]. Cliff [Krä05]. clinical [BYW07, BBP06, BD08c, CH00, Che03b, CLS09, CG00, EMP06, FS01a, Gov05, Gov09, HA08, LMD^{+}01, Lee01b, LC09a, LP06, LZ09b, MBA^{+}08, MPS07, Pot07, PMS^{+}01, SCZ07, Wie05, WY07b]. clinician [KTB01]. clinicians [Ber01a, Lin01a, Mat01b, Sas01]. clones [FI05]. close [For09, Wal04]. closed [CYLC01, Gre08, Sch05a]. closed-form [Gre08]. closest [CN05]. CLT [KS07b]. cluster [CP03, EN09, FM00a, GKM07, KK09c, Rob09a, dE02]. cluster-size [Rob09a]. clustered [GACS07, JL09b, LL09]. Clustering [Ser09, KP09a, MT06, QT07, Qui06, SW05, TCD08, vdLP03]. clusters [NMB02, SA03, SW05]. co [LT02]. Cochran [HBK03, KL07, TS06]. Cochrane [ST00]. Cochrane-Orcutt-estimator [ST00]. Codes [BRS00, BGL01, CGZ00, CK01, Col01, DR02, FS01b, FS01c, GK05, HH02, KdR01, KNTR01, KM00a, RA06, ST08, SvTW00, Zha01a, Zha01b]. coding [PR08b]. codings [PR08a]. coefficient [BM06d, CLHY01, DR04, GG08, GC09, GK00, HB00, Hon04, Hös08, Len09, LFRL07, Lu08, LNS07, Nag09, PGLS08, PY06, SM05b, Wan09, WM08, WW02c, WGCI09, XY06, Zha09a, ZM06, ZBD07]. coefficients [Ahm02, BL02a, Cao09, DMP04, GTT^{+}07, GGS03, KX07, KV02, KV05, MTC05b, SG05, Sub00, VV08, YW09c]. coherent [DM08a, KS07a]. cohort [CZ09c, GPC01, IFC07]. coin [Ant08, Che06d, DH09c, LZN06]. coin-weighing [LZN06]. coincidences [MO07]. coins [Len02]. cointegration [JS08, Vin02]. Colin [Con02b]. Collapsibility [VV08, GGS03]. collision [Hei00]. Colored [BRS00]. column [CM03, CG08a, MU03, MP07f, Udd05]. columns [HJ02, Mut05]. COM [RdCCB09]. COM-Poisson [RdCCB09]. combination [GG08, Hay07, LJC07, TR08]. combinations [DFW08, KK02, Ma00, MZ07, Wit01]. Combinatorial [Ano09-47, GK05, Sri05b, BT04, KD08, MT00, SK06, Sza02, TAGF07, WC00]. combinatorics [Ano05-40, Csá05, SA02]. Combined [BKZ09, EKK06, MS02b, TVYS09, Wu03]. Combining [GT09, GP03, HML00, HY03b, Kis02, YHZ05, CGY07, ES03, HBK03, Pad06, SJ08, YR03]. Comment [Tho02]. Comments [Cha09, LZ09b, MH09, Smi09, Sud01, dL06a, CL01, MRU00, Atk09]. common [BPR09, FL08, LRZ02, LL05b, LLW07, MCDC08, MS07a, MS07b, Nak00, OM07, RBV00, SR02, TK06]. communication [HDM09]. commutative [GPR03a]. compact [Cav06, Maj07]. Comparability [Ver02]. comparative [GB00, GC07, QT07, SR03]. compared [BR00, FHW08, Kwo01]. Comparing [DNN07, FS01a, GB00, SS01, Zha07, DHV01, GRP02, HML00, HLW07, Jac02, Jac03, KT04, Lia03, Llo05, MA05, MS02e, NPF09, SBKF01, WMK07, YP07, YS08]. Comparison [Atk09, BB00, BS06, CLA09, DL03, DDS00, HC06a, JL09b, MRaPGVVV02, MR01, MK08, PdSPS05, SOUS03, ACBM09a, ACBM09b, AGP09, BM02b, Bor09, BDH05, BH05, Cha02c, Cho08, CL09b, CCS07, Cow03, GACL08, GA08, GS00, GJ05, Goo09, GGHS04, GS07c, GK06b, Jon09, KA08, Khu09, LK00, LL07a, LLC09, LLP02, LJZ^{+}07, Luc09, MMM07, Nak09, NA09, NdC08, Pal08, PVHK09, Par09b, Pie09, Qua08, Rob09c, TWH^{+}05, VIBAV06, WYL09b, YHZ05, YSB09, vBP03]. Comparisons [Air00, GT02, CH05a, CC00, FR01, GDL07a, GDL07b, GGS09b, GK06a, Hay07, HR04, HZ06, KK06, KHC02, LT04, Liu00b, LMH00, LK03, LC09b, Mae05, MT07b, MGG09, Nak00, SN08, SSC07, TZ00, WW05, ZB09b]. Compatibility [ACS04]. Competing [Ano06b, BM06e, CMN06, LL06a, BH08a, Bed06, DH09b, DL06b, DG09a, EKMS04, EKT06, FPvdVvH06, KB00, Llo05, TKP06, WG03a]. complementary [Jon02, Qin03, SSY05]. Complete [FHMS02, MP07c, ACS04, Chu01, DÖ01, FWYY03, FR00, FR08, FHJ^{+}04, Mi06, PK01]. completely [MS02c, PW09]. completeness [IP09]. Complex [MDM06, BR02, BL05c, Cha01c, CP02a, CS05, DL00b, Kon07, PR08b]. complexity [BB03b, DEH00a, HP00a, TSN07]. Component [Che03a, FHL03, CD02, DNN07, Hag06, HS05, KK09b, LS04, Rei05]. Component-wise [FHL03, Hag06]. components [Asa06, BW03, BPR09, Che02, Che06b, DS03, Fou00, GS09a, GHSN00, GPV09, KSK00, KN03, LT02, ML03, RH08, SSvRM03, WD04, WYL09b, ZB09a, ZEB07]. composite [AEK09b, BD05, LT02, PCK07, Sta09b]. Compound [BDZ07, Per00, Web07, Zac05b]. compromise [EL04]. compulsory [CS05]. Computation [AFW05, Hei00, HH08, Wit01, vdWD01]. computational [GPR03a]. compute [Ars04]. computed [WL01]. computer [BS06, JCS05, MMM06, RWF04]. computerized [cICY04, Kuh05]. Computing [ND06, BES05, Ken00, Lee01a, LZ07, TMM09]. concave [Pen04, SH02]. concentrated [HK04]. concentration [KRV06]. concept [Bri08, Vin02]. concerning [DLOR02]. conclusions [Gol01, Sud01]. concomitant [Cau05, CS03b]. concomitants [BE06, HN09]. condition [Dav05, GFM05, Pap05]. Conditional [CCP09, FSS07, FS02b, Has09, KR07, Air00, And06, ACS04, BJPZ08, CD07a, FY05, GB09, GSY03, GPL09, GK08, Heu08, IPdUÁ08, KK09a, KS07a, Lee05, LT09a, Nie01, PW02a, QW01, Qua08, Rob09b, Sub00, Sub04b, WPV04, WY07a, ZB09b]. Conditionally [FM05b]. conditions [BLV00, DdHL06, El 04, GKMS02, Kem02a, Nie02b, SC05b, XR09]. cones [HLZ05]. conference [Ano09a]. Confidence [BR03a, BL05b, DH03, FZ03, Fre01, Fre02, FH05, GTT^{+}07, GRC09, HHZ08, KG09, KLX09, QRR08, SZW08, TW05, TVYS09, AP07, AKK09, BLBR00, Bur07, BQI06, Cai05, cIC01, Che06b, CG00, CW05, DK04a, DBG06, Din03, Dub07, EHM09, Fre08, FMN09, GT08a, HK05, HLW07, HLAK09, HCW06, JI07, KP08a, KM03, LF07, LL04, LSW06, LMH00, Liu04, LC09b, LM03, Mae05, MS02b, MS02a, MR01, MS02d, OD06, Par08, PQ07, Pfa00, RHL04, RM05b, SSW09, SN08, Ser09, Sin06a, Sri01, ST05, TR08, VG07, Wan05, Wan06, WW06, WWN05, XM09, ZW03a, Zha09d, ZQ05, ZQ07, ZEB07]. configurations [FCW00, Gly00]. confirmatory [DTM08]. confluent [Kem05]. Confounded [CG08a]. confounders [WGCX09]. Confounding [GPR03a, LuV09]. Congdon [Roy07]. conjecture [LL07b, Qiu01]. Conjectures [Bed98, Bed00, BK09d]. conjugate [Lij03, SZ06, YO05b]. connected [JTS02]. Connection [QL06, Iva06]. Connections [BDHS06, KMB08]. connectivity [Wyn08]. Connie [Bor09, Jon09, Khu09, Luc09, Par09b]. Consecutive [GGS03, CH08, JHCH05]. consecutive- [CH08, JHCH05]. Consequences [ST06]. conservative [Che04a, JI07, SN08]. Considerations [Sch07a]. Consistency [BK07, KMV05, Mey06, PWW08, SL00, AGHMPV07, BGM03, BLV00, CC07b, Cha00, Che00a, Che04b, Cho07b, DL04, DR08a, GB07, Heu08, JI04, Maj07, Sag05, TG07a, TT09, TG07b, TC08, WPV04, XR09]. Consistent [BWA06, CP07, DS01, GB04, JL04, Mei05, MSW09, Sch05b, Soe02, Det02, FL03, GB00, Hag06, IL04, MN08a, SW07, SW05]. consortia [RDS08]. constant [Ber04, BW05, CSW02, FS01c, RD08, WJ08]. constants [BNW05, BS09]. Constrained [CC02, GKK08, LT08a, Bir09, GKK04, LB06, Sen07a, TV04, UB07, XW08]. constraint [MS05, SH05, WO00]. constraints [AC02, AH08, ENAZ09, FSS07, HS07a, Hos07a, Hua02, MTC05a, PPZ02, Páz02, SS02b, SSB02, TSH06, Wil07]. construct [HLW07, RHL04]. constructed [Bul07b, NB08]. Constructing [Cua09, LBS06, MTC05a, Mor09, Atk08b, HP07, JCS05, Neu00, NS00, Roy02, Tha08]. Construction [Bag04, CG03, DR01b, LZ00, LRD07, LLZ07, Ma09, MT06, MW02, SST02, ySD03, yS05, TD03, Tan07, ZDLZ02, AG04, BJ06, BW01, BW00b, FLM00, HLAK09, Huz05, Jac02, KM08, KMS09a, LM00b, LHZ03, MS02b, MRC03, Pro06, ySK05, XM09, YMJ08, YLS00]. Constructions [Bur01, HJ02, MJK01, SK06, Chu01, CD01, DÖ01, Ngu08, OJKM02, vT01]. cont [Ano03x, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano05-38, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. containing [HP07, YLLZ06, YB08]. contaminated [CC06, LY05, YC02]. contamination [AKK09]. Contemporary [FTY04, Das05]. Contents [Ano01j, Ano01k, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g]. Contents [Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m]. context [DdRB00, DGJ06, IP09]. contiguous [FJT09]. contingency [AC02, AG07, Bor00, DTW06, KL07, KB08, RZ09]. Contingent [TW03, GT06]. continual [O’Q06]. continuation [PZ08]. continued [Ano07p]. continuity [AGH07]. continuous [BB04a, DH00, FN07, GS07a, GACS07, HH07, IBL06, Iva06, KD01, KZ03a, KR07, MY00, SL05, ZH09, Zog08]. continuous-time [HH07, SL05]. continuously [UB07]. contoured [BZ06, DGLS05, DGL07, Kon07, PC02]. contours [ACS08]. contrast [ALS04, ALS07, LC09b, Wyn08]. contrasted [BR00]. contrasts [NN09]. contribution [Csá05, NPA06]. Contributions [GG01, EKL07, Zha00]. control [AKN04, AKN06, AK08, AIK00, ALW07, BTK09, BM05a, BDH05, CS04a, Che08, CC04, CC06, CC08, CS02d, DH09a, DHV01, GS00, GP04, GR08b, GL02b, HML00, LWL^{+}09, LP01, MS02e, NaX03, PZ08, SW01, SR05, SS01, YP07, YS08, Yek08, Zha06, Zha07]. controlled [JI07, MS09a]. Controlling [KTMS04, SZ08a, FS07, KHC02, LJC07, Sar08, Tro00, YSB09, Sar04]. controls [GRP02]. controversies [Sen08]. Convergence [Ars04, EPdRV02, LGL05, Nie07, Par09a, RGR09, SV00, AGK00, Can06, CL06, DFS08, ETS02, GB09, GH02, HSV09, KKW09, KvdV08, LT06, LG06, MI08a, MP02, Pen03, SC05b]. conversion [TW03]. Convex [DKSZ07, Ma00, AAEK06, BC03c, Fer02a, GK08, KS02b, PK01, WG01]. convicted [PW06a, PW06b, dL06a]. convolution [KRS07, MW00b, Sen00]. convolutions [Bon09, SW07]. Cook [Atk09, Bor09, Jon09, Khu09, Lee01a, Luc09, Par09b, Pie09, DGGF04]. coordination [MS09a]. Copula [KP09b, ANN08, BNW05, Cua09, KK09a, MJD09, Rüs09]. Copula-based [KP09b, KK09a]. Copulas [Bed06, NK09, RL09, Zez09]. correct [HH01, XM09, vdH05]. Corrected [CG00, CW05, Liu04, Pen04, YL07, Zha06]. correcting [BGL01]. correction [Bat06, KA06, Sub04a, UF09, YSF03]. correlated [AYH09, DKSZ07, GN03, GG09b, HY08, HA08, Hur05, JM08, JL09b, KX07, KM00b, LH00, LTI00, Lin08, MEC04, MW08a, MJS00, SBKF01, SA07, TRJ07, TL07, Udd00, Udd07, Udd08, VFVF02, YC07]. correlation [Cha03a, DH06, DHT07, DF00, DPA08, GHJ^{+}07, GGS09a, GL09, GTT^{+}07, HB00, HLH09, Krä05, KBG09, KX07, LR09, LBCM06, M.04, Ma09, Sch07b, SG05, TL08, YW09c, ZM06]. correlations [FG02, Par02]. Correspondence [Beh08, Cua02, NBH05]. Corrigenda [Gov09]. Corrigendum [Ani10, EN07, FCCN05, FR08, LT00]. corrosion [RWF04]. coset [Xin01]. coset-leader [Xin01]. Cost [DTM08, LdC06, SH05]. Cost-cautious [DTM08]. costs [HS07a, TW03]. Count [CC07c, Con02b, KKB09b, MN08a, Sah08]. countable [KSS01]. Counter [Lia00b]. Counter-examples [Lia00b]. counterargument [Goo08]. counterexample [BK09d]. counterexamples [HMS04]. counting [Gil06, GS09b]. countries [OO08]. country [Ver02]. Course [Sin06b]. Court [Goo08]. covariables [Wan08]. Covariance [FZ00, GP07a, Swi02, ATS02, BSP02, BGH09, CSW02, Fue06, IPS08, Jon04, KZ09, Kon07, KvR08, LW09, Ma03, Oga09, QZF09, Sch01, SL09, Vas03, VKO00, WC01, YA09, YW09b, ZS05]. covariate [Dup05, DDM09, KI09, SN09, Shi00, YXC09]. covariates [AL09, BV01, BV03, DMS03, FBK03, Lar03, OTB06, PW06a, PW06b, Sch00b, Sun06b, WG03a, WGCI09, YL07, ZR05, dL06a]. coverage [BLBR00, CL06, HYCY03, PPM09]. coverages [Pal09]. covering [WC09a]. Cox [BN03, CL06, CMN06, DL09b, FWB08, KI09, LLLH05, Sen02]. crack [CLE09]. Craig [OO08]. Cramér [IL04]. Cramér-von [IL04]. credal [Zaf02a, Zaf02b]. Credible [ZW03a, EC03, FED09, Fox09, GH03]. criteria [BB03b, CLA09, CN07, FTY08, GP07b, HC06b, KC05, KK03, SFT01, SR03, Tha08, VG03, WH01, WWEL08, vdH05]. Criterion [ZT04, CJK04, DZ06, Jac04, MLK09, Ots08, OTB06, Qu06]. Criterion-robust [ZT04]. critical [CS06, LSL06, Sar00, vBLO04]. critique [Goo08]. CRM [CLS09]. cropping [PGS07]. Cross [GRR07, Sen01, ÅvR06, BJ06, BJ08, DLM02, EPdRV02, GD03, KM00b, RX03, RW05, WJ09, vEKO00]. Cross-cumulants [GRR07]. Cross-over [Sen01, ÅvR06, BJ06, BJ08, KM00b, RX03, WJ09]. cross-sectional [DLM02, vEKO00]. cross-validation [RW05]. cross-validatory [EPdRV02]. crossings [HB04]. Crossover [CH00, HY08, BS07b, GF00, GC05, YP07, YS08]. cryptographic [RSS04]. cube [Ker04]. cubic [LNS07]. cumulants [GRR07, IVM03]. Cumulative [IFC07, AZ07, DL09a, EKMS04, EKT06, GLS09, PP06a]. cure [KY07, RdCCB09]. cure-rate [KY07]. current [DLS09, GKM07]. curse [Pav03]. curse-of-dimensionality [Pav03]. curve [ÅvR06, Blu09, CL02, GG09a, KZ09, Moj07, RV03, TT09, WYL09a, WC01, Wu00, YW09b]. curves [BBSZ05, Cha03a, DKSZ07, HHZ08, KS03b, LLP02, NPF09, Oye03, PZ04, VBF03]. cusp [DK03]. CUSUM [Din03, DW06, LJC07, Wu06]. cut [JTS02, Sul05]. cutaneous [RdCCB09]. cycle [VdBS01]. cycles [LT08b]. Cyclic [Mis02, ST08, BM01c, CW07, FS01b, KMS08, KMS09a, LMM01, LZ00, SV01b, ZC07, ZC09]. Cyclically [LMM01]. Cyclicly [Bul07b]. cylindrical [Udd00]. cylindrical-block [Udd00].

D [McF01, Pie09, BRO^{+}09]. Damage [Ano09-48, AFF09, Fin09, PP06a]. Dantzig [PPX09]. Data [AKN06, Con02a, DTW06, HF00, Jup09, WBvdL07, AP07, AJMP09, ABQG09, AC04a, AFW05, ARV03, BBKN07, BZZH09, BX07b, BX07a, BH09a, BGP00, BF07, Bat04, BZ06, Beh08, BI02, BCVV09, Ber02a, BG09, Boc07, BGMPG09, Bor00, BH05, BPAK04, BBR04, BL05c, CV09, CG01, Can06, CN02a, Cha01a, CC00, Che03a, CGY07, CIC08, CZ09b, CC04, CC06, Con09, CL01, CY04b, CA00b, Dan05b, DLM02, DLS09, DG04, Dav05, DFG04, DKSZ07, DMN01, DHT07, DS06, DS07, DDvdL08, Efr04, EN06, EN07, EZ09, EHM09, FMT05, FTY04, FR07, FS07, FS02b, GHJ^{+}07, GKM07, Gov05, Gov09, GPV09, GHL03, HY03a, HCS04, HMW04, HYP02, Heu08, HT08, HC06a, HLP06, HSV09, HN04, IJK08, JI04, JT01, JMJ09, Jin07, JL09b, KP09a, KK09b, KN02, Kir07a, KHL07]. data [Kle03, KKW09, KKB09b, KTMS04, KS07c, Kra09, KG05, KLE^{+}04c, KB00, KN03, LL09, LSL08, Lai09, LL06a, LNO08, Lar03, LH00, Lee01b, Leh09, LR09, LX07b, Li07, Lie06, Lin08, LLC09, LWB09, LW09, LY05, LP06, Llo05, LZ09b, MMM07, MT07a, MEC04, MJD09, MS01, MW00b, MSW09, Mi06, MP07e, Moj07, MH00, Muk06, Mus02, ND05, NK09, PW09, PFF00, PM08a, Par00b, PW06c, PG07a, PLLS08, PZ08, Pre07, PMS^{+}01, QJ00, QZF09, RV07, RL06, RDS08, Rob07, RdCCB09, RK05, RO05b, Sah08, SS02a, SSO07, SY08a, SYM05, SSY05, Sub01, Sun09, TWH^{+}05, TKP06, Tsa06, TZ00, VV05, Vie09, WHS07, Wan08, WYL09a, WL01, XW08, Xu09, Yan05, YXC09, Yao08, YL09, Yin09, YZ09, Zaf02a, ZKW01, Zha06, ZZF09, ZjLS05, ZQ05, ZQ07, ZZF08, ZMJW08, Con02b, Ko00]. Data-adaptive [WBvdL07]. Data-driven [HF00, Jup09, FS07, KLE^{+}04c]. data-on [Dav05]. databases [ABQG09]. Daum [Kem02b]. David [Ko00]. DD [Bie08]. DDA [Bie08]. De-aliasing [BY09]. Dean [YHA09a]. death [BD04, Dau04, DB01, KM07c]. December [Ano01s, Ano03-52, Ano04-60, Ano05-61, Ano06-46, Ano08-42, Ano09-44]. decision [Bul07a, EL04, FG06, Fou00, Har02, LK03, RGL07, Tha08, VBF03]. decisions [BBP06]. deck [SR02]. decoding [Xin01]. decomposable [CL09b, HT09a]. Decomposing [FHJ^{+}04, GC04]. Decomposition [KDT06, AM04c, HF00, Kak06, WYL09b]. Deconvolution [ZK09a]. Deconvolving [YC02, Fot00]. decreasing [Raq07]. deductive [GP04]. Default [KI00, CC07a, ML03]. defense [PW02b]. defined [AK03, ACS08, Bor00, DS07, Pro04]. Defining [But08, CWW03]. Definitions [Sei01, vdVD09]. degenerate [Bed06]. degradation [BBKN07, Leh09, LuV09, Yu06, Ano09-48]. degree [Kle04b]. Delannoy [BS05b]. delay [DG09a, KV07]. delayed [HOS06a]. delays [KS07b]. delete [MG08]. delete- [MG08]. deleting [PW09]. deletion [DDvdL08]. deletion/substitution/addition [DDvdL08]. Dennis [Lee01a]. Denoising [KRV06]. densities [Efr09, KvdV08]. Density [ANR00, AdL05, AH03, AH01, BCC02, Bir09, Bla06, BL05c, CS02a, Che06a, Che04b, Che05, DK04a, Efr07a, EN02, Fot00, GRU01, HYP02, HSV09, KA06, KLA08, KL05, LOPV06, Li03, Lou00, MdUÁ04, MCDC08, MSW09, MPS09, PLK^{+}06, Raq07, SW07, Skö09, Soe02, TG07a, TG07b, Ton07, Wan08, Wit01, Xue04, YO09, Yan05, YC02, ZK09a]. departure [ST06]. Dependence [HC08, ANN08, Ben03, BdM09, BSW08, CS06, Cíz06, Dan05b, FT09, FARMA08, GKDN06, Hil09, HMS04, IL08, LLLH05, LSL06, LZ05, PQ07, RO05b, Sch04a, Wei08, XL08b, Zog00, Zog08]. dependencies [Gen01]. dependency [GR08b]. Dependent [KS08a, And07b, AL09, Boc07, BL08c, Can06, CR08, DH03, EZ09, EPdRV02, GB09, Hai07, Hil09, KOS05, Kir07a, Koh06, Kul09, Let08, LJ09, LP06, ME04, Pap01, RL06, Ryc01, Sch07b, SV00, Ste05, Sun06b, TT07, Yek08]. depth [AMY02, CT04, Mas09]. derivation [CY04a, Lin07]. Derivative [YSH03, Cai02]. derivatives [LC01a, PQ07]. derived [AAEK06, MA05, SSO07, Sri05b, Sri07a]. description [CGG09, SV01b]. descriptive [Ser04]. Design [Ano09-47, Atk09, BB01a, Don04, GGHS07, HKP08, LDT04, RW06, WWEL08, Aic01, ACBM09a, ACBM09b, Bar08, BM02b, Ber04, BDZ07, BW05, Bor09, CN02b, Che06d, CKM00, lCV06, CD03b, CC02, DH00, DLRP06, DB01, DDM09, EOPR05, EKTR00, EKK06, FL09, FG06, GMPZ06, GTT^{+}07, GJ05, Gli08, Goo09, Gra09, HK04, HM02b, Hir02, HW08, IFC07, JCS05, Jon09, KM08, Kee05, Khu09, KKL05, Koh03, Koh08, KS09b, LRZ02, LP06, LLZ07, LHZ03, Luc09, M.04, MP02, MPS07, MCM^{+}07, MU03, MR07, MP07f, MBG^{+}07, NaX03, OP04, Ots08, OGACR08, PK06, PCK07, Par09b, PRBAC08, Pie09, Pro02, Qu06, RM05a, Rob09c, RVZRZ07, Sel00, SK02, TV04, TBM04, UB07, XF00, YCB06, Yua08, ZN08, Zho08, ZS05, Zie03, dHD^{+}09]. design-adaptive [Aic01]. design-based [Bar08]. designated [KL06]. designed [BL02a]. Designing [Yu06]. Designs [BRS00, GP06, PGS07, ÅvR06, ABZ00, AGGZ01, AW09b, AH02, AG04, AZ04b, AYZ06, AFH07, AKJ09, AYH09, Air00, AEK09b, ADW03, Ant08, AT09, Atk08a, Atk08b, Atk09, Bag04, BCR08, BCS07, BJ06, BJ08, Bat04, BH08b, Ben01, BGZZ02, BDZ07, BDP07, BM06d, BH01a, BM01c, BH01b, BJJV08, BDH05, BS07b, BBP06, BC07, BW00b, BP02b, Bul07b, BR08, BR09, BL08b, BS05c, BS01, BS06, But01, But05, But06, But07, But08, But09, CK08, CH00, CM00a, CDD01, Cha01b, CHLY01, cIC01, CL02, CG03, CFQ05, CSL08, CKM06, CGS02, CCW05, CLLZ06, CW07, CL08, CM03, CG08a, CLS09, CKMS02, CJK04, DMS03, DDCC08, DL03, DNK02, DH08, DM02, DMP04, DP08a, Dey09, DDS00, DTM08, DF06, DFW08, DH09c, Duc00, DR01b, DDM09, DR02, ES00]. designs [FC03, FC04, FLM00, Fan00, FW03, FZ04, FWW06, FTY08, dSBFG07, FHKMS02, GACL08, GKMS02, GK05, GKM06, GHKTR07, Geo09, GF06, GT02, GD03, GDL07a, GTV01, GGHS04, GGS09b, GHSN00, GT06, GRP02, HC03, HS02, HOS06a, HS07a, Har08, HYP02, HZN04, HP07, HY08, HJ02, HK01, HKSTR05, HC07a, HS00b, IS01, IS02a, ITA^{+}09, IT10, IBL06, Iva06, Jac02, Jac03, Jac04, Jac06, Jac09, Jeo00, JM08, JR03b, JLNY07, JLN08, JLNY09, KM01, Kag08, KD08, Kle04b, KM00a, KTB01, KS06, KMS08, KMS09a, KS07e, Kus03, Kuw02, LT96, LT00, LK00, LC01b, LJ07a, LM08, LA08, LC09a, LTI00, Lia06c, LS02, LMS08, LS08, LR02, LT08b, LZ00, LRD07, LBS06, LT08c, LFTT08, LM00b, LNS07, MM02, MTC05a, MT06, ME04, MW08a, MS01, McL08, MP07c, Mel05a, MP07d, MJK01]. designs [Mis07, MRC02, MRC03, MPS05, MS02c, MMM06, MK08, MB00b, Mut05, NB08, NS00, OK00, Orr08, Oye03, OJKM02, PS04a, PM08b, Pav01, PPM09, PPX09, PR08a, PR08b, PM01, Pro06, PS02, Qin03, QL06, QW05, Qu07, QKO08, RX03, RRS06, Ree00, RS04b, RO05a, RGR09, Roy02, Ruk06, Ruk07, RB07, ST08, Sch07a, Sch00b, Sei01, SFT01, SYZ07, SSKS02, SK06, SR03, SR05, SBKF01, Sta06, SH05, TGV02, TD03, Tan07, TAGF07, TSR03, TWH^{+}05, Tom09, Top01, TMM09, TZ05, TGM07, Udd00, Udd05, Udd08, VG07, Wah09, WH05, WMSY06, WSY06, WJ09, Wie00, WX08a, WZ08, WX08b, Wu03, WW08, XL06a, Xu09, YM02, Yan05, YLLZ06, YP07, YB08, Yan08, YS08, YLS00, YZ07, Yin01, ZT04, ZP00, ZZL07, ZCS09, ZF01, Zho08, vBP03, vBHvdVW05, vT01, vdV08, vdVD09]. Detecting [DP08b, Dup09, WGCX09, AdHN09, BB03b, Ste07]. Detection [SC08, BM03, BD05, CHP08, Cho01, FWYY03, GP07a, GS09b, HPS07, HKPS08, MPU06, MP07e, PFF00, Sch04b, SY08b, Ste05]. determinant [Han08b, PR06]. determinantal [And07a]. Determination [WRV05, BD08c, CD02, De 04, DG02, DZ06, DHV01, Hös08, Jac02, KD08, NDF03, TSS06, TTC06]. determine [YA09]. determining [SW05]. Deterministic [GA05, CLE09, Hei00, KS09b]. Development [GC07, Sim08, Sen01]. developments [AG07, BB01b, GK07b]. deviates [CS00]. deviation [Che06a, SSE04, Tan01a, WZ09]. deviations [BDGV01, Hui09, Koh00, Lou00, MaL09, MP05]. Dey [Dea00]. DFR [Ryc01]. DFRA [Ryc01]. diagnostic [NdC08, PS03, ZC03]. diagnostics [WLCI09]. diagonal [Cua02, SS02e]. diagonalize [Sri07a]. diallel [GD03]. dice [TB07]. dichotomization [WG05]. dichotomized [GL09]. dichotomous [Cha01c, TTC06, VdCO00]. Dickey [Ste07]. diesel [LWL^{+}09]. Difference [Bie00, Kuh05, AM01, Cao01, DS02b, FHMS02, KS01, KT08a, LF07, LTW08, TVYS09, VV05, WW09, Wu07, ZTQ04, ZQ05, ZQ07]. differences [GMPZ06, KKM09]. different [ASC05, DH06, ES03, Kir07b]. differential [AA08, CIC08, HCW06, KV07, Rao00, WC00]. differentially [DR08b, GBC^{+}05]. Diffuse [KS01]. diffusion [DK03, DK04a, Fue00, RSM05, Shi09a, SGJM02, SH03]. diffusions [Hui09, Kal07, PvZ09]. digraph [Gil06]. dilemma [KSS01]. Dimension [NN09, Ben03, BGM03, CK01, DHT07, FHL03, FHW08, JMJ09, WC07, WC09b, YB06]. dimensional [AT09, BK05a, BWA06, CH05a, Cow03, Cox03, Csá02, DH03, Hai07, HO03, KTMS04, LLFF00, PG07a, Tha08, Wri08, YH09]. dimensionality [Har09, Pav03]. dimensions [DR03a, GPL09, KO03]. Diophantine [NSV02]. Direct [WPV04, BS07b, Lee01b, Pad06]. Directed [CS02c, CS03a, Fer02a, KKM09, SSP03, Yin00]. direction [SK08]. directional [LK03, LZ05, MFSW07, Mus02, SZ08a]. directions [CCL06, PD00]. Dirichlet [BH09b, Ber02a, Dor09, GMM03b, IOQ09, KvdV08, PY06, TG07a]. disaggregation [LOPV06]. discarding [Fou00]. discontinuous [AD01]. discordant [ZG05]. discoveries [And07a, KTMS04]. discovery [Bel07, GCR06, GR08b, KHC02, SZ08a, Tro00, Yek08, YSB09]. discrepancy [Bra06, HL06, ZN08]. Discrete [Pip09, SYZ07, TZ05, Ano05-40, ALH07, BD08a, BD09, Cai05, Cha05, CS08, CL09b, Con04, EKT06, GH07, Hai07, IK06, Kag01a, KP08a, Kem03, LB00, LBM06, LBR08, NK09, NSV02, Ong05, Pap05, Shi09a, SV00, Zho08, Zog08]. discrete-stochastic [SV00]. discrete-time [ALH07, Con04]. discretely [HKW00, Hyr07, Skö09]. discriminant [FHW08, Pav03]. Discriminating [KGM05, GS00, LFTT08, Ots08]. discrimination [Atk08a, BDP07, GDL07a, GDL07b, JLNY07, PSV06, RK05, Tom09, UB07, WWEL08, ZT04]. discrimination-another [JLNY07]. discrimination/estimation [WWEL08]. discussants [PW06a]. Discussion [Ano00l, Bor09, Bro06, Goo09, Jon09, Khu09, Luc09, O’H09, Par09b, Pie09, Rao02a, Sri02a, GR09b, Rob09c, SS02d, LHL09]. disease [BBSZ05, Mez03, RHL04]. diseases [BB01a, LQH03]. disjoint [BKLZ01, BM01c, DS02b, FHMS02]. disorder [LT09b, LT09c]. disparities [Gas07]. disparity [MS00]. Dispersion [Sch04b, BLL06, BCR00, BP02a, DLOR02, FCCN01, Lan01, LK00, Lia06c, Sah08, ZS08, vdV08]. Dispersion-effects [Sch04b]. Dispersive [KK02]. distance [AL08, AD01, BR03a, lCV06, DGGF04, Duc00, KL08, KW09a, KN04, KA09, LSW06, Lou00, MPGS09, Mei09, PdSPS05, Sun03, Tom08, VFVF02, Zha01b, Zog00]. distances [Bak09, SOUS03]. distinct [CL09a]. distinguishability [IL04]. distortions [KS09a]. distributed [BH05, HDM09]. Distribution [Aga02, And07a, Cha07, Dan05a, DGGJ06, Efr04, FL08, GG09b, HN09, SGK09, Sin06b, dE02, Ahs00, ASD09, AC07, AFW05, And05, Arc01, AG09, BCC02, BR03a, BW00a, BX07b, BX07a, BH08a, BTK08, BH09a, Ban03, BLZ08, BP01, BK09c, BL05b, CC07a, Che05, CH05b, CS00, Cua02, CCL06, DK04a, DB08, DvdW05, DGGJ07, DGR07, DdHL06, DL00b, EJ07, EJ09, Eri02, FZ06, Fie04, FG02, Fre08, FMN09, GB09, GT08b, GD02, GC07, GR01, GK07b, GH00, HH01, Hai07, HK04, HO03, HB04, HP02, IPdUÁ08, Ili03, IK06, IU01, IVM03, Jar07, Jon02, JB02, KK09a, Kag01a, Kem02c, Kem02b, KK09c, Ken00, KK06, KHP06, KM04, KSS08, KP08b, Kon07, KP05, KLX09, Kun07, KT08b, LLC02, LCM07, Lai09, Lee05, LY09, LB06]. distribution [LL07c, LH08, LWB09, LL06b, LGMPS03, Lon09, LB08b, LT09b, LT09c, MI08a, MP08, MN08a, MW07a, MG08, Mi06, Min07, MH00, MT02, MW05, MSW04, Oga09, OP00, PWW08, PF01, Pew05, PvdL04, PS09, Pro04, PC02, Raq02, RM05b, RMMA07, Ruk02, Ruk06, SD01, SA02, SGU05, Son01, Sør08, SLS02, SvTW00, Sul02, SZW08, SP05, Tan08b, TS08, TBM04, UG02, WZ04, WH02, WG06, Xue04, YO05a, YH09, ZBZ00, ZM06, ZW03b, dUÁGMCS00, vdWD01, vdW04]. Distribution-free [Dan05a, dE02, DvdW05, EJ07, EJ09, Fre08, IVM03, Oga09, vdW04]. Distributional [BE00, HKZ08, Rüs09]. Distributions [CR08, Hos07a, Ahm01, AF03, AM04b, AJMP09, AH08, Ano05-40, Ano09a, AVGQ05, ACS08, ARS01, BD08a, BD09, BD08b, BNA04, Bed06, BJS09, BH09b, BS09, BBIAV00, BJZ09, Cai05, CD06, Cha05, Cha02c, CIS00, CT04, Che04a, Che05, CT09, DS03, Dat07, DLS09, DLP03, DvdW05, DGLS05, DGJ06, DGL07, DGGJ09, DGGG08, EOGQ09, EHM09, FGT08, Fan04, Fan03, Fan05, FM05b, FCL08, Fuj02, GS07a, Gup03, GHL03, GK06b, GH07, Han08a, Has09, HLZ05, HLZ07, HNW08, Hod04, Hos06b, HS08b, HC07b, IJK08, JB08, JB09, JK03, Jon06, Kag06, KTK08, KM07a, KP08a, Kem02a, Kem03, Kem05, KS02b, Kle04a, KKB09b, KR07, Kot06, KM03, KT08a, KGM05, Kur08, LDC09, LX07a, Lee05, LT06, Let00, Lij03, LZ07, LB00, LBM06, LBR08]. distributions [MC08, MO07, Mar07, MP06, Mei07a, Mei07b, MW07a, MDM06, Min07, Moh05b, MZ07, NSV02, Obi06, Oga09, PNRdA03, Pap05, Pip09, RSD06, Rob09a, Roo03, Sar06, SZ07, Sen02, SJ08, SA07, TW05, Tom08, WWN05, XD05, Zac05b, ZH08, dCD09, vdW04]. disturbance [Krä05]. disturbances [VFVF02]. divergence [BDGV01, Eks08, GNP07, JW09, MLK09, PPZ02, PM08a, Zog08]. divergences [CP02b, Gil05, GP07c, PP05, PPdCP06, PP09]. diversity [SVO05]. divisibility [Das07a]. divisible [AM01, Bag04, Jac03, Kag08, MRC03, RSSR09, SK02, SK06]. DMRL [AM04b]. DNA [CIC08]. Does [LP09, BL09]. domain [BR02, BG07, Mee05, ST09, Vel03a]. Domains [SS06b, Kak02, NPA06]. dominance [AB09, Liu00a, Nam00]. dominant [GPC01]. dominating [EC08]. donors [SR02]. Doob [KM00a]. dose [BDZ07, DF06, GF06, GHSS02, IFC07, LLP02, SR05, TSS06, WL02]. dose-finding [DF06, GF06, IFC07]. dose-response [BDZ07, LLP02, SR05]. doses [BK07]. dosing [MCH^{+}05]. Double [HH09, YA09, YvdL06, DH00, Gov05, LZ09b, NvdL05, Oht00, Viš03]. double-blind [Gov05, LZ09b]. Double-smoothing [HH09]. doubleblind [Gov09]. Doubly [DH09c, AF03, GMPZ06, KHP06, SZW08, Sun03, Zhu01]. Douglas [Bor09, Jon09, Khu09, Luc09, Par09b]. down [BG05, DHV01, GF06, Gor07, Sch05a]. Downton [ASD09, Ili03]. Draper [But06]. driven [AKN06, FS07, HF00, Jup09, KLE^{+}04c]. drop [SCZ07]. drop-the-loser [SCZ07]. dropout [SJ09]. dropouts [KK09a]. drug [DFW08, Sen01]. DT [Atk08a]. DT-optimum [Atk08a]. duadic [Zha01a]. Dual [JG00, KRM^{+}05, CHLY01, KdR01]. Dunnett [Kwo01]. Dunstan [LGMPS03]. duration [Mi08b, Ver01]. Durbin [Sha08]. Durbin-Whittle [Sha08]. during [Aga02]. dyadic [Mas07]. Dyck [MSV02]. dynamic [AZ07, BB03b, GD08a, HT08, HW08, KS07c, LSL08, PGLS08, PP06b, Tar07, Tri08, XW07]. dynamical [DS07]. dynamics [BBSZ05, HC03].

E-optimal [PM08b]. Early [GPV09, BNP06, FS01a]. Easy [HLW07]. Easy-to-construct [HLW07]. econometric [RS08]. econometrics [Ans01, Lin01b]. economic [Jia06, RW06]. EDA [BPAK04]. EDC [DZ06]. edge [KS06]. edges [KDT06]. Edgeworth [AL08, Che03b, LF07]. Edition [McF01]. Editor [Kag01b, Sri04]. Editorial [Ano00m, Ano07p, Mel07a, Ano07o]. Editors [Ano03w, Ano03x, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-40, Ano03-41, Ano03-42, Ano03-43, Ano03-44, Ano03-45, Ano03-46, Ano03-47, Ano03-48, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano04-44, Ano04-45, Ano04-46, Ano04-47, Ano04-48, Ano04-49, Ano04-50, Ano04-51, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano05-38, Ano05-41, Ano05-42, Ano05-43, Ano05-44, Ano05-45, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x]. Editors [Ano06y, Ano06z, Ano06-27, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano06-37, Ano06-38, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano07-38, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano08-36, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36, Ano09-37]. Edward [Lu02]. Effect [Kir07b, BM02b, Che06c, CB06, KM08, NdC08, SST02, WS09, WH09]. effect-sizes [BM02b]. effective [BK07]. effectiveness [LP01]. Effects [GPC01, GDL07a, AA06, AH02, AYH09, Atk08b, BY09, BS07b, BL05d, CLLZ06, DS02a, EBZ08, dSBFG07, GIA04, Gan00, Gen01, GKK08, GGHS04, HB00, HK05, HP07, HY08, HT08, KG05, Kuw02, LRZ02, LT04, LQH03, Lia06c, LMH00, LK03, MCH^{+}05, NvdL07, Nie07, PW06c, RZ05, SZ08a, Sch04b, Sør08, SS08, SB04, TCD08, TGM07, Vid06, WD04, WYL09b, WB05, YLLZ06, YW09c, ZZL07, vdV08]. efficacy [BLU07, DF06, DFW08]. efficacy-toxicity [DF06, DFW08]. efficiencies [LLC09, Udd07]. Efficiency [BDH05, Hal07, Lin08, Liu00b, SBKF01, VdCO00, WH05, Aic01, ADW03, BR03b, BN03, BBP06, CIPS07, CL01, DNK02, DN03, IL04, IL06, Jan04, JL09a, JR03b, KSW05, KD07, LLC02, ST00, WA07, WY07b, YW09a, Zha04b]. Efficient [BM06d, Cow03, DS02a, DP08a, MEC04, Sán03, SZ08b, SV05, Sub00, TGV02, YP07, AEK09a, AEK09b, BRCK03, BW00a, BW01, BJJV08, But05, BL09, Cho07a, DLRP06, DMMS07, HZN04, JCS05, KR03, LRD07, Mal02, MP07d, PCK07, YS08, YW08]. effort [CM00b]. Efron [Che06d, TB07]. Eggenberger [SAC02]. eigenvalue [KT08b]. element [Wie02]. elements [CBD05]. Elfving [Stu05]. elicitation [CISW03]. elimination [LL07a]. ellipse [CF04]. ellipsoidal [Wil07]. elliptical [AVGRZ00, BBIAV00, DGJ06, DGGJ07, EOGQ09, Fan05, Hod04, MDM06, VIBAV06]. elliptically [BZ06, DGLS05, DGL07, Kon07, PC02]. Embeddings [TS01, BGG00]. emesis [VdBS01]. emissions [CGG09]. Empirical [BL02b, CG08b, DLM02, GS07b, Jia03a, Jia03b, KZ03a, Kul09, Lia06a, Lia09, MB00a, Mez03, SH06, SK07b, TP03, WLCI09, YXC09, YZ06, Bab04, Ban03, BKZ09, Bor00, Bra06, BL03, BL05d, Che08, DH03, DK04b, DdHL06, FK09, FWW09, GKK08, GL02b, GL05b, Hor00, HKZ08, HC06b, KZ03b, KLA08, Lai09, LGL05, LG06, Lia00b, Lia00c, Mas03, Mei05, Mei07a, Moj07, NLP06, OM08, Pen00, Pen03, Rüs09, Sen00, WD04]. employing [EOPR05]. empty [Udd05]. endogeneity [DD08]. endpoint [CS04b, LP09]. endpoints [EM07, Fre01, Fre02, HA08, MBA^{+}08]. engineering [Huz05]. enlarged [GKDN06]. entropic [DR03b]. entropies [DL09a]. entropy [AZ07, BAA08, Hos07a, MT02, NP06, Zog08]. Enumeration [JTS02, Fer02a, MP07c, Orr08]. environment [KKLD07]. environments [DNN07]. epidemic [Gua07, RS04a, Vex06]. epidemiologic [Gre08]. epsilon [MH00]. epsilon-skew-normal [MH00]. equal [Yin01]. equal-sized [Yin01]. equality [BNP06, BD08b, Che04d, CDH06, GD08b, KK01, Liu09, MTC05b, Páz02, SN03, Sch01, Sch07b, YLH04]. equation [BDHS06, Con08, SN03, SK08, WH08, WZ01a, WZ01b, ZjLS05]. equations [And07a, But08, CLHY01, DdRB00, FM05a, KV07, Nek00, NSV02, Rao00, War08]. equilibrium [MA05]. Equivalence [Har08, Ste04a, TSR03, Atk08b, Bet01, BD08c, CC09, FS01b, Gil06, KD08, Kwo01, Lia06b, MP08, Mun00, YLH04, Zha04a, dC02, vBLO04, vdV08, vdVD09]. equivalent [GS00]. equivariant [MS05, VOK^{+}03]. Ergodic [Kak06, ALRV01, DK03, DK04a, PvZ09]. Erratum [BD09, Bed00, DGL07, Fox09, Fre02, FM00b, IT10, LT09b, TLBJ11, Tan09a, Ten09]. Error [GDL07b, AAJ00, BGL01, BW05, BCGF09, BL02a, CP02a, CP05a, Che04b, Che05, CCRS07, DH03, Dup05, Efr07a, GP06, GRZ08, GKK04, GS07b, Gor07, HK05, JMJ09, JPS^{+}08, Kas00, KSS08, KMS09b, Lia00a, MSW04, Oye03, Pav03, TB09, WO00, WM02, Wil05, Xue04, YL07, YR00, ZK09b, ZY07]. error-correcting [BGL01]. Errors [GHJ^{+}07, APdR02, ARV03, BGS03, BLV00, CCL03, Don04, GP04, GL02a, GK07a, HY08, HH07, KK01, KK05, KZ03a, Koh06, KS00, KS08b, KS09b, KMV05, KSW05, KM00b, Laï05, LC01b, LJ09, LTI00, MY04, MJS00, NLP06, Par09a, RO05a, Sán03, SYZ07, TT08, TRJ07, TL07, Udd07, WW02a, WLCI09, YC07, ZY07]. errors-in-variables [GL02a, KMV05, WLCI09]. essay [BB01b]. Esseen [AL08, Ben03, BW08, LCM07]. establishing [Kwo01]. estimability [Qu06, TT09]. estimands [Rei05]. Estimate [Tar07, BM06d, Cha01c, CBG04, Efr07b, Hei00, LGMPS03, ZM06]. estimated [BK07, FBK03, HK04, Kul09, MT07b]. estimates [AMY02, AZ04a, AK03, AGK00, Ars04, BLV00, BW03, BL05d, DR03a, GHJ^{+}07, GB07, GJ03, GM05, HY03b, Ino03, JL04, Koh03, Koh08, Laï05, LLC09, LZ07, MP02, MY08, Sar06, SZ07, ST05, TN05, BRO^{+}09]. Estimating [ABS00, BLZ08, BLU07, Bos06, BJPZ08, CO08, EC03, GL09, GT08b, Hai07, KSS08, Kun00, KN03, MRLBBCPA00, MSW04, ODY08, SB05, WZ06, YW09b, YW04, ZjLS05, ATS02, BRCK03, BGV08, CYLC01, CBD05, CC07c, CLHY01, Che02, CC08, DMP04, DW08, EN02, FM05a, Fer04, GPC01, GT09, GHSN00, GS07d, HYP02, HS08b, Hua02, IU01, IP09, KRS07, LRZ02, Lia06c, NSV02, Oye03, TG07a, WT06, WH08, War08, Xu07]. Estimation [ASD09, BH07, Car02, Cha02c, DR05, Dav05, DS07, EKMS04, Fan04, FPR00b, FOS08, GG08, GGS02, HY03a, HO03, HT08, HKMS02, Ili03, KD01, KT07, Lar03, MP00, OS05b, PB08, PS04b, PK05, RMMA07, SD01, Sch00a, SJ08, SSE04, SM05b, SGU05, STW08, Sun06a, Tan01b, TS08, dSSdL08, Ver01, Wah09, WYL09a, WC01, WC08, WYL09b, XW07, XY06, AW09a, AB05, AJMP09, Ahm02, AS04, Alv05, AH01, And06, And07b, AFF09, ANR00, ALS07, ARV03, AT02, Atk08a, AWL04, BCC02, BW00a, BM02a, BM05a, BLL06, Bar08, Bat06, BBW06, BGS03, Ber09, Ber05, BP02a, BK09b, BDZ07, Bir09, Bla00, Bla06, BGMPG09, Bor04, BB06b, BH02, BQI06, BL05c, BL02b, CF04, Cai02, CG01, CFT07, cICY04, CYLC01, CS02a, CX07, CP07, Che06a, CISW03, CTL04, Che06c]. estimation [CZ09b, CCP09, Che04c, lCV06, Chi03, Cho07a, DNR09, DK03, DLM02, DS03, DLS09, DHP06, DS02a, DBDG09, DW06, DG09b, DW08, Efr04, Efr07a, Efr09, EN09, Eks08, EBZ08, El 04, EPdRV02, FZL00, Fer02b, FR07, Fot00, FMS04, FZ00, FARMA08, FH07, FE06, GSFBGM04, GSCG09, GRZ08, GJ01, Gho01, GD03, GHMR06, GS07b, GKK08, GB04, GRC09, GM04, GTV01, GD02, Gov05, Gov09, GG09a, HM01, HL06, HT09a, HB00, Hil09, HT09b, HB04, HYCY03, Hui09, Hyr07, IPdUÁ08, Jin07, KZ03b, KA06, KW09a, KM07b, KW09b, KKB09b, KMR01, Kow09, KV07, KvR08, Kuw02, LC01a, Lan01, LH00, LOPV06, LT02, LGL05, LA08, LT09a, LZ09a, Lia09, LM00a, Lin08, LMH00, Lou00, LCZ06, LB08b, LZ09b, LT09b, LT09c, MB00a, Mai01, MdUÁ04, Mas03, MS00, MP07b]. estimation [Mee05, Mei09, MSW09, Mez03, MS07b, Moj07, MPS09, MRCN09, MR00, MZ07, Nay00, NM06, OM08, OS06, OL03, Ozt05, PLK^{+}06, PK06, PF09, PQ09, Pen00, Phi02, PSJ07, PKM08, QJ00, QZF09, RJT05, Rao00, Rao02a, RZ05, RSD06, Rob07, RBV00, SL05, Sag05, Sch05b, Sen00, SM06, SH09, SCW09, sS06a, SZ08b, SY08a, ST04, SR02, Soe02, ST09, Sri02a, SGJM02, Sri02f, Sta09a, SdMSL04, SB04, Sub00, Sub01, SS02c, SS02d, SSvRM03, Sun03, TT08, TL09a, TVYS09, TG07b, Tom09, TSH06, TL09b, TK09, Vel03b, VS00, Vie09, VG07, WO00, WZ04, WCS01, WHS07, WBvdL07, Wan08, WWEL08, WD04, WR02, Wie00, Wil05, Wil07, WGCI09, Wu00, WB05, XW08, Yan00, YSH03, YWS09, YW08, YR03, YC07, YvdL06, YZZ09, YH09, ZT04, ZKW01, Zha04b, ZR05, Zha06, ZK09a]. estimation [Zha00, ZK09b, ZH08, ZS05, ZW03b, ZGH08, dUÁGMCS00, vdG02]. estimator [Ago09, AR05, And06, AS00, BM01a, BMW09, Che04a, Che04b, DdRB00, EC08, Fre07a, FY05, GM01b, GSC07, Haw01, Hlá03, HPV08, Hug08, JW09, KS03a, Kha00, KL05, KQ02, KMV05, KMS09b, LSW06, LP09, LP04, Liu09, MS05, MG08, MaL09, MP05, Nam00, NH09, NWS01, PWW08, PS09, PSW00, QW01, SG05, ST00, WDE05, XL08a, Yun02, ZRM06, ZL08, ZS08]. estimators [AM04a, AAJ00, ALS04, AH00, Arc01, AD01, BL06, BR03b, BM06c, BN03, BFGV08, BSP02, Ber07, Ber02c, BBdR00, BB04c, BK09c, BO04, BC03c, BL08b, BL03, BL09, Can06, Cao09, CC07b, CP02a, CP05a, Che01, Che05, CB06, CA00a, CR02, CA00b, CDH06, DCC01, DGGG08, EY08, FZ06, FM00a, Fer01, FZ00, Ger05, GKK04, GB00, GFM05, GRU01, GHH00, Gre08, HL04, cIWsL05, JL09a, KSL03, KZ09, KS00, KK09f, KSK00, KSW05, LLC02, Li03, LX07b, Li07, Liu00a, Mar09, MCDC08, MS06, MN03, Nay03, Nec06, NvdL05, Oht00, OS06, Pap01, PPZ02, PM08a, PCK07, Páz02, Pen04, Pfa00, Qi07, Rob09a, RMS02, SW07, Seg01, Seg05, ST06, SMS08, SL00, SV00, SGK09, Skö09, SV05, Sub04b, SC05b, Tom08, Tse04, TK06, TK09, Vas03, VV05, VFVF02]. estimators [WvEZ04, Wan05, WZ08, WN06, Xue04, Yun00, ZY01, Zie03, vdV08]. ethical [BB06a]. eukaryote [SYM05]. Euler [Kem02a, KT08b]. Europaischer [Sin06b]. EV [MaL09]. Evaluating [WC00, MMM06]. Evaluation [Atk09, CP05b, RGL07, ACBM09a, ACBM09b, Bor09, GC05, Goo09, HSS08, HDM09, Jon09, KD08, Khu09, Luc09, MBA^{+}08, Par09b, Pie09, Rob09c]. even [Ree00, SF02]. event [AP05, Ko00, Let08, ND05, ZZF09]. event-related [Let08]. events [LHOP01, PSG07]. evidence [De 04, FK09, KB08, KD00, MPS03, Tsa06]. EWOC [CLS09]. EX1 [Ano03-57, Ano03-56, Ano03-55, Ano04-61, Ano04-63, Ano04-62, Ano05-64]. EX1-EX2 [Ano03-57, Ano03-56, Ano03-55, Ano04-61, Ano04-63, Ano04-62, Ano05-64]. EX2 [Ano03-57, Ano03-56, Ano03-55, Ano04-61, Ano04-63, Ano04-62, Ano05-64]. Exact [BX07b, BX07a, BH08a, DZ00, DvdW05, DMMS07, ES04, JS08, LC09b, Rao02a, Raq05, Rob09b, Sri02a, SS02c, Zaf02a, vdW04, AC07, Ber04, GC07, GGS09b, MC08, VG07, WM02, Wit01, ZF01, vdWD01, SS02d]. examination [BD04]. Examples [Atk08b, CIPS07, Lia00b]. exceedance [BE00, Fer02b]. exceedances [BE06]. excess [BK02, DBDG09]. excesses [DGR07, NPA06]. Exchange [Roy02]. exchangeability [GACS07]. exchangeable [ED07, GK02, HNW08]. excluding [WS08, Wri08]. Executive [Sri04]. exhaust [LWL^{+}09]. Existence [BW03, CZ02, Ge01a, KM01, Kem02a, XP01, YZ07, Yin00, Zha01a, ZC07, Ben01, Ber01b, Cha01b, DI09, Ge01b, LK08, Mut05, Pel01, Sag05, WW09, Win00, Zhi03]. Existing [GK07b]. expanded [SS08]. expanders [DR03a]. expansion [AH01, Che03b, IS02b, KT02, LF07, TK09, WJ09]. expansions [Cua02, DH06, Fuj02, KK09b, Mar07, Oga09]. Expectation [Pap01, Sla06, And07b, BCP03, BJPZ08, Kör00, LT09a]. expectations [Bie08, CK00, Dan05a, GO03, GR08a, JS08, Koh00, PR04, Raq07]. expected [CL09b, Hos07a, LB00, LBM06, LBR08, MR01, YLH04]. experiment [DLRP06, DPR06, WRV05, Yu06]. experimental [AEK09a, Atk08b, BB00, BJJV08, DMS03, DP08a, DDS00, FTY08, FG06, Har08, HW08, MT07a, Ots08, Pro02, RVZRZ07, SBKF01, UB07]. Experiments [Ano09-47, BGL01, BG05, BS05c, BS01, BS06, CK08, CKM00, CM03, DR08b, Don04, ES03, GL05a, GBC^{+}05, HBK03, HKP08, Jeo00, JCS05, KKB09a, Kle04b, Maj07, MEC04, MCH^{+}05, NaX03, PS04a, PPM09, Pre01, RW06, Sch04b, SB04, Tat05, WZ08]. explanatory [DR03b, HLM08]. Explicit [Gor07, GM04, KTB01]. exploiting [TGM07]. exploration [SB06]. exploratory [BPAK04]. explore [Gen01]. Exponential [BLL06, DZ06, FWB08, Fer01, Ahs00, ASD09, AGHMPV07, BRCK03, BX07b, BX07a, BH08a, BH09a, BDP07, BK09c, CG06, CFS05, DZ00, DLRP06, Fer04, FG02, GR01, GHL03, GL05b, GK07b, GPR03b, HC03, HL09a, HT09b, Ili03, IJK08, KZ03a, KHP06, KP05, KGM05, Lan01, LGL05, Lia00c, Lia09, Lij03, MS00, MZ07, Raq02, RSD06, RM05b, SM06, SZW08, VM04, VdCO00, Vel03b, Viš03, WG03b, WW05, XY08, Xin09, ZB09a]. exponentiality [Ani10, BC09, MA08, WMK07]. exponentially [XC08]. exponentiated [SGU05]. exponentiated-Weibull [SGU05]. exposure [CBC^{+}07, KG05, LWL^{+}09, WMES07]. expressed [DR08b, GBC^{+}05]. expression [CIC08, PLLS08, ZMJW08]. expressions [FM00a, HCW06, Jon04]. extended [AN00, Pal09, Pip09, SS09, SH02, TT09, Wu00, YW09b, vdH05]. Extending [Bri08, Aug02]. extension [BT04, CH05a, FFFH02, NB05, Xin01]. Extensions [Das07a, KTA06, SF02, LFTT08, VV08]. external [LP09, TSR03]. extraction [ABQG09, DFG04]. Extrapolation [LuV09, BC07, Fan00, WX08a, WX08b, Xu09]. extrema [HPY02]. Extremal [MS09b, QRC01, Bur08, MF04, Par02, Rob09a]. extreme [BNA04, Che07, CWW03, DGGG08, DdHL06, Fie04, GM05, KS07d, LBR08, PQ09, RSSR09, Seg05, SGK09, Yun00, Yun02]. extremely [ZG05]. extremes [Car02, EZ09, Hai07, HL09b, SS09].

F [Sri02a]. Face [DG01]. Face-regular [DG01]. Factor [LC07, AEK09a, BH08b, BQI06, CA03, DS01, Don04, GTT^{+}07, GGHS07, HP07, MPS07, PP06b, Pre01, QW05, RO05a, TMP06, UB07, YB08]. factorial [AZ04b, AYZ06, AKJ09, AT09, BH08b, BGL01, BP02b, BR09, But05, But07, But08, CFQ05, CLLZ06, CGY07, CM03, Dea00, FPR00a, GACL08, HP07, Jac04, Jac06, Jac09, Jeo00, KM08, KD08, Kuw02, LJ07a, Lia06c, LS08, MEC04, McL08, MB00b, PS04a, PR08a, PR08b, PS02, QL06, QKO08, XL06a, YLLZ06, ZP00, vBHvdVW05, vdV08, vdVD09]. factorials [CKM06, SST02]. factorization [CIPS07]. factorizations [Chu01, Das07a, FR00, FR08]. factors [BCR08, BGM03, CG03, DL04, GGS09b, JR03b, KI00, Lia06c, LS02, SVZ07, TD03, VdBS01, Wie05]. fail [Pal08, ZB09a]. fail-safe [Pal08, ZB09a]. Failure [ABQG09, AFF09, BBKN07, BH08a, BLZ08, Bha07, BM06e, GKDN06, Leh09, Let08, LSL06, LuV09, NJP05, PP06a, ZjLS05]. failures [AL09, BNP06]. False [Yek08, GCR06, GR08b, KTMS04, KHC02, Sar04, SZ08a, Tro00, YSB09]. Families [Ma03, Mis07, Ahm01, AGHMPV07, Ber07, Bie08, Car02, DLM02, DS02b, FJT09, FHMS02, GO03, GPR03b, HL09a, Lia09, MO07, MN08b, MNOP05, Raq07, WW09]. family [BR03a, BP01, BH01a, CG06, CGS02, DGLS05, DGL07, Fan03, Gor07, GHL03, GL05b, KZ03a, Kem02a, LGL05, LS07, Lia00c, Lij03, MPGS09, Pip09, WG03b]. family-wise [Gor07]. FANOVA [AA06]. Fast [vdWD01, DG09b]. fatal [NHZ04]. Fatigue [Ano09-48, CLE09, dHD^{+}09]. fault [FWYY03]. favorable [Aug02]. FDR [FS07, LT04, Sar04, Sar08]. FDR- [FS07]. FDR-controlling [Sar04]. feature [CC07c, Pav03]. February [Ano00-36, Ano01r, Ano02-31, Ano02x, Ano04-59, Ano04-53, Ano05-63, Ano06-43, Ano07-44, Ano09-46]. Fechner [CHP08]. Fechner-asymmetry [CHP08]. feedback [NaX03]. felons [PW06a, PW06b, dL06a]. Ferrers [Cha02b]. few [HS02]. few-stage [HS02]. fewer [SK07b]. FGM [Pip09]. Fibonacci [AK05]. field [MNOP05, RM07, Ton07, WZ01a, WZ01b, WZ08]. fields [ALS04, CFT07, GK05, Ma09, RD08, Ste04a]. filter [Lin07]. filtering [KRV06, KS02b]. filters [Pol03, Pol09]. Filtration [ETS02, GKDN06]. find [AC04a]. Finding [CWW03, MM02, Nay03, DF06, GGS09a, GF06, IFC07, SR05, TSS06, WL02]. fine [CL09a]. Finite [KM07a, Krä05, NK09, AH03, ALH07, AGHMPV07, BCP03, BGI03, Cha01c, CK05, DdRB00, Dat07, Dub07, DMMS07, GS07b, Gol01, GM01b, HK04, Haw01, KX01, KZ03b, KM07c, Lee05, Lia06b, Lij03, LRZ08, LGMPS03, Mas03, Mei05, PK01, Pol09, SL05, SS08, Sud01, ySD03, yS05, WZ01a, WZ01b, ZCL09]. finite-sample [Lee05, Pol09]. First [ZBD07, BG05, DZ09, GC09, RBN09, WJ09]. First-order [ZBD07, RBN09, WJ09]. Fisher [BMP02, Das07b, DMMS07, GK06b, IS02a, KR03, KLR07, KL08, KI08, Lan00, LW07, Van05, ZG05, Zog00]. fit [AP05, Bak09, BM06d, Boc07, CV09, CC04, DR04, DN03, FG02, Fre09, Gil04, GH00, JÜ08, JS00, JPS02, KP08b, Mei07b, Mei09, MN08b, MT02, NMB02, PR06, PZ08, PS03, SN03, SVZ07, Sta09b, SNM08, UV07, WM08, Yin09, dUÁGMCS03]. Fitting [CC06, DG04, EN06, EN07, BJZ09, Cow03, Gua08]. five [KM01, Yin00, ZY02]. Fixed [LSW06, Sri01, Ano09a, DTW06, Gan00, GHSS02, HA08, LdC06, WK01, WB05]. fixed-sequence [HA08]. Fixed-width [LSW06]. Flexible [Arn09, BBP06, HA08, SH05, YO05b]. flexibly [Yao08]. Flocks [Tha01]. Florida [KHL07]. flow [VEM07]. flower [CL03]. Flowgraph [Huz05]. focus [Kir07b, RS07a]. fold [LR02, LT08b]. Folded [LMS08]. foldover [LJ07a]. follow [BL08c, KY07]. follow-up [BL08c, KY07]. following [CG00, Liu04, STW08]. food [GP04]. Forecasting [APR02, BCVV09, GD08a, KS09a, TB09]. foregone [Gol01, Sud01]. forest [BF07, FFP09]. Foreword [Sri00, Sri02c, Sri05a, Sri08]. Form [Vas03, DM08b, Gre08, MA05, TSH06, YW09a]. formed [ZMJW08]. forms [KL01, PC02, WC00]. formula [IS02b]. Formulas [Gil06, Gor07]. formulation [BB03a]. Forward [dCM09, SJ09]. foundational [Gol01, Sud01]. foundations [Cha02a, dC04]. four [DLM02, FKL^{+}02, GKMS02, GH01, LM08, LR02, ZC09]. four-person [DLM02]. four-weight [GH01]. Fourier [SFT01, ZT04]. Fractal [KO03]. Fraction [OGACR08]. Fractional [DHW00, YLLZ06, vBHvdVW05, ALS04, BGL01, BR09, But05, But07, CLLZ06, CM03, DS01, GA05, Jac04, Jac06, Jac09, Kuw02, LJ07a, LS08, McL08, PS04a, PR08a, PR08b, PS02, Tan08b, Vin02, XL06a, ZP00, Dea00]. fractionally [KV06]. fractions [FPR00a, vdVD09]. frailties [MW03]. Frailty [Han08a, GK06a, GG09c, MGG09, XL08b, ZGH08]. frame [BL05d, BL09, XP01]. frameproof [SvTW00]. Frames [FKL^{+}02, Ge01b]. framework [MBA^{+}08, NA09, SVO05]. Frankl [QRC01]. Free [SH02, Dan05a, DvdW05, EJ07, EJ09, FG02, Fre08, GSCG09, IVM03, LS02, Oga09, SL09, SH03, Sto05, WC09b, dE02, vdW04]. frequency [BG07, MCH^{+}05, NK06, Pol09, Ruk06, Vel03a]. frequency-selective [Pol09]. frequentist [EL04, FRF09, FR02, RSD06, Wal02]. friendly [GV02]. Frobenius [Pav01]. frontier [Pen04]. frontiers [DFS08]. full [AT09, GHKTR07, Hur05, ZGH08]. full-dimensional [AT09]. Fuller [Ste07]. fully [CG08b]. function [AKI08, And05, And07b, ATS02, AD01, BCC02, BK09b, Bon08, BMW09, Cao09, Che05, DK04a, DM08b, DBDG09, DdHL06, DW08, EPS08, FWB08, Fre08, GB09, Gen01, GKK08, GT08b, GTV01, GK08, GG09b, HK04, HO03, IPdUÁ08, IP09, KI09, KMB08, KN03, LC01a, Lai09, LT04, LGMPS03, Mei05, Mei07a, NBH05, PK06, PR08b, RM05b, RMMA07, SL05, sS06a, SY08a, Shi00, Soe02, TL09a, WH08, Xu07, Xue04, YO09, ZRM06, ZS05, dUÁGMCS00]. Functional [Che06a, DFS08, KK09b, Yan00, Yao08, BH07, CZ09b, DMP04, FZL00, GS07b, HMW04, Hil09, KS09a, KK03, Lia03, Mel05a, Nec06, Pre07, RV07, SSO07, Shi09b, Zha09a]. functionals [AC03, Csá02, GP06, Jan01, KRS07, LX07a, LT06, LB08b, MG08, MSW04, Tan08b, ZK09b]. functions [AJMP09, AFW05, ADW03, AGH07, BLBR00, BBSZ05, BGV08, BC03c, CYLC01, CS02a, CX07, Che06a, CC08, CK00, Cua02, Cua09, Eks06, EKMS04, EKT06, EPS08, Fue06, FHM00, FY05, FH05, GN03, GT09, GRV08, HKMS02, Ili03, JB02, Ken00, KM07c, Laï05, PB08, Pel01, Pen04, PQ07, Pou01, SSW09, SL09, SG05, Ste09, TS08, Ton07, WT06, WW02b, YMJ08]. fundamental [NK06, Pew05]. Further [AM04b, CMyS02, EKL07, Sri07a, ST00]. future [AdL05, KHP06]. futures [LB06]. fuzzy [Kör00, Vie09]. FWE [FS07]. FWE-controlling [FS07].

G [AWL04, Con04, Sen02]. gains [CL01]. Galton [Moh05b]. Gambler [SF02]. games [Hai03]. Gamma [Ruk06, BW00a, Bon09, Che04c, CH05b, HC07b, Kag06, MP06, MvdMV06, WM08]. Gamma-distribution [Ruk06]. GARCH [HKZ08, MY08, Sta09b]. gatekeeper [WK01]. Gauss [IP09]. Gaussian [CGG09, Cha02c, Cho07b, DKSZ07, DP08c, EL04, FJT09, Gar01, HLZ07, HT06, KK09b, Ker04, KM04, KT08a, LCM07, Ma09, Maj07, MP02, Min07, MT02, MT04, MW07b, NMM05, OP00, Pro04, RM07, RD08, SS09, Ste04a, TT07, Tan01b, TL09a, TW05, TG07b, Zez09]. GCV [Fu05]. GDDs [Yin00, Zhu01]. GDPBIBD [HM02b]. GE [GK06b]. Gene [SSO07, CIC08, HCW06, PLLS08, RHL04, ZMJW08]. General [Bos07, GG09c, MN08b, AJMP09, ARV03, AVGQ05, Aug02, BM02a, BM05a, BLV00, Bha07, Cao09, Car02, CG06, CCL06, DS02a, DGGJ07, Fer04, GS07a, GJ01, cIWsL05, IP09, LS07, LL08, Liu09, Mar07, Moj07, PP06a, PSG07, RH08, SVO05, SW09, VM04, WL01, YWS09, ZGH08, ZS00]. Generalised [MMP00, DGGJ06, NDF03, SW01]. generalization [AG09, CG06, FY05, Par09a, XP02]. Generalizations [GM01b, KTK08, dCPC05]. Generalized [AM04a, Ahs00, AN00, BC03c, Bur07, CH01, CT09, FM05a, GC09, GK07b, HBK03, HS08a, ITA^{+}09, IT10, LL04, LL05b, LLW07, Lu08, ME07, Seg05, SAC02, TC08, ZY01, ZWHF07, ACW02, AF03, ARS01, BM06b, Bar07, BB09, BW09, Ber02c, Bie08, Bon09, CZ09a, CC07a, cIC01, CIS00, CISW03, CdA09, CK00, CM00b, DKSZ07, DGR07, DHW00, FS02b, Geo09, GB04, Gor07, GGHS07, HT09b, HL01, JB09, Jia03a, Kem02a, KI00, KM03, KLX09, KGM05, KS07e, Mai01, MS06, Mis07, MS07a, MK08, MN03, NP06, NK06, Nie07, NWS01, OKD06, OGACR08, PP09, Qin03, Raq02, RM05b, Sha08, Shi05, SCZ07, TD03, TP03, WH08, War08, Wie00, WX08b, Win00, XL08a, Xu07, YSH03, Yan08, YHA09b, Yao08, YC07, Yun02, ZDLZ02, ZRM06, ZB09b, ZZF08]. Generalizing [Har02]. generated [HM02b]. generating [Mei07a, WW02b, YCB06]. generations [Kas00]. generators [CWW03]. genes [DR08b, GBC^{+}05]. genetic [BB03b, SOUS03, ZCL09, ZG05]. genome [AKP^{+}07]. genomic [KTMS04, PdSPS05]. Geo [Con04]. Geo/G/1 [Con04]. geometric [CD07a, Cua09, FC03, Gol01, KP05, RBN09, SV01b, Sud01, Tha08]. Geometrical [Cha02a]. geometrically [Cha07]. geometries [ySD03, yS05]. geometry [CMS02]. Geostatistical [BGP00, Cow03]. geostatistics [PD00]. Germain [Nie02a]. GI [Aga02, ALW07]. GI/M/1 [Aga02]. GI/M/c [ALW07]. GIC [SR00]. given [EPS08, KHP06, Ma09, Tam02]. GLM [KA09, LCZ06]. Global [SYM05, Ber02c, HS09, KKW09, Ral02]. Globally [ASBZ04, Aug02, HY03b]. GM [BBdR00]. GM-based [BBdR00]. GMANOVA [TK06]. GMM [Fan03]. gold [CN05]. Goldstein [LHL09, Cha09, KSS01, MH09, O’H09, Smi09]. good [BT04, BR08, FGS06, GS00, GL05b, RB07]. Goodness [Bak09, CV09, Gil04, JS00, KP08b, Mei09, dUÁGMCS03, AP05, Boc07, CC04, DR04, DN03, FG02, Fre09, GH00, JÜ08, JPS02, MT02, PR06, PZ08, PS03, SN03, SVZ07, SNM08, UV07, WM08, Yin09]. Goodness-of-fit [CV09, Gil04, JS00, KP08b, Mei09, dUÁGMCS03, AP05, Boc07, CC04, DR04, DN03, FG02, Fre09, GH00, JÜ08, JPS02, MT02, PZ08, PS03, SN03, SNM08, UV07, WM08, Yin09]. Govindarajulu [LZ09b]. Gower [BCGF09]. GPU [Yua08]. gradient [LZ07]. gradual [HS00c]. Graph [CN02b, CS03a, FCS05, KDT06]. Graph-theoretic [CN02b]. graphical [CL09b, DP08c, Gas07, HT09a, KM08, OK00, PS03]. graphics [Lee01a]. graphs [FR00, FR08, FHJ^{+}04, IS02a, JTS02, KKM09, KM00a, LK00, LS02, yS05, ZY02]. Graubard [Lu02]. Green [BV01, BV03, GB09, Zha04b, ZRM06, dUÁGMCS00]. Grid [YB08, MM06, SV00]. Groß [Sin06b, KK01, KK05]. Group [ES02, GF06, Bag04, BLSvdDS05, HK05, Jac03, JT01, JG00, Kag08, LT08c, MRC03, RRS06, Sch07a, SK02, SK06, SvTW00, Tat05, WLC08, Zha01a, Zhi03]. group-censoring [WLC08]. group-sequential [LT08c]. group-wise [Sch07a]. grouped [ARV03, KP09a, Mi06, ZKW01]. groupings [MN08b]. groups [GF00, GGS09b, JT01, MSW09, Pav01, SGJM02, TKP06, YHZ05]. grows [BGM03]. growth [ÅvR06, Cha03a, CL02, CLE09, Dau04, DP08a, KZ09, RV03, RWF04, TT09, WC01, Wu00, YW09b, ZWHF07]. Guaranteed [Vex06, KV07]. guarded [KD00]. guide [Lah00]. Gupta [Ano01l, Dee06, MR06, Pan06]. Gy [PSW00].

H [Con02a]. Haar [ES02]. Hadamard [Bed00, Bed98, BK09d, DR08b, KKK03, LA08, MMP00, Win00, ZDLZ02]. Haenszel [FY05]. Hájek [Sen00]. Hall [LW07]. halvings [EKTR00]. Hammersley [Kha00]. Hamming [FS01b, PdSPS05, Zha01b]. Handling [PW09]. hardback [Lin01b]. Hardy [CQ08]. Hardy-Weinberg [CQ08]. harmonic [Gil05]. harmonizable [Kak02, Soe02]. Harris [Hyr07]. Hastings [CT02]. hazard [CMN06, DG09b, Dup09, EPdRV02, GG07, KNH09, MdUÁ04, Pal08, Sub00, TKP06, ZY01]. hazards [BB06b, Dup05, FPvdVvH06, JL09a, Sub00, TR05, WH09, Xu07, YL07]. head [Len02]. head-to-tail [Len02]. Health [Lu02, Con08, Gas07]. heat [BDHS06]. heavy [AM04a, AdHN09, BJS09, Fan04, GN03, KP08b]. heavy-tailed [BJS09, GN03]. hedging [LB06]. height [KP02]. Hellinger [lCV06, KW09a, LSW06, Tom08]. HELP [DH03]. heritability [GD03]. Herman [Zac05a, Ano05-65]. Hermitian [PC02, Sri05b, Sri07a]. Heterogeneity [WH09, HM01, Sch05a]. heterogeneous [BB00, DH09c, Hil09, LQH03, Tar07, Wu07, ZB09a]. heteroscedastic [BBdR00, BW00b, CCP09, FWW06, GK07a, Ino03, KSS08, LC01b, MSW09, MTC05b, NW05, Oye03, RO05a, WZ06, YW04, ZY07]. heteroscedasticity [CH01, Det02, TGV02, WLCI09, WW05]. heteroskedasticity [BSP02, FCCN04, FCCN05]. heteroskedasticity-robust [BSP02]. heuristic [KT08b]. hidden [AT02, Arn09, CGG09, Spe09, Whi05].

Hierarchical [Zas06, AGP09, AC06, CBC^{+}07, DJ06, JW09, LL07c, MS03, RM07, SS03, TP03, XC06, YR03, dC02, vdLP03]. hierarchically [HA08]. hierarchies [CK01]. high [AC04a, BWA06, BS01, CSL08, Cow03, DH03, GPL09, JÜ08, JMJ09, KTMS04, PG07a, PF01, SSO07, VBF03, YH09]. high-bias [BS01]. high-dimensional [Cow03, DH03, KTMS04]. high-throughput [SSO07]. Higher [FCCN01, FM00b, FM00c, IVM03, Mae05, Pal03, Sta09a, TT07, FL03, GM01a, IS02b, SC09, WC00]. Higher-order [FCCN01, Mae05, Sta09a, FL03, GM01a, WC00]. highest [BM06d]. highly [CBG04]. Hilbert [Bos07, Pre07]. Hilferty [MW05]. Hill [GM01b, LP04]. him [Sat07]. historical [BB01b, Bha09]. HIV [AKP^{+}07, HW08, OCR^{+}05]. HNBUE [BC09]. Hochberg [KHC02]. Hölder [RS04a]. holdout [HH01]. holes [Yin01]. Holey [BGZZ02, Ge01a]. Holt [BCVV09]. homogeneity [Bak09, BT06, CS08, CQ08, FCL08, Hay01, Kra09, MT04, Par00a, QSL09, SVO05]. honest [CT00]. honor [Ano07-45]. Honour [HKP08]. horizontal [RL09]. Hosmer [Ko00]. hospital [Aus05]. hot [SR02]. hot-deck [SR02]. Hotelling [KI08, SM01]. Hotspot [MP07e]. Huber [Aug02, KM07b, NH09]. hull [Mar09]. human [Ban03, CBC^{+}07]. Huskova [Ano00n]. hybrid [BX07b, BX07a, KMS09a, Kun07, vdLP03]. hybrids [Hor00]. hyper [Kem02c]. hyper-Poisson [Kem02c]. hypercube [Geo09, YLS00]. hypercubes [LS06]. hypergeometric [Kem05, LCM07, LBM00, Obi06, SJ08]. hyperplanes [OP05]. Hypotheses [DK09, Ahm01, BD05, CKS07b, DF07, DLOR02, EMMM00, KKM01, KLE^{+}04c, Li08, Llo05, MPS03, MW02, MM07, Sen07b, dBJM09]. Hypothesis [CO02, AM04b, Con01, De 04, GH03, Kee05, KS09b, SR00, Sta09b, WM02, WH05, WGCI09].

i.i.d [ALRV01, HL01, MW00b]. IBC [Ano02r, Ano02s, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03y, Ano03z, Ano03-27, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano05-38, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u]. IBC [Ano09v, Ano09w, Ano09x, Ano09y]. ideas [GH07]. identical [CB06, Sch07a]. identifiability [XC06]. identifiable [DSG07, RS08]. Identification [CGPP06, RH08, BLSvdDS05, CGG09, DD08, Gar05, GDL07a, GDL07b, MS03]. identification-robust [DD08]. identified [RS08, Zha09d]. identify [Hay07, PS04a]. Identifying [XD05, DR08b, GBC^{+}05]. identity [BDHS06, CSX08, Das07b, Han08b, KL08, WZ01a]. if [HK04]. IFR [Ani10, MA08, Ryc01]. IFRA [AM04b, Ryc01]. IG [MW07b]. IG-symmetry [MW07b]. IGC [ZWHF07]. ignorable [BYW07]. II [Ano02-32, BD09, BE06, BX07b, BS08, BH08a, BD08a, BGI03, Bur08, Csá02, ES03, Fer04, GS09b, HKPS08, Raq05, SGU05, SZW08, WJ08]. iid [PX09]. III [Ano03-59]. IISA [Ano05-39]. illness [BD04, DB01]. illness-death [BD04, DB01]. imbalances [BCR08]. immigration [Iva06, Qi07]. Impact [LN03]. imperfect [BL09, Fre07b]. imperfections [BL05d]. implementations [Gra09]. Implementing [ZH09]. implication [GL02a]. Implicative [Ber02a]. implied [Wal04]. importance [DNN07, FFH00, MMM06]. Imprecise [CvdL01, UG02, Ber02a, Har02, Nau02, Zad02, dC02]. imprecision [dC02]. Improved [BFGV08, FLCN05, GP07b, GH07, IU01, KL06, KD07, RRS06, Sch05a, Vid08, Wan05, FCCN04, FCCN05, OS06, Sri02f, TSH06, WZ04]. improvement [EY08]. Improvements [Roo03]. Improving [KT02, Kon07, Shi00, Skö09, dHD^{+}09, CC02, GO01, MS05]. imputation [FMT05, Nie01, PMS^{+}01, QRR08, SR02]. imputed [Rei05]. inadmissibility [MvdMV06, NWS01]. INARMA [Wei08]. incidence [EKMS04, EKT06]. incidental [NHZ04]. include [KS08a]. including [Lah00, Lee01a]. inclusion [AR05, Bar03, ND06, Ros00]. income [DLM02]. incomplete [AGGZ01, Bag04, BLSvdDS05, CCW05, DS06, DS08, DS07, DDM09, GK01, HM02b, HJ02, HC06a, KB00, MMM06, Mut05, NB08, OJKM02, PM01, SS01, ZCS09]. incorporate [HS07a]. incorporating [MW03, ODY08]. increase [WY07b]. increasing [AAEK06, Cha07, Eks06, EN02, Mar05]. increments [Dan05a, Hai00, Raq04, Raq07, WC08]. indecomposable [Chu01]. Independence [AG07, MML02, BdM09, GRR07, HL09b, Kag06, MW05, PM08a, RA06, WM08]. independent [CT02, GC07, KL01, KX07, KN03, LDC09, MS07b, RTG07, SV00, TW05, WWN05, ZTQ04]. Index [Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02h, Ano02i, Ahm05b, Ano02a, Ano02b, Ano02c, Ano02d, Ano02f, Ano02g, Ano02e, Ano03c, Ano03d, Ano03j, Ano03e, Ano03f, Ano03g, Ano03h, Ano03b, Ano03i, Ano04a, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04b, Ano04h, Ano04i, Ano05b, Ano05c, Ano05d, Ano05e, Ano05j, Ano05f, Ano05g, Ano05h, Ano05i, Ano06a, BBW06, BFGV08, CCW05, DHP06, FKL^{+}02, GM04, Hil09, KT08c, LP04, Lop09, MF04, MP07b, MRC03, Nec06, PQ09, Rob09a, Seg05, WCS01, Yin00, Yun00, Yun02, ZCS09]. indexed [Che06a]. Indicator [PR08b]. indicators [PY06]. indirect [KRS07]. indirectly [CT00]. Individual [GIA04, DMP04, HT08, RDS08, WH01]. individuality [Rao02b]. induce [GHJ^{+}07]. induced [Man02, MdCC05, NSV02, Tse04]. Inequalities [Koh00, Ahm01, AM04b, FR01, GN09, LLP09, MW02, MA05, Sla06, XY08, Xin09]. inequality [AC02, BSW08, Efr07b, ENAZ09, HL01, IS02a, Kag01a, OS05a, SS02b, SSB02, WO00, XW08, QRC01]. inequality-constrained [XW08]. inevitability [FMS04]. infection [LQH03, MJD09]. infectious [BB01a]. Inference [Ani10, BD09, Bed00, BGI03, CS02d, DGL07, EN07, EJ09, Fox09, FR08, Fre02, FM00b, GS07a, GF00, Gov09, IT10, Jan01, LT00, LY09, Lon09, LT09b, Rao02a, RS08, Sri02a, TLBJ11, Tan09a, Ten09, Wu06, YR00, ASBZ04, AVGQ05, AC04b, BCN04, BX07b, BX07a, BH08a, BL08a, BC02, BZ06, BK09a, BS09, CCL03, CS02c, CY04b, CDB^{+}09, Dau04, DLOR02, ES02, FCCN04, FCCN05, FLCN05, FSS07, FR02, FWW09, GR09a, GRV08, GPV09, HKW00, Hir02, Jia03a, Kal07, KX01, KHP06, KS07c, LSL08, LHL09, LP06, MS07a, NMM05, NvdL05, Ong05, Özt02, PT08, PvZ09, PSG07, Rob09b, RM07, Ruk07, Sch00a, SS02b, Sen07a, Shi00, Sun06b, Sun09, VM04, WW02c, YGJ05, ZY07, Cha02a]. Inferences [DJ06, KM03, KX07, KT08a, Raq02, WS09, YW09c, Che03b, HKP08, LL05b, LLW07, WPV04]. inferential [GP03, Man02]. inferior [Hay07, LL07a]. inferiority [LTW08]. Inferred [HM02a]. infinite [BH01a, CCL03, Das07a, GB07]. infinitely [RSSR09]. inflated [GML06, YHA09c]. Influence [NBH05, RS07a, Loy01]. influential [BB03b]. Information [BB03b, Che04d, DS06, DS08, Pro02, VS00, AFW05, ALS07, BGL01, Bha09, BG09, BG07, CY04a, CL01, CDH06, Das07b, DJ06, EMMM00, Fab09, GPV09, GK06b, Har09, HK03, KG09, KLR07, KL08, Kee05, KC05, KK03, Lan00, MS02b, QW01, RM05c, Sin03, Son01, Sub03, Sub04a, TSH06, TGM07, WY07b, ZG05, Zog00]. informative [EN06, EN07, EN09, FM05b, SJ09]. informatively [TR05]. inhomogeneous [Hui09]. initial [Nie02b]. innovation [MT07a]. innovations [TL08]. inputs [Mal02, MMM06]. inspected [CX05, SS00, TLBJ11]. inspection [Lia06b]. Instrument [DD08, CN05]. integer [BM08, GC09, Kha00, RBN09, ZBD07]. integer-valued [GC09, RBN09, ZBD07]. integers [OT02]. Integral [TSN07, BH02]. integrals [Fue00]. integrated [KV06, Nag09]. integration [BM05b, GA05, Hay06, MRaPGVVV02, Tan08a, Vin02]. intensities [KS07d]. intensity [AD01, FH05, Sch05b, dBJM09]. Interaction [Wyn08, BGP00, Gan00, HP07, VS00, YW09c]. Interaction-contrast [Wyn08]. interactions [GF00, GC05, KI09, YB08]. intercalation [Bie00]. intercept [DH08, DPR06, OTB06]. interest [Kuw02]. interesting [Kuw02]. interference [ME04]. intergranular [RWF04]. interim [BBP06, Gov05, Gov09, LZ09b]. interlaboratory [RS07b]. Intermediate [IL06, Fre01, Fre02, IL04, PY09]. internal [TWH^{+}05]. interpolating [SSC07]. interpolation [BCGF09, BC07, CD03b, Skö09, Cha01a]. interpretations [CY04a]. Interrelations [MW07b]. intersection [CL03, Sen07a, Sen07b]. intersection-union [Sen07b]. Intersections [She01]. Interval [HSV09, OM08, Aug02, AC04b, BQI06, CG01, Che06b, Cho08, Dub07, EBZ08, GT08a, LSW06, LWB09, LMH00, SSE04, Sta09a, TVYS09, TSH06, WWN05, Yan00, ZCL09, ZjLS05, ZQ05, ZQ07]. interval-censored [CG01]. interval-censoring [ZjLS05]. intervals [AP07, AKK09, BLBR00, BDZ07, BL05b, Bur07, Cai05, CL06, CG00, CW05, DK04a, DBG06, EC03, FZ03, FED09, Fox09, Fre01, Fre02, GTT^{+}07, HK05, HCW06, KG09, KP08a, KM03, LF07, LL04, LL08, Liu04, LM03, Mae05, MR01, MS02d, OD06, PQ07, QRR08, Rob09a, Sin06a, Ste04b, SZW08, TR08, TW05, Wan06, XM09, ZTQ04, ZEB07]. interview [GGS02]. intraclass [GH03, GTT^{+}07, HB00, YW09c]. Intrinsic [DL05, Das01, Son01]. Invariance [BW09, Fot00, HS00a]. invariant [BK05a, BK09b, Das01, ES02, KT08b, Ten07, Ten09]. inventory [Jia06]. Inverse [Cav06, Cai02, Cha02c, DGGJ06, HLZ07, Haw01, KT08a, LC00, Lue09, Min07, MT02, MT04, MW07b, NMM05, OP00, SC05a, sS06a, SY09, TW05]. inverse-probability-weighted [sS06a]. inversion [Böh02, IS02b]. inversions [BK05b]. investigation [BL05d]. involving [CG03, KKM09, Kem03, MS03, RW06, SR03, ST05]. irregular [MR07, PQ09, RM05a]. irrelevance [dCM09]. ISBN [Cha01a, Cha02a, Con02a, Con02b, Dea00, Ko00, Lah00, Lee01a, Lin01b, Lu02, Mas01, McF01, Roy07, Sin06b, Wes00]. isometric [CGPP06]. isomorphic [FZ04]. isomorphism [LS08]. isotones [WMK07]. isotonic [Che09, Hir02, HH08, KSH02, LLP02, SSW09, SGK09, ZKW01]. isotropic [GSFBGM04]. Issue [Ano07-45, Ano09-47, Ano09-48, HKP08]. Issues [GR09b, BB01a, MS03, Yua08]. István [Csá05]. Item [CC07c, GGCA00, QRR08]. iterated [Nec06, XL08a]. Iterative [Ino03, Ars04, MP02]. IV [But05, But07].

J [BD09, Bed00, DGL07, Fox09, FR08, Fre02, FM00b, Gov09, IT10, Jan01, LT09b, Rao02a, Sri02a, TLBJ11, Tan09a, Ten09, Wes00, BCS07]. J. [Smi09]. Jackknife [MG08, PY09, SR02]. Jackknifing [SW04]. James [Con02a, GKK04]. January [Ano00-29, Ano01p, Ano02-27, Ano03-50, Ano04-61, Ano04-55, Ano05-58, Ano05-57, Ano06-39, Ano07-39]. Jeffreys [DMN01]. Johnson [Con02a, Tom08]. join [YB08]. joinpoint [GRC09]. Joint [Leh09, MS02a, BBKN07, BTK08, Let00, LW09, Moh05b, Ozt05, Raw02]. Jonathan [LHL09]. Journal [Ani10, EN07, EJ09, FCCN05, LT00]. journey [WR02]. Jr [Ko00]. Jr. [LOV08]. JSPI [Sri02b]. Jubilee [Sri02b]. judicial [KMPP02]. July [Ano00x, Ano01u, Ano02-30, Ano03-56, Ano04-58, Ano05-55, Ano07-42, Ano08-38, Ano08-39, Ano09-40]. Jump [SGJM02, Spe09]. Jump-diffusion [SGJM02]. jumps [Shi09a]. Jun [Mas01]. June [Ano00-35, Ano02-29, Ano04-54, Ano07-41, Ano09-39]. Junjiro [HKP08]. Jürgen [Sin06b]. justified [CM00b].

Kalman [Lin07]. Kaplan [Sub04b]. Kass [Cha02a]. Katz [Fan03]. Keeping [CLT07]. Kendall [Che07, FN07, GN09, LL07b]. Kernel [CFT07, DNR09, KL05, Laï05, dUÁGMCS00, Bir09, BH07, BL05c, Che06a, Che04b, Det02, DT08, EPdRV02, Fot00, GSFBGM04, GHH00, HLH09, JPS^{+}08, KA06, Kas00, Lou00, LSD01, MCDC08, MP05, Nec06, Pre07, Ste05, Sub04b, ZK09a, Zie03]. kernel-type [GHH00, Nec06]. kernels [PSJ07]. key [SvTW00]. Khattree [McF01]. Kiefer [MW05, ZS08]. kin [GPC01]. kin-cohort [GPC01]. kinetic [BW05]. Kirkman [BZ00, CL03, She01, TS01]. KMT [Mas07]. knockout [Gli08]. knot [SH02, SH03, Sto05]. Knowledge [Sin03]. known [BB04b, CSW02, KvR08, SM05b]. knuckle [dHD^{+}09]. Kolmogorov [Mei07a]. Korn [Lu02]. Koziol [BV01, BV03, GB09, Zha04b, ZRM06, dUÁGMCS00]. Kreweras [Nie02a]. kriging [Cha01a]. Kronecker [Kle04b, LLZ07]. Kuks [AS00]. Kullback [EY08, Eks08, HL06, KC05, KSK00, KT07, VS00]. kurtosis [FM00b, FM00c, MML02].

L [Ano07-45, Cha01a, Lu02]. label [Spe09]. labeled [CS03a]. laboratory [HSS08]. lack [BM06d, MN08b, NMB02, PF08]. LAD [FZ06, MR00]. lag [PT08]. Lagrange [Böh02, Zha09c]. Laguerre [KV02]. LAM [MP02]. Lambda [AC07]. LAN [FGT08, HT06, MP02]. Lang [Sin06b]. Laplace [IK06, KP04, LLL09]. Large [BDGV01, BKLZ01, Lou00, ZCL09, BM05b, BL05a, BH08b, BLU07, BM01c, Cha01b, DM08a, FFP09, FHW08, Mar05, PLK^{+}06, Pew05, Roy02, Tan01a, TR08, ZN08]. large-sample [BM05b, Pew05, TR08]. large-scale [FFP09]. Larger [CJRV08, Bos06]. largest [KT08b, STW08]. LASSO [PS09]. last [SJ09]. late [FS01a]. latency [MCH^{+}05]. latent [MS03]. later [AG07]. latin [CD01, DHW00, Sri07a, BW01, CL09a, Geo09, LS06, Sri05b, YLS00]. Lattice [Ano05-40, GV02, KdMSY09, Sen02, SA02, Tam02, Aga02, Csá05, FP05, KR07, Lai09, RSS02, Saj00, SM05a, SAC02, Sul05]. Lattices [FP05]. law [Dik00, Mar05, MS09b, Nec06]. Laws [XL08a, DM08a, Deh00b, DGJ06, GLS09, RSSR09]. layer [CS03b]. layout [Dob02, Sah08, Sch00b]. layouts [GA08, Wyn08]. leader [Xin01]. leads [Xin01]. Learning [XC08, KS03a, XC06]. Least [Cíz06, PX09, Wie00, AM04a, AW09a, Aug02, BSP02, CN02a, Cha03a, EC08, FM05b, GB07, Ino03, KMV05, LLC02, Liu09, Páz02, YC07]. least-informative [FM05b]. Least-square [PX09]. least-squares [AM04a, BSP02, EC08, GB07, Ino03]. leaves [PP02a]. left [HY03a, sS06a, SY08a, ZZF09, dCD09]. left-right [dCD09]. left-truncated [SY08a, ZZF09]. left-truncation [sS06a]. legacy [MHM07]. Leibler [EY08, Eks08, HL06, KSK00, KT07, VS00]. Lemeshow [Ko00]. length [CGG09, cICY04, CA00b, FL08, KOS05, MPP07, MdUÁ04, SG08, ST08, SA02, vEKO00]. lengths [CWW03, MR01]. less [Efr09, Hur05]. Letter [Kag01b]. level [AG04, AFH07, AEK09a, Bar08, But05, But07, But08, But09, CG03, CLLZ06, CL08, Dey09, EKL07, FLM00, FPR00a, GKM06, GGS02, HP07, HB04, Jac09, Lia06c, LS08, LLZ07, LM00b, LHZ03, NB08, PP02a, PCK07, PR08a, Sch04b, SB04, TAGF07, TGM07, XL06a, YM02, YR03, ZP00, vdV08, vdVD09]. levels [BH08b, CFQ05, Don04, WJ08]. Levinson [Sha08]. Levy [Pom00, IS02b, Shi09a]. liar [Kat02]. Life [Ano09-48, MHM07, AB05, AM04b, BCN04, BM06e, CX07, DGLS05, DGL07, ES00, FWB08, FL09, HKMS02, LWB09, LS01, MA05, Pan06, RA08, VM04, WJ08, WG01, YZT03, ZP04, ZP05, dHD^{+}09]. life-testing [FWB08]. lifetime [Asa06, IJK08, Muk06, UG02, Vie09]. lifetimes [KS07a, Tar07]. like [EC08, FPR00b, NMM05, OM07]. Likelihood [BC02, EKT06, FW07, Gar01, GS09a, Kal07, Nav08, PSV06, Par00a, WWN05, XL06b, Zha09d, Zha09e, AFW05, BM01a, BKZ09, BK09b, BW03, CP05a, CK05, CC08, CA00a, CS02c, DG09b, EY08, Eks08, FCCN04, FCCN05, FLCN05, FR07, FCL08, FWW09, GPC01, Gar05, GM01a, GRV08, Gre08, GG09a, HS07b, Hyr07, Jar07, KS03a, KSL03, KLA08, KK09d, KZ09, KQ02, LT02, Li01, Lin07, Lin08, LLC09, LS04, LZ07, LCZ06, MY04, MC08, MS02d, Muk06, MN03, ND05, Obi06, PS06, PW06c, PLK^{+}06, PZY09, PQ09, PW02b, PW03, Pla09, RBV00, Sag05, Sar06, SW01, SH06, Shi00, SL00, SV05, Sør08, SY09, TC08, Viš03, Wal02, WL02, WvEZ04, WLCI09, XL08a, YXC09, YL07, YZ06, ZR05, Zha06, ZRM06, ZGH08, dC04, vdH05]. Likelihood-based [Kal07, PSV06, WWN05, Zha09d, CA00a]. likelihood-ratio [vdH05]. likelihoods [BGV08, SSP03]. LIL [LJ07b]. Limit [Bar07, GLS09, GHH00, HS00c, PF01, AAJ00, DGS05, FZL00, GM05, HdGT09, Hil09, KV02, MT00, MS09b, Min07, RM05b, Saj00, TL09b]. limited [BK02]. Limiting [WH02, FC04, HP02, Ros07]. limits [BR03a, Has09, KL06, KLX09, Pfa00]. line [BE01, CDD01, CH08, DG04, DFG04, GB04, JHCH05]. Linear [DL07, IP09, RW05, WL01, AA08, ASBZ04, AKP^{+}07, APdR02, ARV03, ATS02, AVGRZ00, BBKN07, BR03b, BM02a, BM05a, BM06b, Ber07, BI02, BGH09, BK09b, BB04c, Bla00, BW05, Bos07, BB03b, BW00b, Bur01, Bur07, BL08c, Cai02, Cao09, cIC01, CIS00, CK01, CISW03, CB06, CL06, CIPS07, CdA09, CO08, CM00b, DEH00a, Dau04, DS01, DH08, DS02a, DCC01, DM08b, DP08b, DGGJ07, DH03, DK04b, El 04, EPS08, FZL00, FG06, FR07, GG08, GS09a, GP04, GFM05, GGHS04, GGHS07, GR01, GG09b, HLW07, HLAK09, HH09, HP07, HHKS04, HS08a, Hua02, Ino03, cIWsL05, Jia03a, KL01, KZ03b, KK02, KRS07, KI00, KS00, KQ02, KS09b, Krä05, KKM01, KT07, LDC09, LM04, LLLH05, Lee05, Li03, LJC07, Lia00a, LHS00, LJ09, Lia03, LS02, LL07c, LL08, Liu00b, LJZ^{+}07]. linear [Liu09, Lu08, LB08b, Ma00, Mai01, Mal02, Mas03, MW02, MaL09, MPU06, MPS07, MM07, MZ07, MK08, MN03, MSW04, Nie07, NSV02, NWS01, NDF03, OKD06, Oht00, OGACR08, Pap01, PP09, PZY09, PZ04, PKZ06, PSW00, QJ00, QZF09, Rao02a, RZ05, RH08, RS07a, RL09, SB05, Sch00a, Sch07a, SC04, SZ08b, Shi05, Shi09b, Sri02a, Ste09, SH02, yS05, SS02c, SS02d, SSvRM03, SW09, TR08, TS08, TL07, TSH06, Tri08, TB09, TP03, Vex06, WO00, WM02, WW02a, WR02, WX08b, Wil05, Wil07, Wit01, WLCI09, Wu03, WN06, WC08, XL08a, XW08, YO05a, Yan08, YXC09, YWS09, YC07, YZ09, YR00, Zac05b, Zha00, ZK09b, ZZF08, ZH08]. Linearity [DW00, WG01, AZL00, HP06a, PKZ06]. lines [HLW07]. link [FWB08, NDF03]. linkage [PS05, RS07b]. links [CM00b]. Liouville [Lan00]. List [Ano01m, Gho02, Tan09a, Tan09b]. literature [CIPS07]. lives [BCR00].

LM [BB01b]. load [DL00b, YJS09]. load-sharing [YJS09]. loadings [LC07]. Local [DN03, Hwa04, Kee05, PZ04, WY06, WG03b, AAJ00, CC08, CA00a, Deh00b, DCC01, HH09, HS09, HYCY03, KV02, Loy01, MM06, OP04, PP02b, PW06c, PLK^{+}06, PF01, Pfa00, SG08, Zha04a]. localized [CBG04, CBD05]. Locally [FWW06, Ber09, Gli08, HY03b, HT06, KA06, SK07a, SR05, WCS01, WGCI09]. Location [Sri02f, AP07, ASBZ04, Ars04, BP07, BC02, BZ06, BP02a, Car02, EHM09, FM05b, Fre07a, KK05, Kur08, LNO08, Lia06c, Mi06, MZ07, Ozt05, Pen00, PK05, RHL04, Sch04b, SS02e, Ten07, Ten09, TS08, Wan05, WW05, ZS00]. location-scale [Mi06, Ten07, Ten09]. Location-shifts [Sri02f]. locations [PF02]. loci [ZCL09]. locus [ZG05]. Log [AKP^{+}07, FR07, Mal02, AP07, BL08c, Che04c, EHM09, Jar07, KS03a, KI09, Kle04a, KGM05, PR06, RD08, SZ08b, Shi00, SR03]. Log-efficient [Mal02]. log-gamma [Che04c]. log-Gaussian [RD08]. log-likelihood [Jar07, KS03a, Shi00]. Log-linear [AKP^{+}07, FR07, BL08c]. log-location-scale [AP07, EHM09]. log-log [KI09]. log-logistic [Kle04a]. log-normal [KGM05, SZ08b]. logarithm [Nec06, XL08a]. logistic [AW09b, Bon08, CD03a, CC04, CC06, CB06, FS02b, GRC09, GGS03, GP07c, GV05, HH01, HPV08, Kle04a, KMR01, LP01, LM08, LH08, MS01, NS00, OTB06, PPdCP06, PZ08, SM07, TG07b, WMES07, YSF03, YGJ05, Zha06]. logit [TC08]. loglikelihood [Son01]. loglinear [CP02b, GNP07, JW09, PP05]. lognormal [KM03, LB06, LWB09, Lon09]. Long [PSJ07, BGS03, Ber09, CR02, FARMA08, GK07a, HT09b, HW08, IL08, KS00, KS09b, Kul09, LMD^{+}01, LT06, LX07b, MRCN09, Sub01]. long-memory [BGS03, Ber09, CR02, HT09b, KS00, MRCN09]. long-range [FARMA08, IL08, Kul09]. long-term [HW08, LMD^{+}01, Sub01]. longest [FL08, MPP07]. Longitudinal [HCS04, BYW07, BL08c, CN02a, Dan05b, EN06, EN07, FM05a, HW08, LH00, LW09, PW06c, QZF09, VV05, XW08, Yao08, YvdL06, ZZF08]. look [NH09, Wal04, ZP05]. loop [OCR^{+}05]. looser [Vin02]. Lorenz [DKSZ07, Kle04a]. loser [SCZ07]. loss [AJMP09, Cao09, EY08, FOS08, GKK08, KSK00, KT07, LT04, WZ04, ZK09b]. losses [FPR00b, TK06]. low [CvdL01, JÜ08, JMJ09, LLFF00, dBJM09]. low-dimensional [LLFF00]. Lower [AGK00, Efr09, FTY08, GR08a, Kle03, MdCC05, ZN08, BBW06, DM08a, Din03, Efr07b, Lia00b, LBCM06, MS05, Pet07, WY07a]. lower-bound [Efr07b, MS05]. lower-order [WY07a]. LS [MaL09]. LSE [Che00a, LGV09]. LSSOM [Ge01a].

M [Aga02, ALW07, Can06, El 04, Khu09, Lah00, Par09b, Pie09, QKO08, SA02, Sri02a, vdG02, AWL04, Sen02]. M-estimation [El 04, vdG02]. M-estimators [Can06]. M. [Ano07-45]. M/G/1 [AWL04]. M/G/1/ [Sen02]. MA [SYZ07]. magnitude [DW06, Kir07b]. Mahalanobis [KL08].

Mahonian [Raw02]. main [GGHS04, HP07, KM08, Lia06c, SST02, TSR03, TGM07]. main-effects [TGM07]. maintenance [ABQG09, DG09a, LSL06, MCM^{+}07]. majority [Wie02, WA07]. majorization [FZ04]. making [Bul07a, WPV04]. MAMSE [Pla09]. management [ABQG09, Jia06, Sin03]. Mandatory [Ree00]. Mandel [RBV00]. Mangat [KTA06]. manifolds [BP02a, Jup09]. Mann [BB04a, TB07, XP02]. Mann-Whitney-Wilcoxon [XP02]. MANOVA [Fuj02]. Mantel [FY05]. manufacturing [AC06]. many [BL05d, FD02, SK07b]. many-to-many [BL05d]. mapping [Mez03, ZCL09, ZNWS09]. mappings [MdCC05]. March [Ano00-32, Ano01o, Ano02v, Ano04-52, Ano06-47, Ano08-43, Ano09-45]. margin [Par09a]. margin-based [Par09a]. marginal [Bor00, DTW06, GS07a, NvdL07, QRR08, XW08, YvdL06]. marginals [Ano09a, FCM07, KP05]. margins [Che04d]. mark [EKT06, GSC07]. marked [GSC07, Gua08]. market [Vin02]. markets [Pot07]. Markov [ANN08, Alv05, AT02, CS03a, CGG09, CLE09, CT09, DLS09, DZ06, EZ09, FFH00, FI05, Gil06, Har02, IP09, IVM03, Kem05, LC07, MW00a, MM06, OL03, PP02b, RM07, SL05, SYM05, Spe09, SGJM02, Tan08a, TL09b, Whi05, WW02b]. Markov-switching [EZ09]. Markovian [Sen02]. marks [BD08b]. Marshall [Mei07b]. Marshall-Olkin [Mei07b]. Martingale [WW02b]. masked [Muk06]. mass [ABS00, ZZF09]. match [HA03]. matched [GMM03a, IVM03, LP01, LTW08, OM07, WWL08]. matched-pair [LTW08]. matched-pairs [IVM03]. matching [Das05, FR02]. material [LuV09]. materials [PP06a]. Mathematical [LRZ02, Mas01]. matric [DGGJ09]. Matrices [Bed00, ATS02, Bed98, BH09b, BK09d, Cha03a, KLR07, KKK03, Kuw02, MMP00, Pro02, Sch01, Sch07b, Sri07a, yS07, SZ06, VKO00, Win00, YCB06, ZDLZ02, Zog00]. Matrix [LLP09, BSP02, CSW02, CDH06, DGGJ06, DGGJ09, DR08b, FZ00, GP06, IPS08, Kon07, KvR08, KT08b, OS05a, PR06, PSW00, SN03, TK06, WZ01a, WZ01b, WC00, WN06]. matrix-based [PSW00]. matrix-variate [WC00]. matroids [Sza02]. Matsumoto [BW07]. Matsumoto-Yor [BW07]. Matusita [OO08]. Max [LLP02, HdGT09, IL06, NPA06, SS02a, SS09]. max-domains [NPA06]. Max-min [LLP02]. max-semistable [HdGT09]. max-stable [SS09]. max-type [IL06]. maxima [FR01, Hai00, HdGT09, PF01]. maximal [BK09b, KP02, NB05, Ruk06]. maximally [CR08]. Maximin [DLRP06, NMB02, Duc00, MP07d, OTB06]. maximize [War08]. maximizers [FH05]. Maximum [Gre08, Jar07, Muk06, PQ09, SW01, AMY02, BM01a, BW03, Böh02, CP05a, CS00, DK07, EY08, Eks08, EG00, FR07, FG02, GLS09, HPY02, HP00a, Hos07a, HS07b, KS03a, KSL03, KZ09, KQ02, LL07b, LT08b, LZ07, MY04, NPA06, PF02, PW03, Qu06, RJT05, RBV00, Sag05, Sar06, SL00, SY09, TSS06, WvEZ04, XL08a, YB08, ZBZ00, ZR05, Zha06, ZRM06]. Maximum-likelihood [Gre08, Sag05]. May [Ano00-37, Ano00z, Ano02w, Ano03-57, Ano05-59, Ano08-37]. MBMUDs [BT04]. Mean [AB05, Rao02c, Ryc01, ATS02, Asa06, BB04b, BLL06, BM06c, BK05a, BCR00, CX07, CP02a, CP05a, CP07, DBDG09, FOS08, GP06, GKK04, GS07d, GK08, Har09, HS00a, HKMS02, KZ03b, Kha00, KMS09b, Laï05, Lan01, LJC07, Lia06a, LL05b, LLW07, LW09, Mei05, MS07a, MS07b, Nie01, OM08, Pen03, Raq05, RA08, RD08, RBV00, SN08, SZ08b, SM05b, Tse04, dSSdL08, WO00, WZ04, WZ03, WBvdL07, WZ06, WG01, WM08, Wu06, YZT03]. mean-square [CP05a]. mean-squared [WO00]. Mean-variance [Ryc01, KMS09b]. means [ALRV01, ABK00, BR02, BH05, BL02b, Cua09, HT09a, Hei00, KP09a, KM03, KT04, KT08a, LL04, LG06, Mas03, Mas09, MT04, Net02, OS05b, QSL09, QT07, Rao02c, Sch05a, STW08, TW05, TZ00, WZ09]. measure [AB09, ES02, GRR07, Loy01, MLK09, MNOP05, Son01, XF00]. Measurement [HH07, BL02a, Dup05, EOPR05, GS07b, Koh06, KS08b, KSW05, KMS09b, Lia00a, LC09a, LTI00, NLP06, TL07, TB09, YL07, ZY07]. measurements [AH02, BL08c, Che06c, Cho08, DR01b, KRV06, Kus03]. Measures [XC06, Zog00, Aus05, BM03, BDGV01, BL02a, BL03, DH06, FRF09, FT09, Gas07, Gre08, Gup03, HC06a, HLH09, KC05, KA09, LR09, PPdCP06, RK05, RO05b, Sag05, SVO05, Ser04, Shi09a, Ste04a, Vas03, WL01, XD05, Zog08, ZS00]. median [AZ04a, AKK09, ARV03, Cha00, CT04, DLM02, Dub07, GSY03, HPV08, WH02]. median-based [AZ04a, ARV03]. medians [SM07]. medical [Sen08]. Meier [Sub04b]. melanoma [RdCCB09]. Memory [Azo09, BGS03, Ber09, CR02, GK07a, HT09b, KS00, KS09b, LT06, LX07b, MRCN09, WS09]. Mendelsohn [ABZ00, BKLZ01, Ben01, Mis02]. Menten [LFTT08]. Merge [HH08]. meta [Bar03, KKB09a, MBA^{+}08, RDS08]. meta-analysis [Bar03, KKB09a, RDS08]. meta-analytic [MBA^{+}08]. Metaheuristics [Ano09-47]. metamorphosis [LR02, LT08b]. method [AG04, BM05b, BJZ09, CC08, Cho08, CD03b, Det02, DGR07, DBDG09, EC08, Fab00, FC03, FZ06, FMT04, FBK03, GJ01, GO07, HLAK09, HB04, HA08, Jia03b, KA06, Kas00, KT08c, LGV09, Lin07, LM00b, Mar09, MPS09, Neu00, O’Q06, PY09, PMS^{+}01, SJ09, Viš03, WA07, YL07, ZZL07, ZK09a, ZCL09]. methodologies [Hay06, MS02e, YA09]. methodology [ES03]. Methods [HMW04, Aic01, BP07, BH08b, Ber01a, BD05, CG01, DEH00a, DR08b, FTY04, FS07, GPC01, GC05, GKM07, HC06a, JK08, KD08, Lah00, LB08a, Lin01a, LJZ^{+}07, Mat01b, MB00b, PW06c, PLLS08, PW03, Pre07, RMMA07, SSO07, Sas01, SSC07, Tha08, TSR03, VOK^{+}03, YSB09, Zha09e]. metric [KS09a]. Metropolis [CT02]. Michael [Cha01a]. Michaelis [LFTT08]. microarray [CIC08, DR08b, GBC^{+}05]. microdata [Rei05]. mid [YLH04]. middle [IJK08, JI04]. middle-censored [IJK08, JI04]. Midzuno [Sri02f]. Min [SS02a, LLP02]. minimal [BE00]. minimally [CD07b]. Minimax [Ber04, Bla00, GSM00, LM03, SY08b, TK09, Wil05, Wil07, ZK09b, Ago09, BLL06, BW00b, BR08, IL04, Kle03, KMS08, KSK00, Lan01, LX07b, Li07, Lia00b, MS06, NH09, NS00, RB07, SYZ07, Tse04, YWS09, YZ07, Zho08, ZW03b]. minimax-optimal [KMS08]. minimaxity [XF00]. minimization [GP06]. minimizers [BC03c]. Minimum [AYZ06, AK08, ALS07, lCV06, Col01, FZ04, FMN09, JW09, KW09a, KN04, MS00, PPZ02, RS04b, Sun03, Tom08, WO00, XL06a, AZ04b, AEK09a, ALS04, AD01, BS08, But06, CGG09, CP07, HP00a, Jac04, JMJ09, Lan00, LSW06, LDT04, LL07b, MS05, Mei09, PF02, TD03, Tan07, VFVF02]. Minimum-area [FMN09]. Mining [AC04a]. misclassification [Gre08, SSY05]. Mises [DH09b, FCL08, IL04]. missing [BD08b, BZ06, BCVV09, BGMPG09, FBK03, GO07, HdGT09, Lee01b, Moj07, PW09, PM01, PMS^{+}01, Sub03, Sub04a, SW09, TSH06, Wan08, WGCI09, YXC09, Zaf02a, ZR05]. missingness [BYW07]. misspecified [AW09b, BR03a, OKD06, WX08b]. mixed [AA06, AH02, AZ04b, AYZ06, AFH07, AZL00, Bur07, CFQ05, CISW03, CL08, DS02a, FM05a, GS09a, HY08, Hös08, Jia03a, Kas00, KW05, KP05, LDC09, LLLH05, LL07c, Mai01, MP02, MM07, Ngu08, Nie07, PW06c, PF09, PR08b, RZ05, Roo03, SZ07, Sch07a, SK07a, SS08, ySK05, yS07, SSvRM03, TAGF07, Vid06, WRV05, WJ09, Web07, WYL09b, WB05, YM02, ZH08, Zog08, vdVD09]. mixed-effects [AA06, Nie07, PW06c]. mixed-level [AFH07, CL08, TAGF07, YM02, vdVD09]. mixed-type [AZL00]. mixing [ANR00, CZ09b, FZ00, KL05, LZ07, SC05b, XC08]. mixture [AH03, AGP09, AGHMPV07, CK05, CIC08, CQ08, Dor09, FCL08, FE06, Gar01, Gar05, HS05, Han08a, Has09, HC07b, KW09a, Kle04b, KS07d, LS04, MS00, MW07a, ML03, Net02, NK09, Pal03, PPM09, PK05, Phi02, QT07, WG03b, YO09, ZCL09, dCD09]. Mixtures [GMM03b, AA08, Car02, CS08, Dat07, FMS04, KX01, KM07a, KP08a, Kot06, KvdV08, Kur08, Lij03, Nav08, QSL09, TG07a]. ML [MP02, VV05]. mle [FJT09, AGHMPV07, GPL09, Mi06, Nie07]. MLEs [CP02b, PP05]. MLHD [ZN08]. mode [BBKN07, CJRV08, MP05, YO05b, Zie03]. Model [CGY07, CD02, CP02b, HK03, JLNY07, JLNY09, KC05, PK01, Shi09a, Zha09c, AH02, AFF09, AL09, Atk08a, BL06, BCN04, BX07b, BX07a, BH08a, Ban03, BM05a, BM06b, BN03, BGM03, BCVV09, Ber02a, BD04, BPR09, BW05, Bon08, BB06b, BK02, BV01, BV03, BL02b, Cao09, CH05a, CL02, CISW03, CTL04, CQ08, Che04c, CC04, CS02d, Con08, CL09b, CMN06, CDB^{+}09, Cow03, CO02, DLS09, DH08, DM08b, DB01, DL06b, DGGJ07, DG09a, DPR06, DP08c, Dub07, Dup05, Dup09, DL09b, DR02, FL03, FM05a, Fie04, FPvdVvH06, FE06, GB09, GP07b, GS07a, GSCG09, GH03, GDL07a, GDL07b, GKK08, GP04, GG07, GG09c, GV05, GPW01, Hag06, HS05, HB00, HK05, HP07, HY08, HS08a, Hua02, IPdUÁ08, Ino03, cIWsL05, IP09, JL09a, JL04, KW09a]. model [KW04, KW05, KTA06, Kle04b, KZ09, KK03, KMR01, Kot05, KQ02, KN04, KS08b, KP04, KG05, KT07, KMV05, KMS09b, Laï05, LGV09, LK00, LWL^{+}09, Lee05, Len09, Let08, LQH03, LT08a, LHS00, Lia03, LN03, LSL06, LP04, Liu00b, LS04, Liu09, LFTT08, LSD01, LS01, M.04, MW03, Mal02, MP02, MP07a, MLK09, MaL09, MMM06, Nag09, NK06, Net02, NHZ04, NWS01, NS00, OM08, OTB06, PM08a, PP09, PF09, PY06, PS05, PKZ06, PKM08, QZF09, QT07, RW05, RZ05, RH08, SB05, SZ08a, SC04, SK07a, SAC02, SR00, SSC07, SH06, SSM00, SS03, SS09, Sør08, SLS02, SS08, Sta09b, SW09, Sun03, Sun06a, Sun06b, Sun09, TR05, Tan09a, Tan09b, TL07, Tom09, TGM07, TL08, TC08, TP03, dSSdL08, UB07, VdBS01, VM04, VEM07, VR08, Vex06, VIBAV06, Vid08, VG07]. model [VG03, WSY06, Wan09, WD04, WZ06, WC09b, Wil05, WC01, WC08, WWL08, XL08a, XF00, Xu07, XY06, YXC09, YWS09, YHA09b, YHA09c, YW09c, YL07, YZ06, YGJ05, YZT03, ZT04, Zha04b, ZRM06, ZWHF07, ZY07, Zie03, ZGH08, dUÁGMCS00, vdH05]. model-assisted [Dub07]. model-based [Kot05]. model-comparison [Cow03]. Model-complete [PK01]. model-fitting [Cow03]. model-free [GSCG09, WC09b]. Model-robust [JLNY09, JLNY07]. modeler [Wes00]. Modeling [ANN08, BBSZ05, CA03, GN03, KK09a, AKI08, BF07, CBC^{+}07, EOGQ09, KI09, Ko00, KS07d, Leh09, LL07c, LW09, MGB05, NaX03, NK09, RWF04, SK08, SH02, SH03, Zas06, Con02a]. Modelling [EZ09, LSL06, MJD09, AGP09, BGP00, BJS09, CLE09, GS07a, GR09a, GR09b, IOQ09, LHL09, VV05, Yao08, Roy07]. Models [Ano09-48, ÅvR06, AW09a, ANN08, AA06, AW09b, AP07, ASBZ04, ACW02, AdL05, AP05, AKP^{+}07, AH03, APdR02, ARV03, AVGRZ00, AC06, Arn09, AGHMPV07, Atk08b, Azo09, BM02a, BLSvdDS05, BLL06, BZ06, BLV00, BK09b, BBdR00, BB04c, BDZ07, BDP07, BW03, BO04, BGI03, BB06b, BYW07, BB03b, BW00b, Bur07, BQI06, But01, CS03a, CCH08, CGG09, CHLY01, cIC01, CYLC01, CIS00, CLHY01, CISW03, CD03a, CK05, CIC08, CZ09c, CCP09, CKM00, CC06, CL06, CIPS07, CA00a, CW05, CL09b, Con04, CdA09, Cow03, CP02b, CM00b, DMS03, DSG07, DI09, DH09b, DS01, DMN01, DJ06, DS02a, DS06, DS08, DLRP06, DP08a, Dik00, DGS05, DH03, DZ09, DS07, Dor09, DSF02, DR03b, DW08, EN06, EN07, EN09, EOPR05, EBZ08, El 04, EKT06, Ery03, FK09, Fan00]. models [FW03, FWW06, FCS05, FCCN01, FR07, FCW00, FZ00, Fre07b, FS02b, Fu05, GD08a, GP07b, Gan00, Gar02, GB07, GACS07, Ger05, GRZ08, GGCA00, GML06, GS07b, GS09a, Gil04, GB00, GR08a, GGHS04, GGHS07, Gua08, GH01, GL02a, GK06a, GP07c, GNP07, HPY02, HC03, Han08a, HT09a, HOS06a, HLAK09, HZN04, Hon04, HL04, HHKS04, Hös08, HT08, HT09b, HLP06, HS09, Huz05, HH07, Ioa07, Jan04, Jia03a, JK08, Jin07, JW09, JK03, Kan02, KS08a, Kem05, KLA08, KRS07, KI00, KSS08, KP09b, KS09b, KMS09b, KM00b, LDC09, LSL08, Lan01, LLLH05, LM08, Lia00a, LJ09, LL08, LY05, LJZ^{+}07, Lop09, Lu08, LNS07, Ma03, Mai01, MMM07, MJD09, MW00b, MN08a, Mel05a, MW07a, MPU06, MGG09, MM07, MTC05b, MS03, MK08, MY08, MN03, NLP06, NvdL07, NW05, Nie07, NDF03]. models [OKD06, OP00, OP04, Ots08, OGACR08, PSV06, PGLS08, PT08, Pal03, PB08, PP05, PPdCP06, PW06c, PP06a, PSG07, PZY09, PZ08, PK05, Pre07, QJ00, RV03, RS07a, RdCCB09, RO05a, RS08, RM07, Saj00, SN03, Sch00a, Sch07a, SK07a, SH09, SZ08b, SFT01, SYZ07, SC08, Shi05, ST04, SSB02, SYM05, SS03, SR05, SF02, Spe09, Sri01, Ste09, Sub03, Sub04a, SSvRM03, SN04, SNM08, TMP06, Tan01b, TCD08, TRJ07, TT09, TSH06, Tri08, TB09, TZ05, TK06, UV07, Vas03, VdCO00, VIBAV06, Vid06, Wal04, WM02, WW02a, WG03a, WRV05, WMSY06, WJ09, WWEL08, Wei08, Wie00, WX08b, WW02c, WLCI09, WGCI09, Wu00, WN06, WYL09b, WB05, XC06, XW07, XL08b, XW08, YSF03, YO05a, Yan05, YSH03, Yan08, YS08, YW09b, YW04, YR03, YC07, YZ09, YR00, YvdL06, ZT04, Zha06]. models [ZCL09, Zha09a, Zha00, ZZF08, ZH08, ZEB07, dC02, vdH05]. Moderate [Hui09, MaL09, MP05]. modes [BM06e, GKDN06]. modifications [Fab00]. Modified [BGV08, CS06, CK05, DPA08, FCL08, BM05b, FCCN04, FCCN05, Jac04, Kwo01, KHC02, Pal09, SV05, UV07, WRV05, WH08]. Moment [MP07b, MA05, YB06, KT08c, Mei07a]. Moment-based [MP07b, YB06]. Moments [Ahm01, Cha05, LB06, PP02a, Sul05, AF03, AM04b, AC07, BES05, BS07a, BJZ09, CR08, DGR07, ES04, Eri02, Hos06b, Hos07a, Hos07b, Jia03b, JB02, Jon04, Pro04, Raq04, SC09, WY07a, WC00]. Monitoring [HHKS04, DFG04, SG08]. monotone [And05, BZ06, LG06, PB08, Pen04, WH02]. Monotonicity [CS04b, DSF02, Hua02, MP06, Sar00]. Monotonous [TSN07]. Monro [MPS05]. Monte [And05, Aus05, BM05b, CG01, DEH00a, DPR06, MW00a, Tan08a]. Monte-Carlo [And05]. Montgomery [Atk09, Bor09, Jon09, Khu09, Luc09, Par09b, Pie09]. Moors [SSM00]. moral [CS03a]. mortality [Con09, LWL^{+}09, SR05]. Most [BK05a, Ber01b, GL02b, Pew05]. motion [LMD^{+}01, LL05a, LS01]. motions [Tan08b]. Motoo [MT00]. Moving [HPY02, BPAK04, KV06, SW08]. moving-average [KV06]. Moving-maximum [HPY02]. MRL [ZB09a]. MSE [Nam00]. MSEM [Liu00a]. MTD [MW07a]. Mukerjee [Dea00]. multi [AG04, AG07, BF07, Cox03, DTW06, FLM00, FL08, GKM06, GGHS07, KMS08, LLZ07, LHZ03, MdCC05, PCK07, Rei05, RO05a, SFT01, SCZ07, VdBS01, Ver02]. multi-component [Rei05]. multi-country [Ver02]. multi-criteria [SFT01]. multi-cycle [VdBS01]. multi-dimensional [Cox03]. multi-factor [GGHS07, RO05a]. multi-level [AG04, FLM00, GKM06, LLZ07, LHZ03, PCK07]. multi-objective [KMS08]. multi-resolution [BF07]. multi-state [FL08]. multi-treatment [SCZ07]. multi-valued [MdCC05]. multi-way [AG07, DTW06]. multicenter [GK01]. multicriterion [FG06]. multidimensional [Bor00, YC02]. multigraphs [Chu01]. multihalver [FMT04]. multilevel [AdL05, LNO08, PW06a, PW06b, RGL07, SC08, dL06a]. Multinomial [BLSvdDS05, TC08, FW07, GD08b, JW09, LY02, PPZ02, Ruk06, SF02, SNM08]. multinormal [ATS02]. multinormality [LLFF00]. multiparameter [WT08]. Multiple [And07a, CC00, CS04a, GS00, KKM01, LC09a, Nak00, PFF00, SC05a, BBKN07, BKZ09, BH05, BWA06, CKS07a, CB06, CS04b, CKS07b, CSX08, DL05, DNK02, DBG06, DHV01, DPA08, EM07, EKT06, FR01, FWW09, GACL08, GA08, Gor07, GP03, GH07, Hay01, Hay07, HLAK09, HC06b, HA08, KLE^{+}04c, KHC02, LT04, Li08, LC00, LLP02, LK03, LW07, MS02b, MCH^{+}05, Nak09, NK06, PvdL04, SB06, SN08, Tro00, WGCX09, WMES07, WK01, WW05, WC08, ZY01, ZS08, ZH09]. multiple-comparison [Nak09]. Multiple-objective [LC09a]. Multiplicative [AH01, TMM09, CSW02, VS00, vEKO00]. multiplicatively [LB06]. multiplier [CIPS07, Qiu01]. multipliers [HP00a]. multiply [Rei05]. multipopulation [Kis02]. multiresponse [NMB02]. multiscale [Shi08]. multiserver [ALH07]. Multistage [ASAO02, MS02e, BL08b, FWW06]. multistate [dSSdL08]. multitype [Hyr07]. Multivariate [ACS08, BRdCR03, BW07, Böh02, CF04, HL09a, Koh08, Mas09, Min07, NSV02, RDS08, Tri08, ZS08, Ano09a, ACS07, Arn09, Ars04, AG09, BH08b, Ber02a, BB03b, BMP02, BMW09, CC09, Cao09, CN02a, Cha03b, CS02a, CD03a, CSW02, CA03, CD03b, CSX08, DR08a, DS08, Ery03, Fab09, FZL00, FJT09, Fuj02, GP07a, HP06a, HKP08, HR04, HP06b, HP02, cIWsL05, JI07, KI08, Kal07, KM07a, KS09a, KN02, KSS08, KMV05, LM04, LR09, LJC07, LL07c, LLW07, Lue09, MS02a, McF01, MPU06, MFSW07, MGG09, MPS09, MK08, MB00b, NK09, NWS01, PFF00, PG07a, PW02a, RJT05, RK05, Ser04, SR02, SA07, SSvRM03, TRJ07, Tom08, Tse04, VV05, Vel03a, VOK^{+}03, WZ04, WL02, WG06, WN06, YB06, Zez09, Zog00, ZS00]. mutation [GPC01]. mutations [AKP^{+}07]. Mutual [BG07, BdM09, BG09]. Mutually [CD01, BM01c]. MV [Jac02, Jac03, LFRL07]. MV-optimal [Jac03, Jac02]. MV-optimality [LFRL07].

N [AG07, McF01, Sat07]. Naik [McF01]. naive [Zaf02b]. Narayana [Moh02b, Sul02, Sul05]. Nath [MHM07, Sri07b]. natural [TT09]. NBU [DL06b]. NBUC [AM04b]. NBUCA [AAEK06]. NC [McF01]. Near [MC08, AC07, AFW05, AEK09b, BS05a, BS08, BR09, DB08, GC07, MH00, SB04]. Near-exact [MC08, AC07, GC07]. near-normal [MH00]. near-optimal [SB04]. nearest [CM00a]. nearest-neighbor [CM00a]. nearly [BJ06, CKM00, TS01, WH01]. Necessary [BLV00, GKMS02, SC05b]. Negative [BSW08, XL08b, MS00, Obi06, SD01, WZ03, ZBZ00]. negatively [Xin09]. negatives [Sar04]. Neighbor [Udd07, AYH09, CM00a, Mis07]. neighbor-balanced [AYH09]. Neighborhood [DF00]. neighborhoods [AK03]. neighbour [KS07e]. neighbourhood [GGS09a]. nested [DS01, FHKMS02, HJ02, HS09, KM01, KKL05, Mut05, WM02, WWEL08, YR00, vdH05]. network [KS07c]. networks [AKI08, DDvdL08, HDM09]. neural [DDvdL08]. neutrality [KMPP02]. Neyman [Aug02, BB01b, FGT08, Kra09]. NGINAR [RBN09]. nine [KdR01]. Ninth [Goo08]. no [Gan00, GR08b, PSJ07]. no-interaction [Gan00]. nodes [SV00, Udd05]. noise [CR02, GN03, GDL07a, Kun00, NH09]. noised [YZ06]. noisy [DG04, KV07]. nomination [FMT04]. Non [BR00, BW05, Bor04, DSG07, Gra09, Gri05, Kag08, Li03, MM06, Nek00, Ruk07, Ahm05b, BYW07, CA04, Can06, Cav06, CB06, DL04, DW00, EN02, FL03, FZ04, FPR00b, GHSN00, Har09, HT06, Jac09, KMPP02, KR07, LTW08, LY09, LMS08, LLC09, LBS06, LBR08, Ma09, MW00b, NK06, PX09, PKZ06, Pou01, RD08, Sen02, SK08, SJ08, TT07, Tan01b, TL09a, WZ03, Win00, Yek08]. Non- [BR00]. non-adjacent [DW00]. non-central [LY09, SJ08]. non-centrality [FPR00b, LY09]. non-compact [Cav06]. non-constant [BW05, RD08]. non-continuous [KR07]. non-existence [Win00]. non-extreme [LBR08]. non-Gaussian [HT06, Ma09, TT07, Tan01b, TL09a]. non-i.i.d [MW00b]. non-identical [CB06]. Non-identifiable [DSG07]. non-ignorable [BYW07]. non-iid [PX09]. non-increasing [EN02]. non-inferiority [LTW08]. non-isomorphic [FZ04]. non-lattice [KR07]. Non-linear [BW05, Li03]. non-linearity [PKZ06]. non-Markovian [Sen02]. non-negative [WZ03]. non-normal [Ahm05b]. non-normality [GHSN00, Har09, SK08]. non-opinionated [KMPP02]. non-orthogonal [LMS08]. Non-parametric [Bor04, Gri05, Ruk07, CA04, DL04, FL03, KMPP02, Pou01]. Non-Poisson [Nek00]. non-positively [Yek08]. non-regular [Jac09, LBS06]. Non-rejective [Gra09]. non-separable [LLC09]. non-standard [Can06]. non-stationary [NK06]. Non-validity [Kag08]. non-zero [LTW08]. Nonadditivity [PP05]. Noncanonical [CM00b]. noncompliance [Zha09e]. Nonconjugate [FD02]. nonequispaced [Chi03]. nonexistence [AM01, HH02]. nonignorable [ZC03]. Noninformative [GGCA00, GH03, KKL05, NM06, Mee05, SN04, WHS07]. nonisomorphic [EKL07]. Nonlinear [AKI08, Fu05, HLH09, KS02b, Koh03, AC03, Atk08b, CLHY01, CTL04, Cíz06, CC02, CW05, FH07, GP06, HZN04, IL08, KC05, Laï05, Mel05a, NW05, Nie07, Páz02, Tan01b, TRJ07, VG07, WJ09, YMJ08]. nonmonotone [BLBR00]. Nonnegative [SSvRM03, KSK00]. nonnormal [Bat04, WW02c]. nonnormality [Fuj02, KI08]. Nonparametic [BP02a]. Nonparametric [Ano07-45, AC04b, BTK09, BH08b, BGS03, BT06, CG01, CY04b, CCS07, CA00b, DLS09, GSY03, GACL08, GA08, GSFBGM04, Gho01, GK07a, HS09, IPdUÁ08, IOQ09, Koh06, Kot06, KLE^{+}04c, LX07a, LT09a, LZ09a, LB08b, LT09b, LT09c, Moj07, MB00b, NvdL07, Nie01, OL03, PvZ09, RV07, Rob07, RO05b, SL05, Sag05, Ser04, SH09, SY08a, Sto05, TKP06, Vel03a, WZ03, ZC03, Alv05, AKK09, And07b, AGK00, AGP09, BNP06, Ber04, BGMPG09, BMP02, BMW09, Che04a, CZ09b, Che04b, Che05, Cho07b, CD03b, Det02, Efr07a, FCM07, GRZ08, GRU01, GL02b, Hug08, Hwa04, Jan01, KW09a, KRS07, Koh00, KKW09, LC01a, LB08a, Lia00a, LZ07, MSW04, Mus02, NM06, Ong05, Oye03, OD06, PW06c, PG07b, RA06, XF00, YL09, Zha04a, ZK09a, Zie03, ZS00, Lin01b]. nonparametrics [Sen00]. nonperiodic [FS01b]. nonrandomized [KTB01]. nonreplicated [NMB02]. nonresponse [QRR08, Zha09e]. nonserial [HP06a]. nonsmooth [Shi08]. Nonstandard [Deh00b, El 04, LCM07]. Nonstationary [PP06b, SL00, Ste04a, Swi02, Vel03a]. Nonsubjective [GS02]. Norm [Kuw02, ANR00, Ber04, RS04a, Raq04]. Normal [LCM07, Sin06b, Ahm05b, ABK00, ACS07, BB04b, BW03, CC09, CP07, CQ08, CC02, CS00, Dat07, DS01, FMN09, FE06, GL09, GKK08, GFDMG04, HS05, IU01, JB08, JB09, Kag06, KGM05, LDC09, LL04, LG06, Lia06a, LL05b, LLW07, LN03, LL06b, LS04, MS06, MP08, Mei07a, MS07b, MTC05b, MH00, Net02, NK09, OS05b, Pen03, Phi02, QSL09, QT07, Sar06, SK07a, SZ08b, SSE04, STW08, Tro00, Tse04, TK09, WZ04, WW02c, YHZ05, ZM06, dCD09]. normalising [BS09]. normality [AD01, BC03a, BK07, Che06a, DR01a, GHSN00, GM02, Har09, Haw01, KV05, Lia00a, LJ09, MML02, Qi07, ST06, SK08, WT08, Zie03]. Normalizing [BNW05, HW05, WT06, FM00b, FM00c]. normally [BH05]. normals [FMS04]. norms [SW07]. note [Ano01x, Ano03-58, ADW03, Bab04, BS05a, Bos09, Che07, CH05b, CH08, DL04, DGGF04, EN02, GS07c, HR04, Hua02, JL09a, Jia06, Jon06, Kau05, Kha00, KNH09, LF07, LL06b, LC00, LW07, MFL03, Mas03, Pad06, PZ08, Saj00, Sar00, SSP03, Sub04a, WW08, YHA09a, ZP04]. Notes [XY08]. notion [AAEK06, DL06b, ZS00]. novel [JK03]. November [Ano00-27, Ano01n, Ano03-49, Ano04-62, Ano06-45]. NPMLE [DI09, HSV09]. NRBDs [Mor09]. NRDFs [Mor09]. nucleotide [SYM05]. nuisance [Con01, JPS02, Li01]. null [Con01, DL04, DvdW05, EMMM00, Mar07, MW05, Nag09, PvdL04, vdWD01, vdW04]. number [Aga02, ALW07, BS05a, BL05a, BH08b, BM01c, BK05b, BK02, Che02, Che06b, DG02, GLS09, GC07, KTMS04, Kun00, KN03, ML03, OT02, SW05, VBF03, WZ01a, WZ01b, YB08]. numbers [BS08, BS05b, Cha02b, DM08a, DB08, Mar05, Nak00, PP02a, Sul05, Tar07]. numerical [JPS02, SSC07, SV00, Van05].

OA [YMJ08]. OBC [Ano02t, Ano02u, Ano03-40, Ano03-41, Ano03-42, Ano03-43, Ano03-44, Ano03-45, Ano03-46, Ano03-47, Ano03-48, Ano03-37, Ano03-38, Ano03-39, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano04-44, Ano04-45, Ano04-46, Ano04-47, Ano04-48, Ano04-49, Ano04-50, Ano04-51, Ano05-41, Ano05-42, Ano05-43, Ano05-44, Ano05-45, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano06-27, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano06-37, Ano06-38, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano07-38, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano08-36, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36]. OBC [Ano09-37]. object [SGJM02]. objective [KMS08, LC09a]. observable [AdL05]. observation [SJ09]. observations [AYH09, And07b, BS05a, BS08, BB03b, DB08, EMP06, Ery03, JM08, KS08a, KK09d, KV07, ME04, Mas07, MS09b, PM01, RJT05, Shi09a, SK07b, Udd00, Udd08, Viš03, YC02]. observed [AFW05, DL09b, HKW00, Hyr07, KKM09, PW09, RSM05, Skö09]. obtain [SA03]. obtained [yS07]. Occupancy [Gar02, Tar07]. occupational [KG05]. octants [Nie05]. October [Ano00-34, Ano01q, Ano03-53, Ano05-62, Ano06-44, Ano08-41, Ano09-43]. odd [BR08, GC07, Ree00, XP01]. oddities [HS07b]. odds [SF02, Sun06b, Sun09]. off [FG06]. Ogawa [HKP08]. old [Ahm05a, Ver01]. Olkin [Mei07b]. Olman [AS00]. OLSE [Liu00b]. on-line [DFG04]. One [Cai05, KX07, LNO08, QW05, ASC05, BQI06, CMyS02, CSW02, DR03a, Dav04, Dob02, DPA08, EBZ08, Fuj02, GACL08, GGCA00, GHSN00, Hai07, HB00, HK05, HML00, HP07, HK03, JR03a, KI08, KZ03a, KKM01, KG05, KD07, MP02, NPF09, PW02a, PW02b, Sah08, Sch00b, SS01, Wan06, WJ08, WMES07, WR02, Whi05, WX08a, ZEB07, vdH05]. one-dimensional [Hai07]. one-factor [BQI06]. One-factor-at-a-time [QW05]. one-parameter [GGCA00, KZ03a]. one-point [WX08a]. One-sample [KX07, LNO08, KI08]. One-sided [Cai05, CSW02, HML00, HK03, KKM01, NPF09, PW02a, PW02b]. one-step [WR02]. one-way [Dob02, EBZ08, Fuj02, GACL08, HB00, HK05, KG05, KD07, MP02, Sah08, Sch00b, ZEB07]. one-way-classification [GHSN00]. one-way-table [Dav04]. ones [KS08a, KV02, KV05]. online [GSCG09]. Ontology [SSO07]. open [HYCY03, Hug08]. operating [KKLD07, LZ09a, PZ04]. operator [Kak06]. operators [AA08, Cav06]. opinionated [KMPP02]. opportunities [GR09b]. optical [HYP02]. optima [FED09, Fox09]. Optimal [BH09a, Bha09, BDP07, BS07b, BS05c, CHP08, Cau05, CM00a, CHLY01, CL02, CL08, CKM00, DMS03, DH00, DMP04, DZ09, DW08, DR02, FC04, FL09, Fer02b, Fre08, GD03, GTV01, GGHS04, GHSS02, HKW00, HS02, HZN04, HW09, Jac06, JM08, KZ03b, Kle04b, KKW09, Kow09, LL05a, LG06, LJ07a, LTI00, LZN06, LT08c, LFTT08, LNS07, M.04, MS09a, MS01, McL08, MU03, MP07f, NaX03, NHZ04, OP05, Ots08, RM05a, RB07, Sch00b, SFT01, SR03, ST09, Sta06, Ste05, SH05, SB04, TAGF07, Tat05, TN05, Tom09, Udd00, Udd08, VG07, WG05, WJ09, Web07, WC07, WLC08, YJS09, YS08, ZP00, AFH07, AKJ09, And06, ADW03, BB06a, BDZ07, BM06d, BD04, BD05, BW00b, Bul07b, BR08, BR09, But01, CDD01, CN02b, CKM06, Che03b, CH08, DDCC08, DH08, DB01]. optimal [DMMS07, FC03, FWW06, GGS09a, Gli08, GGS09b, GHSN00, GT06, GRP02, HC03, HP07, Hir02, HC07a, Jac02, JHCH05, Jeo00, JCS05, JLN08, KM08, KS06, KMS08, KMS09a, LM08, LZ00, MM02, MP02, MTC05a, MT06, MW08a, Mei09, MSW09, Mel05a, MR07, Nay03, NS00, Orr08, PM08b, Pou01, RVZRZ07, RGR09, SR05, TV01, Udd05, WMSY06, Wu03, Yan05, Yan08, YL09, YLS00, ZT04, Zho08, ZH09, vBP03, Jac03]. Optimality [AYH09, BC07, KMS09b, KM00b, OJKM02, YM02, BV06, BT04, Bur08, GGHS07, Har08, LC01b, LX07b, Li07, MJS00, OP04, Pfa03, QKO08, SY08b, LFRL07]. Optimally [Bet01, WK01, GS00]. Optimising [JR03b]. Optimization [AM00, Ano09-47, EKK06, LS06, Mor09, Ral02, Zha09c, dHD^{+}09]. Optimum [Atk09, CLHY01, DDM09, HYP02, Atk08b, Pro06, RO05a, TV04, TMM09, UB07, Atk08a]. option [LL05a, TT07]. Optional [CS05]. oracle [Efr07b]. oral [Sta06]. orbits [BE01]. Orcutt [ST00]. Ord [Krä05]. Ord-type [Krä05]. Order [BM01a, BK09a, DLOR02, JB09, Mi08b, NMM05, AW09a, AF03, Ahs00, AN00, AH03, BCP03, BS04, BS05a, BD08a, BD09, BLL06, BES05, Bar07, BAA08, BJS09, BRdCR03, BLPR05, BI02, BN09, Bie08, BG05, BZ00, CGG09, CP05a, CP07, Che07, CB06, CK00, CT09, DH09a, DB08, DLP03, Dey09, DZ06, Efr09, Eks06, EKMS04, FL03, FCCN01, FM00b, FM00c, GL05a, GG08, GKMS02, GHKTR07, GM01a, GFM05, GC09, HN09, HK01, HKSTR05, Hos07a, HC08, IS02b, IVM03, JB08, JS00, KdR01, KW09b, Kle04a, KS00, KLE^{+}04c, KNH09, Lan01, LP09, LLL09, LW07, LBCM06, LBR08, Mae05, Nav08, ND06, OS05b, OS06, Pal03, Par00a, Pel01, PW02b, RBN09, Ros00, Ruk06, Ryc01, SAC02, SC09, Sta09a, SSH08, Tak09, TT07, Tan09a, Tan09b, TS08]. order [Top01, WY07a, WJ09, WC00, XL06b, ZB09a, ZB09b, ZBD07]. Order-preserving [BM01a]. order-restricted [CP05a, PW02b]. Ordered [BL08a, BP07, BI02, Bha07, CCH08, CS02c, CFS05, CO08, Dob02, GG09b, HKMS02, HA08, KSH02, LM00a, LMH00, MT04, Net02, SS02a]. Ordering [KS07a, AAEK06, CX07, EKT06, FW07, FSS07, KK02, Kle04a, LY02, Nav08, Pal08]. orderings [FT09, GG09c]. orders [AEK09a, ALS07, GGS09a, LS07, Ma00, TV01, TV04]. ordinal [FS02b, GGS03, dE02, Con02a]. ordinary [Liu09, Tan08b]. origin [PSJ07]. Ornstein [HKW00, TL09a]. Orthogonal [BCN06, CKMS02, Geo09, LOPV06, ZF01, AK05, BM08, CMyS02, CD01, EKK06, EKL07, GKMS02, GHKTR07, Hag06, HY08, HK01, HKSTR05, Koh03, Koh08, KM00b, KV05, LMS08, MJS00, Ngu08, RSS02, ST08, SGJM02, ySD03, ySK05, yS07, WW08]. Orthogonality [Pom00, MFL03, QL06]. Orthogonally [Dey09, KT08b]. orthonormal [EOPR05]. other [Atk09, But09, Cai02, HC03, Kem02b, SF02]. outcome [BL08c]. outcome-dependent [BL08c]. outcomes [Con08, DH00, GIA04, GACS07, IBL06]. Outlier [Cho01, Bat06, BM03, FMT04, PFF00]. Outliers [MPU06, CB06, GNP07, MRCN09, SC08]. output [SSC07]. Overdispersed [dCPC05, SNM08]. overdispersion [KMB08, YHA09a, YHA09b, YHA09c]. overflown [BK02]. Overview [Mel07a, GS02, GP07c]. ozone [KRV06].

P [Con02b, UV07]. P. [Cha09]. p.g.f [Pip09]. P3 [Sen07b]. p=3 [YP07]. packet [GN03]. Packing [Yin01, ZY02]. packings [Ge01a, LT08b]. Pagan [Lin01b]. Page [vdWD01]. pages [Cha01a, Lah00, Mas01, Wes00, Ano00-28, Ano00-30, Ano00-33, Ano00-31, Ano00-36, Ano00-29, Ano00-35, Ano00x, Ano00-32, Ano00-37, Ano00z, Ano00-27, Ano00-34, Ano00y, Ano01s, Ano01r, Ano01p, Ano01u, Ano01o, Ano01n, Ano01q, Ano01t, Ano02-28, Ano02z, Ano02-31, Ano02x, Ano02-27, Ano02-29, Ano02-30, Ano02v, Ano02w, Ano02y, Ano03-51, Ano03-54, Ano03-52, Ano03-57, Ano03-56, Ano03-55, Ano03-50, Ano03-49, Ano03-53, Ano04-57, Ano04-56, Ano04-60, Ano04-61, Ano04-63, Ano04-62, Ano04-59, Ano04-53, Ano04-55, Ano04-54, Ano04-58, Ano04-52, Ano05-60, Ano05-61, Ano05-64, Ano05-63, Ano05-58, Ano05-57, Ano05-55, Ano05-59, Ano05-62, Ano05-56, Ano06-40, Ano06-41, Ano06-46, Ano06-43, Ano06-39, Ano06-47, Ano06-45, Ano06-44, Ano06-42, Ano07-40, Ano07-44, Ano07-39, Ano07-41, Ano07-42, Ano07-43, Ano08-44, Ano08-40, Ano08-42, Ano08-38, Ano08-39, Ano08-43, Ano08-37]. Pages [Ano08-41, Ano09-38, Ano09-41, Ano09-44, Ano09-46, Ano09-39, Ano09-40, Ano09-45, Ano09-43, Ano09-42]. pair [Let08, LP01, LTW08, WWL08]. pair-matched [LP01]. paired [CH05a, GJ05, GGHS04, GGS09b, JL09b, TVYS09, ZQ05, ZQ07, vBP03]. pairs [Aug02, Bie00, FN07, GMM03a, IVM03, SB04, ZG05]. Pairwise [FY05, MT07b, TZ00, ALRV01, CL09a, HL01, KA09]. Pandurang [Suk02]. Panel [DHT07, HT08, LSL08, YZ09]. panic [LT09b, LT09c]. paper [Bro06]. paperback [Dea00, Lah00, Lin01b]. papers [PS02]. paradigm [Wil03]. paradox [FMS04, TB07]. paradoxes [Gol01, Sud01]. parallel [AL09, Asa06, CDD01]. Parameter [AT02, BK09b, Sub01, Atk08a, Aue06, BW00a, BLL06, BK05a, Ber09, BP01, BK09c, CCH08, CK05, Con01, Dor09, EKK06, FM05b, FPR00b, Fre07a, GGCA00, Hay01, HZN04, Hui09, Jar07, JL04, KZ03a, Kau05, KHP06, KV07, Lan00, LP09, LY09, LBS06, LM03, MS05, Pen00, PK05, PRBAC08, RV07, RVZRZ07, RM05b, Ruk02, SZ07, SL09, SH06, SGU05, SZW08, Tom09, VG07, Wu00, YH09, ZT04, ZW03a, ZL08, ZS05, ZW03b]. parameterization [War08]. parameterized [XW07]. parameters [ASD09, BRCK03, BK09a, CO08, DR05, Dob02, DS07, Fer04, For09, GT08a, GD08b, HY03a, JPS02, KK05, Kul09, LLC02, LY02, Li01, LM08, LdC06, LM00a, MB00a, Mi06, MvdMV06, MZ07, NMM05, O’H09, Ozt05, PKM08, QRR08, QZF09, RS08, Sin06a, Sri01, SSY05, Tan01b, Tha08, TS08, Vin02, Viš03, Wan05, WWN05, WW05, Yan08, YW09b, YW04, Zha09d, vdL00]. Parametric [AKN04, AS04, BL06, BLBR00, Ber07, Bor04, Bos09, CV09, CA04, CR02, DFS08, DSG07, DL04, DR04, DM08b, DSF02, DW08, FL03, FR02, GRC09, GM04, Gri05, Ili03, IP09, KMPP02, Lia00a, MM06, PRBAC08, Pou01, Ruk07, Sch05b, Ste09, XY06, ZR05]. parametrics [Sen00]. Pardoe [Bro06, dL06a]. parent [SSE04]. Pareto [Air00, AR05, ARS01, CC07a, CCL06, FGT08, XD05]. Parsimonious [VS00]. part [Lia00a, Nam00, WZ06]. Partial [CY04a, PQ07, SN09, SCW09, Shi09b, APdR02, AM01, ACS04, BV01, BV03, CS00, CCRS07, DJ06, DM08b, GGS09b, Hor00, KK09d, LJ09, LS01, QJ00, QZF09, Rao00, Tha01, Wu07, Zha04b, ZZF08]. Partially [LZ07, Phi02, YZ09, Bar03, DL09b, DDM09, GK02, IS01, JLNY09, KvR08, Lia00a, LHS00, Lu08, Mi06, MNOP05, PKZ06, RL06, RSM05, RS08, SS02a, Sch00a, WLCI09, YXC09, YC07, YC02, Zha09d]. particle [Nek00]. particles [Tar07]. Partitionable [ZC09]. partitioned [CIPS07]. Partitioning [BBR04, CC09, FGS06, RHL04]. partitions [BDGV01]. partly [BB04c, Hua02]. parts [FFFH02]. past [Asa06, GPV09, NP06]. paste [Sul05]. path [Aga02, Ano05-40, Csá05, SAC02, Sen02]. paths [BPAK04, DK07, FP05, GV02, Kal07, KdMSY09, MSV02, SA02, Sul05, Tam02]. patient [RDS08]. patients [FS01a, GHSS02]. patterned [CA04, Cha03a]. patterns [BW09, DL00a, Kow09, SvTW00]. Paul [Cha02a]. Paule [RBV00]. PBIB [DDM09]. peaks [BJS09]. peaks-over-threshold [BJS09]. Pearson [Aug02, BGI03]. Penalized [DG09b, ACW02, Kau05, OP04]. penalties [vdG02]. penalty [KN03]. penetrance [GPC01]. pentagon [BGZZ02]. people [CC07c]. percentile [BDZ07, CLHY01, Che04a, IU01]. perception [Zad02]. perception-based [Zad02]. Perfect [ABZ00, CD07b, CT02, SC04, Ste09, Ben01, BL09, KM00a, LB08a, LK08, SS00, TLBJ11]. Performance [OKD06, BGP00, GC05, GDL07a, GM02, HSS08, JPS02, MR06, ZS00]. perimeter [Fer02a]. period [Aga02, GF00, SA02]. periodic [KdMSY09]. Periodically [Hur05, CX05, SS00, SA07, TLBJ11]. periodogram [LZ05]. Periodograms [SA07]. permanent [AM04c]. permanent-transitory [AM04c]. Permutation [JK08, BCS07, Cha02b, HJR09, JR03a, Kir07a, WC09a, Zha09b]. permutations [BK05b, HH05]. Persistence [PG07a]. persistent [Böh02]. person [DLM02]. personal [GGS02]. perspective [And06, GCR06, Gua08, JLNY07, LDC09, MRaPGVVV02, PS06]. perspectives [Sen08]. Peter [Sin06b, Roy07]. Petersburg [Ano05a]. PG [BGG00]. PH} [KKLD07]. PH}/1 [KKLD07]. pharmacokinetics [Lia03]. phase [AWL04, CLS09, FFP09, GP04, KQ02, WDE05]. phase-type [AWL04]. phases [WDE05]. Phi [GP07c, GNP07, PM08a, PP09]. phi-divergence [GNP07, PM08a]. Phi-divergences [GP07c, PP09]. photococarcinogenicity [MCH^{+}05]. physical [GR09a, LHL09]. physics [Sri07a, Var02]. Pickands [Seg05, Yun00, Yun02]. Pickands-type [Yun00]. Piecewise [CLE09, RL09]. Pilot [MW08b, WY07b]. Pilot-pivotal [MW08b]. Pitman [Fre07a]. Pivotal [HB00, MW08b]. placement [BTK08]. Planar [Nie02b]. planes [KdR01]. Plann [BD09, Bed00, DGL07, Fox09, FR08, Fre02, FM00b, IT10, Jan01, LT09b, Rao02a, Sri02a, TLBJ11, Tan09a, Ten09]. planner [MT07a]. Planning [Ani10, BL05d, EN07, EJ09, FCCN05, Gov09, LT00, LWB09, BCN04]. plans [AEK09a, CMyS02, LJ07a, MPGS09, MMM06, RVZRZ07, SST02, WS08, Wri08, Dea00]. plant [ABQG09]. play [lCV06]. plot [McL08, RX03, WH05, YLLZ06]. plots [KN02, ME07, OGACR08]. Plug [APdR02, BO04, PG07a]. Plug-in [APdR02, BO04, PG07a]. plus [CMyS02, SST02]. Point [RD08, ST05, BLZ08, Bri08, BD05, Bur01, Col01, CMS02, DL04, Din03, DW06, EMMM00, Fer01, FJT09, GJ03, GM05, Gri05, GSC07, Gua08, GV05, GS09b, HN04, MGB05, PK06, PF01, Ruk02, Sch05b, SY08b, WX08a]. point-regular [Bur01]. points [AH00, GM01c, Gua07, GR01, Heu08, Kur08, MN03, NJP05]. Poisson [Air00, AD01, CGG09, DK09, Fin03, GM05, GL02a, GL02b, GH00, HL06, HT09a, Kag01a, KX01, KM07a, KT02, Kem02c, KMB08, KS07d, KT04, Let00, MN08a, Nek00, Qua08, RdCCB09, Roo03, SZ07, SLS02, SSY05, TC08, Tso06, TP03, WMSY06, WSY06, Web07, Wei08, YHA09b, YHA09c, Zac05b, dCPC05]. Poisson-2 [YHA09b]. Polya [SAC02]. Polya-Eggenberger [SAC02]. polygonal [ITA^{+}09, IT10]. polyhedra [DG01]. Polynomial [DLP03, AAJ00, ADW03, BM06d, But01, CHLY01, CA00a, DMP04, Fan00, FW03, FWW06, Fer01, HS09, Hwa04, KSL03, KSW05, OK08, YZ09, Zha04a]. Polynomials [BCC02, AK05, KV02, KV05, Pro04]. polyominoes [Fer02a]. polytomous [GP07c]. pooled [MT07b]. pooling [ES03, RDS08]. Population [CYLC01, HYCY03, ALH07, BCP03, BNP06, BD08b, Bar08, BL02b, BL05d, Cha07, Cha01c, CBG04, Che08, CSW02, DH08, DT00, Dub07, FCM07, GS07b, GL02b, GS07d, Haw01, HC06b, Hug08, KZ03b, LL07a, LGMPS03, Mas03, ODY08, OM08, SM06, SM05b, SS08, TN05, Vel03b, WZ03, WHS07, WA07]. populations [BGI03, CC09, CFS05, FW07, GL05b, HYCY03, LL04, LL07a, Lia06b, LL05b, LLW07, LRZ08, MvdMV06, MS07b, MT04, PK01, PPZ02, Rob09b, SM07, Tar07, Wu07]. portfolio [DZ09]. pose [SGJM02]. positive [CS06, GR08a, GHL03, GL05b, HMS04, Lia00c, Lia09, Nam00, OT02]. positive-part [Nam00]. positively [XY08, Yek08]. possibly [OS05b]. Post [Lee01b, SC05a, Tsa06, Wu06]. post-change [Wu06]. post-data [Tsa06]. Post-stratification [Lee01b, SC05a]. postal [BGP00]. Posterior [KvdV08, SS03, TG07a, TG07b, Cho07b, CL09b, CJRV08, DS03, Gil05, GC04, Ken00, Maj07, Ruk02, WT08, YO05b]. posteriors [FK09]. poststratification [SS06b]. poststratified [KW09b]. potency [SR03]. potential [Sri07a]. Power [CIS00, TSS06, WW02a, ABQG09, Ant08, Che06d, DR03b, Fre01, Fre02, GM01a, GDL07b, HS05, JÜ08, Krä05, MS09b, Mey06, Shi05, War08]. powered [KBG09]. powerful [BK05a, KT04, MW02, SB02]. Powers [CSW02, Eri02]. pp [Cha02a, Gov09, Ko00, Lah00, Roy07]. PPS [WDE05]. Practical [PT08, CP03, Hos07b, Yua08]. practice [Sen01, Zas06]. practitioner [Lah00]. Pre [Oht00, ABK00, EC08, PM08a]. Pre-test [Oht00, ABK00, EC08, PM08a]. precedence [BTK08, NB05]. precise [MPS03]. Precision [dC02, Dor09, GRP02]. Precision-imprecision [dC02]. precisions [TK09]. Predictability [BCR08]. Predicting [Fue00, SS08, AdL05]. Prediction [AC03, LL08, Pot07, SW08, TB09, Bos07, Bos09, CL06, CA03, DH03, DW08, EN09, FL09, GS07a, GSY03, GBC^{+}05, Jia03a, JPS^{+}08, KZ03b, KM04, KLX09, LH00, LRZ08, Nay00, RSD06, RD08, Sán03, SK07a, SdMSL04, Vid06, Web07, WX08b, Xu09, YO09, ZMJW08]. Predictive [KHP06, AdL05, AH03, AC04b, CvdL01, CY04b, CCS07, GP04, GPW01, MW07a, NdC08, Qui06, Shi00, SS03, Vid08]. predictor [Lee05, Liu09]. predictors [GP03, MY00, WC07]. Preface [Ano00-38, Ano00-39, Ano01v, Ano01w, BM06a, BM01b, Ber02b, CD00, Des02, DC06, DRS02, EGMN09, FR06, GM01d, GDM^{+}05, GM07, GAD08, GM08, KS02a, Kol09, KK09e, Lah03a, Lah03b, Mel00, Mel05b, Mel07b, Moh02a, Moh05a, Mud07, PRS07, RHR07, SG02, Sri02d, Sri02e, Sri03b, Sri03a, Sri05a, TW04]. prefer [NvdL05]. preferences [KTB01]. preferred [CMS02]. presence [AKK09, Don04, EN02, GGS09a, GGHS07, HK03, KK01, KK05, Kat02, KB00, Kun00, MM02, QW01, Sah08, ZjLS05, ZC03]. present [BQI06, Wan08]. preservation [SA03]. preserving [BM01a]. Press [Con02b, Lin01b]. prevention [Sar06]. preventive [LSL06]. prevision [KSS01]. previsions [DM08a, MdCC05]. Prevosti [SOUS03]. Price [Lu02, Dea00, Ko00, Lah00, Lin01b, Wes00]. prices [Jia06]. pricing [LL05a]. prime [GK05, OT02]. Principal [CCL06, BPR09, Che06b, Fou00, GM01c, KK09b, Kur08]. principle [FGS06, RHL04, Sen07a, Sub03, Sub04a, Wal02, dC04]. principles [Fot00, Kir07a]. Prior [CISW03, GS08, CIS00, DJ06, Dor09, EMMM00, HNW08, KG09, KBG09, ODY08, Sam07, TSH06, TGM07, YO05b]. priority [KMSC02]. priors [AGP09, BGI03, Cho07b, CD02, CY04a, Con09, CL09b, Dan05b, Das01, DMN01, GGCA00, GK02, GH03, GPR03b, HP06b, KKL05, Lij03, LL06b, Pot07, SN04, TG07b, Ton07, WHS07, vdL00]. Probabilistic [Zhi03, Zad02]. probabilities [AAJ00, AR05, BLBR00, Che04d, Fer02b, GT02, HP00a, Har02, HM02a, KMSC02, LBM00, Nau02, ND06, Pet07, Rob09a, Ros00, YGJ05, Zad02]. Probability [Wan08, ABS00, And05, AK05, ACS04, Aug02, AC04b, CSL08, CS02b, CL06, CN07, DSG07, DT00, DGR07, DZ06, Dub07, HP00a, KS02b, KM07c, MPS09, MNOP05, ME07, Pav03, Qua08, Sag05, sS06a, WPV04, WH01, vdH05]. probability-probability [ME07]. probability-weighted [DGR07]. probable [Ber01b]. Problem [DMMS07, BW05, BBIAV00, BMP02, CL03, Das05, DK04b, EM07, FFFH02, GG01, KI08, KZ03a, Lan00, Lee01b, LZN06, LW07, MRU00, MS02a, Mun00, SB02, Spe09, Sza02, UB07, UF09, Van05, Wyn08]. problems [Bed06, Cai02, Cau05, Cav06, CN02b, CC02, Fab09, Fan04, GV05, Hai03, HA03, Kag02, LRZ08, Man02, Per00, Ral02, RSS04, Sel00, WH01, WT08]. procedure [AdHN09, ARV03, ATS02, BNP06, BB04b, BB06a, Bar08, BH05, CC07a, Che02, CKS07b, FWYY03, GP03, Hay07, Hlá03, KHC02, LP01, Li08, LLP02, LK03, LZN06, LW07, LHZ03, SW05, SY09, TSS06, WL02, Wu06, ZNWS09]. procedures [BL05a, CH01, CS04b, GACL08, GA08, GS00, Gor07, GS07c, GHL03, GL05b, GH07, HML00, HSS08, HRZ06, HLP06, HKPS08, IU01, KLE^{+}04c, LL07a, Lie06, Nak09, PS04a, Sar04, Sar08, Sch05a, SN08, SV00, Tro00, WK01]. process [Ahm05b, AD01, Bab04, Bel07, Cho07b, DK03, DK04a, Dor09, DPA08, FWYY03, GP04, GC09, GSC07, Gua08, Hai00, HS00a, KS07d, Kul09, LdC06, LS01, Moj07, NLP06, Nag09, Pom00, Raw02, RBN09, Rüs09, Sch05b, SS09, Soe02, TG07b, UV07, WW06]. processes [Ahm05b, Alv05, ANR00, ALS04, ALS07, AM04c, AIK00, Ber09, BGH09, BNW05, Bos07, Bri08, BRO^{+}09, CKS07a, CLE09, CR02, CCRS07, Dau04, Deh00b, FZL00, Fin03, Fue00, GJ03, GM05, GMM03b, GH02, Gri05, GS09b, HKW00, Har02, HT06, Hor00, HKZ08, Hyr07, IOQ09, Iva06, Kak02, KD01, KSL03, Ker04, KL05, KV06, KRM^{+}05, KM07c, LT06, MF04, MRCN09, MM06, NaX03, OL03, PP02b, PF08, PF01, Per00, Phi02, Qi07, RSM05, SL05, SZ07, SM05a, SW08, Sel00, Shi09a, Skö09, SA07, SGJM02, Ste05, Ste07, SH03, Swi02, SS02e, TL09a, WCS01, Web07, Whi05, WC08, Zac05b, ZBD07, vBHvdVW05].

Procuring [CT00]. produce [BL09]. product [FR01, FFFH02, GC07, JW09, Jup09, KK09f, LLZ07, PS04a]. product-multinomial [JW09]. product-parts [FFFH02]. product-type [FR01]. production [vBHvdVW05]. products [GMM03b]. Professor [Sat07]. profile [BGV08, FCCN04, FCCN05, FLCN05, Mar07, PS06, SSP03, SV05, WH08]. profiles [GGS09b, Har09, Lia03, SSO07]. profiling [Aus05]. prognostic [BCR08, MPS07, Wie05]. programmes [Ver02]. programming [BM08, CO08, DL07, EPS08]. progress [Ben01]. progressive [BE06, BS08, BTK08, Bha07, Bor04, Bur08, FWB08, Fer04, HCBK05, NB05, WLC08]. Progressively [BD08a, BD09, BH09a, LWB09]. projected [LZ07]. Projection [Bie08, BCGF09, GO03, GJ03, LA08, Raq04, AMY02, Das07b, KT08c, LLFF00, Mar09, PFF00, yS07]. projections [Blu09]. projective [BGG00, Gly00]. projects [RDS08]. Proof [LL07b, CH08, Haw01]. proofs [GN09, PSW00]. proper [FCCN01, Mis07]. properly [MI08a]. Properties [But01, Sha08, Azo09, BE00, BR03b, BS08, BB09, BRdCR03, BR08, BL05c, CD06, CZ09b, CS04b, Cox03, DdRB00, FS02a, For09, GFDMG04, GSC07, Hlá03, HC08, Hug08, KR03, Kem02a, LJC07, LB06, MP07c, MvdM03, Pla09, QJ00, Qu07, RSSR09, RL06, RGR09, RB07, ST06, SL09, SP05, Udd07, Wal02, WvEZ04, XD05, ZL08, ZB09b]. property [BM01a, BW07, CT09, KL01, Kag06, MP06]. proportion [Bar08, Cha01c, CC07c, HH01, TN05, Wan06]. Proportional [GG07, LY05, Zha09a, dUÁGMCS03, BO04, BB06b, BS07b, CMN06, Dup05, FPvdVvH06, Sun06b, Sun09, TR05, Xu07, YL07]. proportionality [Sub00]. proportions [Bur07, Cho07a, CCS07, JL09b, Lar03, LF07, MS00, MT07b, TR08, YLH04, ZTQ04, ZQ05, ZQ07]. proposed [PS03]. Propriety [Das01, FK09, DS03]. prospective [ZGH08]. protective [BLU07]. protocol [ÖW01]. provided [Che08, GL02b]. Pseudo [DGJ06, Hyr07, YR03, ZGH08, DG09b, LLC09, MP07c]. Pseudo-full [ZGH08]. Pseudo-likelihood [Hyr07, DG09b, LLC09]. Pseudo-Wishart [DGJ06]. pseudo-Youden [MP07c]. PT [Nam00]. PT-2SHI [Nam00]. Publications [Ano05-65]. Publisher [Ano01x, Ano03-58]. Publishing [McF01]. pure [BKLZ01]. Puri [Ano07-45]. Purposes [Sta04]. pursuit [Das07b, KT08c, PFF00]. PYD [MP07c].

Q [Smi09, UV07]. QTL [ZNWS09]. quadratic [BH07, DvdW05, FHM00, LRZ08, PC02, Tan08b, TK06, vdW04]. quadrics [Tha01]. quadruples [MJD09]. qualitative [DR03b, DLOR02, FMT05]. quality [AC07, KP09a, PY06, SSC07, dSSdL08]. quality-adjusted [dSSdL08]. quantification [Vin02]. Quantile [Hon04, LK00, Vel03b, Bed06, BFGV08, Cha03b, CTL04, CD07a, DFS08, Deh00b, DGGG08, HP00b, Ken00, NM06, OD06, QW01, SC05b]. quantile-type [DFS08]. quantiles [AMY04, Bor00, Che00b, GSY03, KPTW02, Par08, PY09, Ser04, Ver01, Wah09, Wit01, ZS08]. quantitative [GGHS07, ZG05]. quantization [OCR^{+}05]. quantizations [BV06]. quartic [FOS08]. Quasi [Kag06, LCZ06, YWS09, BGV08, BH01a, CN02a, Cha03a, CD06, GM01a, KMS09b, Lin07, Lin08, LLC09, MS02d, ND05, ST08, XL08a, Fu05]. quasi-binomial [CD06]. quasi-GCV [Fu05]. Quasi-independence [Kag06]. quasi-least [CN02a, Cha03a]. Quasi-likelihood [LCZ06, Lin07, Lin08, LLC09, MS02d, ND05]. quasi-maximum [XL08a]. Quasi-minimax [YWS09]. quasi-profile [BGV08]. quasi-score [KMS09b]. quasi-symmetric [BH01a, ST08]. questions [GGS02]. queue [ALH07, AWL04, SA02]. queueing [ALW07, Con04, KKLD07, KMSC02]. queues [Aga02, AC06, CKM00, MCM^{+}07, Per00, Sen02, SA02]. queues-lattice [Aga02]. quota [SC05a]. quotients [DHW00].

R [Dea00, Lah00, Lee01a, McF01, Sri02a]. RA [BRO^{+}09]. RA-estimates [BRO^{+}09]. radial [AKI08]. railroad [LWL^{+}09]. ramifications [BC03b]. Random [DR04, Ery03, Kem02b, Nie05, Ste04b, Ton07, ABS00, And07b, ALS07, Atk08b, BRdCR03, Ber04, Böh02, BL02a, BQI06, BL02b, CFT07, CR08, CLHY01, Che06c, CB06, Chr04, Csá02, CX05, DH08, DK07, DR01a, DGGJ06, Dik00, DGS05, EBZ08, EG00, FZ06, FMT05, dSBFG07, FR01, Fot00, FN07, Gan00, GL05a, GKK08, GTT^{+}07, GC09, GFDMG04, GLS09, GC07, HK04, HB00, HK05, HL01, HR04, KL01, Kag06, KOS05, KP02, KK02, KKLD07, Koh03, Koh08, KO03, Kör00, KK09f, KG05, KT08b, LK00, LT04, Li03, LQH03, LL08, LCZ06, LNS07, Ma00, Ma09, MW08a, MS09b, Nak00, NM06, ODY08, OTB06, PK06, PY06, Pro02, PC02, RSSR09, RM07, RD08, SZ08a, Sel00, SH06, Sør08, SdMSL04, SS08, Ste04a, Ste02, SW09, TSN07, Wan08, WD04, Wit01, WLC08]. random [XY08, Xin09, YXC09, YW09c, ZBD07, Zie03, ZEB07]. random-effects [Atk08b]. Randomization [Aic01, Kot05, KS07e, DL03, HRZ06, Ruk06, Ruk07, Wah09]. Randomization-assisted [Kot05]. Randomized [BB04b, Chr05, GL05a, PV00, WWL08, Bar08, Cha01c, CS05, lCV06, GIA04, GS07c, KW04, KW05, KTA06, NA09, Pad06, RL06, SSM00, SC09, Zha09e]. randomized-response [KTA06]. randomly [CC00, DS06, DS08, MPS06, PM01, Sun06a, Sun09, ZG05]. randomness [Lai09, RSS04, RM05c]. range [BD08c, FARMA08, IL08, Kul09, KSH02]. Rank [BL05a, CA04, KvR08, Wan09, BTK09, BI02, BP02b, EOPR05, FPvdVvH06, GG09a, HP06a, Han08b, Hur05, Jan01, JS08, KK01, KK05, MT00, Neu00, OP05, ÖW01, RV03, Sla06, TN05, VR08, VKO00, VOK^{+}03, WL01, WA07]. rank-score [BP02b]. rank-sum [BTK09, WA07]. Ranked [Che01, KPTW02, Özt02, OD06, ASAO02, AS04, ASD09, BC03b, BL08a, Che00b, CS03b, Fre07b, GT08b, LB08a, ODY08, ÖW01, Ozt05, PS04b, SM06, ZM06]. Ranked-set [Che01, OD06, BC03b, Che00b]. ranking [BH05, LB08a]. rankings [Fre07b]. Rao [Ano02-32, Ano03-59, Ans01, BB01b, CGS02, FM00a, GO01, GRU01, Hay01, HS00b, Li01, MW00a, Pal09, Pap05]. Rao-Blackwell [FM00a]. Rasch [FM05a]. Rate [HSV09, AGK00, CL06, EPdRV02, GCR06, GR08b, GG07, KY07, KNH09, KHC02, LG06, MSW09, NJP05, Nie07, Pal08, RdCCB09, SZ08a, SSY05, Sub01, Sub04b, Tro00, Yek08, YSB09, Yu06]. rate-optimal [MSW09]. Rates [GH02, Pen03, Bha07, Can06, CS04a, Gor07, JMJ09, KKW09, KvdV08, LQH03, LGL05, Par09a, Sar04, TKP06, XC08, Xue04, ZY01]. rather [BW07]. Ratio [KK09f, CK05, Cho07a, CS02c, EKT06, FW07, Fre01, Fre02, FCL08, Gar01, Gar05, GS09a, Jar07, KLA08, KT08a, LL04, Li01, LS04, LBCM06, MC08, Nav08, Obi06, Par00a, PZ08, PW02b, PW03, Rao02c, SGK09, Sør08, TW05, Viš03, WL02, WWN05, XL06b, vdH05]. rational [JS08]. ratios [DBG06, Len02, Rao02c, Sin06a]. raw [HYP02]. Ray [QRC01, HLAK09]. Rayleigh [HL01]. RCA [Aue06]. RCINAR [GC09]. re [Gov05, Gov09, LZ09b, Pal09, BP07]. re-estimation [Gov05, Gov09, LZ09b]. Re-sampling [BP07]. re-scaling [Pal09]. reaching [Len02]. Realisations [Pol09]. reanalysis [BYW07]. rearrangements [DKSZ07]. reasoning [Zad02]. reassessment [O’Q06]. recapture [CYLC01, HH07, WHS07]. receiver [LZ09a, PZ04]. Rechtschaffner [Qu07]. reciprocal [Jon08, Yu06]. Reconciliation [CS02b]. Reconciling [Wal02]. record [AJMP09, BE00, BAA08, BNW05, BS07a, Cha07, CR08, DW00, GO03, KNH09, Raq02, Raq04, Raq07, SSH08]. records [BS04, Dan05a, Hai00, LBM06, PW09, RA08, Ste04b]. recruitment [BLU07, LMD^{+}01]. rectangular [Bag04, RS04b, SSKS02]. Recurrence [AF03]. recurrent [AP05, ND05, PSG07]. Recursive [ARV03, DÖ01, Hay06, vT01, Nie02b]. Redescending [SMS08]. redistribution [ZZF09]. Reduced [FPvdVvH06, Han08b, RV03, BFGV08, DGGG08, JS08, VR08]. reduced-bias [BFGV08]. Reduced-rank [Han08b, RV03, VR08]. reducible [Wan09]. reduction [FHL03, GM04, HH09, Hei00, LP09, NN09, VdCO00, WC07, WC09b, YB06]. redundancies [HW09]. Referees [Gho02, Ano01m, Ano03a]. Reference [GPR03b, vdL00, CD02, CY04a, LL06b, Sam07]. refined [BJS09, KL06]. refinements [FCCN01]. regime [VEM07]. regimes [LC07]. region [OCR^{+}05]. Regions [JÜ08, cIC01, EHM09, JI07, LM04, MS02b, MS02a, Mus02, Shi08, Sri01, ST05, UF09, VG07]. Regression [GACS07, KS08b, ND05, Pre07, Ste02, WMES07, AW09a, AMY02, AZ04a, ASBZ04, AMY04, And06, AKI08, ADW03, AGH07, AGK00, Arc01, AVGRZ00, AGP09, Atk08b, BB03a, BLV00, Ber04, BBdR00, BB04c, BDP07, BH07, Bla00, Bon08, BWA06, BL08c, CV09, Cao09, CFT07, Cau05, Cha03b, CHLY01, Cha02c, Cha00, Che01, CD03a, CTL04, CCP09, CC04, Che04b, Che05, CC06, Che09, CLR05, Cho07b, CD03b, Cíz06, CC02, CW05, CA00b, DNR09, DdRB00, DZ00, DH08, DHP06, DW00, DCC01, Det02, DMP04, DLRP06, DM08b, DT08, DSF02, Dup09, DL09b, Efr07a, FK09, FD02, FW03, FZ06, FWW06, FZ03, Fou00, FS02b, GP06, Ger05, GRZ08, GML06, GB04, GRC09, GL02a, GGS03, GK07a, GP07c, GV05, HH01, Hag06, HK04, HC03, Han08a, Han08b, HLW07, HLAK09, HY03a, HCS04, HH09]. regression [HY03b, HPV08, Hon04, Hös08, HLP06, Hwa04, Ino03, IL08, JK08, Jin07, JS08, JPS^{+}08, KG09, KI09, KRS07, KC05, Ko00, Koh00, Koh03, Koh06, Koh08, KKW09, KP09b, KMR01, KS00, KQ02, KS03b, KN04, KS09b, Krä05, KT07, KSW05, LC01a, Lee01a, LM04, LWL^{+}09, LP01, LC01b, LT08a, LHS00, LJZ^{+}07, Lop09, Lue09, LNS07, MY04, MY00, MS01, MaL09, MM07, MTC05b, MSW04, Nam00, NPF09, NH09, NS00, OK08, Oht00, PSV06, PB08, PX09, PPdCP06, PK06, PF09, PY06, PZ08, PSJ07, Pou01, RV07, RS04c, SC04, SN09, SW05, SFT01, SYZ07, Shi09b, ST04, SLS02, SK07b, Sta04, Sto05, Sub00, Sub01, Sub04b, SY09, TRJ07, TL07, TSR03, TWH^{+}05, TSH06, TZ05, Tso06, TK06, VBF03, VV08, VR08, Vex06, VG07, VFVF02, WO00, WH02, WMSY06, WSY06]. regression [Wei08, WR02, WY06, WC09b, Wie00, WX08b, Wil07, Wu03, WC08, WH09, XF00, XC08, YSF03, YO05a, YWS09, YHA09b, YL09, YZ07, Yek08, YW04, YB06, YZ06, YC07, YR00, YZT03, ZT04, Zha04a, ZR05, Zha06, ZCL09, ZjLS05, ZY07, ZS08, Zho08, Zie03, Con02b]. regression-the [CLR05]. regression-type [Che01]. regressions [CCL03, DR04, HH08, Liu00a, SL00, WC07]. regressors [Pro02, SK07b, VBF03]. regular [AZ04b, AYZ06, Bur01, CLLZ06, CM03, CKMS02, DG01, GFM05, IS02a, Jac04, Jac09, Jan04, Kag08, LBS06, vdVD09]. regular-fractional [CM03]. Regularization [GPL09]. Regularized [YH09, AKI08, GB07, KK09b, PX09, XC08]. Reid [ZY01]. Reification [O’H09]. Reified [GR09a, GR09b, LHL09]. reinforced [MPS06]. rejection [Li08]. rejective [Gra09]. Rejoinder [ACBM09a, PW06a, SS02d]. Related [Ano07-45, ARV03, AK05, BGZZ02, BC09, Cai02, Cha02b, CGZ00, DGJ06, Gas07, Hai03, JB02, KL01, KLR07, Kle04a, KV05, Let08, LT04, Moh05b, MT02, MN03, Nie05, Obi06, PS02, RSSR09, SK06, SvTW00]. relating [LWL^{+}09]. relation [Zha01b]. Relations [CB06, CK00, AF03, JS00]. relationship [Che07, FN07]. Relationships [AMY02, Gup03, Nay00, ZN08]. Relative [JPS^{+}08, KSW05, AR05, CC08, DNN07, Jan04, JL09a, JK08, Mez03, OT02, Wil05, WA07, ZS00]. release [DK09]. Reliability [Ano09-48, RZ05, Aus05, BB03a, Che03a, DDS00, FL09, GO01, GPV09, HDM09, OL03, RM05b, SL05, UG02, Vie09, XD05, dHD^{+}09]. reliability-based [dHD^{+}09]. reliable [GL02b]. remarks [Var02]. remember [Sat07]. REML [VV05]. removal [FWYY03, LY05]. removals [WLC08]. removing [FM00b, FM00c]. renewal [Fin03, GS09b]. renewals [BBKN07, CCRS07]. Renyi [DR02, BAA08, MM07, RSM05, Son01]. repair [CX05, LSL06, SS00, SH09, TLBJ11]. repairable [LL06a]. reparametrization [DSG07]. reparametrizations [Jon06]. Repeated [AH02, BL02a, Cho08, DR01b, EOPR05, HC06a, Kus03, LR09, LC09a, PG07b, RK05, Vas03, WL01]. replacement [BN09, CS02b, CX05, FFFH02, GR08a, LBCM06, PR04, Qua08]. replicated [BF07, BCS07, Gho01, GBC^{+}05]. reported [Bar03]. reporting [KS07b]. represent [Ban03]. representable [BGI03]. representation [DL00b, Gly00, Ree00, ZS08]. Representations [WY07a, YB08]. representative [LS02]. Representing [EPS08, GP04, MP07d]. reproducing [Pre07]. require [KSS01]. Resampling [dBJM09, FFH00, HS08b, MM06, PvdL04, HKPS08]. resampling-based [PvdL04]. rescalings [Jon06]. research [PGS07, Sen08]. Residual [GK00, GNP07, Seg01, AB05, AZ07, BCR00, CX07, HKMS02, RA08, SG08, UV07, WG01, YZT03]. residuals [DGGJ07]. resolution [BF07, BR09, But05, But07, Mis02, QW05]. resolutions [DL05]. resolvability [Kag08]. Resolvable [AGGZ01, Bag04, CK08, LT96, LT00, LMM01, Zhu01, vT01]. resource [Jia06]. respect [BM05a, CC09, GS00]. Response [ACBM09a, ACBM09b, Bor09, Goo09, HS07a, IBL06, Jon09, Khu09, Luc09, Par09b, Pie09, Rob09c, Vid06, Atk09, BB04a, Bar08, BDZ07, But06, CHLY01, Cha01c, CS05, CS04a, Chr05, DNK02, DS08, DF06, DFW08, FC04, FED09, Fox09, FS02b, GGCA00, GR09b, GS07c, HOS06a, HRZ06, KW04, KW05, KTA06, LC09a, LLP02, Lue09, Mat01a, NA09, NB08, PV00, Pad06, PB08, PP09, Sam07, SSM00, SM05b, SC09, SR05, SLS02, TTC06, VdCO00, WWEL08, WWL08, Yan05, YW08, YB06, YGJ05]. Response-adaptive [IBL06, HRZ06, LC09a]. responses [CT00, CG00, DH09c, KHP06, Lop09, SBKF01, SW09, TSH06, WY06, ZH09, dE02]. responsiveness [Vin02]. restratified [WDE05]. Restricted [RBV00, BK09a, BDZ07, CCH08, Cha02b, CP05a, DF07, DLOR02, GO03, KSL03, Mey06, MML02, NMM05, OS05b, PW02b, RMS02, ZW03a]. restriction [CP07, CS02c, EKMS04, OS06, Özt02, ST09]. restrictions [Bla00, Par00a, TT09]. restrictive [BW07, CBD05]. result [AM01]. Results [But07, Páz02, ARS01, BM05b, BB01b, BBdR00, CZ09a, CMyS02, CLLZ06, Cho07b, CS02c, Csá02, DD08, ES03, FHW08, GK00, GK07b, HP00b, Heu08, KK06, MvdMV06, NP06, Sri07a, ST00, Tan07, Wu00, Zac05b]. retransformed [WZ06]. retrial [ALH07, KKLD07]. retrospectively [Dav05]. returns [TT07, TT08]. Revealing [Sch04a]. reversed [GG07, KNH09]. Reversible [Spe09]. Review [Cha01a, Cha02a, Con02a, Con02b, Dea00, Ko00, Lah00, LHL09, Lee01a, Lin01b, Lu02, Mas01, McF01, Roy07, Sin06b, Wes00, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, BJ08, CIPS07, Das05, GH07, IPS08, Nay01]. Reviewing [Aug02]. revisit [BC03a]. Revisited [Stu05, GPR03a, IPS08, KP02, PV00, TS06]. reweighting [Ars04]. rho [FN07]. ridge [DCC01, EPS08]. Riemannian [BP02a]. Riesz [HLZ05, HLZ07]. right [AP07, BD08a, BD09, CC00, CL01, CY04b, Dup09, ES02, KS07b, Lop09, LT09b, LT09c, Par00b, sS06a, SY08a, ZZF09, dCD09]. right-censored [CC00, CL01, CY04b, Lop09, SY08a, ZZF09]. right-censoring [Dup09, sS06a]. risk [Bed06, CC08, Cho07a, CMN06, FS02b, HL06, JK08, KKM09, Kle03, LTW08, MT07a, MS05, Mar09, MZ07, VdBS01]. risk-averse [MT07a]. risk-efficient [Cho07a]. Risks [Ano06b, BH08a, DH09b, DL06b, DG09a, EKMS04, EKT06, FPvdVvH06, KB00, LL06a, Mez03, TKP06, WG03a]. river [VEM07]. Robbins [MPS05]. Robbins-Monro [MPS05]. Robert [Cha02a]. Roberts [SY08b]. Robust [AW09b, AP07, AMY04, AKK09, BB04c, BO04, BGMPG09, BPR09, BJZ09, BD08c, DFG04, EKK06, Fan00, FW03, FE06, GSCG09, Ger05, GRZ08, GK02, GRV08, HPV08, HJR09, JMJ09, JR03a, KK01, KK05, KX01, KS09a, KMR01, KD07, LDC09, Mai01, MY00, MRCN09, MY08, OM08, Ozt05, QZF09, Tso06, WH08, Wie00, Wie05, WX08a, WZ08, WX08b, Xu09, ZZF08, ASBZ04, AEK09a, BSP02, BBdR00, Bon08, CTL04, CDH06, DD08, FZ03, GD08a, HY03b, Hlá03, JLNY07, JLNY09, KMPP02, LW09, LBS06, NvdL05, PS04a, PSJ07, PRBAC08, RTG07, RVZRZ07, SB05, SM01, TZ05, WR02, YvdL06, ZT04, ZM06, ZS08, ZNWS09, ZS00]. Robustness [BN03, Gov05, Gov09, PM01, AZ04a, AK03, Ber02c, GO01, LZ09b, PZY09]. ROC [PZ04, Blu09, GG09a, HHZ08, WYL09a, ZC03]. role [DHT07]. rook [Cha02b]. Room [DL00a, Ge01b]. Root [SW07, CCL03, DPR06, STP02]. Root- [SW07]. roots [And07a, DHT07, MR00, Sán03]. Rosenblatt [Ten09, LT06, LJ07b, MW05, Ten07]. Rosenblatt-type [LT06]. rotatable [AEK09b]. Rotation [BS01, KA09, Kow09, PCK07]. rotational [BZ00, Mor09]. Rougier [Cha09, LHL09, MH09, O’H09, Smi09]. rounded [BZZH09]. Rounding [HP00a]. row [CM03, CG08a, MU03, MP07f, TCD08, Udd05]. row-column [CM03, CG08a, MU03, MP07f, Udd05]. rows [HJ02, Mut05]. Roy [AG07, MHM07, Sat07, Sri07b]. Rubin [Pap05]. ruin [SF02]. rule [Bel07, CFS05, Fou00, MR06, Nam00, PM08a, PG07a, SY08b, SCZ07, WA07]. rules [EL04, GM02, HC06b, Kus03, LR09, MT07a, ZMJW08]. Run [AEK09a, AH08, BCN06, CMyS02, GL05a, GGS09a, PSJ07, RSS02, SG08, Ste00, yS05, TV01, TV04]. run-length [SG08]. runs [BTK09, But09, ED07, MPP07, Mar05, Moh05b].

S [Dee06, QKO08, Sat07, SV01a, YMJ08, HM02b]. S. [AG07, SV01a]. s2 [KM08]. s=4 [YMJ08]. Saddlepoint [YGJ05, KR07]. safe [Pal08, TSS06, ZB09a]. safety [DZ09]. safety-first [DZ09]. SAGA [KMS09a]. Sakamoto [OO08]. Samarendra [MHM07, Sri07b]. same [DNN07, DT00, Llo05]. Sampford [Gra09]. Sample [CBD05, DHV01, HM01, MCDC08, TTC06, AN00, AM00, BM05b, BM06c, Blu09, Bos06, BL05b, BD08c, BL05d, CLA09, Che00b, Che03b, DdRB00, De 04, DvdW05, DMMS07, Efr07a, ES04, ENAZ09, FHW08, GT08a, GM01b, GTV01, Gov05, Gov09, HH01, HYCY03, JMJ09, JR03a, KK01, KI08, KPTW02, KLA08, KHP06, Kon07, Krä05, KX07, KA09, LNO08, Lee05, LL07b, LN03, LW07, LBCM06, LZ09b, MS09a, MRLBBCPA00, Mun00, Özt02, OD06, PT08, PS04b, Pew05, Pol09, PG07b, RL06, SS02a, SW04, Sch07b, Shi05, SM07, Sør08, Sun03, Sun06a, TSS06, TR08, TBM04, WDE05, WHS07, War08, WW02c, YA09, ZCL09, Zha09b]. Sample-size-restrictive [CBD05]. sampled [MF04, ZG05]. samples [AFF09, BL06, BL08a, Chi03, DT00, DG02, EPdRV02, FWW09, GT08b, HC06a, HJR09, JL09b, KOS05, Kra09, MI08a, NM06, Ozt05, Pap01, PR04, Raq05, RTG07, Ryc01, SM06, SZW08, TB07, WS09, XL06b, dE02, vdWD01]. Sampling [MMM06, TBM04, Air00, AR05, ASAO02, AS04, ASD09, ABS00, BP07, BC03b, Ber05, BN09, Bla06, Bon08, BL02b, BL09, Cau05, CS02b, CP02a, CP03, CBG04, CBD05, Che01, CS03b, Chr05, CT02, DH00, DI09, DT05, Dub07, EL04, EN06, EN07, EN09, FFP09, FM00a, Fre07b, GSM00, GS07b, Gil04, Gra09, Haw01, JW09, KW09b, Kot05, KK09f, LB08a, LC00, LW07, M.04, MPGS09, MZ07, Nay03, ND06, ODY08, Ong05, ÖW01, PX09, PK01, PCK07, Qua08, Rao02a, Ros00, SC05a, SC04, SSM00, SC09, Sri02a, Sri02f, SdMSL04, Ste02, SS02c, TN05, WDE05, WS08, Wri08, ZM06, ZS05, vEKO00, SS02d]. sampling-estimation [EN09]. sampling/ranked [ODY08]. Sanford [Lee01a]. sanogenesis [BBSZ05]. SAR [Wal04]. SAS [McF01]. Sasieni [BN03]. Sasieni-type [BN03]. saturated [CKM06, CKM00, CJK04, FZ04, JM08, KM08, WW08]. Savage [EOPR05]. savings [ZP04]. scalar [LL06b, Sar06]. scale [AP07, BC02, BZ06, BK09a, BK09c, Car02, EHM09, FFP09, FG02, FMS04, GSCG09, GRZ08, Has09, HL09a, Lan00, LM00a, LW09, MS05, Mi06, MvdMV06, MZ07, NMM05, Ozt05, PSV06, Pen00, PK05, Sin06a, SZW08, Sun03, Sun06a, Ten07, Ten09, TS08, WWN05]. scale-free [FG02]. scale-like [NMM05]. Scaled [Yan05, SP05, WZ09]. scales [LCZ06]. scaling [GKM07, Pal09]. scan [Hai07, ZG08]. scatter [Ars04, BO04]. Scheduling [BD04]. scheme [ASD09, BE06, Ber05, LOPV06, Mas07, NSV02, Sri02f]. schemes [Bla06, Bur08, Kuh05, SZ06, Tan07, Wie05, XL06a, dBJM09]. Schröder [KdMSY09]. Schur [Har08]. scleroderma [BYW07]. Score [Con01, Sta09b, Ans01, BB04a, BB01b, BP02b, FCCN01, FBK03, FS02b, FY05, GO01, GRU01, KR03, KMS09b, Li01, Rao01, WG03b, Wu07, YHA09b]. scores [CLT07, Par00b, Tan09a, Tan09b]. scoring [HSS08]. scorings [SS02a]. scrambled [SM05b]. scrambling [SC09]. screening [CJK04, JG00, PS04a, Sch04b, SH05, WG05, ZZL07]. Search [Kat02, Wie02, CD07b, DR02, GGS09a, GT02, GDL07b, LZN06, Mal02, Sza02, TSN07, WC09a]. searches [Kuh05]. searching [CSL08, Lia06c]. seasonal [GA05, HF00, HT09b]. seasonality [GA05]. Second [AW09a, BLL06, BJS09, KW09b, KS00, Lan01, McF01, ALS07, BG05, Dey09, GFM05, JS00, Kle04b, LP09, LW07, Raq07, Tak09]. second-degree [Kle04b]. Second-order [AW09a, BLL06, BJS09, KW09b, KS00, Lan01, GFM05, JS00]. sectional [DLM02, vEKO00]. sections [RL09]. Secure [SvTW00]. Seemingly [VR08, DdRB00, Liu00a]. segment [DG04]. segments [CBG04, LC09b]. select [LL07a, vdH05]. selected [Bet01, MvdMV06, Vel03b]. Selecting [CN05, GL02b, PP09, BNP06, BM06b, CJK04, Dor09, DFW08, GL05b, HC06b, Jac04, QKO08, WA07]. Selection [Che08, DT00, Lie06, Ban03, BM02a, BM05a, BWA06, CV09, CH01, Che02, CISW03, CD03a, CvdL01, CO02, CG08b, DP08c, DR03b, EN02, FGS06, GP07b, GS07a, GCR06, GMM03b, GPL09, GM02, GHL03, GPW01, Hag06, Hay07, HS08a, HLP06, HK03, JL04, Kan02, Kau05, KC05, KK03, KBG09, KT07, Lee05, Len09, LT08a, MS09a, MLK09, MR06, Pav03, Qu06, RV07, RW05, SC04, SK07a, SR00, Shi09a, Sim08, ST05, STW08, TBM04, Vas03, Vid08, WWEL08, Zha09c, ZMJW08]. selection-functional [KK03]. selective [Pol09]. selector [PPX09]. Self [WCS01, AH02, HY08, JI04, ST08, TC08]. self-consistency [TC08]. self-orthogonal [ST08]. self-similar [WCS01]. Self-similarity [WCS01]. Semi [CR02, Alv05, BL06, BW01, BK09b, DW08, GM04, OL03, PRBAC08, XY06]. semi-Latin [BW01]. semi-linear [BK09b]. semi-Markov [Alv05, OL03]. Semi-parametric [CR02, BL06, DW08, GM04, PRBAC08, XY06]. semifoldover [BY09]. Semiparametric [BBW06, BL08c, Che09, FARMA08, Gua07, HT09b, IL08, KKB09b, PSG07, RZ09, Sub03, Sub04a, Sun06b, Sun09, YC07, BB04c, BR00, BB06b, CZ09c, DH09b, DHP06, Dik00, DGS05, Gil04, HKW00, IPdUÁ08, KLA08, Let08, LZ09a, LSD01, MW03, MO07, MW00b, Ots08, SSB02, SV05, Zha06, ZY07, ZGH08]. semiparametrics [Sen00]. semistable [HdGT09]. Sen [PWW08, Sri02f, ZS08]. sensitive [Cha01c, CC07c, FG06]. Sensitivity [CCH08, Li01, PMS^{+}01, AVGRZ00, Bar08, Che04a, GGS02, VIBAV06]. Sentencing [PW06a, PW06b, dL06a]. separability [Fue06, SM05a]. separable [LLC09]. separate [PSV06]. separated [MS09b]. September [Ano00y, Ano01t, Ano02y, Ano03-55, Ano04-63, Ano05-56, Ano06-42, Ano07-43, Ano09-42]. sequence [CR08, ED07, GLS09, HA08, OT02, UV07, WK01, XC08]. sequences [ALRV01, DP08b, Fot00, FL08, Hai07, HdGT09, HNW08, Hur05, Kul09, Let00, MPP07, McL08, PF02, PdSPS05, Pol03, Sha08]. sequencing [Jac06, Tat05]. Sequential [BB06b, cIC01, cICY04, Cho07a, DG02, GBC^{+}05, GD02, NLP06, PKM08, ST04, WSY06, ZEB07, BB06a, BK09a, CG00, Din03, FZL00, GS09b, HC07a, IU01, JT01, Kee05, LL05a, LL07a, Liu04, LW07, LT08c, MPS07, MBG^{+}07, Pro06, Qi07, RW06, RGR09, ST05, Ste05, Whi05, ZH09]. Serial [Wei08, BdM09, DHT07, GGS09a, HP06a, KZ09, MT00, RA06]. series [APR02, AZL00, AH01, Aue06, Bat06, BCVV09, BG09, CCL03, Cho01, DLM02, DNN07, EN06, EN07, Fab09, GD08a, GP07a, GP07b, GSCG09, Gho01, GO07, GP03, HPY02, HF00, HT09b, HP06b, HPS07, HKPS08, KSS08, Koh03, Koh08, LX07a, LOPV06, MP07d, Muk06, NW05, OM07, PVHK09, PR06, Ros07, RA06, ST04, SSKS02, Tan01a, TMP06, Vel03a, Zha09a]. served [Aga02]. server [KMSC02, KRM^{+}05]. servers [ALW07]. service [AC06, BGP00]. Set [Tse04, ASAO02, AS04, ASD09, BC03b, BL08a, Che00b, Che01, CS03b, DDM09, ENAZ09, Fre07b, GT08b, HK04, Jac02, Jac03, KPTW02, LB08a, ODY08, ÖW01, Özt02, Ozt05, OD06, PS04b, Pou01, SM06, TN05, ZM06].

Set-induced [Tse04]. set-ups [DDM09]. SETAR [GP07b]. sets [AM01, BF07, BKLZ01, Bie00, BM01c, Cao01, DBG06, DH03, FRF09, FMN09, FHMS02, FH05, GH03, Kat02, KS02b, KS01, RHL04, Ser09, Wan05, WW06, WMK07, ZW03a, Zha09d, ZC09]. setting [RMS02]. settings [MR07, RM05a, Sta09a]. Setup [LdC06, BB03a, Sla06]. seven [KDT06]. Several [PLLS08, AC07, BL02b, CS02d, DT00, DPA08, FW07, HML00, LL05b, LLW07, MMM07, MS02e, TKP06, Zha07]. severity [BJZ09]. Shanti [Dee06, Pan06]. Shao [Mas01]. Shape [Bir09, BC02, DGJ06, EPS08, Mey06, MDM06]. shape-restricted [Mey06]. shaped [BLZ08, CS02c, KTK08]. Shapiro [BC03a]. shared [ZGH08]. sharing [DL00b, YJS09]. Sharp [KOS05, LBR08, Pol03, Raq07, PK06, PSJ07]. sharpening [YL09]. sharpness [FR01]. sheet [CGPP06]. Sheffer [Pom00]. Shen [Ano01y]. shift [Maj07, Shi00]. shifts [Sri02f]. Shiryayev [SY08b]. shock [Fin03, Let08]. short [Pol03]. Shortening [AL08]. Shortest [MPP07]. should} [Ber01a, Lin01a, Mat01b, Sas01]. showing [BT04]. Shrikhande [SV01a]. shrink [YA09]. shrinkage [Dan05b, Efr07b, FMS04, Fu05, Hag06, SSE04, vdL00]. sib [ZG05]. sided [Cai05, CSW02, HML00, HK03, KKM01, NPF09, PW02a, PW02b]. sieve [APR02]. sieves [vdG02]. sigmoidal [DP08a]. Sign [VKO00, Kas00, KPTW02, SB02]. signal [Che02, DFG04, HL06]. signals [BRCK03, Kun00, NK06]. Signed [AA08, HP06a, OP05, ÖW01]. signed-rank [OP05]. Significance [Rei05, PG07b]. significant [Mal02]. Silver [Sri02b]. Silvey [BB01b]. Simes [BSW08, CS06]. similar [WCS01]. similarity [FCM07, WCS01]. Simple [BW00a, Fin03, ASBZ04, AH02, AB09, BX07b, BX07a, BH08a, BKTR08, BD04, BH01b, CCW05, CW07, Chu01, DH08, DB01, FL09, FW07, Goo08, Hai03, HLW07, HY08, KP02, Len09, Let00, LB08a, Liu04, MaL09, NM06, NPF09, SM06, SdMSL04, TL07, ZCS09]. simplex [DHW00]. simulated [KMS08, MMM07]. Simulating [EG00]. Simulation [ES00, MBG^{+}07, AM00, ALH07, CKM00, Llo05, PS03, RWF04, Wes00]. Simulation-based [ES00, MBG^{+}07]. simulations [GP04, Nak00]. Simultaneous [Ahm02, BM02a, BM05a, DBG06, EHM09, HCW06, Lia06b, LMH00, SSW09, Sin06a, BDH05, FSS07, FS02b, LC09b, SN08]. sinc [BH07]. since [Dob02]. SINful [DP08c]. Singer [BE01]. Single [JI07, Lop09, AF03, DHP06, KMSC02, KRM^{+}05, Nak09, PT08, WR02]. Single-index [Lop09, DHP06]. single-stage [Nak09]. Singular [DGGJ09, FC03, Lan00, Liu00b, WC01]. sinusoidal [PKM08]. situations [BS01, Das01]. six [ABZ00, KDT06]. Size [Ong05, ABZ00, AGGZ01, Blu09, BD08c, CLA09, CYLC01, CBD05, CCW05, De 04, DHV01, Efr07a, FHW08, FKL^{+}02, Gov05, Gov09, HYCY03, Hug08, ITA^{+}09, IT10, JMJ09, KM01, KL07, LR02, LZ09b, MRLBBCPA00, PG07b, Rob09a, Shi05, Sri01, TSS06, TTC06, WRV05, YA09, Yin00, ZC07, ZC09, ZCS09, ZN08]. Size-biased [Ong05]. sized [JT01, Yin01]. SiZer [MdUÁ04, PVHK09]. sizes [AM00, BM02b, Bos06, BCN06, Che03b, SC05a]. skew [DN03, EOGQ09, Fan05, GFDMG04, JB08, JB09, KM04, KP04, LDC09, LL06b, Mei07a, MH00, Sar06, VIBAV06, WG06, ZH08]. skew- [JB08, JB09, ZH08]. skew-elliptical [EOGQ09]. skew-normal [JB08, JB09, LDC09, LL06b]. skew-slash [WG06]. skewed [IOQ09, TZ00]. skewness [FM00b, FM00c, Jon04, MML02]. skewness-kurtosis [MML02]. Sklar [Rüs09]. slash [AG09, WG06]. Sliced [Lue09, SCW09, SY09, YZZ09]. slippage [KK05]. Small [AEK09b, Mus02, Sør08, WW02c, AFF09, BCN06, BL03, CW07, DS03, Efr07a, EN09, Fer02b, FR00, FR08, GT08a, HC06a, Jia03a, Kat02, Lar03, Mai01, Mun00, PT08, WHS07, YR03]. small-area [BL03, EN09, Jia03a, Lar03]. Smallest [Wan06]. Smirnov [Mei07a]. Smith [Sri02a, Smi09, Tho02]. Smooth [CS02a, CX07, SC05b, BCC02, CKS07b, FGT08, Kra09, MG08, PB08]. Smoothed [Alv05, DR08a, ENG00, GM05, LSD01]. smoothers [Hwa04, JPS^{+}08, Zha04a]. Smoothing [PKZ06, But01, CR02, Gri05, HH09, Kau05, LC01a, Mez03, PZ04, RV07, RZ09, Sch00a, vdL00]. smoothness [She02]. Sobolev [Efr09]. social [KS07c]. sojourn [dSSdL08]. SOLS [XP01]. solution [DK04b, SC05a]. solutions [DMMS07, Ral02, WZ01a, WZ01b, Wyn08]. solve [KRM^{+}05]. solving [DEH00a]. Some [ACW02, AT09, ARS01, BR03b, BM05b, BB09, BBdR00, Bul07a, CZ09a, CKM06, CLLZ06, CIPS07, CL01, Csá02, DdRB00, DD08, Duc00, EMMM00, FL03, For09, GM01c, GT06, GH01, HP00b, Hos07b, HMS04, HC06b, Hug08, Jac03, JK03, KR03, KKK03, KK06, LJC07, LB08a, Liu04, LJZ^{+}07, MRU00, Mas07, MSV02, Moh05b, NP06, Ngu08, RSSR09, Sch01, SSM00, SSKS02, ST00, yS07, SP05, Tha08, Var02, Wu00, Yua08, Zac05b, Zha00, Bat04, BB01b, BSP02, BC07, CD06, Cha01a, CSW02, DS06, DS08, DN03, DR01b, FS02a, Gan00, GK05, GS07c, GK00, Gup03, GK07b, HLP06, HC07b, Hui09, IL06, Jon04, KDT06, KN02, LFTT08, Mar07, MvdMV06, MS07b, MP05, Orr08, PM08a, Qin03, Rao00, RSS04, SS02b, Ste04a, Tan01a, TSH06, Udd05, Udd07, UG02, YMJ08, Zog08, ZS00]. sources [CBC^{+}07]. space [BGG00, DNR09, KD01, OGACR08, SL09, Tan01b, Zho08]. space-time [SL09]. spaces [BK05a, Bos07, Bur01, Gly00, Lin08, Pre07, ZW03a]. spacing [BE00, RJT05]. Spacings [BE06, AN00, Eks06, Eks08, GJ01, HZ06, JB02, LB00, MvdM03, XL06b]. sparse [CD03b, Lie06, RZ09]. Spatial [ZS05, AC03, Ans01, Azo09, BGP00, Cha01a, CA03, Con08, DPA08, FARMA08, Fue06, Gen01, Gua08, KS07d, Krä05, LLC02, Lai09, LLC09, LZ05, M.04, MPU06, PF08, Rob07, Roy02, Sch05b, Ser04, SS09, Wal04, ZS08]. spatial-temporal [Fue06, LLC09, PF08, Sch05b]. Spatially [PGLS08, BF07, Lin08, MEC04]. Spatio [MGB05, Gri05, Ma03, XW07]. Spatio-temporal [MGB05, Gri05, Ma03, XW07]. SPDE [ZL08]. Spearman [Che07, FN07, GN09, LL07b, vdWD01]. Special [Ano07-45, Ano09-47, HKP08, AK03, Ano09-48, YW09b]. species [Che04a]. specific [HT08, TKP06, WBvdL07]. Specification [MS03, FL03]. specifications [ACS04, ES03]. Spectra [Ioa07, Swi02]. Spectral [Kak02, LX07a, Ton07, WYL09b]. spectrum [BKTR08, Rob07]. speeds [Fie04]. sphere [ZF01]. spherical [Kle03, Sei01, WW02a, ZF01]. spherically [Kur08]. sphericity [MC08]. spin [GH01]. Spline [Sel00, She02, ACW02, But01, Kau05, PKZ06, YZ09, Zha04a]. splines [HP00b, ND05, Sch00a, SH02, SH03, Sto05, YZZ09]. Split [RX03, Goo08, HM02b, McL08, OJKM02, WH05, YLLZ06, Zha09b]. split-block [HM02b, OJKM02, WH05]. Split-plot [RX03, McL08, WH05, YLLZ06]. spread [DH06]. Springer [Cha01a, Con02a]. square [Bet01, CP02a, CP05a, DL00a, GP06, JK03, Oga09, PX09, SGK09, SV01b]. squared [AAJ00, BH02, GKK04, HL06, LY02, LY09, WO00, Wil05, ZK09b, dUÁGMCS03]. squares [AM04a, AW09a, BW01, BSP02, CN02a, Cha03a, CL09a, Cíz06, CD01, DHW00, EC08, GB07, GDL07b, Ino03, KMV05, LLC02, Liu09, Páz02, Sri05b, Sri07a, Wie00, YC07]. SSA [GO07]. St [Ano05a, YMJ08]. stability [Aue06, KHL07, Mar09, Mei05, Nag09]. stable [CT04, HL09a, RSSR09, SS09]. stage [ATS02, CP05b, EN09, FBK03, HS02, Hlá03, KM07b, KW09b, LP06, LT08c, MZ07, Nak09, Sar08, Wah09, WW05, XW08]. Stam [Kag01a]. standard [CC09, Can06, CN05, CGPP06, Jac02, Jac03, Kwo01, LL09, SSE04]. Standardized [YO05b, LFRL07]. Stanley [Ko00]. Star [KTK08, CS02c]. Star-shaped [KTK08, CS02c]. Starters [CGZ00]. Stat [Gov09]. state [AS04, DLM02, FL08, KD01, Kem05, Nak00, Par00a, Tan01b, VEM07]. state-space [KD01, Tan01b]. States [PW06a, PW06b, dL06a]. stationarity [Ste07]. Stationary [Ker04, Let00, MW07a, ALRV01, Ber09, Cho01, Fot00, Hai07, HdGT09, HT06, Kak06, LT06, Ma03, MM06, Nag09, NK06, PF02, SA07, TRJ07, WW06, XY08, Xin09]. Statist [BD09, Bed00, DGL07, Fox09, FR08, Fre02, FM00b, IT10, Jan01, LT09b, Rao02a, Sri02a, TLBJ11, Tan09a, Ten09]. statistic [AC07, Ani10, BS05a, BBR04, Fan05, FG02, GH02, Hay01, KSH02, LJC07, LJ07b, MC08, MA08, Moh05b, Oga09, War08, XP02, vdWD01]. Statistical [Ahm05a, AFF09, Ani10, AVGQ05, BCN04, BZZH09, BZ06, CDB^{+}09, De 04, EN07, EJ09, FCCN05, HSS08, KB08, LSL08, LT00, LTW08, Qu07, Rao02b, RWF04, RS07b, SSO07, Sen08, SSC07, SH03, TT08, TMP06, ZY07, ANN08, AM00, BK09a, Cox03, Fre01, Fre02, GP04, Goo08, HKP08, Hay06, Jan01, Jan04, LLP09, Pew05, SS02b, Sen07a, TS06, Zas06]. Statistics [Ano07-45, AF03, Ahs00, AN00, AH03, Ano09a, BE00, BCP03, BS04, BTK08, BD08a, BD09, BTK09, BES05, Bar07, BAA08, BRdCR03, BN09, Bet01, BK09a, BR00, Bie08, BdM09, BW08, Bra06, Cha07, Che07, CB06, Chr04, CK00, CT09, CMS02, Csá05, DB08, DLP03, DR01a, DvdW05, EOPR05, EMMM00, Fuj02, FM00b, FM00c, GO03, GG08, Gen01, GM01a, GR08a, GNP07, Hai07, HP06a, HBK03, HN09, HH05, Hos07a, HC08, IL06, JB08, JB09, KL06, KOS05, Kha03, Kis02, Kle04a, KNH09, LBCM06, LBR08, Mar07, MT00, McF01, MSV02, Mi08b, MM07, MR01, Nav08, OM08, PR04, Par08, PR08a, RS04a, Raq02, RS04b, RSM05, Ruk06, Ryc01, SZ07, SW04, SR02, Sla06, Ste00, SSH08, Tan09a, Tan09b, TS08, WL01]. statistics [Wil03, Yek08, ZG08, ZB09b, vdW04, Mas01]. status [BQI06, DLS09]. Steady [AS04, Kem05, Nak00, Par00a]. Steady-state [AS04, Kem05, Nak00, Par00a]. Stein [Cha01a, BDHS06, Das07b, FPR00b, GKK04, Nam00, PM08a, Ral02]. Stein-rule [Nam00, PM08a]. Stein-type-like [FPR00b]. Steiner [BGZZ02, Col01, DS02b, LMM01, OK00]. step [BX07b, BX07a, BH08a, BH09a, CKS07b, DHV01, FWB08, FL09, Gor07, KKB09a, Kwo01, Li08, LS01, Sar00, Sch05a, TTC06, WR02, WW08, WLC08, ZH08, vBHvdVW05]. step-down [DHV01, Gor07, Sch05a]. step-stress [BX07b, BX07a, BH08a, BH09a, FWB08, KKB09a, LS01, WLC08]. step-up [CKS07b, DHV01, Kwo01, Sar00, WW08]. steps [WR02]. stepup [Sar08]. Stepwise [Tro00, CS04a, Sar04, TSS06]. still [Bat04]. Stochastic [GS07a, GK06a, HZ06, MGG09, ZB09b, ALS04, AIK00, BI02, DH09a, FW07, FSS07, FT09, GP06, GG09c, HS00a, HR04, Iva06, Jia06, KK06, KA08, KV07, LY02, LS07, LLL09, LC07, MvdM03, MPS09, Rao00, SV00, Sla06, Tri08, VdBS01]. stochastically [NH09]. Stopping [Per00, Bel07, LL05a, Zac05b]. Storer [Cao01]. Strassen [Aug02]. strategies [CC02, DL03, FS01a, GSM00, PK01, RW06, SSM00, Sta04]. strategy [DPR06, FFP09]. stratification [Lee01b, SC05a]. stratified [Chr05, DL09b, KW04, KK09f, LTW08, ODY08, TB07]. Strawderman [BK09c]. Strawderman-type [BK09c]. strength [BB03a, BCN06, CKMS02, DL00b, MRC03, Ngu08, ySK05]. stress [BX07b, BX07a, BH08a, BH09a, BB03a, DK09, FWB08, FL09, KKB09a, LS01, WJ08, WLC08]. stress-release [DK09]. stress-strength [BB03a]. strictly [Hay07, XY08, Xin09]. Strong [FR02, GB07, Heu08, CC07b, Che04b, Das05, DM08a, Deh00b, Dik00, KL05, Mar05, SC05b]. Strongly [LS02, BJ06, XC08]. Structural [FS02a, CK05, Con08, Kan02, NvdL07, SN03, SK08, YvdL06]. structurally [Udd05]. structure [BGH09, BL05d, CvdL01, Dan05b, DH03, ES04, GC09, HP02, KL01, KZ09, LR09, Ma09, MNOP05, MS03, OM07, SA03, Sch04a, Wal04, WM02]. Structured [HP06b, Con09, FK09, KMV05]. structures [CL09a, DLP03, EZ09, Nie05, Oga09, SK06, SvTW00, SH05, YA09, YW09b]. Studies [Che00a, Bar03, BB01a, BL02a, CC08, Cho08, DH09a, DDS00, EKMS04, FBK03, GPC01, GT06, HW08, LP01, LTW08, LHOP01, LuV09, OM07, RO05b, RS07b, SW01, TTC06, YHZ05]. Study [AC07, ALH07, Ban03, DCC01, DL09b, FJT09, GC07, Huz05, KTB01, Lai09, LL06a, LWL^{+}09, LS06, O’Q06, Par00b, PS03, QT07, SS06b, TSR03, WY07b, WWL08]. study/external [TSR03]. studying [Qiu01]. Sturm [Lan00]. Sub [KLR07, KS09b, LCM07, MF04]. Sub- [KLR07]. sub-Gaussian [LCM07]. sub-hypothesis [KS09b]. sub-sampled [MF04]. subgrouping [GC05]. Subject [GC05, BL08c, Hos07a, SSY05, YL07]. Subject-treatment [GC05]. subjects [GF00]. subplanes [Zhu01]. subpopulation [SC05a]. subsequence [FL08]. subsequences [DP08b]. subsequent [Zha09e]. subset [CH01, GM02, Hay07, MR06, PT08]. subsets [AC04a, Sch05a]. subspace [SCW09]. subspaces [MRaPGVVV02]. substitution [DDvdL08]. Success [ED07, MPP07]. successive [CC00, LMH00, LK03, Sin06a]. Sufficiency [Kag02, BV06, IP09]. Sufficient [cIWsL05, XR09, BLV00, GKMS02, KY07, MNOP05, SC05b, WC07]. Sugden [Sri02a, Tho02]. suggestions [KN02]. sulfamethoxazole [FS01a]. Sum [Eks06, Eri02, BTK09, CS00, GDL07b, NPA06, PKM08, WA07]. Sum-functions [Eks06]. summaries [NN09, Pew05]. summary [Gas07, WL01]. summation [ST09]. sums [CCRS07, Hor00, HL01, Mar05]. sup [Ber04]. sup-norm [Ber04]. Super [BH01b, CCW05, CW07, ZCS09, AdHN09, CJK04]. super-heavy [AdHN09]. Super-simple [BH01b, CCW05, CW07, ZCS09]. superadditivity [KLR07]. superimposed [BRCK03, DR02]. superiority [KS03b, Kwo01, Sin03]. superiority/equivalence [Kwo01]. superposition [EPS08]. Supersaturated [CSL08, AG04, AFH07, Bul07b, BR08, But09, CG03, CL08, DDCC08, FLM00, GKM06, JLN08, JLNY09, KMS08, KMS09a, LZ00, LRD07, LLZ07, LM00b, LHZ03, PPX09, RB07, TAGF07, YM02, ZZL07]. supplement [ZQ07]. support [ABS00]. sure [ALRV01]. surface [ACBM09a, ACBM09b, Atk09, Bor09, But06, FED09, Fox09, Goo09, Jon09, Khu09, LZ09a, Luc09, NB08, Par09b, Pie09, Rob09c, Sam07]. surprise [BM03]. surrogate [Fre01, Fre02, MBA^{+}08]. surrogates [WMES07]. survey [BL05c, BL05d, Cha01c, CP02a, CD01, Dob02, EN06, EN07, GSM00, GGS02, KD08, KP09b, Kot05, Lar03, Nay03, PF09, Rao02a, Sri02a, SS02c, SS02d, Tha01, Ver02, YR03]. surveys [BR02, CS05, NA09, Lu02]. Survival [AL09, Sub04b, vEKO00, BD08b, BMW09, CS02a, CX07, CC00, DMN01, GHMR06, GMM03b, KI09, Ko00, Kot06, Kra09, LSD01, Neu00, RdCCB09, SL05, sS06a, SY08a, Sub01, TR05, dSSdL08, Wah09, Xu07, Yao08, ZRM06, ZGH08]. Survival-rate [Sub04b]. survivor [BBSZ05]. Swarup [Pan06]. switching [EZ09, HS07a, LC07, Spe09, VEM07]. Symmetric [CTL04, Top01, BH09b, Ber07, BH01a, CT04, GM01c, HLZ05, Kur08, Mei05, MS02c, Pav01, ST08, YLS00]. symmetrical [NS00, QL06]. Symmetry [SK02, Bos06, Dat07, EJ07, EJ09, Jon08, KP04, LZ05, MW07b, Özt02, PF08, RS04c, SM05a, SS02e, YZ07, ZS00]. Synchronized [BCS07]. synthesized [FS01a]. synthetic [Rei05]. system [AL09, Asa06, ALW07, CX05, DNN07, DL00b, HDM09, KKLD07, KMSC02, Muk06, Pom00, SS00, TLBJ11, Tom08, McF01]. systematic [BM06c, Kas00]. systematical [LHZ03]. systems [BKLZ01, BGZZ02, BZ00, CL03, Che03a, Col01, CDB^{+}09, DEH00a, DS02b, DS07, GR09a, HW09, KS07a, KRM^{+}05, LMM01, LL06a, LR02, LK08, Mis02, MRC02, Nav08, Nek00, Pal08, PGS07, She01, TS01, YJS09, ZB09a, LHL09].

T [Sri02a, YMJ08]. T3 [RWF04]. table [Beh08, Che04d, Dav04]. tables [AC02, AG07, Bor00, Cox03, DTW06, FS02a, KL07, KB08, RZ09, SJ08]. Tabu [WC09a]. Tadepalli [Moh02b]. Tail [Hod04, BBW06, BFGV08, BJPZ08, DGGG08, DdHL06, GG08, GM04, GFM05, Hil09, Len02, LP04, MP07b, Nec06, PQ07, Pet07, Ros07]. tail-coefficient [GG08]. tail-distributions [DGGG08]. Tail-thickness [Hod04]. tailbounds [Fer01]. tailed [BJS09, GN03, KP08b]. tails [AM04a, AdHN09, DBDG09, Fan04]. taking [SG08]. tale [OO08]. Tamhane [Kwo01]. tampered [LS01]. target [MI08a, PG07b]. targeting [CF04]. Tartu [Ano09a]. tau [FN07]. taxonomy [Sin03]. Taylor [MP07d]. Tchebyshev [IVM03]. technical [GPV09]. Technique [CC07c, SM05b, VdCO00]. techniques [BY09, CN02b, CS05, PFF00]. teletraffic [MMM07]. Temporal [FCS05, Fue06, Gri05, LLC09, Ma03, MGB05, PF08, Sch05b, XW07]. temporally [LH00]. term [BCGF09, DH08, HW08, LMD^{+}01, Sub01]. terminated [WJ08]. terminating [Fin03]. terms [ACS08, Cua02, Hod04, JB09, Qin03, YO09]. Ternary [KdR01, KNTR01, HH02]. Terry [CH05a]. Test [CQ08, GM01a, Mei07b, QSL09, ABK00, AdHN09, AZL00, Ani10, Ans01, BC03a, BNP06, BTK08, Bat04, BK05a, BLPR05, Boc07, BDH05, Bra06, CK05, Det02, DM08b, Din03, DW06, EJ07, EJ09, EC08, FGT08, FL03, FL09, FW07, FCCN04, FCCN05, FCM07, FCL08, Gar01, Gar05, GP04, GO01, GRU01, GRP02, GH07, HP02, IL04, Jac02, Jac03, Jar07, JPS02, KL07, KPTW02, Kör00, KT04, KLE^{+}04c, KA09, LL09, Li01, LB08a, LJ07b, LS04, Liu04, MPS03, MC08, MCDC08, Mar07, Mei07a, MN08a, Mi08b, MA08, ML03, MT02, MT04, MW05, MR01, Mun00, NPF09, Oht00, OP00, OP04, ÖW01, PM08a, Par00a, PZ08, PW02b, Pou01, PG07b, RS04a, Raq05, RL06, RSM05, RSS04, Sar00, Sch05a, SL09, Sør08, Sta09b, SNM08, SS02e, TSS06, TB07, Tro00]. test [WH01, WL02, WRV05, WG05, WM08, Wu07, WW08, WLC08, YP07, YS08, YHA09a, YHA09b, Yek08, Yin09, vBLO04]. Testing [AA06, Aue06, BD08b, BCR00, BGH09, Bha07, BV03, CS08, CC04, CDH06, DSF02, ENAZ09, Fue06, GKDN06, HS00a, HH05, HL09b, KS03b, KS09b, KP04, Lai09, LLFF00, LZ05, MTC05b, NJP05, Nag09, Net02, PPdCP06, PF08, SVO05, SM05a, Sch07b, Shi08, YHA09c, AP05, Ahm01, AM04b, Aug02, BP07, BCN04, BH09a, BB04b, BLSvdDS05, BKZ09, BK05a, BI02, Bet01, BM06e, BWA06, cICY04, CS04a, Che04d, CS04b, CSX08, CO02, DK09, Das01, DL04, De 04, DF07, DPR06, DdHL06, EM07, EMMM00, FWB08, GS02, GH03, Gor07, HA08, JW09, KY07, KKM01, KSH02, Li08, Lia06a, LS01, MP08, MT07b, MM07, MML02, NPA06, NSV02, PR06, PSJ07, PvdL04, RRS06, RM05c, RA06, SB02, SN03, SB06, Sen07b, SR00, SVZ07, SvTW00, Tat05, VM04, Vex06, Viš03]. testing [WH05, WK01, WGCI09, YSH03, YLH04, ZP04, ZP05, Zhi03, ZZF08, dBJM09, Ano09-48, Jan01]. Tests [AGH07, BLPR05, CN07, GD08b, MCH^{+}05, SSB02, SLS02, ABK00, AKK09, Bak09, BCS07, Bed06, BB01b, BI02, BC09, BPR09, BP02b, BT06, CV09, CA04, CSW02, CD07b, CS02c, CKS07b, Con01, DZ00, DH09a, DHT07, DR04, DD08, DHV01, DN03, DPA08, EKT06, ES00, Fan03, FCCN01, For09, Fre09, FS02b, Fuj02, GS09a, Gil04, Gua07, GH00, Har09, HS09, HJR09, HS00c, IVM03, Jan01, Jan04, JÜ08, JT01, JR03a, Jup09, JS00, KK01, KK05, KI08, Kee05, KP08b, Krä05, Kra09, LNO08, LY02, LG06, Lia00b, Lia00c, LWB09, Llo05, MW02, Mei05, Mei09, Mey06, MN08b, MFSW07, MR00, NMB02, Neu00, NHZ04, NB05, NdC08, OKD06, OP05, Pal03, Pen03, PW02a, PKZ06, Rao01, RTG07, Rei05, RS07a, SB02, STP02, SB05, Sán03, Sch01]. tests [Shi05, Tan01a, Ten07, Ten09, TKP06, WM02, WZ03, WJ08, WL01, WMK07, WG03b, YHZ05, Zha04a, Zha09b, ZC03, dE02, FS02a]. th [BS07a, CD07a, Raq04, Tan09a, Tan09b]. Theil [PWW08, ZS08]. their [AK03, BS07a, Che00b, FHMS02, GH02, Gup03, HN09, Hos06b, Hos07a, KTK08, KNTR01, LX07a, Qin03, Sar04, Udd07, XD05, Yan05, dBJM09]. theme [ACS07]. Theodore [LOV08]. theorem [AAJ00, Atk08b, Das07a, DH09b, DGS05, Har08, HT06, KRS07, LCM07, MT00, OO08, Qiu01, Rüs09, Sen00, TS06, TL09b, Xin01, Stu05]. theorems [FZL00, GM05, GHH00, Hil09, HS00c, KV02, MW00b, MS09b, Saj00, Zhi03]. theoretic [CN02b, FG06]. Theoretical [HS05, O’Q06, Ant08, VBF03]. Theory [AZ04b, Ano06b, FK09, Nie07, PS04a, dUÁGMCS03, ACW02, Aug02, BC03b, Bar07, BGL01, Cha01a, CLR05, DGJ06, FL03, FTY04, Gar05, Hos07b, KK03, MW07b, Oga09, RM05c, Sen01, TBM04, Tro00, Zad02, ZGH08]. there [GGS09b]. thickness [AM04a, Hod04]. third [ALS07]. Thompson [Wes00]. Three [FR07, GHKTR07, TGM07, Whi05, ZH08, AZ04a, BW00a, CL09a, CKMS02, Dey09, FKL^{+}02, GKMS02, Hlá03, Jar07, KKM09, KSW05, LM08, LZ09a, MS02c, MP07f, PM08b, SGU05, Udd08, XL06a, ZC07, vdV08]. Three-level [TGM07, Dey09, XL06a]. three-parameter [BW00a, SGU05]. three-stage [Hlá03]. Three-step [ZH08]. Threshold [Che04c, BJS09, DL00b, Ery03, LWL^{+}09, Sri01]. thresholded [LX07b]. thresholding [Chi03, CJRV08]. throughput [SSO07]. Tides [Wil03]. tied [BBR04, KK09d]. ties [GLS09, vdWD01]. Tight [KM00a]. tilting [KT02]. Time [BG09, AH08, APR02, AZL00, AEK09a, AL09, ALH07, Aue06, Bat06, Bel07, BLPR05, BCVV09, CCL03, Cho01, Con09, Con04, DLM02, DLS09, DHT07, EN06, EN07, EZ09, Fab09, GD08a, GP07a, GP07b, GS07a, GKDN06, GSCG09, Gho01, GO07, GHH00, GP03, HPY02, HF00, HT08, HT09b, HP06b, Hui09, HPS07, HKPS08, HH07, Iva06, KSL03, KI09, KSS08, Ko00, LX07a, Leh09, LR09, LC07, Min07, NW05, PVHK09, PR06, QW05, Raq05, RA06, SL05, SL09, SL00, ST04, Sun06b, Tan01a, TMP06, Vel03a, Yao08, ZZF09, Zha09a, ZjLS05, ZP04]. time-covariate [KI09]. time-dependent [EZ09, Sun06b]. time-series [Fab09, KSS08]. time-specific [HT08]. times [BS07a, BD04, CX05, Let08, MS09b, TR05, dSSdL08, Zac05b]. titration [GHSS02, LHOP01]. Tolerance [LM04, Cho08, Mus02]. Tonchev [Ano01z]. tool [Gen01]. Tools [HP07]. top [Tan09a, Tan09b]. topic [Ano09-48]. Topics [Ano07-45]. Topological [FCW00]. torus [TSN07]. total [AAJ00, AYH09, BLPR05, Bur07, KMV05, Raq05]. totals [BL05d, DTW06]. touching [Tam02]. tournaments [Gli08]. toxicity [DF06, DFW08, Sta06]. trace [LLP09]. Trade [FG06]. Trade-off [FG06]. trades [BKTR08, KNTR01]. trait [ZCL09, ZG05]. transform [BB09, BLPR05, LLL09, Rüs09]. transformation [CZ09c, CGPP06, FWB08, KA06, MW03, Mei09, Sub03, Sub04a]. transformations [FM00b, FM00c]. Transformed [CdA09, CL06, GS07d, LLLH05]. Transient [KMSC02, KM07c, Sen02, SA02]. transition [MW07a, TG07a]. transitive [CS03a]. transitory [AM04c]. translations [MMP00]. trapping [LY05]. trapping-removal [LY05]. Treatment [Man02, BM05a, BM06b, CH00, Che06c, DL03, dSBFG07, FS01a, GIA04, GK01, GC05, GL02a, Kwo01, KV05, LRZ02, LMH00, LK03, MT07a, Sim08, SS01, SCZ07, WBvdL07, YP07, YS08, Zaf02a]. treatment-specific [WBvdL07]. treatments [BB00, BL05a, BDH05, CC00, CS02d, DHV01, FGS06, GF00, GS00, GRP02, HML00, Hay07, Jac02, Jac03, Kwo01, MU03, MP07f, MS02e, PM08b, SBKF01, Udd08, YP07, YS08, Zha07, ZH09]. tree [CP07]. tree-order [CP07]. trees [CS03a, Cha00, KHL07, KO03, PP02a, RGL07, SA03]. trend [Gho01, GR01, KSL03, KL07, LS02, NJP05]. trend-free [LS02]. trends [AEK09a, MM02, SL00]. trial [BYW07, CG00, FS01a, FPvdVvH06, GF00, HC07a, LMD^{+}01, MPS07]. trials [BLU07, BBP06, BD08c, CH00, Che03b, CLS09, ED07, GIA04, GK01, GC05, Gov05, Gov09, GHSS02, HA08, Lee01b, LC09a, LP06, LZ09b, MW08b, MBA^{+}08, Pot07, PMS^{+}01, Sen01, ST05, Ste00, SCZ07, Wie05, WY07b, YW08, Zha09e]. triangle [KS01]. triangulation [Das07b]. trigonometric [Wu03]. trimethoprim [FS01a]. trimethoprim-sulfamethoxazole [FS01a]. Trimmed [SM01, WZ09, BW08, BJZ09, Cíz06, Hos07b, Mas09, MN03]. trinomial [FC04]. Triple [LC01a, BKLZ01, BZ00, CL03, Col01, DS02b, LR02, LW07, LK08, Mis02, She01, TS01]. triply [LR09]. trivariate [BNA04, JB09]. Trivedi [Con02b]. true [BR03a, O’H09, vdH05]. Truncated [GS09b, AF03, HY03a, HSV09, LY09, MY04, SY08a, Sun06a, Sun09, ZZF09, dCD09]. truncation [Arn09, BP01, LT09b, LT09c, MB00a, PSJ07, SH06, sS06a]. TTT [BB09]. tube [KT08c]. Tukey [CT04, Mas09]. tumor [MCH^{+}05]. tumours [NHZ04]. tuning [SL09]. tuple [GT08b]. turning [BLZ08]. Tweedie [HL09a]. Two [BS04, But09, CS03b, EN09, FR00, FR08, KM07b, Lia06c, MW05, PSW00, Rao01, Sar08, WMK07, Wri08, WW05, XW08, AEK09a, ATS02, BB00, BK05a, BP01, BDH05, BL05b, But05, But07, But08, CA04, CH00, Che04d, Cho07a, CP05b, Csá02, DR03a, DNN07, DDS00, EKL07, EPS08, FFP09, FBK03, FPR00a, FL08, GA08, GF00, GTT^{+}07, GC07, GGS09b, GM01c, GS07d, HS05, HML00, HLW07, HP07, HO03, HKMS02, HJR09, IS01, Jac09, KI08, KZ03a, KLA08, KHP06, KW09b, KQ02, KS03b, Kra09, KT04, KT08a, LF07, LK00, LL04, LRZ02, LL07a, Li08, Lia03, LS08, Liu00a, LS04, LP06, LJZ^{+}07, LT08c, LM00b, MS07b, MU03, MZ07, NdC08, Obi06, OO08, OS05b, RTG07, RH08, RL06, Rob09b, RM05b, SS02a, SW04, Sch04b, SST02, SBKF01, STW08, ySK05]. two [SSvRM03, Sun03, Sun06a, Tha08, TW05, TK06, VEM07, Wah09, WDE05, WZ06, WWN05, XL06b, XL06a, YCB06, YLH04, YHZ05, YB08, Yan08, YW09c, ZP00, ZC03, ZTQ04, ZQ05, ZQ07, vBHvdVW05, vdV08, vdW04, vdH05]. two- [XL06a]. two-action [KZ03a]. two-component [HS05, LS04]. Two-dimensional [Wri08, BK05a, Csá02, HO03, Tha08]. two-factor [GTT^{+}07, HP07, YB08]. Two-factorizations [FR00, FR08]. Two-layer [CS03b]. Two-level [But09, Lia06c, But05, But07, But08, FPR00a, HP07, Jac09, LS08, LM00b, Sch04b, ZP00, vdV08]. two-parameter [RM05b]. two-part [WZ06]. two-period [GF00]. two-phase [FFP09, KQ02, WDE05]. two-principal [GM01c]. two-sample [KI08, KLA08, SW04]. two-sided [HML00]. Two-stage [EN09, KM07b, Sar08, XW08, ATS02, CP05b, KW09b, LP06, LT08c, MZ07, Wah09]. two-state [VEM07]. two-step [Li08, vBHvdVW05]. two-treatment [CH00]. two-truncation [BP01]. two-variance [RH08]. two-way [CA04, GA08, LK00, YW09c, vdW04]. Type [BD09, AJMP09, AZL00, AH00, AWL04, BE06, BTK09, BN03, BK09c, BGI03, BW08, Bra06, Bur08, Che01, CCRS07, DFS08, DBDG09, Fer04, FR01, FPR00b, Ge01a, Ge01b, GHH00, GHL03, HCS04, HY08, IL06, Jac03, KKK03, Krä05, KM00b, LT06, Mei07a, MJK01, MW05, Nec06, OK00, RX03, Raq05, SGU05, SP05, WLC08, XP01, Yun00, ZL08, Zog08, BX07b, BX07a, BS08, BH08a, BD08a, BH09a, LWB09, SZW08, WJ08]. type-I [GHL03, BX07a, BH09a, LWB09]. Type-II [BD09, SGU05, BX07b, BS08, BH08a, BD08a]. Types [Pre01].

udder [MJD09]. Uhlenbeck [HKW00, TL09a]. UK [Con02b]. Ullah [Lin01b]. UMPU [FS02a]. UMPU-tests [FS02a]. unbalanced [GACL08, GA08, KK09b, KG05, Len02, SSvRM03, YW09c, YZ09]. Unbiased [Fre09, SM06, BR03b, BM06c, GFM05, Liu09, Mun00, Nay00, Nay03, NSV02, NWS01, PSW00, Rao02a, Sri02a, SS02c, SS02d, WZ04, WH01, WT06, WB05]. unbounded [KKW09]. uncertain [ES03]. uncertainty [Bul07a, BL03, OM08, TGM07, dC02]. uncorrelated [WC08]. underdispersed [dCPC05]. underdispersion [KMB08]. understanding [FRF09]. unequal [Bat04, CS02b, DT00, Dub07, GRP02, HK05, OS05b, Qua08, YR00]. unequally [JT01]. unidentified [Con01, For09]. unified [ARS01, KV05, NA09, SdMSL04]. Uniform [Maj07, PPM09, BJ06, BJ08, Che04b, Eri02, Heu08, KK02, Koh00, LY02, LM00a, Rob09b, XF00, ZN08]. Uniformity [CFQ05, Jup09, MFL03, Mei09, Pfa00, QL06]. uniformly [MW02]. Unifying [CLS09]. unimprovable [Gor07]. Union [Sen07a, Sen07b]. Union-intersection [Sen07a]. unique [Ber01b]. uniqueness [DI09, DK07, KL01]. unit [CCL03, DHT07, DPR06, JK03, MR00, STP02, Sán03, YR03]. Unitary [Sri07a]. United [PW06a, Par08, PW06b, dL06a]. units [BB00, WS08, Wri08]. unity [MRC03]. Univariate [QT07, Arn09, Bon09, Gar01, KKM01, LS07, LB08b, MS02b]. Universal [RA06, Tak09]. Universally [Udd05]. University [Con02b, Lin01b]. unknown [ATS02, BLL06, BS09, DR05, GD02, JT01, LdC06, Lia06a, MRLBBCPA00, NJP05, OS05b, PK05, QSL09, SC05a, SS02e, Tan01b, WZ04]. unobserved [ABS00, Kal07]. unreasonable [LP01]. unrelated [DdRB00, Liu00a, VR08]. unreliable [CD07b]. unreplicated [Sch04b, vdV08]. unrestricted [BP07, GGHS07]. Up-and-down [BG05, GF06]. Updating [PH05]. upgraded [GPV09]. upon [LLL09, SA03]. Upper [BLBR00, LBCM06, FS01c, GG08, LBR08, TSN07]. ups [DDM09]. Urn [Iva06, BK02, Gar02, MPS06, NSV02]. urns [BK02]. US$29.95 [Lin01b]. US$74.95 [Lin01b]. US$84.95 [Wes00]. US$89.95 [Lee01a]. Use [KI09, SK08, TSH06, Atk08b, Bha09, Bos06, CP03, CSX08, Fab09, FS01a, GS07d, JI07, KKM01, MS02b, YO05b]. used [Bat04, HC03]. Useful [Iva06, GP03, Han08b, SN03]. uses [Hos07b]. Using [Cha01c, Fie04, HDM09, KTB01, RHL04, RM05c, WY07b, ASD09, AWL04, BCN04, BNP06, BY09, BB04a, BCVV09, Ber02a, BK09b, BH07, BG09, BYW07, BB03b, CG01, CN02a, Cha03a, Cho07b, CN07, CD02, CP02b, DLM02, DZ06, EOGQ09, FCCN04, FCCN05, FS01a, FS07, GJ01, GT02, GT08b, GTV01, GG09a, GPV09, GS07d, HCS04, HH05, HT09b, HS08a, HYCY03, HK03, IOQ09, JR03b, KS03a, KM04, KHL07, Kot06, KBG09, KM03, LGV09, Lar03, LWL^{+}09, Lin07, LL07c, LZ00, LZ05, MT06, MJD09, MS09a, Nak00, NK06, Nay03, PNRdA03, PFF00, PS04a, Pap05, PP05, PP09, PW06a, PW06b, PSJ07, PdSPS05, PMS^{+}01, PS02, Sen02, SM06, SJ08, SM05b, SM07, Sub04b, dSSdL08, VS00, Vin02, WC09a, WH01, WRV05, WHS07, WY07a, WJ09, ZM06, dL06a, vdG02, vdH05]. Utilization [SC09]. utilizing [KG09].

V [Suk02]. V3 [OCR^{+}05]. vaccine [BLU07]. Valen [Con02a]. valid [Nak09]. Validation [Fre01, Fre02, GP04, RW05, Sim08, TSR03, TWH^{+}05]. validatory [EPdRV02]. validity [Aic01, Aus05, Kag08, PW03]. valuation [GT06, TT07]. value [BNA04, DdHL06, GM05, KS07d, Lan01, LB00, LBM06, MP08, MS07a, PQ09, Seg05, YLH04, Yun00, Yun02]. valued [GC09, MdCC05, RBN09, ZBD07]. values [BAA08, BNW05, BS07a, CS06, CR08, CW07, DH06, DW00, GO07, HdGT09, IVM03, KM03, KNH09, LBR08, NdC08, ODY08, SB05, Sar00, SSH08, vBLO04]. VAR [Ioa07, TT08]. variability [Zog08]. Variable [CD03a, DR03b, Kan02, BWA06, cICY04, CG08b, EKT06, GGS09a, GKMS02, GHKTR07, GCR06, GMM03b, GPL09, HLP06, Koh06, KBG09, KLX09, Lia06b, NDF03, SC04, ZG08]. variable-neighbourhood [GGS09a]. variables [And07b, BdM09, Cau05, CR08, CS03b, CB06, Chr04, DR01a, EPS08, FD02, FR01, Fot00, FN07, GL09, GC07, GL02a, GS07d, GG09b, HLM08, HL01, HR04, KL01, Kag06, KZ03a, KKM09, KK02, Kör00, KMV05, Ma00, NaX03, PH05, RSSR09, SGK09, SC09, VdCO00, WG05, Wit01, WLCI09, XY08, Xin09, YZ06, Zog08].

Variance [Ber05, DT05, Hei00, ME04, TL09b, ABK00, And06, BB04b, BM06c, BSP02, BW03, BW05, Bur07, CCL03, CA04, CCP09, CD02, DS03, DM08b, ES04, Gen01, GS09a, GTV01, GD02, GHSN00, GC04, HK05, HS00a, JT01, Jon04, KW09b, KSK00, KMS09b, LT02, LRZ02, Lia06a, MS06, MG08, OS05a, OS06, PS04b, PSJ07, RH08, RS04b, Ryc01, SM06, SCW09, SR02, SSvRM03, VdCO00, VS00, WZ04, WDE05, WD04, WYL09b, YZZ09, ZEB07]. Variance-balanced [ME04]. variance-components [BW03]. variance-covariance [BSP02]. variance-reduction [VdCO00]. variances [Bat04, BL02b, GTV01, HK05, LN03, MP08, OS05b, QSL09, YR00]. variants [CvRZ01]. variate [DGGJ09, WC00]. variation [Ahm02, GFM05, GK00, LFRL07, Net02, SM05b, WM08, WW02c]. Variations [ACS07]. variogram [GSFBGM04, GSC07, LLC02]. various [FTY08, YS08]. varying [AL09, Hon04, Len09, LC07, Lu08, PGLS08, UB07, Wan09, WGCI09, WH09]. varying-coefficient [Len09, Lu08, WGCI09]. Vector [OCR^{+}05, CC09, EMP06, FOS08, GD08a, Kak02, LJC07, LLW07, SN04, TL08, WY06]. vector-autoregressive [SN04]. vectors [ATS02, BRdCR03, CZ02, GFDMG04, KA09, MW08a, PC02, SN08, WC09a]. vehicle [dHD^{+}09]. Venkata [Moh02b]. verification [NdC08, ZC03]. Verlag [Sin06b]. version [CP03, CKS07b, Kag01a]. versions [GH02]. versus [CS05, DK09, DL04, FS01a, GA05, GM01b, NPF09, OM08, Qua08, WH05]. vertices [JTS02, KDT06, Ker04, ZY02]. Vervaat [CCRS07]. Vervaat-error-type [CCRS07]. via [ÅvR06, ANN08, AA08, AKI08, BWA06, DR03b, EPS08, Gil05, GRC09, GP04, HNW08, JPS^{+}08, KK09b, KMS08, KG05, Kuh05, LTW08, LLZ07, Mei09, MW07a, MR01, Phi02, PPX09, SAC02, Shi08, SV05, TK09, WCS01, YA09]. viability [GS08]. vicious [GV02]. view [AS00, BJPZ08, Qui06]. viewed [Hai07]. Vincze [Csá05]. violations [GPC01]. viral [HC03]. visualization [vdLP03]. vol [Gov09]. volatility [Bri08, GS07a, GT09, HKZ08, Laï05, LC07, Tri08]. Volume [Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02h, Ano02i, Ano03j, Sri02b, Ano02b, Ano02c, Ano02d, Ano02f, Ano02g, Ano02e, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03b, Ano03i, Ano04a, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04b, Ano04h, Ano04i, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano06a]. Vos [Cha02a]. vs [CG08b, DL03].

W [Cha02a, Ko00]. Waiting [AH08, DLS09, MS09b, Min07]. Wald [HBK03, Li01, Shi05, War08]. walk [Böh02, Csá02, EG00]. walkers [GV02]. walks [DK07, KP02, Kem02b, KO03, Nie02b, Nie05]. Warner [KW04]. Watson [CIPS07, YW09a]. Wavelet [Boc07, Oye03, PK06, Cai02, Chi03, CJRV08, GN03, KV06, Li03, LX07b, Li07, OK08, Pen00, VBF03, Xue04]. wavelet-based [Li03]. wavelets [WCS01]. way [AG07, CA04, Dav04, Dob02, DTW06, EBZ08, Fuj02, GACL08, GA08, GHSN00, HB00, HK05, KG05, KD07, LK00, LZ09a, MP02, Sah08, Sch00b, Wyn08, YW09c, ZEB07, vdW04]. Weak [Che04b, DM08a, FT09, GB09, LT06, ETS02, FZ00, HS00a, SC05b]. weakly [Can06, SV00]. Weibull [DGGG08, DBDG09, GG08, GK06b, Jar07, KK06, Kot06, KLX09, Kun07, SGU05, WWN05, Yu06]. Weibull-type [DBDG09]. Weidner [Bro06]. weighing [LZN06]. Weight [CK01, Che06a, FS01c, GH01, KMB08, Tsa06, Zha01b]. Weighted [BC03a, FWW09, GRP02, HLM08, NB05, Ste07, And06, ADW03, CD06, Cua09, DGR07, GMPZ06, Hal07, Heu08, HL01, Ino03, KS03a, KKB09b, KLE^{+}04c, KD07, MI08a, Obi06, PNRdA03, SW07, sS06a, TK06, WvEZ04, WK01, WGCI09, XP02, vBP03]. Weighted-bootstrap [HLM08]. weighting [FZ06, GKM07, Shi00]. weights [Col01, FS07, KD00, Pla09, Wie00, YR03]. Weinberg [CQ08]. Weisberg [Lee01a]. Weitner [dL06a]. Welch [Xin01]. well [HK04]. whether [NJP05]. Which [War08, BCR08, KS08a, vdH05]. white [Kun00]. Whiteman [Cao01]. Whitney [XP02, BB04a, TB07]. Whittle [Sha08]. Wicksell [MRU00]. width [LSW06]. Wielandt [LLP09]. Wiener [Hai00]. Wilbur [LOV08]. Wilcoxon [BTK09, BB04a, TB07, XP02]. Wilcoxon-Mann-Whitney [BB04a, TB07]. Wilcoxon-type [BTK09]. Wiley [Cha02a, Dea00, Ko00, Lah00, Lee01a, Lu02, Roy07, Wes00]. Wilks [AC07]. Wilson [QRC01, MW05]. wind [Fie04]. window [Alv05, BJJV08, ZG08]. winner [lCV06]. Winsorized [WZ09]. wise [CKMS02, FHL03, Gor07, Hag06, OK00, Sch07a]. Wishart [DGJ06]. Wissenschaften [Sin06b]. within [BC09, Car02, DL06b, GF00]. without [BN09, CS02b, GR08a, LBCM06, PR04, vdWD01]. without-replacement [GR08a]. WOC [CLA09]. wordlength [BW09]. work [LP09, Pan06, SV01a]. worker [LWL^{+}09]. Workshop [Ano05a]. wrap [ZN08]. wrap-around [ZN08]. wrong [DZ06]. Wu [Pal09].

x [Roy07, Lee01a, Mas01]. xii [Cha02a, Sin06b]. xiii [Ko00]. xiv [Lah00]. xvii [Cha01a]. XXX [Ano05j, Ano06a].

years [Rao01]. Yor [BW07]. York [Cha01a, Cha02a, Con02a, Dea00, Ko00, Lah00, Lee01a, Lu02, Roy07, Wes00]. Youden [KS07e, MP07c]. Yule [XD05].

Z [LZ09b]. Z. [Gov09]. Zarankiewicz [GG01]. Zellner [DdRB00]. zero [GML06, LTW08, Nay03, YHA09c]. zero-inflated [GML06, YHA09c].

References

Abramovich:2006:TME

[AA06] Felix Abramovich and Claudia Angelini. Testing in mixedeffects FANOVA models. Journal of Statistical Planning and Inference, 136(12):4326–4348, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001400.

Adell:2008:SBA

[AA08] José A. Adell and José M. Anoz. Signed binomial approximation of binomial mixtures via differential calculus for linear operators. Journal of Statistical Planning and Inference, 138(12):3687–3695, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001249.

Ahmad:2006:AND

[AAEK06] I. A. Ahmad, A. Ahmed, I. Elbatal, and M. Kayid. An aging notion derived from the increasing convex ordering: the NBUCA class. Journal of Statistical Planning and Inference, 136(3):555–569, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003544.

Aerts:2000:CLT

[AAJ00] Marc Aerts, Ilse Augustyns, and Paul Janssen. Central limit theorem for the total squared error of local polynomial estimators of cell probabilities. Journal of Statistical Planning and Inference, 91(2):181–193, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001774.

Abdous:2005:MRL

[AB05] Belkacem Abdous and Alexandre Berred. Mean residual life estimation. Journal of Statistical Planning and Inference, 132(1–2):3–19, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804001387.

Anderson:2009:SMD

[AB09] Charles L. Anderson and J. Calvin Berry. On a simple measure of dominance. Journal of Statistical Planning and Inference, 139(3):1098–1108, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002930.

Albers:2000:ABT

[ABK00] Willem Albers, Pieta C. Boon, and Wilbert C. M. Kallenberg. The asymptotic behavior of tests for normal means based on a variance pre-test. Journal of Statistical Planning and Inference, 88(1):47–57, July 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002116.

Alkali:2009:FMD

[ABQG09] Babakalli M. Alkali, Tim Bedford, John Quigley, and Jim Gaw. Failure and maintenance data extraction from power plant maintenance management databases. Journal of Statistical Planning and Inference, 139(5):1766–1776, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002516.

Almudevar:2000:EPM

[ABS00] A. Almudevar, R. N. Bhattacharya, and C. C. A. Sastri. Estimating the probability mass of unobserved support in random sampling. Journal of Statistical Planning and Inference, 91(1):91–105, November 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001257.

Abel:2000:PMD

[ABZ00] R. J. R. Abel, F. E. Bennett, and H. Zhang. Perfect Mendelsohn designs with block size six. Journal of Statistical Planning and Inference, 86(2):287–319, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001147.

Agresti:2002:ACT

[AC02] Alan Agresti and Brent A. Coull. The analysis of contingency tables under inequality constraints. Journal of Statistical Planning and Inference, 107(1–2):45–73, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002434.

Aldworth:2003:PNS

[AC03] Jeremy Aldworth and Noel Cressie. Prediction of nonlinear spatial functionals. Journal of Statistical Planning and Inference, 112(1–2):3–41, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200321X.

Amaratunga:2004:MDF

[AC04a] Dhammika Amaratunga and Javier Cabrera. Mining data to find subsets of high activity. Journal of Statistical Planning and Inference, 122(1–2):23–41, ???? 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002465.

Augustin:2004:NPI

[AC04b] T. Augustin and F. P. A. Coolen. Nonparametric predictive inference and interval probability. Journal of Statistical Planning and Inference, 124(2):251–272, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002246.

Armero:2006:BHM

[AC06] Carmen Armero and David Conesa. Bayesian hierarchical models in manufacturing bulk service queues. Journal of Statistical Planning and Inference, 136(2):335–354, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003180.

Alberto:2007:SQS

[AC07] Rui P. Alberto and Carlos A. Coelho. Study of the quality of several asymptotic and near-exact approximations based on moments for the distribution of the Wilks Lambda statistic. Journal of Statistical Planning and Inference, 137(5):1612– 1626, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002400.

Anderson-Cook:2009:RRS

[ACBM09a] Christine M. Anderson-Cook, Connie M. Borror, and Douglas C. Montgomery. Rejoinder for “Response surface design evaluation and comparison”. Journal of Statistical Planning and Inference, 139(2):671–674, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001675.

Anderson-Cook:2009:RSD

[ACBM09b] Christine M. Anderson-Cook, Connie M. Borror, and Douglas C. Montgomery. Response surface design evaluation and comparison. Journal of Statistical Planning and Inference, 139(2):629–641, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0378375808001572. See discussion [Jon09, Par09b, Khu09, Bor09, Pie09, Goo09, Luc09, Rob09c].

Arnold:2004:CPC

[ACS04] Barry C. Arnold, Enrique Castillo, and José Mariía Sarabia. Compatibility of partial or complete conditional probability specifications. Journal of Statistical Planning and Inference, 123(1):133–159, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580300137X.

Arnold:2007:VCM

[ACS07] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Variations on the classical multivariate normal theme. Journal of Statistical Planning and Inference, 137(11):3249–3260, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000948.

Arnold:2008:MDD

[ACS08] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Multivariate distributions defined in terms of contours. Journal of Statistical Planning and Inference, 138(12):4158–4171, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001262.

Aerts:2002:STP

[ACW02] M. Aerts, G. Claeskens, and M. P. Wand. Some theory for penalized spline generalized additive models. Journal of Statistical Planning and Inference, 103(1–2):455–470, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002373.

Aubry:2001:ANM

[AD01] Christophe Aubry and Ali Souleyman Dabye. Asymptotic normality of the minimum distance estimators for a Poisson process with a discontinuous intensity function. Journal of Statistical Planning and Inference, 99(1):3–23, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801000787.

Alves:2009:TPD

[AdHN09] Isabel Fraga Alves, Laurens de Haan, and Cláudia Neves. A test procedure for detecting super-heavy tails. Journal of Statistical Planning and Inference, 139(2):213–227, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001869.

Afshartous:2005:PDA

[AdL05] David Afshartous and Jan de Leeuw. A predictive density approach to predicting a future observable in multilevel models. Journal of Statistical Planning and Inference, 128(1):149–164, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003689.

Antille:2003:NOD

[ADW03] Gérard Antille, Holger Dette, and Anna Weinberg. A note on optimal designs in weighted polynomial regression for the classical efficiency functions. Journal of Statistical Planning and Inference, 113(1):285–292, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002889.

Angelopoulos:2009:ROE

[AEK09a] P. Angelopoulos, H. Evangelaras, and C. Koukouvinos. Run orders for efficient two level experimental plans with minimum factor level changes robust to time trends. Journal of Statistical Planning and Inference, 139(10):3718–3724, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001281.

Angelopoulos:2009:SBE

[AEK09b] P. Angelopoulos, H. Evangelaras, and C. Koukouvinos. Small, balanced, efficient and near rotatable central composite designs. Journal of Statistical Planning and Inference, 139(6): 2010–2013, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003686.

Ahmad:2003:RRS

[AF03] Abd El-Baset A. Ahmad and Mohammad A. Fawzy. Recurrence relations for single moments of generalized order statistics from doubly truncated distributions. Journal of Statistical Planning and Inference, 117(2):241–249, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003853.

Andronov:2009:SEF

[AFF09] Alexander Andronov, Helen Fioshina, and Maxim Fioshin. Statistical estimation for a failure model with damage accumulation in a case of small samples. Journal of Statistical Planning and Inference, 139(5):1685–1692, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002449.

Ai:2007:OML

[AFH07] Mingyao Ai, Kai-Tai Fang, and Shuyuan He. E(χ^{2})-optimal mixed-level supersaturated designs. Journal of Statistical Planning and Inference, 137(1):306–316, January 1, 2007.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580500279X.

Andrews:2005:CDF

[AFW05] D. F. Andrews, D. A. S. Fraser, and A. C. M. Wong. Computation of distribution functions from likelihood information near observed data. Journal of Statistical Planning and Inference, 134(1):180–193, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400206X.

Aggarwal:2004:NMC

[AG04] M. L. Aggarwal and Seema Gupta. A new method of construction of multi-level supersaturated designs. Journal of Statistical Planning and Inference, 121(1):127–134, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802005049.

Agresti:2007:IMW

[AG07] Alan Agresti and Anna Gottard. Independence in multiway contingency tables: S. N. Roy’s breakthroughs and later developments. Journal of Statistical Planning and Inference, 137(11):3216–3226, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000912.

Arslan:2009:GMS

[AG09] Olcay Arslan and Ali I. Genç. A generalization of the multivariate slash distribution. Journal of Statistical Planning and Inference, 139(3):1164–1170, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002966.

Agarwal:2002:DNS

[Aga02] Manju Agarwal. Distribution of number served during a busy period of GI/M/1 N queues-lattice path approach. Journal of Statistical Planning and Inference, 101(1–2):7–21, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001483.

Abel:2001:RBI

[AGGZ01] R. Julian R. Abel, G. Ge, Malcolm Greig, and L. Zhu. Resolvable balanced incomplete block designs with block size 5. Journal of Statistical Planning and Inference, 95(1–2):49– 65, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002779.

Antoch:2007:TCR

[AGH07] Jaromír Antoch, Gérard Grégoire, and Marie Husková. Tests for continuity of regression functions. Journal of Statistical Planning and Inference, 137(3):753–777, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580600139X.

Atienza:2007:CMF

[AGHMPV07] N. Atienza, J. García-Heras, J. M. Muñoz-Pichardo, and R. Villa. On the consistency of MLE in finite mixture models of exponential families. Journal of Statistical Planning and Inference, 137(2):496–505, February 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000425.

Antos:2000:LBR

[AGK00] András Antos, László Györfi, and Michael Kohler. Lower bounds on the rate of convergence of nonparametric regression estimates. Journal of Statistical Planning and Inference, 83(1):91–100, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000841.

Agostinelli:2009:BBM

[Ago09] Claudio Agostinelli. Bias bound for the minimax estimator. Journal of Statistical Planning and Inference, 139(7):2235– 2241, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003960.

Argiento:2009:CNP

[AGP09] Raffaele Argiento, Alessandra Guglielmi, and Antonio Pievatolo. A comparison of nonparametric priors in hierarchical mixture modelling for AFT regression. Journal of Statistical Planning and Inference, 139(12):3989–4005, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001293.

Antoch:2000:BTE

[AH00] Jaromiír Antoch and Marie Husková. Bayesian-type estimators of change points. Journal of Statistical Planning and Inference, 91(2):195–208, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001786.

Andersen:2001:MCD

[AH01] Kim E. Andersen and Martin B. Hansen. Multiplicative censoring: density estimation by a series expansion approach. Journal of Statistical Planning and Inference, 98(1–2):137–155, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002378.

Afsarinejad:2002:RMD

[AH02] K. Afsarinejad and A. S. Hedayat. Repeated measurements designs for a model with self and simple mixed carryover effects. Journal of Statistical Planning and Inference, 106(1–2):449–459, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002276.

Al-Hussaini:2003:BPD

[AH03] Essam K. Al-Hussaini. Bayesian predictive density of order statistics based on finite mixture models. Journal of Statistical Planning and Inference, 113(1):15–24, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100297X.

Aki:2008:WTD

[AH08] Sigeo Aki and Katuomi Hirano. Waiting time distributions for a run with additional constraints. Journal of Statistical Planning and Inference, 138(11):3492–3501, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800092X.

Ahmad:2001:MIA

[Ahm01] Ibrahim A. Ahmad. Moments inequalities of aging families of distributions with hypotheses testing applications. Journal of Statistical Planning and Inference, 92(1–2):121–132, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001397.

Ahmed:2002:SEC

[Ahm02] S. E. Ahmed. Simultaneous estimation of coefficients of variation. Journal of Statistical Planning and Inference, 104(1):31– 51, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001215.

Ahmad:2005:SAO

[Ahm05a] Ibrahim A. Ahmad. Statistical analysis of old age behavior. Journal of Statistical Planning and Inference, 129(1– 2):239–252, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804001880.

Ahmed:2005:APC

[Ahm05b] S. E. Ahmed. Assessing the process capability index for non-normal processes. Journal of Statistical Planning and Inference, 129(1–2):195–206, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001818.

Ahsanullah:2000:GOS

[Ahs00] M. Ahsanullah. Generalized order statistics from exponential distribution. Journal of Statistical Planning and Inference, 85(1–2):85–91, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000683.

Aickin:2001:RBV

[Aic01] Mikel Aickin. Randomization, balance, and the validity and efficiency of design-adaptive allocation methods. Journal of Statistical Planning and Inference, 94(1):97–119, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002287.

Arsenjev:2000:ACS

[AIK00] D. G. Arsenjev, V. M. Ivanov, and O. Yu. Kul’chitsky. Adaptive control of stochastic calculating processes. Journal of Statistical Planning and Inference, 85(1–2):213–226, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000828.

Aires:2000:CBC

[Air00] Nibia Aires. Comparisons between conditional Poisson sampling and Pareto πps sampling designs. Journal of Statistical Planning and Inference, 88(1):133–147, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002050.

Ahmadi:2009:BEB

[AJMP09] Jafar Ahmadi, Mohammad Jafari Jozani, Éric Marchand, and Ahmad Parsian. Bayes estimation based on k-record data from a general class of distributions under balanced type loss functions. Journal of Statistical Planning and Inference, 139(3):1180–1189, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003017.

Ando:2003:CND

[AK03] Masakazu Ando and Miyoshi Kimura. A characterization of the neighborhoods defined by certain special capacities and their applications to bias-robustness of estimates. Journal of Statistical Planning and Inference, 116(1):61–90, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001830.

Antzoulakos:2005:CPR

[AK05] Demetrios L. Antzoulakos and Markos V. Koutras. On a class of polynomials related to classical orthogonal and Fibonacci polynomials with probability applications. Journal of Statistical Planning and Inference, 135(1):18–39, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000522.

Albers:2008:MCC

[AK08] Willem Albers and Wilbert C. M. Kallenberg. Minimum control charts. Journal of Statistical Planning and Inference, 138(3):539–551, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002849.

Ando:2008:NRM

[AKI08] Tomohiro Ando, Sadanori Konishi, and Seiya Imoto. Nonlinear regression modeling via regularized radial basis function networks. Journal of Statistical Planning and Inference, 138(11):3616–3633, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001031.

Ai:2009:BOB

[AKJ09] Mingyao Ai, Lulu Kang, and V. Roshan Joseph. Bayesian optimal blocking of factorial designs. Journal of Statistical Planning and Inference, 139(9):3319–3328, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000743.

Ando:2009:RNC

[AKK09] Masakazu Ando, Itsuro Kakiuchi, and Miyoshi Kimura. Robust nonparametric confidence intervals and tests for the median in the presence of (c, γ)-contamination. Journal of Statistical Planning and Inference, 139(6):1836–1846, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808003856.

Albers:2004:PCC

[AKN04] Willem Albers, Wilbert C. M. Kallenberg, and Sri Nurdiati. Parametric control charts. Journal of Statistical Planning and Inference, 124(1):159–184, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002003.

Albers:2006:DDC

[AKN06] Willem Albers, W. C. M. Kallenberg, and Sri Nurdiati. Data driven choice of control charts. Journal of Statistical Planning and Inference, 136(3):909–941, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003234.

Ahn:2007:LLM

[AKP^{+}07] C. Ahn, G. G. Koch, L. Paynter, J. S. Preisser, and F. Seillier-Moiseiwitsch. Log-linear models for mutations in the HIV genome. Journal of Statistical Planning and Inference, 137(11):3227–3239, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000924.

Adell:2008:SDB

[AL08] José A. Adell and Alberto Lekuona. Shortening the distance between Edgeworth and Berry–Esseen in the classical case. Journal of Statistical Planning and Inference, 138(4):1167–1178, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807002327.

Arasan:2009:SMP

[AL09] Jayanthi Arasan and Mary Lunn. Survival model of a parallel system with dependent failures and time varying covariates. Journal of Statistical Planning and Inference, 139(3):944–951, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800284X.

Artalejo:2007:SSD

[ALH07] J. R. Artalejo and M. J. Lopez-Herrero. A simulation study of a discrete-time multiserver retrial queue with finite population. Journal of Statistical Planning and Inference, 137(8):2536–2542, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000092.

Ahmed:2001:ASB

[ALRV01] S. E. Ahmed, Deli Li, Andrew Rosalsky, and Andrei I. Volodin. Almost sure lim sup behavior of bootstrapped means with applications to pairwise i.i.d. sequences and stationary ergodic sequences. Journal of Statistical Planning and Inference, 98(1–2):1–14, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003220.

Anh:2004:CMC

[ALS04] V. V. Anh, N. N. Leonenko, and L. M. Sakhno. On a class of minimum contrast estimators for fractional stochastic processes and fields. Journal of Statistical Planning and Inference, 123(1):161–185, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001368.

Anh:2007:MCE

[ALS07] V. V. Anh, N. N. Leonenko, and L. M. Sakhno. Minimum contrast estimation of random processes based on information of second and third orders. Journal of Statistical Planning and Inference, 137(4):1302–1331, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000644.

Alvarez:2005:SNE

[Alv05] Enrique E. Alvarez. Smoothed nonparametric estimation in window censored semi-Markov processes. Journal of Statistical Planning and Inference, 131(2):209–229, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000850.

Ausin:2007:BCN

[ALW07] M. Concepción Ausín, Rosa E. Lillo, and Michael P. Wiper. Bayesian control of the number of servers in a GI/M/c queueing system. Journal of Statistical Planning and Inference, 137(10):3043–3057, October 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000730.

Andronov:2000:OSS

[AM00] Alexander Andronov and Yuri Merkuryev. Optimization of statistical sample sizes in simulation. Journal of Statistical Planning and Inference, 85(1–2):93–102, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000725.

Arasu:2001:NRD

[AM01] K. T. Arasu and S. L. Ma. A nonexistence result on difference sets, partial difference sets and divisible difference sets. Journal of Statistical Planning and Inference, 95(1–2): 67–73, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002780.

Aban:2004:GLS

[AM04a] Inmaculada B. Aban and Mark M. Meerschaert. Generalized least-squares estimators for the thickness of heavy tails. Journal of Statistical Planning and Inference, 119(2):341–352, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004196.

Ahmad:2004:FMI

[AM04b] Ibrahim A. Ahmad and A. R. Mugdadi. Further moments inequalities of life distributions with hypothesis testing applications: the IFRA, NBUC and DMRL classes. Journal of Statistical Planning and Inference, 120(1–2):1–12, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004937.

Arino:2004:PTD

[AM04c] Miguel A. Ariño and Francesc Marmol. A permanenttransitory decomposition for ARFIMA processes. Journal of Statistical Planning and Inference, 124(1):87–97, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001903.

Adrover:2002:RBM

[AMY02] Jorge G. Adrover, Ricardo A. Maronna, and Viíctor J. Yohai. Relationships between maximum depth and projection regression estimates. Journal of Statistical Planning and Inference, 105(2):363–375, July 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002646.

Adrover:2004:RRQ

[AMY04] Jorge Adrover, Ricardo A. Maronna, and Víctor J. Yohai. Robust regression quantiles. Journal of Statistical Planning and Inference, 122(1–2):187–202, ???? 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002362.

Ahsanullah:2000:GSO

[AN00] M. Ahsanullah and V. B. Nevzorov. Generalized spacings of order statistics from extended sample. Journal of Statistical Planning and Inference, 85(1–2):75–83, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000671.

Andronov:2005:MCA

[And05] Alexander Andronov. Monte-Carlo approximation for probability distribution of monotone Boolean function. Journal of Statistical Planning and Inference, 132(1–2):21–31, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001399.

Andersson:2006:CPW

[And06] Per Gösta Andersson. A conditional perspective of weighted variance estimation of the optimal regression estimator. Journal of Statistical Planning and Inference, 136(1):221–234, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002794.

Anderson:2007:MDD

[And07a] T. W. Anderson. Multiple discoveries: Distribution of roots of determinantal equations. Journal of Statistical Planning and Inference, 137(11):3240–3248, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000936.

Andronov:2007:NEE

[And07b] Alexander Andronov. On nonparametric estimation of expectation of random variables function under dependent observations. Journal of Statistical Planning and Inference, 137(12):3828–3837, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001528.

Anis:2010:CSA

[Ani10] M. Z. Anis. Corrigendum to: “An L-statistic approach to a test of exponentiality against IFR alternatives” [Journal of Statistical Planning and Inference 138 (2008) 3144–3148]. Journal of Statistical Planning and Inference, 140(6):1618, June 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809004005. See [MA08].

Abegaz:2008:MSD

[ANN08] Fentaw Abegaz and U. V. Naik-Nimbalkar. Modeling statistical dependence of Markov chains via copula models. Journal of Statistical Planning and Inference, 138(4):1131–1146, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002741.

Anonymous:2000:BRa

[Ano00a] Anonymous. Book review. Journal of Statistical Planning and Inference, 83(1):281–282, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000270.

Anonymous:2000:BRb

[Ano00b] Anonymous. Book review. Journal of Statistical Planning and Inference, 83(2):475–476, February 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001056.

Anonymous:2000:BRc

[Ano00c] Anonymous. Book review. Journal of Statistical Planning and Inference, 83(2):477–480, February 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001068.

Anonymous:2000:BRd

[Ano00d] Anonymous. Book review. Journal of Statistical Planning and Inference, 84(1–2):345–346, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001354.

Anonymous:2000:BRe

[Ano00e] Anonymous. Book review. Journal of Statistical Planning and Inference, 84(1–2):347–348, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001366.

Anonymous:2000:BRf

[Ano00f] Anonymous. Book review. Journal of Statistical Planning and Inference, 84(1–2):349–351, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001378.

Anonymous:2000:BRg

[Ano00g] Anonymous. Book review. Journal of Statistical Planning and Inference, 86(1):279–281, April 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000579.

Anonymous:2000:BRh

[Ano00h] Anonymous. Book review. Journal of Statistical Planning and Inference, 87(1):173–175, May 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001810.

Anonymous:2000:BRi

[Ano00i] Anonymous. Book review. Journal of Statistical Planning and Inference, 87(2):367–368, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000690.

Anonymous:2000:BRj

[Ano00j] Anonymous. Book review. Journal of Statistical Planning and Inference, 87(2):369–370, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000884.

Anonymous:2000:BRk

[Ano00k] Anonymous. Book review. Journal of Statistical Planning and Inference, 87(2):371–373, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000896.

Anonymous:2000:D

[Ano00l] Anonymous. Discussion. Journal of Statistical Planning and Inference, 83(1):277–279, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001190.

Anonymous:2000:E

[Ano00m] Anonymous. Editorial. Journal of Statistical Planning and Inference, 85(1–2):1, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000580.

Anonymous:2000:H

[Ano00n] Anonymous. (huskova). Journal of Statistical Planning and Inference, 91(2):??, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:Ia

[Ano00o] Anonymous. Index. Journal of Statistical Planning and Inference, 83(2):481–482, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001901.

Anonymous:2000:Ib

[Ano00p] Anonymous. Index. Journal of Statistical Planning and Inference, 84(1–2):353–354, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899002104.

Anonymous:2000:Ic

[Ano00q] Anonymous. Index. Journal of Statistical Planning and Inference, 85(1–2):227–228, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800000732.

Anonymous:2000:Id

[Ano00r] Anonymous. Index. Journal of Statistical Planning and Inference, 86(2):629–630, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001087.

Anonymous:2000:Ie

[Ano00s] Anonymous. Index. Journal of Statistical Planning and Inference, 87(2):375–376, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001208.

Anonymous:2000:If

[Ano00t] Anonymous. Index. Journal of Statistical Planning and Inference, 88(2):355–356, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001464.

Anonymous:2000:Ig

[Ano00u] Anonymous. Index. Journal of Statistical Planning and Inference, 89(1–2):287–288, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001555.

Anonymous:2000:Ih

[Ano00v] Anonymous. Index. Journal of Statistical Planning and Inference, 90(2):349–350, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001750.

Anonymous:2000:Ii

[Ano00w] Anonymous. Index. Journal of Statistical Planning and Inference, 91(2):577–578, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580000238X.

Anonymous:2000:PJc

[Ano00x] Anonymous. Pages 1-170 (1 July 2000). Journal of Statistical Planning and Inference, 88(1):??, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PS

[Ano00y] Anonymous. Pages 1-176 (1 September 2000). Journal of Statistical Planning and Inference, 90(1):??, September 1, 2000.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PMc

[Ano00z] Anonymous. Pages 1-176 (15 May 2000). Journal of Statistical Planning and Inference, 87(1):??, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PN

[Ano00-27] Anonymous. Pages 1-178 (1 November 2000). Journal of Statistical Planning and Inference, 91(1):??, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PAa

[Ano00-28] Anonymous. Pages 1-228 (1 April 2000). Journal of Statistical Planning and Inference, 85(1–2):??, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PJa

[Ano00-29] Anonymous. Pages 1-282 (1 January 2000). Journal of Statistical Planning and Inference, 83(1):??, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PAb

[Ano00-30] Anonymous. Pages 1-282 (15 April 2000). Journal of Statistical Planning and Inference, 86(1):??, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PAd

[Ano00-31] Anonymous. Pages 1-288 (15 August 2000). Journal of Statistical Planning and Inference, 89(1–2):??, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PMa

[Ano00-32] Anonymous. Pages 1-354 (March 2000). Journal of Statistical Planning and Inference, 84(1–2):??, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PAc

[Ano00-33] Anonymous. Pages 171-356 (1 August 2000). Journal of Statistical Planning and Inference, 88(2):??, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PO

[Ano00-34] Anonymous. Pages 177-350 (1 October 2000). Journal of Statistical Planning and Inference, 90(2):??, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PJb

[Ano00-35] Anonymous. Pages 177-376 (1 June 2000). Journal of Statistical Planning and Inference, 87(2):??, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PF

[Ano00-36] Anonymous. Pages 283-482 (1 February 2000). Journal of Statistical Planning and Inference, 83(2):??, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:PMb

[Ano00-37] Anonymous. Pages 283-630 (1 May 2000). Journal of Statistical Planning and Inference, 86(2):??, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2000:Pb

[Ano00-38] Anonymous. Preface. Journal of Statistical Planning and Inference, 88(2):171, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800000756.

Anonymous:2000:Pc

[Ano00-39] Anonymous. Preface. Journal of Statistical Planning and Inference, 91(2):179–180, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001762.

Anonymous:2001:AIVa

[Ano01a] Anonymous. Author index volume 92(2001). Journal of Statistical Planning and Inference, 92(1–2):307–308, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002731.

Anonymous:2001:AIVb

[Ano01b] Anonymous. Author index volume 93 (2001). Journal of Statistical Planning and Inference, 93(1–2):325–326, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003542.

Anonymous:2001:AIVc

[Ano01c] Anonymous. Author index volume 94 (2001). Journal of Statistical Planning and Inference, 94(2):429–430, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000702.

Anonymous:2001:AIVd

[Ano01d] Anonymous. Author index volume 95 (2001). Journal of Statistical Planning and Inference, 95(1–2):359–360, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000738.

Anonymous:2001:AIVe

[Ano01e] Anonymous. Author index volume 96 (2001). Journal of Statistical Planning and Inference, 96(2):439–440, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001033.

Anonymous:2001:AIVf

[Ano01f] Anonymous. Author index volume 97 (2001). Journal of Statistical Planning and Inference, 97(2):415–416, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001161.

Anonymous:2001:AIVg

[Ano01g] Anonymous. Author index volume 98 (2001). Journal of Statistical Planning and Inference, 98(1–2):311–312, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001173.

Anonymous:2001:AIVh

[Ano01h] Anonymous. Author index volume 99 (2001). Journal of Statistical Planning and Inference, 99(2):223, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002786.

Anonymous:2001:B

[Ano01i] Anonymous. (Bera). Journal of Statistical Planning and Inference, 97(1):??, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:Ca

[Ano01j] Anonymous. Contents. Journal of Statistical Planning and Inference, 93(1–2):v, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000696.

Anonymous:2001:Cb

[Ano01k] Anonymous. Contents. Journal of Statistical Planning and Inference, 94(2):iii–iv, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000714.

Anonymous:2001:G

[Ano01l] Anonymous. (gupta). Journal of Statistical Planning and Inference, 96(1):??, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:LR

[Ano01m] Anonymous. List of referees 2000. Journal of Statistical Planning and Inference, 93(1–2):vii–xiv, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003554.

Anonymous:2001:PNb

[Ano01n] Anonymous. Pages 1-110 (1 November 2001). Journal of Statistical Planning and Inference, 99(1):??, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PM

[Ano01o] Anonymous. Pages 1-120 (1 March 2001). Journal of Statistical Planning and Inference, 94(1):??, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PJa

[Ano01p] Anonymous. Pages 1-308 (January 2001). Journal of Statistical Planning and Inference, 92(1–2):??, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PO

[Ano01q] Anonymous. Pages 1-312 (1 October 2001). Journal of Statistical Planning and Inference, 98(1–2):??, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PF

[Ano01r] Anonymous. Pages 1-326 (February 2001). Journal of Statistical Planning and Inference, 93(1–2):??, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PD

[Ano01s] Anonymous. Pages 111-224 (1 December 2001). Journal of Statistical Planning and Inference, 99(2):??, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PS

[Ano01t] Anonymous. Pages 201-416 (1 September 2001). Journal of Statistical Planning and Inference, 97(2):??, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:PJb

[Ano01u] Anonymous. Pages 301-440 (1 July 2001). Journal of Statistical Planning and Inference, 96(2):??, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:Pa

[Ano01v] Anonymous. Preface. Journal of Statistical Planning and Inference, 94(2):v, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801000726.

Anonymous:2001:Pb

[Ano01w] Anonymous. Preface. Journal of Statistical Planning and Inference, 95(1–2):1, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002743.

Anonymous:2001:PNa

[Ano01x] Anonymous. Publisher’s note. Journal of Statistical Planning and Inference, 97(2):iii, September 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002270.

Anonymous:2001:S

[Ano01y] Anonymous. (shen). Journal of Statistical Planning and Inference, 94(2):??, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2001:T

[Ano01z] Anonymous. (Tonchev). Journal of Statistical Planning and Inference, 95(1–2):??, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:AI

[Ano02a] Anonymous. Author index. Journal of Statistical Planning and Inference, 107(1–2):377–378, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003786.

Anonymous:2002:AIVc

[Ano02b] Anonymous. Author index to volume 102. Journal of Statistical Planning and Inference, 102(2):501–502, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802000745.

Anonymous:2002:AIVd

[Ano02c] Anonymous. Author index to volume 103. Journal of Statistical Planning and Inference, 103(1–2):471–472, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802000861.

Anonymous:2002:AIVe

[Ano02d] Anonymous. Author index to volume 104. Journal of Statistical Planning and Inference, 104(2):503–504, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001295.

Anonymous:2002:AIVh

[Ano02e] Anonymous. Author index to volume 104. Journal of Statistical Planning and Inference, 108(1–2):411–412, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004305.

Anonymous:2002:AIVf

[Ano02f] Anonymous. Author index to volume 105. Journal of Statistical Planning and Inference, 105(2):449–450, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200191X.

Anonymous:2002:AIVg

[Ano02g] Anonymous. Author index to volume 106. Journal of Statistical Planning and Inference, 106(1–2):517–519, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003452.

Anonymous:2002:AIVa

[Ano02h] Anonymous. Author index volume 100 (2002). Journal of Statistical Planning and Inference, 100(2):445–446, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801003160.

Anonymous:2002:AIVb

[Ano02i] Anonymous. Author index volume 101 (2002). Journal of Statistical Planning and Inference, 101(1–2):327–328, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801003226.

Anonymous:2002:Ca

[Ano02j] Anonymous. Contents. Journal of Statistical Planning and Inference, 102(1):v, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802000733.

Anonymous:2002:Cb

[Ano02k] Anonymous. Contents. Journal of Statistical Planning and Inference, 102(2):ix, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802000782.

Anonymous:2002:Cc

[Ano02l] Anonymous. Contents. Journal of Statistical Planning and Inference, 103(1–2):v–vi, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200085X.

Anonymous:2002:Cd

[Ano02m] Anonymous. Contents. Journal of Statistical Planning and Inference, 104(2):iii, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802001258.

Anonymous:2002:Ce

[Ano02n] Anonymous. Contents. Journal of Statistical Planning and Inference, 105(1):v, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802001386.

Anonymous:2002:Cf

[Ano02o] Anonymous. Contents. Journal of Statistical Planning and Inference, 106(1–2):v–vi, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003440.

Anonymous:2002:Cg

[Ano02p] Anonymous. Contents. Journal of Statistical Planning and Inference, 107(1–2):v–vi, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003774.

Anonymous:2002:Ch

[Ano02q] Anonymous. Contents. Journal of Statistical Planning and Inference, 108(1–2):v–vi, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004299.

Anonymous:2002:Ia

[Ano02r] Anonymous. IBC. Journal of Statistical Planning and Inference, 107(1–2):ibc, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003804.

Anonymous:2002:Ib

[Ano02s] Anonymous. IBC. Journal of Statistical Planning and Inference, 108(1–2):ibc, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802004329.

Anonymous:2002:Oa

[Ano02t] Anonymous. OBC. Journal of Statistical Planning and Inference, 107(1–2):obc, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003816.

Anonymous:2002:Ob

[Ano02u] Anonymous. OBC. Journal of Statistical Planning and Inference, 108(1–2):obc, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004330.

Anonymous:2002:PMa

[Ano02v] Anonymous. Pages 1-210 (1 March 2002). Journal of Statistical Planning and Inference, 102(1):??, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PMb

[Ano02w] Anonymous. Pages 1-238 (1 May 2002). Journal of Statistical Planning and Inference, 104(1):??, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PFb

[Ano02x] Anonymous. Pages 1-328 (15 February 2002). Journal of Statistical Planning and Inference, 101(1–2):??, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PS

[Ano02y] Anonymous. Pages 1-378 (1 September 2002). Journal of Statistical Planning and Inference, 107(1–2):??, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PAb

[Ano02z] Anonymous. Pages 1-472 (15 April 2002). Journal of Statistical Planning and Inference, 103(1–2):??, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PJa

[Ano02-27] Anonymous. Pages 1-90 (1 January 2002). Journal of Statistical Planning and Inference, 100(1):??, January 1, 2002.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PAa

[Ano02-28] Anonymous. Pages 211-502 (1 April 2002). Journal of Statistical Planning and Inference, 102(2):??, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PJb

[Ano02-29] Anonymous. Pages 239-504 (1 June 2002). Journal of Statistical Planning and Inference, 104(2):??, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PJc

[Ano02-30] Anonymous. Pages 283-450 (1 July 2002). Journal of Statistical Planning and Inference, 105(2):??, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:PFa

[Ano02-31] Anonymous. Pages 91-446 (1 February 2002). Journal of Statistical Planning and Inference, 100(2):??, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2002:RI

[Ano02-32] Anonymous. (Rao II). Journal of Statistical Planning and Inference, 108(1–2):??, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:AR

[Ano03a] Anonymous. Acknowledgement to the referees of 2001 and 2002. Journal of Statistical Planning and Inference, 115 (1):vii–xxii, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803001307.

Anonymous:2003:AIVh

[Ano03b] Anonymous. Author index to volume 104. Journal of Statistical Planning and Inference, 116(2):537–538, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001824.

Anonymous:2003:AIVa

[Ano03c] Anonymous. Author index to volume 109. Journal of Statistical Planning and Inference, 109(1–2):189, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200441X.

Anonymous:2003:AIVb

[Ano03d] Anonymous. Author index to volume 110. Journal of Statistical Planning and Inference, 110(1–2):179, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004536.

Anonymous:2003:AIVd

[Ano03e] Anonymous. Author index to volume 112. Journal of Statistical Planning and Inference, 112(1–2):259, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802005165.

Anonymous:2003:AIVe

[Ano03f] Anonymous. Author index to volume 113. Journal of Statistical Planning and Inference, 113(2):715–717, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803000892.

Anonymous:2003:AIVf

[Ano03g] Anonymous. Author index to volume 114. Journal of Statistical Planning and Inference, 114(1–2):187, June 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001046.

Anonymous:2003:AIVg

[Ano03h] Anonymous. Author index to volume 115. Journal of Statistical Planning and Inference, 115(2):699–701, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001630.

Anonymous:2003:AIVi

[Ano03i] Anonymous. Author index to volume 117. Journal of Statistical Planning and Inference, 117(2):387–388, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580300257X.

Anonymous:2003:AIVc

[Ano03j] Anonymous. Author index volume 111. Journal of Statistical Planning and Inference, 111(1–2):249–250, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004780.

Anonymous:2003:Ca

[Ano03k] Anonymous. Contents. Journal of Statistical Planning and Inference, 109(1–2):v, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004408.

Anonymous:2003:Cb

[Ano03l] Anonymous. Contents. Journal of Statistical Planning and Inference, 111(1–2):v, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004779.

Anonymous:2003:Cc

[Ano03m] Anonymous. Contents. Journal of Statistical Planning and Inference, 112(1–2):v, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802005153.

Anonymous:2003:Cd

[Ano03n] Anonymous. Contents. Journal of Statistical Planning and Inference, 113(1):v–vi, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803000806.

Anonymous:2003:Ce

[Ano03o] Anonymous. Contents. Journal of Statistical Planning and Inference, 113(2):iii, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803000880.

Anonymous:2003:Cf

[Ano03p] Anonymous. Contents. Journal of Statistical Planning and Inference, 114(1–2):v, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001034.

Anonymous:2003:Cg

[Ano03q] Anonymous. Contents. Journal of Statistical Planning and Inference, 115(1):v, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001277.

Anonymous:2003:Ch

[Ano03r] Anonymous. Contents. Journal of Statistical Planning and Inference, 115(2):iii–iv, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001629.

Anonymous:2003:Ci

[Ano03s] Anonymous. Contents. Journal of Statistical Planning and Inference, 116(1):v, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001721.

Anonymous:2003:Cj

[Ano03t] Anonymous. Contents. Journal of Statistical Planning and Inference, 116(2):iii, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001812.

Anonymous:2003:Ck

[Ano03u] Anonymous. Contents. Journal of Statistical Planning and Inference, 117(1):v, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002222.

Anonymous:2003:Cl

[Ano03v] Anonymous. Contents. Journal of Statistical Planning and Inference, 117(2):iii, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002532.

Anonymous:2003:E

[Ano03w] Anonymous. (Editors). Journal of Statistical Planning and Inference, 116(1):obc, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001745.

Anonymous:2003:EC

[Ano03x] Anonymous. (Editors cont.). Journal of Statistical Planning and Inference, 116(1):ibc, September 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001733.

Anonymous:2003:Ia

[Ano03y] Anonymous. IBC. Journal of Statistical Planning and Inference, 110(1–2):ibc, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200455X.

Anonymous:2003:Ib

[Ano03z] Anonymous. IBC. Journal of Statistical Planning and Inference, 111(1–2):ibc, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802004809.

Anonymous:2003:Ic

[Ano03-27] Anonymous. IBC. Journal of Statistical Planning and Inference, 112(1–2):ibc, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802005189.

Anonymous:2003:IECa

[Ano03-28] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 109(1–2):ibc, January 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004433.

Anonymous:2003:IECb

[Ano03-29] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 113(1):ibc, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803000818.

Anonymous:2003:IECc

[Ano03-30] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 113(2):ibc, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803000910.

Anonymous:2003:IECd

[Ano03-31] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 114(1–2):ibc, June 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300106X.

Anonymous:2003:IECe

[Ano03-32] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 115(1):ibc, July 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001289.

Anonymous:2003:IECf

[Ano03-33] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 115(2):ibc, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001654.

Anonymous:2003:IECg

[Ano03-34] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 116(2):ibc, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001848.

Anonymous:2003:IECh

[Ano03-35] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 117(1):ibc, November 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002209.

Anonymous:2003:IECi

[Ano03-36] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 117(2):ibc, December 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002556.

Anonymous:2003:Oa

[Ano03-37] Anonymous. OBC. Journal of Statistical Planning and Inference, 110(1–2):obc, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004561.

Anonymous:2003:Ob

[Ano03-38] Anonymous. OBC. Journal of Statistical Planning and Inference, 111(1–2):obc, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004810.

Anonymous:2003:Oc

[Ano03-39] Anonymous. OBC. Journal of Statistical Planning and Inference, 112(1–2):obc, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802005190.

Anonymous:2003:OEa

[Ano03-40] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 109(1–2):obc, January 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004445.

Anonymous:2003:OEb

[Ano03-41] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 113(1):obc, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580300082X.

Anonymous:2003:OEc

[Ano03-42] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 113(2):obc, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803000922.

Anonymous:2003:OEd

[Ano03-43] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 114(1–2):obc, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001071.

Anonymous:2003:OEe

[Ano03-44] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 115(1):obc, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001290.

Anonymous:2003:OEf

[Ano03-45] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 115(2):obc, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001666.

Anonymous:2003:OEg

[Ano03-46] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 116(2):obc, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300185X.

Anonymous:2003:OEh

[Ano03-47] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 117(1):obc, November 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002210.

Anonymous:2003:OEi

[Ano03-48] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 117(2):obc, December 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002568.

Anonymous:2003:PNa

[Ano03-49] Anonymous. Pages 1-166 (1 November 2003). Journal of Statistical Planning and Inference, 117(1):??, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PJ

[Ano03-50] Anonymous. Pages 1-180 (15 January 2003). Journal of Statistical Planning and Inference, 110(1–2):??, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PAa

[Ano03-51] Anonymous. Pages 1-374 (1 April 2003). Journal of Statistical Planning and Inference, 113(1):??, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PD

[Ano03-52] Anonymous. Pages 167-388 (1 December 2003). Journal of Statistical Planning and Inference, 117(2):??, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PO

[Ano03-53] Anonymous. Pages 333-538 (1 October 2003). Journal of Statistical Planning and Inference, 116(2):??, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PAb

[Ano03-54] Anonymous. Pages 349-702 (1 August 2003). Journal of Statistical Planning and Inference, 115(2):??, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PEEc

[Ano03-55] Anonymous. Pages EX1-EX2, 1-332 (1 September 2003). Journal of Statistical Planning and Inference, 116(1):??, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PEEb

[Ano03-56] Anonymous. Pages EX1-EX2, 1-348 (1 July 2003). Journal of Statistical Planning and Inference, 115(1):??, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PEEa

[Ano03-57] Anonymous. Pages EX1-EX2, 375-718 (1 May 2003). Journal of Statistical Planning and Inference, 113(2):??, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2003:PNb

[Ano03-58] Anonymous. Publisher’s note. Journal of Statistical Planning and Inference, 117(2):167, December 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002581.

Anonymous:2003:RI

[Ano03-59] Anonymous. (Rao III). Journal of Statistical Planning and Inference, 109(1–2):??, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:AIVa

[Ano04a] Anonymous. Author index to volume 104. Journal of Statistical Planning and Inference, 118(1–2):221, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300301X.

Anonymous:2004:AIVg

[Ano04b] Anonymous. Author index to volume 104. Journal of Statistical Planning and Inference, 124(2):489–490, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001326.

Anonymous:2004:AIVb

[Ano04c] Anonymous. Author index to volume 119. Journal of Statistical Planning and Inference, 119(2):449–450, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580300346X.

Anonymous:2004:AIVc

[Ano04d] Anonymous. Author index to volume 120. Journal of Statistical Planning and Inference, 120(1–2):167, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803003574.

Anonymous:2004:AIVd

[Ano04e] Anonymous. Author index to volume 121. Journal of Statistical Planning and Inference, 121(2):339–340, April 1, 2004.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000163.

Anonymous:2004:AIVe

[Ano04f] Anonymous. Author index to volume 122. Journal of Statistical Planning and Inference, 122(1–2):293–294, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580400045X.

Anonymous:2004:AIVf

[Ano04g] Anonymous. Author index to volume 123. Journal of Statistical Planning and Inference, 123(2):415–416, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000928.

Anonymous:2004:AIVh

[Ano04h] Anonymous. Author index to volume 125. Journal of Statistical Planning and Inference, 125(1–2):139, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002897.

Anonymous:2004:AIVi

[Ano04i] Anonymous. Author index to volume 126. Journal of Statistical Planning and Inference, 126(2):591–593, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003295.

Anonymous:2004:Ca

[Ano04j] Anonymous. Contents. Journal of Statistical Planning and Inference, 118(1–2):v, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003008.

Anonymous:2004:Cb

[Ano04k] Anonymous. Contents. Journal of Statistical Planning and Inference, 119(1):v, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003380.

Anonymous:2004:Cc

[Ano04l] Anonymous. Contents. Journal of Statistical Planning and Inference, 119(2):iii, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003458.

Anonymous:2004:Cd

[Ano04m] Anonymous. Contents. Journal of Statistical Planning and Inference, 120(1–2):v, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003562.

Anonymous:2004:Ce

[Ano04n] Anonymous. Contents. Journal of Statistical Planning and Inference, 121(1):v, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003847.

Anonymous:2004:Cf

[Ano04o] Anonymous. Contents. Journal of Statistical Planning and Inference, 121(2):iii, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580400014X.

Anonymous:2004:Cg

[Ano04p] Anonymous. Contents. Journal of Statistical Planning and Inference, 122(1–2):v–vi, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000448.

Anonymous:2004:Ch

[Ano04q] Anonymous. Contents. Journal of Statistical Planning and Inference, 123(1):v, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000734.

Anonymous:2004:Ci

[Ano04r] Anonymous. Contents. Journal of Statistical Planning and Inference, 123(2):iii, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000916.

Anonymous:2004:Cj

[Ano04s] Anonymous. Contents. Journal of Statistical Planning and Inference, 124(1):v, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001028.

Anonymous:2004:Ck

[Ano04t] Anonymous. Contents. Journal of Statistical Planning and Inference, 124(2):iii, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001314.

Anonymous:2004:Cl

[Ano04u] Anonymous. Contents. Journal of Statistical Planning and Inference, 125(1–2):v, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002885.

Anonymous:2004:Cm

[Ano04v] Anonymous. Contents. Journal of Statistical Planning and Inference, 126(1):v–vi, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580400312X.

Anonymous:2004:Cn

[Ano04w] Anonymous. Contents. Journal of Statistical Planning and Inference, 126(2):iii, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003283.

Anonymous:2004:IECa

[Ano04x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 118(1–2):ibc, January 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003033.

Anonymous:2004:IECb

[Ano04y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 119(1):ibc, January 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003392.

Anonymous:2004:IECc

[Ano04z] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 119(2):ibc, February 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003483.

Anonymous:2004:IECd

[Ano04-27] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 120(1–2):ibc, February 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003598.

Anonymous:2004:IECe

[Ano04-28] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 121(1):ibc, March 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003859.

Anonymous:2004:IECf

[Ano04-29] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 121(2):ibc, April 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000187.

Anonymous:2004:IECg

[Ano04-30] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 122(1–2):ibc, ???? 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000473.

Anonymous:2004:IECh

[Ano04-31] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 123(1):ibc, June 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000746.

Anonymous:2004:IECi

[Ano04-32] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 123(2):ibc, July 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000941.

Anonymous:2004:IECj

[Ano04-33] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 124(1):ibc, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400103X.

Anonymous:2004:IECk

[Ano04-34] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 124(2):ibc, September 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400134X.

Anonymous:2004:IECl

[Ano04-35] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 125(1–2):ibc, October 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002915.

Anonymous:2004:IECm

[Ano04-36] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 126(1):ibc, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003131.

Anonymous:2004:IECn

[Ano04-37] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 126(2):ibc, December 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003313.

Anonymous:2004:OEa

[Ano04-38] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 118(1–2):obc, January 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003045.

Anonymous:2004:OEb

[Ano04-39] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 119(1):obc, January 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003409.

Anonymous:2004:OEc

[Ano04-40] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 119(2):obc, February 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003495.

Anonymous:2004:OEd

[Ano04-41] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 120(1–2):obc, February 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003604.

Anonymous:2004:OEe

[Ano04-42] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 121(1):obc, March 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003860.

Anonymous:2004:OEf

[Ano04-43] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 121(2):obc, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000199.

Anonymous:2004:OEg

[Ano04-44] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 122(1–2):obc, ???? 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000485.

Anonymous:2004:OEh

[Ano04-45] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 123(1):obc, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000758.

Anonymous:2004:OEi

[Ano04-46] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 123(2):obc, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000953.

Anonymous:2004:OEj

[Ano04-47] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 124(1):obc, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001041.

Anonymous:2004:OEk

[Ano04-48] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 124(2):obc, September 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001351.

Anonymous:2004:OEl

[Ano04-49] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 125(1–2):obc, October 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002927.

Anonymous:2004:OEm

[Ano04-50] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 126(1):obc, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003143.

Anonymous:2004:OEn

[Ano04-51] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 126(2):obc, December 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003325.

Anonymous:2004:PM

[Ano04-52] Anonymous. Pages 1-162 (1 March 2004). Journal of Statistical Planning and Inference, 121(1):??, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PFb

[Ano04-53] Anonymous. Pages 1-168 (15 February 2004). Journal of Statistical Planning and Inference, 120(1–2):??, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PJb

[Ano04-54] Anonymous. Pages 1-206 (1 June 2004). Journal of Statistical Planning and Inference, 123(1):??, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PJa

[Ano04-55] Anonymous. Pages 1-224 (15 January 2004). Journal of Statistical Planning and Inference, 119(1):??, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PAb

[Ano04-56] Anonymous. Pages 1-250 (1 August 2004). Journal of Statistical Planning and Inference, 124(1):??, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PAa

[Ano04-57] Anonymous. Pages 163-340 (1 April 2004). Journal of Statistical Planning and Inference, 121(2):??, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PJc

[Ano04-58] Anonymous. Pages 207-416 (1 July 2004). Journal of Statistical Planning and Inference, 123(2):??, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PFa

[Ano04-59] Anonymous. Pages 225-450 (1 February 2004). Journal of Statistical Planning and Inference, 119(2):??, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PD

[Ano04-60] Anonymous. Pages 401-594 (1 December 2004). Journal of Statistical Planning and Inference, 126(2):??, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PEEa

[Ano04-61] Anonymous. Pages EX1-EX2, 1-222 (1 January 2004). Journal of Statistical Planning and Inference, 118(1–2):??, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PEEc

[Ano04-62] Anonymous. Pages EX1-EX2, 1-400 (1 November 2004). Journal of Statistical Planning and Inference, 126(1):??, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2004:PEEb

[Ano04-63] Anonymous. Pages EX1-EX2, 251-490 (1 September 2004). Journal of Statistical Planning and Inference, 124(2):??, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:SPW

[Ano05a] Anonymous. (4th St. Petersburg Workshop). Journal of Statistical Planning and Inference, 132(1–2):??, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:AIVa

[Ano05b] Anonymous. Author index to volume 127. Journal of Statistical Planning and Inference, 127(1–2):347–348, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003738.

Anonymous:2005:AIVb

[Ano05c] Anonymous. Author index to volume 128. Journal of Statistical Planning and Inference, 128(2):661–663, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004069.

Anonymous:2005:AIVc

[Ano05d] Anonymous. Author index to volume 129. Journal of Statistical Planning and Inference, 129(1–2):447–448, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004343.

Anonymous:2005:AIVd

[Ano05e] Anonymous. Author index to volume 130. Journal of Statistical Planning and Inference, 130(1–2):413–414, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004501.

Anonymous:2005:AIVf

[Ano05f] Anonymous. Author index to volume 132. Journal of Statistical Planning and Inference, 132(1–2):203, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000467.

Anonymous:2005:AIVg

[Ano05g] Anonymous. Author index to volume 133. Journal of Statistical Planning and Inference, 133(2):561–562, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001114.

Anonymous:2005:AIVh

[Ano05h] Anonymous. Author index to volume 134. Journal of Statistical Planning and Inference, 134(2):604–606, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001473.

Anonymous:2005:AIVi

[Ano05i] Anonymous. Author index to volume 135. Journal of Statistical Planning and Inference, 135(2):528–529, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805002363.

Anonymous:2005:AIVe

[Ano05j] Anonymous. Author index to Volume XXX. Journal of Statistical Planning and Inference, 131(2):417–418, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000224.

Anonymous:2005:Ca

[Ano05k] Anonymous. Contents. Journal of Statistical Planning and Inference, 127(1–2):v–vi, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003726.

Anonymous:2005:Cb

[Ano05l] Anonymous. Contents. Journal of Statistical Planning and Inference, 128(1):v–vi, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003866.

Anonymous:2005:Cc

[Ano05m] Anonymous. Contents. Journal of Statistical Planning and Inference, 128(2):iii–iv, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004057.

Anonymous:2005:Cd

[Ano05n] Anonymous. Contents. Journal of Statistical Planning and Inference, 129(1–2):v–vi, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004331.

Anonymous:2005:Ce

[Ano05o] Anonymous. Contents. Journal of Statistical Planning and Inference, 130(1–2):vii–viii, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004495.

Anonymous:2005:Cf

[Ano05p] Anonymous. Contents. Journal of Statistical Planning and Inference, 131(1):v–vi, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580500008X.

Anonymous:2005:Cg

[Ano05q] Anonymous. Contents. Journal of Statistical Planning and Inference, 131(2):iii, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000212.

Anonymous:2005:Ch

[Ano05r] Anonymous. Contents. Journal of Statistical Planning and Inference, 132(1–2):v, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000455.

Anonymous:2005:Ci

[Ano05s] Anonymous. Contents. Journal of Statistical Planning and Inference, 133(1):v, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000996.

Anonymous:2005:Cj

[Ano05t] Anonymous. Contents. Journal of Statistical Planning and Inference, 133(2):iii–iv, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001102.

Anonymous:2005:Ck

[Ano05u] Anonymous. Contents. Journal of Statistical Planning and Inference, 134(1):v–vi, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001242.

Anonymous:2005:Cl

[Ano05v] Anonymous. Contents. Journal of Statistical Planning and Inference, 134(2):iii–iv, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001461.

Anonymous:2005:Cm

[Ano05w] Anonymous. Contents. Journal of Statistical Planning and Inference, 135(1):v–vi, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001631.

Anonymous:2005:Cn

[Ano05x] Anonymous. Contents. Journal of Statistical Planning and Inference, 135(2):iii–iv, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002351.

Anonymous:2005:IECa

[Ano05y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 127(1–2):ibc, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003751.

Anonymous:2005:IECb

[Ano05z] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 128(1):ibc, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003878.

Anonymous:2005:IECc

[Ano05-27] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 128(2):ibc, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004082.

Anonymous:2005:IECd

[Ano05-28] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 129(1–2):ibc, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004367.

Anonymous:2005:IECe

[Ano05-29] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 130(1–2):ibc, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004525.

Anonymous:2005:IECf

[Ano05-30] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 131(1):ibc, April 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000091.

Anonymous:2005:IECg

[Ano05-31] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 131(2):ibc, May 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000248.

Anonymous:2005:IECh

[Ano05-32] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 132(1–2):ibc, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000480.

Anonymous:2005:IECi

[Ano05-33] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 133(1):CO3, July 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580500100X.

Anonymous:2005:IECj

[Ano05-34] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 133(2):CO3, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001138.

Anonymous:2005:IECk

[Ano05-35] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 134(1):CO3, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001229.

Anonymous:2005:IECl

[Ano05-36] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 134(2):CO3, October 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001497.

Anonymous:2005:IECm

[Ano05-37] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 135(1):CO3, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001643.

Anonymous:2005:IECn

[Ano05-38] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 135(2):CO3, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002387.

Anonymous:2005:I

[Ano05-39] Anonymous. (IISA 2002). Journal of Statistical Planning and Inference, 129(1–2):??, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:LPC

[Ano05-40] Anonymous. Lattice path combinatorics and discrete distributions. Journal of Statistical Planning and Inference, 135 (1):??, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:OEa

[Ano05-41] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 127(1–2):obc, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003763.

Anonymous:2005:OEb

[Ano05-42] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 128(1):obc, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400388X.

Anonymous:2005:OEc

[Ano05-43] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 128(2):obc, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004094.

Anonymous:2005:OEd

[Ano05-44] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 129(1–2):obc, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004379.

Anonymous:2005:OEe

[Ano05-45] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 130(1–2):obc, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004537.

Anonymous:2005:OEf

[Ano05-46] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 131(1):obc, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000108.

Anonymous:2005:OEg

[Ano05-47] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 131(2):obc, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580500025X.

Anonymous:2005:OEh

[Ano05-48] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 132(1–2):obc, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000492.

Anonymous:2005:OEi

[Ano05-49] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 133(1):CO4, July 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001011.

Anonymous:2005:OEj

[Ano05-50] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 133(2):CO4, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580500114X.

Anonymous:2005:OEk

[Ano05-51] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 134(1):obc, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001230.

Anonymous:2005:OEl

[Ano05-52] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 134(2):CO4, October 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001503.

Anonymous:2005:OEm

[Ano05-53] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 135(1):CO4, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001655.

Anonymous:2005:OEn

[Ano05-54] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 135(2):CO4, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002399.

Anonymous:2005:PJc

[Ano05-55] Anonymous. Pages 1-222 (1 July 2005). Journal of Statistical Planning and Inference, 133(1):??, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PS

[Ano05-56] Anonymous. Pages 1-302 (1 September 2005). Journal of Statistical Planning and Inference, 134(1):??, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PJb

[Ano05-57] Anonymous. Pages 1-326 (15 January 2005). Journal of Statistical Planning and Inference, 128(1):??, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PJa

[Ano05-58] Anonymous. Pages 1-348 (1 January 2005). Journal of Statistical Planning and Inference, 127(1–2):??, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PM

[Ano05-59] Anonymous. Pages 209-418 (1 May 2005). Journal of Statistical Planning and Inference, 131(2):??, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PA

[Ano05-60] Anonymous. Pages 223-562 (1 August 2005). Journal of Statistical Planning and Inference, 133(2):??, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PD

[Ano05-61] Anonymous. Pages 245-530 (1 December 2005). Journal of Statistical Planning and Inference, 135(2):??, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PO

[Ano05-62] Anonymous. Pages 303-606 (1 October 2005). Journal of Statistical Planning and Inference, 134(2):??, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PF

[Ano05-63] Anonymous. Pages 327-664 (1 February 2005). Journal of Statistical Planning and Inference, 128(2):??, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PEE

[Ano05-64] Anonymous. Pages EX1-EX2, 1-208 (1 April 2005). Journal of Statistical Planning and Inference, 131(1):??, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2005:PHC

[Ano05-65] Anonymous. Publications of Herman Chernoff. Journal of Statistical Planning and Inference, 130(1–2):13–19, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002770.

Anonymous:2006:AIV

[Ano06a] Anonymous. Author index to Volume XXX. Journal of Statistical Planning and Inference, 136(12):4459–4473, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001893.

Anonymous:2006:CRT

[Ano06b] Anonymous. Competing risks: Theory and applications. Journal of Statistical Planning and Inference, 136(5):??, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:Ca

[Ano06c] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(1):v–vi, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002533.

Anonymous:2006:Cb

[Ano06d] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(2):iii, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002600.

Anonymous:2006:Cc

[Ano06e] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(3):iii–v, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002740.

Anonymous:2006:Cd

[Ano06f] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(4):iii–iv, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002879.

Anonymous:2006:Ce

[Ano06g] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(5):iii, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806000061.

Anonymous:2006:Cf

[Ano06h] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(6):iii–iv, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000152.

Anonymous:2006:Cg

[Ano06i] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(7):iii–v, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000486.

Anonymous:2006:Ch

[Ano06j] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(8):iii–iv, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000693.

Anonymous:2006:Ci

[Ano06k] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(9):iii–iv, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000966.

Anonymous:2006:Cj

[Ano06l] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(10):iii–iv, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001248.

Anonymous:2006:Ck

[Ano06m] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(11):iii–iv, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001686.

Anonymous:2006:Cl

[Ano06n] Anonymous. Contents. Journal of Statistical Planning and Inference, 136(12):iii–iv, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001881.

Anonymous:2006:IECa

[Ano06o] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(1):CO3, January 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002545.

Anonymous:2006:IECb

[Ano06p] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(2):CO3, February 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002612.

Anonymous:2006:IECc

[Ano06q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(3):CO3, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002752.

Anonymous:2006:IECd

[Ano06r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(4):CO3, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002880.

Anonymous:2006:IECe

[Ano06s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(5):CO3, May 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000073.

Anonymous:2006:IECf

[Ano06t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(6):CO3, June 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000164.

Anonymous:2006:IECg

[Ano06u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(7):CO3, July 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000498.

Anonymous:2006:IECh

[Ano06v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(8):CO3, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580600070X.

Anonymous:2006:IECi

[Ano06w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(9):CO3, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000942.

Anonymous:2006:IECj

[Ano06x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(10):CO3, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580600125X.

Anonymous:2006:IECk

[Ano06y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(11):CO3, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001698.

Anonymous:2006:IECl

[Ano06z] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 136(12):CO3, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001911.

Anonymous:2006:OEa

[Ano06-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(1):CO4, January 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002557.

Anonymous:2006:OEb

[Ano06-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(2):CO4, February 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002624.

Anonymous:2006:OEc

[Ano06-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(3):CO4, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002764.

Anonymous:2006:OEd

[Ano06-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(4):CO4, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002892.

Anonymous:2006:OEe

[Ano06-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(5):CO4, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000085.

Anonymous:2006:OEf

[Ano06-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(6):CO4, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000176.

Anonymous:2006:OEg

[Ano06-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(7):CO4, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000504.

Anonymous:2006:OEh

[Ano06-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(8):CO4, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000711.

Anonymous:2006:OEi

[Ano06-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(9):CO4, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000954.

Anonymous:2006:OEj

[Ano06-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(10):CO4, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001261.

Anonymous:2006:OEk

[Ano06-37] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(11):CO4, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001704.

Anonymous:2006:OEl

[Ano06-38] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 136(12):CO4, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001923.

Anonymous:2006:PJ

[Ano06-39] Anonymous. Pages 1-310 (1 January 2006). Journal of Statistical Planning and Inference, 136(1):??, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PAa

[Ano06-40] Anonymous. Pages 1197-1568 (1 April 2006). Journal of Statistical Planning and Inference, 136(4):??, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PAb

[Ano06-41] Anonymous. Pages 2407-2846 (1 August 2006). Journal of Statistical Planning and Inference, 136(8):??, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PS

[Ano06-42] Anonymous. Pages 2847-3318 (1 September 2006). Journal of Statistical Planning and Inference, 136(9):??, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PF

[Ano06-43] Anonymous. Pages 311-526 (1 February 2006). Journal of Statistical Planning and Inference, 136(2):??, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PO

[Ano06-44] Anonymous. Pages 3319-3734 (1 October 2006). Journal of Statistical Planning and Inference, 136(10):??, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PN

[Ano06-45] Anonymous. Pages 3735-4118 (1 November 2006). Journal of Statistical Planning and Inference, 136(11):??, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PD

[Ano06-46] Anonymous. Pages 4119-4474 (1 December 2006). Journal of Statistical Planning and Inference, 136(12):??, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2006:PM

[Ano06-47] Anonymous. Pages 527-1196 (1 March 2006). Journal of Statistical Planning and Inference, 136(3):??, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:Aa

[Ano07a] Anonymous. Announcement. Journal of Statistical Planning and Inference, 137(4):I, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002989.

Anonymous:2007:Ab

[Ano07b] Anonymous. Announcement. Journal of Statistical Planning and Inference, 137(6):I, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000328.

Anonymous:2007:Ca

[Ano07c] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(1):v–vi, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002151.

Anonymous:2007:Cb

[Ano07d] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(2):iii–iv, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002357.

Anonymous:2007:Cc

[Ano07e] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(3):iii–v, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002667.

Anonymous:2007:Cd

[Ano07f] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(4):iii–iv, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002953.

Anonymous:2007:Ce

[Ano07g] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(5):iii–iv, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000067.

Anonymous:2007:Cf

[Ano07h] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(6):iii–iv, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000298.

Anonymous:2007:Cg

[Ano07i] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(7):iii–iv, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000687.

Anonymous:2007:Ch

[Ano07j] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(8):iii, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001474.

Anonymous:2007:Ci

[Ano07k] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(9):iii–iv, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001784.

Anonymous:2007:Cj

[Ano07l] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(10):iii, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700239X.

Anonymous:2007:Ck

[Ano07m] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(11):iii–v, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807002844.

Anonymous:2007:Cl

[Ano07n] Anonymous. Contents. Journal of Statistical Planning and Inference, 137(12):iii–iv, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003059.

Anonymous:2007:EB

[Ano07o] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 137(6):CO4, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000316.

Anonymous:2007:EBC

[Ano07p] Anonymous. Editorial board continued. Journal of Statistical Planning and Inference, 137(6):CO3, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000304.

Anonymous:2007:IECa

[Ano07q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(1):CO3, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002163.

Anonymous:2007:IECb

[Ano07r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(2):ibc, February 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002369.

Anonymous:2007:IECc

[Ano07s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(3):CO3, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002679.

Anonymous:2007:IECd

[Ano07t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(4):CO3, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002965.

Anonymous:2007:IECe

[Ano07u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(5):CO3, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000079.

Anonymous:2007:IECf

[Ano07v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(7):ibc, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000699.

Anonymous:2007:IECg

[Ano07w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(8):ibc, August 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001486.

Anonymous:2007:IECh

[Ano07x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(9):ibc, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001796.

Anonymous:2007:IECi

[Ano07y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(10):ibc, October 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002406.

Anonymous:2007:IECj

[Ano07z] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(11):ibc, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002856.

Anonymous:2007:IECk

[Ano07-27] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 137(12):ibc, December 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003060.

Anonymous:2007:OEa

[Ano07-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(1):CO4, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002175.

Anonymous:2007:OEb

[Ano07-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(2):obc, February 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002370.

Anonymous:2007:OEc

[Ano07-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(3):CO4, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002680.

Anonymous:2007:OEd

[Ano07-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(4):CO4, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002977.

Anonymous:2007:OEe

[Ano07-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(5):CO4, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000080.

Anonymous:2007:OEf

[Ano07-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(7):obc, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000705.

Anonymous:2007:OEg

[Ano07-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(8):obc, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001498.

Anonymous:2007:OEh

[Ano07-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(9):obc, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001802.

Anonymous:2007:OEi

[Ano07-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(10):obc, October 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002418.

Anonymous:2007:OEj

[Ano07-37] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(11):obc, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002868.

Anonymous:2007:OEk

[Ano07-38] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 137(12):obc, December 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003072.

Anonymous:2007:PJa

[Ano07-39] Anonymous. Pages 1-362 (1 January 2007). Journal of Statistical Planning and Inference, 137(1):??, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:PA

[Ano07-40] Anonymous. Pages 1099-1514 (1 April 2007). Journal of Statistical Planning and Inference, 137(4):??, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:PJb

[Ano07-41] Anonymous. Pages 1711-2088 (1 June 2007). Journal of Statistical Planning and Inference, 137(6):??, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:PJc

[Ano07-42] Anonymous. Pages 2089-2534 (1 July 2007). Journal of Statistical Planning and Inference, 137(7):??, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:PS

[Ano07-43] Anonymous. Pages 2733-3040 (1 September 2007). Journal of Statistical Planning and Inference, 137(9):??, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:PF

[Ano07-44] Anonymous. Pages 363-628 (1 February 2007). Journal of Statistical Planning and Inference, 137(2):??, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2007:SIN

[Ano07-45] Anonymous. Special issue on nonparametric statistics and related topics: In honor of M. L. Puri. Journal of Statistical Planning and Inference, 137(3):??, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:Ca

[Ano08a] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(1):v–vi, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003424.

Anonymous:2008:Cb

[Ano08b] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(2):iii–iv, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003539.

Anonymous:2008:Cc

[Ano08c] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(3):iii–iv, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003977.

Anonymous:2008:Cd

[Ano08d] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(4):iii–iv, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700417X.

Anonymous:2008:Ce

[Ano08e] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(5):iii–iv, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000098.

Anonymous:2008:Cf

[Ano08f] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(6):iii–iv, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000372.

Anonymous:2008:Cg

[Ano08g] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(7):iii–iv, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001122.

Anonymous:2008:Ch

[Ano08h] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(8):iii–iv, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001729.

Anonymous:2008:Ci

[Ano08i] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(9):iii–iv, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800219X.

Anonymous:2008:Cj

[Ano08j] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(10):iii–iv, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002735.

Anonymous:2008:Ck

[Ano08k] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(11):iii–iv, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003078.

Anonymous:2008:Cl

[Ano08l] Anonymous. Contents. Journal of Statistical Planning and Inference, 138(12):iii–v, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003200.

Anonymous:2008:IECa

[Ano08m] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(1):ibc, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003436.

Anonymous:2008:IECb

[Ano08n] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(2):ibc, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003540.

Anonymous:2008:IECc

[Ano08o] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(3):CO3, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003989.

Anonymous:2008:IECd

[Ano08p] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(4):CO3, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004181.

Anonymous:2008:IECe

[Ano08q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(5):CO3, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000104.

Anonymous:2008:IECf

[Ano08r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(6):ibc, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000384.

Anonymous:2008:IECg

[Ano08s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(7):ibc, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001134.

Anonymous:2008:IECh

[Ano08t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(8):ibc, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001730.

Anonymous:2008:IECi

[Ano08u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(9):ibc, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002206.

Anonymous:2008:IECj

[Ano08v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(10):ibc, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002747.

Anonymous:2008:IECk

[Ano08w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(11):ibc, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800308X.

Anonymous:2008:IECl

[Ano08x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 138(12):ibc, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003212.

Anonymous:2008:OEa

[Ano08y] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(1):obc, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003448.

Anonymous:2008:OEb

[Ano08z] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(2):obc, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003552.

Anonymous:2008:OEc

[Ano08-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(3):CO4, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003990.

Anonymous:2008:OEd

[Ano08-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(4):CO4, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004193.

Anonymous:2008:OEe

[Ano08-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(5):CO4, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000116.

Anonymous:2008:OEf

[Ano08-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(6):obc, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000396.

Anonymous:2008:OEg

[Ano08-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(7):obc, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001146.

Anonymous:2008:OEh

[Ano08-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(8):obc, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001742.

Anonymous:2008:OEi

[Ano08-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(9):obc, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002218.

Anonymous:2008:OEj

[Ano08-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(10):obc, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002759.

Anonymous:2008:OEk

[Ano08-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(11):obc, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003091.

Anonymous:2008:OEl

[Ano08-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 138(12):obc, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003224.

Anonymous:2008:PMb

[Ano08-37] Anonymous. Pages 1227-1520 (1 May 2008). Journal of Statistical Planning and Inference, 138(5):??, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PJa

[Ano08-38] Anonymous. Pages 1521-1918 (1 July 2008). Journal of Statistical Planning and Inference, 138(6):??, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PJb

[Ano08-39] Anonymous. Pages 1919-2252 (1 July 2008). Journal of Statistical Planning and Inference, 138(7):??, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PAb

[Ano08-40] Anonymous. Pages 2253-2592 (1 August 2008). Journal of Statistical Planning and Inference, 138(8):??, August 1, 2008.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PO

[Ano08-41] Anonymous. Pages 2847-3292 (1 October 2008). Journal of Statistical Planning and Inference, 138(10):??, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PD

[Ano08-42] Anonymous. Pages 3635-4198 (1 December 2008). Journal of Statistical Planning and Inference, 138(12):??, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PMa

[Ano08-43] Anonymous. Pages 539-816 (1 March 2008). Journal of Statistical Planning and Inference, 138(3):??, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2008:PAa

[Ano08-44] Anonymous. Pages 817-1226 (1 April 2008). Journal of Statistical Planning and Inference, 138(4):??, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:TCM

[Ano09a] Anonymous. The 8th Tartu conference on multivariate statistics and the 6th conference on multivariate distributions with fixed marginals. Journal of Statistical Planning and Inference, 139(11):3739, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001360.

Anonymous:2009:Ca

[Ano09b] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(1):iii, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003765.

Anonymous:2009:Cb

[Ano09c] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(2):iii–vi, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004084.

Anonymous:2009:Cc

[Ano09d] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(3):iii–vi, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800431X.

Anonymous:2009:Cd

[Ano09e] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(4):iii–iv, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800462X.

Anonymous:2009:Ce

[Ano09f] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(5):iii–iv, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000214.

Anonymous:2009:Cf

[Ano09g] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(6):iii–iv, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000433.

Anonymous:2009:Cg

[Ano09h] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(7):iii–v, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000639.

Anonymous:2009:Ch

[Ano09i] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(8):iii–iv, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001128.

Anonymous:2009:Ci

[Ano09j] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(9):iii–v, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001748.

Anonymous:2009:Cj

[Ano09k] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(10):iii–iv, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580900202X.

Anonymous:2009:Ck

[Ano09l] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(11):iii–iv, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809002390.

Anonymous:2009:Cl

[Ano09m] Anonymous. Contents. Journal of Statistical Planning and Inference, 139(12):iii–iv, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900264X.

Anonymous:2009:I

[Ano09n] Anonymous. IBC. Journal of Statistical Planning and Inference, 139(2):ibc, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808004096.

Anonymous:2009:IECa

[Ano09o] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(1):ibc, January 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003777.

Anonymous:2009:IECb

[Ano09p] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(3):ibc, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004321.

Anonymous:2009:IECc

[Ano09q] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(4):ibc, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004631.

Anonymous:2009:IECd

[Ano09r] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(5):ibc, May 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000226.

Anonymous:2009:IECe

[Ano09s] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(6):ibc, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000445.

Anonymous:2009:IECf

[Ano09t] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(7):ibc, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000640.

Anonymous:2009:IECg

[Ano09u] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(8):ibc, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900113X.

Anonymous:2009:IECh

[Ano09v] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(9):ibc, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900175X.

Anonymous:2009:IECi

[Ano09w] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(10):ibc, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002031.

Anonymous:2009:IECj

[Ano09x] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(11):ibc, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002407.

Anonymous:2009:IECk

[Ano09y] Anonymous. IBC (Editors cont.). Journal of Statistical Planning and Inference, 139(12):ibc, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002651.

Anonymous:2009:OEa

[Ano09z] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(1):obc, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003789.

Anonymous:2009:OEb

[Ano09-27] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(2):obc, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004102.

Anonymous:2009:OEc

[Ano09-28] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(3):obc, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004333.

Anonymous:2009:OEd

[Ano09-29] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(4):obc, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004643.

Anonymous:2009:OEe

[Ano09-30] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(5):obc, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000238.

Anonymous:2009:OEf

[Ano09-31] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(6):obc, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000457.

Anonymous:2009:OEg

[Ano09-32] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(7):obc, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000652.

Anonymous:2009:OEh

[Ano09-33] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(8):obc, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001141.

Anonymous:2009:OEi

[Ano09-34] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(9):obc, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001761.

Anonymous:2009:OEj

[Ano09-35] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(10):obc, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002043.

Anonymous:2009:OEk

[Ano09-36] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(11):obc, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002419.

Anonymous:2009:OEl

[Ano09-37] Anonymous. OBC (Editors). Journal of Statistical Planning and Inference, 139(12):obc, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809002663.

Anonymous:2009:PAa

[Ano09-38] Anonymous. Pages 1257-1574 (1 April 2009). Journal of Statistical Planning and Inference, 139(4):??, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PJa

[Ano09-39] Anonymous. Pages 1821-2116 (1 June 2009). Journal of Statistical Planning and Inference, 139(6):??, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PJb

[Ano09-40] Anonymous. Pages 2117-2510 (1 July 2009). Journal of Statistical Planning and Inference, 139(7):??, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PAb

[Ano09-41] Anonymous. Pages 2511-2862 (1 August 2009). Journal of Statistical Planning and Inference, 139(8):??, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PS

[Ano09-42] Anonymous. Pages 2863-3392 (1 September 2009). Journal of Statistical Planning and Inference, 139(9):??, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PO

[Ano09-43] Anonymous. Pages 3393-3738 (1 October 2009). Journal of Statistical Planning and Inference, 139(10):??, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PD

[Ano09-44] Anonymous. Pages 3947-4240 (1 December 2009). Journal of Statistical Planning and Inference, 139(12):??, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PM

[Ano09-45] Anonymous. Pages 679-1256 (1 March 2009). Journal of Statistical Planning and Inference, 139(3):??, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:PF

[Ano09-46] Anonymous. Pages 81-678 (1 February 2009). Journal of Statistical Planning and Inference, 139(2):??, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2009:SIM

[Ano09-47] Anonymous. Special issue on metaheuristics, combinatorial optimization and design of experiments. Journal of Statistical Planning and Inference, 139(1):1, January 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800222X.

Anonymous:2009:SIT

[Ano09-48] Anonymous. A special issue on the topic of Degradation, Damage, Fatigue and Accelerated Life Models in Reliability Testing. Journal of Statistical Planning and Inference, 139(5): 1575, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002322.

Ango-Nze:2000:DEN

[ANR00] Patrick Ango-Nze and Ricardo Rios. Density estimation in L^{∞} norm for mixing processes. Journal of Statistical Planning and Inference, 83(1):75–90, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001712.

Anselin:2001:RST

[Ans01] Luc Anselin. Rao’s score test in spatial econometrics. Journal of Statistical Planning and Inference, 97(1):113–139, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003499.

Antognini:2008:TAP

[Ant08] Alessandro Baldi Antognini. A theoretical analysis of the power of biased coin designs. Journal of Statistical Planning and Inference, 138(6):1792–1798, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003217.

Agustin:2005:BAG

[AP05] Ma. Zenia N. Agustin and Edsel A. Peña. A basis approach to goodness-of-fit testing in recurrent event models. Journal of Statistical Planning and Inference, 133(2):285–303, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001697.

Adimari:2007:RCI

[AP07]  Gianfranco Adimari and Paolo Preo.  Robust confi-

dence intervals for log-location-scale models with right censored data. Journal of Statistical Planning and Inference, 137(9):2832–2849, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002886.

Aneiros-Perez:2002:PBC

[APdR02] Germán Aneiros-Pérez and Alejandro Quintela del Riío. Plug-in bandwidth choice in partial linear models with autoregressive errors. Journal of Statistical Planning and Inference, 100(1):23–48, January 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000945.

Alonso:2002:FTS

[APR02] Andre’es M. Alonso, Daniel Peña, and Juan Romo. Forecasting time series with sieve bootstrap. Journal of Statistical Planning and Inference, 100(1):1–11, January 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000921.

Aires:2005:IPR

[AR05] Nibia Aires and Bengt Rosén. On inclusion probabilities and relative estimator bias for Pareto πps sampling. Journal of Statistical Planning and Inference, 128(2):543–567, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000242.

Arcones:2001:ADR

[Arc01] Miguel A. Arcones. Asymptotic distribution of regression M-estimators. Journal of Statistical Planning and Inference, 97(2):235–261, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580000224X.

Arnold:2003:BB

[Arn03] Barry C. Arnold. Back to Bayesics. Journal of Statistical Planning and Inference, 109(1–2):179–187, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003105.

Arnold:2009:FUM

[Arn09] Barry C. Arnold. Flexible univariate and multivariate models based on hidden truncation. Journal of Statistical Planning and Inference, 139(11):3741–3749, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001384.

Asadi:2001:SUC

[ARS01] M. Asadi, C. R. Rao, and D. N. Shanbhag. Some unified characterization results on generalized Pareto distributions. Journal of Statistical Planning and Inference, 93(1–2):29–50, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001713.

Arslan:2004:CBI

[Ars04] Olcay Arslan. Convergence behavior of an iterative reweighting algorithm to compute multivariate M-estimates for location and scatter. Journal of Statistical Planning and Inference, 118(1–2):115–128, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004020.

Anido:2003:REL

[ARV03] Carmen Anido, Carlos Rivero, and Teófilo Valdés. Recursive estimation in linear models with general errors and grouped data: a median-based procedure and related asymptotics. Journal of Statistical Planning and Inference, 115(1):85–102, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001143.

Arnold:2000:AVK

[AS00] Bernhard F. Arnold and Peter Stahlecker. Another view of the Kuks–Olman estimator. Journal of Statistical Planning and Inference, 89(1–2):169–174, August 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000987.

Al-Saleh:2004:SSR

[AS04] Mohammad Fraiwan Al-Saleh. Steady-state ranked set sampling and parametric estimation. Journal of Statistical Planning and Inference, 123(1):83–95, June 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001393.

Asadi:2006:MPL

[Asa06] Majid Asadi. On the mean past lifetime of the components of a parallel system. Journal of Statistical Planning and Inference, 136(4):1197–1206, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003994.

Al-Saleh:2002:MRS

[ASAO02] Mohammad Fraiwan Al-Saleh and Amer Ibrahim Al-Omari. Multistage ranked set sampling. Journal of Statistical Planning and Inference, 102(2):273–286, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000866.

Adrover:2004:GRI

[ASBZ04] Jorge Adrover, Matias Salibian-Barrera, and Ruben Zamar. Globally robust inference for the location and simple linear regression models. Journal of Statistical Planning and Inference, 119(2):353–375, February 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004901.

Assaf:2005:WOC

[ASC05] David Assaf and Ester Samuel-Cahn. Why is one choice different? Journal of Statistical Planning and Inference, 130(1–2):127–132, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804002575.

Al-Saleh:2009:EPD

[ASD09] Mohammad Fraiwan Al-Saleh and Yusef Abdallah Diab. Estimation of the parameters of Downton’s bivariate exponential distribution using ranked set sampling scheme. Journal of Statistical Planning and Inference, 139(2):277–286, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001808.

Archer:2002:PEH

[AT02] G. E. B. Archer and D. M. Titterington. Parameter estimation for hidden Markov chains. Journal of Statistical Planning and Inference, 108(1–2):365–390, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200318X.

Aoki:2009:SCA

[AT09] Satoshi Aoki and Akimichi Takemura. Some characterizations of affinely full-dimensional factorial designs. Journal of Statistical Planning and Inference, 139(10):3525–3532, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000937.

Atkinson:2008:DOD

[Atk08a] A. C. Atkinson. DT-optimum designs for model discrimination and parameter estimation. Journal of Statistical Planning and Inference, 138(1):56–64, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001954.

Atkinson:2008:EUE

[Atk08b] Anthony C. Atkinson. Examples of the use of an equivalence theorem in constructing optimum experimental designs for random-effects nonlinear regression models. Journal of Statistical Planning and Inference, 138(9):2595–2606, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000529.

Atkinson:2009:OOR

[Atk09] Anthony C. Atkinson. Optimum and other response surface designs. Comments on “Response Surface Design Evaluation and Comparison” by Anderson-Cook, Borror and Montgomery. Journal of Statistical Planning and Inference, 139(2):662–668, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001651.

Aoshima:2002:TSP

[ATS02] Makoto Aoshima, Yoshikazu Takada, and Muni S. Srivastava. A two-stage procedure for estimating a linear function of k multinormal mean vectors when covariance matrices are unknown. Journal of Statistical Planning and Inference, 100(2):109–119, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001264.

Aue:2006:TPS

[Aue06] Alexander Aue. Testing for parameter stability in RCA(1) time series. Journal of Statistical Planning and Inference, 136(9):3070–3089, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000340.

Augustin:2002:NPT

[Aug02] Thomas Augustin. Neyman–Pearson testing under interval probability by globally least favorable pairs: Reviewing Huber–Strassen theory and extending it to general interval probability. Journal of Statistical Planning and Inference, 105(1):149–173, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801002087.

Austin:2005:RVB

[Aus05] Peter C. Austin. The reliability and validity of Bayesian measures for hospital profiling: a Monte Carlo assessment. Journal of Statistical Planning and Inference, 128(1):109–122, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003732.

Arellano-Valle:2005:SIG

[AVGQ05] Reinaldo B. Arellano-Valle, Héctor W. Gómez, and Fernando A. Quintana. Statistical inference for a general class of asymmetric distributions. Journal of Statistical Planning and Inference, 128(2):427–443, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000254.

Arellano-Valle:2000:BSA

[AVGRZ00] R. B. Arellano-Valle, M. Galea-Rojas, and P. Iglesias Zuazola. Bayesian sensitivity analysis in elliptical linear regression models. Journal of Statistical Planning and Inference, 86(1):175–199, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001664.

Aasenblad:2006:ACD

[ÅvR06] Nils Åsenblad and Dietrich von Rosen. Analysis of crossover designs via growth curve models. Journal of Statistical Planning and Inference, 136(2):475–497, February 1, 2006.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002939.

Abarin:2009:SOL

[AW09a] Taraneh Abarin and Liqun Wang. Second-order least squares estimation of censored regression models. Journal of Statistical Planning and Inference, 139(2):125–135, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001766.

Adewale:2009:RDM

[AW09b] Adeniyi J. Adewale and Douglas P. Wiens. Robust designs for misspecified logistic models. Journal of Statistical Planning and Inference, 139(1):3–15, January 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002243.

Ausien:2004:BEM

[AWL04] M. C. Ausien, M. P. Wiper, and R. E. Lillo. Bayesian estimation for the M/G/1 queue using a phase-type approximation. Journal of Statistical Planning and Inference, 118(1–2):83–101, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003981.

Ai:2009:OCN

[AYH09] Mingyao Ai, Yunlong Yu, and Shuyuan He. Optimality of circular neighbor-balanced designs for total effects with autoregressive correlated observations. Journal of Statistical Planning and Inference, 139(7):2293–2304, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004709.

Ai:2006:MAB

[AYZ06] Mingyao Ai, Guijun Yang, and Runchu Zhang. Minimum aberration blocking of regular mixed factorial designs. Journal of Statistical Planning and Inference, 136(4):1493–1511, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580400391X.

Adrover:2004:BRT

[AZ04a] Jorge Adrover and Ruben H. Zamar. Bias robustness of three median-based regression estimates. Journal of Statistical Planning and Inference, 122(1–2):203–227, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002337.

Ai:2004:TMA

[AZ04b] Mingyao Ai and Runchu Zhang. Theory of minimum aberration blocked regular mixed factorial designs. Journal of Statistical Planning and Inference, 126(1):305–323, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002635.

Asadi:2007:DCR

[AZ07] Majid Asadi and Younes Zohrevand. On the dynamic cumulative residual entropy. Journal of Statistical Planning and Inference, 137(6):1931–1941, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001807.

An:2000:MTT

[AZL00] Hong-Zhi An, Li-Xing Zhu, and Run-Ze Li. A mixed-type test for linearity in time series. Journal of Statistical Planning and Inference, 88(2):339–353, August 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000872.

Azomahou:2009:MPA

[Azo09] Théophile T. Azomahou. Memory properties and aggregation of spatial autoregressive models. Journal of Statistical Planning and Inference, 139(8):2581–2597, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004473.

Baratpour:2008:CBR

[BAA08] S. Baratpour, J. Ahmadi, and N. R. Arghami. Characterizations based on Rényi entropy of order statistics and record values. Journal of Statistical Planning and Inference, 138(8):2544–2551, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000256.

Babu:2004:NBE

[Bab04] Gutti Jogesh Babu. A note on the bootstrapped empirical process. Journal of Statistical Planning and Inference, 126(2):587–589, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002830.

Bagchi:2004:CGD

[Bag04] Sunanda Bagchi. Construction of group divisible designs and rectangular designs from resolvable and almost resolvable balanced incomplete block designs. Journal of Statistical Planning and Inference, 119(2):401–410, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004858.

Bakshaev:2009:GFH

[Bak09] Aleksej Bakshaev. Goodness of fit and homogeneity tests on the basis of N-distances. Journal of Statistical Planning and Inference, 139(11):3750–3758, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001396.

Bandyopadhyay:2003:MSR

[Ban03] Debasis Bandyopadhyay. On the model selection to represent human capital distribution — an empirical study. Journal of Statistical Planning and Inference, 112(1–2):213–220, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003348.

Barry:2003:IPR

[Bar03] Daniel Barry. The inclusion of partially reported studies in a Bayesian meta-analysis. Journal of Statistical Planning and Inference, 115(2):365–378, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001696.

Barakat:2007:LTG

[Bar07] H. M. Barakat. Limit theory of generalized order statistics. Journal of Statistical Planning and Inference, 137(1):1–11, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002934.

Barabesi:2008:DBR

[Bar08] L. Barabesi. A design-based randomized response procedure for the estimation of population proportion and sensitivity level. Journal of Statistical Planning and Inference, 138(8):2398–2408, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003862.

Bathke:2004:ATC

[Bat04] Arne Bathke. The ANOVA F test can still be used in some balanced designs with unequal variances and nonnormal data. Journal of Statistical Planning and Inference, 126(2):413–422, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002787.

Battaglia:2006:BCO

[Bat06] Francesco Battaglia. Bias correction for outlier estimation in time series. Journal of Statistical Planning and Inference, 136(11):3904–3930, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001254.

Bandyopadhyay:2000:CTT

[BB00] Uttam Bandyopadhyay and Atanu Biswas. Comparison of two treatments with heterogeneous experimental units. Journal of Statistical Planning and Inference, 83(2):441–448, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899000993.

Becker:2001:DIS

[BB01a] Niels G. Becker and Tom Britton. Design issues for studies of infectious diseases. Journal of Statistical Planning and Inference, 96(1):41–66, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003232.

Bera:2001:RSN

[BB01b] Anil K. Bera and Yannis Bilias. Rao’s score, Neyman’s C(α) and Silvey’s LM tests: an essay on historical developments and some new results. Journal of Statistical Planning and Inference, 97(1):9–44, August 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003438.

Banerjee:2003:NFS

[BB03a] T. Banerjee and Atanu Biswas. A new formulation of stressstrength reliability in a regression setup. Journal of Statistical Planning and Inference, 112(1–2):147–157, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003294.

Bozdogan:2003:ICC

[BB03b] Hamparsum Bozdogan and Peter Bearse. Information complexity criteria for detecting influential observations in dynamic multivariate linear models using the genetic algorithm. Journal of Statistical Planning and Inference, 114(1–2):31– 44, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004615.

Bandyopadhyay:2004:AAC

[BB04a] Uttam Bandyopadhyay and Atanu Biswas. An adaptive allocation for continuous response using Wilcoxon-Mann-Whitney score. Journal of Statistical Planning and Inference, 123(2):207–224, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001472.

Bandyopadhyay:2004:RAP

[BB04b] Uttam Bandyopadhyay and Atanu Biswas. Randomized allocation procedure for testing a normal mean with known variance. Journal of Statistical Planning and Inference, 121(2):215–229, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001204.

Bianco:2004:RES

[BB04c] Ana Bianco and Graciela Boente. Robust estimators in semiparametric partly linear regression models. Journal of Statistical Planning and Inference, 122(1–2):229–252, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002398.

Bandyopadhyay:2006:OSP

[BB06a] Uttam Bandyopadhyay and Rahul Bhattacharya. An optimal sequential procedure in ethical allocation. Journal of Statistical Planning and Inference, 136(2):430–446, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003192.

Bordes:2006:SES

[BB06b] L. Bordes and C. Breuils. Sequential estimation for semiparametric models with application to the proportional hazards model. Journal of Statistical Planning and Inference, 136(11):3735–3759, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001321.

Bartoszewicz:2009:SPG

[BB09] Jaroslaw Bartoszewicz and Magdalena Benduch. Some properties of the generalized TTT transform. Journal of Statistical Planning and Inference, 139(7):2208–2217, July 1, 2009.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004011.

Bianco:2000:SRR

[BBdR00] Ana Bianco, Graciela Boente, and Julio di Rienzo. Some results for robust GM-based estimators in heteroscedastic regression models. Journal of Statistical Planning and Inference, 89(1–2):215–242, August 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000938.

Branco:2000:BAC

[BBIAV00] M. Branco, H. Bolfarine, P. Iglesias, and R. B. Arellano-Valle. Bayesian analysis of the calibration problem under elliptical distributions. Journal of Statistical Planning and Inference, 90(1):69–85, September 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001105.

Bagdonavicius:2007:AJM

[BBKN07] V. Bagdonavicius, A. Bikelis, V. Kazakevicius, and Mikhail Nikulin. Analysis of joint multiple failure mode and linear degradation data with renewals. Journal of Statistical Planning and Inference, 137(7):2191–2207, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002047.

Brannath:2006:EAD

[BBP06] Werner Brannath, Peter Bauer, and Martin Posch. On the efficiency of adaptive designs for flexible interim decisions in clinical trials. Journal of Statistical Planning and Inference, 136(6):1956–1961, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001953.

Brockhoff:2004:PAS

[BBR04] P. B. Brockhoff, D. J. Best, and J. C. W. Rayner. Partitioning Anderson’s statistic for tied data. Journal of Statistical Planning and Inference, 121(1):93–111, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802005086.

Bart:2005:MDD

[BBSZ05] A. G. Bart, V. A. Bart, A. Steland, and M. L. Zaslavskiy. Modeling disease dynamics and survivor functions by sanogenesis curves. Journal of Statistical Planning and Inference, 132(1–2):33–51, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001405.

Beirlant:2006:SLB

[BBW06] Jan Beirlant, Christel Bouquiaux, and Bas J. M. Werker. Semiparametric lower bounds for tail index estimation. Journal of Statistical Planning and Inference, 136(3):705–729, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003593.

Barnard:2002:LIL

[BC02] George A. Barnard and John B. Copas. Likelihood inference for location, scale, and shape. Journal of Statistical Planning and Inference, 108(1–2):71–83, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002719.

Bai:2003:WTN

[BC03a] Z. D. Bai and Ling Chen. Weighted W test for normality and asymptotics a revisit of Chen-Shapiro test for normality. Journal of Statistical Planning and Inference, 113(2):485–503, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001106.

Bai:2003:TRS

[BC03b] Zhidong Bai and Zehua Chen. On the theory of rankedset sampling and its ramifications. Journal of Statistical Planning and Inference, 109(1–2):81–99, January 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003026.

Bose:2003:GBE

[BC03c] Arup Bose and Snigdhansu Chatterjee. Generalized bootstrap for estimators of minimizers of convex functions. Journal of Statistical Planning and Inference, 117(2):225–239, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003865.

Broniatowski:2007:OBS

[BC07] M. Broniatowski and G. Celant. Optimality and bias of some interpolation and extrapolation designs. Journal of Statistical Planning and Inference, 137(3):858–868, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001467.

Berrendero:2009:CEW

[BC09] José R. Berrendero and Javier Cárcamo. Characterizations of exponentiality within the HNBUE class and related tests. Journal of Statistical Planning and Inference, 139(7):2399– 2406, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004448.

Babu:2002:ABP

[BCC02] G. Jogesh Babu, Angelo J. Canty, and Yogendra P. Chaubey. Application of Bernstein polynomials for smooth estimation of a distribution and density function. Journal of Statistical Planning and Inference, 105(2):377–392, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002658.

Boj:2009:PET

[BCGF09] Eva Boj, M. Mercè Claramunt, Aurea Grané, and Josep Fortiana. Projection error term in Gower’s interpolation. Journal of Statistical Planning and Inference, 139(6):1867–1878, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003467.

Bagdonavicius:2004:SPI

[BCN04] Vilijandas Bagdonavicius, Olga Cheminade, and Mikhail Nikulin. Statistical planning and inference in accelerated life testing using the CHSS model. Journal of Statistical Planning and Inference, 126(2):535–551, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002775.

Brouwer:2006:OAS

[BCN06] Andries E. Brouwer, Arjeh M. Cohen, and Man V. M. Nguyen. Orthogonal arrays of strength 3 and small run sizes. Journal of Statistical Planning and Inference, 136(9):3268–3280, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000327.

Balakrishnan:2003:BEO

[BCP03] N. Balakrishnan, C. Charalambides, and N. Papadatos. Bounds on expectation of order statistics from a finite population. Journal of Statistical Planning and Inference, 113 (2):569–588, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801003214.

Belzunce:2000:TMR

[BCR00] Félix Belzunce, José Candel, and José M. Ruiz. Testing mean residual alternatives by dispersion of residual lives. Journal of Statistical Planning and Inference, 86(1):113–127, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001676.

Barbachano:2008:PDW

[BCR08] Yolanda Barbáchano, D. Stephen Coad, and Derek R. Robinson. Predictability of designs which adjust for imbalances in prognostic factors. Journal of Statistical Planning and Inference, 138(3):756–767, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000468.

Basso:2007:SPT

[BCS07] Dario Basso, Marco Chiarandini, and Luigi Salmaso. Synchronized permutation tests in replicated I×J designs. Journal of Statistical Planning and Inference, 137(8):2564–2578, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000110.

Bermudez:2009:FTS

[BCVV09] José D. Bermúdez, Ana Corberán-Vallet, and Enriqueta Vercher. Forecasting time series with missing data using Holt’s model. Journal of Statistical Planning and Inference, 139(8):2791–2799, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000032.

Biswas:2004:SOE

[BD04] Atanu Biswas and Anup Dewanji. Scheduling optimal examination times in a simple illness-death model. Journal of Statistical Planning and Inference, 126(1):173–187, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580300226X.

Brodsky:2005:AOM

[BD05] Boris Brodsky and Boris Darkhovsky. Asymptotically optimal methods of change-point detection for composite hypotheses. Journal of Statistical Planning and Inference, 133(1):123–138, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001090.

Balakrishnan:2008:PTI

[BD08a] N. Balakrishnan and Anna Dembińska. Progressively Type-II right censored order statistics from discrete distributions. Journal of Statistical Planning and Inference, 138(4):845–856, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000584. See erra-

tum [BD09].

Bandyopadhyay:2008:TES

[BD08b] Dipankar Bandyopadhyay and Somnath Datta. Testing equality of survival distributions when the population marks are missing. Journal of Statistical Planning and Inference, 138(6):1722–1732, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003175.

Brutti:2008:RBS

[BD08c] Pierpaolo Brutti and Fulvio De Santis. Robust Bayesian sample size determination for avoiding the range of equivalence in clinical trials. Journal of Statistical Planning and Inference, 138(6):1577–1591, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700300X.

Balakrishnan:2009:EPT

[BD09] N. Balakrishnan and Anna Dembińska. Erratum to “Progressively Type-II right censored order statistics from discrete distributions” [J. Statist. Plann. Inference 138 (2008) 845–856]. Journal of Statistical Planning and Inference, 139(4):1572–1574, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003704. See [BD08a].

Beirlant:2001:LDD

[BDGV01] Jan Beirlant, Luc Devroye, László Györfi, and Igor Vajda. Large deviations of divergence measures on partitions. Journal of Statistical Planning and Inference, 93(1–2):1–16, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002020.

Bortnick:2005:EBD

[BDH05] Steven M. Bortnick, Angela M. Dean, and Jason C. Hsu. Efficiency of block designs for simultaneous comparison of two test treatments with a control. Journal of Statistical Planning and Inference, 129(1–2):109–124, February 2005.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001570.

Brown:2006:HES

[BDHS06] L. Brown, A. DasGupta, L. R. Haff, and W. E. Strawderman. The heat equation and Stein’s identity: Connections, applications. Journal of Statistical Planning and Inference, 136(7):2254–2278, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805003022.

Bilodeau:2009:DSM

[BdM09] M. Bilodeau and P. Lafaye de Micheaux. A-dependence statistics for mutual and serial independence of categorical variables. Journal of Statistical Planning and Inference, 139(7):2407–2419, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004370.

Biedermann:2007:ODD

[BDP07] Stefanie Biedermann, Holger Dette, and Andrey Pepelyshev. Optimal discrimination designs for exponential regression models. Journal of Statistical Planning and Inference, 137(8):2579–2592, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000122.

Biedermann:2007:COD

[BDZ07] Stefanie Biedermann, Holger Dette, and Wei Zhu. Compound optimal designs for percentile estimation in dose-response models with restricted design intervals. Journal of Statistical Planning and Inference, 137(12):3838–3847, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580700153X.

Bairamov:2000:DPS

[BE00] Ismihan G. Bairamov and Serkan N. Eryilmaz. Distributional properties of statistics based on minimal spacing and record exceedance statistics. Journal of Statistical Planning and Inference, 90(1):21–33, September 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001075.

Baker:2001:SLO

[BE01] R. D. Baker and G. L. Ebert. Singer line orbits in PG(3, q). Journal of Statistical Planning and Inference, 95(1–2):75–88, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002792.

Bairamov:2006:SEC

[BE06] Ismihan Bairamov and Serkan Eryilmaz. Spacings, exceedances and concomitants in progressive type II censoring scheme. Journal of Statistical Planning and Inference, 136(3):527–536, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580400360X.

Beder:1998:CAH

[Bed98] Jay H. Beder. Conjectures about Hadamard matrices. Journal of Statistical Planning and Inference, 72(1–2):7–14, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375898000202. See erra-

tum [Bed00].

Beder:2000:ECA

[Bed00] Jay H. Beder. Erratum to “Conjectures About Hadamard Matrices” [J Statist. Plann. Inference 72 (1998) 7–14]. Journal of Statistical Planning and Inference, 84(1–2):343, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001482. See [Bed98].

Bedford:2006:CDD

[Bed06] Tim Bedford. Copulas, degenerate distributions and quantile tests in competing risk problems. Journal of Statistical Planning and Inference, 136(5):1572–1587, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001680.

Beh:2008:CAA

[Beh08] Eric J. Beh. Correspondence analysis of aggregate data: the 2×2 table. Journal of Statistical Planning and Inference, 138(10):2941–2952, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807004077.

Bellout:2007:ARS

[Bel07] Djamel Bellout. An adaptive rule for stopping a discovery process under time censorship. Journal of Statistical Planning and Inference, 137(7):2328–2335, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002710.

Bennett:2001:RPE

[Ben01] F. E. Bennett. Recent progress on the existence of perfect Mendelsohn designs. Journal of Statistical Planning and Inference, 94(2):121–138, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002457.

Bentkus:2003:DBE

[Ben03] V. Bentkus. On the dependence of the Berry–Esseen bound on dimension. Journal of Statistical Planning and Inference, 113(2):385–402, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802000940.

Berger:2001:WSC

[Ber01a] James Berger. Why should clinicians care about Bayesian methods? Journal of Statistical Planning and Inference, 94 (1):65–67, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002354.

Berry:2001:EUM

[Ber01b] J. Calvin Berry. On the existence of a unique most probable category. Journal of Statistical Planning and Inference, 99(2):175–182, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000829.

Bernard:2002:IAM

[Ber02a] Jean-Marc Bernard. Implicative analysis for multivariate binary data using an imprecise Dirichlet model. Journal of Statistical Planning and Inference, 105(1):83–103, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002051.

Bernard:2002:P

[Ber02b] Jean-Marc Bernard. Preface. Journal of Statistical Planning and Inference, 105(1):1–4, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801002002.

Berrendero:2002:GRG

[Ber02c] José R. Berrendero. On the global robustness of generalized S-estimators. Journal of Statistical Planning and Inference, 102(2):287–302, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580100088X.

Bertin:2004:MEC

[Ber04] Karine Bertin. Minimax exact constant in sup-norm for nonparametric regression with random design. Journal of Statistical Planning and Inference, 123(2):225–242, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300154X.

Berger:2005:VEC

[Ber05] Yves G. Berger. Variance estimation with Chao’s sampling scheme. Journal of Statistical Planning and Inference, 127(1–2):253–277, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002933.

Beran:2007:APF

[Ber07] Rudolf Beran. Adaptation over parametric families of symmetric linear estimators. Journal of Statistical Planning and Inference, 137(3):684–696, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001340.

Beran:2009:PEL

[Ber09] Jan Beran. On parameter estimation for locally stationary long-memory processes. Journal of Statistical Planning and Inference, 139(3):900–915, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002644.

Barakat:2005:CMB

[BES05] H. M. Barakat and M. A. El-Shandidy. On the computing of the moments of bivariate order statistics. Journal of Statistical Planning and Inference, 134(1):15–25, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580400182X.

Betensky:2001:OSC

[Bet01] Rebecca A. Betensky. Optimally selected chi square statistics for equivalence testing. Journal of Statistical Planning and Inference, 93(1–2):247–257, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002044.

Banerjee:2007:BMR

[BF07] Sudipto Banerjee and Andrew Finley. Bayesian multiresolution modeling for spatially replicated data sets with application to forest biomass data. Journal of Statistical Planning and Inference, 137(10):3193–3205, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000857.

Beirlant:2008:IRB

[BFGV08] Jan Beirlant, Fernanda Figueiredo, M. Ivette Gomes, and Björn Vandewalle. Improved reduced-bias tail index and quantile estimators. Journal of Statistical Planning and Inference, 138(6):1851–1870, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003485.

Bortot:2005:EFS

[BG05] Paola Bortot and Alessandra Giovagnoli. Up-and-down experiments of first and second order. Journal of Statistical Planning and Inference, 134(1):236–253, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001946.

Brillinger:2007:MIF

[BG07] David R. Brillinger and Apratim Guha. Mutual information in the frequency domain. Journal of Statistical Planning and Inference, 137(3):1076–1084, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001625.

Biswas:2009:TSA

[BG09] Atanu Biswas and Apratim Guha. Time series analysis of categorical data using auto-mutual information. Journal of Statistical Planning and Inference, 139(9):3076–3087, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000330.

Bennett:2000:BEP

[BGG00] G. K. Bennett, M. J. Grannell, and T. S. Griggs. Biembeddings of the projective space PG(3,2). Journal of Statistical Planning and Inference, 86(2):321–329, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001159.

Berkes:2009:TCC

[BGH09] István Berkes, Edit Gombay, and Lajos Horváth. Testing for changes in the covariance structure of linear processes. Journal of Statistical Planning and Inference, 139(6):2044–2063, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003807.

Bolfarine:2003:IUR

[BGI03] Heleno Bolfarine, Loretta B. Gasco, and Pilar L. Iglesias. Inference under representable priors for Pearson type II models in finite populations. Journal of Statistical Planning and Inference, 111(1–2):23–36, ???? 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002835.

Ben-Gal:2001:AIT

[BGL01] Irad Ben-Gal and Lev B. Levitin. An application of information theory and error-correcting codes to fractional factorial experiments. Journal of Statistical Planning and Inference, 92(1–2):267–282, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001658.

Berger:2003:ACB

[BGM03] James O. Berger, Jayanta K. Ghosh, and Nitai Mukhopadhyay. Approximations and consistency of Bayes factors as model dimension grows. Journal of Statistical Planning and Inference, 112(1–2):241–258, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003361.

Boente:2009:RNE

[BGMPG09] Graciela Boente, Wenceslao González-Manteiga, and Ana Pérez-González. Robust nonparametric estimation with missing data. Journal of Statistical Planning and Inference, 139(2):571–592, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002000.

Banerjee:2000:GMS

[BGP00] Sudipto Banerjee, Alan E. Gelfand, and Wolfgang Polasek. Geostatistical modelling for spatial interaction data with application to postal service performance. Journal of Statistical Planning and Inference, 90(1):87–105, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001117.

Beran:2003:NEL

[BGS03] Jan Beran, Sucharita Ghosh, and Philipp Sibbertsen. Nonparametric M-estimation with long-memory errors. Journal of Statistical Planning and Inference, 117(2):199–205, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003919.

Bellio:2008:MQP

[BGV08] Ruggero Bellio, Luca Greco, and Laura Ventura. Modified quasi-profile likelihoods from estimating functions. Journal of Statistical Planning and Inference, 138(10):3059–3068, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000232.

Bennett:2002:HSP

[BGZZ02] F. E. Bennett, G. Ge, H. Zhang, and L. Zhu. Holey Steiner pentagon systems and related designs. Journal of Statistical Planning and Inference, 106(1–2):353–373, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002227.

Blokhuis:2001:IFQ

[BH01a] Aart Blokhuis and Willem H. Haemers. An infinite family of quasi-symmetric designs. Journal of Statistical Planning and Inference, 95(1–2):117–119, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002810.

Bluskov:2001:SSD

[BH01b] Iliya Bluskov and Katherine Heinrich. Super-simple designs with v ≤ 32. Journal of Statistical Planning and Inference, 95(1–2):121–131, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002822.

Brown:2002:EIS

[BH02] B. M. Brown and T. P. Hettmansperger. On the estimation of integral f-squared. Journal of Statistical Planning and Inference, 107(1–2):365–375, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002641.

Bratcher:2005:BMC

[BH05] Tom Bratcher and Cody Hamilton. A Bayesian multiple comparison procedure for ranking the means of normally distributed data. Journal of Statistical Planning and Inference, 133(1):23–32, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001077.

Bissantz:2007:EQR

[BH07] Nicolai Bissantz and Hajo Holzmann. Estimation of a quadratic regression functional using the sinc kernel. Journal of Statistical Planning and Inference, 137(3):712–719, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001364.

Balakrishnan:2008:EIS

[BH08a] N. Balakrishnan and Donghoon Han. Exact inference for a simple step-stress model with competing risks for failure from exponential distribution under Type-II censoring. Journal of Statistical Planning and Inference, 138(12):4172–4186, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001523.

Bathke:2008:NMM

[BH08b] Arne C. Bathke and Solomon W. Harrar. Nonparametric methods in multivariate factorial designs for large number of factor levels. Journal of Statistical Planning and Inference, 138(3):588–610, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001863.

Balakrishnan:2009:OSS

[BH09a] N. Balakrishnan and Donghoon Han. Optimal step-stress testing for progressively Type-I censored data from exponential distribution. Journal of Statistical Planning and Inference, 139(5):1782–1798, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800253X.

BenFarah:2009:DDS

[BH09b] Mohamed Ben Farah and Abdelhamid Hassairi. On the Dirichlet distributions on symmetric matrices. Journal of Statistical Planning and Inference, 139(8):2559–2570, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580800445X.

Bhattacharya:2007:TOF

[Bha07] Bhaskar Bhattacharya. Testing for ordered failure rates under general progressive censoring. Journal of Statistical Planning and Inference, 137(6):1775–1786, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001108.

Bhattacharya:2009:OUH

[Bha09] Bhaskar Bhattacharya. Optimal use of historical information. Journal of Statistical Planning and Inference, 139(12):4051–4063, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001359.

Berger:2002:BLR

[BI02] Vance W. Berger and Anastasia Ivanova. The bias of linear rank tests when testing for stochastic order in ordered categorical data. Journal of Statistical Planning and Inference, 107(1–2):237–247, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002550.

Bier:2000:DSI

[Bie00] Thomas Bier. Difference sets and intercalation pairs. Journal of Statistical Planning and Inference, 86(2):331–348, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001160.

Bieniek:2008:PBE

[Bie08] Mariusz Bieniek. Projection bounds on expectations of generalized order statistics from DD and DDA families. Journal of Statistical Planning and Inference, 138(4):971–981, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002224.

Birke:2009:SCK

[Bir09] Melanie Birke. Shape constrained kernel density estimation. Journal of Statistical Planning and Inference, 139 (8):2851–2862, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580900010X.

Bate:2006:CNB

[BJ06] S. T. Bate and B. Jones. The construction of nearly balanced and nearly strongly balanced uniform crossover designs. Journal of Statistical Planning and Inference, 136(9):3248–3267, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004690.

Bate:2008:RUC

[BJ08] S. T. Bate and B. Jones. A review of uniform cross-over designs. Journal of Statistical Planning and Inference, 138 (2):336–351, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807002455.

Bogacka:2008:EWE

[BJJV08] B. Bogacka, P. Johnson, B. Jones, and O. Volkov. Defficient window experimental designs. Journal of Statistical Planning and Inference, 138(1):160–168, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002066.

Brazauskas:2008:ECT

[BJPZ08] Vytaras Brazauskas, Bruce L. Jones, Madan L. Puri, and Ricardas Zitikis. Estimating conditional tail expectation with actuarial applications in view. Journal of Statistical Planning and Inference, 138(11):3590–3604, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001018.

Beirlant:2009:SOR

[BJS09] Jan Beirlant, Elisabeth Joossens, and Johan Segers. Secondorder refined peaks-over-threshold modelling for heavytailed distributions. Journal of Statistical Planning and Inference, 139(8):2800–2815, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000068.

Brazauskas:2009:RFC

[BJZ09] Vytaras Brazauskas, Bruce L. Jones, and Ricardas Zitikis. Robust fitting of claim severity distributions and the method of trimmed moments. Journal of Statistical Planning and Inference, 139(6):2028–2043, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003716.

Boutsikas:2002:NOU

[BK02] M. V. Boutsikas and M. V. Koutras. On the number of overflown urns and excess balls in an allocation model with limited urn capacity. Journal of Statistical Planning and Inference, 104(2):259–286, June 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002634.

Begum:2005:MMP

[BK05a] Nelufa Begum and Maxwell L. King. Most mean powerful invariant test for testing two-dimensional parameter spaces. Journal of Statistical Planning and Inference, 134(2):536–548, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002113.

Bohm:2005:NIB

[BK05b] W. Böhm and W. Katzenbeisser. On the number of inversions in bimodal permutations. Journal of Statistical Planning and Inference, 135(1):55–63, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000546.

Bhattacharya:2007:CAN

[BK07] Rabi Bhattacharya and Maiying Kong. Consistency and asymptotic normality of the estimated effective doses in bioassay. Journal of Statistical Planning and Inference, 137(3):643– 658, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001315.

Beutner:2009:ORS

[BK09a] E. Beutner and U. Kamps. Order restricted statistical inference for scale parameters based on sequential order statistics. Journal of Statistical Planning and Inference, 139(9):2963–2969, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000287.

Bhowmik:2009:PES

[BK09b] Jahar L. Bhowmik and Maxwell L. King. Parameter estimation in semi-linear models using a maximal invariant likelihood function. Journal of Statistical Planning and Inference, 139(4):1276–1285, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800325X.

Bobotas:2009:STE

[BK09c] Panayiotis Bobotas and Stavros Kourouklis. Strawdermantype estimators for a scale parameter with application to the exponential distribution. Journal of Statistical Planning and Inference, 139(9):3001–3012, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000329.

Bulutoglu:2009:CBC

[BK09d] Dursun A. Bulutoglu and David M. Kaziska. A counterexample to Beder’s conjectures about Hadamard matrices. Journal of Statistical Planning and Inference, 139(9): 3381–3383, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809000317.

Bennett:2001:LSD

[BKLZ01] Frank E. Bennett, Qingde Kang, Jianguo Lei, and Hantao Zhang. Large sets of disjoint pure Mendelsohn triple systems. Journal of Statistical Planning and Inference, 95(1–2): 89–115, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002809.

Behbahani:2008:SST

[BKTR08] M. Behbahani, G. B. Khosrovshahi, and B. Tayfeh-Rezaie. On the spectrum of simple T (2, 3, v) trades. Journal of Statistical Planning and Inference, 138(7):2236–2242, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003837.

Bathke:2009:CMT

[BKZ09] Arne Bathke, Mi-Ok Kim, and Mai Zhou. Combined multiple testing by censored empirical likelihood. Journal of Statistical Planning and Inference, 139(3):814–827, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002619.

Buonaccorsi:2002:BME

[BL02a] John P. Buonaccorsi and Chii-Dean Lin. Berkson measurement error in designed repeated measures studies with random coefficients. Journal of Statistical Planning and Inference, 104(1):53–72, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001185.

Butar:2002:EBE

[BL02b] Ferry Butar Butar and P. Lahiri. Empirical Bayes estimation of several population means and variances under random sampling variances model. Journal of Statistical Planning and Inference, 102(1):59–69, March 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001689.

Butar:2003:MUE

[BL03] Ferry Butar Butar and P. Lahiri. On measures of uncertainty of empirical Bayes small-area estimators. Journal of Statistical Planning and Inference, 112(1–2):63–76, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003233.

Bathke:2005:RPL

[BL05a] Arne Bathke and Dörte Lankowski. Rank procedures for a large number of treatments. Journal of Statistical Planning and Inference, 133(2):223–238, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001740.

Brown:2005:CIT

[BL05b] Lawrence Brown and Xuefeng Li. Confidence intervals for two sample binomial distribution. Journal of Statistical Planning and Inference, 130(1–2):359–375, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400271X.

Buskirk:2005:APK

[BL05c] Trent D. Buskirk and Sharon L. Lohr. Asymptotic properties of kernel density estimation with complex survey data. Journal of Statistical Planning and Inference, 128(1):165–190, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003690.

Byczkowski:2005:PSS

[BL05d] Terri L. Byczkowski and Martin S. Levy. Planning a sample survey with a many-to-many frame structure: an empirical investigation of the effects of frame imperfections on estimates of population totals. Journal of Statistical Planning and Inference, 135(2):498–515, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002411.

Baek:2006:AES

[BL06] Jong-Il Baek and Han-Ying Liang. Asymptotics of estimators in semi-parametric model under NA samples. Journal of Statistical Planning and Inference, 136(10):3362–3382, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805000698.

Balakrishnan:2008:ORS

[BL08a] N. Balakrishnan and T. Li. Ordered ranked set samples and applications to inference. Journal of Statistical Planning and Inference, 138(11):3512–3524, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000943.

Bunouf:2008:BEM

[BL08b] Pierre Bunouf and Bruno Lecoutre. On Bayesian estimators in multistage binomial designs. Journal of Statistical Planning and Inference, 138(12):3915–3926, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001444.

Buzkova:2008:SLL

[BL08c] Petra Buzková and Thomas Lumley. Semiparametric loglinear regression for longitudinal measurements subject to outcome-dependent follow-up. Journal of Statistical Planning and Inference, 138(8):2450–2461, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003904.

Byczkowski:2009:WDI

[BL09] Terri L. Byczkowski and Martin S. Levy. When does an imperfect sampling frame produce more efficient estimators than a perfect frame? Journal of Statistical Planning and Inference, 139(10):3679–3689, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001219.

Blaker:2000:MEL

[Bla00] Helge Blaker. Minimax estimation in linear regression under restrictions. Journal of Statistical Planning and Inference, 90(1):35–55, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001014.

Blanke:2006:ASS

[Bla06] Delphine Blanke. Adaptive sampling schemes for density estimation. Journal of Statistical Planning and Inference, 136(9):2898–2917, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000169.

Bar-Lev:2000:UBC

[BLBR00] Shaul K. Bar-Lev, Daoud Bshouty, and Benjamin Reiser. Upper bounds for coverage probabilities of confidence intervals for nonmonotone parametric functions. Journal of Statistical Planning and Inference, 89(1–2):109–118, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800000999.

Bar-Lev:2006:EDM

[BLL06] Shaul K. Bar-Lev and Zinoviy Landsman. Exponential dispersion models: Second-order minimax estimation of the mean for unknown dispersion parameter. Journal of Statistical Planning and Inference, 136(11):3837–3851, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001357.

Belzunce:2005:TTT

[BLPR05] Félix Belzunce, Xiaohu Li, José-Francisco Pinar, and José-María Ruiz. Tests for the total time on test transform order. Journal of Statistical Planning and Inference, 133(1):111–121, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001107.

Bar-Lev:2005:MGT

[BLSvdDS05] Shaul K. Bar-Lev, Wolfgang Stadje, and Frank A. van der Duyn Schouten. Multinomial group testing models with incomplete identification. Journal of Statistical Planning and Inference, 135(2):384–401, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002435.

Becker:2007:EPV

[BLU07] Niels Becker, Claude Lefèvre, and Sergey Utev. Estimating protective vaccine efficacy from large trials with recruitment. Journal of Statistical Planning and Inference, 137(3):907–914, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001509.

Blume:2009:BSS

[Blu09] Jeffrey D. Blume. Bounding sample size projections for the area under a ROC curve. Journal of Statistical Planning and Inference, 139(3):711–721, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002012.

Berlinet:2000:NSC

[BLV00] Alain Berlinet, Friedrich Liese, and Igor Vajda. Necessary and sufficient conditions for consistency of M-estimates in regression models with general errors. Journal of Statistical Planning and Inference, 89(1–2):243–267, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002189.

Bebbington:2008:ETP

[BLZ08] Mark Bebbington, Chin-Diew Lai, and Ricardas Zitikis. Estimating the turning point of a bathtub-shaped failure distribution. Journal of Statistical Planning and Inference, 138(4):1157–1166, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700273X.

Balakrishnan:2001:OPP

[BM01a] N. Balakrishnan and Jie Mi. Order-preserving property of maximum likelihood estimator. Journal of Statistical Planning and Inference, 98(1–2):89–99, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002421.

Bera:2001:P

[BM01b] Anil K. Bera and Rahul Mukerjee. Preface. Journal of Statistical Planning and Inference, 97(1):1, August 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003414.

Bluskov:2001:NMD

[BM01c] Iliya Bluskov and Spyros S. Magliveras. On the number of mutually disjoint cyclic designs and large sets of designs. Journal of Statistical Planning and Inference, 95(1–2):133– 142, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002834.

Bansal:2002:SSE

[BM02a] Naveen K. Bansal and Klaus J. Miescke. Simultaneous selection and estimation in general linear models. Journal of Statistical Planning and Inference, 104(2):377–390, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002622.

Bayarri:2002:BAD

[BM02b] M. J. Bayarri and A. M. Mayoral. Bayesian analysis and design for comparison of effect-sizes. Journal of Statistical Planning and Inference, 103(1–2):225–243, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002233.

Bayarri:2003:BMS

[BM03] M. J. Bayarri and J. Morales. Bayesian measures of surprise for outlier detection. Journal of Statistical Planning and Inference, 111(1–2):3–22, ???? 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002823.

Bansal:2005:SSE

[BM05a] Naveen K. Bansal and Klaus J. Miescke. Simultaneous selection and estimation of the best treatment with respect to a control in a general linear model. Journal of Statistical Planning and Inference, 129(1–2):387–404, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002010.

Barabesi:2005:SLS

[BM05b] L. Barabesi and M. Marcheselli. Some large-sample results on a modified Monte Carlo integration method. Journal of Statistical Planning and Inference, 135(2):420–432, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002423.

Balakrishnan:2006:P

[BM06a] N. Balakrishnan and Klaus J. Miescke. Preface. Journal of Statistical Planning and Inference, 136(7):1985, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002041.

Bansal:2006:SBT

[BM06b] Naveen K. Bansal and Klaus J. Miescke. On selecting the best treatment in a generalized linear model. Journal of Statistical Planning and Inference, 136(7):2070–2086, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002119.

Bartolucci:2006:NCU

[BM06c] Francesco Bartolucci and Giorgio E. Montanari. A new class of unbiased estimators of the variance of the systematic sample mean. Journal of Statistical Planning and Inference, 136(4):1512–1525, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003891.

Bischoff:2006:ELF

[BM06d] Wolfgang Bischoff and Frank Miller. Efficient lack of fit designs that are optimal to estimate the highest coefficient of a polynomial. Journal of Statistical Planning and Inference, 136(12):4239–4249, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002016.

Bunea:2006:CFM

[BM06e] C. Bunea and T. A. Mazzuchi. Competing failure modes in accelerated life testing. Journal of Statistical Planning and Inference, 136(5):1608–1620, May 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001709.

Bulutoglu:2008:COA

[BM08] D. A. Bulutoglu and F. Margot. Classification of orthogonal arrays by integer programming. Journal of Statistical Planning and Inference, 138(3):654–666, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000481.

Brunner:2002:MNB

[BMP02] Edgar Brunner, Ullrich Munzel, and Madan L. Puri. The multivariate nonparametric Behrens–Fisher problem. Journal of Statistical Planning and Inference, 108(1–2):37–53, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002690.

Bulla:2009:BNE

[BMW09] P. Bulla, P. Muliere, and S. Walker. A Bayesian nonparametric estimator of a multivariate survival function. Journal of Statistical Planning and Inference, 139(10):3639–3648, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001037.

Bednarski:2003:RES

[BN03] T. Bednarski and M. Nowak. Robustness and efficiency of Sasieni-type estimators in the Cox model. Journal of Statistical Planning and Inference, 115(1):261–272, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001155.

Berred:2009:CBO

[BN09] Alexandre Berred and Valery Nevzorov. Characterizations based on order statistics under sampling without replacement. Journal of Statistical Planning and Inference, 139(2):547–557, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002085.

Barakat:2004:TEV

[BNA04] H. M. Barakat, E. M. Nigm, and M. M. Askar. On the trivariate extreme value distributions. Journal of Statistical Planning and Inference, 118(1–2):19–35, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004044.

Balakrishnan:2006:NPB

[BNP06] N. Balakrishnan, H. K. T. Ng, and S. Panchapakesan. A nonparametric procedure based on early failures for selecting the best population using a test for equality. Journal of Statistical Planning and Inference, 136(7):2087–2111, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805002120.

Berred:2005:NCR

[BNW05] A. Berred, V. B. Nevzorov, and S. Wey. Normalizing constants for record values in Archimedean copula processes. Journal of Statistical Planning and Inference, 133(1):159–172, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001065.

Boente:2004:RPE

[BO04] Graciela Boente and Liliana Orellana. Robust plug-in estimators in proportional scatter models. Journal of Statistical Planning and Inference, 122(1–2):95–110, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002386.

Bochkina:2007:WGF

[Boc07] N. A. Bochkina. Wavelet goodness-of-fit test for dependent data. Journal of Statistical Planning and Inference, 137(8):2593–2612, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000134.

Bohm:2002:MLI

[Böh02] W. Böhm. Multivariate Lagrange inversion and the maximum of a persistent random walk. Journal of Statistical Planning and Inference, 101(1–2):23–31, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001495.

Bondell:2008:CFA

[Bon08] Howard D. Bondell. A characteristic function approach to the biased sampling model, with application to robust logistic regression. Journal of Statistical Planning and Inference, 138(3):742–755, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000572.

Bondesson:2009:UBG

[Bon09] Lennart Bondesson. On univariate and bivariate generalized gamma convolutions. Journal of Statistical Planning and Inference, 139(11):3759–3765, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001402.

Borkowf:2000:MCT

[Bor00] Craig B. Borkowf. On multidimensional contingency tables with categories defined by the empirical quantiles of the marginal data. Journal of Statistical Planning and Inference, 91(1):33–51, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001270.

Bordes:2004:NPE

[Bor04] Laurent Bordes. Non-parametric estimation under progressive censoring. Journal of Statistical Planning and Inference, 119(1):171–189, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004147.

Borkowski:2009:DRS

[Bor09] John J. Borkowski. Discussion of “Response surface design evaluation and comparison” by Christine Anderson-Cook, Connie Borror, and Douglas Montgomery. Journal of Statistical Planning and Inference, 139(2):650–652, February 1, 2009.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001614. See [ACBM09b].

Bose:2006:ECS

[Bos06] Sudip Bose. Estimating the center of symmetry: Is it always better to use larger sample sizes? Journal of Statistical Planning and Inference, 136(12):4194–4203, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001412.

Bosq:2007:GLP

[Bos07] Denis Bosq. General linear processes in Hilbert spaces and prediction. Journal of Statistical Planning and Inference, 137 (3):879–894, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001480.

Bosq:2009:NAP

[Bos09] Denis Bosq. A note on asymptotic parametric prediction. Journal of Statistical Planning and Inference, 139(4): 1506–1513, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003522.

Bhatt:2001:ADT

[BP01] M. B. Bhatt and S. R. Patel. On asymptotic distribution for two-truncation parameter family. Journal of Statistical Planning and Inference, 92(1–2):39–46, January 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000152X.

Bhattacharya:2002:NEL

[BP02a] Rabi Bhattacharya and Vic Patrangenaru. Nonparametic estimation of location and dispersion on Riemannian manifolds. Journal of Statistical Planning and Inference, 108(1–2):23–35, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002689.

Brunner:2002:CRS

[BP02b] Edgar Brunner and Madan L. Puri. A class of rankscore tests in factorial designs. Journal of Statistical Planning and Inference, 103(1–2):331–360, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002300.

Babu:2007:RSM

[BP07] Gutti Jogesh Babu and A. R. Padmanabhan. Re-sampling methods for testing for location against unrestricted and ordered alternatives. Journal of Statistical Planning and Inference, 137(11):3261–3267, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700095X.

Brillinger:2004:EDA

[BPAK04] David R. Brillinger, Haiganoush K. Preisler, Alan A. Ager, and John G. Kie. An exploratory data analysis (EDA) of the paths of moving animals. Journal of Statistical Planning and Inference, 122(1–2):43–63, ???? 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002404.

Boente:2009:RTC

[BPR09] Graciela Boente, Ana M. Pires, and Isabel M. Rodrigues. Robust tests for the common principal components model. Journal of Statistical Planning and Inference, 139(4):1332–1347, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003297.

Burdick:2006:PSC

[BQI06] Richard K. Burdick, Jorge Quiroz, and Hariharan K. Iyer. The present status of confidence interval estimation for one-factor random models. Journal of Statistical Planning and Inference, 136(12):4307–4325, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001977.

Bickel:2000:NSS

[BR00] P. J. Bickel and Y. Ritov. Non- and semiparametric statistics: compared and contrasted. Journal of Statistical Planning and Inference, 91(2):209–228, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001798.

Bellhouse:2002:ADM

[BR02] D. R. Bellhouse and J. N. K. Rao. Analysis of domain means in complex surveys. Journal of Statistical Planning and Inference, 102(1):47–58, March 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001677.

Babu:2003:CLD

[BR03a] G. Jogesh Babu and C. R. Rao. Confidence limits to the distance of the true distribution from a misspecified family by bootstrap. Journal of Statistical Planning and Inference, 115(2):471–478, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001763.

Balakrishnan:2003:SEP

[BR03b] N. Balakrishnan and C. R. Rao. Some efficiency properties of best linear unbiased estimators. Journal of Statistical Planning and Inference, 113(2):551–555, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001076.

Bulutoglu:2008:OSD

[BR08] Dursun A. Bulutoglu and Kenneth J. Ryan. E(s^{2})- optimal supersaturated designs with good minimax properties when N is odd. Journal of Statistical Planning and Inference, 138(6):1754–1762, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003151.

Bulutoglu:2009:ONO

[BR09] Dursun A. Bulutoglu and Kenneth J. Ryan. D-optimal and near D-optimal 2^{k} fractional factorial designs of resolution V. Journal of Statistical Planning and Inference, 139(1):16–22, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002255.

Bravo:2006:BTA

[Bra06] Francesco Bravo. Bartlett-type adjustments for empirical discrepancy test statistics. Journal of Statistical Planning and Inference, 136(3):537–554, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003477.

Bai:2003:EAE

[BRCK03] Z. D. Bai, C. R. Rao, M. Chow, and D. Kundu. An efficient algorithm for estimating the parameters of superimposed exponential signals. Journal of Statistical Planning and Inference, 110(1–2):23–34, January 15, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002877.

Belzunce:2003:MPR

[BRdCR03] Félix Belzunce, José-Mariía Ruiz, and Mariía del Carmen Ruiz. Multivariate properties of random vectors of order statistics. Journal of Statistical Planning and Inference, 115(2):413–424, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802001611.

Brillinger:2008:EVC

[Bri08] David R. Brillinger. Extending the volatility concept to point processes. Journal of Statistical Planning and Inference, 138(9):2607–2614, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000530.

Browne:2006:DPP

[Bro06] William J. Browne. Discussion on the paper by Pardoe and Weidner. Journal of Statistical Planning and Inference, 136(4):1462–1465, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004148. See [PW06b].

Bustos:2009:ABR

[BRO^{+}09] Oscar H. Bustos, Marcelo Ruiz, Silvia Ojeda, Ronny Vallejos, and Alejandro C. Frery. Asymptotic behavior of RAestimates in autoregressive 2D processes. Journal of Statistical Planning and Inference, 139(10):3649–3664, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001207.

Bonnecaze:2000:CDC

[BRS00] A. Bonnecaze, E. Rains, and P. Solé. 3-colored 5designs and Z_{4}-codes. Journal of Statistical Planning and Inference, 86(2):349–368, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001172.

Bursztyn:2001:RDE

[BS01] Dizza Bursztyn and David M. Steinberg. Rotation designs for experiments in high-bias situations. Journal of Statistical Planning and Inference, 97(2):399–414, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002445.

Balakrishnan:2004:TCB

[BS04] N. Balakrishnan and A. Stepanov. Two characterizations based on order statistics and records. Journal of Statistical Planning and Inference, 124(2):273–287, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002234.

Balakrishnan:2005:NNO

[BS05a] N. Balakrishnan and A. Stepanov. A note on the number of observations near an order statistic. Journal of Statistical Planning and Inference, 134(1):1–14, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001764.

Banderier:2005:WDN

[BS05b] Cyril Banderier and Sylviane Schwer. Why Delannoy numbers? Journal of Statistical Planning and Inference, 135 (1):40–54, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805000534.

Burgess:2005:ODC

[BS05c] Leonie Burgess and Deborah J. Street. Optimal designs for choice experiments with asymmetric attributes. Journal of Statistical Planning and Inference, 134(1):288–301, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001892.

Bursztyn:2006:CDC

[BS06] Dizza Bursztyn and David M. Steinberg. Comparison of designs for computer experiments. Journal of Statistical Planning and Inference, 136(3):1103–1119, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400343X.

Bieniek:2007:RTR

[BS07a] Milena Bieniek and Dominik Szynal. On k-th record times, record values and their moments. Journal of Statistical Planning and Inference, 137(1):12–22, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805003034.

Bose:2007:OCD

[BS07b] Mausumi Bose and John Stufken. Optimal crossover designs when carryover effects are proportional to direct effects. Journal of Statistical Planning and Inference, 137(11):3291–3302, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000973.

Balakrishnan:2008:APN

[BS08] N. Balakrishnan and A. Stepanov. Asymptotic properties of numbers of near minimum observations under progressive Type-II censoring. Journal of Statistical Planning and Inference, 138(4):1010–1020, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700225X.

Bhattacharya:2009:BIC

[BS09] Sourabh Bhattacharya and Ashis SenGupta. Bayesian inference for circular distributions with unknown normalising constants. Journal of Statistical Planning and Inference, 139(12):4179–4192, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001906.

Bera:2002:SHR

[BSP02] Anil K. Bera, Totok Suprayitno, and Gamini Premaratne. On some heteroskedasticity-robust estimators of variancecovariance matrix of the least-squares estimators. Journal of Statistical Planning and Inference, 108(1–2):121–136, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002744.

Block:2008:NDS

[BSW08] Henry W. Block, Thomas H. Savits, and Jie Wang. Negative dependence and the Simes inequality. Journal of Statistical Planning and Inference, 138(12):4107–4110, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001067.

Bofill:2004:MCE

[BT04] Pau Bofill and Carme Torras. MBMUDs: a combinatorial extension of BIBDs showing good optimality behaviour. Journal of Statistical Planning and Inference, 124(1):185–204, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803001927.

Butucea:2006:NHT

[BT06] Cristina Butucea and Karine Tribouley. Nonparametric homogeneity tests. Journal of Statistical Planning and Inference, 136(3):597–639, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003374.

Balakrishnan:2008:JDP

[BTK08] N. Balakrishnan, Ram C. Tripathi, and Nandini Kannan. On the joint distribution of placement statistics under progressive censoring and applications to precedence test. Journal of Statistical Planning and Inference, 138(5):1314–1324, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002637.

Balakrishnan:2009:NCC

[BTK09] N. Balakrishnan, I. S. Triantafyllou, and M. V. Koutras. Nonparametric control charts based on runs and Wilcoxontype rank-sum statistics. Journal of Statistical Planning and Inference, 139(9):3177–3192, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000548.

Bulinskaya:2007:SAD

[Bul07a] Ekaterina V. Bulinskaya. Some aspects of decision making under uncertainty. Journal of Statistical Planning and Inference, 137(8):2613–2632, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000146.

Bulutoglu:2007:CCO

[Bul07b] Dursun A. Bulutoglu. Cyclicly constructed E(s^{2})-optimal supersaturated designs. Journal of Statistical Planning and Inference, 137(7):2413–2428, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002734.

Buratti:2001:CPR

[Bur01] Marco Buratti. Constructions for point-regular linear spaces. Journal of Statistical Planning and Inference, 94(2):139–146, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002470.

Burch:2007:GCI

[Bur07] Brent D. Burch. Generalized confidence intervals for proportions of total variance in mixed linear models. Journal of Statistical Planning and Inference, 137(7):2394–2404, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806002618.

Burkschat:2008:OES

[Bur08] M. Burkschat. On optimality of extremal schemes in progressive type II censoring. Journal of Statistical Planning and Inference, 138(6):1647–1659, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003096.

Butler:2001:POP

[But01] Neil A. Butler. Properties of optimal polynomial designs under smoothing spline models. Journal of Statistical Planning and Inference, 92(1–2):259–266, January 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001452.

Butler:2005:CET

[But05] Neil A. Butler. Classification of efficient two-level fractional factorial designs of resolution IV or more. Journal of Statistical Planning and Inference, 131(1):145–159, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000333.

Butler:2006:MBR

[But06] Neil A. Butler. On the minimum bias response surface designs of Box and Draper. Journal of Statistical Planning and Inference, 136(9):3221–3230, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004689.

Butler:2007:RTL

[But07] Neil A. Butler. Results for two-level fractional factorial designs of resolution IV or more. Journal of Statistical Planning and Inference, 137(1):317–323, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000279.

Butler:2008:DET

[But08] Neil A. Butler. Defining equations for two-level factorial designs. Journal of Statistical Planning and Inference, 138(10):3157–3163, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000153.

Butler:2009:TLS

[But09] Neil A. Butler. Two-level supersaturated designs for 2k runs and other cases. Journal of Statistical Planning and Inference, 139(1):23–29, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002267.

Braekers:2001:PKG

[BV01] Roel Braekers and Noël Veraverbeke. The partial Koziol– Green model with covariates. Journal of Statistical Planning and Inference, 92(1–2):55–71, January 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001634.

Braekers:2003:TPK

[BV03] Roel Braekers and Noël Veraverbeke. Testing for the partial Koziol–Green model with covariates. Journal of Statistical Planning and Inference, 115(1):181–192, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001441.

Berlinet:2006:ASO

[BV06] A. Berlinet and I. Vajda. On asymptotic sufficiency and optimality of quantizations. Journal of Statistical Planning and Inference, 136(12):4217–4238, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001540.

Balakrishnan:2000:SEE

[BW00a] N. Balakrishnan and Jun Wang. Simple efficient estimation for the three-parameter gamma distribution. Journal of Statistical Planning and Inference, 85(1–2):115–126, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000749.

Brown:2000:ACO

[BW00b] Lawrence D. Brown and Weng Kee Wong. An algorithmic construction of optimal minimax designs for heteroscedastic linear models. Journal of Statistical Planning and Inference, 85(1–2):103–114, April 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000737.

Bedford:2001:NCE

[BW01] David Bedford and Roger M. Whitaker. A new construction for efficient semi-Latin squares. Journal of Statistical Planning and Inference, 98(1–2):287–292, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003037.

Birkes:2003:EML

[BW03] David Birkes and Shaun S. Wulff. Existence of maximum likelihood estimates in normal variance-components models. Journal of Statistical Planning and Inference, 113(1):35–47, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801003020.

Bogacka:2005:NLD

[BW05] Barbara Bogacka and Francis Wright. Non-linear design problem in a chemical kinetic model with non-constant error variance. Journal of Statistical Planning and Inference, 128(2):633–648, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000308.

Bobecka:2007:MMY

[BW07] Konstancja Bobecka and Jacek Wesolowski. Multivariate Matsumoto-Yor property is rather restrictive. Journal of Statistical Planning and Inference, 137(4):1424–1432, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806000863.

Borovskikh:2008:BET

[BW08] Yuri V. Borovskikh and N. C. Weber. Berry–Esseen type bounds for trimmed U-statistics. Journal of Statistical Planning and Inference, 138(4):1059–1071, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002297.

Beder:2009:IGW

[BW09] Jay H. Beder and Jeb F. Willenbring. Invariance of generalized wordlength patterns. Journal of Statistical Planning and Inference, 139(8):2706–2714, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004540.

Bunea:2006:CVS

[BWA06] Florentina Bunea, Marten H. Wegkamp, and Anna Auguste. Consistent variable selection in high dimensional regression via multiple testing. Journal of Statistical Planning and Inference, 136(12):4349–4364, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001576.

Balakrishnan:2007:EISb

[BX07a] N. Balakrishnan and Qihao Xie. Exact inference for a simple step-stress model with Type-I hybrid censored data from the exponential distribution. Journal of Statistical Planning and Inference, 137(11):3268–3290, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000961.

Balakrishnan:2007:EISa

[BX07b] N. Balakrishnan and Qihao Xie. Exact inference for a simple step-stress model with Type-II hybrid censored data from the exponential distribution. Journal of Statistical Planning and Inference, 137(8):2543–2563, August 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000109.

Balakrishnan:2009:AEU

[BY09] N. Balakrishnan and Po Yang. De-aliasing effects using semifoldover techniques. Journal of Statistical Planning and Inference, 139(9):3102–3111, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000494.

Boscardin:2007:RLS

[BYW07] W. John Boscardin, Xiaohong Yan, and Weng Kee Wong. A reanalysis of a longitudinal scleroderma clinical trial using non-ignorable missingness models. Journal of Statistical Planning and Inference, 137(12):3848–3858, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001541.

Buratti:2000:RKT

[BZ00] Marco Buratti and Fulvio Zuanni. The 1-rotational Kirkman triple systems of order 33. Journal of Statistical Planning and Inference, 86(2):369–377, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001184.

Batsidis:2006:SIL

[BZ06] A. Batsidis and K. Zografos. Statistical inference for location and scale of elliptically contoured models with monotone missing data. Journal of Statistical Planning and Inference, 136(8):2606–2629, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004252.

Bai:2009:SAR

[BZZH09] Zhidong Bai, Shurong Zheng, Baoxue Zhang, and Guorong Hu. Statistical analysis for rounded data. Journal of Statistical Planning and Inference, 139(8):2526–2542, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004722.

Claeskens:2000:BLP

[CA00a] Gerda Claeskens and Marc Aerts. Bootstrapping local polynomial estimators in likelihood-based models. Journal of Statistical Planning and Inference, 86(1):63–80, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001548.

Cristobal:2000:NRE

[CA00b] J. A. Cristóbal and J. T. Alcalá. Nonparametric regression estimators for length biased data. Journal of Statistical Planning and Inference, 89(1–2):145–168, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800000926.

Christensen:2003:MPM

[CA03] William F. Christensen and Yasuo Amemiya. Modeling and prediction for multivariate spatial factor analysis. Journal of Statistical Planning and Inference, 115(2):543–564, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001738.

Callegari:2004:RTP

[CA04] Francesca Callegari and Michael G. Akritas. Rank tests for patterned alternatives in two-way non-parametric analysis of variance. Journal of Statistical Planning and Inference, 126(1):1–23, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002726.

Cai:2002:AWE

[Cai02] T. Tony Cai. On adaptive wavelet estimation of a derivative and other related linear inverse problems. Journal of Statistical Planning and Inference, 108(1–2):329–349, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003166.

Cai:2005:OSC

[Cai05] T. Tony Cai. One-sided confidence intervals in discrete distributions. Journal of Statistical Planning and Inference, 131 (1):63–88, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804000679.

Caner:2006:MEN

[Can06] Mehmet Caner. M-estimators with non-standard rates of convergence and weakly dependent data. Journal of Statistical Planning and Inference, 136(4):1207–1219, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003647.

Cao:2001:WSD

[Cao01] Zhenfu Cao. On Whiteman’s and Storer’s difference sets. Journal of Statistical Planning and Inference, 94(2):147–154, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002482.

Cao:2009:ALE

[Cao09] Mingxiang Cao. Φadmissibility for linear estimators on regression coefficients in a general multivariate linear model under balanced loss function. Journal of Statistical Planning and Inference, 139(9):3354–3360, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000809.

Carpenter:2002:ELE

[Car02] Mark Carpenter. Estimation of location extremes within general families of scale mixtures. Journal of Statistical Planning and Inference, 100(2):197–208, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001343.

Causeur:2005:OSC

[Cau05] David Causeur. Optimal sampling from concomitant variables for regression problems. Journal of Statistical Planning and Inference, 128(1):289–301, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003768.

Cavalier:2006:IPN

[Cav06] L. Cavalier. Inverse problems with non-compact operators. Journal of Statistical Planning and Inference, 136(2):390–400, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003039.

Childs:2006:ROS

[CB06] Aaron Childs and N. Balakrishnan. Relations for order statistics from non-identical logistic random variables and assessment of the effect of multiple outliers on the bias of linear estimators. Journal of Statistical Planning and Inference, 136(7):2227–2253, July 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580500220X.

Cressie:2007:SBH

[CBC^{+}07] Noel Cressie, Bruce E. Buxton, Catherine A. Calder, Peter F. Craigmile, Crystal Dong, Nancy J. McMillan, Michele Morara, Thomas J. Santner, Ke Wang, Gregory Young, and Jian Zhang. From sources to biomarkers: a hierarchical Bayesian approach for human exposure modeling. Journal of Statistical Planning and Inference, 137(11):3361–3379, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001024.

Chaudhuri:2005:SSR

[CBD05] Arijit Chaudhuri, Mausumi Bose, and Kajal Dihidar. Samplesize-restrictive adaptive sampling: an application in estimating localized elements. Journal of Statistical Planning and Inference, 134(1):254–267, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002083.

Chaudhuri:2004:AAS

[CBG04] Arijit Chaudhuri, Mausumi Bose, and J. K. Ghosh. An application of adaptive sampling to estimate highly localized population segments. Journal of Statistical Planning and Inference, 121(2):175–189, April 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001174.

Chen:2000:MCB

[CC00] Yuh-Ing Chen and Yunchan Chi. Multiple comparisons between successive treatments for randomly rightcensored survival data. Journal of Statistical Planning and Inference, 83(1):137–151, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000506.

Clyde:2002:CDS

[CC02] Merlise Clyde and Kathryn Chaloner. Constrained design strategies for improving normal approximations in nonlinear regression problems. Journal of Statistical Planning and Inference, 104(1):175–196, May 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002397.

Cheng:2004:TGF

[CC04] K. F. Cheng and L. C. Chen. Testing goodness-of-fit of a logistic regression model with case-control data. Journal of Statistical Planning and Inference, 124(2):409–422, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002076.

Cheng:2006:FLR

[CC06] K. F. Cheng and L. C. Chen. Fitting logistic regression models with contaminated case-control data. Journal of Statistical Planning and Inference, 136(12):4147–4160, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580500203X.

Castellanos:2007:DBP

[CC07a] M. Eugenia Castellanos and Stefano Cabras. A default Bayesian procedure for the generalized Pareto distribution. Journal of Statistical Planning and Inference, 137(2):473–483, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000395.

Chafai:2007:SCA

[CC07b] Djalil Chafaï and Didier Concordet. On the strong consistency of asymptotic M-estimators. Journal of Statistical Planning and Inference, 137(9):2774–2783, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000432.

Chaudhuri:2007:ICT

[CC07c] Arijit Chaudhuri and Tasos C. Christofides. Item count technique in estimating the proportion of people with a sensitive feature. Journal of Statistical Planning and Inference, 137(2):589–593, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000371.

Cheng:2008:LLM

[CC08] K. F. Cheng and C. K. Chu. A local likelihood method for estimating relative risk functions in case-control studies. Journal of Statistical Planning and Inference, 138(12):3733–3748, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001316.

Cai:2009:PMN

[CC09] Weixing Cai and Pinyuen Chen. Partitioning k multivariate normal populations according to equivalence with respect to a standard vector. Journal of Statistical Planning and Inference, 139(7):2227–2234, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003996.

Castillo:2008:SAO

[CCH08] Enrique Castillo, Carmen Castillo, and Ali S. Hadi. Sensitivity analysis in ordered and restricted parameter models. Journal of Statistical Planning and Inference, 138(6):1556–1576, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002996.

Callegari:2003:AIT

[CCL03] Francesca Callegari, Nunzio Cappuccio, and Diego Lubian. Asymptotic inference in time series regressions with a unit root and infinite variance errors. Journal of Statistical Planning and Inference, 116(1):277–303, September 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002367.

Cuadras:2006:PDG

[CCL06] Carles M. Cuadras, Daniel Cuadras, and Younes Lahlou. Principal directions of the general Pareto distribution with applications. Journal of Statistical Planning and Inference, 136(8):2572–2583, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004276.

Chen:2009:CVE

[CCP09] Lu-Hung Chen, Ming-Yen Cheng, and Liang Peng. Conditional variance estimation in heteroscedastic regression models. Journal of Statistical Planning and Inference, 139(2): 236–245, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808001791.

Csaki:2007:VVE

[CCRS07] Endre Csáki, Miklós Csörgő, Zdzisław Rychlik, and Josef Steinebach. On Vervaat and Vervaat-error-type processes for partial sums and renewals. Journal of Statistical Planning and Inference, 137(3):953–966, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001546.

Coolen:2007:NPC

[CCS07] F. P. A. Coolen and P. Coolen-Schrijner. Nonparametric predictive comparison of proportions. Journal of Statistical Planning and Inference, 137(1):23–33, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805003071.

Chen:2005:SSB

[CCW05] Kejun Chen, Zhenfu Cao, and Ruizhong Wei. Supersimple balanced incomplete block designs with block size 4 and index 6. Journal of Statistical Planning and Inference, 133(2):537–554, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001661.

Colbourn:2000:P

[CD00] Charles J. Colbourn and Michel Deza. Preface. Journal of Statistical Planning and Inference, 86(2):285–286, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001342.

Colbourn:2001:MOL

[CD01] Charles J. Colbourn and Jeffrey H. Dinitz. Mutually orthogonal latin squares: a brief survey of constructions. Journal of Statistical Planning and Inference, 95(1–2):9–48, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002767.

Chung:2002:MDV

[CD02] Younshik Chung and Dipak K. Dey. Model determination for the variance component model using reference priors. Journal of Statistical Planning and Inference, 100(1):49–65, January 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801000908.

Chen:2003:VSM

[CD03a] Ming-Hui Chen and Dipak K. Dey. Variable selection for multivariate logistic regression models. Journal of Statistical Planning and Inference, 111(1–2):37–55, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002847.

Chu:2003:IMA

[CD03b] C. K. Chu and Wen-Shuenn Deng. An interpolation method for adapting to sparse design in multivariate nonparametric regression. Journal of Statistical Planning and Inference, 116(1):91–111, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802001842.

Chakraborty:2006:SPC

[CD06] Subrata Chakraborty and Kishore K. Das. On some properties of a class of weighted quasi-binomial distributions. Journal of Statistical Planning and Inference, 136(1):159–182, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002940.

Cheng:2007:TGC

[CD07a] Yebin Cheng and Jan G. De Gooijer. On the u th geometric conditional quantile. Journal of Statistical Planning and Inference, 137(6):1914–1930, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001716.

Cicalese:2007:PMA

[CD07b] Ferdinando Cicalese and Christian Deppe. Perfect minimally adaptive q-ary search with unreliable tests. Journal of Statistical Planning and Inference, 137(1):162–175, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580500296X.

Cordeiro:2009:TGL

[CdA09] Gauss M. Cordeiro and Marinho G. de Andrade. Transformed generalized linear models. Journal of Statistical Planning and Inference, 139(9):2970–2987, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000299.

Couallier:2009:SIT

[CDB^{+}09] Vincent Couallier, Laurent Denis, Vilijandas Bagdonavicius, Mikhail Nikulin, and Léo Gerville-Réache. Statistical inference through AFT model for biotechnical systems. Journal of Statistical Planning and Inference, 139(5):1649–1656, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002413.

Chai:2001:OBD

[CDD01] Feng-Shun Chai, Ashish Das, and Aloke Dey. A-optimal block designs for parallel line assays. Journal of Statistical Planning and Inference, 96(2):403–414, July 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002147.

Croux:2006:TIM

[CDH06] Christophe Croux, Geert Dhaene, and Dirk Hoorelbeke. Testing the information matrix equality with robust estimators. Journal of Statistical Planning and Inference, 136(10):3583–3613, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000844.

Cabrera:2004:MTA

[CF04] Javier Cabrera and Luisa Turrin Fernholz. Multivariate targeting with applications to ellipse estimation. Journal of Statistical Planning and Inference, 122(1–2):79–94, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002325.

Chatterjee:2005:UFD

[CFQ05] Kashinath Chatterjee, Kai-Tai Fang, and Hong Qin. Uniformity in factorial designs with mixed levels. Journal of Statistical Planning and Inference, 128(2):593–607, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580400028X.

Conde:2005:CRO

[CFS05] David Conde, Miguel A. Fernández, and Bonifacio Salvador. A classification rule for ordered exponential populations. Journal of Statistical Planning and Inference, 135(2):339–356, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002319.

Carbon:2007:KRE

[CFT07] Michel Carbon, Christian Francq, and Lanh Tat Tran. Kernel regression estimation for random fields. Journal of Statistical Planning and Inference, 137(3):778–798, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001406.

Coad:2000:CCI

[CG00] D. S. Coad and Z. Govindarajulu. Corrected confidence intervals following a sequential adaptive clinical trial with binary responses. Journal of Statistical Planning and Inference, 91(1):53–64, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001294.

Calle:2001:NBE

[CG01] M. Luz Calle and Guadalupe Gómez. Nonparametric Bayesian estimation from interval-censored data using Monte Carlo methods. Journal of Statistical Planning and Inference, 98(1–2):73–87, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003207.

Chatterjee:2003:CSD

[CG03] Kashinath Chatterjee and Sudhir Gupta. Construction of supersaturated designs involving s-level factors. Journal of Statistical Planning and Inference, 113(2):589–595, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580200109X.

Chakrabarti:2006:GBG

[CG06] Arijit Chakrabarti and Jayanta K. Ghosh. A generalization of BIC for the general exponential family. Journal of Statistical Planning and Inference, 136(9):2847–2872, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000388.

Choi:2008:CRC

[CG08a] Kuey Chung Choi and Sudhir Gupta. Confounded rowcolumn designs. Journal of Statistical Planning and Inference, 138(1):196–202, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807002091.

Cui:2008:EBV

[CG08b] Wen Cui and Edward I. George. Empirical Bayes vs. fully Bayes variable selection. Journal of Statistical Planning and Inference, 138(4):888–900, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001425.

Chambaz:2009:MDL

[CGG09] A. Chambaz, A. Garivier, and E. Gassiat. A minimum description length approach to hidden Markov models with Poisson and Gaussian emissions. Application to order identification. Journal of Statistical Planning and Inference, 139(3):962–977, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002851.

Cohen:2006:IIT

[CGPP06] Serge Cohen, Xavier Guyon, Olivier Perrin, and Monique Pontier. Identification of an isometric transformation of the standard Brownian sheet. Journal of Statistical Planning and Inference, 136(4):1317–1330, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003908.

Chaudhry:2002:FBR

[CGS02] Ghulam R. Chaudhry, Malcolm Greig, and Jennifer Seberry. On the (v, 5, λ)-family of Bhaskar Rao designs. Journal of Statistical Planning and Inference, 106(1–2):303–327, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002203.

Chen:2007:MCF

[CGY07] Lihua Chen, Panayotis Giannakouros, and Yuhong Yang. Model combining in factorial data analysis. Journal of Statistical Planning and Inference, 137(9):2920–2934, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806002850.

Chen:2000:SRC

[CGZ00] K. Chen, G. Ge, and L. Zhu. Starters and related codes. Journal of Statistical Planning and Inference, 86(2):379–395, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899001196.

Carriere:2000:CDT

[CH00] K. C. Carriere and Rong Huang. Crossover designs for two-treatment clinical trials. Journal of Statistical Planning and Inference, 87(1):125–134, May 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001858.

Chang:2001:GSS

[CH01] Yi-Ping Chang and Wen-Tao Huang. Generalized subset selection procedures under heteroscedasticity. Journal of Statistical Planning and Inference, 98(1–2):239–258, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003050.

Causeur:2005:DEB

[CH05a] David Causeur and François Husson. A 2-dimensional extension of the Bradley-Terry model for paired comparisons. Journal of Statistical Planning and Inference, 135(2):245–259, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002393.

Chou:2005:NCB

[CH05b] Chao-Wei Chou and Wen-Jang Huang. A note on characterizations of the bivariate gamma distribution. Journal of Statistical Planning and Inference, 128(1):259–269, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803003276.

Cui:2008:NPO

[CH08] Lirong Cui and Alan G. Hawkes. A note on the proof for the optimal consecutive-k-out-of-n:G line for n ≤ 2k. Journal of Statistical Planning and Inference, 138(5):1516–1520, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580700290X.

Chaudhuri:2000:ACM

[Cha00] Probal Chaudhuri. Asymptotic consistency of median regression trees. Journal of Statistical Planning and Inference, 91(2):229–238, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001804.

Chang:2001:BRI

[Cha01a] Ted Chang. Book review: Interpolation of spatial data: some theory for kriging: Michael L. Stein; Springer, New York, 1999, 247 + xvii pages, ISBN 0-387-98629-4. Journal of Statistical Planning and Inference, 96(2):431–435, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003098.

Chang:2001:EBD

[Cha01b] Yanxun Chang. On the existence of BIB designs with large λ. Journal of Statistical Planning and Inference, 94(2):155– 165, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002494.

Chaudhuri:2001:URR

[Cha01c] Arijit Chaudhuri. Using randomized response from a complex survey to estimate a sensitive proportion in a dichotomous finite population. Journal of Statistical Planning and Inference, 94(1):37–42, March 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000210X.

Chang:2002:BRG

[Cha02a] Ted Chang. Book review: Geometrical foundations of asymptotic inference: Robert E. Kass and Paul W. Vos; Wiley, New York, 1997, pp. 355 + xii, ISBN: 0-471-82668-5. Journal of Statistical Planning and Inference, 100(1):81–84, January 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001069.

Charalambides:2002:RNF

[Cha02b] Ch. A. Charalambides. The rook numbers of Ferrers boards and the related restricted permutation numbers. Journal of Statistical Planning and Inference, 101(1–2):33–48, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001513.

Chaubey:2002:EIG

[Cha02c] Yogendra P. Chaubey. Estimation in inverse Gaussian regression: comparison of asymptotic and bootstrap distributions. Journal of Statistical Planning and Inference, 100(2):135–143, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001288.

Chaganty:2003:AGC

[Cha03a] N. Rao Chaganty. Analysis of growth curves with patterned correlation matrices using quasi-least squares. Journal of Statistical Planning and Inference, 117(1):123–139, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003622.

Chakraborty:2003:MQR

[Cha03b] Biman Chakraborty. On multivariate quantile regression. Journal of Statistical Planning and Inference, 110(1–2):109– 132, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801002774.

Charalambides:2005:MCD

[Cha05] Ch. A. Charalambides. Moments of a class of discrete q-distributions. Journal of Statistical Planning and Inference, 135(1):64–76, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000558.

Charalambides:2007:DRS

[Cha07] Ch. A. Charalambides. Distribution of record statistics in a geometrically increasing population. Journal of Statistical Planning and Inference, 137(7):2214–2225, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002448.

Challenor:2009:CPC

[Cha09] P. G. Challenor. Comments by P. Challenor on Goldstein and Rougier. Journal of Statistical Planning and Inference, 139(3):1248, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003340.

Chen:2000:SCL

[Che00a] X. R. Chen. Studies on consistency of LSE in China. Journal of Statistical Planning and Inference, 88(2):181–188, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580000077X.

Chen:2000:RSS

[Che00b] Zehua Chen. On ranked-set sample quantiles and their applications. Journal of Statistical Planning and Inference, 83(1):125–135, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899000713.

Chen:2001:RSS

[Che01] Zehua Chen. Ranked-set sampling with regression-type estimators. Journal of Statistical Planning and Inference, 92(1–2):181–192, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001403.

Chen:2002:SPE

[Che02] Pinyuen Chen. A selection procedure for estimating the number of signal components. Journal of Statistical Planning and Inference, 105(2):299–311, July 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002683.

Chen:2003:CRA

[Che03a] Zehua Chen. Component reliability analysis of k-out-of-n systems with censored data. Journal of Statistical Planning and Inference, 116(1):305–315, September 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001799.

Cheng:2003:AEE

[Che03b] Yi Cheng. An application of Edgeworth expansion in Bayesian inferences: optimal sample sizes in clinical trials. Journal of Statistical Planning and Inference, 115(2):379–396, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001623.

Chen:2004:CNE

[Che04a] Ling Chen. A conservative, nonparametric estimator for the 5th percentile of the species sensitivity distributions. Journal of Statistical Planning and Inference, 123(2):243–258, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001484.

Cheng:2004:WSU

[Che04b] Fuxia Cheng. Weak and strong uniform consistency of a kernel error density estimator in nonparametric regression. Journal of Statistical Planning and Inference, 119(1):95–107, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004172.

Cheng:2004:TEL

[Che04c] R. C. H. Cheng. Threshold estimation in the loggamma model. Journal of Statistical Planning and Inference, 119(2):241–262, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200486X.

Chernoff:2004:ITE

[Che04d] Herman Chernoff. Information, for testing the equality of two probabilities, from the margins of the 2×2 table. Journal of Statistical Planning and Inference, 121(2):209–214, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001356.

Cheng:2005:ADE

[Che05] Fuxia Cheng. Asymptotic distributions of error density and distribution function estimators in nonparametric regression. Journal of Statistical Planning and Inference, 128(2):327–349, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000205.

Chebana:2006:FAN

[Che06a] Fateh Chebana. Functional asymptotic normality of the L_{2}-deviation of the kernel density estimation indexed by classes of weight functions. Journal of Statistical Planning and Inference, 136(8):2470–2505, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004598.

Chen:2006:CIN

[Che06b] Pinyuen Chen. A confidence interval for the number of principal components. Journal of Statistical Planning and Inference, 136(8):2630–2639, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004185.

Chen:2006:ARB

[Che06c] Xun Chen. The adjustment of random baseline measurements in treatment effect estimation. Journal of Statistical Planning and Inference, 136(12):4161–4175, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805002466.

Chen:2006:PEB

[Che06d] Yung-Pin Chen. The power of Efron’s biased coin design. Journal of Statistical Planning and Inference, 136(6):1824– 1835, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805001874.

Chen:2007:NRB

[Che07] Yung-Pin Chen. A note on the relationship between Spearman’s ρ and Kendall’s τ for extreme order statistics. Journal of Statistical Planning and Inference, 137(7):2165–2171, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002096.

Chen:2008:SBB

[Che08] Lee-Shen Chen. Selection of the best Bernoulli population provided it is better than a control: an empirical Bayes approach. Journal of Statistical Planning and Inference, 138(8):2366–2379, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003746.

Cheng:2009:SAI

[Che09] Guang Cheng. Semiparametric additive isotonic regression. Journal of Statistical Planning and Inference, 139(6): 1980–1991, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003650.

Chicken:2003:BTW

[Chi03] Eric Chicken. Block thresholding and wavelet estimation for nonequispaced samples. Journal of Statistical Planning and Inference, 116(1):113–129, September 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002380.

Chang:2001:ODD

[CHLY01] Fu-Chuen Chang, Mong-Na Lo Huang, Dennis K. J. Lin, and Huie-Ching Yang. Optimal designs for dual response polynomial regression models. Journal of Statistical Planning and Inference, 93(1–2):309–322, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001622.

Choy:2001:ODS

[Cho01] Kokyo Choy. Outlier detection for stationary time series. Journal of Statistical Planning and Inference, 99(2):111–127, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801000817.

Cho:2007:SRE

[Cho07a] Hokwon Cho. Sequential risk-efficient estimation for the ratio of two binomial proportions. Journal of Statistical Planning and Inference, 137(7):2336–2346, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002606.

Choi:2007:APC

[Cho07b] Taeryon Choi. Alternative posterior consistency results in nonparametric binary regression using Gaussian process priors. Journal of Statistical Planning and Inference, 137(9):2975–2983, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000407.

Choudhary:2008:TIA

[Cho08] Pankaj K. Choudhary. A tolerance interval approach for assessment of agreement in method comparison studies with repeated measurements. Journal of Statistical Planning and Inference, 138(4):1102–1115, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002248.

Cassart:2008:ODF

[CHP08] Delphine Cassart, Marc Hallin, and Davy Paindaveine. Optimal detection of Fechner-asymmetry. Journal of Statistical Planning and Inference, 138(8):2499–2525, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700376X.

Christofides:2004:SAR

[Chr04] Tasos C. Christofides. U-statistics on associated random variables. Journal of Statistical Planning and Inference, 119(1):1–15, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802004184.

Christofides:2005:RRS

[Chr05] Tasos C. Christofides. Randomized response in stratified sampling. Journal of Statistical Planning and Inference, 128(1):303–310, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003744.

Chu:2001:NCS

[Chu01] Wensong Chu. New constructions of simple and indecomposable 1-factorizations of complete multigraphs. Journal of Statistical Planning and Inference, 94(2):181–196, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002512.

Chang:2001:SCR

[cIC01] Yuan chin Ivan Chang. Sequential confidence regions of generalized linear models with adaptive designs. Journal of Statistical Planning and Inference, 93(1–2):277–293, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002007.

Chen:2008:NCM

[CIC08] Ming-Hui Chen, Joseph G. Ibrahim, and Yueh-Yun Chi. A new class of mixture models for differential gene expression in DNA microarray data. Journal of Statistical Planning and Inference, 138(2):387–404, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002509.

Chang:2004:SEV

[cICY04] Yuan chin Ivan Chang and Zhiliang Ying. Sequential estimation in variable length computerized adaptive testing. Journal of Statistical Planning and Inference, 121(2):249–264, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001198.

Chu:2007:EFM

[CIPS07] Ka Lok Chu, Jarkko Isotalo, Simo Puntanen, and George P. H. Styan. The efficiency factorization multiplier for the Watson efficiency in partitioned linear models: Some examples and a literature review. Journal of Statistical Planning and Inference, 137(11):3336–3351, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001000.

Chen:2000:PPD

[CIS00] Ming-Hui Chen, Joseph G. Ibrahim, and Qi-Man Shao. Power prior distributions for generalized linear models. Journal of Statistical Planning and Inference, 84(1–2):121–137, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001408.

Chen:2003:PEM

[CISW03] Ming-Hui Chen, Joseph G. Ibrahim, Qi-Man Shao, and Robert E. Weiss. Prior elicitation for model selection and estimation in generalized linear mixed models. Journal of Statistical Planning and Inference, 111(1–2):57–76, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002859.

Ip:2005:SAE

[cIWsL05] Wai cheung Ip, Heung Wong, and Jin shan Liu. Sufficient and admissible estimators in general multivariate linear model. Journal of Statistical Planning and Inference, 135(2):371–383, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002332.

Cizek:2006:LTS

[Cíz06] Pavel Cízek. Least trimmed squares in nonlinear regression under dependence. Journal of Statistical Planning and Inference, 136(11):3967–3988, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001369.

Coleman:2004:BCS

[CJK04] Daniel Coleman, Yousceek Jeong, and Robert W. Keener. A Bayesian criterion for selecting super saturated screening designs. Journal of Statistical Planning and Inference, 120(1–2):137–153, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802005025.

Cutillo:2008:LPM

[CJRV08] Luisa Cutillo, Yoon Young Jung, Fabrizio Ruggeri, and Brani Vidakovic. Larger posterior mode wavelet thresholding and applications. Journal of Statistical Planning and Inference, 138(12):3758–3773, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800133X.

Cramer:2000:REF

[CK00] Erhard Cramer and Udo Kamps. Relations for expectations of functions of generalized order statistics. Journal of Statistical Planning and Inference, 89(1–2):79–89, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800000744.

Chen:2001:WHL

[CK01] Wende Chen and Torleiv Kløve. Weight hierarchies of linear codes of dimension 3. Journal of Statistical Planning and Inference, 94(2):167–179, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002500.

Chen:2005:MLR

[CK05] Jiahua Chen and John D. Kalbfleisch. Modified likelihood ratio test in finite mixture models with a structural parameter. Journal of Statistical Planning and Inference, 129(1–2):93–107, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001569.

Calinski:2008:AEA

[CK08] Tadeusz Caliński and Sanpei Kageyama. On the analysis of experiments in affine resolvable designs. Journal of Statistical Planning and Inference, 138(11):3350–3356, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000827.

Cheng:2000:ODE

[CKM00] R. C. H. Cheng, J. P. C. Kleijnen, and V. B. Melas. Optimal design of experiments with simulation models of nearly saturated queues. Journal of Statistical Planning and Inference, 85(1–2):19–26, April 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000610.

Chatzopoulos:2006:SOS

[CKM06] St. A. Chatzopoulos and F. Kolyva-Machera. Some D-optimal saturated designs for 3 × m_{2} × m_{3} factorials. Journal of Statistical Planning and Inference, 136(8):2820–2830, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004239.

Colbourn:2002:OAS

[CKMS02] C. J. Colbourn, D. L. Kreher, J. P. McSorley, and D. R. Stinson. Orthogonal arrays of strength three from regular 3-wise balanced designs. Journal of Statistical Planning and Inference, 100(2):191–195, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001331.

Chang:2007:BMC

[CKS07a] Iva Chang, Alan Krinik, and Randall J. Swift. Birthmultiple catastrophe processes. Journal of Statistical Planning and Inference, 137(5):1544–1559, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002278.

Cohen:2007:SVS

[CKS07b] Arthur Cohen, John Kolassa, and Harold B. Sackrowitz. A smooth version of the step-up procedure for multiple tests of hypotheses. Journal of Statistical Planning and Inference, 137(11):3352–3360, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001012.

Cook:2001:SCE

[CL01] R. J. Cook and J. F. Lawless. Some comments on efficiency gains from auxiliary information for right-censored data. Journal of Statistical Planning and Inference, 96(1):191–202, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003359.

Chang:2002:ODG

[CL02] Fu-Chuen Chang and Ching-Fu Lay. Optimal designs for a growth curve model. Journal of Statistical Planning and Inference, 104(2):427–438, June 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002555.

Chang:2003:FIP

[CL03] Yanxun Chang and Giovanni Lo Faro. The flower intersection problem for Kirkman triple systems. Journal of Statistical Planning and Inference, 110(1–2):159–177, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002828.

Cho:2006:BCR

[CL06] Kwanho Cho and Wei-Yin Loh. Bias and convergence rate of the coverage probability of prediction intervals in Box– Cox transformed linear models. Journal of Statistical Planning and Inference, 136(10):3614–3624, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000820.

Chen:2008:OML

[CL08] Jie Chen and Min-Qian Liu. Optimal mixed-level k-circulant supersaturated designs. Journal of Statistical Planning and Inference, 138(12):4151–4157, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001055.

Chang:2009:FST

[CL09a] Yanxun Chang and Giovanni Lo Faro. The fine structures of three pairwise distinct Latin squares. Journal of Statistical Planning and Inference, 139(9):2863–2869, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000500.

Consonni:2009:BMC

[CL09b] Guido Consonni and Monia Lupparelli. Bayesian model comparison based on expected posterior priors for discrete decomposable graphical models. Journal of Statistical Planning and Inference, 139(12):4154–4164, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001840.

Cao:2009:CBS

[CLA09] Jing Cao, J. Jack Lee, and Susan Alber. Comparison of Bayesian sample size criteria: ACC, ALC, and WOC. Journal of Statistical Planning and Inference, 139(12):4111–4122, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001670.

Chiquet:2009:PDM

[CLE09] Julien Chiquet, Nikolaos Limnios, and Mohamed Eid. Piecewise deterministic Markov processes applied to fatigue crack growth modelling. Journal of Statistical Planning and Inference, 139(5):1657–1667, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002425.

Chen:2001:OPE

[CLHY01] Di Chen, Jye-Chyi Lu, Xiaoming Huo, and Ming Yin. Optimum percentile estimating equations for nonlinear random coefficient models. Journal of Statistical Planning and Inference, 97(2):275–292, September 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002196.

Chen:2006:SRB

[CLLZ06] Bao-Jiang Chen, Peng-Fei Li, Min-Qian Liu, and Run-Chu Zhang. Some results on blocked regular 2-level fractional factorial designs with clear effects. Journal of Statistical Planning and Inference, 136(12):4436–4449, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001990.

Chiu:2005:ATB

[CLR05] Grace Chiu, Richard Lockhart, and Richard Routledge. Asymptotic theory for bent-cable regression-the basic case. Journal of Statistical Planning and Inference, 127(1–2):143– 156, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803003148.

Chu:2009:UCE

[CLS09] Pei-Ling Chu, Yong Lin, and Weichung Joe Shih. Unifying CRM and EWOC designs for phase I cancer clinical trials. Journal of Statistical Planning and Inference, 139 (3):1146–1163, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002978.

Collins:2007:KS

[CLT07] David Collins, Tamás Lengyel, and Gregory P. Tollisen. Keeping scores. Journal of Statistical Planning and Inference, 137(7):2208–2213, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002217.

Chai:2000:ODN

[CM00a] Feng-Shun Chai and Dibyen Majumdar. Optimal designs for nearest-neighbor analysis. Journal of Statistical Planning and Inference, 86(1):265–275, April 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900155X.

Czado:2000:NLG

[CM00b] Claudia Czado and Axel Munk. Noncanonical links in generalized linear models — when is the effort justified? Journal of Statistical Planning and Inference, 87(2):317–345, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001950.

Cheng:2003:RFF

[CM03] Ching-Shui Cheng and Rahul Mukerjee. On regular-fractional factorial experiments in row-column designs. Journal of Statistical Planning and Inference, 114(1–2):3–20, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004597.

Cooke:2006:CRC

[CMN06] Roger M. Cooke and Oswald Morales-Napoles. Competing risk and the Cox proportional hazard model. Journal of Statistical Planning and Inference, 136(5):1621–1637, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001710.

Critchley:2002:PPG

[CMS02] Frank Critchley, Paul Marriott, and Mark Salmon. On preferred point geometry in statistics. Journal of Statistical Planning and Inference, 102(2):229–245, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100115X.

Chai:2002:FRO

[CMyS02] Feng Shun Chai, Rahul Mukerjee, and Chung yi Suen. Further results on orthogonal array plus one run plans. Journal of Statistical Planning and Inference, 106(1–2):287–301, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002197.

Chaganty:2002:AML

[CN02a] N. Rao Chaganty and Dayanand N. Naik. Analysis of multivariate longitudinal data using quasi-least squares. Journal of Statistical Planning and Inference, 103(1–2):421–436, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580100235X.

Chatterjee:2002:GTT

[CN02b] Kashinath Chatterjee and Giri Narasimhan. Graph-theoretic techniques in D-optimal design problems. Journal of Statistical Planning and Inference, 102(2):377–387, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000970.

Choudhary:2005:SIC

[CN05] Pankaj K. Choudhary and H. N. Nagaraja. Selecting the instrument closest to a gold standard. Journal of Statistical Planning and Inference, 129(1–2):229–237, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001855.

Choudhary:2007:TAA

[CN07] Pankaj K. Choudhary and H. N. Nagaraja. Tests for assessment of agreement using probability criteria. Journal of Statistical Planning and Inference, 137(1):279–290, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805003083.

Cubedo:2002:HTM

[CO02] M. Cubedo and J. M. Oller. Hypothesis testing: a model selection approach. Journal of Statistical Planning and Inference, 108(1–2):3–21, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002677.

Cottle:2008:EOP

[CO08] Richard W. Cottle and Ingram Olkin. Estimating ordered parameters by linear programming. Journal of Statistical Planning and Inference, 138(9):2622–2633, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000554.

Colbourn:2001:MWP

[Col01] Charles J. Colbourn. Minimum weights of point codes of Steiner triple systems. Journal of Statistical Planning and Inference, 95(1–2):161–166, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002858.

Conniffe:2001:STW

[Con01] Denis Conniffe. Score tests when a nuisance parameter is unidentified under the null hypothesis. Journal of Statistical Planning and Inference, 97(1):67–83, August 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003463.

Conaway:2002:BRO

[Con02a] Mark R. Conaway. Book review: Ordinal Data Modeling: Valen E. Johnson and James H. Albert; Springer, New York, 1999, 258pp., ISBN 0-387-98718-5. Journal of Statistical Planning and Inference, 100(1):87–88, January 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001057.

Conaway:2002:BRR

[Con02b] Mark R. Conaway. Book review: Regression Analysis of Count Data: A. Colin Cameron and P. K. Trivedi; Cambridge University Press, Cambridge, UK, 1998, 411pp., ISBN 0-521-63201-3. Journal of Statistical Planning and Inference, 100(1):85–86, January 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001045.

Conti:2004:BAB

[Con04] Pier Luigi Conti. Bootstrap approximations for Bayesian analysis of Geo/G/1 discrete-time queueing models. Journal of Statistical Planning and Inference, 120(1–2):65–84, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004998.

Congdon:2008:SSE

[Con08] Peter Congdon. A spatial structural equation model for health outcomes. Journal of Statistical Planning and Inference, 138(7):2090–2105, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003369.

Congdon:2009:AAP

[Con09] Peter Congdon. Adaptive autoregressive priors for area and time structured mortality data. Journal of Statistical Planning and Inference, 139(9):2870–2884, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000111.

Cowles:2003:EMF

[Cow03] Mary Kathryn Cowles. Efficient model-fitting and modelcomparison for high-dimensional Bayesian geostatistical models. Journal of Statistical Planning and Inference, 112(1– 2):221–239, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580200335X.

Cox:2003:PMD

[Cox03] Lawrence H. Cox. On properties of multi-dimensional statistical tables. Journal of Statistical Planning and Inference, 117(2):251–273, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003920.

Chaudhuri:2002:CAM

[CP02a] Arijit Chaudhuri and Sanghamitra Pal. On certain alternative mean square error estimators in complex survey sampling. Journal of Statistical Planning and Inference, 104(2):363–375, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002476.

Cressie:2002:MCL

[CP02b] Noel Cressie and Leandro Pardo. Model checking in loglinear models using ϕ-divergences and MLEs. Journal of Statistical Planning and Inference, 103(1–2):437–453, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002361.

Chaudhuri:2003:VCS

[CP03] Arijit Chaudhuri and Sanghamitra Pal. On a version of cluster sampling and its practical use. Journal of Statistical Planning and Inference, 113(1):25–34, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201002048.

Chaudhuri:2005:BMS

[CP05a] Sanjay Chaudhuri and Michael D. Perlman. On the bias and mean-square error of order-restricted maximum likelihood estimators. Journal of Statistical Planning and Inference, 130(1–2):229–250, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002630.

Clayton:2005:EAA

[CP05b] Murray K. Clayton and A. John Petkau. Evaluation of asymptotic approximations for a two-stage Bernoulli bandit. Journal of Statistical Planning and Inference, 130(1–2):133–148, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002587.

Chaudhuri:2007:CEM

[CP07] Sanjay Chaudhuri and Michael. D. Perlman. Consistent estimation of the minimum normal mean under the tree-order restriction. Journal of Statistical Planning and Inference, 137(11):3317–3335, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000997.

Chen:2008:THH

[CQ08] Jiahua Chen and Yongsong Qin. Test for homogeneity in Hardy-Weinberg normal mixture model. Journal of Statistical Planning and Inference, 138(12):3774–3788, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001328.

Crato:2002:SPS

[CR02] Nuno Crato and Bonnie K. Ray. Semi-parametric smoothing estimators for long-memory processes with added noise. Journal of Statistical Planning and Inference, 105(2):283–297, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002750.

Charalambides:2008:DMR

[CR08] Ch. A. Charalambides and Tomasz Rychlik. Distributions and moments of record values in a sequence of maximally dependent random variables. Journal of Statistical Planning and Inference, 138(8):2253–2266, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800044X.

Conniffe:2000:ADM

[CS00] Denis Conniffe and John E. Spencer. Approximating the distribution of the maximum partial sum of normal deviates. Journal of Statistical Planning and Inference, 88(1):19–27, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899002037.

Chaubey:2002:SEM

[CS02a] Yogendra P. Chaubey and Pranab K. Sen. Smooth estimation of multivariate survival and density functions. Journal of Statistical Planning and Inference, 103(1–2):361–376, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002312.

Chaudhary:2002:RAU

[CS02b] Mohammad Ashraf Chaudhary and Pranab Kumar Sen. Reconciliation of asymptotics for unequal probability sampling without replacement. Journal of Statistical Planning and Inference, 102(1):71–81, March 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001690.

Cohen:2002:DLR

[CS02c] Arthur Cohen and H. B. Sackrowitz. Directed likelihood ratio tests for ordered alternatives and inference results for the star-shaped restriction. Journal of Statistical Planning and Inference, 103(1–2):181–189, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002208.

Cohen:2002:IMS

[CS02d] Arthur Cohen and H. B. Sackrowitz. Inference for the model of several treatments and a control. Journal of Statistical Planning and Inference, 107(1–2):89–101, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002458.

Castelo:2003:CMT

[CS03a] Robert Castelo and Arno Siebes. A characterization of moral transitive acyclic directed graph Markov models as labeled trees. Journal of Statistical Planning and Inference, 115(1):235–259, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200143X.

Chen:2003:TLR

[CS03b] Zehua Chen and Liang Shen. Two-layer ranked set sampling with concomitant variables. Journal of Statistical Planning and Inference, 115(1):45–57, July 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001453.

Chen:2004:MTR

[CS04a] Jie Chen and Sanat K. Sarkar. Multiple testing of response rates with a control: a Bayesian stepwise approach. Journal of Statistical Planning and Inference, 125(1–2):3–16, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000527.

Cohen:2004:MPM

[CS04b] Arthur Cohen and H. B. Sackrowitz. Monotonicity properties of multiple endpoint testing procedures. Journal of Statistical Planning and Inference, 125(1–2):17–30, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000539.

Chaudhuri:2005:OVC

[CS05] Arijit Chaudhuri and Amitava Saha. Optional versus compulsory randomized response techniques in complex surveys. Journal of Statistical Planning and Inference, 135(2):516–527, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002368.

Cai:2006:MSC

[CS06] Gengqian Cai and Sanat K. Sarkar. Modified Simes’ critical values under positive dependence. Journal of Statistical Planning and Inference, 136(12):4129–4146, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001539.

Charnigo:2008:THD

[CS08] Richard Charnigo and Jiayang Sun. Testing homogeneity in discrete mixtures. Journal of Statistical Planning and Inference, 138(5):1368–1388, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002686.

Csaki:2002:SRT

[Csá02] Endre Csáki. Some results for two-dimensional random walk, II. Additive functionals. Journal of Statistical Planning and Inference, 101(1–2):49–56, February 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001537.

Csaki:2005:IVH

[Csá05] Endre Csáki. István Vincze (1912–1999) and his contribution to lattice path combinatorics and statistics. Journal of Statistical Planning and Inference, 135(1):3–17, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000510.

Chatterjee:2008:SDH

[CSL08] Kashinath Chatterjee, Angshuman Sarkar, and Dennis K. J. Lin. Supersaturated designs with high searching probability. Journal of Statistical Planning and Inference, 138(1):272–277, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002108.

Chongcharoen:2002:PSO

[CSW02] Samruam Chongcharoen, Bahadur Singh, and F. T. Wright. Powers of some one-sided multivariate tests with the population covariance matrix known up to a multiplicative constant. Journal of Statistical Planning and Inference, 107(1–2):103–121, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200246X.

Cohen:2008:UIA

[CSX08] Arthur Cohen, H. B. Sackrowitz, and Minya Xu. The use of an identity in Anderson for multivariate multiple testing. Journal of Statistical Planning and Inference, 138(9):2615–2621, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000542.

Chua:2000:PHR

[CT00] T. C. Chua and Albert K. Tsui. Procuring honest responses indirectly. Journal of Statistical Planning and Inference, 90(1):107–116, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001099.

Corcoran:2002:PSI

[CT02] J. N. Corcoran and R. L. Tweedie. Perfect sampling from independent Metropolis–Hastings chains. Journal of Statistical Planning and Inference, 104(2):297–314, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002439.

Chen:2004:BTD

[CT04] Zhiqiang Chen and David E. Tyler. On the behavior of Tukey’s depth and median under symmetric stable distributions. Journal of Statistical Planning and Inference, 122(1–2):111–124, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002477.

Cramer:2009:GOS

[CT09] Erhard Cramer and Trinh-Thai-Hang Tran. Generalized order statistics from arbitrary distributions and the Markov chain property. Journal of Statistical Planning and Inference, 139(12):4064–4071, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580900161X.

Chen:2004:SRQ

[CTL04] Lin-An Chen, Lanh Tat Tran, and Li-Ching Lin. Symmetric regression quantile and its application to robust estimation for the nonlinear regression model. Journal of Statistical Planning and Inference, 126(2):423–440, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002817.

Cuadras:2002:CAD

[Cua02] Carles M. Cuadras. Correspondence analysis and diagonal expansions in terms of distribution functions. Journal of Statistical Planning and Inference, 103(1–2):137–150, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002166.

Cuadras:2009:CCF

[Cua09] Carles M. Cuadras. Constructing copula functions with weighted geometric means. Journal of Statistical Planning and Inference, 139(11):3766–3772, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001414.

Cabrera:2009:GFT

[CV09] Jorge L. Ojeda Cabrera and Ingrid Van Keilegom. Goodnessof-fit tests for parametric regression with selection biased data. Journal of Statistical Planning and Inference, 139(8):2836–2850, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000093.

Coolen:2001:IPS

[CvdL01] F. P. A. Coolen and P. van der Laan. Imprecise predictive selection based on low structure assumptions. Journal of Statistical Planning and Inference, 98(1–2):259–277, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580000313X.

Chen:2001:V

[CvRZ01] K. Chen, G. H. J. van Rees, and L. Zhu. v(m, t) and its variants. Journal of Statistical Planning and Inference, 95 (1–2):143–160, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002846.

Coad:2005:CCI

[CW05] D. S. Coad and M. B. Woodroofe. Corrected confidence intervals for adaptive nonlinear regression models. Journal of Statistical Planning and Inference, 130(1–2):63–83, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001260.

Chen:2007:SSC

[CW07] Kejun Chen and Ruizhong Wei. Super-simple cyclic designs with small values. Journal of Statistical Planning and Inference, 137(6):2034–2044, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001170.

Cheng:2003:FDG

[CWW03] Shao-Wei Cheng, C. F. J. Wu, and Huaiqing Wu. Finding defining generators with extreme lengths. Journal of Statistical Planning and Inference, 113(1):315–321, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003093.

Cui:2005:API

[CX05] Lirong Cui and M. Xie. Availability of a periodically inspected system with random repair or replacement times. Journal of Statistical Planning and Inference, 131(1):89–100, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000321.

Chaubey:2007:SES

[CX07] Yogendra P. Chaubey and Haipeng Xu. Smooth estimation of survival functions under mean residual life ordering. Journal of Statistical Planning and Inference, 137(11):3303–3316, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000985.

Clarke:2004:PIR

[CY04a] B. Clarke and A. Yuan. Partial information reference priors: derivation and interpretations. Journal of Statistical Planning and Inference, 123(2):313–345, July 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001575.

Coolen:2004:NPI

[CY04b] F. P. A. Coolen and K. J. Yan. Nonparametric predictive inference with right-censored data. Journal of Statistical Planning and Inference, 126(1):25–54, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002258.

Chao:2001:PSE

[CYLC01] Anne Chao, Paul S. F. Yip, Shen-Ming Lee, and Wenten Chu. Population size estimation based on estimating functions for closed capture-recapture models. Journal of Statistical Planning and Inference, 92(1–2):213–232, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001518.

Chen:2002:EV

[CZ02] K. Chen and L. Zhu. Existence of v(m, t) vectors. Journal of Statistical Planning and Inference, 106(1–2):461–471, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002288.

Cai:2009:SRG

[CZ09a] Nanlian Cai and Yaohui Zheng. Some results on generalized ageing classes. Journal of Statistical Planning and Inference, 139(12):4223–4230, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001645.

Chen:2009:APN

[CZ09b] Jia Chen and Lixin Zhang. Asymptotic properties of nonparametric M-estimation for mixing functional data. Journal of Statistical Planning and Inference, 139(2):533–546, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002115.

Chen:2009:CCA

[CZ09c] Yi-Hau Chen and David M. Zucker. Case-cohort analysis with semiparametric transformation models. Journal of Statistical Planning and Inference, 139(10):3706–3717, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001244.

Danielak:2005:DFB

[Dan05a] Katarzyna Danielak. Distribution-free bounds for expectations of increments of records. Journal of Statistical Planning and Inference, 133(2):239–247, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001624.

Daniels:2005:CSP

[Dan05b] Michael J. Daniels. A class of shrinkage priors for the dependence structure in longitudinal data. Journal of Statistical Planning and Inference, 127(1–2):119–130, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803003185.

Dass:2001:PIP

[Das01] Sarat C. Dass. Propriety of intrinsic priors in invariant testing situations. Journal of Statistical Planning and Inference, 92(1–2):147–162, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001579.

DasGupta:2005:MBS

[Das05] Anirban DasGupta. The matching, birthday and the strong birthday problem: a contemporary review. Journal of Statistical Planning and Inference, 130(1–2):377–389, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002721.

DasGupta:2007:EBT

[Das07a] Anirban DasGupta. Extensions to Basu’s theorem, factorizations, and infinite divisibility. Journal of Statistical Planning and Inference, 137(3):945–952, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001534.

DasGupta:2007:SIF

[Das07b] Anirban DasGupta. Stein’s identity, Fisher information, and projection pursuit: a triangulation. Journal of Statistical Planning and Inference, 137(11):3394–3409, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001048.

Datta:2007:SFM

[Dat07] Gauri Sankar Datta. On symmetry of finite mixtures of normal distributions. Journal of Statistical Planning and Inference, 137(9):2993–2995, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000614.

Dauxois:2004:BIL

[Dau04] J.-Y. Dauxois. Bayesian inference for linear growth birth and death processes. Journal of Statistical Planning and Inference, 121(1):1–19, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802005207.

Davies:2004:OWT

[Dav04] Laurie Davies. The one-way-table. Journal of Statistical Planning and Inference, 122(1–2):3–13, ???? 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002374.

Davidov:2005:ERA

[Dav05] Ori Davidov. Estimation from retrospectively ascertained data-on what to condition? Journal of Statistical Planning and Inference, 134(1):49–63, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400179X.

Dewanji:2001:ODS

[DB01] Anup Dewanji and Atanu Biswas. An optimal design for simple illness-death model. Journal of Statistical Planning and Inference, 96(1):289–300, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003402.

Dembinska:2008:ADN

[DB08] Anna Dembińska and N. Balakrishnan. The asymptotic distribution of numbers of observations near order statistics. Journal of Statistical Planning and Inference, 138(8):2552–2562, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000244.

Dierckx:2009:NEM

[DBDG09] G. Dierckx, J. Beirlant, D. De Waal, and A. Guillou. A new estimation method for Weibull-type tails based on the mean excess function. Journal of Statistical Planning and Inference, 139(6):1905–1920, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003601.

Dilba:2006:SCS

[DBG06] Gemechis Dilba, Frank Bretz, and Volker Guiard. Simultaneous confidence sets and confidence intervals for multiple ratios. Journal of Statistical Planning and Inference, 136(8):2640–2658, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580400463X.

delBarrio:2009:RSL

[dBJM09] Eustasio del Barrio, Arnold Janssen, and Carlos Matrán. Resampling schemes with low resampling intensity and their applications in testing hypotheses. Journal of Statistical Planning and Inference, 139(2):184–202, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001547.

deCooman:2002:PIE

[dC02] Gert de Cooman. Precision-imprecision equivalence in a broad class of imprecise hierarchical uncertainty models. Journal of Statistical Planning and Inference, 105(1):175–198, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002099.

deCristofaro:2004:FLP

[dC04] Rodolfo de Cristofaro. On the foundations of likelihood principle. Journal of Statistical Planning and Inference, 126(2):401–411, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002854.

Deshpande:2006:P

[DC06] Jayant V. Deshpande and Roger M. Cooke. Preface. Journal of Statistical Planning and Inference, 136(5):1569–1571, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001783.

Deng:2001:SLL

[DCC01] Wen-Shuenn Deng, Chih-Kang Chu, and Ming-Yen Cheng. A study of local linear ridge regression estimators. Journal of Statistical Planning and Inference, 93(1–2):225–238, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001610.

delCastillo:2009:MLR

[dCD09] Joan del Castillo and Jalila Daoudi. The mixture of leftright truncated normal distributions. Journal of Statistical Planning and Inference, 139(10):3543–3551, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001013.

deCooman:2009:FI

[dCM09] Gert de Cooman and Enrique Miranda. Forward irrelevance. Journal of Statistical Planning and Inference, 139(2):256–276, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001341.

delCastillo:2005:OUP

[dCPC05] Joan del Castillo and Marta Pérez-Casany. Overdispersed and underdispersed Poisson generalizations. Journal of Statistical Planning and Inference, 134(2):486–500, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002289.

Doko:2008:IEI

[DD08] Firmin Doko and Jean-Marie Dufour. Instrument endogeneity and identification-robust tests: Some analytical results. Journal of Statistical Planning and Inference, 138(9):2649–2661, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000578.

Das:2008:OSD

[DDCC08] Ashish Das, Aloke Dey, Ling-Yau Chan, and Kashinath Chatterjee. On E(s^{2})-optimal supersaturated designs. Journal of Statistical Planning and Inference, 138(12):3749–3757, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001365.

Drees:2006:ATE

[DdHL06] Holger Drees, Laurens de Haan, and Deyuan Li. Approximations to the tail empirical distribution function with application to testing extreme value conditions. Journal of Statistical Planning and Inference, 136(10):3498–3538, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000753.

Dutta:2009:OCD

[DDM09] Ganesh Dutta, Premadhis Das, and Nripes K. Mandal. Optimum covariate designs in partially balanced incomplete block (PBIB) design set-ups. Journal of Statistical Planning and Inference, 139(8):2823–2835, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000081.

Dao-de:2000:SFS

[DdRB00] Gao Dao-de and Huang Rong-Bing. Some finite sample properties of Zellner estimator in the context of m seemingly unrelated regression equations. Journal of Statistical Planning and Inference, 88(2):267–283, August 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000835.

Dharmadhikari:2000:CTE

[DDS00] A. A. Dharmadhikari, A. D. Dharmadhikari, and Kirti R. Shah. Comparison of two experimental designs in reliability studies. Journal of Statistical Planning and Inference, 84(1–2):237–248, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001111.

Durbin:2008:DSA

[DDvdL08] Blythe Durbin, Sandrine Dudoit, and Mark J. van der Laan. A deletion/substitution/addition algorithm for classification neural networks, with applications to biomedical data. Journal of Statistical Planning and Inference, 138(2):464–488, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002558.

DeSantis:2004:SES

[De 04] Fulvio De Santis. Statistical evidence and sample size determination for Bayesian hypothesis testing. Journal of Statistical Planning and Inference, 124(1):121–144, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001988.

deBEdwardes:2002:DFT

[dE02] Michael D. deB. Edwardes. Distribution-free tests for cluster samples of ordinal responses. Journal of Statistical Planning and Inference, 105(2):393–404, July 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002518.

Dean:2000:BRF

[Dea00] A. M. Dean. Book review: Fractional factorial plans: A. Dey and R. Mukerjee: Wiley, New York, 1999, price $94.95 paperback, ISBN 0-471-29414-4. Journal of Statistical Planning and Inference, 91(1):177–178, November 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001312.

Deely:2006:SG

[Dee06] John J. Deely. Shanti S. Gupta. Journal of Statistical Planning and Inference, 136(7):1999–2003, July 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002065.

Danilov:2000:ACM

[DEH00a] D. L. Danilov, S. M. Ermakov, and J. H. Halton. Asymptotic complexity of Monte Carlo methods for solving linear systems. Journal of Statistical Planning and Inference, 85(1–2):5–18, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899000609.

Deheuvels:2000:NSL

[Deh00b] Paul Deheuvels. Nonstandard strong laws for local quantile processes. Journal of Statistical Planning and Inference, 91(2):239–266, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001816.

Deshpande:2002:P

[Des02] J. V. Deshpande. Preface. Journal of Statistical Planning and Inference, 102(1):1, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001574.

Dette:2002:CTH

[Det02] Holger Dette. A consistent test for heteroscedasticity in nonparametric regression based on the kernel method. Journal of Statistical Planning and Inference, 103(1–2):311–329, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002294.

Dey:2009:OBT

[Dey09] Aloke Dey. Orthogonally blocked three-level second order designs. Journal of Statistical Planning and Inference, 139(10):3698–3705, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001256.

Doksum:2000:NC

[DF00] Kjell A. Doksum and Sorana M. Froda. Neighborhood correlation. Journal of Statistical Planning and Inference, 91(2): 267–294, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001828.

Dragalin:2006:ADD

[DF06] Vladimir Dragalin and Valerii Fedorov. Adaptive designs for dose-finding based on efficacy-toxicity response. Journal of Statistical Planning and Inference, 136(6):1800–1823, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001862.

Delmas:2007:TCR

[DF07] Céline Delmas and Jean-Louis Foulley. On testing a class of restricted hypotheses. Journal of Statistical Planning and Inference, 137(4):1343–1361, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001078.

Davies:2004:RSE

[DFG04] P. L. Davies, R. Fried, and U. Gather. Robust signal extraction for on-line monitoring data. Journal of Statistical Planning and Inference, 122(1–2):65–78, ???? 1, 2004.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002416.

Daouia:2008:FCQ

[DFS08] Abdelaati Daouia, Jean-Pierre Florens, and Léopold Simar. Functional convergence of quantile-type frontiers with application to parametric approximations. Journal of Statistical Planning and Inference, 138(3):708–725, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001395.

Dragalin:2008:ADS

[DFW08] Vladimir Dragalin, Valerii Fedorov, and Yuehui Wu. Adaptive designs for selecting drug combinations based on efficacy-toxicity response. Journal of Statistical Planning and Inference, 138(2):352–373, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807002480.

Deza:2001:FRB

[DG01] Michel Deza and Viacheslav Grishukhin. Face-regular bifaced polyhedra. Journal of Statistical Planning and Inference, 95(1–2):175–195, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002871.

Dmitrienko:2002:SDN

[DG02] Alexei Dmitrienko and Z. Govindarajulu. Sequential determination of the number of bootstrap samples. Journal of Statistical Planning and Inference, 100(2):349–363, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001458.

Davidov:2004:FLS

[DG04] O. Davidov and A. Goldenshluger. Fitting a line segment to noisy data. Journal of Statistical Planning and Inference, 119(1):191–206, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004093.

Dijoux:2009:ADC

[DG09a] Yann Dijoux and Olivier Gaudoin. The alert-delay competing risks model for maintenance analysis. Journal of Statistical Planning and Inference, 139(5):1587–1603, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002358.

Du:2009:PPL

[DG09b] Pang Du and Chong Gu. Penalized pseudo-likelihood hazard estimation: a fast alternative to penalized likelihood. Journal of Statistical Planning and Inference, 139(3):891–899, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002607.

Diaz-Garcia:2004:NCD

[DGGF04] José A. Díaz-García and Graciela González-Farías. A note on the Cook’s distance. Journal of Statistical Planning and Inference, 120(1–2):119–136, February 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004949.

Diebolt:2008:BRE

[DGGG08] Jean Diebolt, Laurent Gardes, Stéphane Girard, and Armelle Guillou. Bias-reduced extreme quantile estimators of Weibull tail-distributions. Journal of Statistical Planning and Inference, 138(5):1389–1401, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807002698.

Diaz-Garcia:2006:DGI

[DGGJ06] José A. Díaz-García and Ramón Gutiérrez-Jáimez. Distribution of the generalised inverse of a random matrix and its applications. Journal of Statistical Planning and Inference, 136(1):183–192, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580400299X.

Diaz-Garcia:2007:DRG

[DGGJ07] José A. Díaz-García and Ramón Gutiérrez-Jáimez. The distribution of residuals from a general elliptical linear model. Journal of Statistical Planning and Inference, 137(7):2347– 2354, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580600259X.

Diaz-Garcia:2009:SMM

[DGGJ09] José A. Díaz-García and Ramón Gutiérrez-Jáimez. Singular matric and matrix variate t distributions. Journal of Statistical Planning and Inference, 139(7):2382–2387, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004242.

Diaz-Garcia:2006:WPW

[DGJ06] José A. Díaz-García and Ramón Gutiérrez Jáimez. Wishart and pseudo-Wishart distributions under elliptical laws and related distributions in the shape theory context. Journal of Statistical Planning and Inference, 136(12):4176–4193, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002442.

Diaz-Garcia:2007:ENF

[DGL07] José A. Díaz-García and Víctor Leiva. Erratum to “A new family of life distributions based on the elliptically contoured distributions”: [J. Statist. Plann. Inference 128(2) (2005) 445–457]. Journal of Statistical Planning and Inference, 137(4):1512–1513, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001984. See [DGLS05].

Diaz-Garcia:2005:NFL

[DGLS05] José A. Díaz-García and Víctor Leiva-Sánchez. A new family of life distributions based on the elliptically contoured distributions. Journal of Statistical Planning and Inference, 128(2):445–457, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000072. See erratum [DGL07].

Diebolt:2007:ADE

[DGR07] Jean Diebolt, Armelle Guillou, and Imen Rached. Approximation of the distribution of excesses through a generalized probability-weighted moments method. Journal of Statistical Planning and Inference, 137(3):841–857, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001455.

Dikta:2005:CLT

[DGS05] Gerhard Dikta, Jugal Ghorai, and Christian Schmidt. The central limit theorem under semiparametric random censorship models. Journal of Statistical Planning and Inference, 127(1–2):23–51, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003100.

Davidov:2000:ODD

[DH00] Ori Davidov and Yoel Haitovsky. Optimal design for double sampling with continuous outcomes. Journal of Statistical Planning and Inference, 86(1):253–263, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001603.

Ding:2003:CSH

[DH03] A. Adam Ding and J. T. Gene Hwang. Confidence sets for high-dimensional empirical linear prediction (HELP) models with dependent error structure. Journal of Statistical Planning and Inference, 113(1):189–213, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003111.

DasGupta:2006:AVE

[DH06] Anirban DasGupta and L. R. Haff. Asymptotic values and expansions for the correlation between different measures of spread. Journal of Statistical Planning and Inference, 136(7):2197–2212, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002181.

Debusho:2008:OPD

[DH08] Legesse Kassa Debusho and Linda M. Haines. v- and Doptimal population designs for the simple linear regression model with a random intercept term. Journal of Statistical Planning and Inference, 138(4):1116–1130, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002340.

Davidov:2009:NTS

[DH09a] Ori Davidov and Amir Herman. New tests for stochastic order with application to case control studies. Journal of Statistical Planning and Inference, 139(8):2614–2623, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004497.

DeBlasi:2009:BMT

[DH09b] Pierpaolo De Blasi and Nils Lid Hjort. The Bernstein-von Mises theorem in semiparametric competing risks models. Journal of Statistical Planning and Inference, 139(7):2316– 2328, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004205.

Duan:2009:DAB

[DH09c] Liangliang Duan and Feifang Hu. Doubly adaptive biased coin designs with heterogeneous responses. Journal of Statistical Planning and Inference, 139(9):3220–3230, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900069X.

dIppolito:2009:IFL

[dHD^{+}09] R. d’Ippolito, M. Hack, S. Donders, L. Hermans, N. Tzannetakis, and D. Vandepitte. Improving the fatigue life of a vehicle knuckle with a reliability-based design optimization approach. Journal of Statistical Planning and Inference, 139(5):1619–1632, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002383.

Delecroix:2006:SES

[DHP06] Michel Delecroix, Marian Hristache, and Valentin Patilea. On semiparametric M-estimation in single-index regression. Journal of Statistical Planning and Inference, 136(3):730–769, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003611.

DeWachter:2007:PDU

[DHT07] Stefan De Wachter, Richard D. F. Harris, and Elias Tzavalis. Panel data unit roots tests: the role of serial correlation and the time dimension. Journal of Statistical Planning and Inference, 137(1):230–244, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805003010.

Dunnett:2001:SSD

[DHV01] Charles W. Dunnett, Manfred Horn, and Rüdiger Vollandt. Sample size determination in step-down and step-up multiple tests for comparing treatments with a control. Journal of Statistical Planning and Inference, 97(2):367–384, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002275.

Dugdale:2000:FLS

[DHW00] J. K. Dugdale, A. J. W. Hilton, and J. Wojciechowski. Fractional latin squares, simplex algebras, and generalized quotients. Journal of Statistical Planning and Inference, 86 (2):457–504, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001251.

Davidov:2009:EUN

[DI09] Ori Davidov and George Iliopoulos. On the existence and uniqueness of the NPMLE in biased sampling models. Journal of Statistical Planning and Inference, 139(2):176–183, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001560.

Dikta:2000:SLU

[Dik00] Gerhard Dikta. The strong law under semiparametric random censorship models. Journal of Statistical Planning and Inference, 83(1):1–10, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000865.

Ding:2003:LCB

[Din03] Keyue Ding. A lower confidence bound for the change point after a sequential CUSUM test. Journal of Statistical Planning and Inference, 115(1):311–326, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001477.

Deely:2006:IHM

[DJ06] John J. Deely and Wesley O. Johnson. Inferences for hierarchical models with partial prior information. Journal of Statistical Planning and Inference, 136(7):2327–2339, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805002259.

Dachian:2003:CEE

[DK03] Serguei Dachian and Yury A. Kutoyants. On cusp estimation of ergodic diffusion process. Journal of Statistical Planning and Inference, 117(1):153–166, November 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003658.

Dehay:2004:CID

[DK04a] D. Dehay and Yu. A. Kutoyants. On confidence intervals for distribution function and density of ergodic diffusion process. Journal of Statistical Planning and Inference, 124(1):63–73, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803001952.

Ding:2004:LEB

[DK04b] Keyue Ding and Rohana J. Karunamuni. A linear empirical Bayes solution for the calibration problem. Journal of Statistical Planning and Inference, 119(2):421–447, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004925.

Denker:2007:UMP

[DK07] Manfred Denker and Susanne Koch. On the uniqueness of the maximum of the paths of random walks. Journal of Statistical Planning and Inference, 137(3):659–668, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001327.

Dachian:2009:HTP

[DK09] Sergueï Dachian and Yury A. Kutoyants. Hypotheses testing: Poisson versus stress-release. Journal of Statistical Planning and Inference, 139(5):1668–1684, May 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002437.

Davydov:2007:CRG

[DKSZ07] Youri Davydov, Davar Khoshnevisan, Zhan Shi, and Ricardas Zitikis. Convex rearrangements, generalized Lorenz curves, and correlated Gaussian data. Journal of Statistical Planning and Inference, 137(3):915–934, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001510.

Dinitz:2000:RSP

[DL00a] J. H. Dinitz and E. R. Lamken. Room square patterns. Journal of Statistical Planning and Inference, 86(2):397–410, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001202.

Durham:2000:TRS

[DL00b] S. D. Durham and J. D. Lynch. A threshold representation for the strength distribution of a complex load sharing system. Journal of Statistical Planning and Inference, 83(1):25–46, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899000944.

Dawson:2003:CDA

[DL03] Ree Dawson and Philip W. Lavori. Comparison of designs for adaptive treatment strategies: baseline vs. adaptive randomization. Journal of Statistical Planning and Inference, 117(2):365–385, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003695.

Dass:2004:NCB

[DL04] Sarat C. Dass and Jaeyong Lee. A note on the consistency of Bayes factors for testing point null versus nonparametric alternatives. Journal of Statistical Planning and Inference, 119(1):143–152, January 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004135.

DeIorio:2005:IAM

[DL05] M. De Iorio and M. Lavine. Intrinsic autoregressions at multiple resolutions. Journal of Statistical Planning and Inference, 134(1):102–115, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001843.

deLeeuw:2006:CPW

[dL06a] Jan de Leeuw. Comments on Pardoe and Weitner: sentencing convicted felons in the United States: a Bayesian analysis using multilevel covariates. Journal of Statistical Planning and Inference, 136(4):1466–1468, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004124. See [PW06b].

DiCrescenzo:2006:NAN

[DL06b] Antonio Di Crescenzo and Maria Longobardi. On the NBU ageing notion within the competing risks model. Journal of Statistical Planning and Inference, 136(5):1638–1654, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001722.

Dukes:2007:LPB

[DL07] Peter Dukes and Alan C. H. Ling. Linear programming bounds for balanced arrays. Journal of Statistical Planning and Inference, 137(1):324–330, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805003058.

DiCrescenzo:2009:CE

[DL09a] Antonio Di Crescenzo and Maria Longobardi. On cumulative entropies. Journal of Statistical Planning and Inference, 139(12):4072–4087, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001621.

Dupuy:2009:SRC

[DL09b] Jean-François Dupuy and Eve Leconte. A study of regression calibration in a partially observed stratified Cox model. Journal of Statistical Planning and Inference, 139(2):317–328, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001845.

Datta:2002:EBE

[DLM02] G. S. Datta, P. Lahiri, and T. Maiti. Empirical Bayes estimation of median income of four-person families by state using time series and cross-sectional data. Journal of Statistical Planning and Inference, 102(1):83–97, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001732.

Dykstra:2002:ORI

[DLOR02] Richard L. Dykstra, Joseph B. Lang, Myongsik Oh, and Tim Robertson. Order restricted inference for hypotheses concerning qualitative dispersion. Journal of Statistical Planning and Inference, 107(1–2):249–265, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002562.

Denuit:2003:PSO

[DLP03] M. Denuit, Cl. Lefèvre, and Ph. Picard. Polynomial structures in order statistics distributions. Journal of Statistical Planning and Inference, 113(1):151–178, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002920.

Dette:2006:MED

[DLRP06] Holger Dette, Ignacio Martinez Lopez, Isabel M. Ortiz Rodriguez, and Andrey Pepelyshev. Maximin efficient design of experiment for exponential regression models. Journal of Statistical Planning and Inference, 136(12):4397–4418, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805001813.

Datta:2009:NEW

[DLS09] Somnath Datta, Ling Lan, and Rajeshwari Sundaram. Nonparametric estimation of waiting time distributions in a Markov model based on current status data. Journal of Statistical Planning and Inference, 139(9):2885–2897, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000135.

Denny:2002:CD

[DM02] Paul C. Denny and Rudolf Mathon. A census of t − (t + 8, t + 2, 4) designs, 2 ≤ t ≤ 4. Journal of Statistical Planning and Inference, 106(1–2):5–19, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001970.

DeCooman:2008:WSL

[DM08a] Gert De Cooman and Enrique Miranda. Weak and strong laws of large numbers for coherent lower previsions. Journal of Statistical Planning and Inference, 138(8):2409–2432, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003850.

Dette:2008:TPF

[DM08b] Holger Dette and Mareen Marchlewski. A test for the parametric form of the variance function in a partial linear regression model. Journal of Statistical Planning and Inference, 138(10):3005–3021, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004119.

Dudewicz:2007:ESB

[DMMS07] Edward J. Dudewicz, Yan Ma, Enping (Shirley) Mai, and Haiyan Su. Exact solutions to the Behrens–Fisher problem: Asymptotically optimal and finite sample efficient choice among. Journal of Statistical Planning and Inference, 137(5):1584–1605, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002382.

DeSantis:2001:JPS

[DMN01] Fulvio De Santis, Julia Mortera, and Alessandra Nardi. Jeffreys priors for survival models with censored data. Journal of Statistical Planning and Inference, 99(2):193–209, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801000805.

Dette:2004:ODE

[DMP04] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Optimal designs for estimating individual coefficients in polynomial regression — a functional approach. Journal of Statistical Planning and Inference, 118(1–2):201–219, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580200397X.

Das:2003:OED

[DMS03] Kalyan Das, Nripes K. Mandal, and Bikas K. Sinha. Optimal experimental designs for models with covariates. Journal of Statistical Planning and Inference, 115(1):273–285, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001519.

Durio:2003:LBE

[DN03] A. Durio and Ya. Yu Nikitin. Local Bahadur efficiency of some goodness-of-fit tests under skew alternatives. Journal of Statistical Planning and Inference, 115(1):171–179, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001556.

Dean:2002:BEM

[DNK02] A. M. Dean, W. I. Notz, and Lie-Jane Kao. A bound for efficiency in multiple response designs. Journal of Statistical Planning and Inference, 106(1–2):235–244, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002173.

Deshpande:2007:CRI

[DNN07] J. V. Deshpande and U. V. Naik-Nimbalkar. Comparing relative importance of the same component in a series system in two environments. Journal of Statistical Planning and Inference, 137(11):3410–3415, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700105X.

Dabo-Niang:2009:KRE

[DNR09] Sophie Dabo-Niang and Noureddine Rhomari. Kernel regression estimation in a Banach space. Journal of Statistical Planning and Inference, 139(4):1421–1434, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003443.

Davydov:2001:RCC

[DÖ01] Alexander A. Davydov and Patric R. J. Östergård. Recursive constructions of complete caps. Journal of Statistical Planning and Inference, 95(1–2):167–173, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000286X.

Dobler:2002:OWL

[Dob02] Carolyn Pillers Dobler. The one-way layout with ordered parameters: a survey of advances since 1988. Journal of Statistical Planning and Inference, 107(1–2):75–88, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002446.

Donev:2004:DEP

[Don04] A. N. Donev. Design of experiments in the presence of errors in factor levels. Journal of Statistical Planning and Inference, 126(2):569–585, December 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002799.

Dorazio:2009:SPP

[Dor09] Robert M. Dorazio. On selecting a prior for the precision parameter of Dirichlet process mixture models. Journal of Statistical Planning and Inference, 139(9):3384–3390, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000755.

Dette:2008:EED

[DP08a] Holger Dette and Andrey Pepelyshev. Efficient experimental designs for sigmoidal growth models. Journal of Statistical Planning and Inference, 138(1):2–17, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001917.

Di:2008:DLS

[DP08b] Yanming Di and Michael D. Perlman. Detecting linear sequences and subsequences. Journal of Statistical Planning and Inference, 138(9):2634–2648, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000566.

Drton:2008:SAG

[DP08c] Mathias Drton and Michael D. Perlman. A SINful approach to Gaussian graphical model selection. Journal of Statistical Planning and Inference, 138(4):1179–1200, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002303.

Dutilleul:2008:MTA

[DPA08] Pierre Dutilleul, Bernard Pelletier, and Gülhan Alpargu. Modified F tests for assessing the multiple correlation between one spatial process and several others. Journal of Statistical Planning and Inference, 138(5):1402–1415, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002704.

Dios-Palomares:2006:STU

[DPR06] Rafaela Dios-Palomares and Jose A. Roldan. A strategy for testing the unit root in AR(1) model with intercept: a Monte Carlo experiment. Journal of Statistical Planning and Inference, 136(8):2685–2705, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004227.

Dewan:2001:ANS

[DR01a] I. Dewan and B. L. S. Prakasa Rao. Asymptotic normality for U-statistics of associated random variables. Journal of Statistical Planning and Inference, 97(2):201–225, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002263.

Dutta:2001:CSR

[DR01b] Tridib K. Dutta and Bimal K. Roy. Construction of some repeated measurements designs. Journal of Statistical Planning and Inference, 95(1–2):283–291, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002950.

Dyachkov:2002:OSC

[DR02] Arkadii G. D’yachkov and Vyacheslav V. Rykov. Optimal superimposed codes and designs for Renyi’s search model. Journal of Statistical Planning and Inference, 100(2):281–302, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001409.

DasGupta:2003:BEE

[DR03a] Anirban DasGupta and Herman Rubin. Bayes estimates as expanders in one and two dimensions. Journal of Statistical Planning and Inference, 113(1):1–14, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003019.

Dupuis:2003:VSQ

[DR03b] Jérome A. Dupuis and Christian P. Robert. Variable selection in qualitative models via an entropic explanatory power. Journal of Statistical Planning and Inference, 111(1–2):77– 94, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002860.

Delicado:2004:RCR

[DR04] Pedro F. Delicado and Juan Romo. Random coefficient regressions: parametric goodness-of-fit tests. Journal of Statistical Planning and Inference, 119(2):377–400, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004846.

DasGupta:2005:EBP

[DR05] A. DasGupta and Herman Rubin. Estimation of binomial parameters when both n, p are unknown. Journal of Statistical Planning and Inference, 130(1–2):391–404, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002733.

DeMartini:2008:MSB

[DR08a] Daniele De Martini and Fabio Rapallo. On multivariate smoothed bootstrap consistency. Journal of Statistical Planning and Inference, 138(6):1828–1835, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003242.

Ding:2008:HMM

[DR08b] Yu Ding and Damaraju Raghavarao. Hadamard matrix methods in identifying differentially expressed genes from microarray experiments. Journal of Statistical Planning and Inference, 138(1):47–55, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001942.

Dykstra:2002:P

[DRS02] Richard L. Dykstra, Tim Robertson, and Mervyn J. Silvapulle. Preface. Journal of Statistical Planning and Inference, 107(1–2):1–2, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002410.

DeSantis:2001:CFB

[DS01] Fulvio De Santis and Fulvio Spezzaferri. Consistent fractional Bayes factor for nested normal linear models. Journal of Statistical Planning and Inference, 97(2):305–321, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002408.

Demidenko:2002:EEG

[DS02a] Eugene Demidenko and Thérèse A. Stukel. Efficient estimation of general linear mixed effects models. Journal of Statistical Planning and Inference, 104(1):197–219, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001239.

Dinitz:2002:BDD

[DS02b] Jeffrey H. Dinitz and Nabil Shalaby. Block disjoint difference families for Steiner triple systems: v ≡ 3 mod 6. Journal of Statistical Planning and Inference, 106(1–2):77–86, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002045.

Datta:2003:PPD

[DS03] Gauri Sankar Datta and David Daniel Smith. On propriety of posterior distributions of variance components in small area estimation. Journal of Statistical Planning and Inference, 112(1–2):175–183, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003312.

Desai:2006:IAS

[DS06] Tejas A. Desai and Pranab K. Sen. Information attainable in some randomly incomplete data models. Journal of Statistical Planning and Inference, 136(7):2309–2326, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002247.

Donnet:2007:EPI

[DS07] Sophie Donnet and Adeline Samson. Estimation of parameters in incomplete data models defined by dynamical systems. Journal of Statistical Planning and Inference, 137(9):2815–2831, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000353.

Desai:2008:IAS

[DS08] Tejas A. Desai and Pranab K. Sen. Information attainable in some randomly incomplete multivariate response models. Journal of Statistical Planning and Inference, 138(11):3467–3482, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000906.

Filho:2007:BDR

[dSBFG07] Júlio S. de S. Bueno Filho and Steven G. Gilmour. Block designs for random treatment effects. Journal of Statistical Planning and Inference, 137(4):1446–1451, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000619.

Doveh:2002:TMP

[DSF02] E. Doveh, A. Shapiro, and P. D. Feigin. Testing of monotonicity in parametric regression models. Journal of Statistical Planning and Inference, 107(1–2):289–306, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002598.

Dasgupta:2007:NIP

[DSG07] Abhijit Dasgupta, Steven G. Self, and Somesh Das Gupta. Non-identifiable parametric probability models and reparametrization. Journal of Statistical Planning and Inference, 137 (11):3380–3393, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807001036.

Tunes-da-Silva:2008:EMQ

[dSSdL08] Gisela Tunes da Silva, Pranab K. Sen, and Antonio Carlos Pedroso de Lima. Estimation of the mean quality-adjusted survival using a multistate model for the sojourn times. Journal of Statistical Planning and Inference, 138(8):2267–2282, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003825.

Deville:2000:SSU

[DT00] Jean-Claude Deville and Yves Tillé. Selection of several unequal probability samples from the same population. Journal of Statistical Planning and Inference, 86(1):215–227, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001640.

Deville:2005:VAU

[DT05] Jean-Claude Deville and Yves Tillé. Variance approximation under balanced sampling. Journal of Statistical Planning and Inference, 128(2):569–591, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000035.

DiMarzio:2008:BKR

[DT08] Marco Di Marzio and Charles C. Taylor. On boosting kernel regression. Journal of Statistical Planning and Inference, 138(8):2483–2498, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003771.

Donev:2008:CCD

[DTM08] Alexander N. Donev, Randy Tobias, and Farinaz Monadjemi. Cost-cautious designs for confirmatory bioassay. Journal of Statistical Planning and Inference, 138(12):3805–3812, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001353.

Dobra:2006:DAM

[DTW06] Adrian Dobra, Claudia Tebaldi, and Mike West. Data augmentation in multi-way contingency tables with fixed marginal totals. Journal of Statistical Planning and Inference, 136(2):355–372, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003040.

deUna-Alvarez:2000:KDF

[dUÁGMCS00] Jacobo de Uña-Álvarez, Wenceslao González-Manteiga, and Carmen Cadarso-Suárez. Kernel distribution function estimation under the Koziol–Green model. Journal of Statistical Planning and Inference, 87(2):199–219, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001925.

deUna-Alvarez:2003:CSG

[dUÁGMCS03] Jacobo de Uña-Álvarez, Wenceslao González-Manteiga, and Carmen Cadarso-Suárez. Chi-squared goodness-of-fit theory under proportional censorship. Journal of Statistical Planning and Inference, 109(1–2):101–124, January 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003063.

Dubnicka:2007:CIM

[Dub07] Suzanne R. Dubnicka. A confidence interval for the median of a finite population under unequal probability sampling: a model-assisted approach. Journal of Statistical Planning and Inference, 137(7):2429–2438, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002771.

Duckworth:2000:SBM

[Duc00] William Maury Duckworth. Some binary maximin distance designs. Journal of Statistical Planning and Inference, 88 (1):149–170, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001834.

Dupuy:2005:PHM

[Dup05] Jean-François Dupuy. The proportional hazards model with covariate measurement error. Journal of Statistical Planning and Inference, 135(2):260–275, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002307.

Dupuy:2009:DCH

[Dup09] Jean-François Dupuy. Detecting change in a hazard regression model with right-censoring. Journal of Statistical Planning and Inference, 139(5):1578–1586, May 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002346.

DiBucchianico:2005:END

[DvdW05] A. Di Bucchianico and M. A. van de Wiel. Exact null distributions of distribution-free quadratic t-sample statistics. Journal of Statistical Planning and Inference, 127(1–2):1–21, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003161.

Dembinska:2000:LRN

[DW00] Anna Dembińska and Jacek Wesolowski. Linearity of regression for non-adjacent record values. Journal of Statistical Planning and Inference, 90(2):195–205, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001130.

Ding:2006:BCP

[DW06] Keyue Ding and Yanhong Wu. On the biases of change point and change magnitude estimation after CUSUM test. Journal of Statistical Planning and Inference, 136(4):1258–1280, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003660.

Durairajan:2008:OEF

[DW08] T. M. Durairajan and Martin L. William. Optimal estimating function for estimation and prediction in semiparametric models. Journal of Statistical Planning and Inference, 138(10):3283–3292, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800030X.

Davidov:2000:ETE

[DZ00] Ori Davidov and Marvin Zelen. Exact tests for exponential regression. Journal of Statistical Planning and Inference, 88(1):87–97, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899002025.

Dorea:2006:EBP

[DZ06] C. C. Y. Dorea and L. C. Zhao. Exponential bounds for the probability of wrong determination of the order of a Markov chain by using the EDC criterion. Journal of Statistical Planning and Inference, 136(10):3689–3697, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000741.

Ding:2009:OPS

[DZ09] Yuanyao Ding and Bo Zhang. Optimal portfolio of safetyfirst models. Journal of Statistical Planning and Inference, 139(9):2952–2962, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000263.

El-Bassiouni:2008:AIE

[EBZ08] M. Y. El-Bassiouni and Taoufik Zoubeidi. Adaptive interval estimation in one-way random effects models. Journal of Statistical Planning and Inference, 138(4):857–872, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000559.

Eberly:2003:EBC

[EC03] Lynn E. Eberly and George Casella. Estimating Bayesian credible intervals. Journal of Statistical Planning and Inference, 112(1–2):115–132, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802003270.

Eldar:2008:PTL

[EC08] Yonina C. Eldar and Jacob Slava Chernoi. A pre-test like estimator dominating the least-squares method. Journal of Statistical Planning and Inference, 138(10):3069–3085, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807004223.

Eryilmaz:2007:SRS

[ED07] Serkan Eryilmaz and Sevcan Demir. Success runs in a sequence of exchangeable binary trials. Journal of Statistical Planning and Inference, 137(9):2954–2963, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000456.

Efromovich:2004:DEB

[Efr04] Sam Efromovich. Distribution estimation for biased data. Journal of Statistical Planning and Inference, 124(1):1–43, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002027.

Efromovich:2007:AEE

[Efr07a] Sam Efromovich. Adaptive estimation of error density in nonparametric regression with small sample size. Journal of Statistical Planning and Inference, 137(2):363–378, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806000383.

Efromovich:2007:LBO

[Efr07b] Sam Efromovich. A lower-bound oracle inequality for a blockwise-shrinkage estimate. Journal of Statistical Planning and Inference, 137(1):176–183, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002971.

Efromovich:2009:LBE

[Efr09] Sam Efromovich. Lower bound for estimation of Sobolev densities of order less 12. Journal of Statistical Planning and Inference, 139(7):2261–2268, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004187.

Ensor:2000:SMR

[EG00] Katherine Bennett Ensor and Peter W. Glynn. Simulating the maximum of a random walk. Journal of Statistical Planning and Inference, 85(1–2):127–135, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000750.

Escobar:2009:P

[EGMN09] Luis A. Escobar, Fabrice Guerin, William Q. Meeker, and Mikhail Nikulin. Preface. Journal of Statistical Planning and Inference, 139(5):1576–1577, May 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002334.

Escobar:2009:SCB

[EHM09] Luis A. Escobar, Yili Hong, and William Q. Meeker. Simultaneous confidence bands and regions for log-location-scale distributions with censored data. Journal of Statistical Planning and Inference, 139(9):3231–3245, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000573.

Ekstrom:2007:ADF

[EJ07] Magnus Ekström and S. Rao Jammalamadaka. An asymptotically distribution-free test of symmetry. Journal of Statistical Planning and Inference, 137(3):799–810, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001418.

Ekstrom:2009:AAD

[EJ09] Magnus Ekström and S. Rao Jammalamadaka. Addendum to “An asymptotically distribution-free test of symmetry” [Journal of Statistical Planning and Inference 137 (2007) 799–810]. Journal of Statistical Planning and Inference, 139(4):1569–1571, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003820.

Evangelaras:2006:RPD

[EKK06] H. Evangelaras, E. Kolaiti, and C. Koukouvinos. Robust parameter design: Optimization of combined array approach with orthogonal arrays. Journal of Statistical Planning and Inference, 136(10):3698–3709, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000893.

Evangelaras:2007:FCN

[EKL07] H. Evangelaras, C. Koukouvinos, and E. Lappas. Further contributions to nonisomorphic two level orthogonal arrays. Journal of Statistical Planning and Inference, 137(6): 2080–2086, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580600108X.

ElBarmi:2004:ECI

[EKMS04] Hammou El Barmi, Subhash C. Kochar, Hari Mukerjee, and Francisco J. Samaniego. Estimation of cumulative incidence functions in competing risks studies under an order restriction. Journal of Statistical Planning and Inference, 118(1–2):145–165, January 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004019.

Ekstrom:2006:SFS

[Eks06]  Magnus Ekström.  Sum-functions of spacings of in-

creasing order. Journal of Statistical Planning and Inference, 136(8):2535–2546, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004641.

Ekstrom:2008:AML

[Eks08] Magnus Ekström. Alternatives to maximum likelihood estimation based on spacings and the Kullback–Leibler divergence. Journal of Statistical Planning and Inference, 138(6): 1778–1791, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807003199.

ElBarmi:2006:LRT

[EKT06] Hammou El Barmi, Subhash Kochar, and John Tsimikas. Likelihood ratio tests for and against ordering of the cumulative incidence functions in multiple competing risks and discrete mark variable models. Journal of Statistical Planning and Inference, 136(5):1588–1607, May 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001692.

Eslami:2000:HD

[EKTR00] Z. Eslami, G. B. Khosrovshahi, and B. Tayfeh-Rezaie. On halvings of the 2-(10,3,8) design. Journal of Statistical Planning and Inference, 86(2):411–419, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001214.

ElBantli:2004:MEL

[El 04] Faouzi El Bantli. M-estimation in linear models under nonstandard conditions. Journal of Statistical Planning and Inference, 121(2):231–248, April 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001137.

Eberly:2004:BFC

[EL04] Lynn E. Eberly and Thomas A. Louis. Bayes/frequentist compromise decision rules for Gaussian sampling. Journal of Statistical Planning and Inference, 121(2):191–207, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001344.

Eaton:2007:MET

[EM07] Morris L. Eaton and Robb J. Muirhead. On a multiple endpoints testing problem. Journal of Statistical Planning and Inference, 137(11):3416–3429, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001061.

Esteban:2000:SNS

[EMMM00] M. D. Esteban, A. M. Mayoral, D. Morales, and J. Morales. Some new statistics for testing point null hypotheses with prior information. Journal of Statistical Planning and Inference, 87(2):251–271, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001998.

Eaton:2006:AVC

[EMP06] Morris L. Eaton, Robb J. Muirhead, and Eve H. Pickering. Assessing a vector of clinical observations. Journal of Statistical Planning and Inference, 136(10):3383–3414, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000790.

ElBarmi:2002:NEN

[EN02] Hammou El Barmi and Paul I. Nelson. A note on estimating a non-increasing density in the presence of selection bias. Journal of Statistical Planning and Inference, 107(1–2):353–364, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200263X.

Eideh:2006:FTS

[EN06] Abdulhakeem A. H. Eideh and Gad Nathan. Fitting time series models for longitudinal survey data under informative sampling. Journal of Statistical Planning and Inference, 136(9):3052–3069, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004677. See corrigendum [EN07].

Eideh:2007:CFT

[EN07] Abdulhakeem A. H. Eideh and Gad Nathan. Corrigendum to “Fitting time series models for longitudinal survey data under informative sampling”: [journal of statistical planning and inference 136 (2006) 3052–3069]. Journal of Statistical Planning and Inference, 137(2):628, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001935. See [EN06].

Eideh:2009:TSI

[EN09] Abdulhakeem Eideh and Gad Nathan. Two-stage informative cluster sampling-estimation and prediction with applications for small-area models. Journal of Statistical Planning and Inference, 139(9):3088–3101, September 1, 2009.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000834.

Elbarmi:2009:TAS

[ENAZ09] Hammou Elbarmi, Vicente Núñez-Antón, and Dale L. Zimmerman. Testing for and against a set of inequality constraints: the k-sample case. Journal of Statistical Planning and Inference, 139(3):1012–1022, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002863.

El-Nouty:2000:SB

[ENG00] Charles El-Nouty and Armelle Guillou. On the smoothed bootstrap. Journal of Statistical Planning and Inference, 83(1):203–220, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899000907.

Elal-Olivero:2009:BMU

[EOGQ09] David Elal-Olivero, Héctor W. Gómez, and Fernando A. Quintana. Bayesian modeling using a class of bimodal skew-elliptical distributions. Journal of Statistical Planning and Inference, 139(4):1484–1492, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003509.

Einmahl:2005:ARS

[EOPR05] J. H. J. Einmahl, B. O. Omolo, M. L. Puri, and F. H. Ruymgaart. Aligned rank statistics for repeated measurement models with orthonormal design, employing a Chernoff-Savage approach. Journal of Statistical Planning and Inference, 130(1–2):167–182, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002605.

Estevez-Perez:2002:CRC

[EPdRV02] G. Estévez-Pérez, A. Quintela del Riío, and P. Vieu. Convergence rate for cross-validatory bandwidth in kernel hazard estimation from dependent samples. Journal of Statistical Planning and Inference, 104(1):1–30, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002452.

Evans:2008:RFT

[EPS08] Morgan Evans, Robert Perlmutter, and Larry Shepp. Representing a function of two variables as a superposition over angles of ridge functions with a given shape via linear programming. Journal of Statistical Planning and Inference, 138(9):2662–2680, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800058X.

Ericksen:2002:SPU

[Eri02] Larry Ericksen. Sum of powers from uniform distribution moments. Journal of Statistical Planning and Inference, 101(1–2):57–79, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001549.

Eryilmaz:2003:RTM

[Ery03] S. Eryilmaz. Random threshold models based on multivariate observations. Journal of Statistical Planning and Inference, 113(2):557–568, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801003196.

Erkanli:2000:SBD

[ES00] Alaattin Erkanli and Refik Soyer. Simulation-based designs for accelerated life tests. Journal of Statistical Planning and Inference, 90(2):335–348, October 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001269.

Eaton:2002:GII

[ES02] Morris L. Eaton and William D. Sudderth. Group invariant inference and right Haar measure. Journal of Statistical Planning and Inference, 103(1–2):87–99, April 15, 2002.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001999.

Evans:2003:BMC

[ES03] Richard Evans and J. Sedransk. Bayesian methodology for combining the results from different experiments when the specifications for pooling are uncertain: II. Journal of Statistical Planning and Inference, 111(1–2):95–100, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002872.

Elamir:2004:EVS

[ES04] Elsayed A. H. Elamir and Allan H. Seheult. Exact variance structure of sample L-moments. Journal of Statistical Planning and Inference, 124(2):337–359, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002131.

El-Taha:2002:FAW

[ETS02] Muhammad El-Taha and Shaler Stidham. Filtration of ASTA: a weak convergence approach. Journal of Statistical Planning and Inference, 100(2):171–183, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001318.

Eguchi:2008:AIM

[EY08] Shinto Eguchi and Takemi Yanagimoto. Asymptotical improvement of maximum likelihood estimators on Kullback– Leibler loss. Journal of Statistical Planning and Inference, 138(11):3502–3511, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000931.

Elek:2009:MET

[EZ09] Péter Elek and András Zempléni. Modelling extremes of time-dependent data by Markov-switching structures. Journal of Statistical Planning and Inference, 139(6):1953–1967, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003625.

Fabian:2000:NMB

[Fab00] Václav Fabian. New modifications of the Bechhofer method. Journal of Statistical Planning and Inference, 91(2):313–322, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001841.

Fabian:2009:IUM

[Fab09] Zdenek Fabián. The t-information and its use in multivariate problems and time-series analysis. Journal of Statistical Planning and Inference, 139(11):3773–3778, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001426.

Fang:2000:RED

[Fan00] Zhide Fang. Robust extrapolation designs for biased polynomial models. Journal of Statistical Planning and Inference, 87(1):135–147, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001792.

Fang:2003:GTK

[Fan03] Yue Fang. GMM tests for the Katz family of distributions. Journal of Statistical Planning and Inference, 110(1–2):55–73, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002841.

Fan:2004:EPD

[Fan04] Zhaozhi Fan. Estimation problems for distributions with heavy tails. Journal of Statistical Planning and Inference, 123(1):13–40, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803001423.

Fang:2005:SSE

[Fan05] B. Q. Fang. The t statistic of the skew elliptical distributions. Journal of Statistical Planning and Inference, 134(1):140–157, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002101.

Frias:2008:SES

[FARMA08] M. P. Frías, F. J. Alonso, M. D. Ruiz-Medina, and J. M. Angulo. Semiparametric estimation of spatial long-range dependence. Journal of Statistical Planning and Inference, 138(5):1479–1495, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807002923.

Flanders:2003:ESM

[FBK03] W. Dana Flanders, Huiman X. Barnhart, and Andrzej S. Kosinski. An estimated score method for missing and auxiliary covariates in two stage studies. Journal of Statistical Planning and Inference, 115(1):145–170, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001490.

Fan:2003:GMS

[FC03] Shenghua Kelly Fan and Kathryn Chaloner. A geometric method for singular c-optimal designs. Journal of Statistical Planning and Inference, 113(1):249–257, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002890.

Fan:2004:ODL

[FC04] Shenghua Kelly Fan and Kathryn Chaloner. Optimal designs and limiting optimal designs for a trinomial response. Journal of Statistical Planning and Inference, 126(1):347–360, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002696.

Ferrari:2001:HOA

[FCCN01] Silvia L. P. Ferrari, Gauss M. Cordeiro, and Francisco Cribari-Neto. Higher-order asymptotic refinements for score tests in proper dispersion models. Journal of Statistical Planning and Inference, 97(1):177–190, August 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003529.

Ferrari:2004:ITH

[FCCN04] Silvia L. P. Ferrari, Audrey H. M. A. Cysneiros, and Francisco Cribari-Neto. An improved test for heteroskedasticity using adjusted modified profile likelihood inference. Journal of Statistical Planning and Inference, 124(2):423–437, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002106. See corri-

gendum [FCCN05].

Ferrari:2005:CIT

[FCCN05] Silvia L. P. Ferrari, Audrey H. M. A. Cysneiros, and Francisco Cribari-Neto. Corrigendum to “An improved test for heteroskedasticity using adjusted modified profile likelihood inference” [Journal of Statistical Planning and Inference 124 (2004) 423–437]. Journal of Statistical Planning and Inference, 131(1):207–208, April 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400415X. See [FCCN04].

Fu:2008:MLR

[FCL08] Yuejiao Fu, Jiahua Chen, and Pengfei Li. Modified likelihood ratio test for homogeneity in a mixture of von Mises distributions. Journal of Statistical Planning and Inference, 138(3):667–681, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000602.

Freitag:2007:NTS

[FCM07] Gudrun Freitag, Claudia Czado, and Axel Munk. A nonparametric test for similarity of marginals — with applications to the assessment of population bioequivalence. Journal of Statistical Planning and Inference, 137(3):697–711, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001352.

Ferrandiz:2005:TAC

[FCS05] Juan Ferrándiz, Enrique F. Castillo, and Pilar Sanmartín. Temporal aggregation in chain graph models. Journal of Statistical Planning and Inference, 133(1):69–93, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001156.

Figueroa-Centeno:2000:TMC

[FCW00] Ramón M. Figueroa-Centeno and Arthur T. White. Topological models for classical configurations. Journal of Statistical Planning and Inference, 86(2):421–434, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001226.

Fang:2002:NBR

[FD02] B. Q. Fang and A. P. Dawid. Nonconjugate Bayesian regression on many variables. Journal of Statistical Planning and Inference, 103(1–2):245–261, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002245.

Fujisawa:2006:REN

[FE06] Hironori Fujisawa and Shinto Eguchi. Robust estimation in the normal mixture model. Journal of Statistical Planning and Inference, 136(11):3989–4011, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001023.

Fox:2009:BCI

[FED09] Richard J. Fox, David Elgart, and S. Christopher Davis. Bayesian credible intervals for response surface optima. Journal of Statistical Planning and Inference, 139(7):2498–2501, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003947. See erra-

tum [Fox09].

Ferger:2001:EPT

[Fer01] Dietmar Ferger. Exponential and polynomial tailbounds for change-point estimators. Journal of Statistical Planning and Inference, 92(1–2):73–109, January 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001488.

Feretic:2002:BPE

[Fer02a] Svjetlan Feretić. A bijective perimeter enumeration of directed convex polyominoes. Journal of Statistical Planning and Inference, 101(1–2):81–94, February 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001550.

Ferreira:2002:OAE

[Fer02b] Ana Ferreira. Optimal asymptotic estimation of small exceedance probabilities. Journal of Statistical Planning and Inference, 104(1):83–102, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801002403.

Fernandez:2004:EEP

[Fer04] Arturo J. Fernández. On estimating exponential parameters with general type II progressive censoring. Journal of Statistical Planning and Inference, 121(1):135–147, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803000831.

Fujikoshi:2002:EPP

[FFFH02] Yasunori Fujikoshi, Kahoru Futagami, Junzou Fukunaga, and Takanori Hirata. An extension of the product-parts replacement problem. Journal of Statistical Planning and Inference, 106(1–2):473–484, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200229X.

Fuh:2000:BIR

[FFH00] C. D. Fuh, T. H. Fan, and W. L. Hung. Balanced importance resampling for Markov chains. Journal of Statistical Planning and Inference, 83(1):221–241, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000889.

Fattorini:2009:TPS

[FFP09] L. Fattorini, S. Franceschi, and C. Pisani. A two-phase sampling strategy for large-scale forest carbon budgets. Journal of Statistical Planning and Inference, 139(3):1045–1055, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002875.

Fortiana:2002:SFG

[FG02] J. Fortiana and A. Grané. A scale-free goodness-of-fit statistic for the exponential distribution based on maximum correlations. Journal of Statistical Planning and Inference, 108(1–2):85–97, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002720.

Farrow:2006:TSE

[FG06] Malcolm Farrow and Michael Goldstein. Trade-off sensitive experimental design: a multicriterion, decision theoretic, Bayes linear approach. Journal of Statistical Planning and Inference, 136(2):498–526, February 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003179.

Finner:2006:PPS

[FGS06] H. Finner, G. Giani, and K. Straßburger. Partitioning principle and selection of good treatments. Journal of Statistical Planning and Inference, 136(7):2053–2069, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002107.

Falk:2008:LBN

[FGT08] Michael Falk, Armelle Guillou, and Gwladys Toulemonde. A LAN based Neyman smooth test for Pareto distributions. Journal of Statistical Planning and Inference, 138 (10):2867–2886, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807003795.

Futschik:2005:CSM

[FH05] Andreas Futschik and Wen-Tao Huang. Confidence sets for the maximizers of intensity functions. Journal of Statistical Planning and Inference, 134(2):549–567, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002149.

Friedman:2007:BNE

[FH07] Jerome H. Friedman and Peter Hall. On bagging and nonlinear estimation. Journal of Statistical Planning and Inference, 137(3):669–683, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001339.

Fu:2004:DCG

[FHJ^{+}04] H. L. Fu, F. K. Hwang, M. Jimbo, Y. Mutoh, and C. L. Shiue. Decomposing complete graphs into K_{r} × K_{c}’s. Journal of Statistical Planning and Inference, 119(2):225–236, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004214.

Fuji-Hara:2002:BND

[FHKMS02] Ryoh Fuji-Hara, Shinji Kuriki, Ying Miao, and Satoshi Shinohara. Balanced nested designs and balanced n-ary designs. Journal of Statistical Planning and Inference, 106(1–2):57– 67, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002021.

Fedorov:2003:CWD

[FHL03] Valerii V. Fedorov, Agnes M. Herzberg, and Sergei L. Leonov. Component-wise dimension reduction. Journal of Statistical Planning and Inference, 114(1–2):81–93, June 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004640.

Fuji-Hara:2000:BAQ

[FHM00] Ryoh Fuji-Hara and Nobuko Miyamoto. Balanced arrays from quadratic functions. Journal of Statistical Planning and Inference, 84(1–2):285–293, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001123.

Fuji-Hara:2002:CSD

[FHMS02] R. Fuji-Hara, Y. Miao, and S. Shinohara. Complete sets of disjoint difference families and their applications. Journal of Statistical Planning and Inference, 106(1–2):87–103, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002057.

Fujikoshi:2008:ARC

[FHW08] Yasunori Fujikoshi, Tetsuto Himeno, and Hirofumi Wakaki. Asymptotic results in canonical discriminant analysis when the dimension is large compared to the sample size. Journal of Statistical Planning and Inference, 138(11):3457–3466, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800089X.

Fuh:2005:BBB

[FI05] Cheng-Der Fuh and Edward H. Ip. Bootstrap and Bayesian bootstrap clones for censored Markov chains. Journal of Statistical Planning and Inference, 128(2):459–474, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000059.

Field:2004:UDM

[Fie04] C. A. Field. Using the gh distribution to model extreme wind speeds. Journal of Statistical Planning and Inference, 122(1–2):15–22, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580300243X.

Finkelstein:2003:SBT

[Fin03] M. S. Finkelstein. Simple bounds for terminating Poisson and renewal shock processes. Journal of Statistical Planning and Inference, 113(2):541–549, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001118.

Finkelstein:2009:DAB

[Fin09] Maxim Finkelstein. On damage accumulation and biological aging. Journal of Statistical Planning and Inference, 139 (5):1643–1648, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002401.

Fotopoulos:2009:ASC

[FJT09] Stergios B. Fotopoulos, Venkata K. Jandhyala, and Li Tan. Asymptotic study of the change-point mle in multivariate Gaussian families under contiguous alternatives. Journal of Statistical Planning and Inference, 139(3):1190–1202, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808003005.

Fahrmeir:2009:PPS

[FK09] Ludwig Fahrmeir and Thomas Kneib. Propriety of posteriors in structured additive regression models: Theory and empirical evidence. Journal of Statistical Planning and Inference, 139(3):843–859, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002565.

Furino:2002:FBS

[FKL^{+}02] Steven Furino, Sanpei Kageyama, Alan C. H. Ling, Ying Miao, and Jianxing Yin. Frames with block size four and index three. Journal of Statistical Planning and Inference, 106(1–2):117–124, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002070.

Fan:2003:SHO

[FL03] Yanqin Fan and Oliver Linton. Some higher-order theory for a consistent non-parametric model specification test. Journal of Statistical Planning and Inference, 109(1–2):125–154, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003075.

Fu:2008:DLL

[FL08] James C. Fu and W. Y. Wendy Lou. Distribution of the length of the longest common subsequence of two multistate biological sequences. Journal of Statistical Planning and Inference, 138(11):3605–3615, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800102X.

Fard:2009:OSS

[FL09] Nasser Fard and Chenhua Li. Optimal simple step stress accelerated life test design for reliability prediction. Journal of Statistical Planning and Inference, 139(5):1799–1808, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002541.

Ferrari:2005:IPL

[FLCN05] Silvia L. P. Ferrari, Fernando Lucambio, and Francisco Cribari-Neto. Improved profile likelihood inference. Journal of Statistical Planning and Inference, 134(2):373–391, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580400223X.

Fang:2000:CML

[FLM00] Kai-Tai Fang, Dennis K. J. Lin, and Chang-Xing Ma. On the construction of multi-level supersaturated designs. Journal of Statistical Planning and Inference, 86(1):239–252, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001639.

Felix-Medina:2000:AER

[FM00a] Martín H. Félix-Medina. Analytical expressions for Rao-Blackwell estimators in adaptive cluster sampling. Journal of Statistical Planning and Inference, 84(1–2):221–236, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837589900138X.

Fujioka:2000:EHO

[FM00b] Yumi Fujioka and Yoshihiko Maesono. Erratum to: “Higher order normalizing transformations of asymptotic U-statistics for removing bias, skewness and kurtosis”: [J. Statist. Plann. Inference 83 (2000) 47–74]. Journal of Statistical Planning and Inference, 87(2):365–366, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000720. See [FM00c].

Fujioka:2000:HON

[FM00c] Yumi Fujioka and Yoshihiko Maesono. Higher order normalizing transformations of asymptotic U-statistics for removing bias, skewness and kurtosis. Journal of Statistical Planning and Inference, 83(1):47–74, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000877. See erratum [FM00b].

Feddag:2005:GEE

[FM05a] Mohand L. Feddag and Mounir Mesbah. Generalized estimating equations for longitudinal mixed Rasch model. Journal of Statistical Planning and Inference, 129(1–2):159–179, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001600.

Fernholz:2005:CLI

[FM05b] Luisa T. Fernholz and Stephan Morgenthaler. Conditionally least-informative distributions: the case of the location parameter. Journal of Statistical Planning and Inference, 135(2):357–370, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002447.

Frey:2009:MAC

[FMN09] Jesse Frey, Osvaldo Marrero, and Douglas Norton. Minimumarea confidence sets for a normal distribution. Journal of Statistical Planning and Inference, 139(3):1023–1032, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002814.

Fourdrinier:2004:IPS

[FMS04] Dominique Fourdrinier, Eric Marchand, and William E. Strawderman. On the inevitability of a paradox in shrinkage estimation for scale mixtures of normals. Journal of Statistical Planning and Inference, 121(1):37–51, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001095.

Fernholz:2004:ONM

[FMT04] Luisa T. Fernholz, Stephan Morgenthaler, and John W. Tukey. An outlier nomination method based on the multihalver. Journal of Statistical Planning and Inference, 122(1–2):125–139, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002349.

Favre:2005:CRI

[FMT05] Anne-Catherine Favre, Alina Matei, and Yves Tillé. Calibrated random imputation for qualitative data. Journal of Statistical Planning and Inference, 128(2):411–425, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000047.

Fredricks:2007:RBS

[FN07] Gregory A. Fredricks and Roger B. Nelsen. On the relationship between Spearman’s rho and Kendall’s tau for pairs of continuous random variables. Journal of Statistical Planning and Inference, 137(7):2143–2150, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002588.

Forchini:2009:SPT

[For09] G. Forchini. Some properties of tests for parameters that can be arbitrarily close to being unidentified. Journal of Statistical Planning and Inference, 139(9):3193–3199, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000846.

Fourdrinier:2008:EMV

[FOS08] Dominique Fourdrinier, Idir Ouassou, and William E. Strawderman. Estimation of a mean vector under quartic loss. Journal of Statistical Planning and Inference, 138(12):3841–3857, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001390.

Fotopoulos:2000:IPD

[Fot00] Stergios B. Fotopoulos. Invariance principles for deconvolving kernel density estimation for stationary sequences of random variables. Journal of Statistical Planning and Inference, 86(1):31–50, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001627.

Foucart:2000:DRD

[Fou00] T. Foucart. A decision rule for discarding principal components in regression. Journal of Statistical Planning and Inference, 89(1–2):187–195, August 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002165.

Fox:2009:EBC

[Fox09] Richard Fox. Erratum “Bayesian credible intervals for response surface optima” [J. Statist. Plann. Inference 139 (2009) 2498–2501]. Journal of Statistical Planning and Inference, 139(9):3391, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000913. See [FED09].

Ferrari:2005:LLP

[FP05] Luca Ferrari and Renzo Pinzani. Lattices of lattice paths. Journal of Statistical Planning and Inference, 135(1):77–92, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580500056X.

Fontana:2000:CTL

[FPR00a] Roberto Fontana, Giovanni Pistone, and Maria Piera Rogantin. Classification of two-level factorial fractions. Journal of Statistical Planning and Inference, 87(1):149–172, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001731.

Fourdrinier:2000:ENC

[FPR00b] D. Fourdrinier, A. Philippe, and C. P. Robert. Estimation of a non-centrality parameter under Stein-type-like losses. Journal of Statistical Planning and Inference, 87(1):43–54, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001767.

Fiocco:2006:RRP

[FPvdVvH06] M. Fiocco, H. Putter, C. J. H. van de Velde, and J. C. van Houwelingen. Reduced rank proportional hazards model for competing risks: an application to a breast cancer trial. Journal of Statistical Planning and Inference, 136(5):1655–1668, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001734.

Franek:2000:TFS

[FR00] F. Franek and A. Rosa. Two-factorizations of small complete graphs. Journal of Statistical Planning and Inference, 86(2):435–442, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001238.

See corrigendum [FR08].

Finner:2001:ASP

[FR01] H. Finner and M. Roters. Asymptotic sharpness of producttype inequalities for maxima of random variables with applications in multiple comparisons. Journal of Statistical Planning and Inference, 98(1–2):39–56, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003049.

Fraser:2002:SMF

[FR02] D. A. S. Fraser and N. Reid. Strong matching of frequentist and Bayesian parametric inference. Journal of Statistical Planning and Inference, 103(1–2):263–285, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002257.

Flournoy:2006:P

[FR06] Nancy Flournoy and W. F. Rosenberger. Preface. Journal of Statistical Planning and Inference, 136(6):1747–1748, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001825.

Fienberg:2007:TCC

[FR07] Stephen E. Fienberg and Alessandro Rinaldo. Three centuries of categorical data analysis: Log-linear models and maximum likelihood estimation. Journal of Statistical Planning and Inference, 137(11):3430–3445, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001073.

Franek:2008:CTF

[FR08] F. Franek and A. Rosa. Corrigendum to: “ Two-factorizations of small complete graphs”: J. Statist. Plann. Inference 86 (2000), 435–442. Journal of Statistical Planning and Inference, 138(12):4198, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002632. See [FR00].

Freedman:2001:CIS

[Fre01] Laurence S. Freedman. Confidence intervals and statistical power of the ‘validation’ ratio for surrogate or intermediate endpoints. Journal of Statistical Planning and Inference, 96(1):143–153, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580000330X.

See erratum [Fre02].

Freedman:2002:ECI

[Fre02] Laurence S. Freedman. Erratum to “Confidence intervals and statistical power of the ‘Validation’ ratio for surrogate or intermediate endpoints”: [J. Statist. Plann. Inference 96 (2001) 143–153]. Journal of Statistical Planning and Inference, 100(2):443, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100307X. See [Fre01].

Freue:2007:PEC

[Fre07a] Gabriela V. Cohen Freue. The Pitman estimator of the Cauchy location parameter. Journal of Statistical Planning and Inference, 137(6):1900–1913, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001285.

Frey:2007:NIR

[Fre07b]  Jesse C. Frey.  New imperfect rankings models for

ranked set sampling. Journal of Statistical Planning and Inference, 137(4):1433–1445, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001054.

Frey:2008:ODF

[Fre08] Jesse Frey. Optimal distribution-free confidence bands for a distribution function. Journal of Statistical Planning and Inference, 138(10):3086–3098, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004235.

Frey:2009:UGF

[Fre09] Jesse Frey. Unbiased goodness-of-fit tests. Journal of Statistical Planning and Inference, 139(10):3690–3697, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001189.

Fierens:2009:FUS

[FRF09] P. I. Fierens, L. C. Rêgo, and T. L. Fine. A frequentist understanding of sets of measures. Journal of Statistical Planning and Inference, 139(6):1879–1892, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800356X.

Finkelstein:2001:CTS

[FS01a] Dianne M. Finkelstein and David A. Schoenfeld. Comparing treatment strategies using a synthesized clinical trial: an analysis of late versus early use of trimethoprimsulfamethoxazole for AIDS patients. Journal of Statistical Planning and Inference, 96(1):179–189, June 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003347.

Fu:2001:NCE

[FS01b] Fang-Wei Fu and Shi-Yi Shen. On the nonperiodic cyclic equivalence classes of Hamming codes and BCH codes. Journal of Statistical Planning and Inference, 94(2):205–209, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002536.

Fu:2001:UBB

[FS01c] Fang-Wei Fu and Shi-Yi Shen. An upper bound for binary constant weight codes. Journal of Statistical Planning and Inference, 94(2):197–203, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002524.

Finner:2002:SPU

[FS02a] H. Finner and K. Straßburger. Structural properties of UMPU-tests for 2×2 tables and some applications. Journal of Statistical Planning and Inference, 104(1):103–120, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001227.

Fu:2002:CRM

[FS02b] Limin Fu and Douglas G. Simpson. Conditional risk models for ordinal response data: simultaneous logistic regression analysis and generalized score tests. Journal of Statistical Planning and Inference, 108(1–2):201–217, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002793.

Finos:2007:FFC

[FS07] Livio Finos and Luigi Salmaso. FDR- and FWE-controlling methods using data-driven weights. Journal of Statistical Planning and Inference, 137(12):3859–3870, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001553.

Finos:2007:CIU

[FSS07] Livio Finos, Luigi Salmaso, and Aldo Solari. Conditional inference under simultaneous stochastic ordering constraints. Journal of Statistical Planning and Inference, 137 (8):2633–2641, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000158.

Fountain:2009:WSO

[FT09] R. L. Fountain and A. Talebi. “Weak” stochastic orderings and dependence measures. Journal of Statistical Planning and Inference, 139(11):3779–3784, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001438.

Fernholz:2004:CDA

[FTY04] Luisa Turrin Fernholz, David E. Tyler, and Victor Yohai. Contemporary data analysis: theory and methods. Journal of Statistical Planning and Inference, 122(1–2):1–2, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002489.

Fang:2008:LBV

[FTY08] Kai-Tai Fang, Yu Tang, and Jianxing Yin. Lower bounds of various criteria in experimental designs. Journal of Statistical Planning and Inference, 138(1):184–195, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700208X.

Fu:2005:NGQ

[Fu05] Wenjiang J. Fu. Nonlinear GCV and quasi-GCV for shrinkage models. Journal of Statistical Planning and Inference, 131(2):333–347, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804000825.

Fuentes:2000:PID

[Fue00] Montserrat Fuentes. Predicting integrals of diffusion processes. Journal of Statistical Planning and Inference, 90 (2):183–193, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580000121X.

Fuentes:2006:TSS

[Fue06] Montserrat Fuentes. Testing for separability of spatialtemporal covariance functions. Journal of Statistical Planning and Inference, 136(2):447–466, February 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003210.

Fujikoshi:2002:AED

[Fuj02] Yasunori Fujikoshi. Asymptotic expansions for the distributions of multivariate basic statistics and one-way MANOVA tests under nonnormality. Journal of Statistical Planning and Inference, 108(1–2):263–282, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003130.

Fang:2003:RRD

[FW03] Zhide Fang and Douglas P. Wiens. Robust regression designs for approximate polynomial models. Journal of Statistical Planning and Inference, 117(2):305–321, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003877.

Feng:2007:LRT

[FW07] Yanqin Feng and Jinde Wang. Likelihood ratio test against simple stochastic ordering among several multinomial populations. Journal of Statistical Planning and Inference, 137(4):1362–1374, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000899.

Fan:2008:EPS

[FWB08] Tsai-Hung Fan, Wan-Lun Wang, and N. Balakrishnan. Exponential progressive step-stress life-testing with link function based on Box–Cox transformation. Journal of Statistical Planning and Inference, 138(8):2340–2354, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003722.

Fang:2006:LOD

[FWW06] Zhide Fang, Douglas P. Wiens, and Zheyang Wu. Locally D-optimal designs for multistage models and heteroscedastic polynomial regression models. Journal of Statistical Planning and Inference, 136(11):4059–4070, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001163.

Fu:2009:WEL

[FWW09] Yuejiao Fu, Xiaogang Wang, and Changbao Wu. Weighted empirical likelihood inference for multiple samples. Journal of Statistical Planning and Inference, 139(4):1462–1473, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003480.

Fang:2003:PCF

[FWYY03] Xiangzhong Fang, Ray Watson, Wang Yan, and Paul S. F. Yip. A procedure for complete fault detection with a removal process. Journal of Statistical Planning and Inference, 117(1):1–14, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003634.

Fujii:2005:PCS

[FY05] Yoshinori Fujii and Takemi Yanagimoto. Pairwise conditional score functions: a generalization of the Mantel–Haenszel estimator. Journal of Statistical Planning and Inference, 128(1):1–12, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803003318.

Francq:2000:CME

[FZ00] Christian Francq and Jean-Michel Zakoïan. Covariance matrix estimation for estimators of mixing weak ARMA models. Journal of Statistical Planning and Inference, 83(2):369–394, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899001093.

Field:2003:CIB

[FZ03] Christopher Field and Julie Zhou. Confidence intervals based on robust regression. Journal of Statistical Planning and Inference, 115(2):425–439, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001684.

Fang:2004:MAM

[FZ04] Kai-Tai Fang and Aijun Zhang. Minimum aberration majorization in non-isomorphic saturated designs. Journal of Statistical Planning and Inference, 126(1):337–346, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002611.

Fang:2006:ADL

[FZ06] Yixin Fang and Lincheng Zhao. Approximation to the distribution of LAD estimators for censored regression by random weighting method. Journal of Statistical Planning and Inference, 136(4):1302–1316, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003787.

Fakhre-Zakeri:2000:FLT

[FZL00] Issa Fakhre-Zakeri and Sangyeol Lee. On functional limit theorems for multivariate linear processes with applications to sequential estimation. Journal of Statistical Planning and Inference, 83(1):11–23, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000543.

Gil-Alana:2005:DSV

[GA05] Luis A. Gil-Alana. Deterministic seasonality versus seasonal fractional integration. Journal of Statistical Planning and Inference, 134(2):445–461, October 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002174.

Gao:2008:NMCb

[GA08] Xin Gao and Mayer Alvo. Nonparametric multiple comparison procedures for unbalanced two-way layouts. Journal of Statistical Planning and Inference, 138(12):3674–3686, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001237.

Gao:2008:NMCa

[GACL08] Xin Gao, Mayer Alvo, Jie Chen, and Gang Li. Nonparametric multiple comparison procedures for unbalanced one-way factorial designs. Journal of Statistical Planning and Inference, 138(8):2574–2591, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003886.

George:2007:RMA

[GACS07] E. Olusegun George, Dale Armstrong, Paul J. Catalano, and Deo Kumar Srivastava. Regression models for analyzing clustered binary and continuous outcomes under an assumption of exchangeability. Journal of Statistical Planning and Inference, 137(11):3462–3474, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001097.

Gupta:2008:Pa

[GAD08] Sudhir Gupta, Manohar Aggarwal, and Lih-Yuan Deng. Preface. Journal of Statistical Planning and Inference, 138(1):1, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002194.

Ganju:2000:CBF

[Gan00] Jitendra Ganju. On choosing between fixed and random block effects in some no-interaction models. Journal of Statistical Planning and Inference, 90(2):323–334, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001221.

Garel:2001:LRT

[Gar01] Bernard Garel. Likelihood ratio test for univariate Gaussian mixture. Journal of Statistical Planning and Inference, 96(2):325–350, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002160.

Gardy:2002:OUM

[Gar02] Danièle Gardy. Occupancy urn models in the analysis of algorithms. Journal of Statistical Planning and Inference, 101(1–2):95–105, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001562.

Garel:2005:ATL

[Gar05] Bernard Garel. Asymptotic theory of the likelihood ratio test for the identification of a mixture. Journal of Statistical Planning and Inference, 131(2):271–296, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000801.

Gastwirth:2007:GSD

[Gas07] Joseph L. Gastwirth. A graphical summary of disparities in health care and related summary measures. Journal of Statistical Planning and Inference, 137(3):1059–1065, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001601.

Gimenez:2000:CCE

[GB00] Patricia Gimenez and Heleno Bolfarine. Comparing consistent estimators in comparative calibration models. Journal of Statistical Planning and Inference, 86(1):143–155, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001718.

Gill:2004:CEG

[GB04] Ryan Gill and Michael Baron. Consistent estimation in generalized broken-line regression. Journal of Statistical Planning and Inference, 126(2):441–460, December 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002763.

Gel:2007:SCR

[GB07] Yulia R. Gel and Andrey Barabanov. Strong consistency of the regularized least-squares estimates of infinite autoregressive models. Journal of Statistical Planning and Inference, 137(4):1260–1277, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000620.

Gaddah:2009:WCC

[GB09] Auguste Gaddah and Roel Braekers. Weak convergence for the conditional distribution function in a Koziol–Green model under dependent censoring. Journal of Statistical Planning and Inference, 139(3):930–943, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002681.

Gibbons:2005:SPB

[GBC^{+}05] Robert D. Gibbons, Dulal K. Bhaumik, David R. Cox, Dennis R. Grayson, John M. Davis, and Rajiv P. Sharma. Sequential prediction bounds for identifying differentially expressed genes in replicated microarray experiments. Journal of Statistical Planning and Inference, 129(1–2):19–37, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001521.

Gustafson:2004:DPV

[GC04] Paul Gustafson and Bertrand Clarke. Decomposing posterior variance. Journal of Statistical Planning and Inference, 119(2):311–327, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004913.

Ghosh:2005:STI

[GC05] Subir Ghosh and Heather R. Crosby. Subject-treatment interactions in crossover trials: performance evaluation of subgrouping methods. Journal of Statistical Planning and Inference, 132(1–2):63–73, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001429.

Grilo:2007:DCS

[GC07] Luís M. Grilo and Carlos A. Coelho. Development and comparative study of two near-exact approximations to the distribution of the product of an odd number of independent beta random variables. Journal of Statistical Planning and Inference, 137(5):1560–1575, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580600228X.

Gomes:2009:GIV

[GC09] Dulce Gomes and Luísa Canto e Castro. Generalized integervalued random coefficient for a first order structure autoregressive (RCINAR) process. Journal of Statistical Planning and Inference, 139(12):4088–4097, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001633.

Ghosh:2006:FDR

[GCR06] Debashis Ghosh, Wei Chen, and Trivellore Raghunathan. The false discovery rate: a variable selection perspective. Journal of Statistical Planning and Inference, 136(8):2668–2684, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804004380.

Govindarajulu:2002:SEV

[GD02] Z. Govindarajulu and Alexei Dmitrienko. Sequential estimation of the variance of an unknown distribution. Journal of Statistical Planning and Inference, 100(2):221–237, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001367.

Ghosh:2003:ODC

[GD03] Himadri Ghosh and Ashish Das. Optimal diallel cross designs for estimation of heritability. Journal of Statistical Planning and Inference, 116(1):185–196, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001805.

Gagne:2008:RFD

[GD08a] Christian Gagné and Pierre Duchesne. On robust forecasting in dynamic vector time series models. Journal of Statistical Planning and Inference, 138(12):3927–3938, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001468.

Giordan:2008:TEM

[GD08b] Marco Giordan and Giancarlo Diana. Tests for the equality of multinomial parameters. Journal of Statistical Planning and Inference, 138(12):3789–3796, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001377.

Ghosh:2007:ENP

[GDL07a] Subir Ghosh, Hongjie Deng, and Ying Luan. Effects of noise in performance comparisons of designs for model identification and discrimination. Journal of Statistical Planning and Inference, 137(12):3871–3881, December 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001565.

Ghosh:2007:ESS

[GDL07b] Subir Ghosh, Hongjie Deng, and Ying Luan. Error sum of squares and power comparisons for model search, identification and discrimination. Journal of Statistical Planning and Inference, 137(11):3475–3482, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001103.

Gupta:2005:P

[GDM^{+}05] Sudhir Gupta, Dipak Dey, Glen Meeden, Rahul Mukerjee, and Alan Polansky. Preface. Journal of Statistical Planning and Inference, 129(1–2):1, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002046.

Ge:2001:EHL

[Ge01a] Gennian Ge. Existence of holey LSSOM of type 2^{n} with application to G_{7}-packings of k_{v}. Journal of Statistical Planning and Inference, 94(2):211–218, April 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002548.

Ge:2001:ERF

[Ge01b] Gennian Ge. On the existence of Room frames of type 2^{n}u^{1}. Journal of Statistical Planning and Inference, 94(2):219–230, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580000255X.

Genton:2001:CVF

[Gen01] Marc G. Genton. The change-of-variance function: a tool to explore the effects of dependencies in spatial statistics. Journal of Statistical Planning and Inference, 98(1–2):191–209, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002986.

Georgiou:2009:OLH

[Geo09] Stelios D. Georgiou. Orthogonal Latin hypercube designs from generalized orthogonal designs. Journal of Statistical Planning and Inference, 139(4):1530–1540, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003571.

Gervini:2005:RAE

[Ger05]  Daniel Gervini.  Robust adaptive estimators for bi-

nary regression models. Journal of Statistical Planning and Inference, 131(2):297–311, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000849.

Ghosh:2000:IIB

[GF00] Subir Ghosh and Lisa D. Fairchild. Inference on interactions between treatments and subjects within groups in a two-period crossover trial. Journal of Statistical Planning and Inference, 88(2):301–317, August 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000859.

Gezmu:2006:GDD

[GF06] Misrak Gezmu and Nancy Flournoy. Group up-and-down designs for dose-finding. Journal of Statistical Planning and Inference, 136(6):1749–1764, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001837.

Gonzalez-Farias:2004:APS

[GFDMG04] Graciela González-Farías, Armando Domínguez-Molina, and Arjun K. Gupta. Additive properties of skew normal random vectors. Journal of Statistical Planning and Inference, 126(2):521–534, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002805.

Gomes:2005:ABL

[GFM05] M. Ivette Gomes, Fernanda Figueiredo, and Sandra Mendonça. Asymptotically best linear unbiased tail estimators under a second-order regular variation condition. Journal of Statistical Planning and Inference, 134(2):409–433, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002198.

Godbole:2001:CPZ

[GG01] Anant P. Godbole and Heidi C. Graziano. Contributions to the problem of Zarankiewicz. Journal of Statistical Planning and Inference, 95(1–2):197–208, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002883.

Gupta:2007:PRH

[GG07] Ramesh C. Gupta and Rameshwar D. Gupta. Proportional reversed hazard rate model and its applications. Journal of Statistical Planning and Inference, 137(11):3525–3536, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001140.

Gardes:2008:EWT

[GG08] Laurent Gardes and Stéphane Girard. Estimation of the Weibull tail-coefficient with linear combination of upper order statistics. Journal of Statistical Planning and Inference, 138(5):1416–1427, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807002716.

Gu:2009:BRC

[GG09a] Jiezhun Gu and Subhashis Ghosal. Bayesian ROC curve estimation under binormality using a rank likelihood. Journal of Statistical Planning and Inference, 139(6):2076–2083, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003819.

Gupta:2009:DLF

[GG09b] Pushpa L. Gupta and Ramesh C. Gupta. Distribution of a linear function of correlated ordered variables. Journal of Statistical Planning and Inference, 139(7):2490–2497, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808003959.

Gupta:2009:GFM

[GG09c] Ramesh C. Gupta and Rameshwar D. Gupta. General frailty model and stochastic orderings. Journal of Statistical Planning and Inference, 139(9):3277–3287, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900072X.

Ghosh:2000:NPO

[GGCA00] Malay Ghosh, Atalanta Ghosh, Ming-Hui Chen, and Alan Agresti. Noninformative priors for one-parameter item response models. Journal of Statistical Planning and Inference, 88(1):99–115, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899002013.

Grasshoff:2004:ODM

[GGHS04] Ulrike Graßhoff, Heiko Großmann, Heinz Holling, and Rainer Schwabe. Optimal designs for main effects in linear paired comparison models. Journal of Statistical Planning and Inference, 126(1):361–376, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002283.

Grasshoff:2007:DOM

[GGHS07] Ulrike Graßhoff, Heiko Großmann, Heinz Holling, and Rainer Schwabe. Design optimality in multi-factor generalized linear models in the presence of an unrestricted quantitative factor. Journal of Statistical Planning and Inference, 137(12):3882–3893, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001577.

Gupta:2002:ESL

[GGS02] Sat Gupta, Bhisham Gupta, and Sarjinder Singh. Estimation of sensitivity level of personal interview survey questions. Journal of Statistical Planning and Inference, 100(2):239–247, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001379.

Guo:2003:CCL

[GGS03] Jianhua Guo, Zhi Geng, and Ning-Zhong Shi. Consecutive collapsibility of logistic regression coefficients over an ordinal background. Journal of Statistical Planning and Inference, 115(1):59–67, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802001179.

Garroi:2009:VNS

[GGS09a] Jean-Jacques Garroi, Peter Goos, and Kenneth Sörensen. A variable-neighbourhood search algorithm for finding optimal run orders in the presence of serial correlation. Journal of Statistical Planning and Inference, 139(1):30–44, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002279.

Grossmann:2009:AEO

[GGS09b] Heiko Großmann, Ulrike Graßhoff, and Rainer Schwabe. Approximate and exact optimal designs for paired comparisons of partial profiles when there are two groups of factors. Journal of Statistical Planning and Inference, 139(3):1171–1179, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002991.

Gurtler:2000:RCG

[GH00] Nora Gürtler and Norbert Henze. Recent and classical goodness-of-fit tests for the Poisson distribution. Journal of Statistical Planning and Inference, 90(2):207–225, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001142.

Guo:2001:SCF

[GH01] Haitao Guo and Tayuan Huang. Some classes of four-weight spin models. Journal of Statistical Planning and Inference, 94(2):231–247, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002561.

Gombay:2002:RCS

[GH02] Edit Gombay and Lajos Horváth. Rates of convergence for Ustatistic processes and their bootstrapped versions. Journal of Statistical Planning and Inference, 102(2):247–272, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801000854.

Ghosh:2003:NPC

[GH03] Malay Ghosh and Jungeun Heo. Noninformative priors, credible sets and Bayesian hypothesis testing for the intraclass model. Journal of Statistical Planning and Inference, 112(1–2):133–146, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003282.

Gutman:2007:IMT

[GH07] Roee Gutman and Yosef Hochberg. Improved multiple test procedures for discrete distributions: New ideas and analytical review. Journal of Statistical Planning and Inference, 137(7):2380–2393, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002746.

Grabovsky:2000:LTK

[GHH00] Irina Grabovsky, Lajos Horváth, and Marie Husková. Limit theorems for kernel-type estimators for the time of change. Journal of Statistical Planning and Inference, 89(1–2):25–56, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001002.

Gaile:2007:ECA

[GHJ^{+}07] Daniel P. Gaile, Alan Hutson, James Java, Jeffrey M. Conroy, Devin McQuaid, Michael Rossi, and Norma J. Nowak. Errors in centering of array data can induce biases in correlation estimates. Journal of Statistical Planning and Inference, 137(11):3446–3461, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001085.

Georgiou:2007:TVF

[GHKTR07] S. Georgiou, W. H. Holzmann, H. Kharaghani, and B. Tayfeh-Rezaie. Three variable full orthogonal designs of order 56. Journal of Statistical Planning and Inference, 137(2):611–618, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000346.

Gupta:2003:SPP

[GHL03] Shanti S. Gupta, Shuyuan He, and Jianjun Li. On selection procedures for positive exponential family distributions based on type-I censored data. Journal of Statistical Planning and Inference, 110(1–2):11–21, January 15, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002865.

Ghosh:2006:BBS

[GHMR06] J. K. Ghosh, N. L. Hjort, C. Messan, and R. V. Ramamoorthi. Bayesian bivariate survival estimation. Journal of Statistical Planning and Inference, 136(7):2297–2308, July 2006.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002235.

Ghosh:2001:NTE

[Gho01] Sucharita Ghosh. Nonparametric trend estimation in replicated time series. Journal of Statistical Planning and Inference, 97(2):263–274, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002226.

Ghosh:2002:LR

[Gho02] Subir Ghosh. List of referees 2001. Journal of Statistical Planning and Inference, 102(2):iii, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802000769.

Guiard:2000:AOD

[GHSN00] V. Guiard, G. Herrendörfer, D. Sumpf, and G. Nürnberg. About optimal designs for estimating variance components with ANOVA in one-way-classification under non-normality. Journal of Statistical Planning and Inference, 89(1–2):269– 285, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899002177.

Gupta:2002:OAP

[GHSS02] Sudhir Gupta, Balakrishna Hosmane, Murari Singh, and Rajendra Singh. Optimal allocation of patients in titration to fixed dose trials. Journal of Statistical Planning and Inference, 106(1–2):215–223, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002148.

Gadbury:2004:ITE

[GIA04] Gary L. Gadbury, Hari K. Iyer, and Jeffrey M. Albert. Individual treatment effects in randomized trials with binary outcomes. Journal of Statistical Planning and Inference, 121(2):163–174, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001150.

Gilbert:2004:GFT

[Gil04] Peter B. Gilbert. Goodness-of-fit tests for semiparametric biased sampling models. Journal of Statistical Planning and Inference, 118(1–2):51–81, January 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004056.

Gilardoni:2005:APA

[Gil05] Gustavo L. Gilardoni. Accuracy of posterior approximations via χ^{2} and harmonic divergences. Journal of Statistical Planning and Inference, 128(2):475–487, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000266.

Gillispie:2006:FCA

[Gil06] Steven B. Gillispie. Formulas for counting acyclic digraph Markov equivalence classes. Journal of Statistical Planning and Inference, 136(4):1410–1432, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003982.

Ghosh:2001:GEM

[GJ01] Kaushik Ghosh and S. Rao Jammalamadaka. A general estimation method using spacings. Journal of Statistical Planning and Inference, 93(1–2):71–82, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001609.

Girard:2003:PEP

[GJ03] Stéphane Girard and Pierre Jacob. Projection estimates of point processes boundaries. Journal of Statistical Planning and Inference, 116(1):1–15, September 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001829.

Glickman:2005:APC

[GJ05] Mark E. Glickman and Shane T. Jensen. Adaptive paired comparison design. Journal of Statistical Planning and Inference, 127(1–2):279–293, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003215.

Gupta:2000:RCV

[GK00] Ramesh C. Gupta and S. N. U. A. Kirmani. Residual coefficient of variation and some characterization results. Journal of Statistical Planning and Inference, 91(1):23–31, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001348.

Gansky:2001:ABM

[GK01] Stuart A. Gansky and Gary G. Koch. Assessing bias of multicenter trials with incomplete treatment allocation. Journal of Statistical Planning and Inference, 96(1):83–107, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580000327X.

Ghosh:2002:RBA

[GK02] Malay Ghosh and Dal Ho Kim. Robust Bayesian analysis with partially exchangeable priors. Journal of Statistical Planning and Inference, 102(1):99–107, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001744.

Georgiou:2005:CDC

[GK05] S. Georgiou and C. Koukouvinos. Combinatorial designs and codes over some prime fields. Journal of Statistical Planning and Inference, 135(1):93–106, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000571.

Gupta:2006:SCF

[GK06a] Ramesh C. Gupta and S. N. U. A. Kirmani. Stochastic comparisons in frailty models. Journal of Statistical Planning and Inference, 136(10):3647–3658, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000819.

Gupta:2006:CFI

[GK06b] Rameshwar D. Gupta and Debasis Kundu. On the comparison of Fisher information of the Weibull and GE distributions. Journal of Statistical Planning and Inference, 136(9):3130–3144, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000029.

Guo:2007:NRH

[GK07a] Hongwen Guo and Hira L. Koul. Nonparametric regression with heteroscedastic long memory errors. Journal of Statistical Planning and Inference, 137(2):379–404, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580600084X.

Gupta:2007:GED

[GK07b] Rameshwar D. Gupta and Debasis Kundu. Generalized exponential distribution: Existing results and some recent developments. Journal of Statistical Planning and Inference, 137(11):3537–3547, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001152.

Gupta:2008:CBC

[GK08] Ramesh C. Gupta and S. N. U. A. Kirmani. Characterization based on convex conditional mean function. Journal of Statistical Planning and Inference, 138(4):964–970, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002777.

Gasbarra:2006:TDB

[GKDN06] Dario Gasbarra, S. B. Kulathinal, Isha Dewan, and Aulikki Nissinen. Testing dependence between the failure time and failure modes: an application of enlarged filtration. Journal of Statistical Planning and Inference, 136(5):1669–1686, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001746.

Ghosh:2004:AMS

[GKK04] Malay Ghosh, Dal Ho Kim, and Myung Joon Kim. Asymptotic mean squared error of constrained James– Stein estimators. Journal of Statistical Planning and Inference, 126(1):107–118, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002672.

Ghosh:2008:CBE

[GKK08] Malay Ghosh, Myung Joon Kim, and Dal Ho Kim. Constrained Bayes and empirical Bayes estimation under random effects normal ANOVA model with balanced loss function. Journal of Statistical Planning and Inference, 138(7):2017– 2028, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580700359X.

Georgiou:2006:MLS

[GKM06] S. Georgiou, C. Koukouvinos, and P. Mantas. On multilevel supersaturated designs. Journal of Statistical Planning and Inference, 136(8):2805–2819, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004562.

Gnanadesikan:2007:BAC

[GKM07] R. Gnanadesikan, J. R. Kettenring, and Srinivas Maloor. Better alternatives to current methods of scaling and weighting data for cluster analysis. Journal of Statistical Planning and Inference, 137(11):3483–3496, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001115.

Georgiou:2002:NSC

[GKMS02] S. Georgiou, C. Koukouvinos, M. Mitrouli, and Jennifer Seberry. Necessary and sufficient conditions for three and four variable orthogonal designs in order 36. Journal of Statistical Planning and Inference, 106(1–2):329–352, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002215.

Guo:2002:PRM

[GL02a] Jie Q. Guo and Tong Li. Poisson regression models with errors-in-variables: implication and treatment. Journal of Statistical Planning and Inference, 104(2):391–401, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002506.

Gupta:2002:SMR

[GL02b] Shanti S. Gupta and TaChen Liang. Selecting the most reliable Poisson population provided it is better than a control: a nonparametric empirical Bayes approach. Journal of Statistical Planning and Inference, 103(1–2):191–203, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580100221X.

Ganju:2005:RRR

[GL05a] Jitendra Ganju and James M. Lucas. Randomized and random run order experiments. Journal of Statistical Planning and Inference, 133(1):199–210, July 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001119.

Gupta:2005:EBP

[GL05b] Shanti S. Gupta and Jianjun Li. On empirical Bayes procedures for selecting good populations in a positive exponential family. Journal of Statistical Planning and Inference, 129(1–2):3–18, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580400151X.

Genest:2009:ECD

[GL09] Christian Genest and Jean-Marc Lévesque. Estimating correlation from dichotomized normal variables. Journal of Statistical Planning and Inference, 139(11):3785–3794, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900144X.

Glickman:2008:BLO

[Gli08] Mark E. Glickman. Bayesian locally optimal design of knockout tournaments. Journal of Statistical Planning and Inference, 138(7):2117–2127, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003655.

Gouet:2009:LLC

[GLS09] Raúl Gouet, F. Javier López, and Gerardo Sanz. Limit laws for the cumulative number of ties for the maximum in a random sequence. Journal of Statistical Planning and Inference, 139(9):2988–3000, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000305.

Glynn:2000:RCP

[Gly00] David G. Glynn. On the representation of configurations in projective spaces. Journal of Statistical Planning and Inference, 86(2):443–456, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900124X.

Ghosh:2001:TSA

[GM01a] J. K. Ghosh and Rahul Mukerjee. Test statistics arising from quasi likelihood: Bartlett adjustment and higher-order power. Journal of Statistical Planning and Inference, 97(1):45–55, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580000344X.

Gomes:2001:GHE

[GM01b] M. Ivette Gomes and M. João Martins. Generalizations of the Hill estimator — asymptotic versus finite sample behaviour. Journal of Statistical Planning and Inference, 93(1–2):161–180, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002019.

Gu:2001:SCS

[GM01c] Xiaonong Gu and Thomas Mathew. Some characterizations of symmetric two-principal points. Journal of Statistical Planning and Inference, 98(1–2):29–37, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002433.

Gupta:2001:P

[GM01d] Sudhir Gupta and Rahul Mukerjee. Preface. Journal of Statistical Planning and Inference, 96(1):1, June 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003165.

Gupta:2002:PSS

[GM02] Shanti S. Gupta and Klaus J. Miescke. On the performance of subset selection rules under normality. Journal of Statistical Planning and Inference, 103(1–2):101–115, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002142.

Gomes:2004:BRE

[GM04] M. Ivette Gomes and M. João Martins. Bias reduction and explicit semi-parametric estimation of the tail index. Journal of Statistical Planning and Inference, 124(2):361–378, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002052.

Girard:2005:CLT

[GM05] Stéphane Girard and Ludovic Menneteau. Central limit theorems for smoothed extreme value estimates of Poisson point processes boundaries. Journal of Statistical Planning and Inference, 135(2):433–460, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002290.

Gupta:2007:P

[GM07] Sat Gupta and Satya Mishra. Preface. Journal of Statistical Planning and Inference, 137(5):1515, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002242.

Gupta:2008:Pb

[GM08] Sudhir Gupta and Rahul Mukerjee. Preface. Journal of Statistical Planning and Inference, 138(2):307, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002625.

Ghosh:2006:BAZ

[GML06] Sujit K. Ghosh, Pabak Mukhopadhyay, and Jye-Chyi (JC) Lu. Bayesian analysis of zero-inflated regression models. Journal of Statistical Planning and Inference, 136(4):1360–1375, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804004008.

Giron:2003:BAM

[GMM03a] F. J. Girón, M. L. Martiínez, and E. Moreno. Bayesian analysis of matched pairs. Journal of Statistical Planning and Inference, 113(1):49–66, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002993.

Giudici:2003:MPD

[GMM03b] Paolo Giudici, Maura Mezzetti, and Pietro Muliere. Mixtures of products of Dirichlet processes for variable selection in survival analysis. Journal of Statistical Planning and Inference, 111(1–2):101–115, ???? 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002914.

Geraldes:2006:DAW

[GMPZ06] Margarida Geraldes, Vincent Melfi, Connie Page, and Hui Zhang. The doubly adaptive weighted differences design. Journal of Statistical Planning and Inference, 136(6):1923– 1939, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580500193X.

Gendron:2003:MHT

[GN03] Paul Gendron and Balgobin Nandram. Modeling heavy-tailed correlated noise with wavelet packet basis functions. Journal of Statistical Planning and Inference, 112(1–2):99–114, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003269.

Genest:2009:APC

[GN09] Christian Genest and Johanna Neslehová. Analytical proofs of classical inequalities between Spearman’s ρ and Kendall’s τ. Journal of Statistical Planning and Inference, 139(11):3795–3798, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001451.

Gupta:2007:RAO

[GNP07] A. K. Gupta, T. Nguyen, and L. Pardo. Residual analysis and outliers in loglinear models based on phidivergence statistics. Journal of Statistical Planning and Inference, 137(4):1407–1423, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001042.

Godfrey:2001:IRR

[GO01] L. G. Godfrey and C. D. Orme. On improving the robustness and reliability of Rao’s score test. Journal of Statistical Planning and Inference, 97(1):153–176, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003517.

Gajek:2003:PBE

[GO03] L. Gajek and A. Okolewski. Projection bounds on expectations of record statistics from restricted families. Journal of Statistical Planning and Inference, 110(1–2):97–108, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580100283X.

Golyandina:2007:CSM

[GO07] N. Golyandina and E. Osipov. The “Caterpillar”-SSA method for analysis of time series with missing values. Journal of Statistical Planning and Inference, 137(8):2642–2653, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580700016X.

Goldstein:2001:AFC

[Gol01] Michael Goldstein. Avoiding foregone conclusions: geometric and foundational analysis of paradoxes of finite additivity. Journal of Statistical Planning and Inference, 94(1):73–87, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002299.

Goodman:2008:SSU

[Goo08] Leo A. Goodman. To split or not to split the US Ninth Circuit Court of Appeals: a simple statistical argument, counterargument, and critique. Journal of Statistical Planning and Inference, 138(9):2681–2687, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000591.

Goos:2009:DRS

[Goo09] Peter Goos. Discussion of “Response surface design evaluation and comparison”. Journal of Statistical Planning and Inference, 139(2):657–659, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001638. See [ACBM09b].

Gordon:2007:EFG

[Gor07] Alexander Y. Gordon. Explicit formulas for generalized family-wise error rates and unimprovable step-down multiple testing procedures. Journal of Statistical Planning and Inference, 137(11):3497–3512, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001127.

Govindarajulu:2005:RSS

[Gov05] Z. Govindarajulu. Robustness of sample size re-estimation with interim binary data for double-blind clinical trials. Journal of Statistical Planning and Inference, 129(1–2):125–144, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001582. See corri-

genda [Gov09].

Govindarajulu:2009:CZG

[Gov09] Z. Govindarajulu. Corrigenda to Z. Govindarajulu (2005), “Robustness of sample size re-estimation with interim binary data for doubleblind clinical trials”. J. Stat. Planning and Inference, vol. 129, pp. 125–144. Journal of Statistical Planning and Inference, 139(2):677–678, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001985. See [Gov05].

Guerrero:2003:CMT

[GP03] Viíctor M. Guerrero and Daniel Peña. Combining multiple time series predictors: a useful inferential procedure. Journal of Statistical Planning and Inference, 116(1):249– 276, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802001866.

Girard:2004:DPS

[GP04] Philippe Girard and Eric Parent. The deductive phase of statistical analysis via predictive simulations: test, validation and control of a linear model with autocorrelated errors representing a food process. Journal of Statistical Planning and Inference, 124(1):99–120, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001915.

Gauchi:2006:DNR

[GP06] J.-P. Gauchi and A. Pázman. Designs in nonlinear regression by stochastic minimization of functionals of the mean square error matrix. Journal of Statistical Planning and Inference, 136(3):1135–1152, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003453.

Galeano:2007:CCD

[GP07a] Pedro Galeano and Daniel Peña. Covariance changes detection in multivariate time series. Journal of Statistical Planning and Inference, 137(1):194–211, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002673.

Galeano:2007:IMS

[GP07b] Pedro Galeano and Daniel Peña. Improved model selection criteria for SETAR time series models. Journal of Statistical Planning and Inference, 137(9):2802–2814, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000444.

Gupta:2007:PDP

[GP07c] A. K. Gupta and L. Pardo. Phi-divergences and polytomous logistic regression models: an overview. Journal of Statistical Planning and Inference, 137(11):3513–3524, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001139.

Gail:2001:EVA

[GPC01] Mitchell H. Gail, David Pee, and Raymond Carroll. Effects of violations of assumptions on likelihood methods for estimating the penetrance of an autosomal dominant mutation from kin-cohort studies. Journal of Statistical Planning and Inference, 96(1):167–177, June 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003335.

Greenshtein:2009:RTV

[GPL09] Eitan Greenshtein, Junyong Park, and Guy Lebanon. Regularization through variable selection and conditional MLE with application to classification in high dimensions. Journal of Statistical Planning and Inference, 139(2):385–395, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001833.

Galetto:2003:CRC

[GPR03a] Fausto Galetto, Giovanni Pistone, and Maria Piera Rogantin. Confounding revisited with commutative computational algebra. Journal of Statistical Planning and Inference, 117(2):345–363, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003907.

Gutierrez-Pena:2003:RPE

[GPR03b] E. Gutiérrez-Peña and R. Rueda. Reference priors for exponential families. Journal of Statistical Planning and Inference, 110(1–2):35–54, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801002816.

Guida:2009:EIR

[GPV09] Maurizio Guida, Gianpaolo Pulcini, and Mario Vianello. Early inference on reliability of upgraded automotive components by using past data and technical information. Journal of Statistical Planning and Inference, 139(5):1604–1618, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580800236X.

Gutierrez-Pena:2001:BPA

[GPW01] E. Gutiérrez-Peña and S. G. Walker. A Bayesian predictive approach to model selection. Journal of Statistical Planning and Inference, 93(1–2):259–276, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001725.

Gupta:2001:CPL

[GR01] Arjun K. Gupta and Asoka Ramanayake. Change points with linear trend for the exponential distribution. Journal of Statistical Planning and Inference, 93(1–2):181–195, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002056.

Goroncy:2008:LBE

[GR08a] Agnieszka Goroncy and Tomasz Rychlik. Lower bounds on expectations of positive L-statistics from withoutreplacement models. Journal of Statistical Planning and Inference, 138(12):3647–3659, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001213.

Guo:2008:CFD

[GR08b] Wenge Guo and M. Bhaskara Rao. On control of the false discovery rate under no assumption of dependency. Journal of Statistical Planning and Inference, 138(10):3176–3188, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000165.

Goldstein:2009:RBMa

[GR09a] Michael Goldstein and Jonathan Rougier. Reified Bayesian modelling and inference for physical systems. Journal of Statistical Planning and Inference, 139(3):1221–1239, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808003303. See discussion

[O’H09] and response [GR09b].

Goldstein:2009:RBMb

[GR09b] Michael Goldstein and Jonathan Rougier. Reified Bayesian modelling: Issues and opportunities (response to the discussion). Journal of Statistical Planning and Inference, 139(3):1251–1256, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003364.

Grafstrom:2009:NRI

[Gra09] Anton Grafström. Non-rejective implementations of the Sampford sampling design. Journal of Statistical Planning and Inference, 139(6):2111–2114, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800387X.

Gill:2009:CEP

[GRC09] Ryan Gill, Grzegorz A. Rempala, and Michal Czajkowski. Confidence estimation via the parametric bootstrap in logistic joinpoint regression. Journal of Statistical Planning and Inference, 139(9):3132–3141, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900055X.

Greenland:2008:MLC

[Gre08] Sander Greenland. Maximum-likelihood and closed-form estimators of epidemiologic measures under misclassification. Journal of Statistical Planning and Inference, 138(2):528–538, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002595.

Grillenzoni:2005:NPS

[Gri05] Carlo Grillenzoni. Non-parametric smoothing of spatiotemporal point processes. Journal of Statistical Planning and Inference, 128(1):61–78, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003227.

Gupta:2002:WOB

[GRP02] V. K. Gupta, D. V. V. Ramana, and Rajender Parsad. Weighted A-optimal block designs for comparing test treatments with controls with unequal precision. Journal of Statistical Planning and Inference, 106(1–2):159–175, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002100.

Georgiev:2007:CCM

[GRR07] Pando Georgiev, Anca Ralescu, and Dan Ralescu. Crosscumulants measure for independence. Journal of Statistical Planning and Inference, 137(3):1085–1098, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001637.

Gonzalez-Rivera:2001:RST

[GRU01] Gloria González-Rivera and Aman Ullah. Rao’s score test with nonparametric density estimators. Journal of Statistical Planning and Inference, 97(1):85–100, August 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003475.

Greco:2008:RLF

[GRV08] Luca Greco, Walter Racugno, and Laura Ventura. Robust likelihood functions in Bayesian inference. Journal of Statistical Planning and Inference, 138(5):1258–1270, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001899.

Ghement:2008:REE

[GRZ08] Isabella Rodica Ghement, Marcelo Ruiz, and Ruben Zamar. Robust estimation of error scale in nonparametric regression models. Journal of Statistical Planning and Inference, 138(10):3200–3216, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000177.

Giani:2000:MCP

[GS00] G. Giani and K. Straßburger. Multiple comparison procedures for optimally discriminating between good, equivalent, and bad treatments with respect to a control. Journal of Statistical Planning and Inference, 83(2):413–440, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001044.

Ghosh:2002:NBT

[GS02] Jayanta K. Ghosh and Tapas Samanta. Nonsubjective Bayes testing — an overview. Journal of Statistical Planning and Inference, 103(1–2):205–223, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002221.

Gander:2007:SVM

[GS07a] Matthew P. S. Gander and David A. Stephens. Stochastic volatility modelling in continuous time with general marginal distributions: Inference, prediction and model selection. Journal of Statistical Planning and Inference, 137(10):3068–3081, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000754.

Ghosh:2007:EBE

[GS07b] Malay Ghosh and Karabi Sinha. Empirical Bayes estimation in finite population sampling under functional measurement error models. Journal of Statistical Planning and Inference, 137(9):2759–2773, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000377.

Guerriero:2007:NCS

[GS07c] Massimo Guerriero and Marco F. Sandri. A note on the comparison of some randomized response procedures. Journal of Statistical Planning and Inference, 137(7):2184–2190, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002072.

Gupta:2007:UTA

[GS07d] Sat Gupta and Javid Shabbir. On the use of transformed auxiliary variables in estimating population mean by using two auxiliary variables. Journal of Statistical Planning and Inference, 137(5):1606–1611, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002394.

Goldstein:2008:PVA

[GS08] Michael Goldstein and Allan Seheult. Prior viability assessment for Bayesian analysis. Journal of Statistical Planning and Inference, 138(5):1271–1286, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001905.

Giampaoli:2009:LRT

[GS09a] Viviana Giampaoli and Julio M. Singer. Likelihood ratio tests for variance components in linear mixed models. Journal of Statistical Planning and Inference, 139(4):1435–1448, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003455.

Gut:2009:TSC

[GS09b] Allan Gut and Josef Steinebach. Truncated sequential changepoint detection based on renewal counting processes II. Journal of Statistical Planning and Inference, 139(6):1921–1936, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003613.

Guan:2007:APM

[GSC07] Yongtao Guan, Michael Sherman, and James A. Calvin. On asymptotic properties of the mark variogram estimator of a marked point process. Journal of Statistical Planning and Inference, 137(1):148–161, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002958.

Gelper:2009:ROS

[GSCG09] Sarah Gelper, Karen Schettlinger, Christophe Croux, and Ursula Gather. Robust online scale estimation in time series: a model-free approach. Journal of Statistical Planning and Inference, 139(2):335–349, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800178X.

Garcia-Soidan:2004:NKE

[GSFBGM04] Pilar H. García-Soidán, Manuel Febrero-Bande, and Wenceslao González-Manteiga. Nonparametric kernel estimation of an isotropic variogram. Journal of Statistical Planning and Inference, 121(1):65–92, March 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802005074.

Gabler:2000:MSS

[GSM00] Siegfried Gabler, Horst Stenger, and X. Mannheim. Minimax strategies in survey sampling. Journal of Statistical Planning and Inference, 90(2):305–321, October 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001178.

Gannoun:2003:NPC

[GSY03] Ali Gannoun, Jérôme Saracco, and Keming Yu. Nonparametric prediction by conditional median and quantiles. Journal of Statistical Planning and Inference, 117(2):207–223, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003841.

Ghosh:2002:CSD

[GT02] Subir Ghosh and Lance Teschmacher. Comparisons of search designs using search probabilities. Journal of Statistical Planning and Inference, 104(2):439–458, June 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002580.

Gunduz:2006:SAO

[GT06] Necla Gunduz and Ben Torsney. Some advances in optimal designs in contingent valuation studies. Journal of Statistical Planning and Inference, 136(3):1153–1165, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003362.

Gallagher:2008:SSC

[GT08a] Colin Gallagher and Ferebee Tunno. A small sample confidence interval for autoregressive parameters. Journal of Statistical Planning and Inference, 138(12):3858–3868, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000293.

Ghosh:2008:EDF

[GT08b] Kaushik Ghosh and Ram C. Tiwari. Estimating the distribution function using k-tuple ranked set samples. Journal of Statistical Planning and Inference, 138(4):929–949, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001371.

Ghahramani:2009:CEF

[GT09] M. Ghahramani and A. Thavaneswaran. Combining estimating functions for volatility. Journal of Statistical Planning and Inference, 139(4):1449–1461, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003479.

Gilder:2007:CII

[GTT^{+}07] Kye Gilder, Naitee Ting, Lili Tian, Joseph C. Cappelleri, and R. Choudary Hanumara. Confidence intervals on intraclass correlation coefficients in a balanced twofactor random design. Journal of Statistical Planning and Inference, 137(4):1199–1212, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000851.

Goos:2001:ODV

[GTV01] P. Goos, L. Tack, and M. Vandebroek. Optimal designs for variance function estimation using sample variances. Journal of Statistical Planning and Inference, 92(1–2):233–252, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001567.

Guan:2007:STC

[Gua07] Zhong Guan. Semiparametric tests for change-points with epidemic alternatives. Journal of Statistical Planning and Inference, 137(6):1748–1764, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001789.

Guan:2008:MPP

[Gua08] Yongtao Guan. A marked point process perspective in fitting spatial point process models. Journal of Statistical Planning and Inference, 138(7):2143–2153, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003667.

Gupta:2003:SAM

[Gup03] Ramesh C. Gupta. On some association measures in bivariate distributions and their relationships. Journal of Statistical Planning and Inference, 117(1):83–98, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003671.

Guttmann:2002:LPV

[GV02] Anthony J. Guttmann and Markus Vöge. Lattice paths: vicious walkers and friendly walkers. Journal of Statistical Planning and Inference, 101(1–2):107–131, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001586.

Gurevich:2005:CPP

[GV05] Gregory Gurevich and Albert Vexler. Change point problems in the model of logistic regression. Journal of Statistical Planning and Inference, 131(2):313–331, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000965.

Hirano:2003:MP

[HA03] Katuomi Hirano and Sigeo Aki. On k-match problems. Journal of Statistical Planning and Inference, 109(1–2):67–79, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002999.

Huque:2008:FFS

[HA08] Mohammad F. Huque and Mohamed Alosh. A flexible fixedsequence testing method for hierarchically ordered correlated multiple endpoints in clinical trials. Journal of Statistical Planning and Inference, 138(2):321–335, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002431.

Hagiwara:2006:CMS

[Hag06] Katsuyuki Hagiwara. A consistent model selection for orthogonal regression under component-wise shrinkage. Journal of Statistical Planning and Inference, 136(3):1181–1195, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003635.

Haiman:2000:BRM

[Hai00] George Haiman. Block records and maxima of the increments of the Wiener process. Journal of Statistical Planning and Inference, 85(1–2):137–144, April 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000762.

Haines:2003:ASB

[Hai03] Linda M. Haines. An approach to simple bargaining games and related problems. Journal of Statistical Planning and Inference, 116(2):353–366, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003555.

Haiman:2007:EDO

[Hai07] George Haiman. Estimating the distribution of onedimensional discrete scan statistics viewed as extremes of 1dependent stationary sequences. Journal of Statistical Planning and Inference, 137(3):821–828, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001431.

Hall:2007:EWA

[Hal07] W. J. Hall. Efficiency of weighted averages. Journal of Statistical Planning and Inference, 137(11):3548–3556, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001164.

Hanagal:2008:FRM

[Han08a] David D. Hanagal. Frailty regression models in mixture distributions. Journal of Statistical Planning and Inference, 138(8):2462–2468, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003898.

Hansen:2008:RRR

[Han08b] Peter Reinhard Hansen. Reduced-rank regression: a useful determinant identity. Journal of Statistical Planning and Inference, 138(9):2688–2697, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000608.

Harmanec:2002:GMD

[Har02] David Harmanec. Generalizing Markov decision processes to imprecise probabilities. Journal of Statistical Planning and Inference, 105(1):199–213, June 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002105.

Harman:2008:ETS

[Har08] Radoslav Harman. Equivalence theorem for Schur optimality of experimental designs. Journal of Statistical Planning and Inference, 138(4):1201–1209, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002285.

Harrar:2009:ATM

[Har09] Solomon W. Harrar. Asymptotics for tests on mean profiles, additional information and dimensionality under non-normality. Journal of Statistical Planning and Inference, 139(8):2685–2705, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004552.

Hashorva:2009:CLS

[Has09] Enkelejd Hashorva. Conditional limits of W_{p} scale mixture distributions. Journal of Statistical Planning and Inference, 139(10):3501–3511, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000950.

Hawkins:2001:PAN

[Haw01] D. L. Hawkins. A proof of the asymptotic normality of the inverse sampling estimator in finite population sampling. Journal of Statistical Planning and Inference, 92(1–2):175–180, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580000118X.

Hayakawa:2001:RSH

[Hay01] Takesi Hayakawa. Rao’s statistic for homogeneity of multiple parameter. Journal of Statistical Planning and Inference, 97(1):101–111, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800003487.

Hayter:2006:RIM

[Hay06] A. J. Hayter. Recursive integration methodologies with statistical applications. Journal of Statistical Planning and Inference, 136(7):2284–2296, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002223.

Hayter:2007:CMC

[Hay07] A. J. Hayter. A combination multiple comparisons and subset selection procedure to identify treatments that are strictly inferior to the best. Journal of Statistical Planning and Inference, 137(7):2115–2126, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002059.

Harris:2000:PEA

[HB00] Ian R. Harris and Brent D. Burch. Pivotal estimation with applications for the intraclass correlation coefficient in the balanced one-way random effects model. Journal of Statistical Planning and Inference, 83(1):257–276, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000956.

Huang:2004:DEM

[HB04] M. L. Huang and P. H. Brill. A distribution estimation method based on level crossings. Journal of Statistical Planning and Inference, 124(1):45–62, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001976.

Hartung:2003:GCW

[HBK03] Joachim Hartung, Annette Böckenhoff, and Guido Knapp. Generalized Cochran–Wald statistics in combining of experiments. Journal of Statistical Planning and Inference, 113(1):215–237, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201002036.

Han:2003:ODE

[HC03] Cong Han and Kathryn Chaloner. D- and c-optimal designs for exponential regression models used in viral dynamics and other applications. Journal of Statistical Planning and Inference, 115(2):585–601, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001751.

Huang:2006:CMI

[HC06a] R. Huang and K. C. Carriere. Comparison of methods for incomplete repeated measures data analysis in small samples. Journal of Statistical Planning and Inference, 136(1):235–247, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002526.

Huang:2006:SEB

[HC06b] Wen-Tao Huang and Yi-Ping Chang. Some empirical Bayes rules for selecting the best population with multiple criteria. Journal of Statistical Planning and Inference, 136(7):2129– 2143, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002144.

Hsiao:2007:BOS

[HC07a] Chin-Fu Hsiao and Murray K. Clayton. Bayes optimal sequential trial designs. Journal of Statistical Planning and Inference, 137(4):1129–1137, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002758.

Huang:2007:SCM

[HC07b] Wen-Jang Huang and Shu-Huey Chang. On some characterizations of the mixture of gamma distributions. Journal of Statistical Planning and Inference, 137(9):2964–2974, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000419.

Hu:2008:DPO

[HC08] Taizhong Hu and Huaihou Chen. Dependence properties of order statistics. Journal of Statistical Planning and Inference, 138(7):2214–2222, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003801.

Hofmann:2005:AAP

[HCBK05] Glenn Hofmann, Erhard Cramer, N. Balakrishnan, and Gerd Kunert. An asymptotic approach to progressive censoring.

Journal of Statistical Planning and Inference, 130(1–2):207– 227, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002629.

He:2004:LDA

[HCS04] Xuming He, Hengjian Cui, and Douglas G. Simpson. Longitudinal data analysis using t-type regression. Journal of Statistical Planning and Inference, 122(1–2):253–269, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002453.

Hsu:2006:SCI

[HCW06] Jason C. Hsu, Jane Y. Chang, and Tao Wang. Simultaneous confidence intervals for differential gene expressions. Journal of Statistical Planning and Inference, 136(7):2182– 2196, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580500217X.

Hall:2009:MSL

[HdGT09] Andreia Hall and Maria da Graça Temido. On the max-semistable limit of maxima of stationary sequences with missing values. Journal of Statistical Planning and Inference, 139(3):875–890, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002577.

Honari:2009:UBN

[HDM09] Bahman Honari, John Donovan, and Eamonn Murphy. Using Bayesian Networks in reliability evaluation for an (r, s)- out-of-(m, n):F distributed communication system. Journal of Statistical Planning and Inference, 139(5):1756–1765, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002504.

Heinrich:2000:VRM

[Hei00] Stefan Heinrich. Variance reduction by means of deterministic computation: collision estimate. Journal of Statistical Planning and Inference, 85(1–2):159–169, April 1, 2000.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000786.

Heuchenne:2008:SUC

[Heu08] Cédric Heuchenne. Strong uniform consistency results of the weighted average of conditional artificial data points. Journal of Statistical Planning and Inference, 138(5):1496–1515, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002935.

Heiler:2000:DDD

[HF00] Siegfried Heiler and Yuanhua Feng. Data-driven decomposition of seasonal time series. Journal of Statistical Planning and Inference, 91(2):351–363, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001877.

Hadjicostas:2001:ADP

[HH01] Petros Hadjicostas and George C. Hadjinicola. The asymptotic distribution of the proportion of correct classifications for a holdout sample in logistic regression. Journal of Statistical Planning and Inference, 92(1–2):193–211, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001427.

Hamada:2002:NTC

[HH02] Noboru Hamada and Tor Helleseth. The nonexistence of ternary [97,6,63] codes. Journal of Statistical Planning and Inference, 106(1–2):485–507, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002306.

Horvath:2005:TCU

[HH05] Lajos Horváth and Marie Husková. Testing for changes using permutations of U-statistics. Journal of Statistical Planning and Inference, 128(2):351–371, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000096.

Hwang:2007:MEC

[HH07] Wen-Han Hwang and Yih-Huei Huang. Measurement errors in continuous-time capture-recapture models. Journal of Statistical Planning and Inference, 137(6):1888–1899, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001273.

Hu:2008:MCC

[HH08] Xiaomi Hu and Jürgen Hansohm. Merge and chop in the computation for isotonic regressions. Journal of Statistical Planning and Inference, 138(10):3099–3106, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004260.

He:2009:DSB

[HH09] Hua He and Li-Shan Huang. Double-smoothing for bias reduction in local linear regression. Journal of Statistical Planning and Inference, 139(3):1056–1072, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002905.

Horvath:2004:MCL

[HHKS04] Lajos Horváth, Marie Husková, Piotr Kokoszka, and Josef Steinebach. Monitoring changes in linear models. Journal of Statistical Planning and Inference, 126(1):225–251, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002647.

Horvath:2008:CBR

[HHZ08] Lajos Horváth, Zsuzsanna Horváth, and Wang Zhou. Confidence bands for ROC curves. Journal of Statistical Planning and Inference, 138(6):1894–1904, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003291.

Hill:2009:FCL

[Hil09] Jonathan B. Hill. On functional central limit theorems for dependent, heterogeneous arrays with applications to tail index and tail dependence estimation. Journal of Statistical Planning and Inference, 139(6):2091–2110, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003832.

Hirotsu:2002:ODI

[Hir02] Chihiro Hirotsu. On an optimal design for an isotonic inference. Journal of Statistical Planning and Inference, 106 (1–2):205–213, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802002136.

Hishida:2002:CBI

[HJ02] Takaaki Hishida and Masakazu Jimbo. Constructions of balanced incomplete block designs with nested rows and columns. Journal of Statistical Planning and Inference, 106 (1–2):47–56, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580200201X.

Huang:2009:RPT

[HJR09] Alan Huang, Rungao Jin, and John Robinson. Robust permutation tests for two samples. Journal of Statistical Planning and Inference, 139(8):2631–2642, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004539.

Holzmann:2001:ODO

[HK01] W. H. Holzmann and H. Kharaghani. On the orthogonal designs of order 40. Journal of Statistical Planning and Inference, 96(2):415–429, July 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002135.

Hughes:2003:MSU

[HK03] Anthony W. Hughes and Maxwell L. King. Model selection using AIC in the presence of one-sided information. Journal of Statistical Planning and Inference, 115(2):397–411, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001593.

Hamers:2004:HWC

[HK04] Michael Hamers and Michael Kohler. How well can a regression function be estimated if the distribution of the (random) design is concentrated on a finite set? Journal of Statistical Planning and Inference, 123(2):377–394, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001551.

Hartung:2005:CIA

[HK05] Joachim Hartung and Guido Knapp. On confidence intervals for the among-group variance in the one-way random effects model with unequal error variances. Journal of Statistical Planning and Inference, 127(1–2):157–177, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803003136.

Hu:2002:ETO

[HKMS02] Xiaomi Hu, Subhash C. Kochar, Hari Mukerjee, and Francisco J. Samaniego. Estimation of two ordered mean residual life functions. Journal of Statistical Planning and Inference, 107(1–2):321–341, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002616.

Hayakawa:2008:SIH

[HKP08] T. Hayakawa, S. Kageyama, and M. L. Puri. Special issue in honour of Junjiro Ogawa (1915–2000): Design of experiments, multivariate analysis and statistical inferences. Journal of Statistical Planning and Inference, 138(11):3293, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000748.

Huskova:2008:DCA

[HKPS08] Marie Husková, Claudia Kirch, Zuzana Prásková, and Josef Steinebach. On the detection of changes in autoregressive time series, II. Resampling procedures. Journal of Statistical Planning and Inference, 138(6):1697–1721, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003114.

Holzmann:2005:ODO

[HKSTR05] W. H. Holzmann, H. Kharaghani, Jennifer Seberry, and B. Tayfeh-Rezaie. On orthogonal designs in order 48. Journal of Statistical Planning and Inference, 128(1):311–325, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003677.

Hallin:2000:OID

[HKW00] Marc Hallin, Christophe Koell, and Bas J. M. Werker. Optimal inference for discretely observed semiparametric Ornstein–Uhlenbeck processes. Journal of Statistical Planning and Inference, 91(2):323–340, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001853.

Horvath:2008:DAE

[HKZ08] Lajos Horváth, Piotr Kokoszka, and Ricardas Zitikis. Distributional analysis of empirical volatility in GARCH processes. Journal of Statistical Planning and Inference, 138 (11):3578–3589, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808001006.

Hu:2001:IWS

[HL01] Chin-Yuan Hu and Gwo Dong Lin. An inequality for the weighted sums of pairwise i.i.d. generalized Rayleigh random variables. Journal of Statistical Planning and Inference, 92(1–2):1–5, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580000149X.

Horvath:2004:EAM

[HL04] Lajos Horváth and Friedrich Liese. L_{p}-estimators in ARCH models. Journal of Statistical Planning and Inference, 119(2):277–309, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004883.

Hannig:2006:PSE

[HL06] Jan Hannig and Thomas C. M. Lee. On Poisson signal estimation under Kullback–Leibler discrepancy and squared risk. Journal of Statistical Planning and Inference, 136(3):882–908, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003490.

Hassairi:2009:MSE

[HL09a] A. Hassairi and M. Louati. Multivariate stable exponential families and Tweedie scale. Journal of Statistical Planning and Inference, 139(2):143–158, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001559.

Husler:2009:TAI

[HL09b] Jürg Hüsler and Deyuan Li. Testing asymptotic independence in bivariate extremes. Journal of Statistical Planning and Inference, 139(3):990–998, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002796.

Hlavka:2003:APR

[Hlá03] Zdenek Hlávka. Asymptotic properties of robust three-stage procedure based on bootstrap for M-estimator. Journal of Statistical Planning and Inference, 115(2):637–656, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580200160X.

Hayter:2009:RMC

[HLAK09] A. J. Hayter, W. Liu, and P. Ah-Kine. A ray method of confidence band construction for multiple linear regression models.

Journal of Statistical Planning and Inference, 139(2):329–334, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001870.

Huang:2009:NMA

[HLH09] Su-Yun Huang, Mei-Hsien Lee, and Chuhsing Kate Hsiao. Nonlinear measures of association with kernel canonical correlation analysis and applications. Journal of Statistical Planning and Inference, 139(7):2162–2174, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003923.

Hall:2008:WBA

[HLM08] Peter Hall, Xiaoyan Leng, and Hans-Georg Müller. Weightedbootstrap alignment of explanatory variables. Journal of Statistical Planning and Inference, 138(6):1817–1827, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003205.

Huang:2006:SVS

[HLP06] Deng-Yuan Huang, Ren-Fen Lee, and S. Panchapakesan. On some variable selection procedures based on data for regression models. Journal of Statistical Planning and Inference, 136(7):2020–2034, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002089.

Hayter:2007:ECC

[HLW07] A. J. Hayter, W. Liu, and H. P. Wynn. Easy-to-construct confidence bands for comparing two simple linear regression lines. Journal of Statistical Planning and Inference, 137(4):1213–1225, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000759.

Hassairi:2005:BRD

[HLZ05] A. Hassairi, S. Lajmi, and R. Zine. Beta-Riesz distributions on symmetric cones. Journal of Statistical Planning and Inference, 133(2):387–404, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001223.

Hassairi:2007:RIG

[HLZ07] A. Hassairi, S. Lajmi, and R. Zine. Riesz inverse Gaussian distributions. Journal of Statistical Planning and Inference, 137(6):2024–2033, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001728.

Hallin:2001:SHE

[HM01] Marc Hallin and Ivan Mizera. Sample heterogeneity and M-estimation. Journal of Statistical Planning and Inference, 93(1–2):139–160, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001749.

Hartigan:2002:IP

[HM02a] J. A. Hartigan and T. B. Murphy. Inferred probabilities. Journal of Statistical Planning and Inference, 105(1):23–34, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002026.

Hering:2002:ISB

[HM02b] Franz Hering and Stanislaw Mejza. An incomplete splitblock design generated by GDPBIBD(2)s. Journal of Statistical Planning and Inference, 106(1–2):125–134, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002082.

Hayter:2000:CAO

[HML00] A. J. Hayter, Tetsuhisa Miwa, and Wei Liu. Combining the advantages of one-sided and two-sided procedures for comparing several treatments with a control. Journal of Statistical Planning and Inference, 86(1):81–99, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001615.

Hu:2004:SCP

[HMS04] Taizhong Hu, Alfred Müller, and Marco Scarsini. Some counterexamples in positive dependence. Journal of Statistical Planning and Inference, 124(1):153–158, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002015.

He:2004:MCA

[HMW04] Guozhong He, Hans-Georg Müller, and Jane-Ling Wang. Methods of canonical analysis for functional data. Journal of Statistical Planning and Inference, 122(1–2):141–159, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002428.

Huskova:2004:CPA

[HN04] Marie Husková and Georg Neuhaus. Change point analysis for censored data. Journal of Statistical Planning and Inference, 126(1):207–223, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002684.

He:2009:DCO

[HN09] Qinying He and H. N. Nagaraja. Distribution of concomitants of order statistics and their order statistics. Journal of Statistical Planning and Inference, 139(8):2643–2655, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004503.

Hatjispyros:2008:BPD

[HNW08] S. J. Hatjispyros, Theodoros Nicoleris, and Stephen G. Walker. Bivariate prior distributions via branching exchangeable sequences. Journal of Statistical Planning and Inference, 138(6):1799–1816, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003229.

Henriques:2003:ETD

[HO03] Carla Henriques and Paulo Eduardo Oliveira. Estimation of a two-dimensional distribution function under association. Journal of Statistical Planning and Inference, 113(1):137–150, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002968.

Hodoshima:2004:TTT

[Hod04] Jiro Hodoshima. Tail-thickness in terms of COV(X_{j}^{2}, X_{p}^{2}) in the class of elliptical distributions. Journal of Statistical Planning and Inference, 121(2):285–289, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001186.

Honda:2004:QRV

[Hon04] Toshio Honda. Quantile regression in varying coefficient models. Journal of Statistical Planning and Inference, 121(1): 113–125, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803001101.

Horvath:2000:AHE

[Hor00] Lajos Horváth. Approximations for hybrids of empirical and partial sums processes. Journal of Statistical Planning and Inference, 88(1):1–18, July 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002074.

Hardwick:2006:NAD

[HOS06a] Janis Hardwick, Robert Oehmke, and Quentin F. Stout. New adaptive designs for delayed response models. Journal of Statistical Planning and Inference, 136(6):1940–1955, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001941.

Hosking:2006:CDT

[Hos06b] J. R. M. Hosking. On the characterization of distributions by their L-moments. Journal of Statistical Planning and Inference, 136(1):193–198, January 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002514.

Hosking:2007:DME

[Hos07a] J. R. M. Hosking. Distributions with maximum entropy subject to constraints on their L-moments or expected order statistics. Journal of Statistical Planning and Inference, 137(9):2870–2891, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000341.

Hosking:2007:STP

[Hos07b] J. R. M. Hosking. Some theory and practical uses of trimmed L-moments. Journal of Statistical Planning and Inference, 137(9):3024–3039, September 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000389.

Hossjer:2008:CDM

[Hös08] Ola Hössjer. On the coefficient of determination for mixed regression models. Journal of Statistical Planning and Inference, 138(10):3022–3038, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004107.

Happacher:2000:RPM

[HP00a] Max Happacher and Friedrich Pukelsheim. Rounding probabilities: maximum probability and minimum complexity multipliers. Journal of Statistical Planning and Inference, 85(1–2):145–158, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000774.

He:2000:SAR

[HP00b] Xuming He and Stephen Portnoy. Some asymptotic results on bivariate quantile splines. Journal of Statistical Planning and Inference, 91(2):341–349, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001865.

Huffer:2002:LDT

[HP02] Fred W. Huffer and Cheolyong Park. The limiting distribution of a test for multivariate structure. Journal of Statistical Planning and Inference, 105(2):417–431, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002762.

Hallin:2006:ALS

[HP06a] Marc Hallin and Davy Paindaveine. Asymptotic linearity of serial and nonserial multivariate signed rank statistics. Journal of Statistical Planning and Inference, 136(1):1–32, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002460.

Huerta:2006:SPM

[HP06b] Gabriel Huerta and Raquel Prado. Structured priors for multivariate time series. Journal of Statistical Planning and Inference, 136(11):3802–3821, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000911.

Hedayat:2007:TCO

[HP07] A. S. Hedayat and H. Pesotan. Tools for constructing optimal two-level factorial designs for a linear model containing main effects and one two-factor interaction. Journal of Statistical Planning and Inference, 137(4):1452–1463, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001066.

Huskova:2007:DCA

[HPS07] Marie Husková, Zuzana Prásková, and Josef Steinebach. On the detection of changes in autoregressive time series I. Asymptotics. Journal of Statistical Planning and Inference, 137(4):1243–1259, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000826.

Hobza:2008:RME

[HPV08] T. Hobza, L. Pardo, and I. Vajda. Robust median estimator in logistic regression. Journal of Statistical Planning and Inference, 138(12):3822–3840, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001407.

Hall:2002:MMM

[HPY02] Peter Hall, Liang Peng, and Qiwei Yao. Moving-maximum models for extrema of time series. Journal of Statistical Planning and Inference, 103(1–2):51–63, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001975.

Hu:2004:NMS

[HR04] Taizhong Hu and Lingyan Ruan. A note on multivariate stochastic comparisons of Bernoulli random variables. Journal of Statistical Planning and Inference, 126(1):281–288, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580300260X.

Hu:2006:ABR

[HRZ06] Feifang Hu, William F. Rosenberger, and Li-Xin Zhang. Asymptotically best response-adaptive randomization procedures. Journal of Statistical Planning and Inference, 136 (6):1911–1922, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805001928.

Horvath:2000:TCM

[HS00a] Lajos Horváth and Josef Steinebach. Testing for changes in the mean or variance of a stochastic process under weak invariance. Journal of Statistical Planning and Inference, 91(2):365–376, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001889.

Hurd:2000:BRD

[HS00b] Spencer P. Hurd and Dinesh G. Sarvate. On c-Bhaskar Rao designs. Journal of Statistical Planning and Inference, 90(1):161–175, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001129.

Huskova:2000:LTC

[HS00c] Marie Husková and Josef Steinebach. Limit theorems for a class of tests of gradual changes. Journal of Statistical Planning and Inference, 89(1–2):57–77, August 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000094X.

Hardwick:2002:OFS

[HS02] Janis Hardwick and Quentin F. Stout. Optimal few-stage designs. Journal of Statistical Planning and Inference, 104 (1):121–145, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801002427.

Hall:2005:TAP

[HS05] Peter Hall and Michael Stewart. Theoretical analysis of power in a two-component normal mixture model. Journal of Statistical Planning and Inference, 134(1):158–179, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001879.

Hardwick:2007:RAD

[HS07a] Janis P. Hardwick and Quentin F. Stout. Response adaptive designs that incorporate switching costs and constraints. Journal of Statistical Planning and Inference, 137 (8):2654–2665, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000171.

Huang:2007:MML

[HS07b] J. S. Huang and P. S. Shen. More maximum likelihood oddities. Journal of Statistical Planning and Inference, 137(7):2151–2155, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002576.

Hu:2008:GLM

[HS08a] Bo Hu and Jun Shao. Generalized linear model selection using R^{2}. Journal of Statistical Planning and Inference, 138(12):3705–3712, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001250.

Hu:2008:ARA

[HS08b] Jiaqiao Hu and Zheng Su. Adaptive resampling algorithms for estimating bootstrap distributions. Journal of Statistical Planning and Inference, 138(6):1763–1777, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003187.

Huang:2009:NTN

[HS09] Li-Shan Huang and Haiyan Su. Nonparametric F -tests for nested global and local polynomial models. Journal of Statistical Planning and Inference, 139(4):1372–1380, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580800339X.

Hedayat:2008:SSP

[HSS08] A. S. Hedayat, Guoqin Su, and W. Elane Streets. Statistical scoring procedures applicable to laboratory performance evaluation. Journal of Statistical Planning and Inference, 138(11):3336–3349, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000815.

Huber:2009:ICT

[HSV09] Catherine Huber, Valentin Solev, and Filia Vonta. Interval censored and truncated data: Rate of convergence of NPMLE of the density. Journal of Statistical Planning and Inference, 139(5):1734–1749, May 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002486.

Hirukawa:2006:LTN

[HT06] Junichi Hirukawa and Masanobu Taniguchi. LAN theorem for non-Gaussian locally stationary processes and its applications. Journal of Statistical Planning and Inference, 136(3):640–688, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003568.

Hsiao:2008:EDP

[HT08] Cheng Hsiao and A. K. Tahmiscioglu. Estimation of dynamic panel data models with both individual and time-specific effects. Journal of Statistical Planning and Inference, 138(9):2698–2721, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800061X.

Hara:2009:BAE

[HT09a] Hisayuki Hara and Akimichi Takemura. Bayes admissible estimation of the means in Poisson decomposable graphical models. Journal of Statistical Planning and Inference, 139(4):1297–1319, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003273.

Hsu:2009:SES

[HT09b] Nan-Jung Hsu and Henghsiu Tsai. Semiparametric estimation for seasonal long-memory time series using generalized exponential models. Journal of Statistical Planning and Inference, 139(6):1992–2009, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003698.

Huang:2002:NEP

[Hua02] Jian Huang. A note on estimating a partly linear model under monotonicity constraints. Journal of Statistical Planning and Inference, 107(1–2):343–351, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002628.

Huggins:2008:SPN

[Hug08] Richard Huggins. Some properties of a nonparametric estimator of the size of an open population. Journal of Statistical Planning and Inference, 138(3):618–623, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700047X.

Hui:2009:MDP

[Hui09] Jiang Hui. Moderate deviations for parameter estimation in some time inhomogeneous diffusions. Journal of Statistical Planning and Inference, 139(10):3665–3674, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001190.

Hurd:2005:PCS

[Hur05] H. Hurd. Periodically correlated sequences of less than full rank. Journal of Statistical Planning and Inference, 129 (1–2):279–303, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804001958.

Huzurbazar:2005:FMB

[Huz05] Aparna V. Huzurbazar. Flowgraph models: a Bayesian case study in construction engineering. Journal of Statistical Planning and Inference, 129(1–2):181–193, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001612.

Hinich:2005:NB

[HW05] Melvin J. Hinich and Murray Wolinsky. Normalizing bispectra. Journal of Statistical Planning and Inference, 130 (1–2):405–411, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804002745.

Huang:2008:BED

[HW08] Yangxin Huang and Hulin Wu. Bayesian experimental design for long-term longitudinal HIV dynamic studies. Journal of Statistical Planning and Inference, 138(1):105–113, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002029.

Hu:2009:OAA

[HW09] Taizhong Hu and Yashi Wang. Optimal allocation of active redundancies in r-out-of-n systems. Journal of Statistical Planning and Inference, 139(10):3733–3737, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000962.

Hwang:2004:LPS

[Hwa04] Ruey-Ching Hwang. Local polynomial M-smoothers in nonparametric regression. Journal of Statistical Planning and Inference, 126(1):55–72, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002593.

He:2003:ERP

[HY03a] Shuyuan He and Grace L. Yang. Estimation of regression parameters with left truncated data. Journal of Statistical Planning and Inference, 117(1):99–122, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003609.

Hernandez:2003:CLG

[HY03b] Sonia Hernández and Viíctor J. Yohai. Combining locally and globally robust estimates for regression. Journal of Statistical Planning and Inference, 113(2):633–661, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003147.

Hedayat:2008:CDB

[HY08] A. S. Hedayat and Zhiwu Yan. Crossover designs based on type I orthogonal arrays for a self and simple mixed carryover effects model with correlated errors. Journal of Statistical Planning and Inference, 138(7):2201–2213, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003680.

Huggins:2003:PSE

[HYCY03] Richard Huggins, Hsin-Chou Yang, Anne Chao, and Paul S. F. Yip. Population size estimation using local sample coverage for open populations. Journal of Statistical Planning and Inference, 113(2):699–714, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802000939.

Hedayat:2002:ODE

[HYP02] A. S. Hedayat, Bo Yan, and J. M. John M. Pezzuto. Optimum designs for estimating ED_{p} based on raw optical density data. Journal of Statistical Planning and Inference, 104(1):161–174, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001203.

Hyrien:2007:PLE

[Hyr07] Ollivier Hyrien. Pseudo-likelihood estimation for discretely observed multitype Bellman-Harris branching processes. Journal of Statistical Planning and Inference, 137 (4):1375–1388, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806000723.

Hu:2006:SCS

[HZ06] Taizhong Hu and Weiwei Zhuang. Stochastic comparisons of m-spacings. Journal of Statistical Planning and Inference, 136(1):33–42, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002976.

Hedayat:2004:OED

[HZN04] A. S. Hedayat, J. Zhong, and L. Nie. Optimal and efficient designs for 2-parameter nonlinear models. Journal of Statistical Planning and Inference, 124(1):205–217, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001964.

Ivanova:2006:RAD

[IBL06] Anastasia Ivanova, Atanu Biswas, and Anna Lurie. Responseadaptive designs for continuous outcomes. Journal of Statistical Planning and Inference, 136(6):1845–1852, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001898.

Ivanova:2007:CCD

[IFC07] Anastasia Ivanova, Nancy Flournoy, and Yeonseung Chung. Cumulative cohort design for dose-finding. Journal of Statistical Planning and Inference, 137(7):2316–2327, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002461.

Iyer:2008:AMC

[IJK08] Srikanth K. Iyer, S. Rao Jammalamadaka, and Debasis Kundu. Analysis of middle-censored data with exponential lifetime distributions. Journal of Statistical Planning and Inference, 138(11):3550–3560, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000979.

Ishwaran:2009:AB

[IJZ09] Hemant Ishwaran, Lancelot F. James, and Mahmoud Zarepour. An alternative to the m out of n bootstrap. Journal of Statistical Planning and Inference, 139(3):788–801, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002097.

Inusah:2006:DAL

[IK06] Seidu Inusah and Tomasz J. Kozubowski. A discrete analogue of the Laplace distribution. Journal of Statistical Planning and Inference, 136(3):1090–1102, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003519.

Inglot:2004:CMD

[IL04] Tadeusz Inglot and Teresa Ledwina. On consistent minimax distinguishability and intermediate efficiency of Cramérvon Mises test. Journal of Statistical Planning and Inference, 124(2):453–474, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002039.

Inglot:2006:IES

[IL06] Tadeusz Inglot and Teresa Ledwina. Intermediate efficiency of some max-type statistics. Journal of Statistical Planning and Inference, 136(9):2918–2935, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004744.

Ivanov:2008:SAL

[IL08] A. V. Ivanov and N. N. Leonenko. Semiparametric analysis of long-range dependence in nonlinear regression. Journal of Statistical Planning and Inference, 138(6):1733–1753, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003163.

Iliopoulos:2003:EPF

[Ili03] George Iliopoulos. Estimation of parametric functions in Downton’s bivariate exponential distribution. Journal of Statistical Planning and Inference, 117(2):169–184, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003890.

Inoue:2003:IWL

[Ino03] Kiyoshi Inoue. Iterative weighted least-squares estimates in a heteroscedastic linear regression model. Journal of Statistical Planning and Inference, 110(1–2):133–146, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002853.

Ioannidis:2007:SBV

[Ioa07] E. E. Ioannidis. Spectra of bivariate VAR(p) models. Journal of Statistical Planning and Inference, 137(2):554–566, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000413.

Iglesias:2009:NBM

[IOQ09] Pilar L. Iglesias, Yasna Orellana, and Fernando A. Quintana. Nonparametric Bayesian modelling using skewed Dirichlet processes. Journal of Statistical Planning and Inference, 139(3):1203–1214, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003029.

Isotalo:2009:LSC

[IP09] Jarkko Isotalo and Simo Puntanen. Linear sufficiency and completeness in the context of estimating the parametric function in the general Gauss–Markov model. Journal of Statistical Planning and Inference, 139(3):722–733, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002620.

Iglesias-Perez:2008:NEC

[IPdUÁ08] Maria Carmen Iglesias-Pérez and Jacobo de Uña-Álvarez. Nonparametric estimation of the conditional distribution function in a semiparametric censorship model. Journal of Statistical Planning and Inference, 138(10):3044–3058, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807004041.

Isotalo:2008:BCM

[IPS08] Jarkko Isotalo, Simo Puntanen, and George P. H. Styan. The BLUE’s covariance matrix revisited: a review. Journal of Statistical Planning and Inference, 138(9):2722–2737, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000621.

Ionin:2001:CTC

[IS01] Yury J. Ionin and Mohan S. Shrikhande. On classification of two class partially balanced designs. Journal of Statistical Planning and Inference, 95(1–2):209–228, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002895.

Ionin:2002:FID

[IS02a] Yury J. Ionin and Mohan S. Shrikhande. Fisher’s inequality for designs on regular graphs. Journal of Statistical Planning and Inference, 100(2):185–190, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100132X.

Iwashita:2002:ALI

[IS02b] Toshiya Iwashita and Takashi Seo. An application of Lévy’s inversion formula for higher order asymptotic expansion. Journal of Statistical Planning and Inference, 104(2):403–416, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002543.

Iqbal:2010:EGP

[IT10] Ijaz Iqbal and M. H. Tahir. Erratum “Generalized polygonal designs with block size 3 and λ = 1” by I. Iqbal et al. [J. Statist. Plann. Inference 139 (2009) 3200–3219]. Journal of Statistical Planning and Inference, 140(1):334, January 1, 2010. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809002195. See [ITA^{+}09].

Iqbal:2009:GPD

[ITA^{+}09] Ijaz Iqbal, M. H. Tahir, Munir Akhtar, S. S. A. Ghazali, Javid Shabbir, and N. S. Bukhari. Generalized polygonal designs with block size 3 and λ = 1. Journal of Statistical Planning and Inference, 139(9):3200–3219, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000706. See erratum [IT10].

Isogai:2001:ISP

[IU01] Eiichi Isogai and Chikara Uno. Improved sequential procedures for estimating the percentile of a normal distribution. Journal of Statistical Planning and Inference, 92(1– 2):163–174, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001506.

Ivanova:2006:UDI

[Iva06] Anastasia Ivanova. Urn designs with immigration: Useful connection with continuous time stochastic processes. Journal of Statistical Planning and Inference, 136(6):1836–1844, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001886.

Iyer:2003:HOC

[IVM03] Hari K. Iyer, D. F. Dominic F. Vecchia, and Paul W. Mielke. Higher order cumulants and Tchebyshev–Markov bounds for P -values in distribution-free matched-pairs tests. Journal of Statistical Planning and Inference, 116(1):131–147, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001787.

Jacroux:2002:DCM

[Jac02] Mike Jacroux. On the determination and construction of A- and MV-optimal block designs for comparing a set of test treatments to a set of standard treatments. Journal of Statistical Planning and Inference, 106(1–2):191–204, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002124.

Jacroux:2003:SMO

[Jac03] Mike Jacroux. Some MV-optimal group divisible type block designs for comparing a set of test treatments with a set of standard treatments. Journal of Statistical Planning and Inference, 113(2):597–615, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003159.

Jacroux:2004:MMA

[Jac04] Mike Jacroux. A modified minimum aberration criterion for selecting regular 2^{m−k} fractional factorial designs. Journal of Statistical Planning and Inference, 126(1):325–336, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002271.

Jacroux:2006:OSB

[Jac06] Mike Jacroux. Optimal sequencing of blocked 2m-k fractional factorial designs. Journal of Statistical Planning and Inference, 136(4):1473–1492, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003775.

Jacroux:2009:BTL

[Jac09] Mike Jacroux. Blocking in two-level non-regular fractional factorial designs. Journal of Statistical Planning and Inference, 139(3):1215–1220, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003108.

Janssen:2001:TNS

[Jan01] Arnold Janssen. Testing nonparametric statistical functionals with applications to rank tests [J. Statist. Plann. Inference 81 (1999) 71–93]. Journal of Statistical Planning and Inference, 92(1–2):297, January 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000719.

Janssen:2004:ARE

[Jan04] Arnold Janssen. Asymptotic relative efficiency of tests at the boundary of regular statistical models. Journal of Statistical Planning and Inference, 126(2):461–477, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002829.

Jaruskova:2007:MLL

[Jar07] Daniela Jarusková. Maximum log-likelihood ratio test for a change in three parameter Weibull distribution. Journal of Statistical Planning and Inference, 137(6):1805–1815, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580600111X.

Jones:2002:HMM

[JB02] M. C. Jones and N. Balakrishnan. How are moments and moments of spacings related to distribution functions? Journal of Statistical Planning and Inference, 103(1–2):377–390, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002324.

Jamalizadeh:2008:OSB

[JB08] A. Jamalizadeh and N. Balakrishnan. On order statistics from bivariate skew-normal and skew-t_{ν} distributions. Journal of Statistical Planning and Inference, 138(12):4187–4197, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800150X.

Jamalizadeh:2009:OST

[JB09] A. Jamalizadeh and N. Balakrishnan. Order statistics from trivariate normal and t_{ν}-distributions in terms of generalized skew-normal and skew-t_{ν} distributions. Journal of Statistical Planning and Inference, 139(11):3799–3819, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001463.

Jin:2005:EAC

[JCS05] Ruichen Jin, Wei Chen, and Agus Sudjianto. An efficient algorithm for constructing optimal design of computer experiments. Journal of Statistical Planning and Inference, 134(1):268–287, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001922.

Jeong:2000:AOA

[Jeo00] Yousceek Jeong. Asymptotically optimal adaptive designs in factorial experiments. Journal of Statistical Planning and Inference, 84(1–2):263–283, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002572.

Johnson:2000:DGS

[JG00] Wesley O. Johnson and Joseph L. Gastwirth. Dual group screening. Journal of Statistical Planning and Inference, 83(2):449–473, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001007.

Jalali:2005:OCL

[JHCH05] A. Jalali, A. G. Hawkes, L. R. Cui, and F. K. Hwang. The optimal consecutive-k-out-of-n: G line for n ≤ 2k. Journal of Statistical Planning and Inference, 128(1):281–287, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003264.

Jammalamadaka:2004:ASC

[JI04] S. Rao Jammalamadaka and Srikanth K. Iyer. Approximate self consistency for middle-censored data. Journal of Statistical Planning and Inference, 124(1):75–86, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300199X.

Jose:2007:SUC

[JI07] K. K. Jose and Jikcey Isaac. Single use conservative confidence regions in multivariate controlled calibration. Journal of Statistical Planning and Inference, 137(4):1226–1235, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000760.

Jiang:2003:EBP

[Jia03a] Jiming Jiang. Empirical best prediction for small-area inference based on generalized linear mixed models. Journal of Statistical Planning and Inference, 111(1–2):117–127, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002938.

Jiang:2003:EMM

[Jia03b] Jiming Jiang. Empirical method of moments and its applications. Journal of Statistical Planning and Inference, 115 (1):69–84, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802001180.

Jia:2006:NEM

[Jia06] Siwei Jia. A note on the economic management of inventory or resource under stochastic prices. Journal of Statistical Planning and Inference, 136(3):1020–1038, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003386.

Jin:2007:ERM

[Jin07] Zhezhen Jin. M-estimation in regression models for censored data. Journal of Statistical Planning and Inference, 137(12):3894–3903, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001589.

Johnson:2003:SNM

[JK03] Norman Johnson and Samuel Kotz. Some novel models of distributions over the unit square. Journal of Statistical Planning and Inference, 109(1–2):3–30, January 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002926.

Jiang:2008:PMR

[JK08] Wenyu Jiang and John D. Kalbfleisch. Permutation methods in relative risk regression models. Journal of Statistical Planning and Inference, 138(2):416–431, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002522.

Jiang:2004:CMS

[JL04] Wenxin Jiang and Xiangyang Liu. Consistent model selection based on parameter estimates. Journal of Statistical Planning and Inference, 121(2):265–283, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001125.

Jeong:2009:NAR

[JL09a] Jong-Hyeon Jeong and Seung-Yeoun Lee. A note on the asymptotic relative efficiency of estimators from the additive hazards model. Journal of Statistical Planning and Inference, 139(12):4231–4239, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001657.

Jin:2009:CCP

[JL09b] Hua Jin and Ying Lu. Comparison of correlated proportions based on paired binary data from clustered samples. Journal of Statistical Planning and Inference, 139(12):4206–4212, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001931.

Jones:2008:BOS

[JLN08] Bradley Jones, Dennis K. J. Lin, and Christopher J. Nachtsheim. Bayesian D-optimal supersaturated designs. Journal of Statistical Planning and Inference, 138(1):86–92, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001991.

Jones:2007:MDA

[JLNY07] Bradley A. Jones, William Li, Christopher J. Nachtsheim, and Kenny Q. Ye. Model discrimination-another perspective on model-robust designs. Journal of Statistical Planning and Inference, 137(5):1576–1583, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002291.

Jones:2009:MRS

[JLNY09] Bradley A. Jones, William Li, Christopher J. Nachtsheim, and Kenny Q. Ye. Model-robust supersaturated and partially supersaturated designs. Journal of Statistical Planning and Inference, 139(1):45–53, January 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002280.

Jin:2008:OSB

[JM08] Bo Jin and J. P. Morgan. Optimal saturated block designs when observations are correlated. Journal of Statistical Planning and Inference, 138(11):3299–3308, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000773.

Jiang:2009:RCB

[JMJ09] Jiancheng Jiang, J. S. Marron, and Xuejun Jiang. Robust centroid based classification with minimum error rates for high dimension, low sample size data. Journal of Statistical Planning and Inference, 139(8):2571–2580, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004680.

Jones:2002:CBD

[Jon02] M. C. Jones. The complementary beta distribution. Journal of Statistical Planning and Inference, 104(2):329–337, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002609.

Jones:2004:SEV

[Jon04] M. C. Jones. On some expressions for variance, covariance, skewness and L-moments. Journal of Statistical Planning and Inference, 126(1):97–106, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002738.

Jones:2006:NRR

[Jon06] M. C. Jones. A note on rescalings, reparametrizations and classes of distributions. Journal of Statistical Planning and Inference, 136(10):3730–3733, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000765.

Jones:2008:RS

[Jon08] M. C. Jones. On reciprocal symmetry. Journal of Statistical Planning and Inference, 138(10):3039–3043, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807004053.

Jones:2009:DRS

[Jon09] Bradley Jones. Discussion of “Response surface design evaluation and comparison” by Christine Anderson-Cook, Connie Borror and Douglas Montgomery. Journal of Statistical Planning and Inference, 139(2):642–644, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001584. See [ACBM09b].

Jureckova:2002:GFT

[JPS02] Jana Jurecková, Jan Picek, and Pranab Kumar Sen. A goodness-of-fit test with nuisance parameters: numerical performance. Journal of Statistical Planning and Inference, 102(2):337–347, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001082.

Jones:2008:REP

[JPS^{+}08] M. C. Jones, Heungsun Park, Key-Il Shin, S. K. Vines, and Seok-Oh Jeong. Relative error prediction via kernel regression smoothers. Journal of Statistical Planning and Inference, 138(10):2887–2898, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003928.

Jin:2003:RPT

[JR03a] Rungao Jin and John Robinson. Robust permutation tests for one sample. Journal of Statistical Planning and Inference, 116(2):475–487, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003567.

John:2003:OCD

[JR03b] J. A. John and K. G. Russell. Optimising changeover designs using the average efficiency factors. Journal of Statistical Planning and Inference, 113(1):259–268, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003044.

Jureckova:2000:GFT

[JS00] Jana Jurecková and Pranab Kumar Sen. Goodness-of-fit tests and second-order asymptotic relations. Journal of Statistical Planning and Inference, 91(2):377–397, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001890.

Johansen:2008:ERE

[JS08] Søren Johansen and Anders Rygh Swensen. Exact rational expectations, cointegration, and reduced rank regression. Journal of Statistical Planning and Inference, 138(9):2738–2748, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000633.

Jennison:2001:GST

[JT01] Christopher Jennison and Bruce W. Turnbull. On group sequential tests for data in unequally sized groups and with unknown variance. Journal of Statistical Planning and Inference, 96(1):263–288, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003396.

Jin:2002:ECG

[JTS02] Ying-Lie Jin, Shinsei Tazawa, and Teruhiro Shirakura. Enumeration of connected graphs with cut vertices. Journal of Statistical Planning and Inference, 106(1–2):409–418, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002252.

Janssen:2008:RAH

[JÜ08] Arnold Janssen and Hülya Ünlü. Regions of alternatives with high and low power for goodness-of-fit tests. Journal of Statistical Planning and Inference, 138(8):2526–2543, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003758.

Jupp:2009:DDT

[Jup09] P. E. Jupp. Data-driven tests of uniformity on product manifolds. Journal of Statistical Planning and Inference, 139(11):3820–3829, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001475.

Jin:2009:MDE

[JW09] Yinghua Jin and Yaohua Wu. Minimum ϕ-divergence estimator and hierarchical testing in loglinear models under product-multinomial sampling. Journal of Statistical Planning and Inference, 139(10):3488–3500, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001220.

Karunamuni:2006:LAT

[KA06] R. J. Karunamuni and T. Alberts. A locally adaptive transformation method of boundary correction in kernel density estimation. Journal of Statistical Planning and Inference, 136(9):2936–2960, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580500039X.

Khaledi:2008:SCB

[KA08] Baha-Eldin Khaledi and S. Saeed Ahmadi. On stochastic comparison between aggregate claim amounts. Journal of Statistical Planning and Inference, 138(7):2243–2251, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003849.

Kropf:2009:RTP

[KA09] Siegfried Kropf and Daniela Adolf. Rotation test with pairwise distance measures of sample vectors in a GLM. Journal of Statistical Planning and Inference, 139(11):3857–3864, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001529.

Kagan:2001:DVS

[Kag01a] Abram Kagan. A discrete version of the Stam inequality and a characterization of the Poisson distribution. Journal of Statistical Planning and Inference, 92(1–2):7–12, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800001580.

Kagan:2001:LE

[Kag01b] Abram Kagan. Letter to the Editor. Journal of Statistical Planning and Inference, 99(1):1, November 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002579.

Kagan:2002:SAC

[Kag02] Abram Kagan. Sufficiency and ancillarity in characterization problems. Journal of Statistical Planning and Inference, 102(2):223–228, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801001094.

Kagan:2006:QIR

[Kag06] Abram Kagan. Quasi-independence of random variables and a property of the normal and gamma distributions. Journal of Statistical Planning and Inference, 136(1):199–208, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002952.

Kageyama:2008:NVA

[Kag08] Sanpei Kageyama. Non-validity of affine α-resolvability in regular group divisible designs. Journal of Statistical Planning and Inference, 138(11):3295–3298, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000761.

Kakihara:2002:SDV

[Kak02] Yûichirô Kakihara. Spectral domains of vector harmonizable processes. Journal of Statistical Planning and Inference, 100(2):93–108, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001252.

Kakihara:2006:EDS

[Kak06] Yûichirô Kakihara. Ergodic decomposition of a stationary channel operator. Journal of Statistical Planning and Inference, 136(3):770–779, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003404.

Kalogeropoulos:2007:LBI

[Kal07] Konstantinos Kalogeropoulos. Likelihood-based inference for a class of multivariate diffusions with unobserved paths. Journal of Statistical Planning and Inference, 137(10):3092–3102, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000778.

Kano:2002:VSS

[Kan02] Yutaka Kano. Variable selection for structural models. Journal of Statistical Planning and Inference, 108(1–2):173–187, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580200277X.

Kashtanov:2000:SEG

[Kas00] Yuri N. Kashtanov. The systematic error in the generations method for a kernel of mixed sign. Journal of Statistical Planning and Inference, 85(1–2):27–35, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000634.

Katona:2002:SSS

[Kat02] Gyula O. H. Katona. Search with small sets in presence of a liar. Journal of Statistical Planning and Inference, 100(2):319–336, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001434.

Kauermann:2005:NSP

[Kau05] Göran Kauermann. A note on smoothing parameter selection for penalized spline smoothing. Journal of Statistical Planning and Inference, 127(1–2):53–69, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003124.

Kundu:2000:AID

[KB00] Debasis Kundu and Sankarshan Basu. Analysis of incomplete data in presence of competing risks. Journal of Statistical Planning and Inference, 87(2):221–239, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001937.

Kateri:2008:SEC

[KB08] M. Kateri and N. Balakrishnan. Statistical evidence in contingency tables analysis. Journal of Statistical Planning and Inference, 138(4):873–887, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000560.

Krishna:2009:BVS

[KBG09] Arun Krishna, Howard D. Bondell, and Sujit K. Ghosh. Bayesian variable selection using an adaptive powered correlation prior. Journal of Statistical Planning and Inference, 139(8):2665–2674, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004515.

Kim:2005:MSC

[KC05] Hyun-Joo Kim and Joseph E. Cavanaugh. Model selection criteria based on Kullback information measures for nonlinear regression. Journal of Statistical Planning and Inference, 134(2):332–349, October 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002241.

Kulinskaya:2000:AGW

[KD00] Elena Kulinskaya and Michael B. Dollinger. Asymptotics of guarded weights of evidence. Journal of Statistical Planning and Inference, 86(1):11–29, April 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001573.

Kale:2001:ECS

[KD01] Mohan Kale and S. R. Deshmukh. Estimation in continuous state-space branching processes. Journal of Statistical Planning and Inference, 92(1–2):111–119, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001415.

Kulinskaya:2007:RWO

[KD07] Elena Kulinskaya and Michael B. Dollinger. Robust weighted one-way ANOVA: Improved approximation and efficiency. Journal of Statistical Planning and Inference, 137(2):462–472, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000577.

Katsaounis:2008:SEM

[KD08] T. I. Katsaounis and A. M. Dean. A survey and evaluation of methods for determination of combinatorial equivalence of factorial designs. Journal of Statistical Planning and Inference, 138(1):245–258, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002182.

Kung:2009:LSP

[KdMSY09] Joseph P. S. Kung, Anna de Mier, Xinyu Sun, and Catherine Yan. Lattice and Schröder paths with periodic boundaries. Journal of Statistical Planning and Inference, 139(6):2014–2027, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003674.

Key:2001:TDC

[KdR01] J. D. Key and M. J. de Resmini. Ternary dual codes of the planes of order nine. Journal of Statistical Planning and Inference, 95(1–2):229–236, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002901.

Kang:2006:DSG

[KDT06] Qingde Kang, Yanke Du, and Zihong Tian. Decomposition of λK_{v} into some graph with six vertices and seven edges. Journal of Statistical Planning and Inference, 136(4):1394–1409, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003970.

Keener:2005:LID

[Kee05] Robert W. Keener. Local information and the design of sequential hypothesis tests. Journal of Statistical Planning and Inference, 130(1–2):111–125, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002563.

Kemp:2002:ECP

[Kem02a] Adrienne W. Kemp. Existence conditions and properties for the generalized Euler family of distributions. Journal of Statistical Planning and Inference, 101(1–2):169–178, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001641.

Kemp:2002:RWO

[Kem02b] Adrienne W. Kemp. Random walks and other aspects of the Bailey-daum distribution. Journal of Statistical Planning and Inference, 101(1–2):163–168, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100163X.

Kemp:2002:AHP

[Kem02c] C. D. Kemp. q-analogues of the hyper-Poisson distribution. Journal of Statistical Planning and Inference, 101(1–2):179–183, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001665.

Kemp:2003:CIC

[Kem03] Adrienne W. Kemp. Characterizations involving U|(U + v = m) for certain discrete distributions. Journal of Statistical Planning and Inference, 109(1–2):31–41, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002963.

Kemp:2005:SSM

[Kem05] Adrienne W. Kemp. Steady-state Markov chain models for certain q-confluent hypergeometric distributions. Journal of Statistical Planning and Inference, 135(1):107–120, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000583.

Kennedy:2000:BAC

[Ken00] Marc C. Kennedy. A Bayesian approach to computing posterior distribution and quantile functions. Journal of Statistical Planning and Inference, 83(1):183–201, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000269.

Kerr:2004:SGP

[Ker04] M. Kathleen Kerr. Stationary Gaussian processes on the vertices of the k-cube. Journal of Statistical Planning and Inference, 118(1–2):1–8, January 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004007.

Krishnamoorthy:2005:AOE

[KG05] K. Krishnamoorthy and Huizhen Guo. Assessing occupational exposure via the one-way random effects model with unbalanced data. Journal of Statistical Planning and Inference, 128(1):219–229, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300329X.

Kabaila:2009:CIR

[KG09] Paul Kabaila and Khageswor Giri. Confidence intervals in regression utilizing prior information. Journal of Statistical Planning and Inference, 139(10):3419–3429, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000901.

Kundu:2005:DBL

[KGM05] Debasis Kundu, Rameshwar D. Gupta, and Anubhav Manglick. Discriminating between the log-normal and generalized exponential distributions. Journal of Statistical Planning and Inference, 127(1–2):213–227, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300288X.

Khan:2000:NHE

[Kha00] Rasul A. Khan. A note on Hammersley’s estimator of an integer mean. Journal of Statistical Planning and Inference, 88(1):37–45, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899002049.

Khattree:2003:AAS

[Kha03] Ravindra Khattree. Antieigenvalues and antieigenvectors in statistics. Journal of Statistical Planning and Inference, 114 (1–2):131–144, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802004676.

Kwong:2002:MBH

[KHC02] Koon Shing Kwong, Burt Holland, and Siu Hung Cheung. A modified Benjamini–Hochberg multiple comparisons procedure for controlling the false discovery rate. Journal of Statistical Planning and Inference, 104(2):351–362, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580100252X.

Kitsantas:2007:ASC

[KHL07] Panagiota Kitsantas, Myles Hollander, and Lei M. Li. Assessing the stability of classification trees using Florida birth data. Journal of Statistical Planning and Inference, 137(12):3917–3929, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001607.

Khan:2006:PIF

[KHP06] Hafiz M. R. Khan, M. Safiul Haq, and Serge B. Provost. Predictive inference for future responses given a doubly censored sample from a two parameter exponential distribution. Journal of Statistical Planning and Inference, 136(9):3156–3172, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004719.

Khuri:2009:DRS

[Khu09] André I. Khuri. Discussion of “Response surface design evaluation and comparison” by Christine M. Anderson-Cook, Connie M. Borror, Douglas C. Montgomery. Journal of Statistical Planning and Inference, 139(2):647–649, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001602. See [ACBM09b].

Kim:2000:DBF

[KI00] Seong W. Kim and Joseph G. Ibrahim. Default Bayes factors for generalized linear models. Journal of Statistical Planning and Inference, 87(2):301–315, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001949.

Kakizawa:2008:HOS

[KI08] Yoshihide Kakizawa and Toshiya Iwashita. Hotelling’s one-sample and two-sample T^{2} tests and the multivariate Behrens–Fisher problem under nonnormality. Journal of Statistical Planning and Inference, 138(11):3379–3404, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000840.

Khondoker:2009:ULL

[KI09] Mizanur R. Khondoker and M. Ataharul Islam. Use of log-log survival function in modeling time-covariate interactions in Cox regression. Journal of Statistical Planning and Inference, 139(6):1968–1973, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003649.

Kirch:2007:BPP

[Kir07a] Claudia Kirch. Block permutation principles for the change analysis of dependent data. Journal of Statistical Planning and Inference, 137(7):2453–2474, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002825.

Kirk:2007:EMD

[Kir07b] Roger E. Kirk. Effect magnitude: a different focus. Journal of Statistical Planning and Inference, 137(5):1634–1646, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806002424.

Kish:2002:CMS

[Kis02] Leslie Kish. Combining multipopulation statistics. Journal of Statistical Planning and Inference, 102(1):109–118, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001756.

Kakiuchi:2001:RRT

[KK01]  Itsuro Kakiuchi and Miyoshi Kimura.  Robust rank

tests for k-sample approximate equality in the presence of gross errors. Journal of Statistical Planning and Inference, 93(1–2):117–138, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001701.

Khaledi:2002:DOA

[KK02] Baha-Eldin Khaledi and Subhash Kochar. Dispersive ordering among linear combinations of uniform random variables. Journal of Statistical Planning and Inference, 100(1):13–21, January 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580100091X.

Konishi:2003:ATI

[KK03] Sadanori Konishi and Genshiro Kitagawa. Asymptotic theory for information criteria in model selection-functional approach. Journal of Statistical Planning and Inference, 114 (1–2):45–61, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802004627.

Kakiuchi:2005:RSR

[KK05] Itsuro Kakiuchi and Miyoshi Kimura. Robust slippage rank tests for k location parameters in the presence of gross errors. Journal of Statistical Planning and Inference, 128(1):13–29, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003240.

Khaledi:2006:WDS

[KK06] Baha-Eldin Khaledi and Subhash Kochar. Weibull distribution: Some stochastic comparisons results. Journal of Statistical Planning and Inference, 136(9):3121–3129, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000352.

Kaarik:2009:MDC

[KK09a] Ene Käärik and Meelis Käärik. Modeling dropouts by conditional distribution, a copula-based approach. Journal of Statistical Planning and Inference, 139(11):3830–3835, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001487.

Kayano:2009:FPC

[KK09b] Mitsunori Kayano and Sadanori Konishi. Functional principal component analysis via regularized Gaussian basis expansions and its application to unbalanced data. Journal of Statistical Planning and Inference, 139(7):2388–2398, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004266.

Kemp:2009:CD

[KK09c]  Adrienne W. Kemp and C. David Kemp.  The q-

cluster distribution. Journal of Statistical Planning and Inference, 139(6):1856–1866, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003893.

Kim:2009:BPL

[KK09d] Yongdai Kim and Dohyun Kim. Bayesian partial likelihood approach for tied observations. Journal of Statistical Planning and Inference, 139(2):469–477, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002139.

Koukouvinos:2009:P

[KK09e] Christos Koukouvinos and Ilias Kotsireas. Preface. Journal of Statistical Planning and Inference, 139(1):2, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002231.

Koyuncu:2009:RPE

[KK09f] Nursel Koyuncu and Cem Kadilar. Ratio and product estimators in stratified random sampling. Journal of Statistical Planning and Inference, 139(8):2552–2558, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004424.

Kateri:2009:MAA

[KKB09a] M. Kateri, U. Kamps, and N. Balakrishnan. A meta-analysis approach for step-stress experiments. Journal of Statistical Planning and Inference, 139(9):2907–2919, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000159.

Kokonendji:2009:SEC

[KKB09b] C. C. Kokonendji, T. Senga Kiessé, and N. Balakrishnan. Semiparametric estimation for count data through weighted distributions. Journal of Statistical Planning and Inference, 139(10):3625–3638, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001049.

Kamali:2003:SBT

[KKK03] F. Kamali, Hadi Kharaghani, and G. B. Khosrovshahi. Some Bush-type Hadamard matrices. Journal of Statistical Planning and Inference, 113(2):375–384, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003184.

Kim:2005:NPN

[KKL05] Dal Ho Kim, Sang Gil Kang, and Woo Dong Lee. Noninformative priors for the nested design. Journal of Statistical Planning and Inference, 133(2):453–462, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001703.

Kim:2007:BPR

[KKLD07] Che Soong Kim, Valentina Klimenok, Sang Cheon Lee, and Alexander Dudin. The BMAP/PH /1 retrial queueing system operating in random environment. Journal of Statistical Planning and Inference, 137(12):3904–3916, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001590.

Krishnamoorthy:2001:MUO

[KKM01] K. Krishnamoorthy, Pandurang M. Kulkarni, and Thomas Mathew. Multiple use one-sided hypotheses testing in univariate linear calibration. Journal of Statistical Planning and Inference, 93(1–2):211–223, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001646.

Kaufman:2009:ABC

[KKM09] Sol Kaufman, Jay S. Kaufman, and Richard F. MacLehose. Analytic bounds on causal risk differences in directed acyclic graphs involving three observed binary variables. Journal of Statistical Planning and Inference, 139(10):3473–3487, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000925.

Kohler:2009:OGR

[KKW09] Michael Kohler, Adam Krzyzak, and Harro Walk. Optimal global rates of convergence for nonparametric regression with unbounded data. Journal of Statistical Planning and Inference, 139(4):1286–1296, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003261.

Kagan:2001:PLF

[KL01] Abram Kagan and R. G. Laha. A property of linear forms of independent random variables related to uniqueness of linear structure. Journal of Statistical Planning and Inference, 92(1–2):13–20, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001592.

Kim:2005:KDE

[KL05] Tae Yoon Kim and Sangyeol Lee. Kernel density estimator for strong mixing processes. Journal of Statistical Planning and Inference, 133(2):273–284, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001193.

Kabaila:2006:IBL

[KL06] Paul Kabaila and Chris J. Lloyd. Improved Buehler limits based on refined designated statistics. Journal of Statistical Planning and Inference, 136(9):3145–3155, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004707.

Kang:2007:SCA

[KL07] Seung-Ho Kang and Jae-Won Lee. The size of the Cochran– Armitage trend test in 2 × C contingency tables. Journal of Statistical Planning and Inference, 137(6):1851–1861, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001121.

Kagan:2008:IFI

[KL08] Abram Kagan and Bing Li. An identity for the Fisher information and Mahalanobis distance. Journal of Statistical Planning and Inference, 138(12):3950–3959, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001481.

Keziou:2008:ELS

[KLA08] Amor Keziou and Samuela Leoni-Aubin. On empirical likelihood for semiparametric two-sample density ratio models. Journal of Statistical Planning and Inference, 138(4):915–928, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001383.

Klemela:2003:LBA

[Kle03] Jussi Klemelä. Lower bounds for the asymptotic minimax risk with spherical data. Journal of Statistical Planning and Inference, 113(1):113–136, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003032.

Kleiber:2004:LOO

[Kle04a] Christian Kleiber. Lorenz ordering of order statistics from log-logistic and related distributions. Journal of Statistical Planning and Inference, 120(1–2):13–19, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004950.

Klein:2004:ODS

[Kle04b] Thomas Klein. Optimal designs for second-degree Kronecker model mixture experiments. Journal of Statistical Planning and Inference, 123(1):117–131, June 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001459.

Kropf:2004:NMT

[KLE^{+}04c] Siegfried Kropf, Jürgen Läuter, Markus Eszlinger, Knut Krohn, and Ralf Paschke. Nonparametric multiple test procedures with data-driven order of hypotheses and with weighted hypotheses. Journal of Statistical Planning and Inference, 125(1–2):31–47, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000540.

Kagan:2007:SSC

[KLR07] Abram Kagan, Zinoviy Landsman, and C. Radhakrishna Rao. Sub- and superadditivity à la Carlen of matrices related to the Fisher information. Journal of Statistical Planning and Inference, 137(1):291–298, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002697.

Krishnamoorthy:2009:CLP

[KLX09] K. Krishnamoorthy, Yin Lin, and Yanping Xia. Confidence limits and prediction limits for a Weibull distribution based on the generalized variable approach. Journal of Statistical Planning and Inference, 139(8):2675–2684, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004655.

Koolen:2000:TDP

[KM00a] J. H. Koolen and A. Munemasa. Tight 2-designs and perfect 1-codes in Doob graphs. Journal of Statistical Planning and Inference, 86(2):505–513, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001263.

Kunert:2000:OTO

[KM00b] J. Kunert and R. J. Martin. Optimality of type I orthogonal arrays for cross-over models with correlated errors. Journal of Statistical Planning and Inference, 87(1):119–124, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001822.

Kageyama:2001:END

[KM01] Sanpei Kageyama and Ying Miao. Existence of nested designs with block size five. Journal of Statistical Planning and Inference, 94(2):249–254, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002573.

Krishnamoorthy:2003:IML

[KM03] K. Krishnamoorthy and Thomas Mathew. Inferences on the means of lognormal distributions using generalized p-values and generalized confidence intervals. Journal of Statistical Planning and Inference, 115(1):103–121, July 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001532.

Kim:2004:BPU

[KM04] Hyoung-Moon Kim and Bani K. Mallick. A Bayesian prediction using the skew Gaussian distribution. Journal of Statistical Planning and Inference, 120(1–2):85–101, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802005013.

Karlis:2007:FMM

[KM07a] Dimitris Karlis and Loukia Meligkotsidou. Finite mixtures of multivariate Poisson distributions with application. Journal of Statistical Planning and Inference, 137(6):1942–1960, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001753.

Kim:2007:TSH

[KM07b] Tae-Hwan Kim and Christophe Muller. Two-stage Huber estimation. Journal of Statistical Planning and Inference, 137(2):405–418, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000437.

Krinik:2007:TPF

[KM07c] Alan Krinik and Carrie Mortensen. Transient probability functions of finite birth-death processes with catastrophes. Journal of Statistical Planning and Inference, 137(5):1530– 1543, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002266.

Karagiannis:2008:GCO

[KM08] Vassilis Karagiannis and Chronis Moyssiadis. A graphical construction of the D-optimal saturated, 3× s2 main effect, factorial design. Journal of Statistical Planning and Inference, 138(6):1679–1696, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003473.

Kokonendji:2008:CPW

[KMB08] Célestin C. Kokonendji, Dominique Mizère, and N. Balakrishnan. Connections of the Poisson weight function to overdispersion and underdispersion. Journal of Statistical Planning and Inference, 138(5):1287–1296, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002200.

Kadane:2002:ANP

[KMPP02] Joseph B. Kadane, Elias Moreno, Maria Eglee Perez, and Luis Raul Pericchi. Applying non-parametric robust Bayesian analysis to non-opinionated judicial neutrality. Journal of Statistical Planning and Inference, 102(2):425–439, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001100.

Kordzakhia:2001:REL

[KMR01] N. Kordzakhia, G. D. Mishra, and L. Reiersølmoen. Robust estimation in the logistic regression model. Journal of Statistical Planning and Inference, 98(1–2):211–223, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800003128.

Koukouvinos:2008:OMO

[KMS08] Christos Koukouvinos, Kalliopi Mylona, and Dimitris E. Simos. E(s^{2})-optimal and minimax-optimal cyclic supersaturated designs via multi-objective simulated annealing. Journal of Statistical Planning and Inference, 138(6):1639–1646, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003126.

Koukouvinos:2009:HSA

[KMS09a] Christos Koukouvinos, Kalliopi Mylona, and Dimitris E. Simos. A hybrid SAGA algorithm for the construction of E(s^{2})-optimal cyclic supersaturated designs. Journal of Statistical Planning and Inference, 139(2):478–485, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001997.

Kukush:2009:OQS

[KMS09b] Alexander Kukush, Andrii Malenko, and Hans Schneeweiss. Optimality of the quasi-score estimator in a mean-variance model with applications to measurement error models. Journal of Statistical Planning and Inference, 139(10):3461–3472, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000895.

Krinik:2002:TPS

[KMSC02] Alan Krinik, Dan Marcus, Ray Shiflett, and LiPing Chu. Transient probabilities of a single server priority queueing system. Journal of Statistical Planning and Inference, 101(1–2):185–190, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001768.

Kukush:2005:CST

[KMV05] A. Kukush, I. Markovsky, and S. Van Huffel. Consistency of the structured total least squares estimator in a multivariate errors-in-variables model. Journal of Statistical Planning and Inference, 133(2):315–358, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001752.

Khattree:2002:APM

[KN02] Ravindra Khattree and Dayanand N. Naik. Andrews plots for multivariate data: some new suggestions and applications. Journal of Statistical Planning and Inference, 100(2):411–425, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001501.

Kundu:2003:ENC

[KN03] Debasis Kundu and Swagata Nandi. Estimating the number of components by a data independent penalty function approach. Journal of Statistical Planning and Inference, 116(2):437–450, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003506.

Koul:2004:MDR

[KN04] Hira L. Koul and Pingping Ni. Minimum distance regression model checking. Journal of Statistical Planning and Inference, 119(1):109–141, January 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004159.

Kundu:2009:NRH

[KNH09] Chanchal Kundu, Asok K. Nanda, and Taizhong Hu. A note on reversed hazard rate of order statistics and record values. Journal of Statistical Planning and Inference, 139(4): 1257–1265, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003133.

Khosrovshahi:2001:TTT

[KNTR01] G. B. Khosrovshahi, Reza Naserasr, and B. Tayfeh-Rezaie. Ternary trades and their codes. Journal of Statistical Planning and Inference, 95(1–2):237–243, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002913.

Ko:2000:BRA

[Ko00] Daijin Ko. Book review: Applied survival analysis: regression modeling of time to event data: David W. Hosmer, Jr. and Stanley Lemeshow, Wiley, New York, 1999, pp. 386 + xiii, price $84.95 ISBN 0-471-15410-5. Journal of Statistical Planning and Inference, 91(1):173–175, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001300.

Konsowa:2003:FDR

[KO03] Mokhtar H. Konsowa and Tamer F. Oraby. Fractal dimensions and random walks on random trees. Journal of Statistical Planning and Inference, 116(2):333–342, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580200352X.

Kohler:2000:IUD

[Koh00] Michael Kohler. Inequalities for uniform deviations of averages from expectations with applications to nonparametric regression. Journal of Statistical Planning and Inference, 89(1–2):1–23, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899002153.

Kohler:2003:NOS

[Koh03] Michael Kohler. Nonlinear orthogonal series estimates for random design regression. Journal of Statistical Planning and Inference, 115(2):491–520, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001581.

Kohler:2006:NRA

[Koh06] Michael Kohler. Nonparametric regression with additional measurement errors in the dependent variable. Journal of Statistical Planning and Inference, 136(10):3339–3361, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805000704.

Kohler:2008:MOS

[Koh08] Michael Kohler. Multivariate orthogonal series estimates for random design regression. Journal of Statistical Planning and Inference, 138(10):3217–3237, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000475.

Kollo:2009:P

[Kol09] Tõnu Kollo. Preface. Journal of Statistical Planning and Inference, 139(11):3740, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001372.

Konno:2007:ISC

[Kon07] Yoshihiko Konno. Improving on the sample covariance matrix for a complex elliptically contoured distribution. Journal of Statistical Planning and Inference, 137(7):2475–2486, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580600276X.

Korner:2000:ATE

[Kör00] Ralf Körner. An asymptotic α-test for the expectation of random fuzzy variables. Journal of Statistical Planning and Inference, 83(2):331–346, February 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900107X.

Kaluszka:2005:SBS

[KOS05] M. Kaluszka, A. Okolewski, and K. Szymanska. Sharp bounds for L-statistics from dependent samples of random length. Journal of Statistical Planning and Inference, 127(1–2):71–89, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002908.

Kott:2005:RAM

[Kot05]  Phillip S. Kott.  Randomization-assisted model-based

survey sampling. Journal of Statistical Planning and Inference, 129(1–2):263–277, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001934.

Kottas:2006:NBS

[Kot06] Athanasios Kottas. Nonparametric Bayesian survival analysis using mixtures of Weibull distributions. Journal of Statistical Planning and Inference, 136(3):578–596, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003465.

Kowalski:2009:OER

[Kow09] Jan Kowalski. Optimal estimation in rotation patterns. Journal of Statistical Planning and Inference, 139(4):1405–1420, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800342X.

Katzenbeisser:2002:MHS

[KP02] Walter Katzenbeisser and Wolfgang Panny. The maximal height of simple random walks revisited. Journal of Statistical Planning and Inference, 101(1–2):149–161, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001628.

Kozubowski:2004:TSU

[KP04] Tomasz J. Kozubowski and Anna K. Panorska. Testing symmetry under a skew Laplace model. Journal of Statistical Planning and Inference, 120(1–2):41–63, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802005037.

Kozubowski:2005:MBD

[KP05]  Tomasz J. Kozubowski and Anna K. Panorska.  A

mixed bivariate distribution with exponential and geometric marginals. Journal of Statistical Planning and Inference, 134(2):501–520, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002162.

Karlis:2008:BCI

[KP08a] Dimitris Karlis and Valentin Patilea. Bootstrap confidence intervals in mixtures of discrete distributions. Journal of Statistical Planning and Inference, 138(8):2313–2329, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000219.

Koning:2008:GFT

[KP08b] Alex J. Koning and Liang Peng. Goodness-of-fit tests for a heavy tailed distribution. Journal of Statistical Planning and Inference, 138(12):3960–3981, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001432.

Kaarik:2009:QMC

[KP09a] Meelis Käärik and Kalev Pärna. On the quality of k-means clustering based on grouped data. Journal of Statistical Planning and Inference, 139(11):3836–3841, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001499.

Kolev:2009:CBR

[KP09b] Nikolai Kolev and Delhi Paiva. Copula-based regression models: a survey. Journal of Statistical Planning and Inference, 139(11):3847–3856, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001517.

Kaur:2002:RSS

[KPTW02] A. Kaur, G. P. Patil, C. Taillie, and J. Wit. Ranked set sample sign test for quantiles. Journal of Statistical Planning and Inference, 100(2):337–347, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001446.

Koul:2002:AML

[KQ02] Hira L. Koul and Lianfen Qian. Asymptotics of maximum likelihood estimator in a two-phase linear regression model. Journal of Statistical Planning and Inference, 108 (1–2):99–119, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802002732.

Kagan:2003:SPA

[KR03] Abram Kagan and C. Radhakrishna Rao. Some properties and applications of the efficient Fisher score. Journal of Statistical Planning and Inference, 116(2):343–352, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003518.

Kolassa:2007:CSA

[KR07] John E. Kolassa and John Robinson. Conditional saddlepoint approximations for non-continuous and non-lattice distributions. Journal of Statistical Planning and Inference, 137(1):133–147, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805003009.

Kramer:2005:FSP

[Krä05] Walter Krämer. Finite sample power of Cliff–Ord-type tests for spatial disturbance correlation in linear regression. Journal of Statistical Planning and Inference, 128(2):489–496, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000060.

Kraus:2009:ANS

[Kra09] David Kraus. Adaptive Neyman’s smooth tests of homogeneity of two samples of survival data. Journal of Statistical Planning and Inference, 139(10):3559–3569, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000998.

Krinik:2005:DPS

[KRM^{+}05] Alan Krinik, Gerardo Rubino, Daniel Marcus, Randall J. Swift, Hassan Kasfy, and Holly Lam. Dual processes to solve single server systems. Journal of Statistical Planning and Inference, 135(1):121–147, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000595.

Khoujmane:2007:CTE

[KRS07] Ali Khoujmane, Frits Ruymgaart, and Mikail Shubov. The convolution theorem for estimating linear functionals in indirect nonparametric regression models. Journal of Statistical Planning and Inference, 137(3):811–820, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580600142X.

Katul:2006:DOC

[KRV06] Gabriel Katul, Fabrizio Ruggeri, and Brani Vidakovic. Denoising ozone concentration measurements with BAMS filtering. Journal of Statistical Planning and Inference, 136 (7):2395–2405, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805002302.

Koul:2000:SOB

[KS00] Hira L. Koul and Donatas Surgailis. Second-order behavior of M-estimators in linear regression with long-memory errors. Journal of Statistical Planning and Inference, 91(2):399–412, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001907.

Klove:2001:DDT

[KS01] Torleiv Kløve and Yuri V. Svirid. Diffuse difference triangle sets. Journal of Statistical Planning and Inference, 94(2):255– 276, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002585.

Kageyama:2002:P

[KS02a] Sanpei Kageyama and Teruhiro Shirakura. Preface. Journal of Statistical Planning and Inference, 106(1–2):3–4, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001969.

Kenney:2002:NFC

[KS02b] John Kenney and Wynn Stirling. Nonlinear filtering of convex sets of probability distributions. Journal of Statistical Planning and Inference, 105(1):123–147, June 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002075.

Kanamori:2003:ALA

[KS03a]  Takafumi Kanamori and Hidetoshi Shimodaira.  Ac-

tive learning algorithm using the maximum weighted loglikelihood estimator. Journal of Statistical Planning and Inference, 116(1):149–162, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002343.

Koul:2003:TSA

[KS03b] Hira L. Koul and Anton Schick. Testing for superiority among two regression curves. Journal of Statistical Planning and Inference, 117(1):15–33, November 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003646.

Koukouvinos:2006:NCO

[KS06] C. Koukouvinos and S. Stylianou. New classes of Doptimal edge designs. Journal of Statistical Planning and Inference, 136(1):248–259, January 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002502.

Khaledi:2007:OCL

[KS07a] Baha-Eldin Khaledi and Moshe Shaked. Ordering conditional lifetimes of coherent systems. Journal of Statistical Planning and Inference, 137(4):1173–1184, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000735.

Kopperschmidt:2007:CUR

[KS07b] Kai Kopperschmidt and Winfried Stute. The CLT under right censorship and reporting delays. Journal of Statistical Planning and Inference, 137(3):1035–1042, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001583.

Koskinen:2007:BID

[KS07c] Johan H. Koskinen and Tom A. B. Snijders. Bayesian inference for dynamic social network data. Journal of Statistical Planning and Inference, 137(12):3930–3938, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001619.

Kottas:2007:BMM

[KS07d] Athanasios Kottas and Bruno Sansó. Bayesian mixture modeling for spatial Poisson process intensities, with applications to extreme value analysis. Journal of Statistical Planning and Inference, 137(10):3151–3163, October 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000821.

Kunert:2007:RNB

[KS07e] Joachim Kunert and Oliver Sailer. Randomization of neighbour balanced generalized Youden designs. Journal of Statistical Planning and Inference, 137(6):2045–2055, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806001157.

Kato:2008:DMO

[KS08a] Shogo Kato and Kunio Shimizu. Dependent models for observations which include angular ones. Journal of Statistical Planning and Inference, 138(11):3538–3549, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000967.

Koul:2008:RMC

[KS08b] Hira L. Koul and Weixing Song. Regression model checking with Berkson measurement errors. Journal of Statistical Planning and Inference, 138(6):1615–1628, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003266.

Kharin:2009:RMB

[KS09a] Alexey Yu. Kharin and Pavel A. Shlyk. Robust multivariate Bayesian forecasting under functional distortions in the χ^{2}-metric. Journal of Statistical Planning and Inference, 139(11):3842–3846, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001505.

Koul:2009:TSH

[KS09b] Hira L. Koul and Donatas Surgailis. Testing of a subhypothesis in linear regression models with long memory errors and deterministic design. Journal of Statistical Planning and Inference, 139(8):2715–2730, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004783.

Kuriki:2002:IRS

[KSH02] Satoshi Kuriki, Hidetoshi Shimodaira, and Tony Hayter. On the isotonic range statistic for testing against an ordered alternative. Journal of Statistical Planning and Inference, 105(2):347–362, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801002749.

Kubokawa:2000:BMN

[KSK00] Tatsuya Kubokawa, A. K. Md. Ehsanes Saleh, and Yoshihiko Konno. Bayes, minimax and nonnegative estimators of variance components under Kullback–Leibler loss. Journal of Statistical Planning and Inference, 86(1):201–214, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001706.

Kang:2003:BRM

[KSL03] Weechang Kang, Dong Wan Shin, and Youngjo Lee. Biases of the restricted maximum likelihood estimators for ARMA processes with polynomial time trend. Journal of Statistical Planning and Inference, 116(1):163–176, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002392.

Kadane:2001:GDR

[KSS01] Joseph B. Kadane, Mark J. Schervish, and Teddy Seidenfeld. Goldstein’s dilemma: require countable additivity or abandon “prevision of prevision”. Journal of Statistical Planning and Inference, 94(1):89–91, March 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002305.

Kim:2008:EED

[KSS08] Gunky Kim, Mervyn J. Silvapulle, and Paramsothy Silvapulle. Estimating the error distribution in multivariate heteroscedastic time-series models. Journal of Statistical Planning and Inference, 138(5):1442–1458, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002790.

Kukush:2005:RET

[KSW05] Alexander Kukush, Hans Schneeweiss, and Roland Wolf. Relative efficiency of three estimators in a polynomial regression with measurement errors. Journal of Statistical Planning and Inference, 127(1–2):179–203, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003112.

Kathman:2002:IPA

[KT02] Steven J. Kathman and George R. Terrell. Improving Poisson approximations through tilting and expansion. Journal of Statistical Planning and Inference, 104(2):315–327, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002610.

Krishnamoorthy:2004:MPT

[KT04] K. Krishnamoorthy and Jessica Thomson. A more powerful test for comparing two Poisson means. Journal of Statistical Planning and Inference, 119(1):23–35, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004081.

Kubokawa:2007:ELR

[KT07] Tatsuya Kubokawa and Hisayuki Tsukuma. Estimation in a linear regression model under the Kullback–Leibler loss and its application to model selection. Journal of Statistical Planning and Inference, 137(7):2487–2508, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002813.

Krishnamoorthy:2008:IDR

[KT08a] K. Krishnamoorthy and Lili Tian. Inferences on the difference and ratio of the means of two inverse Gaussian distributions.

Journal of Statistical Planning and Inference, 138(7):2082– 2089, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003618.

Kuriki:2008:ECH

[KT08b] Satoshi Kuriki and Akimichi Takemura. Euler characteristic heuristic for approximating the distribution of the largest eigenvalue of an orthogonally invariant random matrix. Journal of Statistical Planning and Inference, 138(11):3357–3378, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000839.

Kuriki:2008:TMM

[KT08c] Satoshi Kuriki and Akimichi Takemura. The tube method for the moment index in projection pursuit. Journal of Statistical Planning and Inference, 138(9):2749–2762, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000645.

Kim:2006:EMR

[KTA06] Jong-Min Kim, Joshua M. Tebbs, and Seung-Won An. Extensions of Mangat’s randomized-response model. Journal of Statistical Planning and Inference, 136(4):1554–1567, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003969.

Korn:2001:UEC

[KTB01] Edward L. Korn, David M. Teeter, and Sheldon Baumrind. Using explicit clinician preferences in nonrandomized study designs. Journal of Statistical Planning and Inference, 96(1): 67–82, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800003244.

Kamiya:2008:SSD

[KTK08] Hidehiko Kamiya, Akimichi Takemura, and Satoshi Kuriki. Star-shaped distributions and their generalizations. Journal of Statistical Planning and Inference, 138(11):3429–3447, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000876.

Korn:2004:CNF

[KTMS04] Edward L. Korn, James F. Troendle, Lisa M. McShane, and Richard Simon. Controlling the number of false discoveries: application to high-dimensional genomic data. Journal of Statistical Planning and Inference, 124(2):379–398, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002118.

Kuhfeld:2005:DSC

[Kuh05] Warren F. Kuhfeld. Difference schemes via computerized searches. Journal of Statistical Planning and Inference, 127(1–2):341–346, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003203.

Kulik:2009:EPL

[Kul09] Rafal Kulik. Empirical process of long-range dependent sequences when parameters are estimated. Journal of Statistical Planning and Inference, 139(2):287–294, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001900.

Kundu:2000:ENS

[Kun00] D. Kundu. Estimating the number of signals in the presence of white noise. Journal of Statistical Planning and Inference, 90(1):57–68, September 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001026.

Kundu:2007:HCW

[Kun07] Debasis Kundu. On hybrid censored Weibull distribution. Journal of Statistical Planning and Inference, 137(7):2127– 2142, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002060.

Kurata:2008:PPL

[Kur08] Hiroshi Kurata. On principal points for location mixtures of spherically symmetric distributions. Journal of Statistical Planning and Inference, 138(11):3405–3418, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000852.

Kushner:2003:ARA

[Kus03] H. B. Kushner. Allocation rules for adaptive repeated measurements designs. Journal of Statistical Planning and Inference, 113(1):293–313, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580100310X.

Kuwada:2002:NAM

[Kuw02] Masahide Kuwada. Norm of alias matrices for balanced fractional 2^{m} factorial designs when interesting factorial effects are not aliased with effects not of interest in estimation. Journal of Statistical Planning and Inference, 106(1–2):271–286, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002185.

Kyriakoussis:2002:CLL

[KV02] A. Kyriakoussis and M. G. Vamvakari. Central and local limit theorems for the coefficients of polynomials associated with the Laguerre ones. Journal of Statistical Planning and Inference, 101(1–2):191–202, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100177X.

Kyriakoussis:2005:UTA

[KV05] A. Kyriakoussis and M. G. Vamvakari. A unified treatment for the asymptotic normality of the coefficients of polynomials related to orthogonal ones. Journal of Statistical Planning and Inference, 135(1):148–157, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000601.

Ko:2006:BWA

[KV06] Kyungduk Ko and Marina Vannucci. Bayesian wavelet analysis of autoregressive fractionally integrated movingaverage processes. Journal of Statistical Planning and Inference, 136(10):3415–3434, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000716.

Kuchler:2007:PES

[KV07] Uwe Küchler and Vyacheslav A. Vasil’iev. On parameter estimation of stochastic delay differential equations with guaranteed accuracy by noisy observations. Journal of Statistical Planning and Inference, 137(9):3007–3023, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000365.

Kruijer:2008:PCR

[KvdV08] Willem Kruijer and Aad van der Vaart. Posterior convergence rates for Dirichlet mixtures of beta densities. Journal of Statistical Planning and Inference, 138(7):1981–1992, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003321.

Kuljus:2008:RCM

[KvR08] Kristi Kuljus and Dietrich von Rosen. Rank covariance matrix estimation of a partially known covariance matrix. Journal of Statistical Planning and Inference, 138(12):3667–3673, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001201.

Kim:2004:SWR

[KW04] Jong-Min Kim and William D. Warde. A stratified Warner’s randomized response model. Journal of Statistical Planning and Inference, 120(1–2):155–165, February 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802005001.

Kim:2005:MRR

[KW05] Jong-Min Kim and William D. Warde. A mixed randomized response model. Journal of Statistical Planning and Inference, 133(1):211–221, July 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001132.

Karunamuni:2009:MHD

[KW09a] R. J. Karunamuni and J. Wu. Minimum Hellinger distance estimation in a nonparametric mixture model. Journal of Statistical Planning and Inference, 139(3):1118–1133, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580800298X.

Kim:2009:SOV

[KW09b] Kyongryun Kim and Suojin Wang. Second-order variance estimation in poststratified two-stage sampling. Journal of Statistical Planning and Inference, 139(7):2502–2510, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004382.

Kwong:2001:MDT

[Kwo01] Koon-Shing Kwong. A modified Dunnett and Tamhane stepup approach for establishing superiority/equivalence of a new treatment compared with k standard treatments. Journal of Statistical Planning and Inference, 97(2):359–366, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002238.

Karlis:2001:RIF

[KX01] Dimitris Karlis and Evdokia Xekalaki. Robust inference for finite Poisson mixtures. Journal of Statistical Planning and Inference, 93(1–2):93–115, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000207X.

Krishnamoorthy:2007:ICC

[KX07] K. Krishnamoorthy and Yanping Xia. Inferences on correlation coefficients: One-sample, independent and correlated cases. Journal of Statistical Planning and Inference, 137(7):2362–2379, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002564.

Klebanov:2007:NAT

[KY07] Lev B. Klebanov and Andrei Y. Yakovlev. A new approach to testing for sufficient follow-up in cure-rate analysis. Journal of Statistical Planning and Inference, 137(11):3557–3569, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001176.

Karunamuni:2003:EBT

[KZ03a] Rohana J. Karunamuni and Shunpu Zhang. Empirical Bayes two-action problem for the continuous one-parameter exponential family with errors in variables. Journal of Statistical Planning and Inference, 113(2):437–449, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001088.

Karunamuni:2003:OLB

[KZ03b] Rohana J. Karunamuni and Shunpu Zhang. Optimal linear Bayes and empirical Bayes estimation and prediction of the finite population mean. Journal of Statistical Planning and Inference, 113(2):505–525, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201002061.

Klein:2009:MLE

[KZ09] Daniel Klein and Ivan Zezula. The maximum likelihood estimators in the growth curve model with serial covariance structure. Journal of Statistical Planning and Inference, 139(9):3270–3276, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000688.

Li:2008:PEC

[LA08] Yingfu Li and M. L. Aggarwal. Projection estimation capacity of Hadamard designs. Journal of Statistical Planning and Inference, 138(1):154–159, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002054.

Lahiri:2000:BRB

[Lah00] S. Lahiri. Book review: Bootstrap methods: a practitioner’s guide: M. R. Chernick, Wiley, New York, 1999, pp. xiv + 264 (including 84 pages of Bibliography), price $74.95 paperback, ISBN 0-471-34912-7. Journal of Statistical Planning and Inference, 91(1):171–172, November 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001324.

Lahiri:2003:Pa

[Lah03a] P. Lahiri. Preface. Journal of Statistical Planning and Inference, 111(1–2):1, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002811.

Lahiri:2003:Pb

[Lah03b] P. Lahiri. Preface. Journal of Statistical Planning and Inference, 112(1–2):1, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802003208.

Laib:2005:KEM

[Laï05] Naâmane Laïb. Kernel estimates of the mean and the volatility functions in a nonlinear autoregressive model with ARCH errors. Journal of Statistical Planning and Inference, 134(1):116–139, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001776.

Lai:2009:TSR

[Lai09] Dejian Lai. Testing spatial randomness based on empirical distribution function: a study on lattice data. Journal of Statistical Planning and Inference, 139(2):136–142, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001754.

Landsman:2000:MFI

[Lan00] Zinoviy Landsman. On the minimum of the Fisher information about the scale parameter and the singular Sturm–Liouville problem. Journal of Statistical Planning and Inference, 88(1):29–35, July 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837589900213X.

Landsman:2001:SOM

[Lan01] Zinoviy Landsman. Second-order minimax estimation of the mean value for exponential dispersion models. Journal of Statistical Planning and Inference, 98(1–2):57–71, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003153.

Larsen:2003:ESA

[Lar03] Michael D. Larsen. Estimation of small-area proportions using covariates and survey data. Journal of Statistical Planning and Inference, 112(1–2):89–98, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003257.

Lopez-Blazquez:2000:BEV

[LB00] F. López-Blázquez. Bounds for the expected value of spacings from discrete distributions. Journal of Statistical Planning and Inference, 84(1–2):1–9, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001500.

Lien:2006:MPM

[LB06] Donald Lien and N. Balakrishnan. Moments and properties of multiplicatively constrained bivariate lognormal distribution with applications to futures hedging. Journal of Statistical Planning and Inference, 136(4):1349–1359, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003945.

Li:2008:SSN

[LB08a] T. Li and N. Balakrishnan. Some simple nonparametric methods to test for perfect ranking in ranked set sampling. Journal of Statistical Planning and Inference, 138(5):1325–1338, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002649.

Lu:2008:NEL

[LB08b] Xuewen Lu and M. D. Burke. Nonparametric estimation of linear functionals of a bivariate distribution under univariate censoring. Journal of Statistical Planning and Inference, 138(10):3238–3256, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000207.

Lopez-Blazquez:2006:ULB

[LBCM06] Fernando López-Blázquez and Antonia Castaño-Martínez. Upper and lower bounds for the correlation ratio of order statistics from a sample without replacement. Journal of Statistical Planning and Inference, 136(1):43–52, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002800.

Lopez-Blazquez:2000:BAH

[LBM00] F. López-Blázquez and B. Salamanca Miño. Binomial approximation to hypergeometric probabilities. Journal of Statistical Planning and Inference, 87(1):21–29, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001871.

Lopez-Blazquez:2006:BEV

[LBM06] F. López-Blázquez and B. Salamanca Miño. Bounds for the expected value of records from discrete distributions. Journal of Statistical Planning and Inference, 136(2):467–474, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003209.

Lopez-Blazquez:2008:SUB

[LBR08] Fernando López-Blázquez and Tomasz Rychlik. Sharp upper bounds for the expected values of non-extreme order statistics from discrete distributions. Journal of Statistical Planning and Inference, 138(12):3635–3646, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001225.

Loeppky:2006:CNR

[LBS06] J. L. Loeppky, D. Bingham, and R. R. Sitter. Constructing non-regular robust parameter designs. Journal of Statistical Planning and Inference, 136(10):3710–3729, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000789.

Liu:2000:NMI

[LC00] Jeng-Fu Liu and Kuang-Chao Chang. A note on multiple inverse sampling. Journal of Statistical Planning and Inference, 87(2):347–352, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001962.

Lai:2001:TSE

[LC01a] Chia-Jui Lai and C. K. Chu. Triple smoothing estimation of the regression function and its derivatives in nonparametric regression. Journal of Statistical Planning and Inference, 98(1–2):157–175, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003190.

Li:2001:ORD

[LC01b] Kim-Hung Li and Nai N. Chan. L_{p}-optimality for regression designs with heteroscedastic errors. Journal of Statistical Planning and Inference, 92(1–2):253–257, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001439.

Lopes:2007:FSV

[LC07] Hedibert Freitas Lopes and Carlos Marinho Carvalho. Factor stochastic volatility with time varying loadings and Markov switching regimes. Journal of Statistical Planning and Inference, 137(10):3082–3091, October 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000766.

Liang:2009:MOR

[LC09a] Yuanyuan Liang and Keumhee Chough Carrière. Multipleobjective response-adaptive repeated measurement designs for clinical trials. Journal of Statistical Planning and Inference, 139(3):1134–1145, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002954.

Lu:2009:ESC

[LC09b] Xiaojing Lu and John T. Chen. Exact simultaneous confidence segments for all contrast comparisons. Journal of Statistical Planning and Inference, 139(8):2816–2822, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900007X.

Lahiri:2007:NAH

[LCM07] S. N. Lahiri, A. Chatterjee, and T. Maiti. Normal approximation to the hypergeometric distribution in nonstandard cases and a sub-Gaussian Berry–Esseen theorem. Journal of Statistical Planning and Inference, 137(11):3570–3590, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001188.

Cheng:2006:MHD

[lCV06] An lin Cheng and Anand N. Vidyashankar. Minimum Hellinger distance estimation for randomized play the winner design. Journal of Statistical Planning and Inference, 136 (6):1875–1910, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805001916.

Lu:2006:QLE

[LCZ06] Jye-Chyi Lu, Di Chen, and Weixing Zhou. Quasi-likelihood estimation for GLM with random scales. Journal of Statistical Planning and Inference, 136(2):401–429, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002964.

Lian:2006:SAU

[LdC06] Zilong Lian and Enrique del Castillo. Setup adjustment under unknown process parameters and fixed adjustment cost. Journal of Statistical Planning and Inference, 136(3):1039–1060, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003556.

Lachos:2009:RLM

[LDC09] Victor H. Lachos, Dipak K. Dey, and Vicente G. Cancho. Robust linear mixed models with skew-normal independent distributions from a Bayesian perspective. Journal of Statistical Planning and Inference, 139(12):4098–4110, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001669.

Li:2004:DCB

[LDT04] Yingfu Li, Lih-Yuan Deng, and Boxin Tang. Design catalog based on minimum G-aberration. Journal of Statistical Planning and Inference, 124(1):219–230, August 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001897.

Lee:2001:BRA

[Lee01a] Carl Lee. Book review: Applied regression including computing and graphics: R. Dennis Cook and Sanford Weisberg; Wiley, New York, 1999, 593pp, US$89.95, ISBN 0-471-31711-X. Journal of Statistical Planning and Inference, 92(1–2):301–303, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001531.

Lee:2001:PSB

[Lee01b] Young Jack Lee. Post-stratification based direct adjustment approach to a missing data problem in clinical trials. Journal of Statistical Planning and Inference, 96(1):247–262, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003384.

Leeb:2005:DLP

[Lee05] Hannes Leeb. The distribution of a linear predictor after model selection: conditional finite-sample distributions and asymptotic approximations. Journal of Statistical Planning and Inference, 134(1):64–89, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002071.

Lehmann:2009:JMD

[Leh09] Axel Lehmann. Joint modeling of degradation and failure time data. Journal of Statistical Planning and Inference, 139(5):1693–1706, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002450.

Lengyel:2002:RHT

[Len02] Tamás Lengyel. On reaching head-to-tail ratios for balanced and unbalanced coins. Journal of Statistical Planning and Inference, 101(1–2):203–210, February 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001793.

Leng:2009:SAV

[Len09] Chenlei Leng. A simple approach for varying-coefficient model selection. Journal of Statistical Planning and Inference, 139(7):2138–2146, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800390X.

Letac:2000:SSS

[Let00] Gérard Letac. Stationary sequences with simple joint Poisson distributions. Journal of Statistical Planning and Inference, 90(1):1–20, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001063.

Letue:2008:SSM

[Let08] Frédérique Letué. A semiparametric shock model for a pair of event-related dependent censored failure times. Journal of Statistical Planning and Inference, 138(12):3869–3884, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000323.

Lecoutre:2007:NNC

[LF07] Bruno Lecoutre and Sébastien Faure. A note on new confidence intervals for the difference between two proportions based on an Edgeworth expansion. Journal of Statistical Planning and Inference, 137(1):355–356, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000243.

Lopez-Fidalgo:2007:MOS

[LFRL07] J. López-Fidalgo and M. J. Rivas-López. MV-optimality standardized through the coefficient of variation. Journal of Statistical Planning and Inference, 137(8):2666–2680, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000183.

Lopez-Fidalgo:2008:ODD

[LFTT08] Jesús López-Fidalgo, Chiara Tommasi, and Paula Camelia Trandafir. Optimal designs for discriminating between some extensions of the Michaelis–Menten model. Journal of Statistical Planning and Inference, 138(12):3797–3804, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001389.

Li:2006:ORC

[LG06] Jianjun Li and Shanti S. Gupta. Optimal rate of convergence of monotone empirical Bayes tests for normal means. Journal of Statistical Planning and Inference, 136(7):2352– 2366, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002272.

Li:2005:CRE

[LGL05] Jianjun Li, Shanti S. Gupta, and Friedrich Liese. Convergence rates of empirical Bayes estimation in exponential family. Journal of Statistical Planning and Inference, 131(1):101– 115, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000643.

Lombardia:2003:BCD

[LGMPS03] M. J. Lombardía, W. González-Manteiga, and J. M. Prada-Sánchez. Bootstrapping the Chambers–Dunstan estimate of a finite population distribution function. Journal of Statistical Planning and Inference, 116(2):367–388, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002409.

Lantieri:2009:LMU

[LGV09] P. Lantiéri, F. Guérin, and S. Voiculescu. LSE method using the CHSS model. Journal of Statistical Planning and Inference, 139(5):1809–1820, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002553.

Lee:2000:EPT

[LH00] Jack C. Lee and R. C. Hwang. On estimation and prediction for temporally correlated longitudinal data. Journal of Statistical Planning and Inference, 87(1):87–104, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001846.

Lin:2008:CLD

[LH08] Gwo Dong Lin and Chin-Yuan Hu. On characterizations of the logistic distribution. Journal of Statistical Planning and Inference, 138(4):1147–1156, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002789.

Lavine:2009:BRD

[LHL09] Michael Lavine, Gabriele C. Hegerl, and Susan Lozier. Book review: Discussion of reified Bayesian modelling and inference for physical systems by Michael Goldstein and Jonathan Rougier. Journal of Statistical Planning and Inference, 139(3):1243–1245, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003327.

Lin:2001:AAE

[LHOP01] C. Thomas Lin, Balakrishna S. Hosmane, Peggy J. Olson, and R. J. Robert J. Padley. Analysis of adverse events in titration studies. Journal of Statistical Planning and Inference, 96(1):129–142, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003293.

Liang:2000:BAP

[LHS00] Hua Liang, Wolfgang Härdle, and Volker Sommerfeld. Bootstrap approximation in a partially linear regression model. Journal of Statistical Planning and Inference, 91(2):413–426, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001919.

Lu:2003:SPC

[LHZ03] Xuan Lu, Wenbiao Hu, and Yan Zheng. A systematical procedure in the construction of multi-level supersaturated design. Journal of Statistical Planning and Inference, 115(1):287–310, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001167.

Li:2001:SRS

[Li01] Bing Li. Sensitivity of Rao’s score test, the Wald test and the likelihood ratio test to nuisance parameters. Journal of Statistical Planning and Inference, 97(1):57–66, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003451.

Li:2003:NLW

[Li03] Linyuan Li. Non-linear wavelet-based density estimators under random censorship. Journal of Statistical Planning and Inference, 117(1):35–58, November 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200366X.

Li:2007:MOW

[Li07] Linyuan Li. On the minimax optimality of wavelet estimators with censored data. Journal of Statistical Planning and Inference, 137(4):1138–1150, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000607.

Li:2008:TSR

[Li08] Jianjun (David) Li. A two-step rejection procedure for testing multiple hypotheses. Journal of Statistical Planning and Inference, 138(6):1521–1527, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002960.

Liang:2000:ANP

[Lia00a] Hua Liang. Asymptotic normality of parametric part in partially linear models with measurement error in the nonparametric part. Journal of Statistical Planning and Inference, 86(1):51–62, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000932.

Liang:2000:CEM

[Lia00b] TaChen Liang. Counter-examples to “minimax lower bounds of empirical Bayes tests”. Journal of Statistical Planning and Inference, 87(2):273–286, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001883.

Liang:2000:EBT

[Lia00c] TaChen Liang. On empirical Bayes tests in a positive exponential family. Journal of Statistical Planning and Inference, 83(1):169–181, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000518.

Liao:2003:FLM

[Lia03] Jason J. Z. Liao. A functional linear model for comparing two pharmacokinetics profiles. Journal of Statistical Planning and Inference, 116(2):521–535, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003543.

Liang:2006:EBT

[Lia06a] TaChen Liang. Empirical Bayes testing for a normal mean: variance unknown case. Journal of Statistical Planning and Inference, 136(4):1376–1393, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003957.

Liang:2006:SIV

[Lia06b] TaChen Liang. Simultaneous inspection of variable equivalence for finite populations. Journal of Statistical Planning and Inference, 136(7):2112–2128, July 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002132.

Liao:2006:TLF

[Lia06c] Chen-Tuo Liao. Two-level factorial designs for searching dispersion factors and estimating location main effects. Journal of Statistical Planning and Inference, 136(11):4071–4087, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580500131X.

Liang:2009:EBE

[Lia09] TaChen Liang. Empirical Bayes estimation of θb in positive exponential families. Journal of Statistical Planning and Inference, 139(2):411–424, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001948.

Liese:2006:SPS

[Lie06] Friedrich Liese. Selection procedures for sparse data. Journal of Statistical Planning and Inference, 136(7):2035–2052, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805002090.

Lijoi:2003:APF

[Lij03] Antonio Lijoi. Approximating priors by finite mixtures of conjugate distributions for an exponential family. Journal of Statistical Planning and Inference, 113(2):419–435, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520100205X.

Lindley:2001:WSC

[Lin01a] Dennis V. Lindley. Why should clinicians care about Bayesian methods? Journal of Statistical Planning and Inference, 94(1):59–60, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002330.

Linton:2001:BRN

[Lin01b] O. B. Linton. Book review: Nonparametric econometrics: Adrian Pagan and Aman Ullah; Cambridge University Press, Cambridge, 1999, price US$74.95 hardback, ISBN 0-521-35564-8, price US$29.95 paperback, ISBN 0-521-58611-9. Journal of Statistical Planning and Inference, 92(1–2):299–300, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800000707.

Lin:2007:ADK

[Lin07] Yan-Xia Lin. An alternative derivation of the Kalman filter using the quasi-likelihood method. Journal of Statistical Planning and Inference, 137(5):1627–1633, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002412.

Lin:2008:EQL

[Lin08] Pei-Sheng Lin. Efficiency of quasi-likelihood estimation for spatially correlated binary data on L_{p} spaces. Journal of Statistical Planning and Inference, 138(6):1528–1541, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002984.

Liu:2000:MDE

[Liu00a] Jinshan Liu. MSEM dominance of estimators in two seemingly unrelated regressions. Journal of Statistical Planning and Inference, 88(2):255–266, August 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000823.

Liu:2000:ECB

[Liu00b] Shuangzhe Liu. Efficiency comparisons between the OLSE and the BLUE in a singular linear model. Journal of Statistical Planning and Inference, 84(1–2):191–200, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001494.

Liu:2004:SSC

[Liu04] Wei Liu. Some simple corrected confidence intervals following a sequential test. Journal of Statistical Planning and Inference, 126(1):189–205, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002702.

Liu:2009:EOL

[Liu09] Yonghui Liu. On equality of ordinary least squares estimator, best linear unbiased estimator and best linear unbiased predictor in the general linear model. Journal of Statistical Planning and Inference, 139(4):1522–1529, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003546.

Li:2007:OFP

[LJ07a] Fei Li and Mike Jacroux. Optimal foldover plans for blocked 2m-k fractional factorial designs. Journal of Statistical Planning and Inference, 137(7):2439–2452, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002837.

Liang:2007:LBR

[LJ07b] Han-Ying Liang and Bing-Yi Jing. The LIL for the Bickel–Rosenblatt test statistic. Journal of Statistical Planning and Inference, 137(6):1829–1837, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001819.

Liang:2009:ANP

[LJ09] Han-Ying Liang and Bing-Yi Jing. Asymptotic normality in partial linear models based on dependent errors. Journal of Statistical Planning and Inference, 139(4):1357–1371, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003388.

Li:2007:SPL

[LJC07] Ruojia Li, Richard A. Johnson, and Sang-Hoon Cho. Some properties of a linear combination CUSUM statistic for controlling a multivariate mean vector. Journal of Statistical Planning and Inference, 137(11):3591–3600, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700119X.

Liu:2007:SNM

[LJZ^{+}07] Wei Liu, Mortaza Jamshidian, Ying Zhang, Frank Bretz, and Xiaoliang Han. Some new methods for the comparison of two linear regression models. Journal of Statistical Planning and Inference, 137(1):57–67, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002806.

Lee:2000:QDG

[LK00] Juneyoung Lee and André I. Khuri. Quantile dispersion graphs for the comparison of designs for a random two-way model. Journal of Statistical Planning and Inference, 91(1):123–137, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580000135X.

Liu:2003:MDD

[LK03] Wei Liu and Koon-Shing Kwong. A multiple directional decision procedure for successive comparisons of treatment effects. Journal of Statistical Planning and Inference, 116(1):49–59, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002379.

Liu:2008:EPT

[LK08] Yuanyuan Liu and Qingde Kang. The existence of perfect T (K_{1,2p})-triple systems. Journal of Statistical Planning and Inference, 138(12):4111–4131, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001171.

Lee:2004:GCI

[LL04] Jack C. Lee and Shu-Hui Lin. Generalized confidence intervals for the ratio of means of two normal populations. Journal of Statistical Planning and Inference, 123(1):49–60, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001411.

Lai:2005:OSB

[LL05a] Tze Leung Lai and Tiong Wee Lim. Optimal stopping for Brownian motion with applications to sequential analysis and option pricing. Journal of Statistical Planning and Inference, 130(1–2):21–47, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001247.

Lin:2005:GIC

[LL05b] Shu-Hui Lin and Jack C. Lee. Generalized inferences on the common mean of several normal populations. Journal of Statistical Planning and Inference, 134(2):568–582, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002137.

Langseth:2006:CRR

[LL06a] Helge Langseth and Bo Henry Lindqvist. Competing risks for repairable systems: a data study. Journal of Statistical Planning and Inference, 136(5):1687–1700, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001758.

Liseo:2006:NRP

[LL06b] Brunero Liseo and Nicola Loperfido. A note on reference priors for the scalar skew-normal distribution. Journal of Statistical Planning and Inference, 136(2):373–389, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003052.

Levin:2007:CTP

[LL07a] Bruce Levin and Cheng-Shiun Leu. A comparison of two procedures to select the best binomial population with sequential elimination of inferior populations. Journal of Statistical Planning and Inference, 137(1):245–263, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580500306X.

Li:2007:PCS

[LL07b] Xiaohu Li and Zhouping Li. Proof of a conjecture on Spearman’s ρ and Kendall’s τ for sample minimum and maximum. Journal of Statistical Planning and Inference, 137(1):359–361, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002636.

Lin:2007:BAH

[LL07c] Tsung I. Lin and Jack C. Lee. Bayesian analysis of hierarchical linear mixed modeling using the multivariate t distribution. Journal of Statistical Planning and Inference, 137(2):484–495, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000358.

Lin:2008:PIG

[LL08] T. Y. Lin and C. T. Liao. Prediction intervals for general balanced linear random models. Journal of Statistical Planning and Inference, 138(10):3164–3175, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000141.

Lahiri:2009:NAS

[LL09] P. Lahiri and Yan Li. A new alternative to the standard F test for clustered data. Journal of Statistical Planning and Inference, 139(10):3430–3441, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000883.

Lahiri:2002:ADA

[LLC02] Soumendra Nath Lahiri, Yoondong Lee, and Noel Cressie. On asymptotic distribution and asymptotic efficiency of least squares estimators of spatial variogram parameters. Journal of Statistical Planning and Inference, 103(1–2):65–85, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001987.

Lin:2009:CEB

[LLC09] Pei-Sheng Lin, Hui-Yun Lee, and Murray Clayton. A comparison of efficiencies between quasi-likelihood and pseudolikelihood estimates in non-separable spatial-temporal binary data. Journal of Statistical Planning and Inference, 139(9):3310–3318, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000767.

Liang:2000:TMB

[LLFF00] Jiajuan Liang, Runze Li, Hongbin Fang, and Kai-Tai Fang. Testing multinormality based on low-dimensional projection. Journal of Statistical Planning and Inference, 86(1):129–141, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899001688.

Li:2009:NSO

[LLL09] Xiaohu Li, Xiaoliang Ling, and Ping Li. A new stochastic order based upon Laplace transform with applications. Journal of Statistical Planning and Inference, 139(8):2624–2630, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004485.

Lee:2005:BAB

[LLLH05] Jack C. Lee, Tsung I. Lin, Kuo J. Lee, and Ying L. Hsu. Bayesian analysis of Box–Cox transformed linear mixed models with ARMA(p, q) dependence. Journal of Statistical Planning and Inference, 133(2):435–451, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001235.

Lloyd:2005:CAC

[Llo05] Chris J. Lloyd. On comparing the accuracy of competing tests of the same hypotheses from simulation data. Journal of Statistical Planning and Inference, 128(2):497–508, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000229.

Liu:2002:MMM

[LLP02] Lin Liu, Chu-In Charles Lee, and Jianan Peng. Maxmin multiple comparison procedure for isotonic doseresponse curves. Journal of Statistical Planning and Inference, 107(1–2):133–141, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002483.

Liu:2009:MTW

[LLP09] Shuangzhe Liu, Changyu Lu, and Simo Puntanen. Matrix trace Wielandt inequalities with statistical applications. Journal of Statistical Planning and Inference, 139(7):2254–2260, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004692.

Ling:2000:X

[LLvRZ00] C. H. A. Ling, Y. Lu, G. H. J. van Rees, and L. Zhu. V (m, t)’s for m = 4, 5, 6. Journal of Statistical Planning and Inference, 86(2):515–525, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001275.

Lin:2007:GIC

[LLW07] S. H. Lin, Jack C. Lee, and R. S. Wang. Generalized inferences on the common mean vector of several multivariate normal populations. Journal of Statistical Planning and Inference, 137(7):2240–2249, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002205.

Liu:2007:CML

[LLZ07] Yukun Liu, Minqian Liu, and Runchu Zhang. Construction of multi-level supersaturated design via Kronecker product. Journal of Statistical Planning and Inference, 137(9):2984–2992, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000390.

Lillo:2000:BAE

[LM00a] Rosa E. Lillo and Miguel Martiín. Bayesian approach to estimation of ordered uniform scale parameters. Journal of Statistical Planning and Inference, 87(1):105–118, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589900186X.

Lu:2000:NMC

[LM00b] Xuan Lu and Yun Meng. A new method in the construction of two-level supersaturated designs. Journal of Statistical Planning and Inference, 86(1):229–238, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900169X.

Lutsenko:2003:MCI

[LM03] M. M. Lutsenko and S. G. Maloshevskii. Minimax confidence intervals for the binomial parameter. Journal of Statistical Planning and Inference, 113(1):67–77, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002981.

Lee:2004:TRM

[LM04] Yi-Tzu Lee and Thomas Mathew. Tolerance regions in multivariate linear regression. Journal of Statistical Planning and Inference, 126(1):253–271, November 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002295.

Li:2008:ODL

[LM08] Gang Li and Dibyen Majumdar. D-optimal designs for logistic models with three and four parameters. Journal of Statistical Planning and Inference, 138(7):1950–1959, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003308.

Lai:2001:BML

[LMD^{+}01] Dejian Lai, Lemuel A. Moyé, Barry R. Davis, Lisa E. Brown, and Frank M. Sacks. Brownian motion and longterm clinical trial recruitment. Journal of Statistical Planning and Inference, 93(1–2):239–246, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002032.

Liu:2000:SCI

[LMH00] Wei Liu, Tetsuhisa Miwa, and A. J. Hayter. Simultaneous confidence interval estimation for successive comparisons of ordered treatment effects. Journal of Statistical Planning and Inference, 88(1):75–86, July 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002098.

Lam:2001:CRC

[LMM01] Clement Lam, Ying Miao, and Miwako Mishima. Cyclically resolvable cyclic Steiner 2-systems S(2, 4, 52). Journal of Statistical Planning and Inference, 95(1–2):245–256, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002925.

Lin:2008:FNO

[LMS08] C. Devon Lin, A. Miller, and R. R. Sitter. Folded over non-orthogonal designs. Journal of Statistical Planning and Inference, 138(10):3107–3124, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004247.

Linder:2003:ICS

[LN03] R. S. Linder and H. N. Nagaraja. Impact of censoring on sample variances in a bivariate normal model. Journal of Statistical Planning and Inference, 114(1–2):145–160, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004688.

Larocque:2008:OSL

[LNO08] Denis Larocque, Jaakko Nevalainen, and Hannu Oja. Onesample location tests for multilevel data. Journal of Statistical Planning and Inference, 138(8):2469–2482, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003783.

Luoma:2007:ODR

[LNS07] A. Luoma, T. Nummi, and Bikas K. Sinha. Optimal designs in random coefficient cubic regression models. Journal of Statistical Planning and Inference, 137(11):3611–3617, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001218.

Longford:2009:ILD

[Lon09] Nicholas T. Longford. Inference with the lognormal distribution. Journal of Statistical Planning and Inference, 139(7):2329–2340, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808004199.

Lopez:2009:SIR

[Lop09] Olivier Lopez. Single-index regression models with rightcensored responses. Journal of Statistical Planning and Inference, 139(3):1082–1097, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002917.

Leipus:2006:OSD

[LOPV06] Remigijus Leipus, George Oppenheim, Anne Philippe, and Marie-Claude Viano. Orthogonal series density estimation in a disaggregation scheme. Journal of Statistical Planning and Inference, 136(8):2547–2571, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004215.

Louani:2000:LDD

[Lou00] Djamal Louani. Large deviations for the L_{1}-distance in kernel density estimation. Journal of Statistical Planning and Inference, 90(2):177–182, October 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001245.

Lai:2008:TWA

[LOV08] Tze Leung Lai, Ingram Olkin, and Raja Velu. Theodore Wilbur Anderson, Jr. Journal of Statistical Planning and Inference, 138(9):2593–2594, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000517.

Loynes:2001:NML

[Loy01] R. M. Loynes. A new measure in local influence. Journal of Statistical Planning and Inference, 92(1–2):47–53, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800001440.

Levin:2001:UEB

[LP01] B. Levin and Myunghee Cho Paik. The unreasonable effectiveness of a biased logistic regression procedure in the analysis of pair-matched case-control studies. Journal of Statistical Planning and Inference, 96(2):371–385, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002172.

Ling:2004:HET

[LP04] Shiqing Ling and Liang Peng. Hill’s estimator for the tail index of an ARMA model. Journal of Statistical Planning and Inference, 123(2):279–293, July 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001514.

Liu:2006:DIT

[LP06] Qing Liu and Gordon Pledger. On design and inference for two-stage adaptive clinical trials with dependent data. Journal of Statistical Planning and Inference, 136(6):1962–1984, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001965.

Li:2009:DBR

[LP09] Deyuan Li and Liang Peng. Does bias reduction with external estimator of second order parameter work for endpoint? Journal of Statistical Planning and Inference, 139(6):1937–1952, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003637.

Li:2003:REM

[LQH03] Ning Li, Guoqi Qian, and Richard Huggins. A random effects model for diseases with heterogeneous rates of infection. Journal of Statistical Planning and Inference, 116(1):317–332, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200232X.

Lindner:2002:MFB

[LR02] C. C. Lindner and A. Rosa. The metamorphosis of λ-fold block designs with block size four into λ-fold triple systems. Journal of Statistical Planning and Inference, 106(1–2):69– 76, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002033.

Leiva:2009:CRT

[LR09] Ricardo Leiva and Anuradha Roy. Classification rules for triply multivariate data with an AR(1) correlation structure on the repeated measures over time. Journal of Statistical Planning and Inference, 139(8):2598–2613, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004461.

Liu:2007:CAE

[LRD07] Yufeng Liu, Shiling Ruan, and Angela M. Dean. Construction and analysis of E(s^{2}) efficient supersaturated designs. Journal of Statistical Planning and Inference, 137(5):1516– 1529, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002254.

Levin:2002:MAE

[LRZ02] Bruce Levin, Herbert Robbins, and Cun-Hui Zhang. Mathematical aspects of estimating two treatment effects and a common variance in an assured allocation design. Journal of Statistical Planning and Inference, 108(1–2):255–262, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003129.

Liu:2008:QPP

[LRZ08] Xu-Qing Liu, Jian-Ying Rong, and Yu-Ming Zhuang. On quadratic prediction problems in finite populations. Journal of Statistical Planning and Inference, 138(10):3257–3268, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800027X.

Lu:2001:TBM

[LS01] Yili Lu and Barry Storer. A tampered Brownian motion process model for partial step-stress accelerated life testing. Journal of Statistical Planning and Inference, 94(1):15–24, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002159.

Lin:2002:SLT

[LS02] Win-Chin Lin and John Stufken. Strongly linear trend-free block designs and 1-factors of representative graphs. Journal of Statistical Planning and Inference, 106(1–2):375–386, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002239.

Liu:2004:ALR

[LS04] Xin Liu and Yongzhao Shao. Asymptotics for the likelihood ratio test in a two-component normal mixture model. Journal of Statistical Planning and Inference, 123(1):61–81, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001381.

Liefvendahl:2006:SAO

[LS06] M. Liefvendahl and R. Stocki. A study on algorithms for optimization of Latin hypercubes. Journal of Statistical Planning and Inference, 136(9):3231–3247, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000686.

Li:2007:GFU

[LS07] Xiaohu Li and Moshe Shaked. A general family of univariate stochastic orders. Journal of Statistical Planning and Inference, 137(11):3601–3610, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001206.

Lin:2008:ICT

[LS08] C. Devon Lin and R. R. Sitter. An isomorphism check for two-level fractional factorial designs. Journal of Statistical Planning and Inference, 138(4):1085–1101, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002315.

Lu:2001:KSS

[LSD01] Xuewen Lu, R. S. Singh, and A. F. Desmond. A kernel smoothed semiparametric survival model. Journal of Statistical Planning and Inference, 98(1–2):119–135, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003141.

Lindqvist:2006:MDB

[LSL06] Bo H. Lindqvist, Bård Støve, and Helge Langseth. Modelling of dependence between critical failure and preventive maintenance: the repair alert model. Journal of Statistical Planning and Inference, 136(5):1701–1717, May 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580500176X.

Lai:2008:SID

[LSL08] Tze Leung Lai, Dylan S. Small, and Jia Liu. Statistical inference in dynamic panel data models. Journal of Statistical Planning and Inference, 138(9):2763–2776, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000657.

Lee:2006:FWC

[LSW06] Sangyeol Lee, T. N. Sriram, and Xinyu Wei. Fixedwidth confidence interval based on a minimum Hellinger distance estimator. Journal of Statistical Planning and Inference, 136(12):4276–4292, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002004.

Lam:1996:CAR

[LT96] Clement Lam and Vladimir D. Tonchev. Classification of affine resolvable 2 − (27, 9, 4) designs. Journal of Statistical Planning and Inference, 56(2):187–202, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000183. See corrigendum [LT00].

Lam:2000:CCA

[LT00] Clement Lam and Vladimir D. Tonchev. Corrigendum to “Classification of affine resolvable 2−(27, 9, 4) designs” [Journal of Statistical Planning and Inference 56 (1996) 187–202]. Journal of Statistical Planning and Inference, 86(1):277–278, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899000555.

Lele:2002:CLA

[LT02] Subhash Lele and Mark L. Taper. A composite likelihood approach to (co)variance components estimation. Journal of Statistical Planning and Inference, 103(1–2):117–135, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002154.

Lewis:2004:LFR

[LT04] Charles Lewis and Dorothy T. Thayer. A loss function related to the FDR for random effects multiple comparisons. Journal of Statistical Planning and Inference, 125(1–2):49– 58, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000552.

Leonenko:2006:WCF

[LT06] Nikolai N. Leonenko and Emanuele Taufer. Weak convergence of functionals of stationary long memory processes to Rosenblatt-type distributions. Journal of Statistical Planning and Inference, 136(4):1220–1236, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003921.

Li:2008:CRM

[LT08a] Lexin Li and Chih-Ling Tsai. Constrained regression model selection. Journal of Statistical Planning and Inference, 138(12):3939–3949, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000463.

Lindner:2008:MFD

[LT08b] C. C. Lindner and Antoinette Tripodi. The metamorphosis of λ-fold K_{4} − e designs into maximum packings of λK_{n} with 4-cycles, λ ≥ 2. Journal of Statistical Planning and Inference, 138(11):3316–3325, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000797.

Lokhnygina:2008:OTS

[LT08c] Yuliya Lokhnygina and Anastasios A. Tsiatis. Optimal twostage group-sequential designs. Journal of Statistical Planning and Inference, 138(2):489–499, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700256X.

Li:2009:NEC

[LT09a] Jiexiang Li and Lanh Tat Tran. Nonparametric estimation of conditional expectation. Journal of Statistical Planning and Inference, 139(2):164–175, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001821.

Luo:2009:ENE

[LT09b] Xiaodong Luo and Wei-Yann Tsai. Erratum to “Nonparametric estimation of bivariate distribution under right truncation with application to panic disorder” [J. Statist. Plann. Inference 139 (4) (2009) 1559–1568]. Journal of Statistical Planning and Inference, 139(6):2116, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000020. See [LT09c].

Luo:2009:NEB

[LT09c] Xiaodong Luo and Wei-Yann Tsai. Nonparametric estimation of bivariate distribution under right truncation with application to panic disorder. Journal of Statistical Planning and Inference, 139(4):1559–1568, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003595. See erratum [LT09b].

Liao:2000:OBM

[LTI00] C. T. Liao, C. H. Taylor, and H. K. Iyer. Optimal balanced measurement designs when errors are correlated. Journal of Statistical Planning and Inference, 84(1–2):295–321, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837589900110X.

Li:2008:SAN

[LTW08] Hui-Qiong Li, Nian-Sheng Tang, and Liu-Cang Wu. Statistical analysis of non-inferiority via non-zero risk difference in stratified matched-pair studies. Journal of Statistical Planning and Inference, 138(12):4055–4067, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000505.

Lu:2002:BRA

[Lu02] I-Li Lu. Book review: Analysis of Health Surveys: Edward L. Korn and Barry I. Graubard; Wiley, New York, 1999, 382pp. Price $94.50, ISBN 0-471-13773-1. Journal of Statistical Planning and Inference, 100(1):89, January 1, 2002.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001070.

Lu:2008:GPL

[Lu08] Yiqiang Lu. Generalized partially linear varying-coefficient models. Journal of Statistical Planning and Inference, 138 (4):901–914, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807001413.

Lucas:2009:DRS

[Luc09] James M. Lucas. Discussion of “Response surface design evaluation and comparison” by Christine Anderson-Cook, Connie Borror and Douglas Montgomery. Journal of Statistical Planning and Inference, 139(2):660–661, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800164X. See [ACBM09b].

Lue:2009:SIR

[Lue09] Heng-Hui Lue. Sliced inverse regression for multivariate response regression. Journal of Statistical Planning and Inference, 139(8):2656–2664, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004527.

LuValle:2009:ECA

[LuV09] M. J. LuValle. Extrapolation and confounding in accelerated material degradation and failure studies. Journal of Statistical Planning and Inference, 139(5):1707–1721, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002462.

Liu:2007:NHT

[LW07] Wei Liu and Nan Wang. A note on Hall’s triple sampling procedure: a multiple sample second order sequential analogue of the Behrens–Fisher problem. Journal of Statistical Planning and Inference, 137(1):331–343, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002922.

Lin:2009:RAJ

[LW09] Tsung-I. Lin and Yun-Jen Wang. A robust approach to joint modeling of mean and scale covariance for longitudinal data. Journal of Statistical Planning and Inference, 139(9):3013–3026, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000342.

Lin:2009:PLT

[LWB09] Chien-Tai Lin, Sam J. S. Wu, and N. Balakrishnan. Planning life tests with progressively Type-I interval censored data from the lognormal distribution. Journal of Statistical Planning and Inference, 139(1):54–61, January 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002292.

Lee:2009:CCS

[LWL^{+}09] Mei-Ling Ting Lee, G. A. Whitmore, Francine Laden, Jaime E. Hart, and Eric Garshick. A case-control study relating railroad worker mortality to diesel exhaust exposure using a threshold regression model. Journal of Statistical Planning and Inference, 139(5):1633–1642, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002395.

Lai:2007:NFS

[LX07a] Tze Leung Lai and Haipeng Xing. Nonparametric functionals of spectral distributions and their applications to time series analysis. Journal of Statistical Planning and Inference, 137(3):1066–1075, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001613.

Li:2007:MOB

[LX07b] Linyuan Li and Yimin Xiao. On the minimax optimality of block thresholded wavelet estimators with long memory data. Journal of Statistical Planning and Inference, 137(9):2850–2869, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002898.

Lee:2002:CST

[LY02] Chu-In Charles Lee and Xiaosong Yan. Chi-squared tests for and against uniform stochastic ordering on multinomial parameters. Journal of Statistical Planning and Inference, 107(1–2):267–280, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002574.

Liu:2005:PTR

[LY05] Liping Liu and Paul S. F. Yip. Proportional trapping-removal models with contaminated data. Journal of Statistical Planning and Inference, 127(1–2):131–142, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300315X.

Li:2009:INC

[LY09] Qizhai Li and Kai Yu. Inference of non-centrality parameter of a truncated non-central chi-squared distribution. Journal of Statistical Planning and Inference, 139(7):2431–2444, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004357.

Liu:2000:COS

[LZ00] Minqian Liu and Runchu Zhang. Construction of E(s^{2}) optimal supersaturated designs using cyclic BIBDs. Journal of Statistical Planning and Inference, 91(1):139–150, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800001361.

Lu:2005:TDS

[LZ05] Nelson Lu and Dale L. Zimmerman. Testing for directional symmetry in spatial dependence using the periodogram. Journal of Statistical Planning and Inference, 129(1–2):369–385, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002009.

Liu:2007:PPG

[LZ07] Lei Liu and Yu Zhu. Partially projected gradient algorithms for computing nonparametric maximum likelihood estimates of mixing distributions. Journal of Statistical Planning and Inference, 137(7):2509–2522, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002801.

Li:2009:NSE

[LZ09a] Jialiang Li and Xiao-Hua Zhou. Nonparametric and semiparametric estimation of the three way receiver operating characteristic surface. Journal of Statistical Planning and Inference, 139(12):4133–4142, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001694.

Luo:2009:CZG

[LZ09b] Xiaohui Luo and Peng-Liang Zhao. Comments on “Z. Govindarajulu (2005) robustness of sample size re-estimation with interim binary data for double-blind clinical trials”. Journal of Statistical Planning and Inference, 139(2):675–676, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001973.

Liu:2006:OSP

[LZN06] Wen An Liu, Qi Min Zhang, and Zan Kan Nie. Optimal search procedure on coin-weighing problem. Journal of Statistical Planning and Inference, 136(12):4419–4435, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002417.

M:2004:OSD

[M.04] Mohammad Salehi M. Optimal sampling design under a spatial correlation model. Journal of Statistical Planning and Inference, 118(1–2):9–18, January 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004032.

Ma:2000:COL

[Ma00] Chunsheng Ma. Convex orders for linear combinations of random variables. Journal of Statistical Planning and Inference, 84(1–2):11–25, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001433.

Ma:2003:FST

[Ma03] Chunsheng Ma. Families of spatio-temporal stationary covariance models. Journal of Statistical Planning and Inference, 116(2):489–501, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003531.

Mugdadi:2005:MID

[MA05] A. R. Mugdadi and I. A. Ahmad. Moment inequalities derived from comparing life with its equilibrium form. Journal of Statistical Planning and Inference, 134(2):303–317, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002253.

Mitra:2008:SAT

[MA08] Murari Mitra and M. Z. Anis. An L-statistic approach to a test of exponentiality against IFR alternatives. Journal of Statistical Planning and Inference, 138(10):3144–3148, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807004272.

Ma:2009:CNG

[Ma09] Chunsheng Ma. Construction of non-Gaussian random fields with any given correlation structure. Journal of Statistical Planning and Inference, 139(3):780–787, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002048.

Maesono:2005:HOC

[Mae05] Yoshihiko Maesono. Higher-order comparisons of asymptotic confidence intervals. Journal of Statistical Planning and Inference, 133(2):359–379, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400120X.

Maiti:2001:RGL

[Mai01] Tapabrata Maiti. Robust generalized linear mixed models for small area estimation. Journal of Statistical Planning and Inference, 98(1–2):225–238, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003025.

Majumdar:2007:UPC

[Maj07] Suman Majumdar. Uniform L_{1} posterior consistency in compact Gaussian shift experiments. Journal of Statistical Planning and Inference, 137(7):2102–2114, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002023.

Malyutov:2002:LES

[Mal02] Mikhail B. Malyutov. Log-efficient search for significant inputs of linear model. Journal of Statistical Planning and Inference, 100(2):269–280, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001392.

Miao:2009:MDL

[MaL09] Yu Miao and Wen an Liu. Moderate deviations for LS estimator in simple linear EV regression model. Journal of Statistical Planning and Inference, 139(9):3122–3131, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000822.

Manski:2002:TCU

[Man02] Charles F. Manski. Treatment choice under ambiguity induced by inferential problems. Journal of Statistical Planning and Inference, 105(1):67–82, June 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100204X.

Martikainen:2005:SLL

[Mar05] Alexander Martikainen. On the strong law of large numbers for sums over increasing runs. Journal of Statistical Planning and Inference, 132(1–2):75–92, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001430.

Maruyama:2007:AEN

[Mar07] Yosihito Maruyama. Asymptotic expansions of the null distributions of some test statistics for profile analysis in general distributions. Journal of Statistical Planning and Inference, 137(2):506–526, February 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000590.

Marteau:2009:SRH

[Mar09] C. Marteau. On the stability of the risk hull method for projection estimators. Journal of Statistical Planning and Inference, 139(6):1821–1835, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003662.

Massam:2001:BRM

[Mas01] Hélène Massam. Book review: Mathematical Statistics: Jun Shao, 530 pages, ISBN 0-387-98674-X. Journal of Statistical Planning and Inference, 92(1–2):305–306, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001543.

Mashayekhi:2003:NLE

[Mas03] Mostafa Mashayekhi. A note on linear empirical Bayes estimation of finite population means. Journal of Statistical Planning and Inference, 112(1–2):77–88, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003245.

Mason:2007:SOK

[Mas07] David M. Mason. Some observations on the KMT dyadic scheme. Journal of Statistical Planning and Inference, 137 (3):895–906, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001492.

Masse:2009:MTM

[Mas09] Jean-Claude Massé. Multivariate trimmed means based on the Tukey depth. Journal of Statistical Planning and Inference, 139(2):366–384, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001912.

Matthews:2001:AR

[Mat01a] R. A. J. Matthews. Author’s response. Journal of Statistical Planning and Inference, 94(1):69–71, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002366.

Matthews:2001:WSC

[Mat01b] Robert A. J. Matthews. Why should clinicians care about Bayesian methods? Journal of Statistical Planning and Inference, 94(1):43–58, March 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002329.

Ma:2000:EBE

[MB00a] Yimin Ma and N. Balakrishnan. Empirical Bayes estimation for truncation parameters. Journal of Statistical Planning and Inference, 84(1–2):111–120, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001135.

Munzel:2000:NMM

[MB00b] Ullrich Munzel and Edgar Brunner. Nonparametric methods in multivariate factorial designs. Journal of Statistical Planning and Inference, 88(1):117–132, July 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002128.

Molenberghs:2008:MAF

[MBA^{+}08] Geert Molenberghs, Tomasz Burzykowski, Ariel Alonso, Pryseley Assam, Abel Tilahun, and Marc Buyse. The meta-analytic framework for the evaluation of surrogate endpoints in clinical trials. Journal of Statistical Planning and Inference, 138(2):432–449, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002534.

Muller:2007:SBS

[MBG^{+}07] Peter Müller, Don A. Berry, Andy P. Grieve, Michael Smith, and Michael Krams. Simulation-based sequential Bayesian design. Journal of Statistical Planning and Inference, 137(10):3140–3150, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700081X.

Marques:2008:NED

[MC08] Filipe J. Marques and Carlos A. Coelho. Near-exact distributions for the sphericity likelihood ratio test statistic. Journal of Statistical Planning and Inference, 138(3):726–741, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000596.

Martinez-Camblor:2008:STB

[MCDC08] P. Martínez-Camblor, J. De Uña-Álvarez, and N. Corral. k-sample test based on the common area of kernel density estimators. Journal of Statistical Planning and Inference, 138(12):4006–4020, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000499.

McFarland:2001:BRA

[McF01] Richard McFarland. Book review: Applied multivariate statistics with the SAS system, Second Edition: R. Khattree and D. N. Naik; SAS Publishing, Cary, NC, 1999, 338pp., $39.95 ISBN 1-58025-357-1. Journal of Statistical Planning and Inference, 96(2):437–438, July 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003104.

Molefe:2005:TET

[MCH^{+}05] Daniel F. Molefe, James J. Chen, Paul C. Howard, Barbara J. Miller, Christopher P. Sambuco, P. Donald Forbes, and Ralph L. Kodell. Tests for effects on tumor frequency and latency in multiple dosing photococarcinogenicity experiments. Journal of Statistical Planning and Inference, 129(1–2):39–58, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001533.

McLeod:2008:OBS

[McL08] Robert G. McLeod. Optimal block sequences for blocked fractional factorial split-plot designs. Journal of Statistical Planning and Inference, 138(8):2563–2573, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003916.

Morales:2007:BDQ

[MCM^{+}07] Javier Morales, M. Eugenia Castellanos, Asunción M. Mayoral, Roland Fried, and Carmen Armero. Bayesian design in queues: an application to aeronautic maintenance. Journal of Statistical Planning and Inference, 137(10):3058–3067, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000742.

Miranda:2005:LPI

[MdCC05] Enrique Miranda, Gert de Cooman, and Inés Couso. Lower previsions induced by multi-valued mappings. Journal of Statistical Planning and Inference, 133(1):173–197, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001053.

Micheas:2006:CED

[MDM06] Athanasios C. Micheas, Dipak K. Dey, and Kanti V. Mardia. Complex elliptical distributions with application to shape analysis. Journal of Statistical Planning and Inference, 136(9):2961–2982, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000339.

Marron:2004:SLB

[MdUÁ04] J. S. Marron and Jacobo de Uña-Álvarez. SiZer for length biased, censored density and hazard estimation. Journal of Statistical Planning and Inference, 121(1):149–161, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802005062.

Martin:2004:VBD

[ME04] R. J. Martin and J. A. Eccleston. Variance-balanced designs under interference for dependent observations. Journal of Statistical Planning and Inference, 119(1):207–223, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580200407X.

Mushkudiani:2007:GPP

[ME07] Nino A. Mushkudiani and John H. J. Einmahl. Generalized probability-probability plots. Journal of Statistical Planning and Inference, 137(3):738–752, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001388.

Martin:2004:EFE

[MEC04] R. J. Martin, J. A. Eccleston, and B. S. P. Chan. Efficient factorial experiments when the data are spatially correlated. Journal of Statistical Planning and Inference, 126(1):377–395, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002714.

Meeden:2005:NBA

[Mee05]  Glen Meeden.  A noninformative Bayesian approach

to domain estimation. Journal of Statistical Planning and Inference, 129(1–2):85–92, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001557.

Meintanis:2005:CTS

[Mei05] Simos G. Meintanis. Consistent tests for symmetric stability with finite mean based on the empirical characteristic function. Journal of Statistical Planning and Inference, 128(2):373–380, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580400031X.

Meintanis:2007:KST

[Mei07a] Simos G. Meintanis. A Kolmogorov–Smirnov type test for skew normal distributions based on the empirical moment generating function. Journal of Statistical Planning and Inference, 137(8):2681–2688, August 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000195.

Meintanis:2007:TFM

[Mei07b] Simos G. Meintanis. Test of fit for Marshall-Olkin distributions with applications. Journal of Statistical Planning and Inference, 137(12):3954–3963, December 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001632.

Meintanis:2009:GFT

[Mei09] Simos G. Meintanis. Goodness-of-fit tests and minimum distance estimation via optimal transformation to uniformity. Journal of Statistical Planning and Inference, 139(2):100–108, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001511.

Melas:2000:P

[Mel00] V. Melas. Preface. Journal of Statistical Planning and Inference, 85(1–2):3–4, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000592.

Melas:2005:FAO

[Mel05a] Viatcheslav B. Melas. On the functional approach to optimal designs for nonlinear models. Journal of Statistical Planning and Inference, 132(1–2):93–116, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001442.

Melas:2005:P

[Mel05b] Viatcheslav B. Melas. Preface. Journal of Statistical Planning and Inference, 132(1–2):1, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001508.

Melas:2007:EO

[Mel07a] Viatcheslav B. Melas. Editorial overview. Journal of Statistical Planning and Inference, 137(8):2535, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000249.

Melas:2007:P

[Mel07b] Viatcheslav B. Melas. Preface. Journal of Statistical Planning and Inference, 137(12):3827, December 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001516.

Meyer:2006:CPT

[Mey06] Mary C. Meyer. Consistency and power in tests with shape-restricted alternatives. Journal of Statistical Planning and Inference, 136(11):3931–3947, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001151.

Meza:2003:EBE

[Mez03] Jane L. Meza. Empirical Bayes estimation smoothing of relative risks in disease mapping. Journal of Statistical Planning and Inference, 112(1–2):43–62, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003221.

Martins:2004:EIS

[MF04] A. P. Martins and H. Ferreira. The extremal index of sub-sampled processes. Journal of Statistical Planning and Inference, 124(1):145–152, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001940.

Ma:2003:NUO

[MFL03] Chang-Xing Ma, Kai-Tai Fang, and Dennis K. J. Lin. A note on uniformity and orthogonality. Journal of Statistical Planning and Inference, 113(1):323–334, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002932.

Minhajuddin:2007:BTM

[MFSW07] Abu T. M. Minhajuddin, William H. Frawley, William R. Schucany, and Wayne A. Woodward. Bootstrap tests for multivariate directional alternatives. Journal of Statistical Planning and Inference, 137(7):2302–2315, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002473.

Messer:2008:ADD

[MG08] Karen Messer and Anthony Gamst. The asymptotic distribution of the delete-d jackknife variance estimator for smooth functionals. Journal of Statistical Planning and Inference, 138(7):2154–2162, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003643.

Majumdar:2005:STC

[MGB05] Anandamayee Majumdar, Alan E. Gelfand, and Sudipto Banerjee. Spatio-temporal change-point modeling. Journal of Statistical Planning and Inference, 130(1–2):149–166, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002599.

Misra:2009:SCM

[MGG09] Neeraj Misra, Nitin Gupta, and Rameshwar D. Gupta. Stochastic comparisons of multivariate frailty models. Journal of Statistical Planning and Inference, 139(6):2084–2090, June 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003844.

Mudholkar:2000:ESN

[MH00] Govind S. Mudholkar and Alan D. Hutson. The epsilon-skew-normal distribution for analyzing near-normal data. Journal of Statistical Planning and Inference, 83(2):291–309, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899000968.

Morris:2009:CGR

[MH09] Max D. Morris and Dave Higdon. Comments on Goldstein and Rougier. Journal of Statistical Planning and Inference, 139(3):1249–1250, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003352.

Mudholkar:2007:LLS

[MHM07] Govind S. Mudholkar, Alan D. Hutson, and Michael P. Mc-Dermott. Life and legacy of Samarendra Nath Roy 1906– 1964. Journal of Statistical Planning and Inference, 137 (11):3208–3212, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000882.

Mi:2006:MPL

[Mi06] Jie Mi. MLE of parameters of location-scale distribution for complete and partially grouped data. Journal of Statistical Planning and Inference, 136(10):3565–3582, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000807.

Malefaki:2008:CPW

[MI08a] Sonia Malefaki and George Iliopoulos. On convergence of properly weighted samples to the target distribution. Journal of Statistical Planning and Inference, 138(4):1210–1225, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002261.

Mi:2008:OST

[Mi08b] Jie Mi. Order statistics and test duration. Journal of Statistical Planning and Inference, 138(4):834–844, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000493.

Minami:2007:MIG

[Min07] Mihoko Minami. Multivariate inverse Gaussian distribution as a limit of multivariate waiting time distributions. Journal of Statistical Planning and Inference, 137(11):3626–3633, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001231.

Mishima:2002:CMT

[Mis02] Miwako Mishima. Cyclic Mendelsohn triple systems with a cyclic resolution or a cyclic almost resolution. Journal of Statistical Planning and Inference, 106(1–2):105–115, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002069.

Mishra:2007:FPG

[Mis07] Nutan S. Mishra. Families of proper generalized neighbor designs. Journal of Statistical Planning and Inference, 137 (5):1681–1686, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806002503.

Massonnet:2009:MUI

[MJD09] Goele Massonnet, Paul Janssen, and Luc Duchateau. Modelling udder infection data using copula models for quadruples. Journal of Statistical Planning and Inference, 139(11):3865–3877, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001530.

Mishima:2001:CCT

[MJK01] Miwako Mishima, Masakazu Jimbo, and Sanpei Kageyama. Constructions for a certain type of balanced bipartite block designs. Journal of Statistical Planning and Inference, 95 (1–2):257–268, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002937.

Mishima:2000:OOA

[MJS00] Miwako Mishima, Masakazu Jimbo, and Teruhiro Shirakura. On the optimality of orthogonal arrays in case of correlated errors. Journal of Statistical Planning and Inference, 88(2):319–338, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800000860.

Mukhopadhyay:2008:CDM

[MK08] S. Mukhopadhyay and A. I. Khuri. Comparison of designs for multivariate generalized linear models. Journal of Statistical Planning and Inference, 138(1):169–183, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002078.

Moreno:2003:DBT

[ML03] Elías Moreno and Brunero Liseo. A default Bayesian test for the number of components in a mixture. Journal of Statistical Planning and Inference, 111(1–2):129–142, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580200294X.

Mattheou:2009:MSC

[MLK09] K. Mattheou, S. Lee, and A. Karagrigoriou. A model selection criterion based on the BHHJ measure of divergence. Journal of Statistical Planning and Inference, 139(2):228–235, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800181X.

Majumdar:2002:FOD

[MM02] Dibyen Majumdar and R. J. Martin. Finding optimal designs in the presence of trends. Journal of Statistical Planning and Inference, 106(1–2):177–190, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002112.

Monbet:2006:NPR

[MM06] Valérie Monbet and Pierre-François Marteau. Non parametric resampling for stationary Markov processes: the local grid bootstrap approach. Journal of Statistical Planning and Inference, 136(10):3319–3338, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000637.

Molina:2007:RST

[MM07] I. Molina and D. Morales. Rényi statistics for testing hypotheses in mixed linear regression models. Journal of Statistical Planning and Inference, 137(1):87–102, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805002983.

Mudholkar:2002:ICT

[MML02] Govind S. Mudholkar, Carol E. Marchetti, and C. Thomas Lin. Independence characterizations and testing normality against restricted skewness-kurtosis alternatives. Journal of Statistical Planning and Inference, 104(2):485–501, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002531.

Morris:2006:SPB

[MMM06] Max D. Morris, Leslie M. Moore, and Michael D. McKay. Sampling plans based on balanced incomplete block designs for evaluating the importance of computer model inputs. Journal of Statistical Planning and Inference, 136(9): 3203–3220, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805000157.

Mallor:2007:CBS

[MMM07] F. Mallor, P. Mateo, and J. A. Moler. A comparison between several adjustment models to simulated teletraffic data. Journal of Statistical Planning and Inference, 137(12):3939–3953, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001620.

McDonough:2000:GHM

[MMP00] T. P. McDonough, V. C. Mavron, and C. A. Pallikaros. Generalised Hadamard matrices and translations. Journal of Statistical Planning and Inference, 86(2):527–533, May 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001287.

Muller:2003:BPT

[MN03] Christine H. Müller and Neyko Neykov. Breakdown points of trimmed likelihood estimators and related estimators in generalized linear models. Journal of Statistical Planning and Inference, 116(2):503–519, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002653.

Meintanis:2008:CCM

[MN08a] S. G. Meintanis and Ya. Yu. Nikitin. A class of count models and a new consistent test for the Poisson distribution. Journal of Statistical Planning and Inference, 138(12):3722–3732, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001304.

Miller:2008:GLF

[MN08b] Forrest R. Miller and James W. Neill. General lack of fit tests based on families of groupings. Journal of Statistical Planning and Inference, 138(8):2433–2449, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000268.

Montanero:2005:SPM

[MNOP05] Jesús Montanero, Agustín G. Nogales, José A. Oyola, and Paloma Pérez. On the structure of probability measure families admitting a partially sufficient σ-field. Journal of Statistical Planning and Inference, 134(1):224–235, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001867.

Marshall:2007:CDT

[MO07] Albert W. Marshall and Ingram Olkin. Characterizations of distributions through coincidences of semiparametric families. Journal of Statistical Planning and Inference, 137(11):3618–3625, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700122X.

Mohanty:2002:P

[Moh02a] S. G. Mohanty. Preface. Journal of Statistical Planning and Inference, 101(1–2):1, February 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001707.

Mohanty:2002:TVN

[Moh02b] Sri Gopal Mohanty. Tadepalli Venkata Narayana 1930–1987. Journal of Statistical Planning and Inference, 101(1–2):5, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001720.

Mohanty:2005:P

[Moh05a] S. G. Mohanty. Preface. Journal of Statistical Planning and Inference, 135(1):1–2, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000509.

Mohanty:2005:SJD

[Moh05b] S. G. Mohanty. Some joint distributions related to the Galton statistic and runs. Journal of Statistical Planning and Inference, 135(1):158–164, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000613.

Mojirsheibani:2007:NCE

[Moj07] Majid Mojirsheibani. Nonparametric curve estimation with missing data: a general empirical process approach. Journal of Statistical Planning and Inference, 137(9):2733– 2758, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000511.

Morales:2009:CRN

[Mor09] Luis B. Morales. Constructing 1-rotational NRDFs through an optimization approach: new (46,9,8), (51,10,9) and (55,9,8)-NRBDs. Journal of Statistical Planning and Inference, 139(1):62–68, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002309.

Melfi:2000:EAA

[MP00] Vincent F. Melfi and Connie Page. Estimation after adaptive allocation. Journal of Statistical Planning and Inference, 87(2):353–363, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001986.

Malyutov:2002:LLC

[MP02] M. B. Malyutov and R. S. Protassov. LAN, LAM and convergence of Iterative ML estimates; optimal design in Gaussian one-way mixed model. Journal of Statistical Planning and Inference, 100(2):249–268, February 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001380.

Mokkadem:2005:MDK

[MP05] Abdelkader Mokkadem and Mariane Pelletier. Moderate deviations for the kernel mode estimator and some applications. Journal of Statistical Planning and Inference, 135(2):276–299, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002381.

McDonald:2006:MPG

[MP06] Gary C. McDonald and S. Panchapakesan. A monotonicity property for gamma distributions. Journal of Statistical Planning and Inference, 136(7):2279–2283, July 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002211.

Mas:2007:AM

[MP07a] André Mas and Besnik Pumo. The ARHD model. Journal of Statistical Planning and Inference, 137(2):538–553, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806000310.

McElroy:2007:MBT

[MP07b] Tucker McElroy and Dimitris N. Politis. Moment-based tail index estimation. Journal of Statistical Planning and Inference, 137(4):1389–1406, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001005.

McSorley:2007:CEP

[MP07c] John P. McSorley and Nicholas C. Phillips. Complete enumeration and properties of binary pseudo-Youden designs PYD(9,6,6). Journal of Statistical Planning and Inference, 137(4):1464–1473, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000838.

Melas:2007:RME

[MP07d] Viatcheslav Melas and Andrey Pepelyshev. On representing maximin efficient designs by Taylor series. Journal of Statistical Planning and Inference, 137(8):2689–2702, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000201.

Modarres:2007:HDB

[MP07e] Reza Modarres and G. P. Patil. Hotspot detection with bivariate data. Journal of Statistical Planning and Inference, 137(11):3643–3654, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001255.

Morgan:2007:ORC

[MP07f] J. P. Morgan and Valentin Parvu. Optimal row-column design for three treatments. Journal of Statistical Planning and Inference, 137(4):1474–1487, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000784.

Mathew:2008:VTE

[MP08] Thomas Mathew and Gitanjali Paul. A p-value for testing the equivalence of the variances of a bivariate normal distribution. Journal of Statistical Planning and Inference, 138(12):3982–3992, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000451.

Mandal:2009:FDB

[MPGS09] B. N. Mandal, Rajender Parsad, V. K. Gupta, and U. C. Sud. A family of distance balanced sampling plans. Journal of Statistical Planning and Inference, 139(3):860–874, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002589.

Makri:2007:SLL

[MPP07] Frosso S. Makri, Andreas N. Philippou, and Zaharias M. Psillakis. Shortest and longest length of success runs in binary sequences. Journal of Statistical Planning and Inference, 137(7):2226–2239, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002229.

Madruga:2003:BET

[MPS03] M. R. Madruga, C. A. B. Pereira, and J. M. Stern. Bayesian evidence test for precise hypotheses. Journal of Statistical Planning and Inference, 117(2):185–198, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003683.

Moler:2005:ADR

[MPS05] José A. Moler, Fernando Plo, and Miguel San Miguel. Adaptive designs and Robbins-Monro algorithm. Journal of Statistical Planning and Inference, 131(1):161–174, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000655.

Muliere:2006:RRU

[MPS06] Pietro Muliere, Anna Maria Paganoni, and Piercesare Secchi. A randomly reinforced urn. Journal of Statistical Planning and Inference, 136(6):1853–1874, June 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001904.

Moler:2007:SDC

[MPS07] José Antonio Moler, Fernando Plo, and Miguel San Miguel. A sequential design for a clinical trial with a linear prognostic factor. Journal of Statistical Planning and Inference, 137(12):3964–3974, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001644.

Mokkadem:2009:SAM

[MPS09] Abdelkader Mokkadem, Mariane Pelletier, and Yousri Slaoui. The stochastic approximation method for the estimation of a multivariate probability density. Journal of Statistical Planning and Inference, 139(7):2459–2478, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004412.

Militino:2006:ODM

[MPU06] A. F. Militino, M. B. Palacios, and M. D. Ugarte. Outliers detection in multivariate spatial linear models. Journal of Statistical Planning and Inference, 136(1):125–146, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002988.

Moreno:2000:BTU

[MR00] Marta Moreno and Juan Romo. Bootstrap tests for unit roots based on LAD estimation. Journal of Statistical Planning and Inference, 83(2):347–367, February 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001019.

Mukerjee:2001:CTS

[MR01] Rahul Mukerjee and Nancy Reid. Comparison of test statistics via expected lengths of associated confidence intervals. Journal of Statistical Planning and Inference, 97(1):141–151, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800003505.

Miescke:2006:PGS

[MR06] Klaus J. Miescke and Kenneth J. Ryan. On the performance of Gupta’s subset selection rule. Journal of Statistical Planning and Inference, 136(7):2004–2019, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002077.

Morgan:2007:ODI

[MR07] J. P. Morgan and Brian Reck. E-optimal design in irregular BIBD settings. Journal of Statistical Planning and Inference, 137(5):1658–1668, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002485.

Martin-Rodriguez:2002:CIS

[MRaPGVVV02] Jesús Martín-Rodríguez, M. a Purificación Galindo-Villardón, and José L. Vicente-Villardón. Comparison and integration of subspaces from a biplot perspective. Journal of Statistical Planning and Inference, 102(2):411–423, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100101X.

Mohacsy:2002:CSD

[MRC02] Hedvig Mohácsy and D. K. Ray-Chaudhuri. Candelabra systems and designs. Journal of Statistical Planning and Inference, 106(1–2):419–448, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002264.

Mohacsy:2003:CGD

[MRC03] Hedvig Mohácsy and D. K. Ray-Chaudhuri. A construction for group divisible t-designs with strength t ≥ 2 and index unity. Journal of Statistical Planning and Inference, 109(1–2):167–177, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003099.

Molinares:2009:REL

[MRCN09] Fabio Fajardo Molinares, Valdério Anselmo Reisen, and Francisco Cribari-Neto. Robust estimation in long-memory processes under additive outliers. Journal of Statistical Planning and Inference, 139(8):2511–2525, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004771.

Moreno-Rebollo:2000:EUS

[MRLBBCPA00] J. L. Moreno-Rebollo, F. López-Blázquez, I. Barranco-Chamorro, and A. Pascual-Acosta. Estimating the unknown sample size. Journal of Statistical Planning and Inference, 83(2):311–318, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837589900097X.

Mair:2000:SCW

[MRU00] Bernard A. Mair, Frits H. Ruymgaart, and Theresa Urrabazo. Some comments on Wicksell’s problem. Journal of Statistical Planning and Inference, 87(1):31–42, May 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001780.

McCann:2000:MNE

[MS00] Melinda McCann and Sahadeb Sarkar. Minimum negative exponential disparity estimation of mixture proportions. Journal of Statistical Planning and Inference, 87(2):187–197, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001895.

Mathew:2001:ODB

[MS01] Thomas Mathew and Bikas Kumar Sinha. Optimal designs for binary data under logistic regression. Journal of Statistical Planning and Inference, 93(1–2):295–307, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800001737.

Mathew:2002:JCR

[MS02a] Thomas Mathew and Manoj Kumar Sharma. Joint confidence regions in the multivariate calibration problem. Journal of Statistical Planning and Inference, 100(2):427–441, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001525.

Mathew:2002:CMU

[MS02b] Thomas Mathew and Manoj Kumar Sharma. On the construction of multiple use confidence regions based on combined information in univariate calibration. Journal of Statistical Planning and Inference, 103(1–2):151–172, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002178.

Morgan:2002:CSD

[MS02c] John P. Morgan and Sudesh K. Srivastav. The completely symmetric designs with blocksize three. Journal of Statistical Planning and Inference, 106(1–2):21–30, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001982.

Mukerjee:2002:CIA

[MS02d] Rahul Mukerjee and Brajendra Chandra Sutradhar. On confidence intervals associated with quasi-likelihood. Journal of Statistical Planning and Inference, 102(1):119–128, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001781.

Mukhopadhyay:2002:MMC

[MS02e] Nitis Mukhopadhyay and Tumulesh K. S. Solanky. Multistage methodologies for comparing several treatments with a control. Journal of Statistical Planning and Inference, 100(2):209–220, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001355.

Mouchart:2003:SII

[MS03] M. Mouchart and E. San Martín. Specification and identification issues in models involving a latent hierarchical structure. Journal of Statistical Planning and Inference, 111(1–2):143– 163, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002951.

Marchand:2005:IMR

[MS05] Éric Marchand and William E. Strawderman. On improving on the minimum risk equivariant estimator of a scale parameter under a lower-bound constraint. Journal of Statistical Planning and Inference, 134(1):90–101, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001806.

Maruyama:2006:NCM

[MS06] Yuzo Maruyama and William Edward Strawderman. A new class of minimax generalized Bayes estimators of a normal variance. Journal of Statistical Planning and Inference, 136(11):3822–3836, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001370.

Mitra:2007:GVA

[MS07a] Pranab K. Mitra and Bimal K. Sinha. A generalized p-value approach to inference on common mean. Journal of Statistical Planning and Inference, 137(11):3634–3642, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001243.

Mitra:2007:SAE

[MS07b] Pranab K. Mitra and Bimal K. Sinha. On some aspects of estimation of a common mean of two independent normal populations. Journal of Statistical Planning and Inference, 137(1):184–193, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000280.

Matei:2009:OSC

[MS09a] Alina Matei and Chris Skinner. Optimal sample coordination using controlled selection. Journal of Statistical Planning and Inference, 139(9):3112–3121, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000512.

Meerschaert:2009:ELT

[MS09b] Mark M. Meerschaert and Stilian A. Stoev. Extremal limit theorems for observations separated by random power law waiting times. Journal of Statistical Planning and Inference, 139(7):2175–2188, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003984.

Merlini:2002:SSD

[MSV02] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. Some statistics on Dyck paths. Journal of Statistical Planning and Inference, 101(1–2):211–227, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100180X.

Muller:2004:ELF

[MSW04] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Estimating linear functionals of the error distribution in nonparametric regression. Journal of Statistical Planning and Inference, 119(1):75–93, January 15, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004123.

Meister:2009:CRO

[MSW09] Alexander Meister, Ulrich Stadtmüller, and Christian Wagner. Consistent and rate-optimal density estimation from heteroscedastic data groups. Journal of Statistical Planning and Inference, 139(6):1893–1904, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003558.

Mason:2000:MCC

[MT00] David M. Mason and Tatyana S. Turova. Motoo’s combinatorial central limit theorem for serial rank statistics. Journal of Statistical Planning and Inference, 91(2):427–440, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001920.

Mudholkar:2002:ECI

[MT02] Govind S. Mudholkar and Lili Tian. An entropy characterization of the inverse Gaussian distribution and related goodness-of-fit test. Journal of Statistical Planning and Inference, 102(2):211–221, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000994.

Mudholkar:2004:THO

[MT04] Govind S. Mudholkar and Lili Tian. A test for homogeneity of ordered means of inverse Gaussian populations. Journal of Statistical Planning and Inference, 118(1–2):37–49, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003968.

Mandal:2006:COD

[MT06] S. Mandal and B. Torsney. Construction of optimal designs using a clustering approach. Journal of Statistical Planning and Inference, 136(3):1120–1134, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003520.

Manski:2007:ATR

[MT07a] Charles F. Manski and Aleksey Tetenov. Admissible treatment rules for a risk-averse planner with experimental data on an innovation. Journal of Statistical Planning and Inference, 137(6):1998–2010, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001820.

McCann:2007:PCP

[MT07b] Melinda H. McCann and Joshua M. Tebbs. Pairwise comparisons for proportions estimated by pooled testing. Journal of Statistical Planning and Inference, 137(4):1278–1290, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000796.

Mandal:2005:COD

[MTC05a] S. Mandal, B. Torsney, and K. C. Carriere. Constructing optimal designs with constraints. Journal of Statistical Planning and Inference, 128(2):609–621, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000217.

Moreno:2005:TER

[MTC05b] Eliías Moreno, Francisco Torres, and George Casella. Testing equality of regression coefficients in heteroscedastic normal regression models. Journal of Statistical Planning and Inference, 131(1):117–134, April 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000631.

Morgan:2003:ORC

[MU03] J. P. Morgan and Nizam Uddin. Optimal row-column design for two treatments. Journal of Statistical Planning and Inference, 115(2):603–622, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001726.

Mudholkar:2007:P

[Mud07] Govind S. Mudholkar. Preface. Journal of Statistical Planning and Inference, 137(11):3207, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000870.

Mukhopadhyay:2006:MLA

[Muk06] Chiranjit Mukhopadhyay. Maximum likelihood analysis of masked series system lifetime data. Journal of Statistical Planning and Inference, 136(3):803–838, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003349.

Munk:2000:UTA

[Mun00] Axel Munk. An unbiased test for the average equivalence problem — the small sample case. Journal of Statistical Planning and Inference, 87(1):69–86, May 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001743.

Mushkudiani:2002:SNT

[Mus02] Nino A. Mushkudiani. Small nonparametric tolerance regions for directional data. Journal of Statistical Planning and Inference, 100(1):67–80, January 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000933.

Mutoh:2005:EBI

[Mut05] Yukiyasu Mutoh. On the existence of balanced incomplete block designs with nested rows and columns. Journal of Statistical Planning and Inference, 133(2):489–508, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001739.

Misra:2003:SPS

[MvdM03] Neeraj Misra and Edward C. van der Meulen. On stochastic properties of m-spacings. Journal of Statistical Planning and Inference, 115(2):683–697, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200157X.

Misra:2006:SIR

[MvdMV06] Neeraj Misra, Edward C. van der Meulen, and Karlien Vanden Branden. On some inadmissibility results for the scale parameters of selected gamma populations. Journal of Statistical Planning and Inference, 136(7):2340–2351, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002260.

McKeague:2000:MCM

[MW00a] Ian W. McKeague and Wolfgang Wefelmeyer. Markov chain Monte Carlo and Rao–Blackwellization. Journal of Statistical Planning and Inference, 85(1–2):171–182, April 1, 2000.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000798.

McNeney:2000:ACT

[MW00b] Brad McNeney and Jon A. Wellner. Application of convolution theorems in semiparametric models with non-i.i.d. data. Journal of Statistical Planning and Inference, 91(2):441–480, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001932.

McDermott:2002:CUM

[MW02] Michael P. McDermott and Yining Wang. Construction of uniformly more powerful tests for hypotheses about linear inequalities. Journal of Statistical Planning and Inference, 107(1–2):207–217, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002537.

Mallick:2003:BST

[MW03] Bani K. Mallick and Stephen Walker. A Bayesian semiparametric transformation model incorporating frailties. Journal of Statistical Planning and Inference, 112(1–2):159–174, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003300.

Mudholkar:2005:TWH

[MW05] Govind S. Mudholkar and Gregory E. Wilding. Two Wilson– Hilferty type approximations for the null distribution of the Blum, Kiefer and Rosenblatt test of bivariate independence. Journal of Statistical Planning and Inference, 128(1):31–41, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003756.

Mena:2007:SMT

[MW07a] Ramsés H. Mena and Stephen G. Walker. Stationary mixture transition distribution (MTD) models via predictive distributions. Journal of Statistical Planning and Inference, 137(10):3103–3112, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700078X.

Mudholkar:2007:ISS

[MW07b] Govind S. Mudholkar and Hongyue Wang. IG-symmetry and R-symmetry: Interrelations and applications to the inverse Gaussian theory. Journal of Statistical Planning and Inference, 137(11):3655–3671, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001267.

Masaro:2008:ODC

[MW08a] Joe Masaro and Chi Song Wong. D-optimal designs for correlated random vectors. Journal of Statistical Planning and Inference, 138(12):4093–4106, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000669.

Mathew:2008:PPT

[MW08b] Thomas Mathew and Yanping Wu. Pilot-pivotal trials for average bioequivalence. Journal of Statistical Planning and Inference, 138(7):2106–2116, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003679.

Maronna:2000:RRB

[MY00] Ricardo A. Maronna and Viíctor J. Yohai. Robust regression with both continuous and categorical predictors. Journal of Statistical Planning and Inference, 89(1–2):197–214, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899002086.

Marazzi:2004:ATM

[MY04] Alfio Marazzi and Victor J. Yohai. Adaptively truncated maximum likelihood regression with asymmetric errors. Journal of Statistical Planning and Inference, 122(1–2):271–291, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002350.

Muler:2008:REG

[MY08] Nora Muler and Victor J. Yohai. Robust estimates for GARCH models. Journal of Statistical Planning and Inference, 138(10):2918–2940, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807004089.

Mukhopadhyay:2007:BRE

[MZ07] N. Mukhopadhyay and S. Zacks. Bounded risk estimation of linear combinations of the location and scale parameters in exponential distributions under two-stage sampling. Journal of Statistical Planning and Inference, 137(11):3672–3686, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001279.

Nayak:2009:UFA

[NA09] Tapan K. Nayak and Samson A. Adeshiyan. A unified framework for analysis and comparison of randomized response surveys of binary characteristics. Journal of Statistical Planning and Inference, 139(8):2757–2766, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004746.

Nagakura:2009:TCS

[Nag09] Daisuke Nagakura. Testing for coefficient stability of AR(1) model when the null is an integrated or a stationary process. Journal of Statistical Planning and Inference, 139(8):2731–2745, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004758.

Nakayama:2000:MCB

[Nak00] Marvin K. Nakayama. Multiple comparisons with the best using common random numbers for steady-state simulations. Journal of Statistical Planning and Inference, 85(1–2):37–48, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899000646.

Nakayama:2009:AVS

[Nak09] Marvin K. Nakayama. Asymptotically valid single-stage multiple-comparison procedures. Journal of Statistical Planning and Inference, 139(4):1348–1356, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003376.

Namba:2000:MDP

[Nam00] Akio Namba. MSE dominance of the PT-2SHI estimator over the positive-part Stein-rule estimator in regression. Journal of Statistical Planning and Inference, 89(1–2):175–185, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800000951.

Nau:2002:AIP

[Nau02] Robert F. Nau. The aggregation of imprecise probabilities. Journal of Statistical Planning and Inference, 105(1):265–282, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002130.

Navarro:2008:LRO

[Nav08] Jorge Navarro. Likelihood ratio ordering of order statistics, mixtures and systems. Journal of Statistical Planning and Inference, 138(5):1242–1257, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001735.

Nair:2003:ODE

[NaX03] Vijayan N. Nair and Li an Xu. Optimal design of experiments for modeling processes with feedback control variables. Journal of Statistical Planning and Inference, 113(1):269–284, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801003056.

Nayak:2000:BUP

[Nay00] Tapan K. Nayak. On best unbiased prediction and its relationships to unbiased estimation. Journal of Statistical Planning and Inference, 84(1–2):171–189, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001524.

Nayak:2001:BR

[Nay01] Tapan K. Nayak. Book review. Journal of Statistical Planning and Inference, 93(1–2):323–324, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002214.

Nayak:2003:FOE

[Nay03] Tapan K. Nayak. Finding optimal estimators in survey sampling using unbiased estimators of zero. Journal of Statistical Planning and Inference, 114(1–2):21–30, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004603.

Ng:2005:WPM

[NB05] H. K. T. Ng and N. Balakrishnan. Weighted precedence and maximal precedence tests and an extension to progressive censoring. Journal of Statistical Planning and Inference, 135(1):197–221, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000650.

Nguyen:2008:NLR

[NB08] Nam-Ky Nguyen and John J. Borkowski. New 3-level response surface designs constructed from incomplete block designs. Journal of Statistical Planning and Inference, 138(1):294–305, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002157.

Nowak:2005:IFC

[NBH05] Emmanuel Nowak and Avner Bar-Hen. Influence function and correspondence analysis. Journal of Statistical Planning and Inference, 134(1):26–35, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001910.

Nielsen:2005:RSQ

[ND05] J. D. Nielsen and C. B. Dean. Regression splines in the quasi-likelihood analysis of recurrent event data. Journal of Statistical Planning and Inference, 134(2):521–535, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002150.

Ng:2006:CIP

[ND06] Meei Pyng Ng and Martin Donadio. Computing inclusion probabilities for order sampling. Journal of Statistical Planning and Inference, 136(11):4026–4042, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000935.

Nofuentes:2008:EVB

[NdC08] J. A. Roldán Nofuentes and J. D. Luna del Castillo. The effect of verification bias on the comparison of predictive values of two binary diagnostic tests. Journal of Statistical Planning and Inference, 138(4):950–963, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002212.

Ntzoufras:2003:BVL

[NDF03] Ioannis Ntzoufras, Petros Dellaportas, and Jonathan J. Forster. Bayesian variable and link determination for generalised linear models. Journal of Statistical Planning and Inference, 111(1–2):165–180, ???? 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002987.

Necir:2006:FLI

[Nec06] Abdelhakim Necir. A functional law of the iterated logarithm for kernel-type estimators of the tail index. Journal of Statistical Planning and Inference, 136(3):780–802, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003416.

Nekrutkin:2000:NPP

[Nek00] V. Nekrutkin. Non-Poisson particle systems and Boltzmann equations. Journal of Statistical Planning and Inference, 85(1–2):183–197, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000804.

Nettleton:2002:TOM

[Net02] Dan Nettleton. Testing for ordered means in a variation of the normal mixture model. Journal of Statistical Planning and Inference, 107(1–2):143–153, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002495.

Neuhaus:2000:MCR

[Neu00] Georg Neuhaus. A method of constructing rank tests in survival analysis. Journal of Statistical Planning and Inference, 91(2):481–497, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001944.

Nguyen:2008:SNC

[Ngu08] Man V. M. Nguyen. Some new constructions of strength 3 mixed orthogonal arrays. Journal of Statistical Planning and Inference, 138(1):220–233, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002145.

Ni:2009:ALH

[NH09] Xuelei Sherry Ni and Xiaoming Huo. Another look at Huber’s estimator: a new minimax estimator in regression with stochastically bounded noise. Journal of Statistical Planning and Inference, 139(2):503–515, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001936.

Neuhaus:2004:OTC

[NHZ04] Georg Neuhaus, Günter Heimann, and Lixing Zhu. Optimal tests for carcinogenicity in a model with fatal and incidental tumours. Journal of Statistical Planning and Inference, 119(1):153–169, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004160.

Nielsen:2001:NCM

[Nie01] Søren Feodor Nielsen. Nonparametric conditional mean imputation. Journal of Statistical Planning and Inference, 99(2):129–150, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000878.

Niederhausen:2002:GK

[Nie02a] Heinrich Niederhausen. Germain Kreweras, 1918–1998. Journal of Statistical Planning and Inference, 101(1–2):3, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001719.

Niederhausen:2002:PWR

[Nie02b] Heinrich Niederhausen. Planar walks with recursive initial conditions. Journal of Statistical Planning and Inference, 101(1–2):229–253, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001823.

Niederhausen:2005:RWO

[Nie05] Heinrich Niederhausen. Random walks in octants, and related structures. Journal of Statistical Planning and Inference, 135(1):165–196, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000625.

Nie:2007:CRM

[Nie07] Lei Nie. Convergence rate of MLE in generalized linear and nonlinear mixed-effects models: Theory and applications. Journal of Statistical Planning and Inference, 137(6): 1787–1804, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001972.

Na:2005:TWF

[NJP05] Myung Hwan Na, Jongwoo Jeon, and Dong Ho Park. Testing whether failure rate changes its trend with unknown change points. Journal of Statistical Planning and Inference, 129(1–2):317–325, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001971.

Nandi:2006:ANS

[NK06] Swagata Nandi and Debasis Kundu. Analyzing nonstationary signals using generalized multiple fundamental frequency model. Journal of Statistical Planning and Inference, 136(11):3871–3903, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580500128X.

Nikoloulopoulos:2009:FNM

[NK09] Aristidis K. Nikoloulopoulos and Dimitris Karlis. Finite normal mixture copulas for multivariate discrete data modeling. Journal of Statistical Planning and Inference, 139(11):3878–3890, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001542.

Na:2006:SEP

[NLP06] Seongryong Na, Sangyeol Lee, and Hyeonah Park. Sequential empirical process in autoregressive models with measurement errors. Journal of Statistical Planning and Inference, 136(12):4204–4216, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002429.

Nelson:2006:NNQ

[NM06] David Nelson and Glen Meeden. Noninformative nonparametric quantile estimation for simple random samples. Journal of Statistical Planning and Inference, 136(1):53–67, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002484.

Neill:2002:MCN

[NMB02] James W. Neill, Forrest R. Miller, and Duane D. Brown. Maximin clusters for nonreplicated multiresponse lack of fit tests. Journal of Statistical Planning and Inference, 102(2):359–375, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001021.

Natarajan:2005:ORI

[NMM05] Rajeshwari Natarajan, Govind S. Mudholkar, and Michael P. McDermott. Order restricted inference for the scale-like inverse Gaussian parameters. Journal of Statistical Planning and Inference, 127(1–2):229–236, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003094.

Nelson:2009:DRS

[NN09] David Nelson and Siamak Noorbaloochi. Dimension reduction summaries for balanced contrasts. Journal of Statistical Planning and Inference, 139(2):617–628, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002061.

Nanda:2006:SRG

[NP06] Asok K. Nanda and Prasanta Paul. Some results on generalized past entropy. Journal of Statistical Planning and Inference, 136(10):3659–3674, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000649.

Neves:2006:CMS

[NPA06] Cláudia Neves, Jan Picek, and M. I. Fraga Alves. The contribution of the maximum to the sum of excesses for testing max-domains of attraction. Journal of Statistical Planning and Inference, 136(4):1281–1301, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003799.

Neumeyer:2009:STC

[NPF09] Natalie Neumeyer and Juan Carlos Pardo-Fernández. A simple test for comparing regression curves versus onesided alternatives. Journal of Statistical Planning and Inference, 139(12):4006–4016, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900130X.

Noubiap:2000:MAC

[NS00] Roger Fandom Noubiap and Wilfried Seidel. A minimax algorithm for constructing optimal symmetrical balanced designs for a logistic regression model. Journal of Statistical Planning and Inference, 91(1):151–168, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001373.

Nikulin:2002:MDD

[NSV02] M. S. Nikulin, T. I. Smirnov, and V. G. Voinov. Multivariate discrete distributions induced by an urn scheme, linear Diophantine equations, unbiased estimating, testing and applications. Journal of Statistical Planning and Inference, 101(1–2):255–266, February 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001847.

Neugebauer:2005:WPD

[NvdL05] Romain Neugebauer and Mark van der Laan. Why prefer double robust estimators in causal inference? Journal of Statistical Planning and Inference, 129(1–2):405–426, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002022.

Neugebauer:2007:NCE

[NvdL07] Romain Neugebauer and Mark van der Laan. Nonparametric causal effects based on marginal structural models. Journal of Statistical Planning and Inference, 137(2):419–434, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000334.

Ngatchou-Wandji:2005:CNH

[NW05] Joseph Ngatchou-Wandji. Checking nonlinear heteroscedastic time series models. Journal of Statistical Planning and Inference, 133(1):33–68, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001168.

Noda:2001:AIG

[NWS01] Kazuo Noda, Qi-Guang Wu, and Kunio Shimizu. Admissibility and inadmissibility of a generalized Bayes unbiased estimator in a multivariate linear model. Journal of Statistical Planning and Inference, 93(1–2):197–210, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800001683.

Obilade:2006:WLR

[Obi06] Titi Obilade. A weighted likelihood ratio of two related negative hypergeometric distributions. Journal of Statistical Planning and Inference, 136(1):147–158, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002824.

Olshen:2005:VQA

[OCR^{+}05] A. B. Olshen, P. C. Cosman, A. G. Rodrigo, P. J. Bickel, and R. A. Olshen. Vector quantization of amino acids: Analysis of the HIV V3 loop region. Journal of Statistical Planning and Inference, 130(1–2):277–298, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002654.

Ozturk:2006:RSS

[OD06] Omer Ozturk and Jayant V. Deshpande. Ranked-set sample nonparametric quantile confidence intervals. Journal of Statistical Planning and Inference, 136(3):570–577, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003350.

Ohyama:2008:EPC

[ODY08] Tetsuji Ohyama, Jimmy A. Doi, and Takashi Yanagawa. Estimating population characteristics by incorporating prior values in stratified random sampling/ranked set sampling. Journal of Statistical Planning and Inference, 138(12):4021–4032, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000426.

Ogasawara:2009:AED

[Oga09] Haruhiko Ogasawara. Asymptotic expansions of the distributions of the chi-square statistic based on the asymptotically distribution-free theory in covariance structures. Journal of Statistical Planning and Inference, 139(9):3246– 3261, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809000561.

Ozol-Godfrey:2008:FDS

[OGACR08] Ayca Ozol-Godfrey, Christine Anderson-Cook, and Timothy J. Robinson. Fraction of design space plots for generalized linear models. Journal of Statistical Planning and Inference, 138(1):203–219, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700211X.

OHagan:2009:RTP

[O’H09] Anthony O’Hagan. Reification and true parameters: Discussion of Goldstein and Rougier. Journal of Statistical Planning and Inference, 139(3):1240–1242, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003315. See [GR09a, GR09b].

Ohtani:2000:PTD

[Oht00]  Kazuhiro Ohtani.  Pre-test double k-class estimators

in linear regression. Journal of Statistical Planning and Inference, 87(2):287–299, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001974.

Ozawa:2002:OCI

[OJKM02] Kazuhiro Ozawa, Masakazu Jimbo, Sanpei Kageyama, and Stanislaw Mejza. Optimality and constructions of incomplete split-block designs. Journal of Statistical Planning and Inference, 106(1–2):135–157, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002094.

Olsen:2000:SGW

[OK00] Cheryl L. Olsen and Donald L. Kreher. Steiner graphical t-wise balanced designs of type n^{r}. Journal of Statistical Planning and Inference, 86(2):535–566, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001299.

Oh:2008:BAP

[OK08] Hee-Seok Oh and Hyoung-Moon Kim. Bayesian automatic polynomial wavelet regression. Journal of Statistical Planning and Inference, 138(8):2303–2312, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000220.

OBrien:2006:PTA

[OKD06] Sean M. O’Brien, Lawrence L. Kupper, and David B. Dunson. Performance of tests of association in misspecified generalized linear models. Journal of Statistical Planning and Inference, 136(9):3090–3100, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000145.

Ouhbi:2003:NRE

[OL03] Brahim Ouhbi and Nikolaos Limnios. Nonparametric reliability estimation of semi-Markov processes. Journal of Statistical Planning and Inference, 109(1–2):155–165, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003087.

Oliveira:2007:ALA

[OM07] Manuela M. Oliveira and João Mexia. ANOVA-like analysis of matched series of studies with a common structure. Journal of Statistical Planning and Inference, 137(6):1862–1870, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001790.

Ong:2008:IEP

[OM08] S. H. Ong and Rahul Mukerjee. Interval estimation of the population mean under model uncertainty: Robust versus empirical statistics. Journal of Statistical Planning and Inference, 138(12):3696–3704, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001183.

Ongaro:2005:SBS

[Ong05] Andrea Ongaro. Size-biased sampling and discrete nonparametric Bayesian inference. Journal of Statistical Planning and Inference, 128(1):123–148, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003720.

Ogawa:2008:TTC

[OO08] Junjiro Ogawa and Ingram Olkin. A tale of two countries: the Craig–Sakamoto–Matusita theorem. Journal of Statistical Planning and Inference, 138(11):3419–3428, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000864.

Onar:2000:ATM

[OP00] A. Onar and W. J. Padgett. Accelerated test models with the inverse Gaussian distribution. Journal of Statistical Planning and Inference, 89(1–2):119–133, August 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002141.

Onar:2004:PLO

[OP04] Arzu Onar and W. J. Padgett. A penalized local D-optimality approach to design for accelerated test models. Journal of Statistical Planning and Inference, 119(2):411–420, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004895.

Oja:2005:OSR

[OP05] Hannu Oja and Davy Paindaveine. Optimal signed-rank tests based on hyperplanes. Journal of Statistical Planning and Inference, 135(2):300–323, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002356.

OQuigley:2006:TSC

[O’Q06] John O’Quigley. Theoretical study of the continual reassessment method. Journal of Statistical Planning and Inference, 136(6):1765–1780, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001849.

Orrick:2008:ESO

[Orr08] William P. Orrick. On the enumeration of some D-optimal designs. Journal of Statistical Planning and Inference, 138 (1):286–293, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807002133.

Olkin:2005:MVI

[OS05a] Ingram Olkin and Larry Shepp. A matrix variance inequality. Journal of Statistical Planning and Inference, 130(1–2):351– 358, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002708.

Oono:2005:ETO

[OS05b] Youhei Oono and Nobuo Shinozaki. Estimation of two order restricted normal means with unknown and possibly unequal variances. Journal of Statistical Planning and Inference, 131(2):349–363, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000862.

Oono:2006:CIE

[OS06] Youhei Oono and Nobuo Shinozaki. On a class of improved estimators of variance and estimation under order restriction. Journal of Statistical Planning and Inference, 136(8):2584–2605, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004392.

Ohtomo:2002:RPN

[OT02] Masahide Ohtomo and Fumikazu Tamari. On relative prime number in a sequence of positive integers. Journal of Statistical Planning and Inference, 106(1–2):509–515, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002318.

Ouwens:2006:MCL

[OTB06] Mario J. N. M. Ouwens, Frans E. S. Tan, and Martijn P. F. Berger. A maximin criterion for the logistic random intercept model with covariates. Journal of Statistical Planning and Inference, 136(3):962–981, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003532.

Otsu:2008:OED

[Ots08] Taisuke Otsu. Optimal experimental design criterion for discriminating semiparametric models. Journal of Statistical Planning and Inference, 138(12):4141–4150, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001079.

Ozturk:2001:NRS

[ÖW01] Ömer Öztürk and Douglas A. Wolfe. A new ranked set sampling protocol for the signed rank test. Journal of Statistical Planning and Inference, 96(2):351–370, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002111.

Oyet:2003:WDE

[Oye03] Alwell J. Oyet. Wavelet designs for estimating nonparametric curves with heteroscedastic error. Journal of Statistical Planning and Inference, 117(2):323–343, December 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003701.

Ozturk:2002:RSS

[Özt02] Ömer Öztürk. Ranked set sample inference under a symmetry restriction. Journal of Statistical Planning and Inference, 102(2):317–336, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000982.

Ozturk:2005:RJE

[Ozt05] Omer Ozturk. Robust joint estimation of location and scale parameters in ranked set samples. Journal of Statistical Planning and Inference, 127(1–2):295–308, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003173.

Padmawar:2006:NCR

[Pad06] V. R. Padmawar. A note on combining randomized response and direct response. Journal of Statistical Planning and Inference, 136(8):2746–2755, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004173.

Pal:2003:HOT

[Pal03] Chandranath Pal. Higher order C(α) tests with applications to mixture models. Journal of Statistical Planning and Inference, 113(1):179–187, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003123.

Paltanea:2008:CHR

[Pal08] Eugen Paltanea. On the comparison in hazard rate ordering of fail-safe systems. Journal of Statistical Planning and Inference, 138(7):1993–1997, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003084.

Pal:2009:RWR

[Pal09] Sanghamitra Pal. Rao and Wu’s re-scaling bootstrap modified to achieve extended coverages. Journal of Statistical Planning and Inference, 139(10):3552–3558, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001001.

Panchapakesan:2006:SSG

[Pan06]  S. Panchapakesan.  Shanti Swarup Gupta: His life

(1925–2002) and work. Journal of Statistical Planning and Inference, 136(7):1986–1998, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002053.

Papadatos:2001:EBL

[Pap01] Nickos Papadatos. Expectation bounds on linear estimators from dependent samples. Journal of Statistical Planning and Inference, 93(1–2):17–27, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001671.

Papadatos:2005:CDD

[Pap05] Nickos Papadatos. Characterizations of discrete distributions using the Rao–Rubin condition. Journal of Statistical Planning and Inference, 135(1):222–228, November 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000662.

Park:2000:LRT

[Par00a] Chul Gyu Park. Likelihood ratio test for homogeneity of steady-state availabilities against order restrictions. Journal of Statistical Planning and Inference, 84(1–2):159–169, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899001469.

Park:2000:SSR

[Par00b] Hyo-Il Park. A study on the scores for right censored data. Journal of Statistical Planning and Inference, 86(1):101–111, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899001597.

Parthasarathy:2002:EC

[Par02] K. R. Parthasarathy. On extremal correlations. Journal of Statistical Planning and Inference, 103(1–2):173–180, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002191.

Parzen:2008:USC

[Par08] Emanuel Parzen. United statistics, confidence quantiles, Bayesian statistics. Journal of Statistical Planning and Inference, 138(9):2777–2785, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000682.

Park:2009:CRG

[Par09a] Changyi Park. Convergence rates of generalization errors for margin-based classification. Journal of Statistical Planning and Inference, 139(8):2543–2551, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004436.

Parker:2009:DRS

[Par09b] Peter A. Parker. Discussion — “Response surface design evaluation and comparison” by Christine M. Anderson-Cook, Connie M. Borror, Douglas C. Montgomery. Journal of Statistical Planning and Inference, 139(2):645–646, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001596. See [ACBM09b].

Pavcevic:2001:SDA

[Pav01] Mario-Osvin Pavcević. On symmetric (69, 17, 4)-designs admitting an action of Frobenius groups. Journal of Statistical Planning and Inference, 94(2):277–285, April 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002597.

Pavlenko:2003:FSC

[Pav03] Tatjana Pavlenko. On feature selection, curse-of-dimensionality and error probability in discriminant analysis. Journal of Statistical Planning and Inference, 115(2):565–584, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001660.

Pazman:2002:RNL

[Páz02] Andrej Pázman. Results on nonlinear least squares estimators under nonlinear equality constraints. Journal of Statistical Planning and Inference, 103(1–2):401–420, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002348.

Pal:2008:ESR

[PB08]  Jayanta Kumar Pal and Moulinath Banerjee.  Esti-

mation of smooth regression functions in monotone response models. Journal of Statistical Planning and Inference, 138(10):3125–3143, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000189.

Provost:2002:DHQ

[PC02] Serge B. Provost and Young-Ho Cheong. The distribution of Hermitian quadratic forms in elliptically contoured random vectors. Journal of Statistical Planning and Inference, 102(2):303–316, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000969.

Park:2007:BML

[PCK07] Y. S. Park, J. W. Choi, and K. W. Kim. A balanced multi-level rotation sampling design and its efficient composite estimators. Journal of Statistical Planning and Inference, 137(2):594–610, February 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000322.

Ploner:2000:NDG

[PD00] Alexander Ploner and Rudolf Dutter. New directions in geostatistics. Journal of Statistical Planning and Inference, 91(2):499–509, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001956.

Pinheiro:2005:CGS

[PdSPS05] Hildete Prisco Pinheiro, Aluísio de Souza Pinheiro, and Pranab Kumar Sen. Comparison of genomic sequences using the Hamming distance. Journal of Statistical Planning and Inference, 130(1–2):325–339, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400268X.

Pellikaan:2001:EOF

[Pel01] Ruud Pellikaan. On the existence of order functions. Journal of Statistical Planning and Inference, 94(2):287–301, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002603.

Pensky:2000:AWE

[Pen00] Marianna Pensky. Adaptive wavelet empirical Bayes estimation of a location or a scale parameter. Journal of Statistical Planning and Inference, 90(2):275–292, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001233.

Pensky:2003:RCE

[Pen03] Marianna Pensky. Rates of convergence of empirical Bayes tests for a normal mean. Journal of Statistical Planning and Inference, 111(1–2):181–196, ???? 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003002.

Peng:2004:BCE

[Pen04] Liang Peng. Bias-corrected estimators for monotone and concave frontier functions. Journal of Statistical Planning and Inference, 119(2):263–275, February 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004871.

Perry:2000:SPC

[Per00] David Perry. Stopping problems for compound processes with applications to queues. Journal of Statistical Planning and Inference, 91(1):65–75, November 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001038.

Petrov:2007:LBT

[Pet07] Valentin V. Petrov. On lower bounds for tail probabilities. Journal of Statistical Planning and Inference, 137(8): 2703–2705, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000213.

Pewsey:2005:LSD

[Pew05] Arthur Pewsey. The large-sample distribution of the most fundamental of statistical summaries. Journal of Statistical Planning and Inference, 134(2):434–444, October 1, 2005.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002216.

Pereira:2001:LDP

[PF01] L. Pereira and H. Ferreira. Limit distribution for point processes of high local maxima. Journal of Statistical Planning and Inference, 97(2):227–233, September 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002251.

Pereira:2002:ALM

[PF02] L. Pereira and H. Ferreira. The asymptotic locations of the maximum and minimum of stationary sequences. Journal of Statistical Planning and Inference, 104(2):287–295, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002592.

Park:2008:TLS

[PF08] Man Sik Park and Montserrat Fuentes. Testing lack of symmetry in spatial-temporal processes. Journal of Statistical Planning and Inference, 138(10):2847–2866, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004028.

Park:2009:MMS

[PF09] Mingue Park and Wayne A. Fuller. The mixed model for survey regression estimation. Journal of Statistical Planning and Inference, 139(4):1320–1331, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003285.

Pfanzagl:2000:LUE

[Pfa00] J. Pfanzagl. On local uniformity for estimators and confidence limits. Journal of Statistical Planning and Inference, 84(1–2):27–53, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001445.

Pfanzagl:2003:AAA

[Pfa03] J. Pfanzagl. An alternative approach to asymptotic optimality. Journal of Statistical Planning and Inference, 116 (1):17–47, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802002331.

Pan:2000:MOD

[PFF00] Jian-Xin Pan, Wing-Kam Fung, and Kai-Tai Fang. Multiple outlier detection in multivariate data using projection pursuit techniques. Journal of Statistical Planning and Inference, 83(1):153–167, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000919.

Park:2007:PPR

[PG07a] Junyong Park and Jayanta K. Ghosh. Persistence of plug-in rule in classification of high dimensional multivariate binary data. Journal of Statistical Planning and Inference, 137(11):3687–3705, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001280.

Pozdnyakov:2007:NRS

[PG07b] Vladimir Pozdnyakov and Joseph Glaz. A nonparametric repeated significance test with adaptive target sample size. Journal of Statistical Planning and Inference, 137(3):869–878, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001479.

Paez:2008:SVD

[PGLS08] Marina S. Paez, Dani Gamerman, Flávia M. P. F. Landim, and Esther Salazar. Spatially varying dynamic coefficient models. Journal of Statistical Planning and Inference, 138(4):1038–1058, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807002650.

Parsad:2007:DCS

[PGS07] Rajender Parsad, V. K. Gupta, and R. Srivastava. Designs for cropping systems research. Journal of Statistical Planning and Inference, 137(5):1687–1703, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002515.

Papathomas:2005:UBB

[PH05] Michail Papathomas and Anthony O’ Hagan. Updating beliefs for binary variables. Journal of Statistical Planning and Inference, 135(2):324–338, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400237X.

Phillips:2002:PAE

[Phi02] Robert F. Phillips. Partially adaptive estimation of autoregressive processes via a normal mixture. Journal of Statistical Planning and Inference, 105(2):313–326, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002695.

Piepel:2009:DRS

[Pie09] Greg F. Piepel. Discussion of “Response surface design evaluation and comparison” by C. M. Anderson-Cook, C. M. Borror, and D. C. Montgomery. Journal of Statistical Planning and Inference, 139(2):653–656, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001626. See [ACBM09b].

Piperigou:2009:DDE

[Pip09] Violetta E. Piperigou. Discrete distributions in the extended FGM family: the p.g.f. approach. Journal of Statistical Planning and Inference, 139(11):3891–3899, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001554.

Papageorgiou:2001:MCS

[PK01]  Ioulia Papageorgiou and K. X. Karakostas.  Model-

complete strategies for sampling from convex autocorrelated finite populations. Journal of Statistical Planning and Inference, 99(1):71–89, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000763.

Petropoulos:2005:ESP

[PK05] Constantinos Petropoulos and Stavros Kourouklis. Estimation of a scale parameter in mixture models with unknown location. Journal of Statistical Planning and Inference, 128(1):191–218, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003239.

Park:2006:WER

[PK06] Cheolwoo Park and Woo-Chul Kim. Wavelet estimation of a regression function with a sharp change point in a random design. Journal of Statistical Planning and Inference, 136(7):2381–2394, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002296.

Prasad:2008:SES

[PKM08] Anurag Prasad, Debasis Kundu, and Amit Mitra. Sequential estimation of the sum of sinusoidal model parameters. Journal of Statistical Planning and Inference, 138(5):1297–1313, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002339.

Pitrun:2006:SSB

[PKZ06] Ivet Pitrun, Maxwell L. King, and Xibin Zhang. Smoothing spline based tests for non-linearity in a partially linear model. Journal of Statistical Planning and Inference, 136(8):2446–2469, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004161.

Plante:2009:APM

[Pla09] Jean-François Plante. Asymptotic properties of the MAMSE adaptive likelihood weights. Journal of Statistical Planning and Inference, 139(7):2147–2161, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003935.

Park:2006:LLD

[PLK^{+}06] Byeong U. Park, Young Kyung Lee, Tae Yoon Kim, Cheolyong Park, and Shinto Eguchi. On local likelihood density estimation when the bandwidth is large. Journal of Statistical Planning and Inference, 136(3):839–859, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003623.

Park:2008:SBM

[PLLS08] Mira Park, Jae Won Lee, Jung Bok Lee, and Seuck Heun Song. Several biplot methods applied to gene expression data. Journal of Statistical Planning and Inference, 138(2):500–515, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002571.

Prescott:2001:RBI

[PM01] Philip Prescott and Ralph Mansson. Robustness of balanced incomplete block designs to randomly missing observations. Journal of Statistical Planning and Inference, 92(1–2):283–296, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800001476.

Pardo:2008:SPT

[PM08a] L. Pardo and M. L. Menéndez. On some pre-test and Stein-rule phi-divergence test estimators in the independence model of categorical data. Journal of Statistical Planning and Inference, 138(7):2163–2179, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003692.

Parvu:2008:ODT

[PM08b] Valentin Parvu and J. P. Morgan. E-optimal designs for three treatments. Journal of Statistical Planning and Inference, 138(3):642–653, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001887.

Proschan:2001:SAU

[PMS^{+}01] Michael A. Proschan, Robert P. McMahon, Joanna H. Shih, Sally A. Hunsberger, Nancy L. Geller, Genell Knatterud, and Janet Wittes. Sensitivity analysis using an imputation method for missing binary data in clinical trials. Journal of Statistical Planning and Inference, 96(1):155–165, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003323.

Pakes:2003:CUW

[PNRdA03] Anthony G. Pakes, Jorge Navarro, José M. Ruiz, and Yolanda del Aguila. Characterizations using weighted distributions. Journal of Statistical Planning and Inference, 116(2):389–420, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003579.

Pollock:2003:SFS

[Pol03] D. S. G. Pollock. Sharp filters for short sequences. Journal of Statistical Planning and Inference, 113(2):663–683, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802000770.

Pollock:2009:RFS

[Pol09] D. S. G. Pollock. Realisations of finite-sample frequencyselective filters. Journal of Statistical Planning and Inference, 139(4):1541–1558, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003583.

Pommeret:2000:OSS

[Pom00] Denys Pommeret. Orthogonality of the Sheffer system associated to a Levy process. Journal of Statistical Planning and Inference, 86(1):1–10, April 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001585.

Potthoff:2007:PMB

[Pot07] Richard F. Potthoff. Prediction markets, Bayesian priors, and clinical trials. Journal of Statistical Planning and Inference, 137(11):3706–3721, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001292.

Pouet:2001:AOT

[Pou01] Christophe Pouet. An asymptotically optimal test for a parametric set of regression functions against a nonparametric alternative. Journal of Statistical Planning and Inference, 98(1–2):177–189, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003001.

Panholzer:2002:MLN

[PP02a] Alois Panholzer and Helmut Prodinger. Moments of level numbers of leaves in binary trees. Journal of Statistical Planning and Inference, 101(1–2):267–279, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001872.

Paparoditis:2002:LBM

[PP02b] Efstathios Paparoditis and Dimitris N. Politis. The local bootstrap for Markov processes. Journal of Statistical Planning and Inference, 108(1–2):301–328, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003154.

Pardo:2005:NLM

[PP05] L. Pardo and M. C. Pardo. Nonadditivity in loglinear models using ϕ-divergences and MLEs. Journal of Statistical Planning and Inference, 127(1–2):237–252, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003082.

Park:2006:GCC

[PP06a] Chanseok Park and W. J. Padgett. A general class of cumulative damage models for materials failure. Journal of Statistical Planning and Inference, 136(11):3783–3801, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001266.

Pena:2006:NDF

[PP06b] Daniel Peña and Pilar Poncela. Nonstationary dynamic factor analysis. Journal of Statistical Planning and Inference, 136(4):1237–1257, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003659.

Pardo:2009:SCB

[PP09] J. A. Pardo and M. C. Pardo. Selecting and checking a binomial response generalized linear model using phi-divergences. Journal of Statistical Planning and Inference, 139(9):3038–3049, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000378.

Pardo:2006:TLR

[PPdCP06] Julio Angel Pardo, Leandro Pardo, and María del Carmen Pardo. Testing in logistic regression models based on ϕ-divergences measures. Journal of Statistical Planning and Inference, 136(3):982–1006, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003441.

Pepelyshev:2009:UCD

[PPM09] Andrey Pepelyshev, Irene Poli, and Viatcheslav Melas. Uniform coverage designs for mixture experiments. Journal of Statistical Planning and Inference, 139(10):3442–3452, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000871.

Phoa:2009:ASD

[PPX09] Frederick K. H. Phoa, Yu-Hui Pan, and Hongquan Xu. Analysis of supersaturated designs via the Dantzig selector. Journal of Statistical Planning and Inference, 139(7):2362–2372, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004254.

Pardo:2002:MDE

[PPZ02] J. A. Pardo, L. Pardo, and K. Zografos. Minimum ϕdivergence estimators with constraints in multinomial populations. Journal of Statistical Planning and Inference, 104 (1):221–237, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801001136.

Peng:2007:PDC

[PQ07] Liang Peng and Yongcheng Qi. Partial derivatives and confidence intervals of bivariate tail dependence functions. Journal of Statistical Planning and Inference, 137(7):2089–2101, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002011.

Peng:2009:MLE

[PQ09] Liang Peng and Yongcheng Qi. Maximum likelihood estimation of extreme value index for irregular cases. Journal of Statistical Planning and Inference, 139(9):3361–3376, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000780.

Papadatos:2004:BES

[PR04] N. Papadatos and T. Rychlik. Bounds on expectations of L-statistics from without replacement samples. Journal of Statistical Planning and Inference, 124(2):317–336, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002143.

Pena:2006:LDA

[PR06] Daniel Peña and Julio Rodríguez. The log of the determinant of the autocorrelation matrix for testing goodness of fit in time series. Journal of Statistical Planning and Inference, 136(8):2706–2718, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004574.

Pistone:2008:ASL

[PR08a] Giovanni Pistone and Maria-Piera Rogantin. Algebraic statistics of level codings for fractional factorial designs. Journal of Statistical Planning and Inference, 138(1):234–244, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002169.

Pistone:2008:IFC

[PR08b] Giovanni Pistone and Maria-Piera Rogantin. Indicator function and complex coding for mixed fractional factorial designs. Journal of Statistical Planning and Inference, 138(3):787–802, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000638.

Pickle:2008:SPA

[PRBAC08] Stephanie M. Pickle, Timothy J. Robinson, Jeffrey B. Birch, and Christine M. Anderson-Cook. A semi-parametric approach to robust parameter design. Journal of Statistical Planning and Inference, 138(1):114–131, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002030.

Preece:2001:TFE

[Pre01] D. A. Preece. Types of factor in experiments. Journal of Statistical Planning and Inference, 95(1–2):269–282, May 2001.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002949.

Preda:2007:RMF

[Pre07] Cristian Preda. Regression models for functional data by reproducing kernel Hilbert spaces methods. Journal of Statistical Planning and Inference, 137(3):829–840, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001443.

Pronzato:2002:IMR

[Pro02] Luc Pronzato. Information matrices with random regressors. Application to experimental design. Journal of Statistical Planning and Inference, 108(1–2):189–200, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002781.

Prodinger:2004:MDD

[Pro04] Helmut Prodinger. On the moments of a distribution defined by the Gaussian polynomials. Journal of Statistical Planning and Inference, 119(2):237–239, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004226.

Pronzato:2006:SCO

[Pro06] Luc Pronzato. On the sequential construction of optimum bounded designs. Journal of Statistical Planning and Inference, 136(8):2783–2804, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004240.

Pievatolo:2007:P

[PRS07] Antonio M. Pievatolo, Gareth O. Roberts, and Bruno Sansó. Preface. Journal of Statistical Planning and Inference, 137(10):3041–3042, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000869.

Prvan:2002:ABA

[PS02] Tania Prvan and Deborah J. Street. An annotated bibliography of application papers using certain classes of fractional factorial and related designs. Journal of Statistical Planning and Inference, 106(1–2):245–269, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002161.

Peternelli:2003:SSP

[PS03] Luiz A. Peternelli and Carlos H. Osório Silva. A simulation study of a proposed graphical diagnostic for assessing goodness-of-fit. Journal of Statistical Planning and Inference, 112(1–2):185–194, March 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003324.

Pan:2004:TSP

[PS04a] Guohua Pan and Thomas J. Santner. Theory of screening procedures to identify robust product designs using fractional factorial experiments. Journal of Statistical Planning and Inference, 125(1–2):59–84, October 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000564.

Perron:2004:EVB

[PS04b] François Perron and Bimal K. Sinha. Estimation of variance based on a ranked set sample. Journal of Statistical Planning and Inference, 120(1–2):21–28, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004974.

Peng:2005:AML

[PS05] Jie Peng and D. Siegmund. The admixture model in linkage analysis. Journal of Statistical Planning and Inference, 130(1–2):317–324, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804002678.

Pace:2006:APL

[PS06] Luigi Pace and Alessandra Salvan. Adjustments of the profile likelihood from a new perspective. Journal of Statistical Planning and Inference, 136(10):3554–3564, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000881.

Potscher:2009:DAL

[PS09] Benedikt M. Pötscher and Ulrike Schneider. On the distribution of the adaptive LASSO estimator. Journal of Statistical Planning and Inference, 139(8):2775–2790, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000044.

Pena:2007:SIG

[PSG07] Edsel A. Peña, Elizabeth H. Slate, and Juan R. González. Semiparametric inference for a general class of models for recurrent events. Journal of Statistical Planning and Inference, 137(6):1727–1747, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580600173X.

Phillips:2007:LRV

[PSJ07] Peter C. B. Phillips, Yixiao Sun, and Sainan Jin. Long run variance estimation and robust regression testing using sharp origin kernels with no truncation. Journal of Statistical Planning and Inference, 137(3):985–1023, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580600156X.

Pace:2006:LBD

[PSV06] L. Pace, A. Salvan, and L. Ventura. Likelihood-based discrimination between separate scale and regression models. Journal of Statistical Planning and Inference, 136(10):3539–3553, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805000777.

Puntanen:2000:TMB

[PSW00] Simo Puntanen, George P. H. Styan, and Hans Joachim Werner. Two matrix-based proofs that the linear estimator Gy is the best linear unbiased estimator. Journal of Statistical Planning and Inference, 88(2):173–179, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800000768.

Paige:2008:PSS

[PT08] Robert L. Paige and A. Alexandre Trindade. Practical small sample inference for single lag subset autoregressive models. Journal of Statistical Planning and Inference, 138(7):1934– 1949, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002947.

Padmawar:2000:RRR

[PV00] V. R. Padmawar and K. Vijayan. Randomized response revisited. Journal of Statistical Planning and Inference, 90(2):293–304, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800000975.

Pollard:2004:CND

[PvdL04] Katherine S. Pollard and Mark J. van der Laan. Choice of a null distribution in resampling-based multiple testing. Journal of Statistical Planning and Inference, 125(1–2):85–100, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000576.

Park:2009:SAC

[PVHK09] Cheolwoo Park, Amy Vaughan, Jan Hannig, and Kee-Hoon Kang. SiZer analysis for the comparison of time series. Journal of Statistical Planning and Inference, 139(12):3974–3988, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580900127X.

Panzar:2009:NBI

[PvZ09] Laura Panzar and Harry van Zanten. Nonparametric Bayesian inference for ergodic diffusions. Journal of Statistical Planning and Inference, 139(12):4193–4199, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001918.

Perlman:2002:CCT

[PW02a] Michael D. Perlman and Lang Wu. A class of conditional tests for a multivariate one-sided alternative. Journal of Statistical Planning and Inference, 107(1–2):155–171, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002501.

Perlman:2002:DLR

[PW02b] Michael D. Perlman and Lang Wu. A defense of the likelihood ratio test for one-sided and order-restricted alternatives. Journal of Statistical Planning and Inference, 107(1–2):173–186, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002513.

Perlman:2003:VLR

[PW03] Michael D. Perlman and Lang Wu. On the validity of the likelihood ratio and maximum likelihood methods. Journal of Statistical Planning and Inference, 117(1):59–81, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003592.

Pardoe:2006:RDS

[PW06a] Iain Pardoe and Robert R. Weidner. Rejoinder to discussants of “Sentencing convicted felons in the United States: a Bayesian analysis using multilevel covariates”. Journal of Statistical Planning and Inference, 136(4):1456–1461, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004112. See [PW06b,

Bro06, dL06a].

Pardoe:2006:SCF

[PW06b] Iain Pardoe and Robert R. Weidner. Sentencing convicted felons in the United States: a Bayesian analysis using multilevel covariates. Journal of Statistical Planning and Inference, 136(4):1433–1455, April 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004100. See discussion [Bro06,

Bro06, dL06a] and rejoinder [PW06a].

Park:2006:BLL

[PW06c] Jeong-Gun Park and Hulin Wu. Backfitting and local likelihood methods for nonparametric mixed-effects models with longitudinal data. Journal of Statistical Planning and Inference, 136(11):3760–3782, November 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001059.

Paik:2009:HMD

[PW09] Myunghee Cho Paik and Cuiling Wang. Handling missing data by deleting completely observed records. Journal of Statistical Planning and Inference, 139(7):2341–2350, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004734.

Peng:2008:CAD

[PWW08] Hanxiang Peng, Shaoli Wang, and Xueqin Wang. Consistency and asymptotic distribution of the Theil–Sen estimator. Journal of Statistical Planning and Inference, 138(6):1836–1850, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003254.

Pan:2009:LSR

[PX09] Zhi-Wei Pan and Quan-Wu Xiao. Least-square regularized regression with non-iid sampling. Journal of Statistical Planning and Inference, 139(10):3579–3587, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000986.

Pena:2006:DRC

[PY06] Daniel Peña and Victor Yohai. A Dirichlet random coefficient regression model for quality indicators. Journal of Statistical Planning and Inference, 136(3):942–961, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003398.

Peng:2009:JMI

[PY09] Liang Peng and Jingping Yang. Jackknife method for intermediate quantiles. Journal of Statistical Planning and Inference, 139(7):2373–2381, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004230.

Peng:2004:LLS

[PZ04] Liang Peng and Xiao-Hua Zhou. Local linear smoothing of receiver operating characteristic (ROC) curves. Journal of Statistical Planning and Inference, 118(1–2):129–143, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003944.

Peng:2008:NGF

[PZ08] Cheng Peng and Biao Zhang. A note on goodness-of-fit test of continuation ratio logistic regression models under case-control data. Journal of Statistical Planning and Inference, 138(8):2355–2365, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003874.

Pena:2009:BLR

[PZY09] Daniel Peña, Ruben Zamar, and Guohua Yan. Bayesian likelihood robustness in linear models. Journal of Statistical Planning and Inference, 139(7):2196–2207, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004035.

Qi:2007:ANS

[Qi07] Yongcheng Qi. On asymptotic normality of sequential estimators for branching processes with immigration. Journal of Statistical Planning and Inference, 137(9):2892–2902, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806002874.

Qin:2003:CGA

[Qin03] Hong Qin. Characterization of generalized aberration of some designs in terms of their complementary designs. Journal of Statistical Planning and Inference, 117(1):141–151, November 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003610.

Qiu:2001:TSM

[Qiu01] Weisheng Qiu. A theorem for studying the Multiplier Conjecture and its applications. Journal of Statistical Planning and Inference, 94(2):303–312, April 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002615.

Qin:2000:APE

[QJ00] Gengsheng Qin and Bing-Yi Jing. Asymptotic properties for estimation of partial linear models with censored data. Journal of Statistical Planning and Inference, 84(1–2):95–110, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837589900141X.

Qu:2008:MOS

[QKO08] Xianggui Qu, Robert Kushler, and Theophilus Ogunyemi. (m,s)-optimality in selecting factorial designs. Journal of Statistical Planning and Inference, 138(1):65–79, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001978.

Qin:2006:CBU

[QL06] Hong Qin and Dan Li. Connection between uniformity and orthogonality for symmetrical factorial designs. Journal of Statistical Planning and Inference, 136(8):2770–2782, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004604.

Qian:2001:ECF

[QRC01] Jin Qian and D. K. Ray-Chaudhuri. Extremal case of Frankl–Ray–Chaudhuri–Wilson Inequality. Journal of Statistical Planning and Inference, 95(1–2):293–306, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002962.

Qin:2008:CIM

[QRR08] Yongsong Qin, J. N. K. Rao, and Qunshu Ren. Confidence intervals for marginal parameters under imputation for item nonresponse. Journal of Statistical Planning and Inference, 138(8):2283–2302, August 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003813.

Qin:2009:THN

[QSL09] Yongsong Qin, Bruce Smith, and Qingzhu Lei. Test for homogeneity in normal mixtures with unknown means and variances. Journal of Statistical Planning and Inference, 139(12):4165–4178, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001864.

Qiu:2007:CSM

[QT07] Dingxi Qiu and Ajit C. Tamhane. A comparative study of the k-means algorithm and the normal mixture model for clustering: Univariate case. Journal of Statistical Planning and Inference, 137(11):3722–3740, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001309.

Qu:2006:MEC

[Qu06] Xianggui Qu. A maximum estimability criterion for design classification and selection. Journal of Statistical Planning and Inference, 136(8):2756–2769, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004665.

Qu:2007:SPR

[Qu07] Xianggui Qu. Statistical properties of Rechtschaffner designs. Journal of Statistical Planning and Inference, 137(7):2156– 2164, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002035.

Qualite:2008:CCP

[Qua08] Lionel Qualité. A comparison of conditional Poisson sampling versus unequal probability sampling with replacement. Journal of Statistical Planning and Inference, 138(5):1428–1432, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002728.

Quintana:2006:PVB

[Qui06] Fernando A. Quintana. A predictive view of Bayesian clustering. Journal of Statistical Planning and Inference, 136 (8):2407–2429, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804004203.

Qin:2001:ECQ

[QW01] Y. S. Qin and Y. Wu. An estimator of a conditional quantile in the presence of auxiliary information. Journal of Statistical Planning and Inference, 99(1):59–70, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801000799.

Qu:2005:OFT

[QW05] Xianggui Qu and C. F. Jeff Wu. One-factor-at-a-time designs of resolution v. Journal of Statistical Planning and Inference, 131(2):407–416, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804000837.

Qin:2009:REC

[QZF09] Guoyou Qin, Zhongyi Zhu, and Wing K. Fung. Robust estimation of covariance parameters in partial linear model for longitudinal data. Journal of Statistical Planning and Inference, 139(2):558–570, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002036.

Ryabko:2006:UCB

[RA06] Boris Ryabko and Jaakko Astola. Universal codes as a basis for nonparametric testing of serial independence for time series. Journal of Statistical Planning and Inference, 136(12):4119–4128, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805002028.

Raqab:2008:MRL

[RA08] Mohammad Z. Raqab and Majid Asadi. On the mean residual life of records. Journal of Statistical Planning and Inference, 138(12):3660–3666, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001195.

Ralescu:2002:SGS

[Ral02] Stefan Ralescu. On solutions for global Stein optimization problems with applications. Journal of Statistical Planning and Inference, 103(1–2):391–400, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002336.

Rao:2000:BES

[Rao00] B. L. S. Prakasa Rao. Bayes estimation for some stochastic partial differential equations. Journal of Statistical Planning and Inference, 91(2):511–524, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001968.

Rao:2001:TSY

[Rao01] C. Radhakrishna Rao. Two score and 10 years of score tests. Journal of Statistical Planning and Inference, 97(1):3–7, August 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003426.

Rao:2002:DEL

[Rao02a] J. N. K. Rao. Discussion of “Exact linear unbiased estimation in survey sampling” [J. Statist. Plann. Inference 102(1) 25–38 (2002)]. Journal of Statistical Planning and Inference, 102(1):39–40, March 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002701. See [SS02c, SS02d].

Rao:2002:SAB

[Rao02b] Shobha Rao. Statistical appraisal of biological individuality. Journal of Statistical Planning and Inference, 102(1):139–147, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001835.

Rao:2002:MRR

[Rao02c] T. J. Rao. Mean of ratios or ratio of means or both? Journal of Statistical Planning and Inference, 102(1):129–138, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001811.

Raqab:2002:IGE

[Raq02] Mohammad Z. Raqab. Inferences for generalized exponential distribution based on record statistics. Journal of Statistical Planning and Inference, 104(2):339–350, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002464.

Raqab:2004:PNB

[Raq04] Mohammad Z. Raqab. Projection P -norm bounds on the moments of k th record increments. Journal of Statistical Planning and Inference, 124(2):301–315, September 1, 2004.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002040.

Raqab:2005:EBM

[Raq05] Mohammad Z. Raqab. Exact bounds for the mean of total time on test under type II censoring samples. Journal of Statistical Planning and Inference, 134(2):318–331, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002125.

Raqab:2007:SBE

[Raq07] Mohammad Z. Raqab. Sharp bounds on expectations of second record increments from decreasing density families. Journal of Statistical Planning and Inference, 137(4):1291–1301, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000802.

Rawlings:2002:JAM

[Raw02] Don Rawlings. A joint Absorption–Mahonian process. Journal of Statistical Planning and Inference, 101(1–2):281–289, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001902.

Ryan:2007:OSD

[RB07] Kenneth J. Ryan and Dursun A. Bulutoglu. E(s^{2})-optimal supersaturated designs with good minimax properties. Journal of Statistical Planning and Inference, 137(7):2250–2262, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806002230.

Ristic:2009:NGF

[RBN09] Miroslav M. Ristić, Hassan S. Bakouch, and Aleksandar S. Nastić. A new geometric first-order integer-valued autoregressive (NGINAR(1)) process. Journal of Statistical Planning and Inference, 139(7):2218–2226, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800400X.

Rukhin:2000:RML

[RBV00] Andrew L. Rukhin, Brad J. Biggerstaff, and Mark G. Vangel. Restricted maximum likelihood estimation of a common mean and the Mandel–Paule algorithm. Journal of Statistical Planning and Inference, 83(2):319–330, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000981.

Rui:2008:PBP

[RD08] Changxiang Rui and Victor De Oliveira. Point and block prediction in log-Gaussian random fields: the nonconstant mean case. Journal of Statistical Planning and Inference, 138(7):2128–2142, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003370.

Rodrigues:2009:CPC

[RdCCB09] Josemar Rodrigues, Mário de Castro, Vicente G. Cancho, and N. Balakrishnan. COM-Poisson cure rate survival models and an application to a cutaneous melanoma data. Journal of Statistical Planning and Inference, 139(10):3605–3611, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001074.

Ritz:2008:MMA

[RDS08] John Ritz, Eugene Demidenko, and Donna Spiegelman. Multivariate meta-analysis for data consortia, individual patient meta-analysis, and pooling projects. Journal of Statistical Planning and Inference, 138(7):1919–1933, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002959.

Rees:2000:MRD

[Ree00] R. S. Rees. Mandatory representation designs MR(3, k; v) with k even and v odd. Journal of Statistical Planning and Inference, 86(2):567–594, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001305.

Reiter:2005:STM

[Rei05] J. P. Reiter. Significance tests for multi-component estimands from multiply imputed, synthetic microdata. Journal of Statistical Planning and Inference, 131(2):365–377, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000771.

Rogard:2007:EMD

[RGL07] Erwann Rogard, Andrew Gelman, and Hao Lu. Evaluation of multilevel decision trees. Journal of Statistical Planning and Inference, 137(4):1151–1160, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000632.

Roy:2009:CPS

[RGR09] Anindya Roy, Subhashis Ghosal, and William F. Rosenberger. Convergence properties of sequential Bayesian D-optimal designs. Journal of Statistical Planning and Inference, 139(2):425–440, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001857.

Reich:2008:IVC

[RH08] Brian J. Reich and James S. Hodges. Identification of the variance components in the general two-variance linear model. Journal of Statistical Planning and Inference, 138(6):1592– 1604, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580700314X.

Rogers:2004:UPP

[RHL04] James A. Rogers, Jason C. Hsu, and Shili Lin. Using the partitioning principle to construct confidence sets for the location of a disease gene. Journal of Statistical Planning and Inference, 125(1–2):101–118, October 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000588.

Ralescu:2007:P

[RHR07] Stefan Ralescu, George Haiman, and Frits Ruymgaart. Preface. Journal of Statistical Planning and Inference, 137 (3):629–642, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001303.

Ranneby:2005:MSE

[RJT05] Bo Ranneby, S. Rao Jammalamadaka, and Alex Teterukovskiy. The maximum spacing estimation for multivariate observations. Journal of Statistical Planning and Inference, 129 (1–2):427–446, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804002034.

Roy:2005:DCM

[RK05] Anuradha Roy and Ravindra Khattree. On discrimination and classification with multivariate repeated measures data. Journal of Statistical Planning and Inference, 134(2):462–485, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002186.

Rempala:2006:APT

[RL06] Grzegorz A. Rempala and Stephen W. Looney. Asymptotic properties of a two sample randomized test for partially dependent data. Journal of Statistical Planning and Inference, 136(1):68–89, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002496.

Rodriguez-Lallena:2009:CCP

[RL09] José Antonio Rodríguez-Lallena. A class of copulas with piecewise linear horizontal sections. Journal of Statistical Planning and Inference, 139(11):3908–3920, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001578.

Reck:2005:ODI

[RM05a] Brian Reck and J. P. Morgan. Optimal design in irregular BIBD settings. Journal of Statistical Planning and Inference, 129(1–2):59–84, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001545.

Roy:2005:GCL

[RM05b] Anindya Roy and Thomas Mathew. A generalized confidence limit for the reliability function of a two-parameter exponential distribution. Journal of Statistical Planning and Inference, 128(2):509–517, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000023.

Ryabko:2005:UIT

[RM05c] B. Ya. Ryabko and V. A. Monarev. Using information theory approach to randomness testing. Journal of Statistical Planning and Inference, 133(1):95–110, July 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001144.

Rue:2007:ABI

[RM07] HÅvard Rue and Sara Martino. Approximate Bayesian inference for hierarchical Gaussian Markov random field models. Journal of Statistical Planning and Inference, 137(10): 3177–3192, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000845.

Rueda:2007:EDF

[RMMA07] M. Rueda, S. Martínez, H. Martínez, and A. Arcos. Estimation of the distribution function with calibration methods. Journal of Statistical Planning and Inference, 137(2):435–448, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580600036X.

Rueda:2002:BAE

[RMS02] Cristina Rueda, José A. Menéndez, and Bonifacio Salvador. Bootstrap adjusted estimators in a restricted setting. Journal of Statistical Planning and Inference, 107(1– 2):123–131, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802002471.

Rodriguez:2005:ODM

[RO05a] C. Rodríguez and I. Ortiz. D-optimum designs in multifactor models with heteroscedastic errors. Journal of Statistical Planning and Inference, 128(2):623–631, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000278.

Rukhin:2005:NMD

[RO05b] Andrew L. Rukhin and Anna Osmoukhina. Nonparametric measures of dependence for biometric data studies. Journal of Statistical Planning and Inference, 131(1):1–18, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000357.

Robinson:2007:NSE

[Rob07] P. M. Robinson. Nonparametric spectrum estimation for spatial data. Journal of Statistical Planning and Inference, 137(3):1024–1034, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001571.

Robert:2009:ADI

[Rob09a] C. Y. Robert. Asymptotic distributions for the intervals estimators of the extremal index and the clustersize probabilities. Journal of Statistical Planning and Inference, 139(9):3288–3309, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000779.

Roberts:2009:ECI

[Rob09b] R. A. Roberts. Exact conditional inference for two uniform populations. Journal of Statistical Planning and Inference, 139(3):1109–1117, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002942.

Robinson:2009:DRS

[Rob09c] Timothy J. Robinson. A discussion of “Response surface design evaluation and comparison”. Journal of Statistical Planning and Inference, 139(2):669–670, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001663. See [ACBM09b].

Roos:2003:IPA

[Roo03] Bero Roos. Improvements in the Poisson approximation of mixed Poisson distributions. Journal of Statistical Planning and Inference, 113(2):467–483, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802000952.

Rosen:2000:IPO

[Ros00] Bengt Rosén. On inclusion probabilities for order πps sampling. Journal of Statistical Planning and Inference, 90 (1):117–143, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580000104X.

Rosalsky:2007:LBT

[Ros07] Andrew Rosalsky. On the limiting behavior of tail series. Journal of Statistical Planning and Inference, 137(3):935–944, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001522.

Royle:2002:EAC

[Roy02] J. A. Royle. Exchange algorithms for constructing large spatial designs. Journal of Statistical Planning and Inference, 100(2):121–134, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001276.

Roy:2007:BRP

[Roy07] Anindya Roy. Book review: Peter Congdon, Applied Bayesian Modelling, Wiley, New York, ISBN: 0-471-49600-6, 2001 (x + 531 pp., $120.00). Journal of Statistical Planning and Inference, 137(6):2087–2088, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001996.

Rao:2006:IBG

[RRS06] P. S. S. N. V. P. Rao, S. B. Rao, and Bikas K. Sinha. Improved bounds for group testing designs. Journal of Statistical Planning and Inference, 136(1):260–269, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002472.

Rackauskas:2004:HNT

[RS04a] Alfredas Rackauskas and Charles Suquet. Hölder norm test statistics for epidemic change. Journal of Statistical Planning and Inference, 126(2):495–520, December 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002842.

Rempala:2004:MVR

[RS04b] Grzegorz A. Rempala and Sudesh Srivastav. Minimum variance rectangular designs for U-statistics. Journal of Statistical Planning and Inference, 120(1–2):103–118, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802004962.

Rousseeuw:2004:CAS

[RS04c] Peter J. Rousseeuw and Anja Struyf. Characterizing angular symmetry and regression symmetry. Journal of Statistical Planning and Inference, 122(1–2):161–173, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002441.

Ritz:2007:ITF

[RS07a] Christian Ritz and Ib. M. Skovgaard. Influence on tests with focus on linear models. Journal of Statistical Planning and Inference, 137(6):1979–1991, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001741.

Rukhin:2007:SAL

[RS07b] Andrew L. Rukhin and William E. Strawderman. Statistical aspects of linkage analysis in interlaboratory studies. Journal of Statistical Planning and Inference, 137(1):264–278, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002685.

Romano:2008:IIP

[RS08] Joseph P. Romano and Azeem M. Shaikh. Inference for identifiable parameters in partially identified econometric models. Journal of Statistical Planning and Inference, 138(9):2786–2807, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000694.

Ren:2006:BFE

[RSD06] Cuirong Ren, Dongchu Sun, and Dipak K. Dey. Bayesian and frequentist estimation and prediction for exponential distributions. Journal of Statistical Planning and Inference, 136(9):2873–2897, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000376.

Rivas:2005:RTS

[RSM05] M. J. Rivas, M. T. Santos, and D. Morales. Rényi test statistics for partially observed diffusion processes. Journal of Statistical Planning and Inference, 127(1–2):91–102, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803003070.

Rains:2002:LRO

[RSS02] E. M. Rains, N. J. A. Sloane, and John Stufken. The lattice of N-run orthogonal arrays. Journal of Statistical Planning and Inference, 102(2):477–500, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001197.

Ryabko:2004:NTR

[RSS04] B. Ya. Ryabko, V. S. Stognienko, and Yu. I. Shokin. A new test for randomness and its application to some cryptographic problems. Journal of Statistical Planning and Inference, 123(2):365–376, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001496.

Rao:2009:SPE

[RSSR09] C. Radhakrishna Rao, Damodar N. Shanbhag, Theofanis Sapatinas, and M. Bhaskara Rao. Some properties of extreme stable laws and related infinitely divisible random variables. Journal of Statistical Planning and Inference, 139(3):802–813, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002899.

Rasch:2007:HRT

[RTG07] Dieter Rasch, Friedrich Teuscher, and Volker Guiard. How robust are tests for two independent samples? Journal of Statistical Planning and Inference, 137(8):2706–2720, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000225.

Rukhin:2002:ABP

[Ruk02] Andrew L. Rukhin. Asymptotic behavior of posterior distribution of the change-point parameter. Journal of Statistical Planning and Inference, 105(2):327–345, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100266X.

Rukhin:2006:GDO

[Ruk06] Andrew L. Rukhin. Gamma-distribution order statistics, maximal multinomial frequency and randomization designs. Journal of Statistical Planning and Inference, 136(7):2213– 2226, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002193.

Rukhin:2007:NPI

[Ruk07] Andrew L. Rukhin. Non-parametric inference for balanced randomization designs. Journal of Statistical Planning and Inference, 137(3):967–984, March 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001558.

Ruschendorf:2009:DTS

[Rüs09] Ludger Rüschendorf. On the distributional transform, Sklar’s theorem, and the empirical copula process. Journal of Statistical Planning and Inference, 139(11):3921–3927, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580900158X.

Reinsel:2003:RRG

[RV03] Gregory C. Reinsel and Raja P. Velu. Reduced-rank growth curve models. Journal of Statistical Planning and Inference, 114(1–2):107–129, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004664.

Rachdi:2007:NRF

[RV07] M. Rachdi and P. Vieu. Nonparametric regression for functional data: Automatic smoothing parameter selection. Journal of Statistical Planning and Inference, 137(9):2784– 2801, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806002783.

Romero-Villafranca:2007:OEP

[RVZRZ07] R. Romero-Villafranca, L. Zúnica, and R. Romero-Zúnica. D_{s}-optimal experimental plans for robust parameter design. Journal of Statistical Planning and Inference, 137(4): 1488–1495, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806000875.

Rao:2005:LMS

[RW05] C. R. Rao and Y. Wu. Linear model selection by cross-validation. Journal of Statistical Planning and Inference, 128(1):231–240, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803003719.

Raghavarao:2006:DSS

[RW06] Damaraju Raghavarao and James B. Wiley. Design strategies for sequential choice experiments involving economic alternatives. Journal of Statistical Planning and Inference, 136(9):3287–3306, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000261.

Ruan:2004:SMC

[RWF04] Shiling Ruan, Douglas A. Wolfe, and Gerald S. Frankel. Statistical modeling and computer simulation of intergranular corrosion growth in AA2024-T3 aluminum alloy. Journal of Statistical Planning and Inference, 126(2):553–568, December 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002878.

Raghavarao:2003:SPT

[RX03] Damaraju Raghavarao and Yang Xie. Split-plot type cross-over designs. Journal of Statistical Planning and Inference, 116(1):197–207, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802001817.

Rychlik:2001:MVB

[Ryc01] Tomasz Rychlik. Mean-variance bounds for order statistics from dependent DFR, IFR, DFRA and IFRA samples. Journal of Statistical Planning and Inference, 92(1–2):21–38, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800001385.

Rao:2005:RET

[RZ05] Poduri S. R. S. Rao and Nicholas A. Zaino. Reliability estimation through the linear mixed effects model. Journal of Statistical Planning and Inference, 129(1–2):327–337, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001983.

Radavicius:2009:SSS

[RZ09] Marijus Radavicius and Jurgita Zidanaviciute. Semiparametric smoothing of sparse contingency tables. Journal of Statistical Planning and Inference, 139(11):3900–3907, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001566.

Sen:2002:LPC

[SA02] Kanwar Sen and Manju Agarwal. Lattice paths combinatorics applied to transient queue length distribution of C_{2} /M/1 queues and busy period analysis of bulk queues C_{2}^{b} /M/1. Journal of Statistical Planning and Inference, 100(2):365–397, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580100146X.

Salicru:2003:AOA

[SA03] M. Salicrú and A. Arcas. An algorithm to obtain additive trees based upon the preservation of the clusters structure. Journal of Statistical Planning and Inference, 114(1–2):95– 105, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004652.

Soltani:2007:PAD

[SA07] A. R. Soltani and M. Azimmohseni. Periodograms asymptotic distributions in periodically correlated processes and multivariate stationary processes: an alternative approach. Journal of Statistical Planning and Inference, 137(4):1236–1242, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806000772.

Sen:2002:GPE

[SAC02] Kanwar Sen, Manju Lata Agarwal, and Sonali Chakraborty. Generalized Polya-Eggenberger model of order k via lattice path approach. Journal of Statistical Planning and Inference, 102(2):467–476, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001124.

Sagara:2005:NML

[Sag05] Nobusumi Sagara. Nonparametric maximum-likelihood estimation of probability measures: existence and consistency. Journal of Statistical Planning and Inference, 133(2):249–271, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580400165X.

Saha:2008:AOW

[Sah08] Krishna K. Saha. Analysis of one-way layout of count data in the presence of over or under dispersion. Journal of Statistical Planning and Inference, 138(7):2067–2081, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003357.

Sajjan:2000:NCL

[Saj00] S. G. Sajjan. A note on central limit theorems for lattice models. Journal of Statistical Planning and Inference, 83 (2):283–290, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001032.

Sambucini:2007:RPA

[Sam07] Valeria Sambucini. A reference prior for the analysis of a response surface. Journal of Statistical Planning and Inference, 137(4):1119–1128, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000747.

Sanchez:2003:ETU

[Sán03] Ismael Sánchez. Efficient tests for unit roots with prediction errors. Journal of Statistical Planning and Inference, 113(1):341–370, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801003135.

Sarkar:2000:NMC

[Sar00] Sanat K. Sarkar. A note on the monotonicity of the critical values of a step-up test. Journal of Statistical Planning and Inference, 87(2):241–249, June 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002001.

Sarkar:2004:FCS

[Sar04] Sanat K. Sarkar. FDR-controlling stepwise procedures and their false negatives rates. Journal of Statistical Planning and Inference, 125(1–2):119–137, October 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400059X.

Sartori:2006:BPM

[Sar06] Nicola Sartori. Bias prevention of maximum likelihood estimates for scalar skew normal and skew t distributions. Journal of Statistical Planning and Inference, 136(12):4259–4275, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002405.

Sarkar:2008:TSS

[Sar08] Sanat K. Sarkar. Two-stage stepup procedures controlling FDR. Journal of Statistical Planning and Inference, 138 (4):1072–1084, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807002765.

Sasieni:2001:WSC

[Sas01] P. Sasieni. Why should clinicians care about Bayesian methods? Journal of Statistical Planning and Inference, 94(1):61– 63, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002342.

Sathe:2007:PRR

[Sat07] Y. S. Sathe. Professor S. N. Roy: As I remember him. Journal of Statistical Planning and Inference, 137(11):3215, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000900.

Saikali:2002:MPT

[SB02] Khalil G. Saikali and Roger L. Berger. More powerful tests for the sign testing problem. Journal of Statistical Planning and Inference, 107(1–2):187–205, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002525.

Street:2004:ONO

[SB04] Deborah J. Street and Leonie Burgess. Optimal and near-optimal pairs for the estimation of effects in 2-level choice experiments. Journal of Statistical Planning and Inference, 118(1–2):185–199, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003993.

Salibian-Barrera:2005:EVR

[SB05] Matias Salibian-Barrera. Estimating the p-values of robust tests for the linear model. Journal of Statistical Planning and Inference, 128(1):241–257, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003306.

Scott:2006:EAB

[SB06] James G. Scott and James O. Berger. An exploration of aspects of Bayesian multiple testing. Journal of Statistical Planning and Inference, 136(7):2144–2162, July 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002156.

Smythe:2001:EED

[SBKF01] R. T. Smythe, M. G. Bickis, D. Krewski, and K. Y. Fung. Efficiency of experimental designs for comparing two treatments with correlated binary responses. Journal of Statistical Planning and Inference, 99(2):151–173, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000830.

Schneider:2004:PSB

[SC04] U. Schneider and J. N. Corcoran. Perfect sampling for Bayesian variable selection in a linear regression model. Journal of Statistical Planning and Inference, 126(1):153–171, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580300274X.

Salehi:2005:MIS

[SC05a] M. Mohammad Salehi and Kuang-Chao Chang. Multiple inverse sampling in post-stratification with subpopulation sizes unknown: a solution for quota sampling. Journal of Statistical Planning and Inference, 131(2):379–392, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400076X.

Sun:2005:SQE

[SC05b] Shan Sun and Minerva Cordero. Smooth quantile estimators under strong mixing: necessary and sufficient conditions on bandwidth for weak convergence. Journal of Statistical Planning and Inference, 128(2):397–409, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000084.

Shi:2008:DOM

[SC08] Lei Shi and Gemai Chen. Detection of outliers in multilevel models. Journal of Statistical Planning and Inference, 138(10):3189–3199, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000190.

Singh:2009:UHO

[SC09] Sarjinder Singh and Cheng C. Chen. Utilization of higher order moments of scrambling variables in randomized response sampling. Journal of Statistical Planning and Inference, 139(9):3377–3380, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000160.

Schimek:2000:EIP

[Sch00a] Michael G. Schimek. Estimation and inference in partially linear models with smoothing splines. Journal of Statistical Planning and Inference, 91(2):525–540, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000197X.

Schwabe:2000:ODO

[Sch00b] Rainer Schwabe. Optimal designs for a one-way layout with covariates. Journal of Statistical Planning and Inference, 85(1–2):49–59, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000658.

Schott:2001:STE

[Sch01] James R. Schott. Some tests for the equality of covariance matrices. Journal of Statistical Planning and Inference, 94(1):25–36, March 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002093.

Schmitz:2004:RDS

[Sch04a] V. Schmitz. Revealing the dependence structure between X_{(1)} and X_{(n)}. Journal of Statistical Planning and Inference, 123(1):41–47, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803001435.

Schoen:2004:DED

[Sch04b] E. D. Schoen. Dispersion-effects detection after screening for location effects in unreplicated two-level experiments. Journal of Statistical Planning and Inference, 126(1):289–304, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002659.

Schmeidler:2005:IAC

[Sch05a] James Schmeidler. Improved algorithms for closed step-down procedures to test heterogeneity of all subsets of means. Journal of Statistical Planning and Inference, 128(2):519–526, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804000291.

Schoenberg:2005:CPE

[Sch05b] Frederic Paik Schoenberg. Consistent parametric estimation of the intensity of a spatial-temporal point process. Journal of Statistical Planning and Inference, 128(1):79–93, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580300332X.

Schmelter:2007:CGW

[Sch07a] Thomas Schmelter. Considerations on group-wise identical designs for linear mixed models. Journal of Statistical Planning and Inference, 137(12):4003–4010, December 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700167X.

Schott:2007:TEC

[Sch07b] James R. Schott. Testing the equality of correlation matrices when sample correlation matrices are dependent. Journal of Statistical Planning and Inference, 137(6):1992–1997, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001777.

Shao:2009:PCS

[SCW09] Yongwu Shao, R. Dennis Cook, and Sanford Weisberg. Partial central subspace and sliced average variance estimation. Journal of Statistical Planning and Inference, 139(3):952–961, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002760.

Sun:2007:GDL

[SCZ07] Ruibo Sun, Siu Hung Cheung, and Li-Xin Zhang. A generalized drop-the-loser rule for multi-treatment clinical trials. Journal of Statistical Planning and Inference, 137(6): 2011–2023, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001947.

Sathe:2001:ENB

[SD01] Y. S. Sathe and U. J. Dixit. Estimation of P [X ≤ Y] in the negative binomial distribution. Journal of Statistical Planning and Inference, 93(1–2):83–92, February 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002068.

Stanek:2004:UAE

[SdMSL04] Edward J. Stanek III, Julio da Motta Singer, and Viviana Beatriz Lencina. A unified approach to estimation and prediction under simple random sampling. Journal of Statistical Planning and Inference, 121(2):325–338, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001149.

Segers:2001:RE

[Seg01] Johan Segers. Residual estimators. Journal of Statistical Planning and Inference, 98(1–2):15–27, October 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003219.

Segers:2005:GPE

[Seg05] Johan Segers. Generalized Pickands estimators for the extreme value index. Journal of Statistical Planning and Inference, 128(2):381–396, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580300377X.

Seidel:2001:DSD

[Sei01] J. J. Seidel. Definitions for spherical designs. Journal of Statistical Planning and Inference, 95(1–2):307–313, May 2001.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002974.

Seleznjev:2000:SAR

[Sel00] Oleg Seleznjev. Spline approximation of random processes and design problems. Journal of Statistical Planning and Inference, 84(1–2):249–262, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001081.

Sen:2000:HCT

[Sen00] Pranab K. Sen. The Hájek convolution theorem and empirical Bayes estimation: parametrics, semiparametrics and nonparametrics. Journal of Statistical Planning and Inference, 91(2):541–556, December 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001981.

Senn:2001:CTD

[Sen01] Stephen Senn. Cross-over trials in drug development: theory and practice. Journal of Statistical Planning and Inference, 96(1):29–40, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800003189.

Sen:2002:LPA

[Sen02] Kanwar Sen. Lattice path approach to transient analysis of M/G/1/N non-Markovian queues using Cox distributions. Journal of Statistical Planning and Inference, 101(1–2):133–147, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001616.

Sen:2007:UIP

[Sen07a] Pranab Kumar Sen. Union-intersection principle and constrained statistical inference. Journal of Statistical Planning and Inference, 137(11):3741–3752, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001310.

SenGupta:2007:PAI

[Sen07b] Ashis SenGupta. P3 approach to intersection-union testing of hypotheses. Journal of Statistical Planning and Inference, 137(11):3753–3766, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001322.

Sen:2008:BBM

[Sen08] Pranab Kumar Sen. Burden of bioinformatics in medical research: Statistical perspectives and controversies. Journal of Statistical Planning and Inference, 138(2):450–463, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002546.

Serfling:2004:NMD

[Ser04] Robert Serfling. Nonparametric multivariate descriptive measures based on spatial quantiles. Journal of Statistical Planning and Inference, 123(2):259–278, July 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001563.

Serban:2009:CCS

[Ser09] Nicoleta Serban. Clustering confidence sets. Journal of Statistical Planning and Inference, 139(2):109–124, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001298.

Sobel:2002:EGR

[SF02]  Milton Sobel and Krzysztof Frankowski.  Extensions

of Gambler’s ruin with even odds: multinomial and other models. Journal of Statistical Planning and Inference, 101(1–2):291–310, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001914.

Shi:2001:OMC

[SFT01] Peide Shi, Kai-Tai Fang, and Chih-Ling Tsai. Optimal multicriteria designs for Fourier regression models. Journal of Statistical Planning and Inference, 96(2):387–401, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002123.

Sharma:2002:P

[SG02] Bhu Dev Sharma and Sat Gupta. Preface. Journal of Statistical Planning and Inference, 100(2):91, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001240.

Sheena:2005:NEF

[SG05] Yo Sheena and Arjun K. Gupta. New estimator for functions of the canonical correlation coefficients. Journal of Statistical Planning and Inference, 131(1):41–61, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000345.

Salvador:2008:BML

[SG08] Manuel Salvador and Pilar Gargallo. Bayesian monitoring of local residual autocorrelations taking into account the run-length. Journal of Statistical Planning and Inference, 138(12):3885–3898, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001456.

Srivastava:2002:JDM

[SGJM02] Anuj Srivastava, Ulf Grenander, Gary R. Jensen, and Michael I. Miller. Jump-diffusion Markov processes on orthogonal groups for object pose estimation. Journal of Statistical Planning and Inference, 103(1–2):15–37, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001951.

Singh:2009:DRE

[SGK09] Parminder Singh, Amar Nath Gill, and Narendra Kumar. Distribution of ratio of extreme of isotonic estimators of chi-square variables with applications. Journal of Statistical Planning and Inference, 139(9):3262–3269, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000718.

Singh:2005:ETP

[SGU05] Umesh Singh, Pramod K. Gupta, and S. K. Upadhyay. Estimation of three-parameter exponentiated-Weibull distribution under type-II censoring. Journal of Statistical Planning and Inference, 134(2):350–372, October 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002277.

Stone:2002:FKS

[SH02] Charles J. Stone and Jianhua Z. Huang. Free knot splines in concave extended linear modeling. Journal of Statistical Planning and Inference, 108(1–2):219–253, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200280X.

Stone:2003:SMD

[SH03] Charles J. Stone and Jianhua Z. Huang. Statistical modeling of diffusion processes with free knot splines. Journal of Statistical Planning and Inference, 116(2):451–474, October 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580200349X.

Stout:2005:OSD

[SH05] Quentin F. Stout and Janis Hardwick. Optimal screening designs with flexible cost and constraint structures. Journal of Statistical Planning and Inference, 132(1–2):149–162, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001478.

Shen:2006:ELT

[SH06] Junshan Shen and Shuyuan He. Empirical likelihood for truncation parameter in random truncation model. Journal of Statistical Planning and Inference, 136(1):90–107, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003027.

Sethuraman:2009:NBE

[SH09] Jayaram Sethuraman and Myles Hollander. Nonparametric Bayes estimation in repair models. Journal of Statistical Planning and Inference, 139(5):1722–1733, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002474.

Shaman:2008:PGL

[Sha08] Paul Shaman. Properties of generalized Levinson–Durbin-Whittle sequences. Journal of Statistical Planning and Inference, 138(9):2808–2814, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000700.

Shen:2001:IKT

[She01] Hao Shen. Intersections of Kirkman triple systems. Journal of Statistical Planning and Inference, 94(2):313–325, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002627.

Shen:2002:SAS

[She02] Xiaotong Shen. Spline adaptation to smoothness. Journal of Statistical Planning and Inference, 108(1–2):391–409, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802003191.

Shimodaira:2000:IPI

[Shi00] Hidetoshi Shimodaira. Improving predictive inference under covariate shift by weighting the log-likelihood function. Journal of Statistical Planning and Inference, 90(2):227–244, October 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001154.

Shieh:2005:PSS

[Shi05] Gwowen Shieh. On power and sample size calculations for Wald tests in generalized linear models. Journal of Statistical Planning and Inference, 128(1):43–59, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003252.

Shimodaira:2008:TRN

[Shi08] Hidetoshi Shimodaira. Testing regions with nonsmooth boundaries via multiscale bootstrap. Journal of Statistical Planning and Inference, 138(5):1227–1241, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001504.

Shimizu:2009:MSL

[Shi09a] Yasutaka Shimizu. Model selection for Lévy measures in diffusion processes with jumps from discrete observations. Journal of Statistical Planning and Inference, 139(2):516–532, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002140.

Shin:2009:PFL

[Shi09b] Hyejin Shin. Partial functional linear regression. Journal of Statistical Planning and Inference, 139(10):3405–3418, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000585.

Simon:2008:DVB

[Sim08] Richard Simon. Development and validation of biomarker classifiers for treatment selection. Journal of Statistical Planning and Inference, 138(2):308–320, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700242X.

Singpurwalla:2003:KMI

[Sin03] Nozer D. Singpurwalla. Knowledge management and information superiority (a taxonomy). Journal of Statistical Planning and Inference, 115(2):361–364, August 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001647.

Singh:2006:SCI

[Sin06a] Parminder Singh. Simultaneous confidence intervals for the successive ratios of scale parameters. Journal of Statistical Planning and Inference, 136(3):1007–1019, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003428.

Sinha:2006:BRN

[Sin06b] Bimal Sinha. Book review: A Normal Distribution Course, Jürgen Groß. Peter Lang, Europaischer Verlag der Wissenschaften (2004), (xii + 221pp), ISBN: 3-631-52934-1, US ISBN 0-8204-7348-0. Journal of Statistical Planning and Inference, 136(10):3734, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000832.

Sidik:2008:EUN

[SJ08] Kurex Sidik and Jeffrey N. Jonkman. Estimation using non-central hypergeometric distributions in combining 2×2 tables. Journal of Statistical Planning and Inference, 138(12):3993–4005, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000438.

Saha:2009:BLO

[SJ09] Chandan Saha and Michael P. Jones. Bias in the last observation carried forward method under informative dropout. Journal of Statistical Planning and Inference, 139(2):246–255, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001778.

Shimata:2002:SGD

[SK02] Tsutomu Shimata and Sanpei Kageyama. Symmetry of a group divisible design with r = λ_{1} + 1. Journal of Statistical Planning and Inference, 106(1–2):31–37, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001994.

Sinha:2006:CGD

[SK06] Kishore Sinha and Sanpei Kageyama. Constructions of group divisible designs and related combinatorial structures. Journal of Statistical Planning and Inference, 136(9):3281– 3286, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805000315.

Sei:2007:BPM

[SK07a] Tomonari Sei and Fumiyasu Komaki. Bayesian prediction and model selection for locally asymptotically mixed normal models. Journal of Statistical Planning and Inference, 137(7):2523–2534, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002795.

Srivastava:2007:EBR

[SK07b] Muni S. Srivastava and Tatsuya Kubokawa. Empirical Bayes regression analysis with many regressors but fewer observations. Journal of Statistical Planning and Inference, 137(11):3778–3792, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001346.

Shimizu:2008:UNN

[SK08] Shohei Shimizu and Yutaka Kano. Use of non-normality in structural equation modeling: Application to direction of causation. Journal of Statistical Planning and Inference, 138(11):3483–3491, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000918.

Skold:2009:IDE

[Skö09] Martin Sköld. Improving density estimators of discretely observed processes by interpolation. Journal of Statistical Planning and Inference, 139(6):1847–1855, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003431.

Shin:2000:CML

[SL00] Dong Wan Shin and Jong Hyup Lee. Consistency of the maximum likelihood estimators for nonstationary ARMA regressions with time trends. Journal of Statistical Planning and Inference, 87(1):55–68, May 15, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001755.

Sadek:2005:NER

[SL05] Amr Sadek and Nikolaos Limnios. Nonparametric estimation of reliability and survival function for continuoustime finite Markov processes. Journal of Statistical Planning and Inference, 133(1):1–21, July 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001120.

Shao:2009:TPF

[SL09] Xiaofeng Shao and Bo Li. A tuning parameter free test for properties of space-time covariance functions. Journal of Statistical Planning and Inference, 139(12):4031–4038, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001335.

Slaby:2006:ERS

[Sla06] A. Slabý. Expectation of rank statistics under setup of stochastic inequalities. Journal of Statistical Planning and Inference, 136(2):311–319, February 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003064.

Spinelli:2002:TRD

[SLS02] John J. Spinelli, Richard A. Lockhart, and Michael A. Stephens. Tests for the response distribution in a Poisson regression model. Journal of Statistical Planning and Inference, 108(1–2):137–154, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002756.

Srivastava:2001:TRA

[SM01] Deo Kumar Srivastava and Govind S. Mudholkar. Trimmed T^{2}: a robust analog of Hotelling’s T^{2}. Journal of Statistical Planning and Inference, 97(2):343–358, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002391.

Scaccia:2005:TAS

[SM05a] L. Scaccia and R. J. Martin. Testing axial symmetry and separability of lattice processes. Journal of Statistical Planning and Inference, 131(1):19–39, April 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400062X.

Singh:2005:EPM

[SM05b] Housila P. Singh and Nidhi Mathur. Estimation of population mean when coefficient of variation is known using scrambled response technique. Journal of Statistical Planning and Inference, 131(1):135–144, April 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000606.

Sengupta:2006:UVE

[SM06] S. Sengupta and Sujay Mukhuti. Unbiased variance estimation in a simple exponential population using ranked set samples. Journal of Statistical Planning and Inference, 136(4):1526–1553, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804003805.

Singh:2007:CLP

[SM07] Parminder Singh and Satya N. Mishra. Classifying logistic populations using sample medians. Journal of Statistical Planning and Inference, 137(5):1647–1657, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002436.

Smith:2009:CJQ

[Smi09] J. Q. Smith. Comments by J. Q. Smith on Goldstein and Rougier. Journal of Statistical Planning and Inference, 139(3):1246–1247, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003339.

Shevlyakov:2008:RE

[SMS08] Georgy Shevlyakov, Stephan Morgenthaler, and Alexander Shurygin. Redescending M-estimators. Journal of Statistical Planning and Inference, 138(10):2906–2917, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807004120.

Satorra:2003:MEU

[SN03] Albert Satorra and Heinz Neudecker. A matrix equality useful in goodness-of-fit testing of structural equation models. Journal of Statistical Planning and Inference, 114(1–2):63– 80, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004639.

Sun:2004:BAV

[SN04] Dongchu Sun and Shawn Ni. Bayesian analysis of vectorautoregressive models with noninformative priors. Journal of Statistical Planning and Inference, 121(2):291–309, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001162.

Seo:2008:CSC

[SN08] Takashi Seo and Takahiro Nishiyama. On the conservative simultaneous confidence procedures for multiple comparisons among mean vectors. Journal of Statistical Planning and Inference, 138(11):3448–3456, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000888.

Senturk:2009:PCA

[SN09] Damla Sentürk and Danh V. Nguyen. Partial covariate adjusted regression. Journal of Statistical Planning and Inference, 139(2):454–468, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800205X.

Sutradhar:2008:GFT

[SNM08] Santosh C. Sutradhar, Nagaraj K. Neerchal, and Jorge G. Morel. A goodness-of-fit test for overdispersed binomial (or multinomial) models. Journal of Statistical Planning and Inference, 138(5):1459–1471, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002893.

Soedjak:2002:CEB

[Soe02] H. Soedjak. Consistent estimation of the bispectral density function of a harmonizable process. Journal of Statistical Planning and Inference, 100(2):159–170, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001306.

Song:2001:RIL

[Son01] Kai-Sheng Song. Rényi information, loglikelihood and an intrinsic distribution measure. Journal of Statistical Planning and Inference, 93(1–2):51–69, February 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001695.

Sorensen:2008:SSD

[Sør08] Helle Sørensen. Small sample distribution of the likelihood ratio test in the random effects model. Journal of Statistical Planning and Inference, 138(6):1605–1614, July 1, 2008.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003230.

Sanchez:2003:CPG

[SOUS03] Alex Sánchez, Jordi Ocaña, Frederic Utzet, and Lluiís Serra. Comparison of Prevosti genetic distances. Journal of Statistical Planning and Inference, 109(1–2):43–65, January 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002975.

Surles:2005:SPS

[SP05] J. G. Surles and W. J. Padgett. Some properties of a scaled Burr type X distribution. Journal of Statistical Planning and Inference, 128(1):271–280, January 15, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003707.

Spezia:2009:RJL

[Spe09] Luigi Spezia. Reversible jump and the label switching problem in hidden Markov models. Journal of Statistical Planning and Inference, 139(7):2305–2315, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004217.

Shao:2000:GMS

[SR00] Jun Shao and J. Sunil Rao. The GIC for model selection: a hypothesis testing approach. Journal of Statistical Planning and Inference, 88(2):215–231, August 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000080X.

Skinner:2002:JVE

[SR02] C. J. Skinner and J. N. K. Rao. Jackknife variance estimation for multivariate statistics under hot-deck imputation from common donors. Journal of Statistical Planning and Inference, 102(1):149–167, March 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001859.

Smith:2003:ODC

[SR03] D. M. Smith and M. S. Ridout. Optimal designs for criteria involving log(potency) in comparative binary bioassays. Journal of Statistical Planning and Inference, 113(2):617–632, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580200112X.

Smith:2005:AFL

[SR05] D. M. Smith and M. S. Ridout. Algorithms for finding locally and Bayesian optimal designs for binary dose-response models with control mortality. Journal of Statistical Planning and Inference, 133(2):463–478, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001727.

Srivastava:2000:F

[Sri00] Jagdish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 86(2):283, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001330.

Sriram:2001:FSC

[Sri01] T. N. Sriram. Fixed size confidence regions for parameters of threshold AR(1) models. Journal of Statistical Planning and Inference, 97(2):293–304, September 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002469.

Srivastava:2002:DEL

[Sri02a] J. N. Srivastava. Discussion of “Exact linear unbiased estimation in survey sampling” by R. A. Sugden and T. M. F. Smith [J. Statist. Plann. Inference 102(1) 25–38 (2002)]. Journal of Statistical Planning and Inference, 102(1):41–42, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0378375801002713. See [SS02c,

SS02d].

Srivastava:2002:JSJ

[Sri02b] J. N. Srivastava. JSPI: the Silver Jubilee and the 100th volume anniversary. Journal of Statistical Planning and Inference, 102(2):v–vii, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802000757.

Srivastava:2002:F

[Sri02c] Jagdish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 106(1–2):1, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001957.

Srivastava:2002:Pa

[Sri02d] Jagdish N. Srivastava. Preface. Journal of Statistical Planning and Inference, 103(1–2):1, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001938.

Srivastava:2002:Pb

[Sri02e] Jagdish N. Srivastava. Preface. Journal of Statistical Planning and Inference, 108(1–2):1, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002665.

Srivenkataramana:2002:LSI

[Sri02f] T. Srivenkataramana. Location-shifts for improved estimation under Midzuno–Sen sampling scheme. Journal of Statistical Planning and Inference, 102(1):179–187, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001884.

Srivastava:2003:Pb

[Sri03a] Jagdish Srivastava. Preface. Journal of Statistical Planning and Inference, 114(1–2):1–2, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004706.

Srivastava:2003:Pa

[Sri03b] Jagdish N. Srivastava. Preface. Journal of Statistical Planning and Inference, 109(1–2):1, January 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002902.

Srivastava:2004:ACE

[Sri04] Jagdish N. Srivastava. Announcement: Change of Executive Editor. Journal of Statistical Planning and Inference, 119(1):I, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803003665.

Srivastava:2005:FP

[Sri05a] Jagadish N. Srivastava. Foreword (and preface). Journal of Statistical Planning and Inference, 130(1–2):1, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002757.

Srivastava:2005:CHA

[Sri05b] Jagdish N. Srivastava. Combinatorial Hermitian algebras derived from Latin squares. Journal of Statistical Planning and Inference, 129(1–2):305–316, February 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400196X.

Srivastava:2007:FRH

[Sri07a] Jagdish N. Srivastava. Further results on Hermitian algebras derived from latin squares with potential applications to physics and chemistry: Unitary matrices that diagonalize the algebras. Journal of Statistical Planning and Inference, 137(11):3767–3777, November 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001334.

Srivastava:2007:SNR

[Sri07b]  Jagdish N. Srivastava.  Samarendra Nath Roy: an

appreciation. Journal of Statistical Planning and Inference, 137(11):3213–3214, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000894.

Srivastava:2008:F

[Sri08] Jagdish Srivastava. Foreword. Journal of Statistical Planning and Inference, 138(11):3294, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800075X.

Sarkar:2000:API

[SS00] Jyotirmoy Sarkar and Sahadeb Sarkar. Availability of a periodically inspected system under perfect repair. Journal of Statistical Planning and Inference, 91(1):77–90, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800001282. See erra-

tum [TLBJ11].

Solorzano:2001:CMT

[SS01] Eleanne Solorzano and John D. Spurrier. Comparing more than one treatment to more than one control in incomplete blocks. Journal of Statistical Planning and Inference, 97(2):385–398, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580000241X.

Sampson:2002:MMS

[SS02a] Allan R. Sampson and Harshinder Singh. Min and max scorings for two sample partially ordered categorical data. Journal of Statistical Planning and Inference, 107(1–2):219–236, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002549.

Sen:2002:ASA

[SS02b] Pranab K. Sen and Mervyn J. Silvapulle. An appraisal of some aspects of statistical inference under inequality constraints. Journal of Statistical Planning and Inference, 107(1–2):3–43, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802002422.

Sugden:2002:ELU

[SS02c] R. A. Sugden and T. M. F. Smith. Exact linear unbiased estimation in survey sampling. Journal of Statistical Planning and Inference, 102(1):25–38, March 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003256. See discussion [Rao02a,

Sri02a, Tho02] and rejoinder [SS02d].

Sugden:2002:RDE

[SS02d] R. A. Sugden and T. M. F. Smith. Rejoinder to discussion of ‘Exact linear unbiased estimation in survey sampling’. Journal of Statistical Planning and Inference, 102(1):45, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003268. See [SS02c,

Rao02a, Sri02a, Tho02].

Szekely:2002:CDS

[SS02e] Gabor J. Szekely and Pranab K. Sen. Characterization of diagonal symmetry: location unknown, and a test based on allied U-processes. Journal of Statistical Planning and Inference, 102(2):349–358, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001148.

Sinharay:2003:PPM

[SS03] Sandip Sinharay and Hal S. Stern. Posterior predictive model checking in hierarchical models. Journal of Statistical Planning and Inference, 111(1–2):209–221, ???? 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003038.

Shen:2006:IPW

[sS06a] Pao sheng Shen. An inverse-probability-weighted approach to estimation of the bivariate survival function under lefttruncation and right-censoring. Journal of Statistical Planning and Inference, 136(12):4365–4384, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805001382.

Sugden:2006:DSP

[SS06b] R. A. Sugden and T. M. F. Smith. Domains of study and poststratification. Journal of Statistical Planning and Inference, 136(9):3307–3317, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000303.

Stanek:2008:PRE

[SS08] Edward J. Stanek III and Julio M. Singer. Predicting random effects with an expanded finite population mixed model. Journal of Statistical Planning and Inference, 138(10):2991–3004, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807004259.

Smith:2009:EGM

[SS09] Elizabeth L. Smith and Alec G. Stephenson. An extended Gaussian max-stable process model for spatial extremes. Journal of Statistical Planning and Inference, 139 (4):1266–1275, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003145.

Silvapulle:2002:TAI

[SSB02] Mervyn J. Silvapulle, Paramsothy Silvapulle, and Ishwar V. Basawa. Tests against inequality constraints in semiparametric models. Journal of Statistical Planning and Inference, 107(1–2):307–320, September 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002604.

Shao:2007:SCM

[SSC07] Xiaofeng Shao, Michael Stein, and Jason Ching. Statistical comparisons of methods for interpolating the output of a numerical air quality model. Journal of Statistical Planning and Inference, 137(7):2277–2293, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002692.

Singh:2004:ESD

[SSE04] Housila P. Singh, Sharad Saxena, and M. Ruiz Espejo. Estimation of standard deviation in normal parent by shrinkage towards an interval. Journal of Statistical Planning and Inference, 126(2):479–493, December 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002866.

Su:2008:CBR

[SSH08] Jyh-Cherng Su, Nan-Cheng Su, and Wen-Jang Huang. Characterizations based on record values and order statistics. Journal of Statistical Planning and Inference, 138(5):1358– 1367, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807002674.

Sinha:2002:SSR

[SSKS02] K. Sinha, M. K. Singh, S. Kageyama, and R. S. Singh. Some series of rectangular designs. Journal of Statistical Planning and Inference, 106(1–2):39–46, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002008.

Singh:2000:SAS

[SSM00] Sarjinder Singh, Ravindra Singh, and Naurang Singh Mangat. Some alternative strategies to Moors’ model in randomized response sampling. Journal of Statistical Planning and Inference, 83(1):243–255, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000920.

Sanchez:2007:SMA

[SSO07] Alex Sánchez, Miquel Salicrú, and Jordi Ocaña. Statistical methods for the analysis of high-throughput data based on functional profiles derived from the gene ontology. Journal of Statistical Planning and Inference, 137(12):3975–3989, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001656.

Sartori:2003:NDA

[SSP03] Nicola Sartori, Alessandra Salvan, and Luigi Pace. A note on directed adjusted profile likelihoods. Journal of Statistical Planning and Inference, 110(1–2):1–9, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002907.

Shirakura:2002:CME

[SST02] Teruhiro Shirakura, Takeaki Suetsugu, and Takumi Tsuji. Construction of main effect plus two plans for 2^{m} factorials. Journal of Statistical Planning and Inference, 105(2):405–415, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002671.

Sun:2003:NEV

[SSvRM03] Yijun Sun, Bimal K. Sinha, Dietrich von Rosen, and Qingyuan Meng. Nonnegative estimation of variance components in multivariate unbalanced mixed linear models with two variance components. Journal of Statistical Planning and Inference, 115(1):215–234, July 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001507.

Sampson:2009:SCB

[SSW09] Allan R. Sampson, Harshinder Singh, and Lyn R. Whitaker. Simultaneous confidence bands for isotonic functions. Journal of Statistical Planning and Inference, 139(3):828–842, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002127.

Stamey:2005:BAC

[SSY05] J. D. Stamey, J. W. Seaman, and D. M. Young. Bayesian analysis of complementary Poisson rate parameters with data subject to misclassification. Journal of Statistical Planning and Inference, 134(1):36–48, September 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002058.

Stemann:2000:SFR

[ST00] D. Stemann and G. Trenkler. Some further results on the efficiency of the Cochrane-orcutt-estimator. Journal of Statistical Planning and Inference, 88(2):205–214, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000793.

Shiohama:2004:SET

[ST04] Takayuki Shiohama and Masanobu Taniguchi. Sequential estimation for time series regression models. Journal of Statistical Planning and Inference, 123(2):295–312, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001538.

Stallard:2005:PEC

[ST05] Nigel Stallard and Susan Todd. Point estimates and confidence regions for sequential trials involving selection. Journal of Statistical Planning and Inference, 135(2):402–419, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002344.

Shalabh:2006:CDN

[ST06] Shalabh and H. Toutenburg. Consequences of departure from normality on the properties of calibration estimators. Journal of Statistical Planning and Inference, 136(12):4385–4396, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805001989.

Sarami:2008:CQS

[ST08] Chekad Sarami and Vladimir D. Tonchev. Cyclic quasisymmetric designs and self-orthogonal codes of length 63. Journal of Statistical Planning and Inference, 138(1):80–85, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580700198X.

Sostra:2009:ODE

[ST09] Kaja Sõstra and Imbi Traat. Optimal domain estimation under summation restriction. Journal of Statistical Planning and Inference, 139(11):3928–3941, November 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001591.

Stahel:2004:PSR

[Sta04] Werner A. Stahel. Purposes and strategies in regression analysis. Journal of Statistical Planning and Inference, 122(1–2):175–186, ???? 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002313.

Stallard:2006:OAD

[Sta06] Nigel Stallard. Optimal adaptive designs for acute oral toxicity assessment. Journal of Statistical Planning and Inference, 136(6):1781–1799, June 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805001850.

Staicu:2009:HOA

[Sta09a] Ana-Maria Staicu. Higher-order approximations for interval estimation in binomial settings. Journal of Statistical Planning and Inference, 139(10):3393–3404, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900086X.

Stawiarski:2009:STF

[Sta09b] Bartosz Stawiarski. Score test of fit for composite hypothesis in the GARCH(1,1) model. Journal of Statistical Planning and Inference, 139(2):593–616, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002103.

Stefanov:2000:RSB

[Ste00] Valeri T. Stefanov. On run statistics for binary trials. Journal of Statistical Planning and Inference, 87(2):177–185, June 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001913.

Stenger:2002:RAR

[Ste02] Horst Stenger. Regression analysis and random sampling. Journal of Statistical Planning and Inference, 102(1):169–178, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001860.

Stein:2004:EGM

[Ste04a] Michael L. Stein. Equivalence of Gaussian measures for some nonstationary random fields. Journal of Statistical Planning and Inference, 123(1):1–11, June 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001447.

Stepanov:2004:RIB

[Ste04b] A. Stepanov. Random intervals based on records. Journal of Statistical Planning and Inference, 118(1–2):103–113, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003956.

Steland:2005:OSK

[Ste05] Ansgar Steland. Optimal sequential kernel detection for dependent processes. Journal of Statistical Planning and Inference, 132(1–2):131–147, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001466.

Steland:2007:WDF

[Ste07] Ansgar Steland. Weighted Dickey–Fuller processes for detecting stationarity. Journal of Statistical Planning and Inference, 137(12):4011–4030, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700170X.

Stepniak:2009:PLM

[Ste09] Czesław Stepniak. Perfect linear models and perfect parametric functions. Journal of Statistical Planning and Inference, 139(2):159–163, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001535.

Stone:2005:NRF

[Sto05] Charles J. Stone. Nonparametric M-regression with free knot splines. Journal of Statistical Planning and Inference, 130(1–2):183–206, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804002617.

Sakiyama:2002:ATU

[STP02] Kenji Sakiyama, Masanobu Taniguchi, and Madan L. Puri. Asymptotics of tests for a unit root in autoregression. Journal of Statistical Planning and Inference, 108(1–2):351–364, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003178.

Studden:2005:ETR

[Stu05] W. J. Studden. Elfving’s Theorem revisited. Journal of Statistical Planning and Inference, 130(1–2):85–94, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400254X.

Stallard:2008:EFS

[STW08] Nigel Stallard, Susan Todd, and John Whitehead. Estimation following selection of the largest of two normal means. Journal of Statistical Planning and Inference, 138(6):1629–1638, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003138.

Subramanian:2000:EER

[Sub00] Sundarraman Subramanian. Efficient estimation of regression coefficients and baseline hazard under proportionality of conditional hazards. Journal of Statistical Planning and Inference, 84(1–2):81–94, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001536.

Subramanian:2001:PER

[Sub01] Sundarraman Subramanian. Parameter estimation in regression for long-term survival rate from censored data. Journal of Statistical Planning and Inference, 99(2):211–222, December 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801000842.

Subramanian:2003:STM

[Sub03] Sundarraman Subramanian. Semiparametric transformation models and the missing information principle. Journal of Statistical Planning and Inference, 115(1):327–348, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001544. See correction

[Sub04a].

Subramanian:2004:CNS

[Sub04a] Sundarraman Subramanian. A correction note on “Semiparametric transformation models and the missing information principle”. Journal of Statistical Planning and Inference, 126(1):397–399, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002301. See [Sub03].

Subramanian:2004:SRR

[Sub04b] Sundarraman Subramanian. Survival-rate regression using kernel conditional Kaplan–Meier estimators. Journal of Statistical Planning and Inference, 123(1):187–205, June 1, 2004.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300140X.

Sudderth:2001:CAF

[Sud01] W. Sudderth. Comments on “Avoiding foregone conclusions: geometric and foundational analysis of paradoxes of finite additivity”. Journal of Statistical Planning and Inference, 94(1):93–95, March 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002317.

Sukhatme:2002:PVS

[Suk02] Shashikala Sukhatme. Pandurang V. Sukhatme 1911–1997. Journal of Statistical Planning and Inference, 102(1):3–24, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001604.

Sulanke:2002:ND

[Sul02] Robert A. Sulanke. The Narayana distribution. Journal of Statistical Planning and Inference, 101(1–2):311–326, February 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001926.

Sulanke:2005:MNN

[Sul05] Robert A. Sulanke. Moments, Narayana numbers, and the cut and paste for lattice paths. Journal of Statistical Planning and Inference, 135(1):229–244, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000674.

Sundaram:2003:MDE

[Sun03] Rajeshwari Sundaram. Minimum distance estimation in doubly censored two sample scale model. Journal of Statistical Planning and Inference, 115(2):657–681, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200174X.

Sundaram:2006:ETS

[Sun06a] Rajeshwari Sundaram. Estimation in two sample randomly truncated scale model. Journal of Statistical Planning and Inference, 136(9):2983–3006, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000297.

Sundaram:2006:SIP

[Sun06b] Rajeshwari Sundaram. Semiparametric inference for the proportional odds model with time-dependent covariates. Journal of Statistical Planning and Inference, 136(2):320–334, February 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003222.

Sundaram:2009:SIP

[Sun09] Rajeshwari Sundaram. Semiparametric inference of proportional odds model based on randomly truncated data. Journal of Statistical Planning and Inference, 139(4):1381–1393, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808003406.

Shkarupa:2000:CDS

[SV00] E. V. Shkarupa and A. V. Voytishek. Convergence of discretestochastic numerical procedures with independent or weakly dependent estimators at grid nodes. Journal of Statistical Planning and Inference, 85(1–2):199–211, April 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000816.

Singhi:2001:SHW

[SV01a] N. M. Singhi and K. S. Vijayan. S. S. Shrikhande and his work: an appreciation. Journal of Statistical Planning and Inference, 95(1–2):3–7, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002755.

Storme:2001:GDC

[SV01b] L. Storme and S. K. J. Vereecke. The geometric description of a cyclic (q − q + 1)-arc in PG(q − q − 3, q), q square. Journal of Statistical Planning and Inference, 95(1–2):315– 333, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800003062.

Slud:2005:ESE

[SV05] Eric V. Slud and Filia Vonta. Efficient semiparametric estimators via modified profile likelihood. Journal of Statistical Planning and Inference, 129(1–2):339–367, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001995.

Salicru:2005:THD

[SVO05] M. Salicrú, S. Vives, and J. Ocaña. Testing the homogeneity of diversity measures: a general framework. Journal of Statistical Planning and Inference, 132(1–2):117–129, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001454.

Stinson:2000:SFC

[SvTW00] D. R. Stinson, Tran van Trung, and R. Wei. Secure frameproof codes, key distribution patterns, group testing algorithms and related structures. Journal of Statistical Planning and Inference, 86(2):595–617, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001317.

Spezzaferri:2007:BFG

[SVZ07] Fulvio Spezzaferri, Isabella Verdinelli, and Massimo Zeppieri. Bayes factors for goodness of fit testing. Journal of Statistical Planning and Inference, 137(1):43–56, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002661.

Scott:2001:MLG

[SW01] A. J. Scott and C. J. Wild. Maximum likelihood for generalised case-control studies. Journal of Statistical Planning and Inference, 96(1):3–27, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003177.

Schechtman:2004:JTS

[SW04] Edna Schechtman and Suojin Wang. Jackknifing twosample statistics. Journal of Statistical Planning and Inference, 119(2):329–340, February 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802004202.

Shao:2005:CPD

[SW05] Q. Shao and Y. Wu. A consistent procedure for determining the number of clusters in regression clustering. Journal of Statistical Planning and Inference, 135(2):461–476, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002320.

Schick:2007:RCD

[SW07] Anton Schick and Wolfgang Wefelmeyer. Root-n consistent density estimators of convolutions in weighted L_{1}- norms. Journal of Statistical Planning and Inference, 137 (6):1765–1774, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580600200X.

Schick:2008:PMA

[SW08] Anton Schick and Wolfgang Wefelmeyer. Prediction in moving average processes. Journal of Statistical Planning and Inference, 138(3):694–707, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700184X.

Sun:2009:CAG

[SW09] Zhihua Sun and Qihua Wang. Checking the adequacy of a general linear model with responses missing at random. Journal of Statistical Planning and Inference, 139(10):3588–3604, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001062.

Swift:2002:CAA

[Swi02] Randall J. Swift. Covariance analysis and associated spectra for classes of nonstationary processes. Journal of Statistical Planning and Inference, 100(2):145–157, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580100129X.

Shen:2008:NEB

[SY08a] Pao-Sheng Shen and Ya-Fang Yan. Nonparametric estimation of the bivariate survival function with left-truncated and right-censored data. Journal of Statistical Planning and Inference, 138(12):4041–4054, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000487.

Siegmund:2008:MOS

[SY08b] D. O. Siegmund and B. Yakir. Minimax optimality of the Shiryayev–Roberts change-point detection rule. Journal of Statistical Planning and Inference, 138(9):2815–2825, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000712.

Szretter:2009:SIR

[SY09] María Eugenia Szretter and Víctor Jaime Yohai. The sliced inverse regression algorithm as a maximum likelihood procedure. Journal of Statistical Planning and Inference, 139(10):3570–3578, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000974.

Simons:2005:GMM

[SYM05] Gordon Simons, Yi-Ching Yao, and Glenn Morton. Global Markov models for eukaryote nucleotide data. Journal of Statistical Planning and Inference, 130(1–2):251–275, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804002642.

Shi:2007:DMD

[SYZ07] Peilin Shi, Jane Ye, and Julie Zhou. Discrete minimax designs for regression models with autocorrelated MA errors. Journal of Statistical Planning and Inference, 137(8):2721–2731, August 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807000237.

Szczerba-Zubek:2006:ASC

[SZ06] Anna Szczerba-Zubek. Association schemes of conjugate matrices. Journal of Statistical Planning and Inference, 136 (1):270–281, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804003003.

Savani:2007:ADS

[SZ07] V. Savani and A. Zhigljavsky. Asymptotic distributions of statistics and parameter estimates for mixed Poisson processes. Journal of Statistical Planning and Inference, 137(12):3990–4002, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001668.

Sarkar:2008:CBD

[SZ08a] Sanat K. Sarkar and Tianhui Zhou. Controlling Bayes directional false discovery rate in random effects model. Journal of Statistical Planning and Inference, 138(3):682–693, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001838.

Shen:2008:EME

[SZ08b] Haipeng Shen and Zhengyuan Zhu. Efficient mean estimation in log-normal linear models. Journal of Statistical Planning and Inference, 138(3):552–567, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000523.

Szalkai:2002:CSP

[Sza02] Ákos Szalkai. A combinatorial search problem on matroids. Journal of Statistical Planning and Inference, 100(2):303–311, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801001410.

Sun:2008:CIS

[SZW08] Xiaoqian Sun, Xian Zhou, and Jinglong Wang. Confidence intervals for the scale parameter of exponential distribution based on Type II doubly censored samples. Journal of Statistical Planning and Inference, 138(7):2045–2058, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700362X.

Tang:2007:OML

[TAGF07] Yu Tang, Mingyao Ai, Gennian Ge, and Kai-Tai Fang. Optimal mixed-level supersaturated designs and a new class of combinatorial designs. Journal of Statistical Planning and Inference, 137(7):2294–2301, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002552.

Takagi:2009:UAS

[Tak09] Yoshiji Takagi. Universal admissibility in second order asymptotics. Journal of Statistical Planning and Inference, 139(9):2898–2906, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000123.

Tamm:2002:LPT

[Tam02] Ulrich Tamm. Lattice paths not touching a given boundary. Journal of Statistical Planning and Inference, 105(2):433–448, July 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002737.

Taniguchi:2001:LDA

[Tan01a] Masanobu Taniguchi. On large deviation asymptotics of some tests in time series. Journal of Statistical Planning and Inference, 97(1):191–200, August 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003530.

Tanizaki:2001:EUP

[Tan01b] Hisashi Tanizaki. Estimation of unknown parameters in nonlinear and non-Gaussian state-space models. Journal of Statistical Planning and Inference, 96(2):301–323, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002184.

Tang:2007:CRM

[Tan07] Boxin Tang. Construction results on minimum aberration blocking schemes for 2m designs. Journal of Statistical Planning and Inference, 137(7):2355–2361, July 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002722.

Tan:2008:MCI

[Tan08a] Zhiqiang Tan. Monte Carlo integration with Markov chain. Journal of Statistical Planning and Inference, 138(7):1967– 1980, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003333.

Tanaka:2008:DQF

[Tan08b] Katsuto Tanaka. On the distribution of quadratic functionals of the ordinary and fractional Brownian motions. Journal of Statistical Planning and Inference, 138(11):3525–3537, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000955.

Tanil:2009:EOS

[Tan09a] Halil Tanil. Erratum to “An order statistics model based on the list of top m scores after l th change” [J. Statist.

Plann. Inference 139 (2009) 2189–2195]. Journal of Statistical Planning and Inference, 139(9):3392, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001232. See [Tan09b].

Tanil:2009:OSM

[Tan09b] Halil Tanil. An order statistics model based on the list of top m scores after l-th change. Journal of Statistical Planning and Inference, 139(7):2189–2195, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003972. See erratum [Tan09a].

Tardelli:2007:OND

[Tar07] Paola Tardelli. Occupancy numbers for dynamic heterogeneous populations: Estimate of particles lifetimes. Journal of Statistical Planning and Inference, 137(10):3127–3139, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807000808.

Tatsuoka:2005:OSE

[Tat05] Curtis Tatsuoka. Optimal sequencing of experiments in Bayesian group testing. Journal of Statistical Planning and Inference, 133(2):479–488, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001636.

Thangavelu:2007:WMW

[TB07] Karthinathan Thangavelu and Edgar Brunner. Wilcoxon-Mann-Whitney test for stratified samples and Efron’s paradox dice. Journal of Statistical Planning and Inference, 137 (3):720–737, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001376.

Tripodis:2009:PFL

[TB09] Yorghos Tripodis and John P. Buonaccorsi. Prediction and forecasting in linear models with measurement error. Journal of Statistical Planning and Inference, 139(12):4039–4050, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001347.

Traat:2004:SDS

[TBM04] Imbi Traat, Lennart Bondesson, and Kadri Meister. Sampling design and sample selection through distribution theory. Journal of Statistical Planning and Inference, 123(2):395–413, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803001502.

Tsodikov:2008:GSC

[TC08] Alex Tsodikov and Solomon Chefo. Generalized selfconsistency: Multinomial logit model and Poisson likelihood. Journal of Statistical Planning and Inference, 138 (8):2380–2397, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807003734.

Tarantola:2008:BCR

[TCD08] Claudia Tarantola, Guido Consonni, and Petros Dellaportas. Bayesian clustering for row effects models. Journal of Statistical Planning and Inference, 138(7):2223–2235, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003709.

Tang:2003:CGM

[TD03] Boxin Tang and Lih-Yuan Deng. Construction of generalized minimum aberration designs of 3, 4 and 5 factors. Journal of Statistical Planning and Inference, 113(1):335–340, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002956.

Tenreiro:2007:ABL

[Ten07] Carlos Tenreiro. On the asymptotic behaviour of locationscale invariant Bickel–Rosenblatt tests. Journal of Statistical Planning and Inference, 137(1):103–116, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805003046. See erratum [Ten09].

Tenreiro:2009:EAB

[Ten09] Carlos Tenreiro. Erratum to “On the asymptotic behaviour of location-scale invariant Bickel–Rosenblatt tests” [J. Statist. Plann. Inference 137, 103–116, 2007]. Journal of Statistical Planning and Inference, 139(6):2115, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004394. See [Ten07].

Tang:2007:PCD

[TG07a] Yongqiang Tang and Subhashis Ghosal. Posterior consistency of Dirichlet mixtures for estimating a transition density. Journal of Statistical Planning and Inference, 137(6):1711–1726, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001091.

Tokdar:2007:PCL

[TG07b] Surya T. Tokdar and Jayanta K. Ghosh. Posterior consistency of logistic Gaussian process priors in density estimation. Journal of Statistical Planning and Inference, 137(1):34–42, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002648.

Tsai:2007:TLM

[TGM07] Pi-Wen Tsai, Steven G. Gilmour, and Roger Mead. Threelevel main-effects designs exploiting prior information about model uncertainty. Journal of Statistical Planning and Inference, 137(2):619–627, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806000309.

Tack:2002:EBD

[TGV02] Lieven Tack, Peter Goos, and Martina Vandebroek. Efficient Bayesian designs under heteroscedasticity. Journal of Statistical Planning and Inference, 104(2):469–483, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002567.

Thas:2001:FPF

[Tha01] J. A. Thas. Flocks and partial flocks of quadrics: a survey. Journal of Statistical Planning and Inference, 94(2):335– 348, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800002640.

Thall:2008:SGM

[Tha08] Peter F. Thall. Some geometric methods for constructing decision criteria based on two-dimensional parameters. Journal of Statistical Planning and Inference, 138(2):516–527, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002583.

Thompson:2002:CSS

[Tho02] M. E. Thompson. Comment on Sugden and Smith. Journal of Statistical Planning and Inference, 102(1):43, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002725. See [SS02c].

Tsukuma:2006:AEC

[TK06] Hisayuki Tsukuma and Yoshihiko Konno. Alternative estimators of the common regression matrix in two GMANOVA models under weighted quadratic losses. Journal of Statistical Planning and Inference, 136(4):1331–1348, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003933.

Tsukuma:2009:MEN

[TK09] Hisayuki Tsukuma and Tatsuya Kubokawa. Minimax estimation of normal precisions via expansion estimators. Journal of Statistical Planning and Inference, 139(2):295–309, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001882.

Tiwari:2006:NTC

[TKP06] Ram Tiwari, K. B. Kulasekera, and Chanseok Park. Nonparametric tests for cause specific hazard rates with censored data for competing risks among several groups. Journal of Statistical Planning and Inference, 136(5):1718–1745, May 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001771.

Thoresen:2007:SLR

[TL07] Magne Thoresen and Petter Laake. On the simple linear regression model with correlated measurement errors. Journal of Statistical Planning and Inference, 137(1):68–78, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580500265X.

Tsay:2008:CCA

[TL08] Ruey S. Tsay and Shiqing Ling. Canonical correlation analysis for the vector AR(1) model with ARCH innovations. Journal of Statistical Planning and Inference, 138(9):2826–2836, September 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808000724.

Taufer:2009:CFE

[TL09a] Emanuele Taufer and Nikolai Leonenko. Characteristic function estimation of non-Gaussian Ornstein–Uhlenbeck processes. Journal of Statistical Planning and Inference, 139(9):3050–3063, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000366.

Trevezas:2009:VEC

[TL09b] Samis Trevezas and Nikolaos Limnios. Variance estimation in the central limit theorem for Markov chains. Journal of Statistical Planning and Inference, 139(7):2242–2253, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004023.

Tang:2011:EAP

[TLBJ11] Tianqiao Tang, Daming Lin, Dragan Banjevic, and Andrew K. S. Jardine. Erratum to “Availability of a periodically inspected system under perfect repair” [J. Statist.

Plann. Inference 91 (2000) 77–90]. Journal of Statistical Planning and Inference, 141(1):610, January 2011. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810003186. See [SS00].

Torsney:2009:MAC

[TMM09] B. Torsney and R. Martín-Martín. Multiplicative algorithms for computing optimum designs. Journal of Statistical Planning and Inference, 139(12):3947–3961, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001311.

Taniguchi:2006:SAC

[TMP06] Masanobu Taniguchi, Kousuke Maeda, and Madan L. Puri. Statistical analysis of a class of factor time series models. Journal of Statistical Planning and Inference, 136(7):2367– 2380, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805002284.

Terpstra:2005:ORS

[TN05] Jeff T. Terpstra and Eric J. Nelson. Optimal rank set sampling estimates for a population proportion. Journal of Statistical Planning and Inference, 127(1–2):309–321, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803003197.

Toma:2008:MHD

[Tom08] Aida Toma. Minimum Hellinger distance estimators for multivariate distributions from the Johnson system. Journal of Statistical Planning and Inference, 138(3):803–816, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002273.

Tommasi:2009:ODB

[Tom09] C. Tommasi. Optimal designs for both model discrimination and parameter estimation. Journal of Statistical Planning and Inference, 139(12):4123–4132, December 1, 2009.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001682.

Tonellato:2007:RFP

[Ton07] Stefano F. Tonellato. Random field priors for spectral density functions. Journal of Statistical Planning and Inference, 137(10):3164–3176, October 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000833.

Topalova:2001:SDA

[Top01] Svetlana Topalova. Symmetric 2 − (69, 17, 4) designs with automorphisms of order 13. Journal of Statistical Planning and Inference, 95(1–2):335–339, May 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800003074.

Tu:2003:EBA

[TP03] Wanzhu Tu and Walter W. Piegorsch. Empirical Bayes analysis for a hierarchical Poisson generalized linear model. Journal of Statistical Planning and Inference, 111(1–2):235–248, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802003051.

Tanaka:2005:PHM

[TR05] Yoko Tanaka and P. V. Rao. A proportional hazards model for informatively censored survival times. Journal of Statistical Planning and Inference, 129(1–2):253–262, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001909.

Tebbs:2008:NLS

[TR08] Joshua M. Tebbs and Scott A. Roths. New largesample confidence intervals for a linear combination of binomial proportions. Journal of Statistical Planning and Inference, 138(6):1884–1893, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700328X.

Triantafyllopoulos:2008:MSV

[Tri08] K. Triantafyllopoulos. Multivariate stochastic volatility with Bayesian dynamic linear models. Journal of Statistical Planning and Inference, 138(4):1021–1037, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002236.

Terdik:2007:MNR

[TRJ07] Gy. Terdik, T. Subba Rao, and S. Rao Jammalamadaka. On multivariate nonlinear regression models with stationary correlated errors. Journal of Statistical Planning and Inference, 137(11):3793–3814, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001358.

Troendle:2000:SNT

[Tro00] James F. Troendle. Stepwise normal theory multiple test procedures controlling the false discovery rate. Journal of Statistical Planning and Inference, 84(1–2):139–158, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001457.

Tang:2001:ENK

[TS01]  Sheng Tang and Hao Shen.  Embeddings of nearly

Kirkman triple systems. Journal of Statistical Planning and Inference, 94(2):327–333, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002639.

Tian:2006:CST

[TS06] Yongge Tian and George P. H. Styan. Cochran’s statistical theorem revisited. Journal of Statistical Planning and Inference, 136(8):2659–2667, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004550.

Thomas:2008:ELS

[TS08] P. Yageen Thomas and N. V. Sreekumar. Estimation of location and scale parameters of a distribution by U-statistics based on best linear functions of order statistics. Journal of Statistical Planning and Inference, 138(7):2190–2200, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003564.

Tsao:2006:APD

[Tsa06] C. Andy Tsao. Assessing post-data weight of evidence. Journal of Statistical Planning and Inference, 136(11):4012–4025, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805000923.

Tseng:2004:SIM

[Tse04] Yu-Ling Tseng. Set-induced minimax estimators for a multivariate normal mean. Journal of Statistical Planning and Inference, 121(1):53–64, March 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802005050.

Toutenburg:2006:UPI

[TSH06] H. Toutenburg, Shalabh, and C. Heumann. Use of prior information in the form of interval constraints for the improved estimation of linear regression models with some missing responses. Journal of Statistical Planning and Inference, 136(8):2430–2445, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004264.

Tikhomirov:2007:MRS

[TSN07] A. Tikhomirov, T. Stojunina, and V. Nekrutkin. Monotonous random search on a torus: Integral upper bounds for the complexity. Journal of Statistical Planning and Inference, 137(12):4031–4047, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001711.

Tsou:2006:RPR

[Tso06] Tsung-Shan Tsou. Robust Poisson regression. Journal of Statistical Planning and Inference, 136(9):3173–3186, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000285.

Thurston:2003:ERC

[TSR03] Sally W. Thurston, Donna Spiegelman, and David Ruppert. Equivalence of regression calibration methods in main study/external validation study designs. Journal of Statistical Planning and Inference, 113(2):527–539, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003202.

Tamhane:2006:PSS

[TSS06] Ajit C. Tamhane, Kunyang Shi, and Klaus Straßburger. Power and sample size determination for a stepwise test procedure for finding the maximum safe dose. Journal of Statistical Planning and Inference, 136(7):2163–2181, July 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002168.

Tamaki:2007:HOA

[TT07] Kenichiro Tamaki and Masanobu Taniguchi. Higher order asymptotic option valuation for non-Gaussian dependent returns. Journal of Statistical Planning and Inference, 137(3): 1043–1058, March 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001595.

Taniai:2008:SEE

[TT08] Hiroyuki Taniai and Masanobu Taniguchi. Statistical estimation errors of VaR under ARCH returns. Journal of Statistical Planning and Inference, 138(11):3568–3577, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000992.

Tian:2009:CNR

[TT09] Yongge Tian and Yoshio Takane. On consistency, natural restrictions and estimability under classical and extended growth curve models. Journal of Statistical Planning and Inference, 139(7):2445–2458, July 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808004710.

Tang:2006:SSD

[TTC06] Man-Lai Tang, Nian-Sheng Tang, and Vincent J. Carey. Sample size determination for 2-step studies with dichotomous response. Journal of Statistical Planning and Inference, 136(3):1166–1180, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804003581.

Tack:2001:ORO

[TV01] Lieven Tack and Martina Vandebroek. (D_{t}, C)-optimal run orders. Journal of Statistical Planning and Inference, 98(1–2):293–310, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002998.

Tack:2004:BCR

[TV04] Lieven Tack and Martina Vandebroek. Budget constrained run orders in optimum design. Journal of Statistical Planning and Inference, 124(1):231–249, August 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001939.

Tian:2009:CIE

[TVYS09] Lili Tian, Albert Vexler, Li Yan, and Enrique F. Schisterman. Confidence interval estimation of the difference between paired AUCs based on combined biomarkers. Journal of Statistical Planning and Inference, 139(10):3725–3732, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001268.

Thomann:2003:CCA

[TW03] Enrique Thomann and Edward C. Waymire. Contingent claims on assets with conversion costs. Journal of Statistical Planning and Inference, 113(2):403–417, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003172.

Tamhane:2004:P

[TW04] Ajit C. Tamhane and Peter H. Westfall. Preface. Journal of Statistical Planning and Inference, 125(1–2):1, October 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000515.

Tian:2005:CIR

[TW05] Lili Tian and Gregory E. Wilding. Confidence intervals of the ratio of means of two independent inverse Gaussian distributions. Journal of Statistical Planning and Inference, 133(2):381–386, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001648.

Thurston:2005:CRC

[TWH^{+}05] Sally W. Thurston, Paige L. Williams, Russ Hauser, Howard Hu, Mauricio Hernandez-Avila, and Donna Spiegelman. A comparison of regression calibration approaches for designs with internal validation data. Journal of Statistical Planning and Inference, 131(1):175–190, April 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000618.

Tu:2000:PCM

[TZ00] Wanzhu Tu and Xiao-Hua Zhou. Pairwise comparisons of the means of skewed data. Journal of Statistical Planning and Inference, 88(1):59–74, July 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899002062.

Tsai:2005:DRD

[TZ05] Yu-Ling Tsai and Julie Zhou. Discrete M-robust designs for regression models. Journal of Statistical Planning and Inference, 131(2):393–406, May 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000795.

Ucinski:2007:COE

[UB07] Dariusz Uciński and Barbara Bogacka. A constrained optimum experimental design problem for model discrimination with a continuously varying factor. Journal of Statistical Planning and Inference, 137(12):4048–4065, December 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807001723.

Uddin:2000:OCB

[Udd00] Nizam Uddin. Optimal cylindrical-block designs for correlated observations. Journal of Statistical Planning and Inference, 84(1–2):323–336, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899001391.

Uddin:2005:UOS

[Udd05] Nizam Uddin. Universally optimal structurally balanced rowcolumn designs with some empty nodes. Journal of Statistical Planning and Inference, 133(2):509–522, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001715.

Uddin:2007:NPS

[Udd07] Nizam Uddin. Neighbor properties of some classes of BIBRC s and their efficiencies for correlated errors. Journal of Statistical Planning and Inference, 137(5):1669–1680, May 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002497.

Uddin:2008:OBD

[Udd08] Nizam Uddin. Optimal block designs for three treatments when observations are correlated. Journal of Statistical Planning and Inference, 138(7):1960–1966, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003576.

Ueki:2009:BCA

[UF09] M. Ueki and K. Fueda. A bias correction and acceleration approach for the problem of regions. Journal of Statistical Planning and Inference, 139(10):3533–3542, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001025.

Utkin:2002:IRS

[UG02] Lev V. Utkin and Sergey V. Gurov. Imprecise reliability for some new lifetime distribution classes. Journal of Statistical Planning and Inference, 105(1):215–232, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002117.

Ubierna:2007:GFP

[UV07] Andrés Ubierna and Santiago Velilla. A goodness-of-fit process for ARMA (p,q) models based on a modified residual autocorrelation sequence. Journal of Statistical Planning and Inference, 137(9):2903–2919, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002862.

Vangel:2005:NAB

[Van05] Mark G. Vangel. A numerical approach to the Behrens– Fisher problem. Journal of Statistical Planning and Inference, 130(1–2):341–350, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804002691.

Varadarajan:2002:SRA

[Var02] V. S. Varadarajan. Some remarks on arithmetic physics. Journal of Statistical Planning and Inference, 103(1–2):3–13, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580100194X.

Vasdekis:2003:FSC

[Vas03] V. G. S. Vasdekis. Form and selection of covariance adjusted estimators in repeated measures models. Journal of Statistical Planning and Inference, 113(1):239–248, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801003007.

Vannucci:2003:DTA

[VBF03] M. Vannucci, P. J. Brown, and T. Fearn. A decision theoretical approach to wavelet regression on curves with a high number of regressors. Journal of Statistical Planning and Inference, 112(1–2):195–212, March 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003336.

vanBerkum:2005:FFD

[vBHvdVW05] E. E. M. van Berkum, E. J. Habers, P. M. v. d. Ven, and J. T. M. Wijnen. Fractional factorial designs for two-step production processes. Journal of Statistical Planning and Inference, 132(1–2):53–62, June 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001417.

vanBerkum:2004:NCV

[vBLO04] E. E. M. van Berkum, H. N. Linssen, and D. A. Overdijk. New critical values for the average equivalence test. Journal of Statistical Planning and Inference, 121(1):21–35, March 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001083.

vanBerkum:2003:OWP

[vBP03] E. E. M. van Berkum and B. Pauwels. D-optimal weighted paired comparison designs. Journal of Statistical Planning and Inference, 110(1–2):147–157, January 15, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002919.

Vach:2001:SMA

[VdBS01] Werner Vach, Andreas du Bois, and Kirstin Schulz. A stochastic model to analyse risk factors for emesis in multi-cycle chemotherapies. Journal of Statistical Planning and Inference, 96(1):221–246, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003372.

Vegas:2000:EEM

[VdCO00] Esteban Vegas, Joan del Castillo, and Jordi Ocaña. Efficiency and exponential models in a variance-reduction technique for dichotomous response variables. Journal of Statistical Planning and Inference, 85(1–2):61–74, April 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900066X.

vandeGeer:2002:MEU

[vdG02] Sara van de Geer. M-estimation using penalties or sieves. Journal of Statistical Planning and Inference, 108(1–2):55–69, November 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802002707.

vanderHoeven:2005:PSC

[vdH05] Nelly van der Hoeven. The probability to select the correct model using likelihood-ratio based criteria in choosing between two nested models of which the more extended one is true. Journal of Statistical Planning and Inference, 135(2):477–486, December 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580400240X.

vanderLinde:2000:RPS

[vdL00] Angelika van der Linde. Reference priors for shrinkage and smoothing parameters. Journal of Statistical Planning and Inference, 90(2):245–274, October 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001166.

vanderLaan:2003:NAH

[vdLP03] Mark J. van der Laan and Katherine S. Pollard. A new algorithm for hybrid hierarchical clustering with visualization and the bootstrap. Journal of Statistical Planning and Inference, 117(2):275–303, December 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003889.

vandeVen:2008:ETE

[vdV08] P. M. van de Ven. On the equivalence of three estimators for dispersion effects in unreplicated two-level factorial designs. Journal of Statistical Planning and Inference, 138(1):18–29, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807001929.

vandeVen:2009:EDR

[vdVD09] P. M. van de Ven and A. Di Bucchianico. On the equivalence of definitions for regular fractions of mixedlevel factorial designs. Journal of Statistical Planning and Inference, 139(7):2351–2361, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004229.

vandeWiel:2004:END

[vdW04] M. A. van de Wiel. Exact null distributions of quadratic distribution-free statistics for two-way classification. Journal of Statistical Planning and Inference, 120(1–2):29–40, February 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004986.

vandeWiel:2001:FCE

[vdWD01] M. A. van de Wiel and A. Di Bucchianico. Fast computation of the exact null distribution of Spearman’s ρ and Page’s L statistic for samples with and without ties. Journal of Statistical Planning and Inference, 92(1–2):133–145, January 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580000166X.

vanEs:2000:SAU

[vEKO00] Bert van Es, Chris A. J. Klaassen, and Karin Oudshoorn. Survival analysis under cross-sectional sampling: length bias and multiplicative censoring. Journal of Statistical Planning and Inference, 91(2):295–312, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580000183X.

Velasco:2003:NFD

[Vel03a] Carlos Velasco. Nonparametric frequency domain analysis of nonstationary multivariate time series. Journal of Statistical Planning and Inference, 116(1):209–247, September 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802002355.

Vellaisamy:2003:QES

[Vel03b] P. Vellaisamy. Quantile estimation of the selected exponential population. Journal of Statistical Planning and Inference, 115(2):461–470, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802001568.

Vasas:2007:TSR

[VEM07] Krisztina Vasas, Péter Elek, and László Márkus. A twostate regime switching autoregressive model with an application to river flow analysis. Journal of Statistical Planning and Inference, 137(10):3113–3126, October 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000791.

Veraverbeke:2001:EQD

[Ver01] Noël Veraverbeke. Estimation of the quantiles of the duration of old age. Journal of Statistical Planning and Inference, 98(1–2):101–106, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003116.

Verma:2002:CMC

[Ver02] Vijay Verma. Comparability in multi-country survey programmes. Journal of Statistical Planning and Inference, 102(1):189–210, March 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375801001896.

Vexler:2006:GTE

[Vex06] A. Vexler. Guaranteed testing for epidemic changes of a linear regression model. Journal of Statistical Planning and Inference, 136(9):3101–3120, September 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004732.

Vilar-Fernandez:2002:BMD

[VFVF02] Juan M. Vilar-Fernández and José A. Vilar-Fernández. Bootstrap of minimum distance estimators in regression with correlated disturbances. Journal of Statistical Planning and Inference, 108(1–2):283–299, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003142.

Vlachos:2003:CBM

[VG03] Pantelis K. Vlachos and Alan E. Gelfand. On the calibration of Bayesian model choice criteria. Journal of Statistical Planning and Inference, 111(1–2):223–234, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200304X.

Vila:2007:ODB

[VG07] Jean-Pierre Vila and Jean-Pierre Gauchi. Optimal designs based on exact confidence regions for parameter estimation of a nonlinear regression model. Journal of Statistical Planning and Inference, 137(9):2935–2953, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807000420.

Vidal:2006:BSA

[VIBAV06] I. Vidal, P. Iglesias, M. D. Branco, and R. B. Arellano-Valle. Bayesian sensitivity analysis and model comparison for skew elliptical models. Journal of Statistical Planning and Inference, 136(10):3435–3457, October 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805000728.

Vidoni:2006:RPM

[Vid06] Paolo Vidoni. Response prediction in mixed effects models. Journal of Statistical Planning and Inference, 136(11): 3948–3966, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375805001047.

Vidoni:2008:IPM

[Vid08] Paolo Vidoni. Improved predictive model selection. Journal of Statistical Planning and Inference, 138(12):3713–3721, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001286.

Viertl:2009:REB

[Vie09] Reinhard Viertl. On reliability estimation based on fuzzy lifetime data. Journal of Statistical Planning and Inference, 139(5):1750–1755, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002498.

Vinod:2002:LCC

[Vin02] H. D. Vinod. A looser cointegration concept using fractional integration parameters and quantification of market responsiveness. Journal of Statistical Planning and Inference, 100(2):399–410, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801001471.

Viskek:2003:LRM

[Viš03] Tomáš Viškek. The likelihood ratio method for testing changes in the parameters of double exponential observations.

Journal of Statistical Planning and Inference, 113(1):79–111, April 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801003081.

Visuri:2000:SRC

[VKO00] Samuli Visuri, Visa Koivunen, and Hannu Oja. Sign and rank covariance matrices. Journal of Statistical Planning and Inference, 91(2):557–575, December 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001993.

VanDorp:2004:GBE

[VM04] J. René Van Dorp and Thomas A. Mazzuchi. A general Bayes exponential inference model for accelerated life testing. Journal of Statistical Planning and Inference, 119(1):55–74, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802004111.

Visuri:2003:AEM

[VOK^{+}03] S. Visuri, E. Ollila, V. Koivunen, J. Möttönen, and H. Oja. Affine equivariant multivariate rank methods. Journal of Statistical Planning and Inference, 114(1–2):161–185, June 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580200469X.

Velu:2008:SUR

[VR08] Raja Velu and Joseph Richards. Seemingly unrelated reducedrank regression model. Journal of Statistical Planning and Inference, 138(9):2837–2846, September 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000736.

Viele:2000:PEM

[VS00] Kert Viele and C. Srinivasan. Parsimonious estimation of multiplicative interaction in analysis of variance using Kullback–Leibler information. Journal of Statistical Planning and Inference, 84(1–2):201–219, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001512.

vanTrung:2001:RCD

[vT01] Tran van Trung. Recursive constructions for 3-designs and resolvable 3-designs. Journal of Statistical Planning and Inference, 95(1–2):341–358, May 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800003086.

Vasdekis:2005:DBM

[VV05] Vassilis G. S. Vasdekis and Ioannis G. Vlachonikolis. On the difference between ML and REML estimators in the modelling of multivariate longitudinal data. Journal of Statistical Planning and Inference, 134(1):194–205, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001788.

Vellaisamy:2008:CRC

[VV08] P. Vellaisamy and V. Vijay. Collapsibility of regression coefficients and its extensions. Journal of Statistical Planning and Inference, 138(4):982–994, April 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001814.

Wu:2007:AEM

[WA07] Samuel S. Wu and David H. Annis. Asymptotic efficiency of majority rule relative to rank-sum method for selecting the best population. Journal of Statistical Planning and Inference, 137(6):1838–1850, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001145.

Wahed:2009:ESQ

[Wah09] Abdus S. Wahed. Estimation of survival quantiles in twostage randomization designs. Journal of Statistical Planning and Inference, 139(6):2064–2075, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003790.

Walley:2002:RFP

[Wal02] Peter Walley. Reconciling frequentist properties with the likelihood principle. Journal of Statistical Planning and Inference, 105(1):35–65, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801002038.

Wall:2004:CLS

[Wal04] Melanie M. Wall. A close look at the spatial structure implied by the CAR and SAR models. Journal of Statistical Planning and Inference, 121(2):311–324, April 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803001113.

Wang:2005:ICE

[Wan05] Hsiuying Wang. Improved confidence estimators for confidence sets of location parameters. Journal of Statistical Planning and Inference, 128(1):95–107, January 15, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003288.

Wang:2006:SCI

[Wan06] Weizhen Wang. Smallest confidence intervals for one binomial proportion. Journal of Statistical Planning and Inference, 136(12):4293–4306, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002430.

Wang:2008:PDE

[Wan08] Qihua Wang. Probability density estimation with data missing at random when covariables are present. Journal of Statistical Planning and Inference, 138(3):568–587, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003345.

Wang:2009:RRV

[Wan09] Hansheng Wang. Rank reducible varying coefficient model. Journal of Statistical Planning and Inference, 139(3):999– 1011, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002802.

Warton:2008:WWS

[War08] David I. Warton. Which Wald statistic? Choosing a parameterization of the Wald statistic to maximize power in k-sample generalized estimating equations. Journal of Statistical Planning and Inference, 138(10):3269–3282, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000311.

Wulff:2005:BUE

[WB05] Shaun S. Wulff and David Birkes. Best unbiased estimation of fixed effects in mixed ANOVA models. Journal of Statistical Planning and Inference, 133(2):305–313, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001685.

Wang:2007:DAE

[WBvdL07] Yue Wang, Oliver Bembom, and Mark J. van der Laan. Dataadaptive estimation of the treatment-specific mean. Journal of Statistical Planning and Inference, 137(6):1871–1887, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001182.

Wong:2000:EMV

[WC00] Chi Song Wong and Hua Cheng. Evaluating matrixvariate moments through higher-order differential forms and combinatorial algorithms. Journal of Statistical Planning and Inference, 91(1):1–21, November 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001809.

Wong:2001:EGC

[WC01] Chi Song Wong and Hua Cheng. Estimation in a growth curve model with singular covariance. Journal of Statistical Planning and Inference, 97(2):323–342, September 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002202.

Wen:2007:OSD

[WC07] Xuerong Wen and R. Dennis Cook. Optimal sufficient dimension reduction in regressions with categorical predictors. Journal of Statistical Planning and Inference, 137(6):1961–1978, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375806001765.

Wu:2008:EML

[WC08] Tiee-Jian Wu and Huang-Yu Chen. Estimation in multiple linear regression Berkson model for processes with uncorrelated increments. Journal of Statistical Planning and Inference, 138(4):827–833, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000535.

Walker:2009:TSC

[WC09a] Robert A. Walker and Charles J. Colbourn. Tabu search for covering arrays using permutation vectors. Journal of Statistical Planning and Inference, 139(1):69–80, January 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002310.

Wen:2009:NAM

[WC09b] Xuerong Meggie Wen and R. Dennis Cook. New approaches to model-free dimension reduction for bivariate regression. Journal of Statistical Planning and Inference, 139(3):734–748, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808002668.

Wang:2001:SSI

[WCS01] Yazhen Wang, Joseph E. Cavanaugh, and Changyong Song. Self-similarity index estimation via wavelets for locally self-similar processes. Journal of Statistical Planning and Inference, 99(1):91–110, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375801000751.

Wei:2004:EBE

[WD04] Laisheng Wei and Xiao Ding. On empirical Bayes estimation of variance components in random effects model. Journal of Statistical Planning and Inference, 123(2):347–364, July 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803001526.

Wang:2005:NVE

[WDE05] Suojin Wang, Alan H. Dorfman, and Lawrence R. Ernst. A new variance estimator for a two-phase restratified sample with PPS sampling at both phases. Journal of Statistical Planning and Inference, 129(1–2):207–227, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001831.

Weba:2007:OPC

[Web07] Michael Weba. Optimal prediction of compound mixed Poisson processes. Journal of Statistical Planning and Inference, 137(4):1332–1342, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001030.

Weiss:2008:SDR

[Wei08] Christian H. Weiß. Serial dependence and regression of Poisson INARMA models. Journal of Statistical Planning and Inference, 138(10):2975–2990, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580700403X.

West:2000:BRS

[Wes00] Webster West. Book review: Simulation: a modeler’s approach: J. Thompson; Wiley, New York, 306 pages, price US$84.95, ISBN 0-471-25184-4. Journal of Statistical Planning and Inference, 91(1):169–170, November 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000902.

Wesolowski:2001:LCM

[WG01] Jacek Wesolowski and Arjun K. Gupta. Linearity of convex mean residual life. Journal of Statistical Planning and Inference, 99(2):183–191, December 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801000891.

Wang:2003:BAB

[WG03a] Chen-Pin Wang and Malay Ghosh. Bayesian analysis of bivariate competing risks models with covariates. Journal of Statistical Planning and Inference, 115(2):441–459, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001775.

Wu:2003:LST

[WG03b] Y. Wu and A. K. Gupta. Local score tests in mixture exponential family. Journal of Statistical Planning and Inference, 116(2):421–435, October 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802003580.

Wang:2005:ODS

[WG05] Ming-Dauh Wang and Seymour Geisser. Optimal dichotomization of screening test variables. Journal of Statistical Planning and Inference, 131(1):191–206, April 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804000667.

Wang:2006:MSS

[WG06] Jing Wang and Marc G. Genton. The multivariate skewslash distribution. Journal of Statistical Planning and Inference, 136(1):209–220, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002782.

Wong:2009:LWE

[WGCI09] Heung Wong, Shaojun Guo, Min Chen, and Wai-Cheung Ip. On locally weighted estimation and hypothesis testing of varying-coefficient models with missing covariates. Journal of Statistical Planning and Inference, 139(9):2933– 2951, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809000275.

Wang:2009:DMC

[WGCX09] Xueli Wang, Zhi Geng, Hua Chen, and Xianchao Xie. Detecting multiple confounders. Journal of Statistical Planning and Inference, 139(3):1073–1081, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002929.

Wang:2001:NUT

[WH01] Weizhen Wang and J. T. Gene Hwang. A nearly unbiased test for individual bioequivalence problems using probability criteria. Journal of Statistical Planning and Inference, 99(1):41–58, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580100074X.

Wang:2002:LDM

[WH02] Yazhen Wang and Jian Huang. Limiting distribution for monotone median regression. Journal of Statistical Planning and Inference, 107(1–2):281–287, September 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002586.

Wang:2005:ESB

[WH05] Minghui Wang and Franz Hering. Efficiency of split-block designs versus split-plot designs for hypothesis testing. Journal of Statistical Planning and Inference, 132(1–2):163–182, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580400148X.

Wang:2008:RMP

[WH08] Molin Wang and John J. Hanfelt. Robust modified profile estimating function with application to the generalized estimating equation. Journal of Statistical Planning and Inference, 138(7):2029–2044, July 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807003631.

Wu:2009:HVE

[WH09] Hong-Dar Isaac Wu and Fushing Hsieh. Heterogeneity and varying effect in hazards regression. Journal of Statistical Planning and Inference, 139(12):4213–4222, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809001979.

Whittle:2005:TAS

[Whi05] Peter Whittle. Three approaches to sequential analysis and one to hidden Markov processes. Journal of Statistical Planning and Inference, 130(1–2):49–62, March 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001259.

Wang:2007:BPE

[WHS07] Xiaoyin Wang, Chong Z. He, and Dongchu Sun. Bayesian population estimation for small sample capture-recapture data using noninformative priors. Journal of Statistical Planning and Inference, 137(4):1099–1118, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000991.

Wiens:2000:RWD

[Wie00] Douglas P. Wiens. Robust weights and designs for biased regression models: Least squares and generalized M-estimation. Journal of Statistical Planning and Inference, 83(2):395–412, February 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375899001020.

Wiener:2002:SME

[Wie02] Gábor Wiener. Search for a majority element. Journal of Statistical Planning and Inference, 100(2):313–318, February 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001422.

Wiens:2005:RAS

[Wie05] Douglas P. Wiens. Robust allocation schemes for clinical trials with prognostic factors. Journal of Statistical Planning and Inference, 127(1–2):323–340, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300291X.

Wilson:2003:TCB

[Wil03] Greg Wilson. Tides of change: is Bayesianism the new paradigm in statistics? Journal of Statistical Planning and Inference, 113(1):371–374, April 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801003068.

Wilczynski:2005:MEL

[Wil05] Maciej Wilczyński. Minimax estimation in the linear model with a relative squared error. Journal of Statistical Planning and Inference, 127(1–2):205–212, January 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002945.

Wilczynski:2007:MEL

[Wil07] Maciej Wilczyński. Minimax estimation in linear regression with ellipsoidal constraints. Journal of Statistical Planning and Inference, 137(1):79–86, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002995.

Winterhof:2000:NEG

[Win00] Arne Winterhof. On the non-existence of generalized Hadamard matrices. Journal of Statistical Planning and Inference, 84(1–2):337–342, March 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0378375899001470.

Witkovsky:2001:ECD

[Wit01] Viktor Witkovský. On the exact computation of the density and of the quantiles of linear combinations of t and F random variables. Journal of Statistical Planning and Inference, 94(1):1–13, March 1, 2001. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002081.

Watkins:2008:CSA

[WJ08] A. J. Watkins and A. M. John. On constant stress accelerated life tests terminated by Type II censoring at one of the stress levels. Journal of Statistical Planning and Inference, 138(3):768–786, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001851.

Wang:2009:OCD

[WJ09] Jixian Wang and Byron Jones. Optimal cross-over designs for nonlinear mixed models using a first-order expansion. Journal of Statistical Planning and Inference, 139(2):203–212, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808001420.

Westfall:2001:OWF

[WK01] Peter H. Westfall and Alok Krishen. Optimally weighted, fixed sequence and gatekeeper multiple testing procedures. Journal of Statistical Planning and Inference, 99(1):25–40, November 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801000775.

Weinberg:2001:LRT

[WL01] Janice M. Weinberg and Stephen W. Lagakos. Linear rank tests under general alternatives, with application to summary statistics computed from repeated measures data. Journal of Statistical Planning and Inference, 96(1):109–127, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003281.

Wang:2002:DFP

[WL02] Yining Wang and Gang Li. A dose finding procedure based on the multivariate likelihood ratio test. Journal of Statistical Planning and Inference, 104(2):417–425, June 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002440.

Wu:2008:OSS

[WLC08] Shuo-Jye Wu, Ying-Po Lin, and Shyi-Tien Chen. Optimal step-stress test under type I progressive group-censoring with random removals. Journal of Statistical Planning and Inference, 138(4):817–826, April 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807000547.

Wong:2009:ELB

[WLCI09] Heung Wong, Feng Liu, Min Chen, and Wai Cheung Ip. Empirical likelihood based diagnostics for heteroscedasticity in partially linear errors-in-variables models. Journal of Statistical Planning and Inference, 139(3):916–929, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800267X.

Wang:2002:ETL

[WM02] Song-Gui Wang and Wen-Qing Ma. On exact tests of linear hypothesis in linear models with nested error structure. Journal of Statistical Planning and Inference, 106(1–2):225–233, August 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580200215X.

Wilding:2008:GGF

[WM08] Gregory E. Wilding and Govind S. Mudholkar. A gamma goodness-of-fit test based on characteristic independence of the mean and coefficient of variation. Journal of Statistical Planning and Inference, 138(12):3813–3821, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808000281.

Weller:2007:RCL

[WMES07] Edie A. Weller, Donald K. Milton, Ellen A. Eisen, and Donna Spiegelman. Regression calibration for logistic regression with multiple surrogates for one exposure. Journal of Statistical Planning and Inference, 137(2):449–461, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000589.

Wilding:2007:TSI

[WMK07] Gregory E. Wilding, Govind S. Mudholkar, and Georgia D. Kollia. Two sets of isotones for comparing tests of exponentiality. Journal of Statistical Planning and Inference, 137(11):3815–3825, November 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580700136X.

Wang:2006:ODP

[WMSY06] Yanping Wang, Raymond H. Myers, Eric P. Smith, and Keying Ye. D-optimal designs for Poisson regression models. Journal of Statistical Planning and Inference, 136(8):2831–2845, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004197.

Wu:2006:AML

[WN06] Qi-Guang Wu and Kazuo Noda. Admissibilities of matrix linear estimators multivariate linear models. Journal of Statistical Planning and Inference, 136(11):3852–3870, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375805001333.

Wan:2000:MMS

[WO00]  Alan T. K. Wan and Kazuhiro Ohtani.  Minimum

mean-squared error estimation in linear regression with an inequality constraint. Journal of Statistical Planning and Inference, 86(1):157–173, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837589900172X.

Walley:2004:DAC

[WPV04] Peter Walley, Renato Pelessoni, and Paolo Vicig. Direct algorithms for checking consistency and making inferences from conditional probability assessments. Journal of Statistical Planning and Inference, 126(1):119–151, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375803002751.

Welsh:2002:JSS

[WR02] A. H. Welsh and Elvezio Ronchetti. A journey in single steps: robust one-step M-estimation in linear regression. Journal of Statistical Planning and Inference, 103(1–2):287–310, April 15, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002282.

Wright:2008:TDB

[Wri08] James H. Wright. Two-dimensional balanced sampling plans excluding adjacent units. Journal of Statistical Planning and Inference, 138(1):145–153, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002017.

Wang:2005:DSB

[WRV05] Minghui Wang, Dieter Rasch, and Rob Verdooren. Determination of the size of a balanced experiment in mixed ANOVA models using the modified approximate F -test. Journal of Statistical Planning and Inference, 132(1–2):183–201, June 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001491.

Wright:2008:NBS

[WS08] James H. Wright and John Stufken. New balanced sampling plans excluding adjacent units. Journal of Statistical Planning and Inference, 138(11):3326–3335, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000803.

Wang:2009:IBS

[WS09] B. Wang and J. Sun. Inferences from biased samples with a memory effect. Journal of Statistical Planning and Inference, 139(2):441–453, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002073.

Wang:2006:SDP

[WSY06] Yanping Wang, Eric P. Smith, and Keying Ye. Sequential designs for a Poisson regression model. Journal of Statistical Planning and Inference, 136(9):3187–3202, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805000273.

Wang:2006:NUE

[WT06] Jinfang Wang and Nobuhiro Taneichi. Normalizing unbiased estimating functions. Journal of Statistical Planning and Inference, 136(3):689–704, March 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580400357X.

Weng:2008:APN

[WT08] Ruby C. Weng and Wen-Chi Tsai. Asymptotic posterior normality for multiparameter problems. Journal of Statistical Planning and Inference, 138(12):4068–4080, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001493.

Wu:2000:SRP

[Wu00] Qi-Guang Wu. Some results on parameter estimation in extended growth curve models. Journal of Statistical Planning and Inference, 88(2):285–300, August 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800000847.

Wu:2003:ODC

[Wu03] Huaiqing Wu. D-optimal designs for combined linear and trigonometric regression. Journal of Statistical Planning and Inference, 116(1):177–184, September 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001854.

Wu:2006:IPC

[Wu06] Yanhong Wu. Inference for post-change mean by a CUSUM procedure. Journal of Statistical Planning and Inference, 136(10):3625–3646, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000856.

Wu:2007:PST

[Wu07] Hong-Dar Isaac Wu. A partial score test for difference among heterogeneous populations. Journal of Statistical Planning and Inference, 137(2):527–537, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000401.

Wang:2004:APM

[WvEZ04] Xiaogang Wang, Constance van Eeden, and James V. Zidek. Asymptotic properties of maximum weighted likelihood estimators. Journal of Statistical Planning and Inference, 119(1):37–54, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580200410X.

Wang:2002:PAL

[WW02a] Weijing Wang and Martin T. Wells. Power analysis for linear models with spherical errors. Journal of Statistical Planning and Inference, 108(1–2):155–171, November 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002768.

Williams:2002:MGF

[WW02b] E. J. Williams and R. K. Watson. Martingale generating functions for Markov chains. Journal of Statistical Planning and Inference, 103(1–2):39–49, April 15, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001963.

Wong:2002:SSA

[WW02c] A. C. M. Wong and J. Wu. Small sample asymptotic inference for the coefficient of variation: normal and nonnormal models. Journal of Statistical Planning and Inference, 104(1):73–82, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002415.

Wu:2005:TSM

[WW05] Shu-Fei Wu and Chin-Chuan Wu. Two stage multiple comparisons with the average for exponential location parameters under heteroscedasticity. Journal of Statistical Planning and Inference, 134(2):392–408, October 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002228.

Weng:2006:ACS

[WW06] Ruby C. Weng and Michael Woodroofe. Approximate confidence sets for a stationary AR(p) process. Journal of Statistical Planning and Inference, 136(8):2719–2745, August 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004653.

Wu:2008:NST

[WW08] Samuel S. Wu and Weizhen Wang. A note on step-up test in orthogonal saturated designs. Journal of Statistical Planning and Inference, 138(10):3149–3156, October 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000025.

Wang:2009:EAD

[WW09] X. Wang and D. Wu. The existence of almost difference families. Journal of Statistical Planning and Inference, 139(12):4200–4205, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580900192X.

Waterhouse:2008:DSC

[WWEL08] T. H. Waterhouse, D. C. Woods, J. A. Eccleston, and S. M. Lewis. Design selection criteria for discrimination/estimation for nested models and a binomial response. Journal of Statistical Planning and Inference, 138(1):132–144, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001966.

Wu:2008:RRM

[WWL08] Chien-Hua Wu, Shu-Mei Wan, and Mei-Chi Li. Randomized response model in a matched pair study. Journal of Statistical Planning and Inference, 138(7):2059–2066, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003588.

Wu:2005:LBC

[WWN05] J. Wu, A. C. M. Wong, and K. W. Ng. Likelihood-based confidence interval for the ratio of scale parameters of two independent Weibull distributions. Journal of Statistical Planning and Inference, 135(2):487–497, December 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002459.

Wiens:2008:RDO

[WX08a] Douglas P. Wiens and Xiaojian Xu. Robust designs for one-point extrapolation. Journal of Statistical Planning and Inference, 138(5):1339–1357, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002662.

Wiens:2008:RPE

[WX08b] Douglas P. Wiens and Xiaojian Xu. Robust prediction and extrapolation designs for misspecified generalized linear regression models. Journal of Statistical Planning and Inference, 138(1):30–46, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001930.

Welsh:2006:LRV

[WY06] A. H. Welsh and T. W. Yee. Local regression for vector responses. Journal of Statistical Planning and Inference, 136(9):3007–3031, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805000364.

Wang:2007:RMU

[WY07a] Wei Wang and Yannis G. Yatracos. Representations of moments using lower-order conditional moments. Journal of Statistical Planning and Inference, 137(1):299–305, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580500282X.

Wu:2007:UPS

[WY07b] Samuel S. Wu and Mark C. K. Yang. Using pilot study information to increase efficiency in clinical trials. Journal of Statistical Planning and Inference, 137(7):2172–2183, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806002084.

Wang:2009:EAU

[WYL09a] Qihua Wang, Lili Yao, and Peng Lai. Estimation of the area under ROC curve with censored data. Journal of Statistical Planning and Inference, 139(3):1033–1044, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002826.

Wu:2009:EVC

[WYL09b] Mi-Xia Wu, Kai-Fun Yu, and Ai-Yi Liu. Estimation of variance components in the mixed effects models: a comparison between analysis of variance and spectral decomposition. Journal of Statistical Planning and Inference, 139(12):3962–3973, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000792.

Wynn:2008:ASC

[Wyn08] H. P. Wynn. Algebraic solutions to the connectivity problem for m-way layouts: Interaction-contrast aliasing. Journal of Statistical Planning and Inference, 138(1):259–271, January 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002042.

Wei:2001:NSA

[WZ01a] Hongzeng Wei and Yibin Zhang. The number of solutions to the alternate matrix equation over a finite field and a qidentity. Journal of Statistical Planning and Inference, 94 (2):349–358, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002652.

Wei:2001:NSB

[WZ01b] Hongzeng Wei and Xingfen Zheng. The number of solutions to the bilinear matrix equation over a finite field. Journal of Statistical Planning and Inference, 94(2):359–369, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002664.

Wang:2003:NTM

[WZ03] Weizhen Wang and Linda H. Zhao. Nonparametric tests for the mean of a non-negative population. Journal of Statistical Planning and Inference, 110(1–2):75–96, January 15, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002944.

Wan:2004:UIL

[WZ04] Alan T. K. Wan and Guohua Zou. On unbiased and improved loss estimation for the mean of a multivariate normal distribution with unknown variance. Journal of Statistical Planning and Inference, 119(1):17–22, January 15, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802004068.

Welsh:2006:ERM

[WZ06] A. H. Welsh and X. H. Zhou. Estimating the retransformed mean in a heteroscedastic two-part model. Journal of Statistical Planning and Inference, 136(3):860–881, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804003337.

Wiens:2008:RED

[WZ08] Douglas P. Wiens and Julie Zhou. Robust estimators and designs for field experiments. Journal of Statistical Planning and Inference, 138(1):93–104, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002005.

Wu:2009:TWM

[WZ09] Mingxin Wu and Yijun Zuo. Trimmed and Winsorized means based on a scaled deviation. Journal of Statistical Planning and Inference, 139(2):350–365, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001924.

Xie:2006:MBL

[XC06] Yang Xie and Bradley P. Carlin. Measures of Bayesian learning and identifiability in hierarchical models. Journal of Statistical Planning and Inference, 136(10):3458–3477, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375805001278.

Xu:2008:LRR

[XC08] Yong-Li Xu and Di-Rong Chen. Learning rates of regularized regression for exponentially strongly mixing sequence. Journal of Statistical Planning and Inference, 138(7):2180–2189, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003461.

Xekalaki:2005:IPY

[XD05] Evdokia Xekalaki and Caterina Dimaki. Identifying the Pareto and Yule distributions by properties of their reliability measures. Journal of Statistical Planning and Inference, 131(2):231–252, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375804000813.

Xie:2000:AMU

[XF00] Min-Yu Xie and Kai-Tai Fang. Admissibility and minimaxity of the uniform design measure in nonparametric regression model. Journal of Statistical Planning and Inference, 83(1):101–111, January 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375899000890.

Xin:2001:AAC

[Xin01] Dingjia Xin. Analytic approach to coset-leader decoding leads to extension of the Welch–Berlekamp theorem. Journal of Statistical Planning and Inference, 94(2):371–380, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800002676.

Xing:2009:EIS

[Xin09] Guodong Xing. On the exponential inequalities for strictly stationary and negatively associated random variables. Journal of Statistical Planning and Inference, 139(10):3453–3460, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000858.

Xu:2006:MAB

[XL06a] Hongquan Xu and Sovia Lau. Minimum aberration blocking schemes for two- and three-level fractional factorial designs. Journal of Statistical Planning and Inference, 136(11):4088–4118, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001345.

Xu:2006:LRO

[XL06b] Maochao Xu and Xiaohu Li. Likelihood ratio order of mspacings for two samples. Journal of Statistical Planning and Inference, 136(12):4250–4258, December 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002478.

Xiao:2008:LIL

[XL08a] Zhihong Xiao and Luqin Liu. Laws of iterated logarithm for quasi-maximum likelihood estimator in generalized linear model. Journal of Statistical Planning and Inference, 138(3):611–617, March 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807001401.

Xu:2008:NDF

[XL08b] Maochao Xu and Xiaohu Li. Negative dependence in frailty models. Journal of Statistical Planning and Inference, 138 (5):1433–1441, May 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807002753.

Xiong:2009:CAC

[XM09] Shifeng Xiong and Weiyan Mu. On construction of asymptotically correct confidence intervals. Journal of Statistical Planning and Inference, 139(4):1394–1404, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003418.

Xu:2001:EFS

[XP01] Yunqing Xu and Huisheng Pei. Existence of frame SOLS of type a^{n}b^{1} for odd n. Journal of Statistical Planning and Inference, 94(2):381–391, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800002688.

Xie:2002:WGM

[XP02] Jingdong Xie and Carey E. Priebe. A weighted generalization of the Mann-Whitney-Wilcoxon statistic. Journal of Statistical Planning and Inference, 102(2):441–466, April 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801001112.

Xing:2009:SCB

[XR09] Yang Xing and Bo Ranneby. Sufficient conditions for Bayesian consistency. Journal of Statistical Planning and Inference, 139(7):2479–2489, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004400.

Xu:2007:ESF

[Xu07] Jian-Lun Xu. On estimating a survival function under the generalized proportional hazards model. Journal of Statistical Planning and Inference, 137(4):1161–1172, April 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001029.

Xu:2009:RPE

[Xu09] Xiaojian Xu. Robust prediction and extrapolation designs for censored data. Journal of Statistical Planning and Inference, 139(2):486–502, February 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001961.

Xue:2004:ARE

[Xue04] Liu-Gen Xue. Approximation rates of the error distribution of wavelet estimators of a density function under censorship. Journal of Statistical Planning and Inference, 118(1–2):167–183, January 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003932.

Xu:2007:EPS

[XW07] Ke Xu and Christopher K. Wikle. Estimation of parameterized spatio-temporal dynamic models. Journal of Statistical Planning and Inference, 137(2):567–588, February 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375806000292.

Xu:2008:TSE

[XW08] Jing Xu and Jinde Wang. Two-stage estimation of inequalityconstrained marginal linear models with longitudinal data. Journal of Statistical Planning and Inference, 138(6):1905– 1918, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580700331X.

Xue:2006:ESP

[XY06] Lan Xue and Lijian Yang. Estimation of semi-parametric additive coefficient model. Journal of Statistical Planning and Inference, 136(8):2506–2534, August 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804004586.

Xing:2008:NEI

[XY08] Guodong Xing and Shanchao Yang. Notes on the exponential inequalities for strictly stationary and positively associated random variables. Journal of Statistical Planning and Inference, 138(12):4132–4140, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001043.

Yata:2009:DSM

[YA09] Kazuyoshi Yata and Makoto Aoshima. Double shrink methodologies to determine the sample size via covariance structures. Journal of Statistical Planning and Inference, 139(2):81–99, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808001274.

Yang:2000:FEU

[Yan00] Song Yang. Functional estimation under interval censoring case 1. Journal of Statistical Planning and Inference, 89(1–2):135–144, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800000914.

Yanagisawa:2005:SDM

[Yan05] Yukio Yanagisawa. Scaled density models for binary response data and their D- and D_{s}-optimal designs. Journal of Statistical Planning and Inference, 128(2):649–660, February 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000230.

Yang:2008:ODG

[Yan08] Min Yang. A-optimal designs for generalized linear models with two parameters. Journal of Statistical Planning and Inference, 138(3):624–641, March 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001875.

Yao:2008:FAF

[Yao08] Fang Yao. Functional approach of flexibly modelling generalized longitudinal data and survival time. Journal of Statistical Planning and Inference, 138(4):995–1009, April 1, 2008.

CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807001826.

Yin:2006:MBD

[YB06] Xiangrong Yin and Efstathia Bura. Moment-based dimension reduction for multivariate response regression. Journal of Statistical Planning and Inference, 136(10):3675–3688, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580500073X.

Yang:2008:GRJ

[YB08] Guijun Yang and Neil A. Butler. Grid representations for 2_{IV}^{m−(m−k)} join designs containing maximum number of clear two-factor interactions. Journal of Statistical Planning and Inference, 138(5):1472–1478, May 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002911.

Yuan:2002:DMD

[YC02] Ming Yuan and Jiaqin Chen. Deconvolving multidimensional density from partially contaminated observations. Journal of Statistical Planning and Inference, 104(1):147–160, May 1, 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375801002385.

You:2007:SGL

[YC07] Jinhong You and Gemai Chen. Semiparametric generalized least squares estimation in partially linear regression models with correlated errors. Journal of Statistical Planning and Inference, 137(1):117–132, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002818.

Yan:2006:AAD

[YCB06] Xiuyuan Yan, Yu Cheng, and Zhidong Bai. Asymptotics of adaptive design with two alternating generating matrices. Journal of Statistical Planning and Inference, 136(11):4043–4058, November 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375805001035.

Yekutieli:2008:FDR

[Yek08] Daniel Yekutieli. False discovery rate control for nonpositively regression dependent test statistics. Journal of Statistical Planning and Inference, 138(2):405–415, February 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002510.

Yubin:2005:SAI

[YGJ05] Tian Yubin, Li Guoying, and Yang Jie. Saddlepoint approach to inference for response probabilities under the logistic response model. Journal of Statistical Planning and Inference, 133(2):405–416, August 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804001211.

Yuan:2009:RPE

[YH09] Ming Yuan and Jianhua Z. Huang. Regularized parameter estimation of high dimensional t distribution. Journal of Statistical Planning and Inference, 139(7):2284–2292, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808004163.

Yang:2009:NDO

[YHA09a] Zhao Yang, James W. Hardin, and Cheryl L. Addy. A note on Dean’s overdispersion test. Journal of Statistical Planning and Inference, 139(10):3675–3678, October 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001177.

Yang:2009:STO

[YHA09b] Zhao Yang, James W. Hardin, and Cheryl L. Addy. A score test for overdispersion in Poisson regression based on the generalized Poisson-2 model. Journal of Statistical Planning and Inference, 139(4):1514–1521, April 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003534.

Yang:2009:TOZ

[YHA09c] Zhao Yang, James W. Hardin, and Cheryl L. Addy. Testing overdispersion in the zero-inflated Poisson model. Journal of Statistical Planning and Inference, 139(9):3340– 3353, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809000810.

Yang:2005:CAN

[YHZ05] Song Yang, Li Hsu, and Lueping Zhao. Combining asymptotically normal tests: case studies in comparison of two groups. Journal of Statistical Planning and Inference, 133(1):139–158, July 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001089.

Yin:2000:EDG

[Yin00] Jianxing Yin. Existence of directed GDDs with block size five and index λ ≥ 2. Journal of Statistical Planning and Inference, 86(2):619–627, May 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001329.

Yin:2001:PDE

[Yin01] Jianxing Yin. Packing designs with equal-sized holes. Journal of Statistical Planning and Inference, 94(2):393–403, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580000269X.

Yin:2009:BGF

[Yin09] Guosheng Yin. Bayesian goodness-of-fit test for censored data. Journal of Statistical Planning and Inference, 139(4): 1474–1483, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808003492.

Yamamoto:2009:OAL

[YJS09] Watalu Yamamoto, Lu Jin, and Kazuyuki Suzuki. Optimal allocations for load-sharing k-out-of-n:F systems. Journal of Statistical Planning and Inference, 139(5):1777–1781, May 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808002528.

Yi:2007:CLM

[YL07] Grace Y. Yi and Jerald F. Lawless. A corrected likelihood method for the proportional hazards model with covariates subject to measurement error. Journal of Statistical Planning and Inference, 137(6):1816–1828, June 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001133.

Yao:2009:AAO

[YL09] Fang Yao and Thomas C. M. Lee. Automatic and asymptotically optimal data sharpening for nonparametric regression. Journal of Statistical Planning and Inference, 139(12):4017–4030, December 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001323.

Yang:2004:EMV

[YLH04] Ming-Chung Yang, Ding-Wen Lee, and J. T. Gene Hwang. The equivalence of the mid p-value and the expected p-value for testing equality of two balanced binomial proportions. Journal of Statistical Planning and Inference, 126(1):273–280, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002623.

Yang:2006:FFS

[YLLZ06] Jian-Feng Yang, Peng-Fei Li, Min-Qian Liu, and Run-Chu Zhang. 2^{(n_{1}+n_{2})−(k_{1}+k_{2})} fractional factorial split-plot designs containing clear effects. Journal of Statistical Planning and Inference, 136(12):4450–4458, December 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002454.

Ye:2000:ACO

[YLS00] Kenny Q. Ye, William Li, and Agus Sudjianto. Algorithmic construction of optimal symmetric Latin hypercube designs. Journal of Statistical Planning and Inference, 90(1):145–159, September 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375800001051.

Yamada:2002:OML

[YM02] Shu Yamada and Tomomi Matsui. Optimality of mixedlevel supersaturated designs. Journal of Statistical Planning and Inference, 104(2):459–468, June 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002488.

Yagi:2008:COS

[YMJ08] Shintaro Yagi, Kentaro Mimura, and Masakazu Jimbo. A construction of OA(st,t+1,s,t) by nonlinear functions and some classification for s=4. Journal of Statistical Planning and Inference, 138(11):3309–3315, November 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000785.

Yanagihara:2005:DAL

[YO05a] Hirokazu Yanagihara and Chihiro Ohmoto. On distribution of AIC in linear regression models. Journal of Statistical Planning and Inference, 133(2):417–433, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580400117X.

Yanagimoto:2005:SPM

[YO05b] Takemi Yanagimoto and Toshio Ohnishi. Standardized posterior mode for the flexible use of a conjugate prior. Journal of Statistical Planning and Inference, 131(2):253–269, May 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804000783.

Yanagimoto:2009:BPD

[YO09] Takemi Yanagimoto and Toshio Ohnishi. Bayesian prediction of a density function in terms of e-mixture. Journal of Statistical Planning and Inference, 139(9):3064–3075, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000354.

Yang:2007:ECD

[YP07] Min Yang and Mingue Park. Efficient crossover designs for comparing test treatments with a control treatment when p=3. Journal of Statistical Planning and Inference, 137(6):2056–2067, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806001169.

Yu:2000:INE

[YR00] M. Yu and J. N. K. Rao. Inference for nested error linear regression models with unequal error variances. Journal of Statistical Planning and Inference, 88(2):233–253, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375800000811.

You:2003:PHB

[YR03] Yong You and J. N. K. Rao. Pseudo hierarchical Bayes small area estimation combining unit level models and survey weights. Journal of Statistical Planning and Inference, 111(1–2):197–208, ???? 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375802003014.

Suen:2005:CRL

[yS05] Chung yi Suen. Construction of m^{4}-run linear graphs by finite geometries. Journal of Statistical Planning and Inference, 133(2):523–536, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804001181.

Suen:2007:SMO

[yS07] Chung yi Suen. Some mixed orthogonal arrays obtained by orthogonal projection matrices. Journal of Statistical Planning and Inference, 137(5):1704–1710, May 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806002527.

Yang:2008:OEC

[YS08] Min Yang and John Stufken. Optimal and efficient crossover designs for comparing test treatments to a control treatment under various models. Journal of Statistical Planning and Inference, 138(1):278–285, January 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002121.

Yin:2009:CCM

[YSB09] Yaling Yin, Christine E. Soteros, and Mikelis G. Bickis. A clarifying comparison of methods for controlling the false discovery rate. Journal of Statistical Planning and Inference, 139(7):2126–2137, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808003911.

Suen:2003:CAO

[ySD03] Chung yi Suen and Aloke Dey. Construction of asymmetric orthogonal arrays through finite geometries. Journal of Statistical Planning and Inference, 115(2):623–635, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001659.

Yanagihara:2003:BCA

[YSF03] Hirokazu Yanagihara, Risa Sekiguchi, and Yasunori Fujikoshi. Bias correction of AIC in logistic regression models. Journal of Statistical Planning and Inference, 115(2):349–360, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001672.

Yang:2003:DET

[YSH03] Lijian Yang, Stefan Sperlich, and Wolfgang Härdle. Derivative estimation and testing in generalized additive models. Journal of Statistical Planning and Inference, 115(2):521–542, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001635.

Suen:2005:CMO

[ySK05] Chung yi Suen and Warren F. Kuhfeld. On the construction of mixed orthogonal arrays of strength two. Journal of Statistical Planning and Inference, 133(2):555–560, August 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804001673.

Yu:2006:DAD

[Yu06] Hong-Fwu Yu. Designing an accelerated degradation experiment with a reciprocal Weibull degradation rate. Journal of Statistical Planning and Inference, 136(1):282–297, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804003015.

Yuan:2008:SPI

[Yua08] Ao Yuan. Some practical issues of adaptive GPU design. Journal of Statistical Planning and Inference, 138(10):2953–2974, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807004065.

Yun:2000:CPT

[Yun00] Seokhoon Yun. A class of Pickands-type estimators for the extreme value index. Journal of Statistical Planning and Inference, 83(1):113–124, January 1, 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000853.

Yun:2002:GPE

[Yun02] Seokhoon Yun. On a generalized Pickands estimator of the extreme value index. Journal of Statistical Planning and Inference, 102(2):389–409, April 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801001008.

Yu:2006:DRE

[YvdL06] Zhuo Yu and Mark van der Laan. Double robust estimation in longitudinal marginal structural models. Journal of Statistical Planning and Inference, 136(3):1061–1089, March 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804003489.

Yin:2004:EPC

[YW04] Su-Ju Yin and Song-Gui Wang. Estimating the parameters of a circle by heteroscedastic regression models. Journal of Statistical Planning and Inference, 124(2):439–451, September 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375803002064.

Yi:2008:AEE

[YW08] Yanqing Yi and Xikui Wang. Asymptotically efficient estimation in response adaptive trials. Journal of Statistical Planning and Inference, 138(10):2899–2905, October 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807004090.

Yang:2009:AFW

[YW09a] Hu Yang and Litong Wang. An alternative form of the Watson efficiency. Journal of Statistical Planning and Inference, 139(8):2767–2774, August 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000056.

Ye:2009:EPE

[YW09b] Ren-Dao Ye and Song-Gui Wang. Estimating parameters in extended growth curve models with special covariance structures. Journal of Statistical Planning and Inference, 139(8):2746–2756, August 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580800476X.

Ye:2009:IIC

[YW09c] Ren-Dao Ye and Song-Gui Wang. Inferences on the intraclass correlation coefficients in the unbalanced two-way random effects model with interaction. Journal of Statistical Planning and Inference, 139(2):396–410, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580800195X.

Yang:2009:QME

[YWS09] Hu Yang, Litong Wang, and Lijuan Song. Quasi-minimax estimation in the general linear regression model. Journal of Statistical Planning and Inference, 139(7):2117–2125, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808003881.

Yang:2009:ELP

[YXC09] Yiping Yang, Liugen Xue, and Weihu Cheng. Empirical likelihood for a partially linear model with covariate data missing at random. Journal of Statistical Planning and Inference, 139(12):4143–4153, December 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809001839.

You:2006:ELR

[YZ06] Jinhong You and Xian Zhou. Empirical likelihood in a regression model with noised variables. Journal of Statistical Planning and Inference, 136(10):3478–3497, October 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837580500090X.

Ye:2007:ESM

[YZ07] Jane J. Ye and Julie Zhou. Existence and symmetry of minimax regression designs. Journal of Statistical Planning and Inference, 137(1):344–354, January 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375805002946.

You:2009:PLM

[YZ09] Jinhong You and Xian Zhou. Partially linear models and polynomial spline approximations for the analysis of unbalanced panel data. Journal of Statistical Planning and Inference, 139(3):679–695, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808002838.

Yuen:2003:MRL

[YZT03] K. C. Yuen, L. X. Zhu, and N. Y. Tang. On the mean residual life regression model. Journal of Statistical Planning and Inference, 113(2):685–698, May 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802000915.

Yu:2009:SAS

[YZZ09] Zhou Yu, Li-Ping Zhu, and Li-Xing Zhu. On splines approximation for sliced average variance estimation. Journal of Statistical Planning and Inference, 139(4):1493–1505, April 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808003510.

Zacks:2005:HCA

[Zac05a] S. Zacks. Herman Chernoff: an appreciation. Journal of Statistical Planning and Inference, 130(1–2):3–12, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002769.

Zacks:2005:SRR

[Zac05b] S. Zacks. Some recent results on the distributions of stopping times of compound Poisson processes with linear boundaries. Journal of Statistical Planning and Inference, 130(1–2):95– 109, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002551.

Zadeh:2002:TPB

[Zad02] Lotfi A. Zadeh. Toward a perception-based theory of probabilistic reasoning with imprecise probabilities. Journal of Statistical Planning and Inference, 105(1):233–264, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002129.

Zaffalon:2002:ECT

[Zaf02a] Marco Zaffalon. Exact credal treatment of missing data. Journal of Statistical Planning and Inference, 105(1):105– 122, June 2002. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375801002063.

Zaffalon:2002:NCC

[Zaf02b] Marco Zaffalon. The naive credal classifier. Journal of Statistical Planning and Inference, 105(1):5–21, June 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375801002014.

Zaslavsky:2006:HMS

[Zas06] Alan M. Zaslavsky. Hierarchical modeling: into statistical practice. Journal of Statistical Planning and Inference, 136(4):1469–1472, April 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804004136.

Zhao:2009:CMO

[ZB09a] Peng Zhao and N. Balakrishnan. Characterization of MRL order of fail-safe systems with heterogeneous exponential components. Journal of Statistical Planning and Inference, 139(9):3027–3037, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580900038X.

Zhao:2009:SCP

[ZB09b] Peng Zhao and N. Balakrishnan. Stochastic comparisons and properties of conditional generalized order statistics.

Journal of Statistical Planning and Inference, 139(9):2920– 2932, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375809000147.

Zheng:2007:FOR

[ZBD07] Haitao Zheng, Ishwar V. Basawa, and Somnath Datta. Firstorder random coefficient integer-valued autoregressive processes. Journal of Statistical Planning and Inference, 137 (1):212–229, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806000267.

Zhang:2000:MNB

[ZBZ00] Zhongxin Zhang, Barbara A. Burtness, and Daniel Zelterman. The maximum negative binomial distribution. Journal of Statistical Planning and Inference, 87(1):1–19, May 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899001779.

Zhou:2003:NAR

[ZC03] Xiao-Hua Zhou and Pete Castelluccio. Nonparametric analysis for the ROC areas of two diagnostic tests in the presence of nonignorable verification bias. Journal of Statistical Planning and Inference, 115(1):193–213, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001465.

Zhang:2007:EBC

[ZC07] Jing Zhang and Yanxun Chang. Existence of BSAs and cyclic BSAs of block size three. Journal of Statistical Planning and Inference, 137(4):1496–1511, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000814.

Zhang:2009:PSC

[ZC09] Jing Zhang and Yanxun Chang. Partitionable sets and cyclic BSECs with block size four. Journal of Statistical Planning and Inference, 139(6):1974–1979, June 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808003868.

Zhang:2009:LSI

[ZCL09] Hong Zhang, Hanfeng Chen, and Zhaohai Li. Large sample interval mapping method for genetic trait loci in finite regression mixture models. Journal of Statistical Planning and Inference, 139(3):764–779, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002024.

Zhang:2009:SSB

[ZCS09] Yong Zhang, Kejun Chen, and Yigang Sun. Supersimple balanced incomplete block designs with block size 4 and index 9. Journal of Statistical Planning and Inference, 139(10):3612–3624, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809001050.

Zhang:2002:CGH

[ZDLZ02] Yingshan Zhang, Lian Duan, Yiqiang Lu, and Zhongguo Zheng. Construction of generalized Hadamard matrices D(r^{m}(r + 1), r^{m}(r + 1); p). Journal of Statistical Planning and Inference, 104(2):239–258, June 1, 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580100249X.

Zoubeidi:2007:SCI

[ZEB07] Taoufik Zoubeidi and Mohamed Y. El-Bassiouni. Sequential confidence intervals for variance components in one-way random models. Journal of Statistical Planning and Inference, 137(4):1185–1198, April 1, 2007. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806001017.

Zezula:2009:MGC

[Zez09] Ivan Zezula. On multivariate Gaussian copulas. Journal of Statistical Planning and Inference, 139(11):3942–3946, November 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809001608.

Zhao:2001:OED

[ZF01] Yi Zhao and Kai-Tai Fang. Orthogonal exact designs on a sphere, a spherical cap, or a spherical belt. Journal of Statistical Planning and Inference, 98(1–2):279–285, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003311.

Zheng:2005:FIR

[ZG05] Gang Zheng and Joseph L. Gastwirth. Fisher information in randomly sampled sib pairs and extremely discordant sib pairs in genetic analysis for a quantitative trait locus. Journal of Statistical Planning and Inference, 130(1–2):299–315, March 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002666.

Zhang:2008:BVW

[ZG08] Zhenkui Zhang and Joseph Glaz. Bayesian variable window scan statistics. Journal of Statistical Planning and Inference, 138(11):3561–3567, November 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808000980.

Zucker:2008:PFL

[ZGH08] David M. Zucker, Malka Gorfine, and Li Hsu. Pseudo-full likelihood estimation for prospective survival analysis with a general semiparametric shared frailty model: Asymptotic theory. Journal of Statistical Planning and Inference, 138(7):1998– 2016, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375807003606.

Zhou:2008:TSE

[ZH08] Tianyue Zhou and Xuming He. Three-step estimation in linear mixed models with skew- t distributions. Journal of Statistical Planning and Inference, 138(6):1542–1555, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807002972.

Zhu:2009:IOA

[ZH09] Hongjian Zhu and Feifang Hu. Implementing optimal allocation for sequential continuous responses with multiple treatments. Journal of Statistical Planning and Inference, 139(7):2420–2430, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375808004369.

Zhao:2000:SCE

[Zha00] Lincheng Zhao. Some contributions to M-estimation in linear models. Journal of Statistical Planning and Inference, 88(2):189–203, August 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800000781.

Zhang:2001:ECC

[Zha01a] Shengyuan Zhang. Existence of certain class of duadic group algebra codes. Journal of Statistical Planning and Inference, 94(2):405–411, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800002706.

Zhang:2001:RBA

[Zha01b] Zhao Zhi Zhang. A relation between the average Hamming distance and the average Hamming weight of binary codes. Journal of Statistical Planning and Inference, 94(2):413–419, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800002718.

Zhang:2004:AEN

[Zha04a] Chunming Zhang. Assessing the equivalence of nonparametric regression tests based on spline and local polynomial smoothers. Journal of Statistical Planning and Inference, 126(1):73–95, November 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375803002660.

Zhang:2004:AEE

[Zha04b] Haimeng Zhang. Asymptotic efficiency of estimation in the partial Koziol–Green model. Journal of Statistical Planning and Inference, 124(2):399–408, September 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300209X.

Zhang:2006:BCM

[Zha06] Biao Zhang. Bias-corrected maximum semiparametric likelihood estimation under logistic regression models based on case-control data. Journal of Statistical Planning and Inference, 136(1):108–124, January 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002812.

Zhao:2007:CST

[Zha07] H. B. Zhao. Comparing several treatments with a control. Journal of Statistical Planning and Inference, 137(9): 2996–3006, September 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375807000626.

Zhang:2009:PFC

[Zha09a] Riquan Zhang. Proportional functional coefficient time series models. Journal of Statistical Planning and Inference, 139(3):749–763, March 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375808002656.

Zhang:2009:SSP

[Zha09b] Shunpu Zhang. The split sample permutation t-tests. Journal of Statistical Planning and Inference, 139(10):3512–3524, October 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375809000949.

Zhang:2009:MSL

[Zha09c] Yongli Zhang. Model selection: a Lagrange optimization approach. Journal of Statistical Planning and Inference, 139(9):3142–3159, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375809000524.

Zhang:2009:LBC

[Zha09d] Zhiwei Zhang. Likelihood-based confidence sets for partially identified parameters. Journal of Statistical Planning and Inference, 139(3):696–710, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002590.

Zhao:2009:LMR

[Zha09e] Yang Zhao. Likelihood methods in randomized trials with noncompliance and subsequent nonresponse. Journal of Statistical Planning and Inference, 139(2):310–316, February 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375808001894.

Zhigljavsky:2003:PET

[Zhi03] Anatoly Zhigljavsky. Probabilistic existence theorems in group testing. Journal of Statistical Planning and Inference, 115(1):1–43, July 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375802001489.

Zhou:2008:OMR

[Zho08] Julie Zhou. D-optimal minimax regression designs on discrete design space. Journal of Statistical Planning and Inference, 138(12):4081–4092, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808000670.

Zhu:2001:CDR

[Zhu01] L. Zhu. A class of doubly resolvable GDDs from Baer subplanes. Journal of Statistical Planning and Inference, 94 (2):421–427, April 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837580000272X.

Ziegler:2003:ANK

[Zie03] Klaus Ziegler. On the asymptotic normality of kernel regression estimators of the mode in the nonparametric random design model. Journal of Statistical Planning and Inference, 115(1):123–144, July 1, 2003. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802001520.

Zhao:2005:EEA

[ZjLS05] Xingqiu Zhao, Hee jeong Lim, and Jianguo Sun. Estimating equation approach for regression analysis of failure time data in the presence of interval-censoring. Journal of Statistical Planning and Inference, 129(1–2):145–157, February 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804001594.

Zhang:2009:DBK

[ZK09a] Shunpu Zhang and Rohana J. Karunamuni. Deconvolution boundary kernel method in nonparametric density estimation. Journal of Statistical Planning and Inference, 139(7):2269– 2283, July 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375808004175.

Zhao:2009:MEL

[ZK09b] Meng Zhao and K. B. Kulasekera. Minimax estimation of linear functionals under squared error loss. Journal of Statistical Planning and Inference, 139(9):3160–3176, September 1, 2009. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375809000536.

Zhang:2001:AAI

[ZKW01] Rong Zhang, Jeankyung Kim, and Michael Woodroofe. Asymptotic analysis of isotonic estimation for grouped data. Journal of Statistical Planning and Inference, 98(1–2):107– 117, October 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375800003013.

Zhang:2008:APP

[ZL08] Caiya Zhang and Zhengyan Lin. On asymptotic properties of the parameter estimator for a type of SPDE. Journal of Statistical Planning and Inference, 138(12):4033–4040, December 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S037837580800147X.

Zheng:2006:REC

[ZM06] Gang Zheng and Reza Modarres. A robust estimate of the correlation coefficient for bivariate normal distribution using ranked set sampling. Journal of Statistical Planning and Inference, 136(1):298–309, January 1, 2006. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375804002538.

Zhu:2008:SBP

[ZMJW08] J. X. Zhu, G. J. McLachlan, L. Ben-Tovim Jones, and I. A. Wood. On selection biases with prediction rules formed from gene expression data. Journal of Statistical Planning and Inference, 138(2):374–386, February 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375807002492.

Zhou:2008:LBW

[ZN08] Yong-Dao Zhou and Jian-Hui Ning. Lower bounds of the wrap-around L_{2}-discrepancy and relationships between MLHD and uniform design with a large size. Journal of Statistical Planning and Inference, 138(8):2330–2339, August 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375807003710.

Zou:2009:RQM

[ZNWS09] Fei Zou, Lei Nie, Fred. A. Wright, and Pranab K. Sen. A robust QTL mapping procedure. Journal of Statistical Planning and Inference, 139(3):978–989, March 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808002887.

Zografos:2000:MMD

[Zog00] K. Zografos. Measures of multivariate dependence based on a distance between Fisher information matrices. Journal of Statistical Planning and Inference, 89(1–2):91–107, August 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800000963.

Zografos:2008:SED

[Zog08] K. Zografos. On some entropy and divergence type measures of variability and dependence for mixed continuous and discrete variables. Journal of Statistical Planning and Inference, 138(12):3899–3914, December 1, 2008. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375808001419.

Zhang:2000:OBT

[ZP00] Runchu Zhang and DongKwon Park. Optimal blocking of two-level fractional factorial designs. Journal of Statistical Planning and Inference, 91(1):107–121, November 1, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375800001336.

Zheng:2004:NTS

[ZP04] Gang Zheng and Sangun Park. A note on time savings in censored life testing. Journal of Statistical Planning and Inference, 124(2):289–300, September 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803002088.

Zheng:2005:ALL

[ZP05] Gang Zheng and Sangun Park. Another look at life testing. Journal of Statistical Planning and Inference, 127(1–2): 103–117, January 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803002891.

Zhou:2005:NCI

[ZQ05] Xiao-Hua Zhou and Gengsheng Qin. A new confidence interval for the difference between two binomial proportions of paired data. Journal of Statistical Planning and Inference, 128(2):527–542, February 1, 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375803003781.

Zhou:2007:SNC

[ZQ07] Xiao-Hua Zhou and Gengsheng Qin. A supplement to: “A new confidence interval for the difference between two binomial proportions of paired data”. Journal of Statistical Planning and Inference, 137(1):357–358, January 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375806000255.

Zhang:2005:MLE

[ZR05] Zhiwei Zhang and Howard E. Rockette. On maximum likelihood estimation in parametric regression with missing covariates. Journal of Statistical Planning and Inference, 134(1):206–223, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375804002095.

Zhang:2006:GML

[ZRM06] Haimeng Zhang, M. Bhaskara Rao, and Rupa C. Mitra. On the generalized maximum likelihood estimator of survival function under Koziol–Green model. Journal of Statistical Planning and Inference, 136(9):3032–3051, September 1, 2006. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375804004720.

Zuo:2000:PSR

[ZS00] Yijun Zuo and Robert Serfling. On the performance of some robust nonparametric location measures relative to a general notion of multivariate symmetry. Journal of Statistical Planning and Inference, 84(1–2):55–79, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899001421.

Zhu:2005:SSD

[ZS05] Zhengyuan Zhu and Michael L. Stein. Spatial sampling design for parameter estimation of the covariance function. Journal of Statistical Planning and Inference, 134(2):583–603, October 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375804002265.

Zhou:2008:MSQ

[ZS08] Weihua Zhou and Robert Serfling. Multivariate spatial U-quantiles: a Bahadur–Kiefer representation, a Theil– Sen estimator for multiple regression, and a robust dispersion estimator. Journal of Statistical Planning and Inference, 138(6):1660–1678, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003102.

Zen:2004:CRO

[ZT04] Mei-Mei Zen and Min-Hsiao Tsai. Criterion-robust optimal designs for model discrimination and parameter estimation in Fourier regression models. Journal of Statistical Planning and Inference, 124(2):475–487, September 1, 2004. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037837580300212X.

Zhou:2004:NID

[ZTQ04] Xiao-Hua Zhou, Min Tsao, and Gengsheng Qin. New intervals for the difference between two independent binomial proportions. Journal of Statistical Planning and Inference, 123(1):97–115, June 1, 2004. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375803001460.

Zhang:2003:CCS

[ZW03a] Tonglin Zhang and Michael Woodroofe. Credible and confidence sets for restricted parameter spaces. Journal of Statistical Planning and Inference, 115(2):479–490, August 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375802001702.

Zou:2003:AME

[ZW03b] Guohua Zou and Alan T. K. Wan. Admissible and minimax estimation of the parameter n in the binomial distribution. Journal of Statistical Planning and Inference, 113(2):451–466, May 1, 2003. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0378375802001131.

Zhao:2007:GMI

[ZWHF07] Shang Zhao, Douglas A. Wolfe, Tsai-Shang Huang, and Gerald S. Frankel. Generalized model for IGC growth in aluminum alloys. Journal of Statistical Planning and Inference, 137(7):2405–2412, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375806002709.

Zhang:2001:GBR

[ZY01] Zhong-Zhan Zhang and Takashi Yanagawa. Generalized Begun–Reid estimators for multiple hazard rates. Journal of Statistical Planning and Inference, 96(1):203–220, June 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-

1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375800003360.

Zhang:2002:PCG

[ZY02] Shucheng Zhang and Jianxing Yin. Packing of k_{v} with certain graphs of five vertices. Journal of Statistical Planning and Inference, 106(1–2):387–408, August 2002. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375802002240.

Zhou:2007:SIS

[ZY07] Haibo Zhou and Jinhong You. Statistical inference for a semiparametric measurement error regression model with heteroscedastic errors. Journal of Statistical Planning and Inference, 137(7):2263–2276, July 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S037837580600245X.

Zhou:2008:RTG

[ZZF08] Jianhui Zhou, Zhongyi Zhu, and Wing K. Fung. Robust testing with generalized partial linear models for longitudinal data. Journal of Statistical Planning and Inference, 138(6):1871–1883, July 1, 2008. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0378375807003278.

Zhang:2009:MRA

[ZZF09] Xu Zhang, Mei-Jie Zhang, and Jason P. Fine. A mass redistribution algorithm for right-censored and left-truncated time to event data. Journal of Statistical Planning and Inference, 139(9):3329–3339, September 1, 2009. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375809000731.

Zhang:2007:MSA

[ZZL07] Qiao-Zhen Zhang, Run-Chu Zhang, and Min-Qian Liu. A method for screening active effects in supersaturated designs. Journal of Statistical Planning and Inference, 137(6): 2068–2079, June 1, 2007. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375806001832.